A Selected Bibliography of Publications by, and about, Werner Heisenberg

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 August 2025 Version 1.246

Title word cross-reference

(k_{1} − k_{n}) [Tem91]. + [BKBS93]. $1.40 [Ano68]. 1/2 [BHX96, JKS13, LB72, Man91]. $10.95 [Jor78]. $11.50 [Jor78]. $12.00 [Kra07, Lan08]. $129.99 [Cru23]. $18.50 [Jor78]. $2 [Ano29]. 2 [Hen93]. $27.00 [Bed09]. $27.50 [Cas93a, Wil94]. $28.95 [Klo96]. $29.95 [Fra94, Hay90]. 2pπ [SHI73]. 3 [Isa82]. $35 [Lan02]. $35.00 [Neu99, Par06]. $4.75 [Ano54]. $6.95 [Lib68]. $60.00 [Wat01]. $80.00 [Cas11]. = [BKBS93].

∗ [BT96]. ^{235} [DK78]. ^{238} [Tur46a, Tur46b]. ^{239} [DK78]. ^{252} [DK78]. _{2} [BBF^{+}45, DDH41a, DDH42a, DDH42c, DDH89a, DDH89c, DDH89f, Hei42b, HBB^{+}89, Hei89g, VBN95a, VBN95b, vBN95c]. _{4} [VBN95a, VBN95b, vBN95c]. _{q} [Bec02]. β [Gau14]. C^{∗} [Pac88, Sud06]. U_{h,w}(H(4)) [Abd97]. d [VZ09]. η [Hei89k]. H(1)_{q} [CGST91]. H_{0} [CBS20]. K_{0} [Pac88]. L^{2} [NRS71]. L_{2}(R^{d}) [RS94]. L_{p} [LN06]. Λ [Cul05]. ω [DH62]. π [Hei51l]. pq − qp = \sqrt{−1}h̄/(2π) [Ber35]. q [AFW94, GH10, Sch99]. S [Bor70, Bor72, Cus86, NRS71]. S2 [Wan98]. S = 1/2 [MC92]. su(2, 1) [MP92]. SU(3) [CM02, MP92]. u [dB58]. U(1, 3) ⊗_{s} H(1, 3) [Low97]. VN(G) [CZ81].

V θ [Nel72]. XXZ [LFK99, ZG96]. xy [Pug80]. |kq{K̃.}(k_{1} − k_{n}): [λ̃], S_{n}⟩⟩ [Tem91].

-algebra [AFW94, Sud06]. -algebras [BT96, Pac88]. -component [Isa82]. -D [Hen93]. -deformed [Sch99]. -electron [SHI73]. -graded [GH10]. -groups [Pac88]. -Matrix [Hei89k, Hei89k, Cus86]. -Meson [DH62, Hei51l]. -Mesonen [Hei51l]. -orthogonal [VZ09].

/Das [DRF90a, DRF90b]. /Deutsches [DRF90a, DRF90b]. /European [Ano84b]. /Herder [RFS^{+}91]. /Max [vBL96a, vBL96b]. /Rath [GRE^{+}01]. /Weisbuch [RFS^{+}91]. /Wissenschaftsrat [Ano84a, URWB95]. /Zentrum [DRF90a, DRF90b].

0 [Fra94, Hen00, Mar94, Tur06]. 0-394-51411-4 [Wil94]. 0-520-08499-3 [Mar94]. 0-520-21077-8 [Hen00]. 0-674-01519-3 [Tur06]. 0-7167-2243-7 [Fra94]. 0-7923-3794-8 [Lan02]. 0-9672617-2-4 [Kra07, Lan08].

1 [Ano08]. 1-Megabit-Speicherchip [Ano84a]. 1.5 [GRE^{+}01]. 10 [Str62b]. 100 [Hei42a]. 1000 [Hei57b]. 100th [Hei58l, Hei05b, Jac01, KRW05, Rec01, Rec02b, RH02, vM02]. 109 [RDM^{+}12a]. 112 [FW07]. 12. [Hei43a, Str62c]. 13.80 [Hei53b]. 14th [Hei75s]. 15 [BDR84h]. 15.12.1958 [Hei59j]. 16.10.1971 [Hei72f]. 170 [Kra86a]. 18.11.1962 [Hei63h]. 19 [Hei89-36]. 19-24 [vW41a]. 1911 [Meh75]. 1914 [Hei64c]. 1916-1955 [Mar73]. 1925 [AMS04, Hei35e, Hen98]. 1926 [Sta11a]. 1927 [Rig87]. 1928 [Hei29a, HH84]. 1929 [HEB^{+}80, HvMW79, Pye81]. 1930s [Fie80, Hei93e]. 1931 [Hei31g]. 1932 [Hei32c]. 1933 [CCJ^{+}34]. 1933/Chadwick [Ano51]. 1935 [Hei35d, Hei85b]. 1935/1936 [Ano51]. 1939 [Hei89m, OSWR86, vMHW85, Hay90, Rec86, Sei90a]. 1939-1945 [Sei90a]. 1939-1949 [Hay90]. 1940 [DDH89e, DDH89d, Hei89h]. 1940s [MWL^{+}15]. 1941 [Aas02c, DDH89f, DDH89g, Hei41b, Hei89-36, Hei89x, Hei89-34, vW41a]. 1942 [DDH89a, DDH89c, DDH89b, Hei89g, HBF^{+}89a, Hei89n, HBF^{+}89b, HBF^{+}89c, Hei95, KW93, Bar04]. 1943 [Ano44a, Har43, Hei43r, Wal20b]. 1943/1944 [Hei89e]. 1943/44 [Brü64]. 1944 [Ano44a, Hei89-33]. 1945 [BC95, Cas79a, Cas05, HBB^{+}89, HGW89b, HHH03, Hen00, Kan03, Sei90a]. 1946 [Hei47f]. 1946/47 [Hei47f, Hei47f]. 1947 [Sei90a]. 1947-1977 [Sei90a]. 1948 [Hei48a]. 1949 [PvM93, Wal08, vM93, Bro95, Hay90, Kra96]. 1950 [Hei51k, Hei52p]. 1950/51 [Hei52p, Hei52p]. 1950s [Hei86a, Sei90a]. 1952 [vM96]. 1953 [Hei56g]. 1954 [vM99]. 1954/Mladjenovic [ARK^{+}99]. 1956 [vM01]. 1958 [vM05b]. 1960s [ARK^{+}99, Mla98]. 1960s/Frenkel [ARK^{+}99]. 1962/Gray [SPS^{+}64]. 1967 [Paw06]. 1968 [BDR84d, Hei69b]. 1970 [BSP^{+}70]. 1971/Benz [LSK^{+}76]. 1972 [Hei72c, Meh73]. 1973 [Hei73a]. 1974 [CDvW75]. 1976 [BDR85a, MP77, Hah06, Mac78b]. 1977 [Sei90a].

1979 [Rür79]. 1980 [Dür80]. 1984/Zen [Ano84a]. 1984/Zen-Ereignisse [Ano84a]. 1985 [CW85]. 1985/Welch [Ano84b]. 1985/Welch-Stipendium [Ano84b]. 199 [Bok10]. 1993 [BCEP94].

2 [Ben35, BDR84a]. 2. [Ano35a, BDR84a, Hei43b]. 2.50 [Deh37]. 20 [Cas93a]. 2000 [Log00]. 20th [Bro97, Meh73, Sch96a, Bre97, Hei75a]. 20th-Century [Hei75a]. 22 [CCJ^{+}34]. 238 [Hei41f, Hei89x, vW40]. 24.3.41 [vW41a]. 26 [DDH89c, DDH89b, Hei89n, Hei89-34, HBF^{+}89b]. 26.-28 [DDH89b, HBF^{+}89b]. 27 [Har43]. 28 [DDH89f, DDH89g, Har43, Hei72c]. 28.9.43 [Har43]. 29 [Hei89h]. 298.00 [Hen86]. 2nd [Hei31b, Hei93d].

3 [Ano44a, Ano08, Büh95, CDvW79, HBB^{+}89, Mar94, Rec03, Tur06]. 3-05-501585-1 [Ano08]. 3-540-13609-6 [Hen86]. 3-86025-315-8 [Büh95, Rec03]. 3.50 [DS43]. 3.85 [Bro72]. 30 [HBF^{+}89a]. 31 [Hei64c, Hei89g]. 33rd [Ano68]. 35 [DPL96]. 37 [BDR84h]. 38 [DDH40a, DDH41a, DDH41b, Fri98, Hah40, Hei42b]. 38-ball [HBF^{+}89c]. 38-layer [HBF^{+}42, HBF^{+}89a]. 38-Metal [Hei42b, Hei89g]. 38-Metall [BFH^{+}42b, DDH42b, DDH89b, HBF^{+}42, Hei89g, HBF^{+}89a, HBF^{+}89c, Hei42b]. 3d [Tho97]. 3rd [Cas92a].

4.1.1961 [Hei61d]. 4/Pauli [OSWR86]. 40 [HB82a]. 43 [Sch02a, Wit83]. 47 [Hei47f]. 47th [IEE06]. 49/Atmanspacher/Primas [MGG^{+}95]. 4th [RDP^{+}95].

5 [DDH89d, Hei43r]. 5. [Hei43c]. 50 [Ben30, Ben34, Clu37, Hei51a]. 500th [Gin75, Hei86n]. 51 [Hei52p]. 54 [May87]. 55 [BB69]. 5th [Hei48a].

6 [Ano44a, Hei89m, HBF^{+}89c, vRFS^{+}00]. 6.0 [EHR^{+}96]. 60th [Fue61, Hei72c, Hei89-37, MF69, Wig96b]. 65th [Hei44b, PFH20].

7 [DDH89e, Fra94, Hei35d, HGW89b]. 7.80 [Deh36]. 70th [DB71, HBD71, PRW55]. 72 [Got04]. 75th [Hei43r, Hei64h].

8 [Büh95, Hen00, Rec03]. 8. [Hei43d]. 800-Jahrfeier [Hei58k]. 800th [Hei58k, Hei74g]. 80th [Sch82]. 88-389-1450-8 [Pol00].

9.10.1949 [Hei86j]. 90th [Rec02a]. 95h [DPL96]. 978-0-521-82170-4 [Sta11b]. 978-0-521-88484-6 [Bok10, Sta11b]. 978-3-540-69221-8 [Hen11]. 989 [FS93].

= [BD92].

A-Bomb [Rec91d, RW92, LRD^{+}91, Bro93, Mac85]. A-Bombs [Klo93]. A. [BDR89b, Hei51e, Kan03, Wig96j]. A./Kraus [GRE^{+}01]. a.o [BDR86, BDR89b, Hen86, HvMW79, vMHW85, vM93, vM96, vM99, vM01, vM05b]. A/Part [Bro86, Mat87]. år [Boh64]. Abänderung [Hei24d]. Abb [Büh95, Deh36, Rec03]. Abbildungen [HH14]. abermals [Haa34]. abgehörter [WR92]. Abgeschlossene [Hei48e, Hei48d]. abgeschlossenen [Hei73e]. Abhandlungen [Wit83, Bor63a, Sch27]. Abides [Gla00]. Abroad [Hei77f]. Abschied [Dür76]. Abschiedsworte [Hei69a]. Abschluß [Hei70a, Hei71a]. Abschlußbericht [Eic53]. absolute [Hei22a]. absoluten [Hei22a]. Absorption [Hei38d, Fis41, Hei38d]. Absorptionsquerschnittes [Fis41]. Abstract [BSP^{+}70, Hei37d, Hei39f, Hei39g, Hei39h, Hei63d, Hei38b, Hei82c]. Abstraction [Hei74a, Hei08d, Lac02b, Hei70d, Lac02a, Hei61c]. Abstrakt [Hei37d, Hei38b, Hei39f, Hei39h, Hei63d, Hei82c]. Abstraktion [Hei61c, Hei70d]. Abt [BDR84a, BDR84b, BDR85a]. Abteilung [BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93]. Academic [Her76a, Hei31g, Rin69, Rin90]. Academy [Hei86c, Hei89q, Hei89f, LMOS15]. Acceleration [HEB^{+}80]. Acceleration/Brosche [HEB^{+}80]. Accelerators [Hei73b, Hei72d, Hei72l]. accident [Hei72g, Pop16b]. accidents [Döp42]. accommodation [Hof05, HW12b]. according [Hei53b, KH47]. Account [Sei90a]. Achievements [Ble90]. acht [BBH^{+}62]. Achtzehn [BBF^{+}57b, ABH^{+}55, ABH^{+}87]. Achtzig [Hei38g]. achtzigsten [Hei62i]. acquisition [Hei39e]. Across [Hei74b, Hei71l, Hei73g, Hei76g, Hei77h, Hei84-62, Hei89-29, Cas74]. Acta [Wit83]. acting [Dje07]. actions [DDI03]. Active [Hei47a]. Activities [vL48a, vL48b, Eic53]. activity [Hei56g]. Actors [FW20]. acts [Dür62, Dür80, Dür06, Fra05]. actual [Hei83b]. Adam [DT18]. adaptation [HW07]. Adapting [Hei63a]. Addendum [Hei79c, Hei79b, Hei89x]. Adding [Wit12]. Address [Hei56c, Wig96a, Wig96d, Hei86c, Hei89f, Hei89y]. Addressing [CBS20]. Adolf [CL83]. Adrien [KW87]. advanced [DJR93, RS11b]. advancement [Hei89p, Hol86, Hol98]. Advances [Hei84-43, Hei74p, BDR84h, Hei30f, Hei37d, Hei58m, Hei84s, Hei93c]. adventure [MB11a]. advertisement [HB^{+}72]. Adviser [Eck90]. Advisory [Sei90a]. af [Ano44a, BR64, Hei35e, RRK^{+}64]. affair [HW11]. Affairs [Hei73h, Sch11b, Luc04]. Affäre [HW11]. affären [Sch11b]. affine [DDI03, Tor91]. After [Hee95, Kae31, And89, GHH02, Hei35e, Kae48b]. Aftermath [Wal08, Hen07]. against [ABH^{+}55, ABH^{+}87, Hof93, KP07, Mas04, Wei46, Wei99]. Age [Cas11, Del13, Hei66b, Hei70e, Hen10, Lal15, Sta11b, Car10, Hei69k, Hei69l, Dje07, Gil09, Hah11, Sei90a]. Akademie [Ano08, Hei86c, Hei89f, Hei89q, Rec85a, Wit83]. Akademie-Verlag [Ano08]. Akademische [Deh36, Hei31g, Her76a]. Akademischer [Büh95]. AKNS [Bt87]. Akten [Dür80, Dür62, Fra05]. Aktennotiz [Hei53a]. Aktionspapier [RDP^{+}95]. aktive [Hei47a]. aktuell [URWB95]. aktuell/Koordinierung [URWB95]. Aktuelles [Ano84a, DRF90a, DRF90b, Rec85a, RDP^{+}95, URWB95]. al [Bla86, Car97, Hah06]. Alamos [McC24, Bet91, SR92]. Alan [Jor78]. Albert [Adl03, Bes04, Bro72, CS94, Hei53b, Hei75e, Hei77c, Mar73, Wig96g, BB69, BB71a, BB71b, BB05, Hei49a, Hei55a, Hei74o, Hei75e, Hei77c, Hei83f, Hol00b, Meh87, Wig96g, Wig96l]. Albrecht [PFH20, PFH20]. Alemán [Bal07]. Alexander [Ano71, Hei56g, Hei69e, Hei89-28, Hei69e, Hei89-28, Pfe69]. Alfred [Adl03, Hol93a, LH93]. Algebra [Bil03, Sor02, SPS^{+}64, Abd97, AFW94, BMP87, BM94, BP75, CZ81, DLW08, ES95, FT94, Koo89, LFK99, LJ08, LWZ10, MMSV98, Sch02a, Sch02b, Sch99, Sud06, Tem91, VZ09]. Algebra/Bermant [SPS^{+}64]. Algebraic [LC71, Nel72, CRR93, KS92, Liu72, Tru72]. Algebras [FO99, BT96, Bt87, BP75, DMdA06, FO94, GH10, GS12, MP92, Pac88]. algorithm [Bec02]. alle [Cap01, HB^{+}72]. allemande [Lur95]. allemandes [Kar07]. Allgemeinen [KPMN20, Hei84-48, Hei84g]. Allgemeiner [Hei58a, Hei84d]. Allgemeinverständliche [BDR85a, BDR85b, BDR86]. Allied [Irv67c, Log96, Nag16]. Allies [Hof93, Hol93b]. Alliierten [Hof93, Nag16]. Alltag [EHR^{+}96]. also [Ano29]. Alsos [Gou48b, GP92a, Gou47e, Gou47f, Gou47d, Gou48b, Gou83, Gou96, Mor47]. alt [Hei38g, Hei49a]. Alte [Hei61a]. Alteration [Hei82c]. alternative [Hau22]. Alternativen [Rad83, Rad91]. alternativer [Hau22]. alternatives [Rad83, Rad91]. Alwyn [Sei90a]. Am [Got04, May87, AH93, Har43, Hei31g, Hei35d, Hei41a, Hei41b, Hei42c, Hei43r, Hei53l, Hei53q, Hei64b, Hei72f, Hei86m, Hei89i, KPMN20, Wal05, Wal07]. Ambegaokar [HVTK^{+}97]. Ambition [Jon08]. Ambrosius [Wit83]. America [Sei90a, Stu85]. American [CS94, Hei53b, Sei90a, SPS^{+}64, CHW91, Gou47a, Hei86a, WC84]. americani [Ces18]. amerikanischen [Hei53b]. among [Hei46e]. Amy [Dje07]. anaesthetic [Wil99]. Analyses [Bar04, CD94, Mat86b, Mat86a, Mat87]. Analysis [BSP^{+}70, KED^{+}60, SPS^{+}64, WMG^{+}92, DKLN92, Hei89-33, KW76, Sem03]. Analysis/Bopp [SPS^{+}64]. Analysis/Mehlin [KED^{+}60]. analytical [Bes98]. Analyzed [Gou47b]. ancient [Hei37e, Hei52e]. and/or [Mil12]. Andenken [Hei43u]. anderen [OS03, vW41b]. Anders [Sei90a]. Anderson [Ano51]. Änderungen [Hei70b]. Anfange [Hei77e]. Angeles [Neu99]. Angeles/London [Neu99]. angeregten [BH23a, BH63a]. angewandte [Hei56n]. angewandten [Hei56j]. Angst [Hof93, Ano02a, Hei74c]. anharmonic [KW76, Tru75]. Anisotropic [BHX96, Bra97, DC87, JZ11, TBL02]. anisotropy [DKLN92, KGNS10]. Anlage [Hei79c, Hei79b]. Anlaß [Hei89b, Sch82]. anläßlich [Hei58k, Hei69b, Hei89a]. Anmerkungen [OS03]. Anna [Cas05]. annihilation [Hei52s]. Anniversary [Gin75, Hei58k, Hei74g]. Annual [Hei73a, IEE06]. anomalen [Hei22c, Hei24d, Hei25d, HJ26]. anomalies [Mas04]. Anomalous [Hen98, Hei22c, Hei24d, Hei25d, HJ26, Bre23]. anomaly [Gau14]. Anpassung [HW07, Hei63a]. ansatz [KW03, LFK99]. ansätze [GbK77]. Anschauliche [Hei36c]. anschaulichen [Hei27e]. Anschaulichkeit [Lac02a, Lac02b]. Anschauung [Hei10b]. Ansprache [Hei86b, Hei89b, Hei89a]. Answer [Boh62]. anthology [Hen96, HH11]. Antiferromagnet [Her69, DRB86, Man91, TBL02]. antiferromagnetic [CRR93, HWY69, HY72, MC92]. Antiferromagnetism [KI64]. antiken [Hei37e]. Anton [Fis10a, Fis12a, Fis10a, Fis12a]. Antonio [Pol00]. Antony [Sta11a]. Antragsvarianten [URWB95]. Antrittsrede [Hei86c]. Anwendbarkeit [Hei38f]. Anwendung [Hei25b, Hei53k, HJ26]. Anzeige [HB^{+}72]. Apfel [Büh98b]. Apologetics [BC95]. Apparent [LD66]. Appendix [Hei41e, Man11]. apple [Büh98b]. applicability [Hei38f]. Applicable [MWL^{+}11]. Application [Hei47g, Hei47h, Hei58b, Hei84-35, KSSA78, CL05, Hei25b, Hei53k, Man91, HJ26]. Applications [HEB^{+}80, HVTK^{+}97, LSK^{+}76, RFS^{+}91, TV23, Enz83, GKZL96, HW12a, Hei31c]. Applications/Eysenck [LSK^{+}76]. Applications/Kaufmann [HVTK^{+}97]. Applications/Lehmann [HEB^{+}80]. Applied [BSP^{+}70, Cop81, Hei56n, Hei56j, Hei56p]. Applying [Bes98, Har73]. Appraisal [Hei49f, HR02b, RW02]. Appreciation [Wig96b]. approach [BT96, BG96, Flo97]. Approximate [SMMM87]. approximation [Kle78, MRV24]. Apr [DDH89g]. April [DDH89g, Hei41b, Hei85b, Hei85b, Hen98]. Aquinas [Sil13]. arat [DDH41a, DDH41b, DDH89f, DDH89g]. Arbeit [Eic50, Hei24a, Hei31f, Hei34b, Hei36e, HA^{+}48, RDP^{+}95, Sch05]. Arbeiten [ARK^{+}99, Hah40, Hei41g, Hei42c, Hei46d, Hou30, Hei43f, Hei64a]. Arbeitsgemeinschaft [LSK^{+}76]. arbitrary [DPL94, DPL96, Mor72c]. archeology [PFH20]. Archimedes [Pic08]. archive [HHHH11, Aas02b]. Argon [Hei27b]. Argyris [MGG^{+}95]. Argyris/Faust [MGG^{+}95]. Arising [Kor05]. Aristid [Erm89]. Aristole [vW02, vWR04]. Aristoteles [vRFS^{+}00, vW02, vWR04]. Arkadiusz [Lic16]. armaments [Bro92]. arme [Kar07]. Armin [Mac78b, Rec86, Tre78]. armory [Wal05, Wal07]. Arms [Hei76b, Sei90a, vW57]. Arms-Control [Hei76b]. Army [BS47, Gou47b, Sei90a, Jon85]. Arnold [Han62, Hei39a, Hei43g, Hei43r, Hei86m, Huf71, Joh59, LSK^{+}76, Sch59, Sta11a, Hei39a, Hei43g, Hei43r, Hei48a, Hei51b, Hei51c, Hei86m, PS69, Set10]. arrangement [BFH^{+}42a, DDH41a, DDH41b, DDH89f, DDH89g, Hei41a, HBF^{+}89b]. arrangements [HJ43, Hei89i]. arrays [HBF^{+}42, HBF^{+}89a]. art [Hei70d, Hei75p]. Arthur [Adl03, Deh36, Cal14, Cap57a, Cap57b, Kra95]. Article [Ome79, Hei36i]. Articles [Rec84, Sei90a]. Artificial [GR09, Hah43]. Artikel [Hei36i]. artist [Hei83i]. Arts [Hei66b]. Ascent [Bro73a]. ASDEX [RDP^{+}95]. Aspects [EH38, Hei73b, PFH20, SW72, EH84, Hei72d, Meh75, PDL92]. Aspekte [PFH20]. assassinare [Ces18]. assassinate [Ces18]. Assessment [Rec81a]. associated [CRR93, Tor91]. astounding [Haw11]. Astronautenzentrum [DRF90a, DRF90b]. Astronomy [KED^{+}60, WMG^{+}92, FF91].

Astronomy/Born [KED^{+}60]. Astrophysics [WMG^{+}92]. Astrophysik [Hei62j]. asymptotic [Ter82]. Athens [Ano66]. Atmosphäre [Hei39h, Hei84-44]. Atmosphere [Hei39h, Hei84-44]. Atom [Cas79b, Gou47a, Gou49, Hei49f, JTD^{+}98, PSZM02, Rod19, Sei90a, Ser77, vM02, AK15, BH23a, BH63a, CHW91, CL94, Hah43, JLP^{+}02, LH01, Pow93a, WC84, vCH18]. Atomare [URWB95]. Atombau [Hei39a]. Atombombe [Bed03, Bed04, Brü64, Hei50e, Hei67a, NW95, Pop16b, Pop16a, Tel01, Hei48g, Hof93, Irv67a, Pow93a, Wal90b]. Atombombe“ [Wal93]. Atomdynamik [Hei27a]. Atome [DRF90a, DRF90b, Eck12a, Eck12b, Hen11, Jac11b, Rec09, Rec10, URWB95, KH25, Hof11, Jac11a]. Atome/Europäisches [DRF90a, DRF90b]. Atome/Neutrinooszillationen [URWB95]. Atomen [BH23b]. Atomenergie [Hei53k]. atomfizikáról [Hei83h]. Atomforscher [Hof23, Jun56, Jun64, Jun72, Jun90]. Atomforschern [BBF^{+}57b, Hei57e]. Atomforschung [Hei56j, KP07]. Atomforskernes [Jun57]. Atomgesetz [SPS^{+}64]. Atomic [ARK^{+}99, Bed03, Bed04, Bor58, BBSR69, Bør12, Car99a, Cas93a, Cas11, Cas17, Com56, DW52, Del13, Erm89, Ess71, FR13b, Gla00, Hah11, Hei27a, HSD34a, Hei36a, Hei47g, Hei47h, Hei56b, Hei58c, Hei68i, Hei70e, Hei84-35, Hen00, Hen10, Hof93, Jun58, Klo96, Lal15, Lan02, Lib68, Log99, Neu99, Pei71, Pei97, PS69, Pop16c, Pop21, Rho86, Sei90a, Sta11b, Sti47, Tur46a, Tur46b, Wal90a, Wal99, Wal17, Wal20a, Wal22, Wal24a, WHF^{+}96, vW57, Bet91, BDR84f, BDR84h, Boh64, Boh67, BBF^{+}57b, Bal07, Brü64, Car10, CCJ^{+}34, FR08, Geĭ36, Geĭ58, Gou47a, Gou48b, Hei26e, Hei35a, Hei35c, Hei37d, Hei38c, Hei38b, Hei39d, Hei40b, Hei43n, Hei46d, Hei47b, Hei48g, Hei48i, Hei49e, HW49, Hei50e, Hei51f, Hei51j, Hei51k, Hei52c]. atomic [Hei52d, Hei53k, Hei56i, Hei56n, Hei56p, Hei57e, Hei58g, Hei58h, Hei58i, Hei58u, Hei60e, Hei65g, Hei65h, Hei67a, Hei69d, Hei72e, Hei73c, Hei74k, Hei81b, Hei81c, Hei83e, Hei84j, Hei84f, Hei84l, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06, Hou30, Irv67b, Irv67a, Irv67c, Jon85, Jun56, Jun57, Jun82, Kae48a, Log96, Luc05, Luc16, Mac86, Mac87, Otr79, Pop16b, Pop16a, Ros98, SR92, Telxx, Tel01, Wal90b, Wal93, Wal95, Wal24b, Hei39a, Hei52a, Hei58d, Hei65a, Sei90a]. atomica [Hei60e, FR08, Hei36a]. Atómico [Bal07]. atomique [Gou48b, Luc05, Luc16]. atomiques [CCJ^{+}34]. Atomkeller [Fec90]. Atomkeller-Museum [Fec90]. Atomkern [Hei38b, Rec02a, BDR84f, Hei36d, Hei37b, Hei38c, Hei84l]. Atomkerne [BMG^{+}50, Hei32f, Hei32g, Hei33b, Hei35c, Hei43n, Hei47b, HW49]. Atomkerne/Räntsch [BMG^{+}50]. Atomkernenergie [Hei46d]. Atomkerns [BDR84h, Hei37d, Hei51k, Hei35a, Hei51j, Hei84f]. Atomkernspaltung [Hei89l, Hei43l]. Atomlehre [Hei58g]. atomnogo [Geĭ36]. atomnoj [Geĭ58]. Atomowa [Kar06]. Atomphysik [Hei39d, Hei86f, Ess71, Hei49e, Hei51f, Hei52a, Hei56i, Hei56n, Hei58d, Hei58h, Hei58i, Hei65a, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06]. Atomphysiker [Hof92, Otr79, Wal93]. Atompläne [Kra87]. Atoms [CR12, EHR^{+}96, Hei70j, KPS^{+}95, Sal19, BH23b, Eck12a, KH25, Rec10]. Atomskerns [Hou30]. Atomsystemen [Hei26e]. Atomtechnik [Hei53k, Hei53g, Hei89d, Hei89c]. Atomtheorie [Ben34, HSD34a, Wil36b, Hei75g]. Atomumwandlungen [Hah43]. Atomunion [Hei52c]. atomvåbnenes [Boh64]. Atomwaffen [ABH^{+}55, ABH^{+}87, Hei57b]. Atomwirtschaft [Hei56a, Rad83, Rad91]. Atomzeitalters [Hei72a]. atoombom [vC16]. atoomkern [Hei39c, Hei84j]. Attempt [dB58]. attempted [Bro92]. attempts [HW84b, Kar07]. auditorium [Hei60c]. Audretsch [RFS^{+}91]. Aufbruch [MGG^{+}95]. Auffassungen [Hei59k, LBBH69]. Aufgabe [Hei72f]. Aufgaben [Hei38e]. Aufl [Deh36, Deh37, Mos42, Str62b, Str62c, Haa34, Hei31b]. Auflage [Clu37, DS43, Hei53x]. Auflösung [URWB95]. aufmachen [Büh98a]. Aufnahme [Rec91a, Rec91b]. Aufruf [HA^{+}48]. Aufstieg [Rad83, Rad91]. auftretende [Hei38h]. Aug [DDH89e]. Augen [Büh98a]. August [BC95, Hei53q, HGW89b, WR92]. Ausbeute [HJ43]. Auschwitz [Wal24b]. Auseinandergelebt [RDP^{+}95]. Ausgabe [Hei53b]. Ausgewählte [Bor63a, Hei87a]. Ausländische [Hei68a]. Auslösung [Hou41]. Ausnutzung [Hei46d, Hei89w]. auspices [CCJ^{+}34]. Ausschließungsprinzip [Hei31k]. Ausschreibung [Ano84b, URWB95]. Außenwelt [Hei31g]. Aussichten [Hei63k]. Ausstellung [Rec02b, Ano84b]. Austausch [Hei71n]. Auswertung [Hei44c, Hei89-33, Hei43e, Hei89e]. Auswirkung [vW1 ]. Auswirkungen [Hei49b]. Auszüge [Hei69b]. authenticity [Bro09]. Author [Rec84, Dur00, Sch59]. authors [Bes04, SH57]. Autobiographical [Ano61b, Hei35g, Hei48m, BDR86]. autobiographischen [Ano61b]. Autobiographisches [BDR86]. Autobiography [Hei82b]. Automorphisms [VZ09, GS12]. Autonomie [HW07]. Autonomy [Sei90b, Hof05, HW07, HW12b]. Autour [Rhe91]. Auwärter [Ano84b]. aux [Per39]. Aver [Cho02]. averaged [DK78]. avoid [Hei89-40]. Await [Hei49f]. Award [Ano33, LW34, Ano29]. away [Hei89-40]. axis [DRB86].

B [BDR84b, Cru23, Kra86a, vM05b, Ano93a, BDR84e, CP93, GORS09]. B-VIII [GORS09]. B. [Hei43f, Hei48h]. B7 [BBF^{+}45, HBB^{+}89, BBF^{+}45, HBB^{+}89]. Bacciagaluppi [Sta11a]. Back [Hei27b]. background [Hei74j]. backscatter [BBF^{+}45, HBB^{+}89]. Backyard [Sei90a]. Bahnbrecher [Hei72a]. Bahre [Hei86l]. Baker [Kra86a, Mat86b, Rec85b]. balance [Hei89j]. ball [DDH42a, DDH89a, HBF^{+}89c, Sei90a]. ball-layer [DDH89a]. ball-tier [DDH42a]. Ban [Bad95]. Band [ARK^{+}99, BDR84c, BDR84d, BDR85b, BDR86, BDR89b, Eis00, HvMW79, MGG^{+}95, PvM93, Wit83, vMHW85, vM93, vM96, vM99, vM01, vM05b, vRFS^{+}00, Hei39a]. Bandwidth [CP84]. Bar [UM92]. Bar-Hillel [UM92]. Bargmann [BK88]. Barlow [HvL47]. Barlowsche [HvL47]. Barnsley [Lic16, Jad14]. Barr [Dje07]. Barth [Wit83]. base [SME77]. based [Ano11, Haa34, NKD^{+}11, Tiw14]. Basel [Kra86a]. Bases [RS94]. Basic [Los61, Hei27f, Hei39b, Hei58g, Hei56n, Hei59b]. basics [HW12a]. basis [BK88, Hei42e, Hei89n]. Bass [OSWR86]. Bateman [Hei32a]. battle [Hei24b]. Bau [Hei32f, Hei32g, Hei33b, Hei72l]. bauen [Hei67g, Hei67a]. Bausteinen [Hei53o]. baute [Bed03, Bed04]. BCC [Mor72c, TBL02]. Bd [HEB^{+}80, Hei43g, Hei43i]. Be [Hei79b, Büh98a, Hei89-40, Pop16c, SK09]. Beams [Ano84b, JTD^{+}98]. Beams/Bericht [Ano84b]. Beauty [Hei79a, Ome79, Smo93, vRFS^{+}00, BDR85b, Hah10, Hei53r, Hei71d, Hei79e, HHvB10]. Becker [DT18, Hei86l]. Beckers [Hei86l]. Become [Kae31]. becomes [Gla00]. Bedeutung [BDR85b, Hei71d, MGG^{+}95, Smo93]. Bedingung [Mül40]. Beeinflussung [Hei25a]. Been [Kae31]. before [Han79]. Began [Kae48b, FS93]. begann [FS93]. Begegnungen [Hei75e, Hei77c]. Beginning [Mil85, Hei77e]. beginnings [Hei83c]. Begivenhederne [Ano44a]. Begleitbuch [Rec02b]. Begriff [EHR^{+}96, Hei48e, Hei48d, Hei71c]. Begriff/Vogel [EHR^{+}96]. Begriffliche [HVTK^{+}97]. Begründer [Kle17]. Begründung [BHJ62]. Begrüßung [Hei64b]. Begrüßungsansprache [Hei69b, Hei89u, Hei89f]. Begrüßungsworte [Hei73a]. Behandlung [Hei89k]. Behavior [BHX96, Hei48c, TBL02]. behind [Pic08]. Behrens [SCH^{+}78]. beide [Büh98a]. beim [Döp42, Hei24d, Hei72l]. Beisetzung [Hei86b]. Beiträge [Hei05b, KRW05, KED^{+}60, Rür79, SPS^{+}64]. Belief [Hab83, Hei35b]. Bell [DGH^{+}12]. Bellevue [Bed09]. Bellman [KED^{+}60]. below [Mac85]. Bemerkung [SH22b, Hei24a]. Bemerkungen [Hei34b, Bor61, Bor63b, Hei31a, Hei34a, Hei35a, Hei42b, Hei48b, Hei52b, Hei56e, Hei56d, Hei72b, Hei75f, Hei84e, Hei84f, Hei89g]. Ben [Jor78]. Benchmarking [BPP11]. Bending [Klo96]. benennen [URWB95]. benennt [RDP^{+}95]. Benz [LSK^{+}76]. beobachtbaren [Hei43m, Hei43j, Hei44a]. ber [vW1 ]. Berechnung [HJ43]. Bergakademie [DRF90a, DRF90b]. Bergen [Ano29]. Bericht [Ano84b, BFH^{+}42b, Hei41e, Hei64b, Hei89x, Hei89-34, HBF^{+}89c, RDP^{+}95, vW1 , BBF^{+}45, Döp42, DRF90a, DRF90b, Hah40, Har43, Hei30a, Hei40a, Hei41a, Hei84-48, Hei84g, Hei89h, Hei89i, HBB^{+}89, vW41a]. Berichte [Hei05b, KRW05, LSK^{+}76]. Berichtigung [Hei34b]. Berkeley [IEE06, Kra86a, Kra07, Lan02, Lan08, Mat86b, Neu99, Ros79]. Berkeley/Los [Neu99]. Berlin [Cas79a, DRF90a, DRF90b, Eis00, Har43, HMW^{+}84, Hei89i, Hen86, Kra86a, Rec85a, RDP^{+}95, Rür79, URWB95, Cas79a, Har43, Har60, Hei41a, Hei43p, Hei56g, LMOS15, Rür79, Sch15]. Berlin-Dahlem [Hei89i, Hei41a, Hei43p]. Berlin/CERN [DRF90a, DRF90b]. Berlin/Fraunhofer [URWB95]. Berlin/Fraunhofer-Gesellschaft [URWB95]. Berlin/Heidelberg/New [Eis00]. Berliner [DRF90a, DRF90b, Hei56g]. Berlins [LMOS15]. Bermant [SPS^{+}64]. Bernard [Hei72a, Men84]. Bernstein [Got04]. Bertrand [Bro72, Mar73]. Besættelsen [Ano44a]. besøk [Aas02e]. Besprechung [Hei25a, Hei27b, Hei27a, Hei28a, Hei28b, Hei30b, Hei31b, Hei31c, HPE^{+}33, Hei35b, Hei36b, Hei36a, Hei36c, Hei39a, Hei39b, Hei43g, Hei43h, Hei43i, Hei50a, Hei50b]. Besprechungen [Har43, HvLM^{+}35, HJFS36, HSBR36, HGT^{+}39, HSBB43, HF46, HHKG50, KHBF43, MPB^{+}24]. BESSY [URWB95]. BESSY-II [URWB95]. best [Rec02a]. Bestandsaufnahme [GHH02]. besteht [URWB95]. Bestimmung [DDH40a, DDH40b, DDH89e, DDH89d, Fis41]. beszélgetések [Hei83h]. Beta [GR09]. Beth [Bor58]. Bethe [Bet93, KW03, LFK99]. Betrachtungen [DB71, Hei48k, HBD71]. Betreuung [Hei53d]. Betriebsunfall [Hei72g]. Better [Sei90a]. Between [Hei76c, Hof05, Lan08, Wüt16, Bok04, BB71a, BB71b, BB05, BM94, Bro72, Bro09, CM02, DC48, Dör05, Fis15, Gou90, Hei53e, Hei75m, HW07, HW12b, KF02, Kra07, Lau07, LTG^{+}12, LD66, NRS71, Sal01, Sch13, SH57]. Beugungsversuche [Hei36b]. Beute [Nag16]. Bevollmächtigte [Mai07]. Beweise [URWB95]. Bewerber [Ano84a]. Bewertung [Hei43k]. Bewußtseinswandel [vW88, vW91]. Beyerchen [Jor78]. Beyond [Bok08, Bun71, Cas09, Dje07, Hei70e, Hei84-55, Hei00a, Hei08c, Huf71, Rod19, Ano11, BW04, Geĭ87, Hei71i, Hei71j, Hei04, NKD^{+}11, Bed09, Sap10, Sch11a, Sta11a, Kea09]. Beziehungen [Hei53e, Hei75m, Hei25c]. Bibliography [DW52, Kra86a, Rec85b, Rec93, CB84, Cas01, Mat86b]. Bifurkation [MGG^{+}95]. Big [Hei72a, Hür22, Rec01]. Bigelow [HVTK^{+}97]. Bilddokumenten [Her01, Her76c, Her84]. Bilder [Tel93]. Bildung [Hei69k, Hei69l, Hei89-30, Hei49h]. Bildungswärme [Hei24c]. Bill [HBF^{+}89c]. Binary [Gre63]. Biographical [CS94, CR84, OS98, Sei90a, BDR86]. Biographie [Eck12b, Hen11, Rec09, Rec10, Eck12a]. Biographies [FW20]. Biographisches [BDR86]. Biography [Eck12a, Gar91, Sei90a, Wal20a, Wig96c, Pei92b, Rec10]. Biological [Cop81]. biquadratic [PDL92]. Birkhäuser [Kra86a]. Birkhoff [KJ06]. Birth [BH83a, BH83b, Gin75, BD92, BH82]. Birthday [Hei44b, Hei54e, Hei64h, HBD71, Jac01, KRW05, Rec01, RH02, Wig96b, vM02, DB71, Fue61, Hei35d, Hei43r, Hei58l, Hei62i, Hei72c, Hei86n, Hei89-37, Hei05b, MF69, PRW55, PFH20, Rec02a, Rec02b, Sch82]. Birthwistle [Hei28a]. bis [BDJ57, Büh98b, Cas79a, Cas05, Fis10a, Fis12a, Hau22, Hei72a, Hei30a, HHH03, Hen11, HH14, Rec09, Rec10, vRFS^{+}00, vW02, vWR04]. Bisher [Ano84a]. Bisherige [Hei79c, Hei79b, Hei85b]. bisherigen [Hei38f]. bisherighen [Hei84-48]. Bislang [RDP^{+}95, HHHH11]. black [Bes04]. Blätter [Ano84a]. Blätter/Bisher [Ano84a]. Blaue [URWB95]. Blocks [Ano43, Hei53o]. Blokhintsev [Hei56l]. Blomeyer [Hei53b]. Blum [Bro85, Bro86, Kra86a, Mat87, Sch91b, TW91]. Board [Hei64b]. Body [Hei89k, Hei27c, vV78, Hei26a]. Bohmsche [GRE^{+}01]. Bohr [Aas02b, Adl03, Ano02a, ARK^{+}99, Bor58, Bro95, Car97, Eis00, Fay07, Fer07, For09, GRE^{+}01, HEB^{+}80, Hei35d, Hei63h, Hei86g, Hen86, HvMW79, Kra96, Kra07, Lak96, Lan72, Lan08, Lau07, MGG^{+}95, OSWR86, Par06, PHvMW01, PvM93, Paw02a, Pye81, Rec86, Sch11c, Tur06, vMHW85, vM93, vM96, vM99, vM01, vM02, vM05b, Aas00, Aas02a, Aas02b, Aas02d, Aas02c, Aas02e, AK15, Ano02a, Bel92, Ber95, BR64, BH23b, Bro71, Cam07a, Cav14, CG14, Dah02a, Dah02b, Dir28, Dör05, Fav05, FK85, GLSC01, Got02a, Hei28c, Hei35d, Hei62a, Hei86g, Hei10a, HvMW79, Hol05, Lan72, Lin07, Lin08, Los61, Mas04, Meh87, Pai91, PRW55, PHvMW01, PvM93, Paw02a, Plo10, Pow02, Rhe91, RRK^{+}64, Roz67, Sch00b, Shi83, Shi86, Tan04, Telxx, Whi96]. Bohr [Yam02a, vD15, vM02]. Bohr-Dokumente [Paw02a]. Bohr-Heisenberg [Dah02b]. Bohr-Sommerfeldschen [vM02]. Bohrende [Paw02a]. Bohrs [Hei10a, Hei62a]. Bohrschen [BH23b, Hei75g]. Bomb [Ano93a, Ano02a, Bed03, Bed04, Ber92, BC95, Ber95, BL03, Bet93, Bro93, Car99a, Cas93b, Cha92, Cho02, Gla00, Gol92, Gol95, Gou47a, Hei68i, Hei93f, Hei95, Hen00, Hof93, Lan02, Lau58a, Lib68, Log93, Log99, Luc11, Neu99, Pei93b, Rho86, Sch01, Sei90a, Sti47, Wal99, Wal20a, Wal24a, WHF^{+}96, Ano99, Bal07, Brü64, Cas09, Erm89, FR08, GP92a, GP92b, Hau22, Hei48g, Hei48i, Hei50e, Hei67a, Him90, Irv67b, Irv67a, Jon85, Kae48a, Kar05c, Kar05b, KW05, Kar07, KP07, LS93a, LS93b, Log96, Luc05, Luc16, Lur95, Mac85, Mac86, Mac87, NW95, Pei92a, Pop16b, Pop16a, Pow93a, Pow93b, Pow94, Pow00, Ree24, Ros98, San02, SR92, Telxx, Tel01, Wal90b, Wal93, Wal95, Wal24b, Wel92, Zan02a, Zan02b, vCH18, vS14, Pop16c]. Bomb [Bed09, Rec91d, RW92, Sta11a, Wal90a, Wal17, LRD^{+}91, Ber93, Hol93a, Klo96, LH93, Sap10, Sei90a, Swe93, Wil94, Cas93a]. bomba [FR08, Kar06]. Bombe [All06, Ano02a, Cas02a, Hau22, Kar05c, KP07, Luc05, Sch01, Kar07, Luc16, Lur95, Him90, Kar05b, vS14, Hof02]. Bombs [Klo93, Pop21, Wal22, Sei90a]. bond [Hea11]. Bonhoeffer [BMG^{+}50, Hei86b, Hei86b]. Bonn [HR^{+}89, BBH^{+}62, Hei53q, HR^{+}89]. Bonner [BBH^{+}62]. Book [And89, Ano54, Ano68, Ano94, Bed09, Ber93, BDR86, Bok10, BB96, Bor58, Bor53, Bro72, Bro85, Bro86, Bro95, Bro97, Bun71, Car97, Car99a, Cas74, Cas93a, Cas97, Cas02a, Cas03, Cas05, Cas11, DT18, Dje07, Dre92, Fay07, Fel54, Fer07, For09, Fra94, Fra10, Gar91, Gre10, Hab83, Hay90, Hei72a, Hei32a, Hei75g, Hei93b, Hei93d, Hen86, Hen11, Hol93a, Hol00a, Huf71, Joh59, Jor78, Kle54, Kra86a, Kra96, Kra07, Lan02, Lan08, Lau07, Lev95, Lib68, Log99, Mac78b, Mac58, Mar94, Mar73, Mot85, Neu99, Par06, Pau31, Phi92, Pol00, Pye81, Rec85b, Rec86, RS11a, Roh85, Rud58, Sap10, Sch59, Sch91b, Sei90a, Sta11a, Sta11b, Str62a, Swe93, TW91, Tre01, Tre78, Tur06, Wal08, Wat01, Wei85, Wig67, Wil94, Rec02b]. Book [Ess71, Hei25a, Hei27b, Hei27a, Hei28b, Hei31b, Hei31c, Hei35b, Hei36b, Hei36a, Hei36c, Hei39a, Hei39b, Hei43g, Hei43h, Hei43i, HF46, Hei50a, Hei50b]. Books [BDR84e, Hei54a, LH93, Rec84, Sei90a, AFG90, BDR84b]. Boorse [Sei90a]. Bopp [Hei89-37, SPS^{+}64, Hei89-37]. border [Sch05]. borders [Bei13, Ben13, Hei71l, Hei73g, Hei76g, Hei77h, Hei84-62, Hei89-29]. Born [Bro72, Hei72a, Hei27a, Hei62i, Hei71h, KED^{+}60, Mar73, Sch05, Sch07, Sch26, Sch66, Fue61, Bac22, BB69, BB71a, BB71b, BB05, Bro72, Gal01, Geĭ70a, Gre05, Hei62i, Hei70g, Hei71f, Sch05, Sch07, dG14]. Borna [Geĭ70a]. Borries [BMG^{+}50]. Bosković [Hei59f]. Boston [Kra86a, CW85]. both [Büh98a]. Bound [DeG74, Kle78, SZ73, SZ74, Tru72]. Boundary [Dun86, Ni04a, Ni04b, ZG96]. br [Ben35, Hei52f]. Br. [Hei52f]. bracket [Geĭ58]. Bragg [CS94]. braided [BM93]. branches [Hei39d, Hei40b]. Brandmüller [MGG^{+}95]. Brandt [SPS^{+}64]. Braucht [Hei68b]. Braun [LSK^{+}76]. breaker [Log93]. Breakthroughs [Cre10, Gre10, Lig05]. Brennan [Bro97, Cas97]. brev [Sch15]. Brewster [Jor78]. Brief [Hei56a, HGW^{+}89a, Hei89-30]. Briefe [Cas05, Hei79c, HHH03, Hei05b, Hof23, Kan03, KRW05, Som36]. Briefwechsel [ARK^{+}99, BB69, Car97, EB82, Eis00, GRE^{+}01, HEB^{+}80, HHHH11, Hen86, HvMW79, Kra96, MGG^{+}95, OSWR86, PHvMW01, PvM93, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05b]. Brighter [Jun58, Jun82, Jun56, Jun57, Jun64, Jun72, Jun90, Lau58a]. Brillouin [Hei31c]. Bringing [Stu85]. Bristol [Wat01]. British [Jon78]. Broadway [Sch00b]. Brod [MGG^{+}95]. Broglie [Ano35b, Cas75, Deh36, Haa34, RFS^{+}91, Ano35b, DC48, Haa34, Loc07, Mil58, PSZM02]. Broken [Dje07]. bronia [Kar06]. Brooks [CS94]. Brosch [Ano08]. Brosche [HEB^{+}80]. Brownian [Isa82]. Brückner [RFS^{+}91]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34, Far01]. Buchbesprechung [Ano35a, Ano06, Deh36, Deh37, HVTK^{+}97, Hei30c, Jac03, Mos42, RFS^{+}91, Wil36a, Wil36b, Wit83]. Buchbesprechungen [BDR86, EHR^{+}96, MGG^{+}95, OSWR86, WMG^{+}92]. Bücher [Ben30, Ben34, Ben35, Clu37, ARK^{+}99, BMG^{+}50, BDR84b, BDR84e, BSP^{+}70, Büh95, GRE^{+}01, HEB^{+}80, HMSH78, Jac11b, KED^{+}60, LSK^{+}76, SCH^{+}78, SPS^{+}64, vRFS^{+}00]. Build [Gol92, KH19, Hei67a, Hei67g, SR92]. Building [Ano43, Gol95, Hei53o]. Built [Bed03, Bed04]. bulk [Eck70]. Bulletin [FR98]. Bunde [MGG^{+}95]. Bundesregierung [Hei53d]. Bundesrepublik [Hau22, Hei65d, Sch05, Sch07, Hei57d]. Bundestag [Hei53q]. Bundestages [Hei53q]. Bunin [Ano33]. burdens [Hol86, Hol98]. burner [HW84b]. Burns [Sei90a]. Bush [Gol92]. bust [Hei86o]. Büsten [Rec91a, Rec91b]. busts [Rec91a, Rec91b].

C [Adl03, Ano94, Bed09, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89b, Bor53, Büh95, CS94, CD94, Dre92, Fra94, Gar91, Kea09, Kra86a, Mat86a, Phi92, Rec03, Sap10, Sei90a, Sta11a, Swe93, Tho97, Wei85, BDR84c, BDR84d, BDR85b, BDR86, BDR89b]. C. [Erm89, Hah06, Hah11, Hei72c]. CA [Kra07]. Calcul [Per39]. calculation [Bec02, KW02, KW03, Liu72, McC24, HJ43, Per39]. calculational [AMS04]. Calculations [LD66, El 00, TR09]. calculus [DLW08]. Calder [BDJ57, BDJ57]. California [Kra86a, Lan02, Lan08, Neu99, Ano34, IEE06]. Callaghan [WMG^{+}92]. called [Hei89j]. cambiato [Lud01]. Cambridge [Bok10, Cas11, Hay90, Par06, Sta11b, Tur06]. came [Log96]. Camilleri [Bok10, Eng09, Fra10, Sta11b]. Campaign [Bør12]. can [Mas04, SK09, Hei79b]. Canham [CS94]. Canonical [Car71, KF07]. Canonically [Low97]. Cape [Cas93a]. Cappellari [Kra86a, Mat86a, Wei85]. caractéristique [DC48]. carbon [Hei27b]. care [Hei49g]. Carl [Ano02a, BMG^{+}50, GRE^{+}01, Hei02, Rec02a, Ano02a, Cas14a, Cas15, Eck70, Hei02, HH14, KF01, Rec02a, vWL93]. Carlo [HN99]. Carnap [LSK^{+}76]. Carolina [BCEP94]. carried [Hei89-40]. carrying [Hei41g]. Carson [Cas11, Hah11, Hen10, Sta11b, Lal15]. Cascade [Hei43b]. Case [Gou47b, Han79, Joh13, Jor78, SS16, Wüt16, dR99, Gau14, Wis98]. Cassidy [Ano94, Bed09, Büh95, CS94, Dre92, Fra94, Kra86a, MGG^{+}95, Mat86b, Rec85b, Rec03, Sap10, Sta11a, WMG^{+}92, CD94, Gar91, Kea09, Swe93, Phi92]. cast [Ano12]. Cat [Bru80]. Catastrophe [LH93]. Catherine [Bar04]. Cathryn [Lal15, Sta11b, Cas11, Hen10]. causal [Hei58c]. Causality [Bor58, Hei51m, Hei52a, Hei30h, Hei30g, Hei31e]. cause [Hei32h]. Cave [FTF01]. Cayley [Tem91]. CD [Hah10]. CD/DVD [Hah10]. Celebration [KRW05, Hei89a]. Celia [FB00]. Centenary [BCEP94, FK85]. Centennial [BW04, Hei04, PS69]. center [CZ81]. central [BP75]. Centre [Meh73, RDP^{+}95]. centrifuge [HJ43]. Century [Bro72, TW10, Bra09, Hei75k, Lig05, Meh73, Sch96a, Bre97, Hei75a, Bro97, Cas97]. ceremony [Hei64b, Hei86m, Hei89b, Hei89u, Hei60c]. CERN [DRF90a, DRF90b, Hei72f, Hei72f]. Certain [AP22, DM99, GV86, Gre05, Wre75]. certificates [HR02b, RW02]. certifications [Her76c, Her84, Her01]. Cf [DK78]. Ch [Ano08]. chaîne [Per39]. Chadwell [Sei90a]. Chadwick [Ano51]. chain [BG96, Bec02, DKLN92, DC87, DGS99, DL08, Hou41, KSSA78, KKM^{+}09, LFK99, Lun84, Per39, PDL92, Wan98]. Chains [Unr72, CRR93, DM99, DMCF06, JKS13, MC92]. Chairman [Hei64b]. Challenge [Sta09a]. Cham [Cru23]. Chamber [Hei53q]. champ [Hei67k]. Chance [Hei51e]. Chandra [LJ08]. Change [Hee75, Gau14, VB15, vW88, vW91]. changed [Luc04, Lud01, VB15]. Changes [Str62a, Hei34i, Hei34f, Hei34g, Hei34h, Hei35f, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei49l, Hei59i, Hei70b, Hei73j, Som36]. Chaos [BKBS93, Cap01, EHR^{+}96, MGG^{+}95, Sch91a, DKLN92]. Chaos/Plaschko/Brod [MGG^{+}95]. Chaotische [MGG^{+}95]. characteristic [DC48]. characteristics [KF07]. Charge [Sch13, Hei34e]. Charles [HVTK^{+}97, Mar94]. charms [Ber91]. Chast [Geĭ90]. chastits [Geĭ77c]. chastitsa [Geĭ56]. che [Lud01]. Cheating [Ano04a]. chef [Ost11]. Chem [BMG^{+}50]. Chemical [BSP^{+}70, BH24a, BH63b, Hea11, Hei24c, Hei41b]. Chemical-Physical [Hei41b]. Chemie [BSP^{+}70, Hah40, Hei53e, KED^{+}60, MGG^{+}95, Sim07]. Chemie/Dictionary [MGG^{+}95]. Chemie/Fletscher [BSP^{+}70]. Chemielehrer [Ost11]. Chemisch [Hei41b]. Chemisch-Physikalische [Hei41b]. chemische [BH24a, BH63b, Hei24c]. Chemistry [Joh13, Hei53e, Ost11, Hah40, MGG^{+}95, Sim07]. Chemists [Joh13, Har60]. Chevalley [Bar04]. Chief [RS11a]. Chimäre [Hau22]. chimera [Hau22]. Chimerical [Bru80]. Churchman [KED^{+}60]. Ci [Arg09]. Ciencia [FR04]. circle [BR64]. Citizen [Sei90a]. City [Hei74g]. cksprache [vW1 ]. Claim [Gou49]. class [NRS71]. Classical [BHN98, Boh62, Bok08, Cam07b, Bes04, Bok04, Isa82, JLP^{+}02, LH01, MP92, PDL92, TBL02, Wre75, Yeh70]. Clerk [Bes04]. Climate [Gol92]. cloak [BBF^{+}45, HBB^{+}89]. close [SAR^{+}09]. Closed [Bok06, Bok04, Hei48e, Hei48d, Hei74i, Hei83d, LFK99]. Closing [Hei50g, Hei53q]. cloth [Bed09, Cas05, Cas11, Neu99, Par06]. Cloud [Sei90a]. club [Ber96, For22]. cluster [GbK77, KSGB^{+}83]. cm [BBF^{+}45, HBB^{+}89]. Co [Ano94, Fra94, RRK^{+}64]. co-workers [RRK^{+}64]. Code [Pag82]. codes [BPP11]. Cohen [Men84]. coherent [BK88, CM02, KSGB^{+}83]. Coin [FR13a]. Coined [Gal01]. Cold [Sal19]. collaboration [BLW19, Dür05, vM05a]. Collapse [CK02b, CK02a]. colleagues [BR64, RKW05, Roz67]. Collected [BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR84e, BDR85b, BDR86, BDR89b, BDR93, Bro85, Bro86, Kra86a, Mat87, Sch91b, TW91, Wig96m, BDR89a]. Collective [HEB^{+}80, Hei65i, Hei84z, SME77]. college [Hei72j, SM41]. collide [Wal08]. Collisions [Hei84-28, Hei55e, Hei56l, Hei83g, JD06]. Colloquium [CW85, UM92]. color [Hei41d]. colour [Hei52o]. Columbia [Ano29]. Combs [Ano04a]. comedy [Dür62, Dür80, Dür06]. Commemorating [Gin75]. Commemoration [Her76a, Hei69f, Hei71f]. Comment [Got04, Rab48]. commentaries [BB05, Bro72, Yan13a]. Commentary [Got83, Yan13b, Cru23]. Comments [HH75, Hei75h, HS81, OS03, Hei69i, Rob82, Bor63b, Hei42b, Hei56d, Hei72b, Hei84e, Hei84f, Hei89g]. Commission [Sei90a]. Committee [Sei90a]. Communicating [vD15]. communication [Hei28c]. communicazione [Hei28c]. communists [Bro09]. Community [Hab69, Jor78, Bey77, Rin69, Rin90, Mai07]. Como [And89]. compact [Wan98]. comparative [SW05, Hei51l]. Comparison [Los61, Eck33]. Compartmentalization [Gol95]. Compendium [GHW09]. complement [Hei85d]. complémentarité [DC48]. Complementarity [Bel92, Cam07a, DC48, vD15, Meh87]. complementary [Loc07]. Complete [Hei68j, Hei73e, Hof23]. Completed [Hei79b]. Completion [Hei70a, Hei71a]. Complex [RFS^{+}91, LM87]. complexes [OD03]. Complexity [vRFS^{+}00]. component [Hei38d, Hei41c, Hei84n, Isa82]. components [SH22a]. composed [DDH89a]. composition [Hei39c, Hei84j]. compressible [DGS99]. Compton [Wil84]. computation [Ras78]. Computations [SK73]. Computer [DRF90a, DRF90b, IEE06, Wal08]. con [Hei84-55, Hei00a, Hei08c]. concentration [CP84]. Concept [BR96b, Hei65l, Car76, Hei51i, Hei69c, Hei71c, Hei82c]. ConceptDraw [GRE^{+}01]. Conception [Han62, Hei60g, Joh59, Sch59, Hei58r, Hei70i, Meh73]. Conceptions [LBBH69]. Concepts [Boh62, Cam07b, Hei75a, Hei75j, Hei75b, Ome75, Str62a, GHW09, Hei73d, Hei75l, Hei75k]. Conceptually [dlP80]. Concerning [BCL^{+}69, Hei37a, Hei58n, Hei65n, Hei84t, Hei84-38, Sno49, Jes98]. conciencia [FR04]. Concluding [Wig96d]. conclusion [Log96]. Condensed [ARK^{+}99]. condition [Mül40, Ter82]. conditions [Hei59b, MF69, Mon04, Per39, Tru72]. cones [Pac88]. Conf [BSP^{+}70].

Conference [BCEP94, Hei75s, RDP^{+}95, AK15, CDvW75, CDvW79, GZ95, Far01, Sta11a, BV09]. conferences [Meh75]. configurations [Woo77]. Conformist [Jor78]. Congratulations [Hei86g]. Congress [Hei64b]. Congruence [Nel72]. conjecture [LN06]. Connected [Hei34e]. connecting [dB58]. connection [BM94]. Connections [Ano60]. Conrad [MGG^{+}95]. conscience [FR04]. consciousness [Wal08, vW88, vW91]. conseil [CCJ^{+}34]. Consequences [HE06, Hei89v, HE35]. Conservation [GLR14]. Considérations [Hei34c, Hei84b, Hei32e, Hei34c, Hei48k, Hei72l, Hei84b]. conspiracies [Cas93a]. constant [BK04, Bro77, DHYY65, Hei45a, Mie80]. constants [BH24a, BH63b, DH65b, Hei24c, OD03]. Constitution [Gam30]. constrained [Lun84]. Construction [HW48, Hei72d, Hei73b, CM02, Hei72l, HW84b]. contemporanee [Cas14b]. Contemporary [Hei58v, Cas14b, Hei53c, Hei55b, Hei56f, Hei57c, Hei58o, Hei58e, Hei59a, Hei60a, Hei62b, Hei63b, HG65, Hei66a, Hei68c, Hei70c, HnH72, Hei74d, Hei74m, Hei76e, Hei79d, Hof23]. content [Hei27e, Hei83b, Oza03]. contents [Hei77j]. Context [Pan02, Hau22, HW11]. Contextualized [Hen10]. Contingency [Lev95, Cus94]. Continual [MP92]. continued [DR92]. Continues [BSMB99]. Continuous [BMP87, BM94, Lun84]. continuum [DKLN92, Isa82, PDL92, Ter82]. Contracting [Lan59]. contraction [ES95]. Contribution [Hei32b, Hei84c]. Contributions [KRW05, Rür79, Hei05b]. Control [Hei76b]. controversial [Eck01]. Controversy [Gou90, Rad83, Rad91]. conversas [Hei96c]. Conversation [Hei67h, Hei67i, Hei67g]. Conversations [Hei70e, Sch01, Hei71e, Hei71i, Hei71j, Hei75e, Hei77c, Hei83f, Hei96c, Bun71, Huf71]. cooling [SAR^{+}09]. Cooperation [Sal19, Hei56i, Hei89p]. coordinates [KF07]. Copenhagen [And89, Kra07, Lan08, Lev95, Wal08, Aas00, Aas02c, Ano02b, BJN01, BSMB99, Cas00, Cas02b, Cus94, Dör05, Fra98, Fra00, FB01, Fra05, HB63, Hei89r, Hei91b, Hei07a, Kar05a, Lau07, OS03, Paw02a, Pow03, Wal19, vW41a, Dje07, BW00a, BW00b, Hol00a, Ros00, Sch00b]. Copenhagen. [Log00]. Copernicus [Gin75, Hei86n]. Coral [BSP^{+}70]. Core [Cas79b, Pan02, SHI73, JD06]. Core-Context [Pan02]. Cornelius [BCEP94]. Corporation [Ano29]. corpuscles [BDR84f, Hei36d, Hei37b, Hei84l]. corrected [Pop16c]. Correction [Hei34b]. Corrections [Ano93a]. correctness [Hei73e, Hei83d]. Correlation [Fal64, HS70, LL86, May87, Mor71, Mor72b, DC48, Mor72c, Pea76, SL70, DC48]. Correlations [AG88, vRFS^{+}00]. Correspondance [Mar73]. Correspondence [BB69, Bro72, Bro95, Car97, Dah02a, Dah02b, Eis00, HH17, Hen86, HvMW79, LS93b, Rec86, vMHW85, vM93, vM96, vM99, vM01, vM05b, BB71a, BB71b, BB05, HHHH11, JLP^{+}02, LH01, PHvMW01, PvM93, EB82, Hei25b, SH57]. Corrigendum [Ano25]. cose [Cap01]. Cosmic [EH38, Hei39f, Hei39g, Hei39h, Hei43o, Hei46a, Hei51l, Hei53u, Hei75i, Hei75s, Hei76d, Hei77d, Hei84h, Hei84y, MGG^{+}95, Pag82, RFS^{+}91, BDR84g, EH84, GE38, GE39a, GE39b, Hei32h, Hei32e, Hei36g, Hei36j, Hei37a, Hei38d, Hei38a, Hei39j, Hei41c, Hei43r, Hei53t, Hei84n, Hei84i, HW84a, Hei84-44, Rec02a, BBB^{+}84, Hei43p, Hei51f, Hei53l, HMW^{+}84]. cosmic-ray [Hei37a]. Cosmological [ARK^{+}99, CBS20]. Cosmology [Hei68d]. Costandina [Sei90a]. could [Hei67g]. Council [HR53, Hei53q, Hei64b, HR^{+}89, Hei89-38, Eic50, Eic53, Hei49g, Hei50c, Hei50g]. Countdown [Sei90a]. counter [Irv67c]. counter-measures [Irv67c]. coupling [Bro77, DH65b, Hei27b]. Courier [Hei60d]. Course [SPS^{+}64, Luc04]. CPU [BPP11]. Cracow [Wal20b]. Crafting [Set10, Sta11a]. Crazy [Arr06]. Crease [Gre10]. Creation [Hei34e, Oor09, Žup65, Wal93, FS93]. Creator [Dje07, Ful15]. creators [Str11]. crimes [Wei46, Wei99]. Crisis [Hei85a, Rad83, Rad91, Cas76, Rin61]. criteria [Hei73e, Hei83d]. Critical [BHX96, Log96, Rec81a, Sta09a, Ber02, Cas05, El 00, HHH03, Kan03, McC24, TBL02]. criticism [Geĭ70b, Hei89-35]. Croatian [Hei97]. croix [Luc05]. cross [DK78, Fis41, Luc05]. cross-sections [DK78]. crossers [Sch05]. Crossing [Bei13, Ben13, Hei24b]. Crossroads [Sta11a, BV09, ET07]. Cruz [Sei90a]. Cryptography [BBE92]. crystal [KW76]. crystals [Hei33c, Hei84-41, Jau93]. cuántica [FR08, HBS96]. cuánticas [W^{+}87, WdC02]. Cube [KH19]. Cuestiones [W^{+}87, WdC02]. Culture [CS94, Hen00, Lan02, Cas14b, PFH20, Ros98, Neu99, Car99a, Log99, Wal99]. cultures [Cas14b]. Cummings [IM95]. CuO [VBN95a, VBN95b, vBN95c]. cuprous [Man91]. cura [Pol00]. Curies [Kae48b]. curious [Gau14]. current [Hei38e, Hei53t, Hei55c, Hei61b, Hei84-48, Hei84k, Hei84m]. Curriculum [Hei86d, Hei86h]. curvature [Mie80]. Curve [Gam30, Hei31l]. Cushing [Lev95]. Cut [vBW88]. Cyclotrons [Row09]. Czech [Hei69i]. Czechoslovak [Hei89y].

D [BBF^{+}45, Dje07, DDH41a, DDH42a, DDH89a, DDH89c, DDH89f, FS93, HBB^{+}89, Jac03, Jor78, Rec03, Sei90a, Ano44a, BBF^{+}45, DDH41a, DDH42a, DDH42c, DDH89a, DDH89c, DDH89f, Hei42b, HBB^{+}89, Hei89g, Hen93]. D. [FS93, Hei56l, Wig96j]. D2O [Hei42b, Hei89g, DDH42c]. D2O-38-Metallschichtensystem [DDH42c]. Dahl [Wat01]. Dahlem [HMW^{+}84, Hei89i, Hei41a, Hei43p, Hei44c, Hei89-33]. Dahlem-scale [Hei89-33]. Dahlemer [Hei44c, Hei89-33]. Dall’ordine [Cap01]. Dan [Sei90a]. Dancoff [Hei56e, Hei58b, Hei72b, Hei84e]. Dancoff-Method [Hei58b]. Dancoff-Methode [Hei84e]. Dancoff-Methode“ [Hei56e, Hei72b]. Dangers [Hei81a]. Danish [Ano44a, Boh64, BR64, Fav05, Hei35e, Hei64d, Jun57, RRK^{+}64]. Dank [Hei67g]. dann [Büh98a]. dans [Dem15]. darf [URWB95]. dargestellt [Her01]. Data [CR84, Hei52n]. Datenautobahn [URWB95]. David [Ano94, Büh95, CS94, CD94, Gar91, Hei43u, Kea09, Lib68, Mat86b, Phi92, Rec85b, Sta11a, Swe93, Wal08, Bed09, CS94, Dre92, Fer07, For09, Fra94, Gar91, Hei66e, Hei66f, Kra86a, Sap10]. Davis [Ano29]. day [Hei49e, Hei52d, Sei90a]. days [Blo76a, Blo76b, Blo85]. DC [Cas11]. Dean [Sei90a, Sei90a]. Dear [Cas05, HH17, Kan03, HHHH11, HHH03]. death [Hei68f]. Debate [BSMB99, Hei86a, Kra07, Lan08, Lau07, Bro93, DR92, Dör05, GP92b]. Debye [HW11, HW11]. Debye-Affäre [HW11]. Dec [DDH89d, Hei89m]. Decades [Sei90a]. Decay [GR09, HVTK^{+}97, Hei43d, Gau14]. December [Ano44a, BCEP94, DDH89d, FS93, Hei43r, Hei48a, Hei89m, MP77]. deception [Gra00]. Deceptive [RMRP09]. Decision [Gol92, Sti47]. Declaration [BBF^{+}57a, BBF^{+}57b, Hei57e, KF01]. Declassified [GP92b]. decline [Rin69, Rin90]. decoherence [BM11, CL94]. decoherence-enhanced [BM11]. dedicated [Hei43r, PRW55, Rec02a]. Deeper [Sei05]. defeat [Ols63]. defeats [Smo92a, Smo92d]. Defect [Gam30, HEB^{+}80]. Defects [MGG^{+}95]. Defects/Weiss [MGG^{+}95]. defined [MP92]. Defining [ARK^{+}99, Mla98]. Definitions [KED^{+}60]. deformability [BDR84j, BH24a, BH63b, Hei23a, Hei24c]. deformations [GWW^{+}09]. deformed [GH10, Sch99]. Deformierbarkeit [BDR84j, BH24a, BH63b, Hei23a, Hei24c]. degenerate [DGS99, KGNS10]. del [W^{+}87, WdC02]. Delivered [Wig96a]. della [Arg09, Fer99, Hei36a, Hei60e, Hei99, Pol00]. dell’epistemologia [Arg09]. dell’indeterminazione [Cav14]. delusion [Gra00]. demand [Hei64e]. demonstrate [DRB86]. demonstration [EMGM94]. demonstrative [Mas04]. Demtröder [HVTK^{+}97]. Denken [Hei65a, Pfe77]. Denkschrift [Hei49g, Hei53d]. Denkstrukturen [Hei70b]. Denmark [Sch11c]. dens [Hei64d]. Density [KBWJ96, CL05, Sny82]. deny [Bør12]. dependent [BPS74, BP75, CGM^{+}71]. deployment [Bro92]. Derivation [Cop81, LL82, Tya63, KSGB^{+}83, Mar08, Cop81, Man11]. Derivative [RG72]. derzeitige [Hei61b, Hei84k]. Deschamps [BSP^{+}70]. details [AMS04]. detection [STCW94]. Detector [Cop81]. Detectors [SPS^{+}64]. Detectors/Gärtner [SPS^{+}64]. Determinanten [BMG^{+}50]. Determination [DDH40a, DDH40b, DDH89e, DDH89d, Fis41]. Determines [OW10]. deterministic [Hei85d]. Deterministically [Hei79b]. deterministische [Hei79b, Hei85d]. Deterrence [Hei86a]. Deuterium [CR12]. Deutsch [BSP^{+}70, GRE^{+}01]. Deutsch-Englisch [GRE^{+}01]. Deutsch-Englisch/ConceptDraw [GRE^{+}01]. Deutsche [Hei36i, Hei53b, HW07, Pop18, Rec92, Cas93c, Hei48g, Hei50c, Hei56d, Nag16, SM41, Wal90b, Wal93, vS14, Men36, Sch11b]. Deutschen [Ano84b, Eic50, Eic53, Hei30a, Hei41b, Hei49g, Hei50g, Hei53d, Hei53q, Hei64b, HB^{+}72, Hei86o, NW95, Rec91a, Rec91b, Wit83, All06, Hei50d, Hei68a, Hei93a, Him90, Hof92, Hof93, Hof23, Irv67a, Kar05c, Pow93a, Rad83, Rad91, vL47, vL48a]. Deutscher [Hei89-38, Sta36, Wal93]. Deutsches [DRF90a, DRF90b, SPS^{+}64, Sim10]. Deutschland [Brü64, Hau22, MGG^{+}95, Sch07, Hei39d, Hei46d, Hei53k, Hei53g, HR53, Hei56j, Hei67c, KP07]. Deutschlands [Hei51e, Hau22]. Deutung [Bor61, Bor63b, EH38, EH84, Hei56h, HB63, Hei39b]. Develop [WHF^{+}96]. Development [Gou47a, Hei73d, Hei74l, Hei75a, Hei75l, Hei75j, Meh73, Ome75, Bad95, Bro92, Hei29a, Hei29b, Hei33a, Hei34d, Hei39d, Hei40b, Hei55d, Hei56b, Hei56h, Hei63a, Hei63c, Hei65c, Hei71c, Hei75k, Hei84-51, Hei84o, HH84, HBE87, Meh75, PRW55, Hei35e]. Developments [BW04, Hei04, Hau22, Hei56b]. Devil [Cor03]. Dezember [FS93, Hei43r, Hei48a, Hei71h, RDP^{+}95]. Dezember-Heft [FS93]. DFG [DRF90a, DRF90b]. DFR [HR53, HR53]. d’Heisenberg [Ech96, dB82]. d’Hitler [Luc16]. Diagrammatic [SL70]. Dialog [MGG^{+}95]. Dialogue [Bel92, Bel99, Hei74e, Wüt16]. Diaries [Hah45]. Diary [Sei90a, HHHH11]. dice [MB11a]. Dickens [HVTK^{+}97]. Dickens/Krätz [HVTK^{+}97]. dictionaries [MRV24]. Dictionary [DG99, MGG^{+}95]. Did [Ber95, LH93, Pop21, RW92, Wal22, Hei67a, Kat96, MS15, Ols63, Pop16a]. Diebner [vW1 ]. died [Kva76]. Dies [Hei68e]. Difference [Kor05, AFW94, Bt87]. different [HJ43]. Differential [Bt87, Hei32a, LSK^{+}76, Hei22b, Hei84w]. Differential-difference [Bt87]. Differentialgleichungen [Hei22b, Hei84w]. difficult [Luc82]. Difficulties [Ano34, Hei93c, Jes98]. Diffraction [Hei36b]. Diffusion [ARK^{+}99, DDH40a, DDH40b, DDH89e, DDH89d, Hei31i]. Diffusionslänge [DDH40a, DDH40b, DDH89e, DDH89d]. dilemma [Whi96]. Dilute [Rus64, KSSA78]. dimension [Hei22a]. Dimensional [Fis64, BT78, DMdA06, Isa82, PDL92, Woo77, ZG96]. Dimensionen [Hei22a]. dimensions [DPL94, DPL96]. Dingler [BMG^{+}50]. Diplomacy [Hei76b]. Dirac [Ano33, Ano51, Ben34, Deh36, Haa34, Hei30c, Lak96, Smo93, Wig96a, Wil36b, Ano34, Ano35b, Ano75, Bok04, BR87, Haa34, Hei30b, Hei34b, Hei34a, HE35, HDW75, HE06, KW87, LW34, Nis32, Smo93, Wig96d, Wig96k]. Diracschen [Hei34b, Hei34a, HE35]. Direct [EMGM94, KW02, KW03]. Director [Sei90a]. Dirichlet [GV85, GV86]. disappearance [GHH02]. Discontinuous [Sta11a]. discourse [Sal01]. Discoverer [Hei68e]. discoveries [Bra09, Lig05]. discovering [Ano20]. Discovery [CR12, Gin75, GR09, Pan02, Rog13, Shi78, Sim00, Bel85, CHW91, Hei58g, Hei58h, Hei58i, Hei60e, Hei74k, Kae48b, Sim89, WC84]. Discrepancy [LD66]. Discrete [Kor05, Zak96, DM99, DL08, Sud06]. Discussing [Mil12, Fri11]. Discussion [Dir28, Hei53i, Hei56l, Hei62d, Hei62e, Hei62c, Hei75n, Hei84c, Pau34, Hei32b, Hei62d, Hei62e, Hei62c]. Discussione [Hei28c, HSBA60]. Discussions [Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06, Meh87, CCJ^{+}34, Hei28c, HSBA60]. Diskurs [Sal01]. Diskussion [Rec85a]. Diskussion/Neutronen [Rec85a]. dispersion [Hea77]. Display [Sim10]. Disposal [SPS^{+}64]. Dispose [Ano34]. dispute [Sal01]. Disputed [Gou49]. Disputes [Gou47a]. Dissertation [KPMN20]. Distance [Aas02a, Cop81]. distinction [Mon04]. distributed [Jor78]. Distribution [KBWJ96, Mor72a]. Disturbance [RDM^{+}12a, RDM^{+}12b, D’A03, Wis98]. disturbing [JD06]. divergent [Cam07a]. DM [Büh95, Hei53b, Hen86, Str62b, Str62c, Wit83]. Do [Sei05, Mas04]. Dobzhansky [Adl03]. doch [NW95]. Doctor [Dje07]. doctorate [Hei89-27]. doctorates [Hei01b]. documentary [CHW91, WC84]. documentation [Aas02e]. Documents [Kra07, Lan08, Lau07, Pop16c, Aas02b, Aas02d, Ano02a, Dör05, Hof23, Paw02a]. Dodging [Mac78a]. does [MB11a, Hei51e, Hei68b]. Doktor [HVTK^{+}97]. dokumentasjon [Aas02e]. Dokumentation [GZ92b]. Dokumente [Ano02a, Hof23, Paw02a]. dollar [CP93]. Domains [Gre62, Maa05]. Dominique [Jor78]. done [Hei89-40]. Don’t [RMRP09]. Doomed [Eck90]. Dörries [Kra07, Lan08, Lau07]. Double [Hei74e, dB58]. double-solution [dB58]. doublets [SH22b]. doubt [Ano12]. Doubts [Hei53j, Hei84-53]. d’ouvrages [Bar04, CD94, Mat86b, Mat86a, Mat87]. Downwind [Sei90a]. Dr [Ano71, Hei36i, Hei53b, Hei53q, Hei65f, Hei89t, vW1 ]. Dr. [Ano29, Mor48]. draft [Hei62n]. Drama [Aas02e, Cas22, Sch22, Gla00, Sch15]. dramatic [Cas17]. Dramatik [Sch15]. Drawing [Ber03]. dream [Irv67a]. dreamers [VB15]. dreams [Haw11]. Drei [Deh37, RDP^{+}95, Hei39b]. dreiatomiger [Hei24c]. Dreier [HMSH78, Sch05]. dreissiger [Fie80]. Dresdener [URWB95]. Dritte [Bed03, Bed04, Cas93c, Hei67a]. Dritten [Bey82, BDR89b, Hau22, Hei47a, HW07, HW11, Pop18]. Drive [Bro93]. Drogenkoch [Ost11]. drogi [Inf33]. Drop [Kae31]. drug [Ost11]. dual [Per88, Tem91]. Duality [RS94, Cam06]. due [Wei93]. Duffee [BMG^{+}50]. durch [BDR84f, DDH42b, DDH89b, EHR^{+}96, Hei32h, Hei36d, Hei37b, Hei38c, Hei38b, HA^{+}48, Hei64b, Hei84l, HMSH78, KH25, URWB95]. durchdringende [Hei41c, Hei84n]. durchdringenden [Hei38d, Hei43c]. Durchführung [Hei41g]. Durchgang [Hei36d, Hei37b, Hei38c, Hei38b, Hei84l, BDR84f]. During [KH19, Ano44a, HB82b, Hei73a, Hof05, vM05a]. Durr [Bro85, Bro86, GRE^{+}01, Kra86a, TW91, Mat87, Sch91b]. Dutch [Hei39c, Hei84j]. DVD [Hah10]. DVT [DRF90b]. DVT-Preis [DRF90a]. Dynamic [Pan02, TBL02]. dynamical [Kav24]. Dynamics [RFS^{+}91, RG72, BT96, BG96, DMCF06, EHUY92, HP29, HP30, JKS13, PL91, Hei27a]. Dynamics/Audretsch [RFS^{+}91]. Dynamik [MGG^{+}95]. Dyson [LD66].

E. [Hei41e, Hei89-34, Wig96j]. early [Blo76a, Blo76b, Blo85, Sch05, Sch07, Sal19]. Earth [HEB^{+}80]. East [Ano68]. easy [DRB86]. easy-axis [DRB86]. Eavesdropping [Hol93b]. Ebel [SCH^{+}78]. Eckart [Eck70]. Economic [McC94]. Econophysics [vRFS^{+}00]. Ed [Lan08, Wit83, Cru23, Kra07]. Eddington [Adl03, Cal14]. edelgasahnlicher [Hei23a, BDR84j]. edge [Hol18]. Edited [Cas05, Eis00]. Edition [Hei53x, Hei93b, Hei93d, Tre01, Hei31b, Hei53b, Hei68j]. Editor [Ben13, LRD^{+}91, Pow02, Rab48]. Editorial [Rig87]. editors [And89]. Eds [Kra86a]. Education [GHW36, HGW^{+}89a, Hei89-30, Ano61a, Hei49h, Hei69k, Hei69l, Rot93, HGW^{+}89a]. Edward [Dür05, Sei90a, Dür05, MF69, Wig96b]. Effect [BM84, Hen98, HJ43, Hei22c, Hei24d, Hei25d, HJ26, Bre23, Lud01]. effective [DDH42a, KSGB^{+}83]. Effekte [Hei24d]. effektiven [DDH42a, DDH89a]. Effetto [Lud01]. efficiency [HJ43]. Efficient [TR09]. effort [Bro93]. Efforts [Gla00]. efter [Hei35e]. egész [Hei83h]. egg [Rec01]. Ehren [URWB95]. Ehrendoktorgrades [Hei89-27]. Ehrensaal [Rec91a, Rec91b]. ehrt [DRF90a, DRF90b]. Ehrung [Rec91a, Rec91b]. Ehrungen [Hei64b]. Eichler [BMG^{+}50]. Eigenschaften [Hei84-48, Hei84g]. Eigenvalues [KBWJ96]. Eight [Bor53, Hei52k, BBH^{+}62]. eighteen [BBF^{+}57b, ABH^{+}55, ABH^{+}87, Got83]. eightieth [Hei62i]. Eighty [Hei38g]. ein/Gesamtdeutsche [DRF90a, DRF90b]. einem [DDH42a, DDH42c, DDH89a, DDH89c, Fis02, Hei31j, RDP^{+}95]. Einfluß [BH24a, BH63b, Hei24c]. Einführ [Ano35b]. Einführ-ung [Ano35b]. Einführung [Deh36, EHR^{+}96, Haa34, Hei36b, LSK^{+}76, RFS^{+}91, SCH^{+}78, SPS^{+}64, Hei43h, Hei62f, Hei67d]. Einführung/Gaisser [RFS^{+}91]. Einführungstagung [Hei69b]. eingeweiht [DRF90a, DRF90b, URWB95]. eingeweiht/Photonik [URWB95]. eingeweiht/Studienförderpreis [DRF90a, DRF90b]. Einheit [Hei42d, Hei49c, Hei53f, Hei69e, Lan72]. einheitliche [Hei67d]. einheitlichen [Hei62h, Hei63c, Hei64a, Hei84o]. einigen [Hei27b]. Einleitung. [Hei43o]. einmal [Hei64c]. Einsatz [URWB95]. Einstein [Adl03, ARK^{+}99, Bes04, BB69, Bro72, Bro95, Car97, Eis00, Fay07, Fer07, For09, GRE^{+}01, HEB^{+}80, HVTK^{+}97, Hei49a, Hei53b, Hen86, HvMW79, Kra96, Mar73, MGG^{+}95, Mil12, OSWR86, Par06, PHvMW01, PvM93, Pye81, Rec86, RS11a, Tur06, vMHW85, vM93, vM96, vM99, vM01, vM05b, Acz01, BB69, BB71a, BB71b, BB05, Cas75, CG14, FW07, Fin96, Fri11, Hei49a, Hei53b, Hei55a, Hei62n, Hei67e, Hei74o, Hei75e, Hei77c, Hei83f, Hei86o, Hei89o, HvMW79, Hol18, Hol00b, Hol05, Kle17, Lin07, Lin08, Meh87, Mon04, Pan02, PHvMW01, PvM93, Whi96, Wig96g, Wig96e, Wig96j, Wig96l, vD15]. Einsteinbüste [Hei86o]. Einsteins [Hei55a]. Einweihung [Hei72f, Hei86o]. einzelnen [HJ43]. Electric [BSP^{+}70]. electrodynamic [Hei48c]. electrodynamical [SME77]. Electrodynamics [Ano75, BSP^{+}70, Hei62g, Hei65k, HDW75, Hei79c, Hei79b, Hei84-53, Hei84q, Hei84r, Hei84-30, Sch02c, Sch58, Sch03, DHYY65, Dys51a, Dys51b, Hei53j, Hei85b, AH57, Hei63d]. Electromagnetic [Hei47c, RFS^{+}91]. Electron [Ano34, Hei29c, Hei46f, WMG^{+}92, BH23a, Hei30e, Hei38a, Hei52s, Hei84i, HBE87, SHI73, Ano34, BDR84l, BH63a]. Electronic [Dje07]. Electronics [BSP^{+}70]. Electronics/Moelwyn [BSP^{+}70]. Electronics/Moelwyn-Hughes [BSP^{+}70]. Electrons [OSWR86, Hei26e, Arg09]. Elektrodynamik [vRFS^{+}00, AH57, Hei63d, Hei84r]. elektrodynamische [Hei48c]. elektromagnetischer [Hei47c]. Elektron [Hei38a, Hei52s, Hei84i]. Elektronen [Hei26e]. Elektronenbahnen [BH23a, BH63a]. Elektronenmikroskopie [DRF90a, DRF90b]. Elektronenmikroskopische [BSP^{+}70]. Elektronentheorie [BDR84l, Hei46f]. Elektronikindustrie [RDP^{+}95]. Elektrons [Hei30e]. Elektrotechnik [RDP^{+}95]. Element [MA40, vW41b, Rig02]. Elementarteilchen [BDR85b, Hei38h, Hei43j, Hei44a, Hei53o, Hei68d, Hei84-48, Hei84-59, Hei93c, BDR84g, DHM^{+}59, Hei39f, Hei43m, Hei50j, Hei51m, Hei51i, Hei55c, Hei56k, Hei58a, Hei58m, Hei59b, Hei61b, Hei62a, Hei62f, Hei62h, Hei63c, Hei64a, Hei65b, Hei65e, Hei67d, Hei70h, Hei84d, Hei84g, Hei84k, Hei84o, Hei84m, Hei76h]. Elementarteilchen“ [Hei56c]. Elementarteile [Hei54c]. Elementary [BH83a, HEB^{+}80, Hei29c, Hei55g, Hei56c, Hei57g, Hei58a, Hei58m, Hei58n, Hei60h, Hei62f, Hei62k, Hei63i, Hei64a, Hei65n, Hei65e, Hei68d, Hei68g, Hei74l, Hei75c, Hei76a, Hei77a, Hei84-48, Hei84t, Hei84-57, Hei84v, Hei84-31, Hei84-34, Hei84-32, Hei84-38, Hei84-39, Hei84-43, Hei93b, Hei93c, HP93, Hei93d, Pau41, Pau93, Rec81a, Wig67, BDR85b, BH82, DHM^{+}59, Geĭ56, Geĭ77c, Hei38h, Hei43m, Hei43j, Hei44a, Hei50j, Hei51m, Hei51h, Hei51i, Hei55c, Hei56k, Hei59b, Hei61b, Hei62a, Hei63c, Hei63g, Hei65b, Hei66d, Hei67d, Hei70h, Hei74p, Hei76f, Hei76h, Hei84d, Hei84g, Hei84k, Hei84o, Hei84m, Hei84-58, Hei84-59, Hei84-27, BDR84g, Hei39f, Hei54c]. elementary-particle [BH82]. Elemente [URWB95]. Elementen [vW41b]. Elements [Sim00, vW41b, JLP^{+}02, Ras78, CR13, Hei72a]. elettroni [Arg09]. Eliminated [Kas11]. Elisabeth [Kra86a, Mat86a, Roh85, Wei85, Mot85]. Eltern [Cas05, HHH03, Kan03]. Embattled [Sal19]. embeddability [BG96]. Embellishments [Kra93]. Emmet [Los61]. Empfang [Hei89-28]. Empfehlungen [Ano84a]. empfiehlt [URWB95]. Empiricist [Sch86]. Employment [BS47, Gou47b]. encontros [Hei96c]. Encounters [Bun71, Hei70e, Hei71e, Hei83f, Hei89o, Huf71, Hei71i, Hei71j, Hei75e, Hei77c]. Encyclopedia [CF98, WMG^{+}92, OS98]. Encyclopedia/Watts [WMG^{+}92]. End [Hei75b, Sei90a, DF96, FW07, Gre05, Hei74f, Jun57]. endgultig [Hei89-40]. ends [Zan02b]. energetic [BDR84f, Hei84l]. Energie [Hei89n]. Energieabfuhr [RDP^{+}95]. Energieabfuhr/Poröses [RDP^{+}95]. Energieerzeugung [Hei41f, Hei89-36, Hei89x]. Energiegesetz [Hei42a]. Energiegewinnung [Hei39e, Hei40a, Hei42e, Hei43l, Hei89h, Hei89l, vW40, Ano42]. Energiegewunnung [Hei89m]. Energien [LMOS15]. energiereicher [BDR84f, Hei36d, Hei37b, Hei38c, Hei38b, Hei84l]. Energies [LMOS15]. Energieschwankungen [Hei31j]. Energy [Ano29, Bra97, Cop81, Cus86, Dah99, DW52, DGH^{+}12, GLR14, HB82b, Hei47g, Hei47h, Hei84-28, Hei84-35, New35, Sei90a, Tur46a, Tur46b, Wit12, BSP^{+}70, HB82a, Hei30e, Hei36d, Hei37b, Hei38c, Hei38b, Hei39e, Hei40a, Hei41f, Hei42a, Hei42e, Hei43l, Hei46d, Hei53k, Hei55e, Hei56b, Hei56l, Hei83g, Hei89h, Hei89-36, Hei89l, Hei89m, Hei89x, Hei89n, KW03, TR09, vW40, vW41b, Ano42, Hei31j, Wat01]. energy/Kowalski [BSP^{+}70]. engaging [Oor09]. Engine [Dje07]. Engineer [CS94]. England [Ano84a, Hof23, Hof23]. England/Forschungsbudget [Ano84a]. Englisch [BSP^{+}70]. Englisch/ConceptDraw [GRE^{+}01]. englischen [Hof92]. English [Hof92]. enhanced [BM11]. Enrico [Hei36a, Pol00, GR09, Hei36a, Sch02c, Wig96f]. Entangled [LBS^{+}04, Paw02b]. Entanglement [Acz01, Bar07, DMCF06, HBB^{+}07, Gil09]. Entanglement-free [HBB^{+}07]. entdeckt [RDP^{+}95]. Entdeckung [Hei58g, Hei58h, Hei58i]. Enterprise [UM92]. Entgegennahme [Wil36b, HSD34a]. Entgegnung [Hei36i]. entre [DC48, dB58]. Entropie [Tel61]. entropy [DMCF06, Tel61]. Entstehung [HVTK^{+}97, Hei34e, Hei49k, Hei75f, Wal93]. Entwicklung [Hei29a, Hei29b, Hei34d, Hei39d, Hei56h, Hei60b, Hei63c, Hei65c, Hei71c, Hei84o, HH84, Hei86f, Hei63a]. envision [Gau14]. enyémhez [Sch66]. Eopredelennost [Smo92a]. epigoni [Fer99]. episodes [Bra09]. Epistemological [Hei84-54]. Epistemology [BDR85a, Plo10, Arg09]. Epoch [Hei75b]. Epsilon [DF96, FS93, Hof92, Hof93, Hof23, NW95, DF96, Fra93b, Fra93c, Fra93a, Hof92, Hof93, Hof23, NW95]. equality [HR02a]. Equation [ARK^{+}99, FMV07, Hei32a, Kra84, MB11b, Wan01, BN08, KW03, LL82, Mie80, Mie81, SHI73, dB58, dB58]. Equation/Gaponenko [ARK^{+}99]. Equations [Cre10, Gre10, Hak02, Hei55f, Hei57f, LSK^{+}76, AH57, Bt87, Bes98, DGS99, Hei22b, Hei53y, Hei54f, HKM55, Hei84w, Hei84-29, IM95, KJ06]. Equivalence [Dir64, Per64, dG14, DM99, Pac88]. Era [MGG^{+}95]. Ereignisse [Ano84a]. erfanden [NW95]. Erforschung [Fis10a, Fis12a, MGG^{+}95]. ergänzt [Hof23]. Erganzung [Hei85d, Hei79b]. Ergebnisse [BFH^{+}42b, HBF^{+}89c]. Erho [HJ43]. Erinnerung [Hei86f]. Erinnerungen [Dür05, Hei64c, Hei80, Hei83a, Hei91a, LL77, Hei60b, Hei67f, Rec91c, Tel93, Wig96g]. Erkenntnis [BDR85a, BDR84c, BDR84d, BDR85b, Ano61a, Hei43i, HPE^{+}33]. Erkenntnistheoretische [Hei84-54]. Erklärung [GRE^{+}01, KF01, BBF^{+}57b, Hei57e]. Ermenc [Sei90a]. Ernst [FS06]. Eröffnung [SM41]. Errata [Bor72]. Erratum [CJ72a, DPL96, Fri98, SZ74, Sch02a]. erreicht [DRF90a, DRF90b]. Erroneous [Tru72]. erste [Hei30a]. Erstgeburt [Fue61]. erstmals [Hof23]. erw [Mos42]. erweiterte [Clu37, Deh37, DS43]. Erweiterungen [Hei56m]. Erweiterungsbaus [Hei89b]. Erwin [Adl03, Ful15, Han62, Hei36b, HF46, Hei63e, BD92, Ful15, Han79, Hei36b, Hei61d, Meh87]. erzählt [FS06]. Erzeugung [Hei49d, Hei84p]. Erziehung [GHW36, HGW^{+}89a, Ano61a]. escritos [W^{+}87, WdC02]. esperienza [Arg09]. Espionage [Kra86b]. Essay [Hen00, Hen10, Lus10]. Essays [Bab02, CL83, Kra07, Lan08, Lau07, Dör05, Hei89o, Hol86, Men84, PRW55, Sch96a, Sch27]. essential [Rig02]. est [Dem15]. established [Log96]. establishment [Hei30a, Ols63]. Estate [Ano76]. estimation [HBB^{+}07]. Ethics [Log00]. Ettore [GHH02, GR08]. Études [Rhe91]. EU [URWB95, URWB95]. Eugene [Wig96m]. Europa [DRF90a, DRF90b, Hei89-40]. Europa/Risikostudie [DRF90a, DRF90b]. Europäische [Hei54b, Hei56i]. Europäisches [DRF90a, DRF90b]. Europas [Hei68b, Hei52c]. Europe [Hei52c, Hei68b, Hei89-40, Rod19]. European [Ano84b, Hei54b, Hei56i, Hei89p]. evaluated [SHI73]. Evaluation [Hei44c, Flo97, Hei43k, Hei43e, Hei89e]. evangelischen [BBH^{+}62]. even [Lan02]. Evening [BDH70]. events [Ano44a, Hei61e]. eventuelle [vW1 ]. éventuelles [Per39]. Evert [Bor58]. everyday [Mil97]. Evidence [Cas93b]. Evolution [Hei72g, Hei60b, Hei72g]. evolutionary [Nar93]. Exact [CGM^{+}71, Dob77, Hei79a, IM95, LB72, OS07, Ome79, Hei34i, Hei34f, Hei34g, Hei34h, Hei53r, Hei71d, Hei82c, Mie80, Mie81]. exakten [Hei34i, Hei34f, Hei34g, Hei34h, Hei71d, Hei82c]. examination [Hei01b]. Excellence [RDP^{+}95]. Excerpts [Sei90a, Hei69b]. Exchange [Ano29, BL03, CJ72a, CJ72b, vBN95c, Hei71n, OD03, Sla72, VBN95a, VBN95b]. excitations [Pug80]. excited [BH23a, BH63a]. Exclusion [Hei31k, Tho80]. excursions [Hol18]. Exercises [BSP^{+}70]. exhibition [Rec02b]. Exile [Ano33, Kra86a, Mat86a, Mot85, Roh85, Wei85, Hei84a]. exist [Pop16a]. Existence [ST65]. Existenz [RDP^{+}95]. existing [Gau14]. Exner [Han79]. expansion [Fed04]. expansions [GbK77, KSGB^{+}83]. Expelled [MP01]. Expense [Ros00]. experience [Arg09]. Experiment [Hei43e, Hei67c, HS83, BCL^{+}69, BBF^{+}45, CP93, Hei89e, HBB^{+}89, Rob69]. Experimental [Ano34, DJR93, NAZ02, DDH42a]. Experimentalversion [Ano84a]. Experimentalversion/Reparatur [Ano84a]. experimentelle [DDH42a, DDH89a]. Experiments [BSP^{+}70, Hei44c, HW48, TV23, GHW09, Hei36b, Hei41a, Hei89i, vD15, DDH41a, DDH41b, DDH89f, DDH89g]. Experiments/Leighton [BSP^{+}70]. Explain [Kae31]. explanation [Hei33d]. explicit [BPS74]. explicit-time-dependent [BPS74]. Exploded [Kae31]. Exploit [Tau94]. exploitation [Gim90]. Explosion [DeG74, Hei39g, Hei43a, Hei52q, Hei53s, Hei84y, Hei39j]. Explosion-Like [Hei43a]. explosionsartigen [Hei39j, Hei43a]. Explosionsschauer [Hei53s]. Explosive [Wal04, El 00, Mül40]. Explosives [MS15]. Extend [Ano34]. Extended [MSUU80, JU82]. Extending [Wis98]. extension [BP75, Hei89b]. Extensions [Hei56m]. extraction [Hei89m]. Eyes [Bab02, Büh98a]. eynshteynovskomu [Hei62n]. Eysenck [LSK^{+}76].

F [Bor53, CS94, Cas05, Hei24a, Hei31g, Hei72c, Hei86b, Pye81, Rec86, Wat01]. F. [Hei24a]. Fachausstellung [DRF90a, DRF90b]. Fachhochschulen [RDP^{+}95]. Fachwörterbuch [GRE^{+}01]. facility [Hei57d]. Facing [Cle09]. Facsimile [Hei93b]. factorization [KJ06]. factors [KKM^{+}09]. Facts [Pop16c]. Fail [GLR14, Kra87]. Failed [WHF^{+}96]. Failure [Eck90, KH19, Gou47f]. Fall [Dje07, MGG^{+}95]. Faltings [Ano84a]. famosos [W^{+}87, WdC02]. famous [W^{+}87, WdC02, Wit12]. fangen [RDP^{+}95]. Farbe [SPS^{+}64]. Farbe/Disposal [SPS^{+}64]. Farbenlehre [Hei41d]. Farewell [Dür76]. Farm [BL08, FS93, Hof92, Hof93, Hof23, Mar94, OS03, WR92, Wal93, Ber92, Ber93, BC95, Ber96, Cas17, Cas22, Dah22, FS93, Fra93b, Fra93c, Fra93a, GP92c, Hah45, Hof92, Hof93, Hof22, Hof23, Hol22, Klo93, Luc07, McC24, WR92, Wal93]. Farm-Hall-Protokolle [Hof93, Hof23]. Farm-Hall-Tonbänder [Hof92, WR92]. Farmhall [Hof92, OS03]. Fast [DDH42b, DDH89b]. fate [Jun56, Jun57, Jun64, Jun72, Jun90]. fathers [Kra84]. Faulted [BB96]. Faust [MGG^{+}95, Wal08]. Faustu [Dje07]. FCC [Dob77]. Fear [Hof93, Sei90a, Ano02a, Hei74c]. Fearful [vRFS^{+}00]. feasibility [Log96]. Feb [DDH89c, DDH89b, Hei89h, Hei89n, HBF^{+}89b]. Februar [Ano44a]. February [Ano44a, DDH89c, DDH89b, Hei89h, Hei89n, HBF^{+}89b, MP77, Rig87, Hei95]. Federal [Hau22, Hei57d, Hei65d, Hei86a, Sch05, Sch07]. Feier [Hei89a, RDP^{+}95, Hei60c]. feierlich [URWB95]. Feierstunde [Hei89u]. Feinmesstechnik [BMG^{+}50]. Feinmesstechnik/Borries [BMG^{+}50]. Feldafing [CDvW75]. Felder [DB71, HBD71, Hei56e]. Feldtheorie [Hei62h, Hei63c, Hei64a, Hei67d, Hei84o]. féle [Sch66]. Feller [RFS^{+}91]. Fellows [Hei69b, Hei73a, Hei75r]. fenomenologica [Arg09]. fenomenologicheskih [Geĭ67]. Fermi [Sei90a, Wig96f, GR08, GR09, Mac86, Mac87, Sch02c]. Fermion [Gid71]. Ferro [KI64, Sno49]. Ferro- [KI64]. Ferro-magnetism [Sno49]. ferroelectric [SF71]. Ferromagnet [KA71, LB72, LD66, BT78, DPL94, DPL96, Fed04, Pow76, dSL02]. ferromagnetic [DM99, DL08, HN99, PL91]. Ferromagnetics [Rus64]. Ferromagnetism [CJ72a, CJ72b, Gre63, Hei84s, Bro77, Hei28d, Hei28e, Hei30f, Hei84-47, Ing32, Hei31f]. Ferromagnetismus [Hei28d, Hei28e, Hei30f, Hei31f, Hei84s, Hei84-47]. Ferromagnets [PSZM02, HS70, SZ73, SZ74, Tho80, Woo77]. Festakt [Hei64b]. Festival [BDR86, Hei74g]. Festkolloquium [RDP^{+}95]. Festrede [BDR86, Hei58j, Hei58k, Hei58l, Hei89q]. Festschrift [Hei05b, KRW05, PFH20, Sim79, PFH20, Sim79]. Festvortrag [Hei56c]. Few [Hei27c]. Few-body [Hei27c]. Feynman [Mil12, Fri11, Hei62c]. fictional [Sal01]. Field [Bro02a, Bro02b, Ess71, Hei31j, Hei57f, Hei62h, Hei64a, Hei84-57, Hei84v, Hei84-43, Hei93b, Hei93d, KED^{+}60, Pau41, Squ67, Wig67, DKLN92, EHUY92, Hei37c, Hei56i, Hei62n, Hei63c, Hei63k, Hei66d, Hei67d, Hei67k, Hei67m, Hei74p, Hei84o, LM87, MSUU80, Tem91]. Fields [EHR^{+}96, Gre62, Hei50i, Hei57g, Hei84-31, Hei84-46, RG72, Alb73, Car71, DB71, HP29, HP30, Hei38i, Hei46b, Hei47f, Hei50h, Hei56e, Hei56m, HBD71, Hei84-45, KLV80, Hei43h]. Fields/Herrmann [EHR^{+}96]. fiftieth [Hei35d]. Fifty [Cas15, Kae48b, PC77]. Fifty-Year [Cas15]. Fig [Wil36b]. figure [Cas14b]. figures [Cas14b]. fiica [Hei96c]. fiktiver [Sal01]. files [GP92b, Hof92]. Filmbesprechung [HVTK^{+}97]. Filme [HEB^{+}80, HMSH78, SCH^{+}78]. Filosofia [Pol00, Alt06, Cas75, Cir88, Gem87, Hei96c]. filosofici [Hei60e]. filosofico [Hei99, Pol00]. filosofija [Geĭ70b]. filosofiya [Geĭ90]. filosofo [GA02]. filosofskie [Geĭ58]. filozofija [Hei97]. Final [Eic53]. finally [Hei89-40]. Finance [vRFS^{+}00]. Finance/Maple [vRFS^{+}00]. findings [Bro93]. fine [BK04, DHYY65]. fine-structure [DHYY65]. finished [Hei86d]. finite [KS92, SME77]. Finkelstein [HVTK^{+}97]. Finnland [DRF90a, DRF90b]. First [Ano33, Cas79b, MGG^{+}95, Cas78, DRB86, Fue61, GKZL96, Hei30a, Hof23, LTG^{+}12, SR92]. first- [DRB86]. first-born [Fue61]. first-order [LTG^{+}12]. Fischer [Cru23, Hei41e, Hei89-34, FS06, Hei41e, Hei89-34]. Fisica [Alt06, Cas75, Hei84-55, Hei00a, Hei08c, Fer99, HSBA60, Hei60e, Hei99, Pol00, Hei96c, HBS96]. fisico [Gem87, Cir88]. físicos [W^{+}87, WdC02]. Fission [DK78, Hei68e, NBDG40, Sim00, Sim10, Sim12, Stu85, TV23, Bad95, BDJ57, CHW91, Hei39e, Hei40a, Hei42e, Hei43l, Hei89h, Hei89l, Hei89m, Hei89n, KF01, Sim89, Sim07, WC84, Fav05]. five [CW85]. Fixer [MGG^{+}95]. Fixer/Wenske [MGG^{+}95]. Fizika [Geĭ90, Hei97, Hei58p]. fizike [BDH70, Geĭ68]. fiziki [Geĭ67]. Flawed [Bør12, Mar08]. Fletscher [BSP^{+}70]. Flow [Hei84x, Hei23b, Hei24e, Hei52i]. flows [Hei51g]. ≪Spurion^{≫} [Hei84-33]. ^{≪}Spurion^{≫}-Theory [Hei84-33]. Fluctuation [Hei26d]. Fluctuations [Hei31j, Hei34e]. fluid [Hei23b, Hei24e, Hei51g]. fluids [Hei22b, Hei84w]. fluorescence [Hei25b, Hei25a]. Fluoreszenzlichtes [Hei25b]. Flussigkeiten [Hei22b, Hei84w]. Flüssigkeitsströmmen [Hei23b, Hei24e]. FOCS [IEE06]. Focus [Wal04]. focuses [Hei64e]. Folge [Wit83]. Folgerungen [HE35]. follower [Fer99]. following [Dir28, Gou47a]. Fölsing [MGG^{+}95]. Fondamenti [Hei36a]. Footnote [Mac85]. Force [Ano05, Har73]. Forced [Kae31]. Forces [Bri65a, BR96b, Hei47c, Hei50i, Hei54d, Hei84-46, SHD91]. Forderung [Hei64e, Hei31g]. Foreign [Cla23, Wal92, Hei68a, HW11]. Forensic [MWL^{+}15]. Forensics [MWL^{+}11]. Foreword [Bro72, Hei75o, Cas05, Hei39i, Hei43s, Hei53x, Hei53v, Hei53w, Hei59h, Hei60j, Hei62m, Hei63l, Hei64f, Hei64g, Hei65q, Hei66h, Hei66i, Hei67n, Hei67o, Hei68j, Hei68k, Hei69m, Hei69n, Hei71m, Hei72n, Hei73i, Hei74q, Hei75t, Hei89-39]. Forgetting [Sim12]. Forging [Sei90a]. Forgotten [GR08, Fis02]. form [KKM^{+}09, Ter82]. formal [Hei24d]. formalen [Hei24d]. Formalism [Eps71]. Formation [Cas79b, Pol95, Gau14, Hei24c]. Formen [Hei61a]. Formenreichtum [Hei65p]. forms [Hei61a, Hei65p]. Formula [Cas14b, He04, vM02, Hah10, Hea77, HHvB10, Cas14b]. formulas [LL82]. Formulation [Bro02a, Bro02b, Gre63]. Forschen [SPS^{+}64]. Forscher [Hei36e]. Forschung [BDR89b, EHR^{+}96, Hei30a, Hei49g, Hei51e, Hei53d, HR53, Hei56n, Hei63k, Mai07, WMG^{+}92, Hau22, Hei48l, Hei49c, Hei53f, Hei54b, Hei64e, Läs13]. Forschung/Omnès [EHR^{+}96]. Forschungsbudget [Ano84a]. Forschungsminister [RDP^{+}95]. Forschungsprogramm [Rec96]. Forschungsprojekten [Hei65d]. Forschungsrat [Hei50c]. Forschungsrates [Eic53, Hei53d, Hei53q, Eic50, Hei49g, Hei50g]. Forschungsrat“ [Hei89-38]. Forschungsreaktor [DRF90a, DRF90b]. Forschungsstipendiaten [Hei73a, Hei75r]. Forschungszentrum [DRF90a, DRF90b]. fortalt [BR64, RRK^{+}64]. fortolkning [Hei64d]. Fortschritt [Hei70b, Hei74c, Hei34j]. Fortschritte [Hei58a, Hei84d, BDR84h, Hei30f, Hei37c, Hei37d, Hei58m, Hei62h, Hei84s, Hei93c]. Forty [HB82a]. Fotos [HHHH11]. Foundation [Ano71, BDR89b, Hei56g, Hei69e, Hei71n, Hei89u, Hei89-28]. Foundations [And89, BHJ62, GL23, IEE06, Ano35b, Hei34i, Hei35e, Mil90, PC77]. Founders [Kle17]. Four [Ada72, Ano75, HDW75, VB15]. Fourier [Bru76, Gel84]. fra [Ano44a]. från [Sch15]. fractality [Jad14, Lic16].

Fractals [MGG^{+}95, Sch91a, Jad14, Lic16]. Frage [Hei25b, Hei51m, Hei89-32, Hou41]. Fragen [Hei36f, Paw02a]. fragment [BGĬ90]. fragmentation [Hei32h]. frailty [Gra00]. frame [Hei51h]. Frames [RS94]. Framework [Car02a, Car02b, Hei46b]. France [Ano33]. Frank [Mar94]. Franz [Han79]. Fraser [vRFS^{+}00]. Fraunhofer [Rec85a]. Fraunhofer-Gesellschaft [URWB95]. Fraunhofer-Institut [Rec85a]. Frayn [Kra07, Lan08, Lau07, Log00, Dör05, Ros00]. Frederic [Tho97]. Free [Com31, Hei76b, Mar31, HBB^{+}07, KW03, Oor09]. Freeman [Ano94, Fra94, WMG^{+}92]. Freeon [MC97]. Freiberg [DRF90a, DRF90b]. Freiburg [Hei52f, Hei52f]. freigegeben [Paw02a]. Freiherr [Cas15, Hei02]. Fremde [HW11]. French [BH28, BH84, CCJ^{+}34, Dem15, DC48, Ech96, Fra93a, Gou48b, Hei34c, Hei53p, Hei57a, Hei62d, Hei62e, Hei62c, Hei67k, Hei68h, Hei84b, Hei84-61, Kar07, Luc05, Luc16, Lur95, Per39, VB15, dB58, dB82]. Frenkel [ARK^{+}99]. Fresh [SPS^{+}64]. FRG [Eck90]. Friction [HEB^{+}80]. Frieden [Hei89d, Hei89c]. Friedensforscher [Rec02a]. Friedensforschung [HH14]. Friedliche [Hei53k]. Friedrich [Ano02a, ARK^{+}99, Cas14a, Cas15, GRE^{+}01, Hei72a, Hei02, HH14, Rec02a, vWL93, HKvL47]. friends [BR64, RRK^{+}64, Roz67]. Friendship [Cas15, Gla00, BB05]. Frisch [KED^{+}60]. Fritzsch [RS11a, Mil12]. Fröhliche [Hen11, Rec09, Rec10]. Froissart [DeG74]. frontier [Hei36e, Sch07]. Frontiers [Ano66, Hei74b, Cas74]. frühen [Sch05, Sch07]. frustrated [KGNS10, sMDN^{+}11]. Fuchs [Sei90a]. Fueki [RFS^{+}91]. Fueled [BSMB99]. fuerzas [SHD91]. Fues [Hei36b]. Full [FMV07]. fullerene [NAZ02]. Fullerenen [RDP^{+}95]. Function [KA71, LB72, Mor71, Mor72b, Ni04b, Unr72, Fed04, KSSA78, Mor72c, SL70, SF71]. functional [CL05]. functionals [Per88]. functions [Dun86, MRV24, Mor72a, Pea76]. Fundament [URWB95]. Fundamental [BW04, Hei52d, Hei75s, Hei76d, Hei84h, Hei84-56, Hei04, JZ11, Str62a, BSP^{+}70, BN08, Hei53o, Hei75i, Hei77d, SHD91, GZ95, Hei36f, Hei49e]. fundamentales [SHD91]. Fundamentals [Som36, Hei34f, Hei34g, Hei34h, Hei35f, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei49l, Hei59i, Hei73j, Hei36a]. Fundamente [HVTK^{+}97]. funeral [Hei86b]. fünfzigsten [Hei35d]. Funktionen [SPS^{+}64]. Furman [AO44]. Furry [Ano34]. Fürsorge [Hei49g]. Fusionsexperiment [RDP^{+}95]. Future [Hei79c, Hei79b, URWB95, Hei85b]. Fysik [Ano44a].

G [Adl03, Ano08, Hei28a]. G. [Paw08]. gab [Hei48g, Pop16a]. Gables [BSP^{+}70]. Gaisser [RFS^{+}91]. Galilei [Büh98b, MGG^{+}95]. Galileo [Büh98b, Fer99, RS11a]. GALLEX [URWB95]. Galoisian [PC14]. gammée [Luc05]. Ganze [BDR85b, BSP^{+}70, Ess71, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06, Sim07, Bun71]. Ganze/Hewitt [BSP^{+}70]. Gaponenko [ARK^{+}99]. Gärtner [SPS^{+}64]. gas [BDR84j, Hei23a]. Gate [Sei90a]. Gauge [LTG^{+}12, ZG96]. Gauss [Mor97]. Gauverein [Hei41b]. geb [Ben30, Büh95, Deh36, Str62c]. Gebaude [Bor58]. Gebiet [Hei37c, Hei56i, Hei57d, Hei63k]. geboten [RDP^{+}95]. geboten/Drei [RDP^{+}95]. gebunden [Rec03]. Geburtstag [DB71, Fue61, Hei43r, HBD71, Hei72c, Hei86n, Hei89-37, Jac01, Rec01, Rec02a, Rec02b, vM02, Hei44b, Hei54e, Hei62i, Hei64h, Hei05b, KRW05, PFH20, RH02]. Geburtstage [Hei35d]. Geburtstages [Sch82]. gedachte [WMG^{+}92]. Gedanken [Hei37e, Hei75p, HS83]. Gedankenexperiment [URWB95]. Gedenkfeier [Her76a]. Gedenkworte [Hei63e, Hei69f, Hei75q]. Gefahr [Hei89-40]. Gegen [EHR^{+}96, Lan72]. Gegenwart [Hei69e]. gegenwärtige [Hei55c, Hei84m]. gegenwartigen [Hei49e, Hei38e]. geh [Ano35a, Ben30, Ben34]. gehalten [Hei31g, Hei43p, Hei52p, Hei53l, HMW^{+}84]. gehaltenen [HSD34a, Wil36b]. Geheimdienstakten [Hof92]. geheime [All06, Kar05c, Nag16]. Geheimgeschichte [Pow93a]. geheimnisvollen [HVTK^{+}97]. geheimnisvolles [Kar05a]. geht [Ano84a]. Geiger [Cas79a, DRF90a, DRF90b, Cas79a]. Gel′fand [Tem91]. Gelehrten [Hei68f]. Geleit [Hei52r]. gelingt [RDP^{+}95]. Gembillo [Cir88, Pol00]. gemeinsam [Hei41b]. Gemeinschaftsforschung [Mai07]. General [Hei58n, Hei84-48, Hei84t, Mor72a, Sei90a, BJN01, Hei84g, Hei84b, Mil12, SF71, ZG96, Hei34c, Hei58a, Hei84d]. générales [Hei34c, Hei84b]. Generalisation [Ing70]. Generalization [Mes04, San81, Hea77]. Generalizations [DD13, Tri02]. Generalized [AG88, Fri97, Fri98, LL86, May87, SL70, dlP80]. Generalizing [Chi01]. generating [vW40]. Generation [Hür22, Ano42, Hei52g, Hei56l, TR09, Hei72a]. generations [Ada72]. Genesis [Bel92, Hei70e]. Genie [Fis02]. génies [VB15]. Genius [Dje07, Fis02]. geniuses [VB15]. Geodesics [Ni04a]. geöffnet [Hof92]. Geoffrey [CS94]. geometric [DM99]. geometry [AS03, Ano00, Sem03]. geplant [Rec85a]. geplant/Strahlenbelastung [Rec85a]. geplanten [Hei42b, Hei89g]. Gerald [Fay07, Par06, Tur06]. gerettet [Sim07]. Gerhard [URWB95]. Gerlach [Hei64h, Hei64h]. German [Ber93, Car99a, Cas93a, Cas05, Cas11, Eis00, Hei41b, Hei50c, Hei50g, HR53, Hei53b, Hei53q, Hei64b, Hei89-38, Hol93a, Kar07, Klo96, Lan02, LH93, Lib68, Log99, Nag16, Neu99, Pei93b, Swe93, Wal93, Wal99, Wal08, Wil94, ABH^{+}55, ABH^{+}87, Ano35b, Ano36, Ano42, Ano61a, Ano61b, Ano71, Ano73, Ano76, Ano93a, Ano02a, Arr06, AH57, BDJ57, BBH^{+}62, Bed03, Bed04, Ber92, Bes04, BS47, Bet93, Bet00, Bey82, BBB^{+}84, BDR84f, BDR84g, BDR84l, BDR84h, BDR84j, BDR84k, BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93, BFH^{+}42a, BFH^{+}42b, BBF^{+}45, BBF^{+}57a, BBF^{+}57b, B^{+}61, BH23a, BH23b, BH24a, BH24b, BJ25, BHJ26, Bor61, BHJ62, Bor63a, Bor63b, BH63b, BH63a, BHJ63, BB69, Bal07, Bro93, Brü64, Büh98a, Büh98b, Bur62, Cap57b, Cas79a, Cas92a]. German [Cas93c, Cas95, Cas02b, Cas17, DW52, DF96, DDH40a, DDH40b, DDH41a, DDH41b, Döp42, DDH42a, DDH42c, DDH42b, DDH89e, DDH89d, DDH89a, DDH89c, DDH89b, DDH89f, DDH89g, DRF90b, DHM^{+}59, DH61a, DB71, Dür76, Dür05, Dür62, Dür80, Eck12a, Eic50, Eic53, EB82, Ess71, EH38, EH84, Fie80, Fis41, Fis02, FS06, Fis10a, Fis12a, Fis15, For22, FS93, Fra05, Fue61, Ful15, GZ92b, GZ92a, GHH02, GP92a, Got02b, Gou47f, Gou47a, Gou47b, Gou47c, Gou49, Haa34, Hah40, Hah43, Hah45, Hah10, HB82a, Har43, HJ43, Har37a, Har37b, Har81, Hau22, Hei22a, Hei22b, Hei22c, Hei23a, Hei23b, Hei24d, Hei24a, Hei24c, Hei24b, Hei24e, Hei25b, Hei25a, Hei25d, Hei25c, HJ26, Hei26a, Hei26b, Hei26c, Hei26f, Hei26e, Hei26d, Hei27e, Hei27b, Hei27a]. German [Hei27f, Hei27c, Hei27d, Hei28b, Hei28d, Hei28e, Hei29a, Hei29b, HP29, Hei30a, Hei30f, Hei30h, Hei30g, Hei30d, HP30, Hei30e, Hei31b, Hei31c, Hei31a, Hei31j, Hei31i, Hei31e, Hei31f, Hei31k, Hei31h, Hei31g, Hei31d, Hei31l, Hei32f, Hei32g, Hei32h, Hei32i, Hei32j, Hei32d, Hei32e, Hei33b, HPE^{+}33, Hei33d, Hei33c, Hei34b, Hei34a, Hei34e, Hei34d, HSD34a, Hei34i, Hei34f, Hei34g, Hei34h, Hei34j, HvLM^{+}35, Hei35b, Hei35a, HE35, Hei35d, Hei35c, Hei35f, Hei36i, Hei36b, Hei36a, HJFS36, HSBR36, Hei36c, Hei36d, Hei36e, Hei36f, Hei36g, Hei36j, Hei36h, Hei37b, Hei37c, Hei37d, Hei37e, Hei37f, Hei38d, Hei38c, Hei38b, Hei38f, Hei38e, Hei38g, Hei38a, Hei38h, Hei38i, Hei39a]. German [HGT^{+}39, Hei39b, Hei39d, Hei39f, Hei39e, Hei39j, Hei39i, Hei39h, Hei40a, Hei41a, Hei41b, Hei41g, Hei41c, Hei41d, Hei41f, Hei41e, Hei42c, Hei42b, Hei42d, Hei42a, HBF^{+}42, Hei42e, Hei43u, Hei43e, Hei43g, HSBB43, Hei43m, Hei43j, Hei43f, Hei43h, Hei43i, Hei43k, Hei43l, Hei43o, Hei43b, Hei43p, Hei43n, Hei43d, Hei43c, Hei43a, Hei43s, Hei43r, Hei43t, Hei44a, Hei44b, Hei44c, Hei44d, Hei45a, Hei45b, Hei46d, HF46, Hei46f, Hei46b, Hei46c, Hei46e, Hei47a, HvL47, HKvL47, Hei47e, Hei47b, Hei47f, Hei47j, Hei47k, Hei47l, Hei47c, Hei47i, Hei48a, Hei48e, Hei48d, Hei48b, Hei48c, Hei48g, HvW48, HW48, Hei48n, Hei48k, Hei48l, Hei49a, Hei49b, Hei49c, Hei49k, Hei49d, Hei49f, Hei49e]. German [Hei49g, Hei49h, HW49, Hei49l, Hei50e, Hei50a, HHKG50, Hei50b, Hei50c, Hei50d, Hei50j, Hei50g, Hei50h, Hei51c, Hei51e, Hei51m, Hei51a, Hei51f, Hei51i, Hei51d, Hei51j, Hei51k, Hei51l, Hei52a, Hei52b, Hei52c, Hei52r, Hei52f, Hei52g, Hei52p, Hei52s, Hei53a, Hei53t, Hei53e, Hei53f, Hei53k, Hei53l, Hei53g, HR53, Hei53b, Hei53c, Hei53o, Hei53y, Hei53h, Hei53q, Hei53s, Hei53u, Hei53x, Hei53v, Hei53w, Hei54c, Hei54b, Hei54e, Hei54f, Hei55a, Hei55c, Hei55b, HKM55, Hei56a, Hei56g, Hei56e, Hei56d, Hei56h, Hei56m, Hei56i, Hei56c, Hei56n, Hei56j, Hei56f, Hei56k, Hei57e, Hei57b, Hei57c, Hei57d, Hei58d, Hei58a, Hei58k, Hei58l, Hei58m, Hei58e, Hei58g, Hei58h, Hei58i, Hei58t]. German [Hei58f, Hei59b, Hei59a, Hei59d, Hei59e, Hei59g, Hei59h, Hei59i, Hei59k, Hei60b, Hei60c, Hei60f, Hei60a, Hei60i, Hei60j, Hei61c, Hei61a, Hei61b, Hei61d, Hei61e, Hei61h, Hei61i, Hei62a, Hei62f, Hei62h, Hei62i, Hei62b, Hei62m, Hei63a, Hei63c, Hei63d, Hei63e, HB63, Hei63b, Hei63k, Hei63l, Hei64a, Hei64b, Hei64e, Hei64f, Hei64g, Hei64c, Hei64h, Hei65a, Hei65p, Hei65c, Hei65f, Hei65b, Hei65d, HG65, Hei65j, Hei65m, Hei65e, Hei65o, Hei65q, Hei66b, Hei66a, Hei66c, Hei66h, Hei66i, Hei67a, Hei67d, Hei67f, Hei67e, Hei67c, Hei67g, Hei67h, Hei67i, Hei67b, Hei67p, Hei67n, Hei67o, Hei68a, Hei68b, Hei68d, Hei68c, Hei68f, Hei68j, Hei68k, Hei69b, Hei69a, Hei69k, Hei69e, Hei69f]. German [Hei69g, Hei69l, Hei69o, Hei69d, Hei69m, Hei69n, Hei70b, Hei70a, Hei70f, Hei70c, Hei70h, Hei70d, Hei71a, Hei71c, Hei71d, Hei71b, Hei71h, HBD71, Hei71l, Hei71m, Hei71n, HB^{+}72, Hei72f, Hei72b, Hei72c, Hei72g, Hei72j, Hei72i, HnH72, Hei72k, Hei72l, Hei72m, Hei72e, Hei72n, Hei73a, Hei73f, Hei73e, Hei73g, Hei73c, Hei73i, Hei73j, Hei74c, Hei74d, Hei74q, Hei75f, Hei75e, Hei75m, Hei75r, Hei75q, Hei75p, Hei75t, Hei76e, Hei76g, Hei76h, Hei77e, Hei77c, Hei77h, Hei77i, Hei78, Hei79c, Hei79b, Hei79d, Hei79g, Hei79e, Hei80, Hei81b, Hei81c, Hei82b, Hei82c, Hei83a, Hei83e, Hei84d, Hei84-48, Hei84g, Hei84e, Hei84f, Hei84-63, Hei84n, Hei84l, Hei84k, Hei84o, Hei84p, Hei84r, Hei84-54]. German [HH84, Hei84s, Hei84m, HW84b, HMW^{+}84, Hei84w, Hei84-52, Hei84-60, Hei84-59, Hei84-49, HL84, Hei84-47, Hei84-42, Hei84i, Hei84-62, Hei84-41, Hei84-50, HW84a, Hei84-44, Hei84-45, Hei85b, Hei85d, Hei85c, Hei86b, Hei86c, Hei86o, Hei86n, Hei86g, Hei86h, Hei86j, Hei86i, Hei86a, Hei86e, Hei86l, Hei86m, Hei87a, Hei87c, Hei87b, Hei88, Hei89b, Hei89a, Hei89d, Hei89e, Hei89c, Hei89h, Hei89i, HBB^{+}89, Hei89f, Hei89g, Hei89-36, Hei89k, Hei89l, Hei89q, HGW^{+}89a, Hei89s, Hei89u, HGW89b, HR^{+}89, HBF^{+}89a, Hei89w, Hei89m, Hei89x, Hei89y, Hei89z, Hei89-28, Hei89-27, Hei89-30, Hei89-32, Hei89-29, Hei89j, Hei89-33, Hei89n, Hei89-34, HBF^{+}89b, HBF^{+}89c, Hei89-38, Hei89-39, Hei89-37, Hei89-40, Hei90a, Hei90b, Hei91a, Hei93c, Hei93f]. German [Hei93a, Hei96a, Hei00c, Hei01a, Hei01b, Hei02, HR02b, HHH03, Hei03, Hei05b, Hei06, Hei07c, Hei08b, Hei10b, Hei10a, HHvB10, HHHH11, Hei11, Hen00, Hen07, Her76a, Her76c, HvMW79, Her84, Her01, Him90, Hof92, Hof93, HW04, Hof05, HW07, HW11, HW12b, Hof23, Hol93b, Hor66, Hou30, Hou41, Irv67b, Irv67a, Jac01, Joh13, Jun56, Jun64, Jun72, Jun90, Kan02, Kan03, Kar05a, Kar05c, Kar05b, KP07, KHBF43, Kle17, KRW05, KH19, KH47, KH25, Kra87, KF01, LH24, LBBH69, Lan72, Läs13, LL77, LH93, LMOS15, LRD^{+}91, Luc82, Lur95, Mai07, Maj33, MPB^{+}24, MWL^{+}15, Men36, Mey76, MS15, Mor47, Mül40, Nag16, NW95, OS03, Ols63, Ost11, Otr79, PHvMW01, PvM93, Paw02a, Paw02b, Paw06, Pei93a, Pfe69, Pfe77, PFH20, Pop16b]. German [Pop16a, Pop18, PdK23, Pow93a, Pow93b, Pow94, Pow00, Rad83, Rad91, Rec91a, Rec91b, Rec91c, Rec91d, Rec92, Rec96, Rec01, Rec02a, RW02, Rec02b, RKW05, Rec10, Ree19, RH02, Rin61, Rin69, Rin90, Ros98, Rür79, Sal01, San02, Sch01, Sch82, Sch11a, Sch96a, Sch05, Sch07, Sch26, Sch27, Sch30, Sch11b, Sim07, Sim12, Sim79, Smo93, SH22b, SH22a, Som36, Sta36, Sta37, SM41, Tel30, Tel61, Tel71, Tel93, Tel01, Wal89a, Wal89b, Wal90a, Wal90b, Wal93, Wal95, Wal03, Wal05, Wal07, Wal17, Wal20a, Wig96g, vWL93, vBL96b, vL47, vL48a, vL48b, vM02, vM05a, vS14, vW35, vW40, vW41a, vW1 , vW41b, vW48, vW88, vW91, vW02, vWR04, Hay90]. Germans [HB^{+}72, Klo93, NW95, Pei71]. Germany [Cas05, Hei57d, Hei65d, NBH97, Bør12, Bro92, Brü64, Car99b, Cas93b, Gim90, Gla00, Gou46a, Gou46b, Gou47a, Gou48b, Gou48a, GWY05a, GWY05b, Hau22, Hei39d, Hei40b, Hei46d, Hei47g, Hei47h, Hei51e, Hei53k, Hei53g, HR53, Hei56b, Hei56j, Hei67c, Hei84-35, Hei86a, Hei93a, Irv67b, Irv67c, Joh13, KP07, Log96, Mac93, Sch13, Sch07, TW10, Wei46, Wei99, WHF^{+}96, vW57, Lib68]. Gershenfeld [ARK^{+}99]. Gerthsen [EHR^{+}96]. Gesammelte [BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93, Kra86a, TW91, BDR84e, Bro85, Bro86, Mat87, Sch91b]. Gesamtausgabe [Hei68j]. Gesamtdeutsche [DRF90a, DRF90b]. Geschehen [Hei61e]. Geschichte [AH93, All06, Hau22, Hei33d, Hei71b, Hei75g, Kar05c, LMOS15, MGG^{+}95, Rür79, vBL96a, vBL96b]. Gesellschaft [BDR89b, DRF90a, DRF90b, Hei30a, Hei31g, Hei41b, Hei75m, HW07, Mai07, RDP^{+}95, Rür79, URWB95, vBL96a, vBL96b, Tel71]. Gesellschaft/Bergakademie [DRF90a, DRF90b]. Gesellschaft/Die [RDP^{+}95]. Gesichtspunkte [EH38, EH84, Hei72l]. Gesprach [Hei67g, Hei67h, Hei67i, Hei65f]. Gesprache [Hei75e, Hei77c, Ess71, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06, Sch01, Cas02a, HHvB10]. Gestalt [HvW48]. Gestalten [SPS^{+}64]. Gestalten/Kromphardt [SPS^{+}64]. Get [Ano29, Ano33, Tut05, Kar07]. Gets [Ano33, Wit12]. Getting [Lan02, Mon09]. Geĭzenberga [Ano77a, Smo92a, Smo92c]. Gewaltlosigkeit [Hei69g]. gewidmet [Hei43r, Rec02a]. gewinnung [Hei89n]. gewusst [Got02b]. Geyer [EHR^{+}96]. ghost [Hei56o]. Giannetto [Pol00]. Gierer [WMG^{+}92]. Gifford [Cal14, Edd07, Edd12]. Gift [MP01]. Gilmore [HVTK^{+}97]. Gino [Wal08]. Giuseppe [Pol00, Cir88]. Given [Hei75s]. Giving [Wei93]. Glanz [EHR^{+}96]. glass [BPP11]. Gleichgewicht [Hei89j]. Gleichungen [Hei53y]. Glückwunschschreiben [Hei86g]. God [Bab02, Hei67g, MB11a]. Gödel [FW07, Hol18, Mor97]. Goemans [LN06]. Goes [Ano33]. Goethe [Hei41d, Hei52o, Hei67b, Hei74h]. Goethes [Hei67b]. Goethe’sche [Hei41d]. Goetzberger [MGG^{+}95]. Gogh [Bab02]. Going [Sta09a]. Goldman [Sei90a]. Gombás [Hei50a]. good [Hei67e]. Google [Dje07]. Goose [Ano36]. Goose-Steps [Ano36]. Göppingen [Hei89b, Hei89b]. Gordon [Sei90a, Kra84, Mie80]. got [Ols63]. Gott [Hei67g]. Göttingen [Hei89q, Got83, Hei53l, Hei77e, Hei83c, Hei89f, KF01, OS03, OS03]. Göttinger [GRE^{+}01, Hei89f, Hei89q, KF01]. Gottow [Hei43e]. Gottower [Hei43e, Hei89e, Hei89e]. Goudsmit [Hei27b, Lak96, Kae47, LRD^{+}91, Rec91d, Wal90a, vC16, vCH18]. GPU [BPP11]. Grabe [Hei86m]. graded [GH10, MP92]. graduations [HR02b, RW02]. granite [Hei77f]. Grant [Ano33, Sei90a]. graphite [BBF^{+}45, HBB^{+}89]. Grau [LSK^{+}76]. grave [Hei86l, Hei86m]. Gray [SPS^{+}64]. Great [Bro97, Cas97, Cre10, Gre10, Hal09, Hei75b, Bre97, Büh98b, Hei65d, Kae48b, KW87, Lig05, Ols63, Pic08, AFG90, Bes04, vW02, vWR04]. Greatest [Acz01, Hür22, Kra86b]. greedy [MRV24]. Greek [Nic90]. Green [KA71, LB72, Unr72]. Green’s [Ni04b]. Greeting [Hei73a]. Greetings [Hei75r]. Gregg [Wal08]. Gregor [Hei43h, Hei43h]. Greif [Kra87]. Grenze [Hei36e]. Grenzen [Hei38f, Hei84-48, Hei71l, Hei73g, Hei76g, Hei77h, Hei84-62, Hei89-29]. Grenzgänger [Sch05, Sch07]. Greub [SPS^{+}64]. Griffin [Kra86b, Kra87]. Griffiths [Yeh70]. grinding [Hei22b, Hei84w]. Grönblom [Hei43f]. Grönbloms [Hei43f]. Gross [HW84b]. Großbeschleunigern [Hei72l]. Grosse [ARK^{+}99, Hei72a, vRFS^{+}00, Rec01, Bes04, Erm89, vW02, vWR04]. Größen [Hei43j, Hei65d, URWB95, vW48]. Größen“ [Hei43m, Hei44a]. großer [vRFS^{+}00]. Großforschung [MGG^{+}95]. Grosßversuche [Hei44c]. Großversuche [Hei89-33, HW84b, HW48]. Ground [BT78, Bra97, Bec02, Wre75]. Ground-State [Bra97]. Group [Far01, FMV07, He04, HP93, Kor05, Ni04a, Ni04b, Pau93, Wan01, BM93, BK88, CZ81, CGST91, Dun86, ES95, Flo97, GBLW86, GV85, GV86, Gel84, Gre80, KS92, LN06, Low97, RRK^{+}64, Sud06, Tor91, WB03, Hei28b, Hei31b, Ras78, Hei27b]. Group-theoretic [Ras78]. Groups [EHR^{+}96, Wey31b, Wey50, DDI03, JZ11, KSSA78, Pac88, Scu75, Sem03]. Groups/Schmidt [EHR^{+}96]. Groups/Schmidt-Rohr [EHR^{+}96]. Groves [Sei90a, Erm89, Gol95]. Grummt [RFS^{+}91]. Grünbaum [CL83]. Grund [Ano35b, Deh36, Haa34, vW1 ]. Grundfragen [Hei58g]. Grundkurs [MGG^{+}95]. Grundlag [Hei35e]. Grundlage [GRE^{+}01]. Grundlagen [Ano35a, Ben35, Clu37, Deh37, DS43, Hei34i, Hei34f, Hei34g, Hei34h, Hei35f, Hei36h, Hei42e, Hei43t, Hei44d, Hei45b, Hei47i, Hei49l, Hei59i, Hei73j, Hei89n, KED^{+}60, Mos42, Som36, Str62a, Str62c, Wil36a]. Grundlagenforschung [Hei56n]. Grundlegende [Hei59b]. Grundphänomene [Hei39b]. Grundprinzipien [Hei27f]. Grundprobleme [Hei49e]. Gründung [Hei30a, RDP^{+}95]. Gruppentheorie [Hei28b, Hei31b, Wey31a]. Grußwort [Hei75r]. Gu [HVTK^{+}97]. guests [GP92c]. Guide [Rod19, Sei90a]. Guido [Sta11a]. Gun [Dah22]. Gustav [Cas79a]. Gutachten [Cas03, Hei01b, HR02b, RW02, Hei02]. Gutachten- [Cas03, Hei01b, HR02b, RW02]. guter [Hei67e]. Gymnasium [Hei89b, Hei89b].

H [Adl03, Ano94, Bro85, Bro86, Cru23, Fra94, Jac11a, Jac11b, Kra86a, Mat87, Sch91b, TW91, Wit83, JZ11]. H-type [JZ11]. H. [Hof11]. H.-P [Bro85, Bro86, Kra86a, Mat87, Sch91b, TW91]. Haas [Deh36, HEB^{+}80]. Habilitationen [Cas03, Hei01b, HR02b, RW02]. Habilitations [HR02b, RW02]. Hafemeister [Sei90a]. Hahn [GRE^{+}01, Hei54e, Hei68f, Hei69f, KF01, MGG^{+}95, Sim07, Wal03, vS14, BMP87, BM94, Hei54e, Hei68e, Hei68f, Hei69f, KF01, OS03, She83a, She83b, Sim07, Sim12, Wal03, Wal06, vS14]. Hahn-Meitner-Instituts [MGG^{+}95].

Hahn-Meitner-Instituts/Fölsing [MGG^{+}95]. Haigerloch [Läs13, Fec90, FTF01, GORS09, Läs13, Mac85]. Halbleiter [WMG^{+}92]. Halbleiterbauelemente [OSWR86]. Halbleiterbauelemente/Stitch [OSWR86]. Halbleitermarkt [RDP^{+}95]. Halbleitermarkt/Röntgenspiegel [RDP^{+}95]. halbtechnischen [Hei42b, Hei89g]. half [GV86]. half-space [GV86]. Halfway [Bar07]. Hall [BC95, BL08, FS93, Hah45, Hei64b, Hof92, Hof93, Hof23, Mar94, OS03, WR92, Wal93, BDJ57, Ber92, Ber93, Ber96, Cas17, Cas22, Dah22, FS93, Fra93b, Fra93c, Fra93a, GP92c, Hah45, Hof92, Hof93, Hof22, Hol22, Klo93, Luc07, McC24, Rec91a, Rec91b, WR92, Wal93]. Hall-Protokolle [Wal93]. Halle [HH14, Wit83]. halting [Lap91]. Hamiltonian [Fri98, CA76, Fal64, Fri97, Gid71, KBWJ96, TR09]. Hamiltonians [KSGB^{+}83, SMMM87]. Handbook [FBD^{+}22, OSWR86, SPS^{+}64]. Handbook/Brandt [SPS^{+}64]. handing [Hei89b]. handing-over [Hei89b]. handling [Döp42]. Hane [Sei90a]. Hanle [Hei25a]. Hans [Cas79a, Hei75q, BR64, Cas79a, Hei75q]. Happer [BSP^{+}70]. Happy [Rec10]. Harald [Mil12, RS11a]. hard [JD06]. hard-core [JD06]. hardback [Hen11, Sta11b]. Hardbound [Wat01]. Hardy [Adl03]. hare [Hei39c, Hei84j]. Harish [LJ08]. Harish-Chandra [LJ08]. Harmonic [BSP^{+}70, GV86, New35, BDH88, CGM^{+}71, Dun86, Lun84]. harmonics [Gel84, Gre80]. Harmonie [Hei67h, Hei67i]. Harmony [Hei67h, Hei67i]. Harnack [Hei31g]. Harnack-Haus [Hei31g]. Harper [Ano68, Bun71]. Harriet [CS94]. Harteck [Erm89]. Harvard [Par06, Tur06]. harvest [Bra09]. Haus [Hei31g, RDP^{+}95, URWB95]. Havlin [MGG^{+}95]. Hawking [Pic08]. Hayes [Bor53]. hbk [Lan02]. Head [Mon09]. headed [Gou47a]. Health [KED^{+}60]. Heat [Ni04b, Hei24c, KH47]. Heavy [Bør12, Dah99, Hei52n, DDH40b, DDH89e, Hei38a, Hei84i, vW41b, Wat01]. Hechingen [Läs13, Läs13]. Hedwig [Bro72, BB69, BB05]. Heelan [Bab02, HH75]. Heelans [Hei75h]. Heft [FS93]. hegemony [Cus94, Lev95]. Heidegger [Sei90b]. Heidelberg [Büh95, Eis00, Hen86, Hen11, Kra86a, Rec03]. Heimendahl [SPS^{+}64]. Heinrich [URWB95, BHN98, Hen98]. Heinrich-Hertz-Institut [URWB95]. Heirs [Tau94]. Heisenberg [Adl03, Ano35a, Ano35b, Ano51, Ano84a, Ano06, Ano08, ARK^{+}99, Bar04, Bed09, Ben30, Ben34, Ben35, Bes04, BMG^{+}50, BCL^{+}69, BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93, B^{+}61, Bor58, Bor72, Bor53, Bro72, Bro85, Bro86, Bro95, BSP^{+}70, Büh95, Büh98a, Bun71, BW04, Car97, Cas75, Cas74, Cas93c, Cas95, Cas03, Cas05, Cir88, CL94, Clu37, Cru23, Dah02a, DPL96, Deh36, Deh37, DS43, DRF90a, DRF90b, Dür05, Eck12a, Eck12b, EHR^{+}96, Eic50, Eis00, Fay07, Fel54, Fer07, Fis02, Fis10b, Fis12b, Fis15, For09, Fri98, Fue61, Ful15, GZ92b, GZ92a, GRE^{+}01, Got02b, Got04, Gou76, Haa34, HEB^{+}80, Hah06, Hah10, Han62, HVTK^{+}97, Hen86, Hen11, Her76a, Her76c, HMSH78, HvMW79, Her84, Her01, Hof11, Hor66, Huf71]. Heisenberg [Jac01, Jac03, Jac11a, Jac11b, Joh59, Jor71, Kan03, Kle17, KW93, KRW05, KED^{+}60, Kra86a, Kra96, Kra07, KF01, LW34, Lak96, Lic16, LSK^{+}76, Lud01, Mac78b, Mac58, Mar73, MGG^{+}95, Mat86b, Mat86a, Mat87, May87, Mil12, Mil58, Mos42, MP77, Mot85, OSWR86, OS03, Otr79, Par06, Pau31, PHvMW01, PvM93, Paw02a, Paw06, Pol00, Pop18, Pye81, Rec81b, Rec85a, Rec85b, Rec86, RFS^{+}91, Rec91a, Rec91b, Rec92, Rec96, Rec01, RW02, Rec02b, Rec03, RKW05, Rec09, Rec10, Rob82, RDP^{+}95, RDM^{+}12a, SZ74, Sap10, Sch82, Sch05, Sch07, Sch59, Sch02a, Sch26, Sch30, Sch00b, Sch91b, Sei90a, Som36, Sta11a, SCH^{+}78, Str62b, Str62c, SPS^{+}64, Tel93, Tre78, Tur06, URWB95, Wal92, WMG^{+}92, Wei85, Wig67, Wil36a, Wil36b, Wit83]. Heisenberg [vMHW85, vM93, vM96, vM99, vM01, vM05b, vRFS^{+}00, vS14, Aas00, Aas02a, Aas02b, Aas02d, Aas02c, Aas02e, Abd97, Aga02, AMS04, Alt06, AS03, Ano29, Ano33, Ano34, Ano35b, Ano44b, Ano54, Ano61b, Ano68, Ano71, Ano73, Ano75, Ano76, Ano77a, Ano77b, Ano83a, Ano98, Ano99, Ano02a, Ano04a, Ano05, Ano11, Ano12, Ano20, Ano25, Anoxx, Arg09, Ari23, AG88, AFW94, AP22, BT78, BC09, Bac22, Bac52, BT96, BG96, BKBS93, BM93, BK88, Bec02, BDH88, BK04, Bel85, Bel92, BMP87, Bt87, BPP11, Ber95, Ber02, BL03, Ber04, Bes98, Bes04, Bil03, BHX96, Bla86, Ble90, Blo76a, Blo76b, Blo85, BDR84c, BDR84d, BDR84e, BDR85b, BDR86, BDR89b, BRL06, BJLR17, Boh62, Bok04, Bok06, Bok10]. Heisenberg [BB96, B^{+}61, Bor70, Bor49, Bal07, Bra97, BM94, BM11, Bre23, Bre97, Bri65b, Bro93, Bro02a, Bro02b, Bro71, Bro73b, Bro77, BM84, BM86, BR87, BR96a, BN08, Bru76, Büh98a, Bun77, BW04, Cam05, Cam06, Cam07a, Cam07b, Cam09, Cap57a, Cap57b, CL05, CBS20, Cap01, CK95, Car76, CS94, Car99b, Car02a, Car02b, Car03, Car10, Cas76, Cas78, Cas79b, CR84, CB84, Cas92a, Cas92b, Cas92c, Cas93c, Cas95, Cas01, Cas02c, Cas02d, Cas09, Cas11, Cas14a, Cas15, CDvW77, Cav14, CZ81, CGST91, Ces18, Cha92, CM02, CK02a, CK02b, Chi01, CRR93, Cla23, Cle09, CA76, Cle06, Cop81, Cor79, Cor71, CP84, Cre10, CG14, Cul05, Cus86, D’A03, Dah02b, DMdA06, DKLN92, DPL94, DM99, DLW08, DC87, DDI03, DMCF06, DeG74, DJR93, Del13]. Heisenberg [DC48, DGH^{+}12, DGS99, DL08, Dir28, Dir64, Dir76, Dör05, DD13, DRB86, Dun86, DB71, Dür76, D^{+}92, Dür05, Dur00, Dys51a, Dys51b, DLS76, Ech96, Eck33, Eck90, Eck92, Eck01, Eic50, El 00, EMGM94, ES95, Eng09, Enz83, Erm89, Ess71, EHUY92, FO94, FO99, Fal64, Fav05, Fed04, Fed18, FR04, FR08, FR13a, Fer99, Fin96, Fis02, Fis15, Fis64, FT94, Flo97, Foo94, Fra94, Fra10, Fri97, Fri13, Fri11, Fue61, Ful15, FMV07, GBLW86, GbK77, Gar91, GV85, GV86, Gel84, GH10, Gem87, GA02, GZ92b, GZ92a, GLSC01, Got02a, Got02b, Gou90, Gou47c, Gre62, Gre63, GKZL96, Gre80, GS12, GWW^{+}09, HN55, Haa34, Hah10, Hah11, Hak02, HR02a, Han79, Har73, He04, Hea77, Hea11, Hee75, Hei54a, Hei72a, Hei35g, Hei36i, Hei36e]. Heisenberg [Hei47f, Hei48m, Hei51k, Hei56c, Hei65f, Hei67h, Hei67i, Hei70j, HBD71, Hei72m, Hei74e, Hei75n, HDW75, Hei79c, Hei80, Hei83a, Hei84a, Hei89t, Hei91a, Hei01b, Hei02, HR02b, Hei04, Hei05b, Hei08a, Hen93, Hen98, Hen00, Hen10, Her69, Her76a, Her76b, Her76c, HvMW79, Her84, Her01, HBB^{+}07, HN99, HK00, Hof02, Hol84, Hol00b, Hol05, HS81, HS83, Hor66, HS70, Ing70, Ing32, Isa82, IM95, Jac01, Jad14, Jam76, JD06, Jau93, Jau29, Jes98, JLP^{+}02, JZ11, JKS13, JTD^{+}98, Kan02, Kar05b, KBWJ96, KF02, Kas11, KW02, KW03, Kat96, KI64, KSSA78, Kav24, Kem85, KKM^{+}09, Kle54, Kle78, KLV80, KSGB^{+}83, KS92, Kle17, KRW05, Koo89, Kor05, Kra95, Kra93, KF01, KJ06, KF07, KA71, KGNS10, KPS^{+}95, Kva76]. Heisenberg [Lac02a, Lac02b, LM87, Lan59, Lan72, Lan02, Lan08, Lap90, Lap91, LB72, LL77, LN06, LBS^{+}04, LL86, LM29, LTG^{+}12, Lin07, Lin08, LFK99, LL82, LH01, LJ08, LWZ10, Loc07, Loe63, LRD^{+}91, LS93a, LS93b, Log99, LD66, Los61, Low97, Luc82, Lud01, Lun84, Lus10, sMDN^{+}11, Maa05, Mac78a, Mac77, Mac86, Mac87, MP92, Man11, Man91, Mar08, Mas04, Mas14, MC97, MRV24, McC24, MC92, Meh87, Mes04, Mey76, MS15, Mie80, Mie81, ML71, Mil85, MMSV98, Mon09, Mon04, MB11b, Moo93a, Moo93b, Mor71, Mor72a, Mor72b, Mor72c, Mor97, NAZ02, NKD^{+}11, Nar93, Neu99, Ni04a, Ni04b, Nis32, OD03, O’F92, OS07, OS03, Ome75, Ome79, Ore98, Ost11, Otr79, Oza02, Oza03, Pac88, PC14, Pan02, PLS02, Pau34, PHvMW01, PvM93, Paw02a, Paw02b, Paw06]. Heisenberg [Pei71, Pei92a, Pei92b, Pei93a, Per88, Per64, Pfe77, Phi92, Plo10, Pol95, Pop18, PL91, PDL92, Pow76, Pow93a, Pow93b, Pow94, Pow00, Pow02, PSZM02, Pug80, Raz11, Rec81a, Rec84, Rec91a, Rec91b, Rec91c, Rec91d, Rec92, Rec93, Rec96, Rec01, RW02, Rec02b, RKW05, Rec10, Ree24, RH02, Rhe91, Rig87, Rob69, Rob80, Roh85, RS94, RW92, Ros98, Ros60, Ros79, Rot93, Roy78, RDM^{+}12b, Rud58, Rus64, SZ73, SHI73, Sal01, SMMM87, San02, San81, Sch01, Sch82, Sch11a, Sch05, Sch07, SAR^{+}09, Sch02b, Sch99, Sch00a, Sch26, Sch66, Sch15, Sch22, Scu75, Sei90b, SK73, Sem03, Ser77, Set11, Shi02, Shi78, Shi83, Shi86, Sil13, ST65, SK09, Sla72, Smo93, Smo92a, Smo92b, Smo92c, Smo92d, Sno49, Sny82]. Heisenberg [Sor02, Spi98, Squ67, SL70, Sta11b, Str62a, Sud06, SF71, Swe93, Tan04, Tau94, TW91, Tel76, Tel93, Telxx, Tel01, Tem91, Ter82, Tho80, Tiw14, TR09, Tor91, Tre01, Tri02, TBL02, Tut05, Tya63, Unr72, VBN95a, VBN95b, Van41, VZ09, VL97, Vya99, Wag81, Wal90a, Wal20b, Wal22, Wal83, Wan98, Wan01, War90, Wei98, Wei93, Wel92, WF19, WB03, Wig76, Wil84, Wil99, Wis98, Wit12, Wol14, Wre75, Wüt16, Yaf00, Yam02a, Yam02b, Yam02c, Yan13b, Yeh70, Zak96, Zan02a, Zan02b, Zel87, ZG96, Žup65, d’E92, dB58, dB82, dB90, dG14, dR99, dSL02, dlP80, vD15, vBN95c, vBW88, vM02, vM05a, vS14, vWvdW77, vW02, vWR04, Ano94, CD94, Dre92, Gar91, Gre10, Lau07, Mat86a, Mot85, Paw08, Roh85, Wei85]. Heisenberg [Ber93, Bro97, Car99a, Cas91, Cas93a, Cas97, Cas02a, Hol93a, Lal15, LH93, Pei93b, RS11b, Swe93, Wal99, Wil94]. Heisenberg- [Pug80]. Heisenberg-brev [Sch15]. Heisenberg-Limited [LBS^{+}04, BM11, HBB^{+}07]. Heisenberg-Limits [Sch11a]. Heisenberg-Mikroskop [URWB95]. Heisenberg-Programm [Ano84a, Ano84b, Ano83a]. Heisenberg-Programm/1-Megabit-Speicherchip [Ano84a]. Heisenberg-type [HR02a]. Heisenberg/Braun [LSK^{+}76]. Heisenberg/Die [Eic50]. Heisenberg/DVT [DRF90a, DRF90b]. Heisenberg/DVT-Preis [DRF90a]. Heisenberg/Schrödinger [Fri98, Fri97]. Heisenberg/Strobl [HVTK^{+}97]. Heisenberg/The [Eic50]. Heisenberg/Werner [WMG^{+}92]. Heisenberg/Zee [vRFS^{+}00]. Heisenbergs [Hei79c, Paw02b, WR92, vM02, Kar05b, Pow93a, vM05a]. Heisenberg’schen [Lan72, Sch30]. Heitjans [ARK^{+}99]. Heitler [Hei62c]. Held [Meh73, CDvW75, CDvW79, CCJ^{+}34, GZ95, Hei31g, Hei43p, Hei52p, Hei53l, HMW^{+}84, NBH97]. Helgoland [Rov20, Rov21]. helium [BH23a, BH63a, Hei24a]. Heliumatom [BH23a, BH63a]. Heliums“ [Hei24a]. Heller [Jun56, Jun64, Jun72, Jun90]. Helmut [Cas05, Eck12a, Eck12b, Hen11]. Helmuth [PFH20, PFH20]. help [Hei41f, Hei89-36, Hei89x]. Heman [Hei89s, Hei89s]. hemmungslosen [Hei74c]. Henri [Adl03]. Henry [Sei90a]. Hentschel [Wal08]. Heraeus [DRF90a, DRF90b]. Herausgegeben [Eis00]. Herbert [Sim79, Sim79]. Herbig [Cas05]. Herder [RFS^{+}91]. Here [Ano29, Bro97, Cas97, Bre97]. heretics [DT18]. Herewith [Ols63]. Heritage [PFH20]. Hermann [HEB^{+}80, Hei28b, Hei31b, LSK^{+}76, Mac78b, Pye81, Rec86, SCH^{+}78, Hei28b, Hei31b, Tre78]. Hermeneutic [Bab02]. hermeneutics [Hee95]. Hermite [Per88]. Herneck [Hei72a]. Hero [Tho97]. Herold [EHR^{+}96]. Herrmann [EHR^{+}96]. Herrn [HVTK^{+}97, Hei31f, HGW^{+}89a]. Herstellung [vW41b]. Hertz [Cas79a, URWB95, BHN98, Cas79a, Hen98]. hervorgerufenen [Hei32h]. Herzberg [URWB95]. Herzberg/EU [URWB95]. herzlichen [Rec02a]. Hess [Ano51]. Heuristic [dR99]. heuristics [BJLR17]. heutige [Hei65b]. Heutigen [Bor58, Hei53t, Hei53c, Hei53u, Hei55b, Hei56f, Hei57c, Hei58e, Hei59a, Hei60a, Hei62b, Hei63b, HG65, Hei66a, Hei68c, Hei70c, Hei72j, Hei72i, HnH72, Hei74d, Hei76e, Hei79d]. Hewitt [BSP^{+}70]. Hexeneinmaleins [vM02]. Hidden [Mac85, ZG96, BC09]. hierarchies [LTG^{+}12]. Hierarchy [HK00]. High [Cus86, DGH^{+}12, Hei84-28, BSP^{+}70, Hei36d, Hei37b, Hei38c, Hei38b, Hei55e, Hei56l, Hei83g]. High-Energy [Cus86, Hei36d, Hei37b]. Higher [RG72, GWW^{+}09, LH24, GbK77]. Higher-Derivative [RG72]. highly [Hei84l]. hike [Wei98]. Hilbert [Hei56m, Hei43u, Hei56m, Hei56o, Hei58w, Hei84-36, WB03]. Hilbert-Raums [Hei56m]. Hildegard [Hei53b]. Hilfe [Hei41f, Hei89-36, Hei89k, Hei89x]. Hillel [UM92]. Hintertreppe [Fis10a, Fis12a]. Hinweis [RDP^{+}95]. Hinweise [URWB95]. Hiroshima [Wal24b, Sei90a]. Hirsch [Cas05, Kan03]. Hirsch-Heisenberg [Cas05, Kan03]. Hirzel [Ano35a, Ben30, Ben34, Ben35, Clu37, Deh37, DS43, Mos42, Str62b, Str62c, Wil36a, Wil36b]. histogram [dSL02]. histoire [Kar07, Lur95, VB15]. historia [Kar06]. Historian [Cas22]. Historical [Bru80, Cas00, Cas11, GL23, Hak02, Jam74b, Kra07, Lan08, Lau07, Lev95, SS16, Ano99, Cus94, Dör05, Hei29b, Mar08, Mil90]. Histories [Pei97]. Historiography [Wal24a]. historische [Hei29b]. Historischen [AH93]. History [Ber93, Ber03, Dur00, FW20, FBD^{+}22, Hei75j, Hol93a, Jun58, Kar05c, Kra86a, Kra07, Lan08, LH93, LMOS15, Lib68, Mac85, Mat86b, Nic90, Pei93b, Pop16c, Ros00, Rür79, Sch00a, Sei90a, Sim10, Swe93, UM92, Wal17, WP85, Wil94, Wüt16, Ano02b, CHW91, Cas93a, Cas17, ET07, GHW09, Hau22, Hei33d, Hei73d, Hei75l, Irv67b, Jun82, Kar07, KHFA67, Lud01, Lur95, PFH20, Pop16b, Pow93a, Pow93b, Pow94, Pow00, VB15, Wal24b, WC84, vBL96b, Hei71b, Sei90a]. Hit [Ano04b]. Hitherto [O’F92]. Hitler [Bed03, Bed04, Bey82, Cas02a, Hof02, Jor78, Kra87, Sch01, Bal13, Bal14, Bed03, Bed04, Ber96, Bey77, Bey82, Bro92, Cor03, Hol93b, Kar05c, KW05, Kar07, KP07, Luc16, MP01, Pop16b, Pop16a, Pop21, Sch01, Wal24b, Wal24a, Wei46, Wei99, vC16, vCH18, Pop16c, All06]. Hitlera [Kar06]. Hitlers [KP07, Kra87, Kar05c, Pop16b, Pop16a]. HLISP [KSSA78]. hoax [Dur00]. hochauflösende [DRF90a, DRF90b]. Hochschulbau [Ano84a]. Hochschulbau/Faltings [Ano84a]. Hochschule [Cas79a, Hei72j, Hei72i]. Hochschulen [Hei68a]. Hochschullehrer [Hei36e]. Hoffman [FS93]. Hoffmann [FS93]. Höhenstrahlung [Hei32e, Hei36j, Hei38d, Hei38a, Hei41c, Hei84i, Hei84n]. höherer [LH24]. holes [Bes04]. Holton [Fay07, Par06, Tur06]. homogeneous [Bt87]. Honor [Ano29, Bab02, GZ95, MF69, Men84, Rec91a, Rec91b]. honorary [Hei89-27]. Honour [CL83]. Honours [Hei64b]. Hopes [Hei84-53, Hei53j]. horizon [Geĭ87]. Horrifying [Hen00]. Hörsaal [Hei60c]. hos [Aas02e]. House [Hei31g, Hei53b, Irv67c]. Houtermans [ARK^{+}99]. Hove [vV78]. Howard [Sei90a]. Hoyer [Hei75g]. Hrsg [Ano06, Hah06]. Hubbard [CA76]. Hüfner [RFS^{+}91]. Hughes [BSP^{+}70]. Human [Cop81, Gra00, Luc04]. humanism [Ano66]. humanistische [Hei49h]. humanities [Hei49h]. Humboldt [Ano71, BDR89b, Hei56g, Hei69b, Hei69e, Hei89u, Hei89-28, Pfe69, Hei69b, Hei71n, Hei89u, Pfe69]. Humboldt-Stiftung [Ano71, BDR89b, Hei56g, Hei89u, Hei89-28, Hei71n]. Humboldt-Stipendiaten [Hei69b]. hundred [AK15]. hung [HJ43]. Hungarian [Hei58p, Sch66]. hunt [vCH18]. Hupfeld [BM84]. Hydrogen [Lak96, Rig02, JLP^{+}02, LH01, Tel30]. Hypothesis [Gau14].

I. [Men84]. I/Westphal [SPS^{+}64]. Ice [BSP^{+}70]. Ice/Schimmel [BSP^{+}70]. Idea [Kae31, Hei48i, Hei58o, Kae48a]. Idealistic [Kae31]. Ideas [Bro97, Cas97, HVTK^{+}97, Hei52e, Ble90, Bre97, Fin96, Hei53o]. ideological [Hei27e]. IEEE [IEE06]. If [Büh98a, Hei68h]. ignited [Gre05]. Ignored [Pop16c]. ihre [Hei39b, Läs13, NW95, vBL96a, vBL96b]. ihrer [SH22a, vBL96a, vBL96b]. II [ARK^{+}99, BDR89a, BSP^{+}70, Eis00, Hei24c, Hei40a, Hei84-48, Hei89h, OSWR86, Rec86, URWB95, Wig96m, vMHW85, vM99, BS47, BHJ26, BHJ63, BHJ67, BGĬ90, BN08, Cas17, CM02, Dys51b, GV85, Hei24c, Hei27c, HP30, Hei32g, Hei39j, Hei40a, Hei43j, Hei43i, Hei56o, HY72, Hei84-48, Hei89h, KH19, Kra86b, LH93, Moo93b, Mor72b, Sim12, Yan13a]. II/O’Dell [BSP^{+}70]. III [BDR93, Bro95, Hei84-48, PvM93, vM93, vM01, BT78, GV86, Hei33b, Hei44a, HKM55, Hei65h, Hei84-48, Mor72c]. ill [Tur06]. illus [Hay90, Lib68]. Illustrated [Hol93a, LH93]. Illustration [Roy78]. Illustrative [Hei36c]. im [Ano84a, Ano84b, Bey82, BMG^{+}50, BDR89b, Bor58, BH23a, BH63a, DRF90a, DRF90b, Ess71, Hau22, Hei31g, Hei35b, Hei41d, Hei43p, Hei47a, Hei47c, Hei51d, Hei51k, Hei52p, Hei53d, Hei53h, Hei53q, Hei56g, Hei57d, Hei60c, Hei61e, Hei64b, Hei66b, Hei66c, Hei69d, Hei70b, Hei72e, Hei73c, Hei75m, Hei81b, Hei81c, Hei83e, HMW^{+}84, Hei84-50, Hei85c, Hei86o, Hei86e, Hei87b, Hei88, Hei89c, Hei96a, Hei01a, Hei06, Hof92, HW07, HW11, Läs13, Mai07, Pop18, RKW05, RDP^{+}95, Sim07, URWB95]. image [Hei53c, Hei67b]. Images [Sei90a]. Impact [Bro71, MTV22, Yam02b, Yam02c, Hei49b, PLS02]. implementation [OD03]. implied [Mes04]. Importance [Cus86, BDR85b]. important [Hei48l]. improved [DRB86]. Inaugural [Hei75s, Hei86c]. Inauguration [Rec91a, Rec91b, Hei72f, Hei86o]. incertidumbre [FR04, FR08]. incertitudes [dB82]. incoherent [Hei31i]. incontri [Hei84-55, Hei00a, Hei08c]. Inconvenient [Sim10]. increase [HJ43]. Indefinite [Hei72h, Hei84u, Hei84-36, Hei58w]. indeterminación [Bla86]. Indeterminacy [Bac52, BKBS93, Cam07b, Han79, Jam74a, PC14, Pol95, BK04, Bel85, Mes04]. Indetermination [Jau29, Cav14]. Indeterminism [Bru76, Sil13]. Indictment [Lau58a]. Induced [GR09]. induction [Mas04]. industrial [PFH20]. Industrie [Rec85a, Hei70f]. Industrie/Akademie [Rec85a]. Industriearchäologie [PFH20]. Industriekultur [PFH20]. Industrielandschaft [DRF90a, DRF90b]. Industry [Rad83, Rad91, Hei56a, Hei70f]. induzierter [URWB95]. Inequalities [Bun77, DGH^{+}12, Fal64, FMV07, HS70, Pow76]. inequality [CP84, LM87, LL86, May87, dlP80]. Infeld [Cas75]. Infinite [Fis64, ML71, Mor71, Mor72b, Sch02a, Sch02b, BT78, Mor72c]. Infinite-Spin [ML71]. Influence [Dir76, GR09, Bel85, BH24a, BH63b, Hei24c, Hei25a, Pfe77]. Information [GZ92b, Har73, Wal08, D’A03]. information-disturbance [D’A03]. Informationsseminar [URWB95]. Informationstechnik [RDP^{+}95]. Informationstechnologie [URWB95]. Informationstechnologie/Ausschreibung [URWB95]. infrared [Sch02a, Sch02b]. Inhalt [Hei27e]. inherent [Tem91]. inhomogeneous [BG96, BP75, CL05, DPL94, DPL96]. inkohärente [Hei31i]. Inner [Hei84a, Kra86a, Mot85, Mat86a, Roh85, Wei85]. Innere [MGG^{+}95, Hei75p]. innocence [FW07]. Innovationsförderung [RDP^{+}95]. Innovationsförderung/1994 [RDP^{+}95]. inquiries [Mil90]. ins [Hei75p, MGG^{+}95, vRFS^{+}00]. Insight [Ber91, CS94]. Instead [For22]. Institut [AH93, Ano44a, Cas79a, Hei41a, Hei43p, Hei53l, Hei71b, HMW^{+}84, Hei89i, Rec81b, Rec85a, Sim07, URWB95, Wal05, CCJ^{+}34, Hei53l, Hei62j]. Institute [Cas11, CCJ^{+}34, Hah40, Hei41a, Hei43p, Hei71b, Hei89i, Läs13, Rec96, Sim07, SPS^{+}64, Wal05, Wal07, Wat01, vBL96a, vBL96b, Ano44a, CR13, Cas79a, Hei60c, SW05, Sch11c]. Institutes [Läs13]. institution [Hei71b]. Institutions [FW20, vBL96b]. Instituts [Hah40, Rec96, Hei60c, Hei71b]. Instituts/Fölsing [MGG^{+}95]. integrability [DPL94, DPL96, PDL92]. Integrable [GWW^{+}09, DM99]. integral [Sla72]. integrals [LL82, SHI73]. Intellectual [BLW19]. Intelligence [Bør12, Jon78]. intelligent [Ras78]. intelligenten [GRE^{+}01]. Intensität [SH22a]. intensity [SH22a]. Intensivkurs [HMSH78]. Interaction [Ano11, CJ72a, CJ72b, Gid71, HEB^{+}80, Her69, MC97, NKD^{+}11, EHUY92, Hei56k, Hei84-48]. Interaction-based [Ano11, NKD^{+}11]. Interactions [San81, BSP^{+}70, Oza02]. interest [Mil12]. interesting [dlP80]. interferometry [CL94]. interior [Hei83i]. International [BCEP94, Hei75s, Hei76b, Meh73, Rec81a, RDP^{+}95, TW10, BDR89b, CCJ^{+}34, Hau22, Hei52m, Hei65d, NBH97, Pfe69, SW05]. Internationale [BDR89b, DRF90a, DRF90b]. internationalen [Hau22, Hei65d]. interned [Hof23, Hol93b]. internierten [Hof23]. Internierung [Hof92]. internment [Hof92]. Interpretation [And89, Bok10, Bun77, EHR^{+}96, Eng09, EH38, Fra10, Geĭ77b, Hei62a, HB63, Hei77b, PC14, RFS^{+}91, Sta11b, dB90, Ano02b, Bac22, Bor61, Bor63b, Cam09, EH84, Geĭ77a, Hei55d, Hei56h, Hei64d, Hei91b, Hei07a, dB82, dB82, Hei39b]. Interpretations [BSMB99, FBD^{+}22, Jam74b]. Interpreting [Fri13]. interpretive [Rot93]. Interview [Hei36e, Hei48i, Hei50e, Hei50c, Hei51e, Hei67j, Hei68b, Hei68h, Hei69i, Hei70f, Hei74c, Hei89r, Hei89t, Wig96j, Ano71, Ano73, Eic50, Hei51e, Hei65f, Hei89s]. interviews [Erm89]. intrigue [FB01]. Introduced [Mar94]. Introduction [Bro72, CK02b, Hei36b, Hei66d, Hei71g, Hei84-57, Hei84v, Hei05a, KED^{+}60, WMG^{+}92, vRFS^{+}00, Ano35b, CK02a, Cru23, Haa34, Hol98, JKS13, Sem03, Hei43h, Hei43o, Hei62f, Hei67d, Hei93b, Wal08, Wig67, Hei93d]. Introductory [Ome79, Rec02b, BSP^{+}70, Hei69b]. invariance [BPS74, BP75, JU82, ZG96]. Invariants [Fed18, Tem91]. invented [NW95]. Inventing [Gol92]. inventory [KHFA67]. Inverse [Gro96, DC87, JLP^{+}02]. Inversion [He04]. investiert [URWB95]. Investigate [KH19]. investigating [Gou47a]. Investigation [MWL^{+}15, FB00]. Investigations [Hei59g]. Ion [HEB^{+}80, Tel30]. Ionen [BH24a, BH63b, Hei24c, BDR84j, Hei24c]. Ionenlaser [Ano84a]. Ionenlaser-Plasmarohren [Ano84a]. Ionics [RFS^{+}91]. Ionics/Kovetz [RFS^{+}91]. ions [BDR84j, BH24a, BH63b, Hei23a, Hei24c]. Irene [Bro72]. Iron [Luc05]. Irrational [Ber35]. irre. [Büh98a]. Irreducible [MMSV98, KSSA78]. Irreversibility [Bru76]. Irrtümer [Bed03, Bed04]. Irving [Lib68, Hei66e, Hei66f]. Isaac [Bes04]. ISBN [Ano08, Bok10, Büh95, Fra94, Hen86, Hen00, Hen11, Kra07, Lan02, Lan08, Mar94, Pol00, Rec03, Sta11b, Tur06, Wil94]. ISBN-13 [Bok10]. Ising [Mor72a, Rus64, SF71]. islychenia [GE38]. isospectral [LTG^{+}12]. Isospin [HP93, Pau93]. isotope [Hei41f, Hei89-36, Hei89x, vW41b]. Isotopes [NBDG40]. Isotopic [Zel87]. Isotops [Hei41f, Hei89-36, Hei89x]. isotropic [Hei48j, PDL92, Pow76]. issue [FS93]. issues [W^{+}87, WdC02, Sei90a]. istolkovanii [Geĭ77a]. Italian [Alt06, Arg09, Cap01, Cas75, Cas14b, Cav14, Ces18, Fer99, GA02, Hei28c, HSBA60, Hei60e, Hei84-55, Hei00a, Hei08c, Lud01]. Italy [Meh73]. IV [AH57, ARK^{+}99, DRF90a, DRF90b, Eis00, vM96, vM99, vM01, vM05b, AH57, Far01]. IV/Teil [ARK^{+}99]. IX [Kra86a, Mat86b]. izluchenija [GE39a, GE39b].

J [DPL96, Fri98, Got04, Han62, Huf71, Joh59, May87, Sch59, Sch02a, Sei90a, Wit83]. jacht [vC16]. Jadczyk [Lic16]. jadra [Geĭ36]. jadrowa [Kar06]. Jaeger [Wal08]. Jaffe [Hei72a]. Jahr [URWB95]. Jahre [AH93, GHH02, HB82a, Hei38g, Hei42a, Hei49a, Hei51a, Hei56g, Hei67c, RDP^{+}95, Jor71, vM05a]. Jahren [Cas79a, Fie80, Hei53b, Hei86f, Hei53e]. Jahrestag [RDP^{+}95]. Jahrestagung [Ano84b, Hei73a]. Jahrfeier [Hei58k]. Jahrhunderts [KED^{+}60]. Jahrhunderts/Bellman [KED^{+}60]. Jahrhundertwissenschaft [SCH^{+}78, Tre78]. James [Bes04, Lev95]. Jan [HBB^{+}89, HBF^{+}89c]. Januar [Hei71h]. January [HBF^{+}89c]. Japan [RDP^{+}95, GWY05a, GWY05b]. j’avais [Hei68h]. Jaynes [IM95]. Jefferson [Sei90a, Hol86]. jelementarnyh [Geĭ56, Geĭ77c]. Jenseits [Sch11a]. Jeremy [Got04]. jetzigen [Hei43o]. Jewish [Wei46, Hei36i, Men36, Rec92, SM41, Wei99]. Jews’ [Sta37]. Joel [Dje07]. Johann [Wit83]. John [CS94, GZ95, GW96, Wig96c, Wig96h, WG96]. Jones [Sei90a]. Jordan [Hei36c, Sch05, Sch07, Bel85, Sch05, Sch07, Sch26, Sch66, dG14]. Jordan-féle [Sch66]. Jordanschen [Sch26]. Joseph [Sei90a]. Journey [FR13b, KH19, Lac02b, Hei75p, Hei83i, Lac02a]. Joy [Ber91, CS94]. Juden [Sta37]. jüdische [Hei36i, Men36, Rec92, SM41]. Jugend [Hen11, Rec09, Rec10]. Jul [DDH89a, Hei89g]. Juli [Hei64c]. Jülich [DRF90a, DRF90b]. July [CDvW75, CDvW79, Hei64c, Hei89g, NBH97, AMS04, DDH89a]. Jun [Hei89-34]. June [Hei72c, Hei89-34]. Jung [BMG^{+}50, MGG^{+}95, Mil10]. Jungk [Hei72a]. jüngsten [Hei39d]. jüngster [Hei34i, Hei34f, Hei34g, Hei34h]. Juni [Hei72c, HH14]. Jürgen [DRF90b]. Jürgen-Geiger-Preis [DRF90a]. Jurisdiction [Hei50e]. Justice [Sei90a].

K. [Eng09, Hei86b]. K.F [Hei86b]. Kaempffert [Kae47]. Kafka [EHR^{+}96]. Kaiser [BSP^{+}70, Hah40, Hei31g, Hei41a, Hei71b, Hei89i, Läs13, Mai07, Rec81b, Rec96, Sim07, Wal05, Wal07, vBL96a, vBL96b, Hah40, Hei41a, Hei89i, Läs13, Mai07, Rec96, SW05, Sim07, Wal05, Wal07]. Kaiser-Wilhelm- [Rec81b, vBL96a, vBL96b]. Kaiser-Wilhelm-/Max-Planck-Gesellschaft [vBL96a, vBL96b]. Kaiser-Wilhelm-Gesellschaft [Mai07]. Kaiser-Wilhelm-Institut [Hei71b, Sim07, Wal05]. Kaiser-Wilhelm-Institute [Läs13, Wal07]. Kaiser-Wilhelm-Instituts [Rec96]. Källén [Hei56o, Hei62d, Nel72]. Kamlah [Hei50b]. Kampf [Hei24b, Hei35b]. Kant [Mon04]. Kantian [Cam05]. Karajan [Sim79]. Kärger [ARK^{+}99]. Karl [Anoxx, Bla86, Car97, Eis00, Erm89, Kra96, Rec86, Bro95]. Karlsruhe [CR13]. Kármán [Hei22a]. Kármánschen [Hei22a]. kart [Clu37, Deh37, DS43, Mos42, Wil36a, Wil36b]. Kaskadentheorie [Hei43b]. Kassel [EHR^{+}96]. Katze [RFS^{+}91]. Kaufmann [HVTK^{+}97]. Kausalgesetz [Hei30g, Hei31e, Hei52a]. Kausalität [Hei51m, Hei30h]. kein [Hei72g]. keine [URWB95]. Kenntnis [Hei65b]. Kenntnisse [Hei43o, Hei53t, Hei53u]. Kept [Cha92]. Keramik [RDP^{+}95]. Kern [Hou41]. Kern-Kettenreaktionen [Hou41]. Kernel [Gre63, Ni04b]. Kernenergie [HB82a]. Kernenergieerzeugung [vW41b]. Kernforschung [Hei68b]. Kernkraftwerke [DRF90a, DRF90b]. Kernkraftwerke/Computer [DRF90a, DRF90b]. Kernmassen [vW35]. Kernmechanik [Hei32i, Hei32j]. Kernphysik [BDR86]. kernphysikalische [Hei54b]. Kerntechnik [Rad83, Rad91]. Kerntheorie [Maj33]. Kernwaffen [Wal05, Wal07]. Kernwaffen- [Wal05, Wal07]. Kernwaffenentwicklungen [Hau22]. Kernwaffenversuche [Hei89-32, Kar05c, All06]. Kettenreaktionen [Hou41]. Kienle [Hei75q]. Kilometer [Hei57b]. Kinematic [Her69, Geĭ77a, Hei25c]. kinematicheskih [Geĭ77a]. kinematics [Hei26f, Hei27e, Hei77j, Hei83b, Hei84-42]. Kinematik [Hei26f, Hei27e, Hei84-42]. kinematiki [Geĭ77b]. kinematischer [Hei25c]. Kinetic [Geĭ77b, Hei77b]. Kirche [MGG^{+}95]. Kirschner [OSWR86]. Kirsten [Dje07]. Klassiker [Bes04]. Klatt [BSP^{+}70]. Klaus [Wal08, Sei90a]. Klein [Mie80, GS12, Kra84]. Kleines [SPS^{+}64]. Kleinknecht [Cru23]. kleinsten [Fis10a, Fis12a, Hei53o, Hei71c]. Kleint [Ano08, Hah06]. Kline [CS94]. km [Hei57b]. Knall [URWB95]. Knall/The [URWB95]. Knew [Klo93]. Knobloch [MGG^{+}95]. Knopf [Cas93a, Pei93b, Wil94, Hol93a, LH93]. Knopf/Cape [Cas93a]. know [Got02b, Kat96]. Knowing [Wig96l]. Knowledge [Ano34, Hei43o, Hei53u, BDR84c, BDR84d, BDR85b, Hei53t, Hei65b, Mai07, Ano61a, HPE^{+}33, Hei43i]. known [Brü64]. Knows [GLSC01, Hei48l]. Koch [HEB^{+}80]. Kofler [EHR^{+}96]. Kohlenstoffgruppe [Hei27b]. Kohlernückstreumantel [BBF^{+}45]. Kohlerückstreumantel [HBB^{+}89]. Kohn [JZ11]. Kollegen [RKW05]. Kollegiums [SM41]. Kolloquium [Hei52f]. Köln [DRF90a, DRF90b]. Köln/20 [DRF90a, DRF90b]. Kommentar [SPS^{+}64]. Kommission [URWB95]. Kommissionsverlag [Wit83]. Komödie [Dür62, Dür80]. Komponente [Hei38d, Hei41c, Hei84n]. konchine [Ano77a]. konec [DF96]. Konferenz [URWB95]. Kongreßsaal [Hei64b]. konnten [Hei67g]. Konrad [Cru23, Cru23]. Konstanten [BH24a, BH63b, Hei24c]. Konstruktion [HW48, HW84b]. Kontext [Hau22, HW11]. kontrolliertes [HMSH78]. Kontroverse [Rad83, Rad91]. Koordinierung [URWB95]. Kopenhagen [EHR^{+}96, OS03, Paw02a, vW41a, Fra05, Kar05a]. Kopenhagen/Luchner [EHR^{+}96]. Kopenhagener [HB63]. Kopernikus [Hei86n]. Kopplung [Hei27b]. Korpuskeln [BDR84f, Hei36d, Hei37b, Hei38c, Hei38b, Hei84l]. Korrespondenzprinzips [Hei25b]. kosmicheskogo [GE39a, GE39b]. kosmicheskovo [GE38]. Kosmische [BBB^{+}84, HMW^{+}84, BDR84g, Hei39f, Hei43r, Hei43p, Hei51f, Hei53l].

Kosmischen [Hei36g, EH38, EH84, Hei39j, Hei39h, Hei43o, Hei53t, Hei53u, Hei84-44, Rec02a]. Kosmologie [vRFS^{+}00, Hei68d]. Kosmologie/Meschede [vRFS^{+}00]. Kovetz [RFS^{+}91]. Kowalski [BSP^{+}70]. Kowarski [Erm89]. Krafft [vRFS^{+}00]. Kräfte [Hei47c]. Kramers [Hea77]. Kranzniederlegung [Hei86m]. Krätz [HVTK^{+}97]. Krätz/Rang [HVTK^{+}97]. Kraus [GRE^{+}01]. kreds [BR64, RRK^{+}64]. Krein [MMSV98]. Krieg [Pow93a]. Krieger [BMG^{+}50]. Kriegsaktivitätkeit [vL48a]. kriegsrelevanter [Mai07]. Kriegstätigkeit [vL47]. Krigens [Boh64]. krigstidens [Sch15]. kring [Sch15]. Krise [Rad83, Rad91]. Kristallen [Hei33c, Hei84-41]. Kristian [Bok10, Bok10, Fra10, Sta11b]. kritika [Geĭ70b]. kritischer [Cas05, HHH03, Kan03]. Kromphardt [SPS^{+}64]. Kudo [RFS^{+}91]. Kugel [DDH42a, DDH89a]. Kugel-Schichten [DDH89a]. Kugel-Schichten-System [DDH42a]. Kugellager/Bislang [RDP^{+}95]. Kuhn [Bok06]. Kundgebung [Hei53q]. künftigen [RDP^{+}95]. Kunst [Hei70d, Hei75p]. Künste [Hei66b]. Künstliche [Hah43]. künstlicher [URWB95]. Kunz [HEB^{+}80]. Kurzbericht [Hei61h, Hei61i, Hei67p]. Kurze [Hei89u]. Kurzer [vW1 ]. Kurzman [Sei90a]. Kvanteteorien [Hei64d]. Kvanteteoriens [Hei35e]. kvantovoteoreticheskoj [Geĭ77b]. kvantovoteoreticheskom [Geĭ77a]. kvantummechanika [Sch66]. kwanty [Inf33].

L [Hei48h, Lan02, Pol00, Sei90a]. lab [Mac85]. Laboratory [For22, DJR93]. Ladungsschwankungen [Hei34e]. Lady [BM84]. Lahti [And89]. L’Allemagne [Gou48b]. Lambrecht [NBH97]. Laminar [Hei84x, Hei52i, Hei84w]. Lanczos [BCEP94, MC92]. land [Enz83]. Landau [RDP^{+}95]. Landau-Programm [RDP^{+}95]. Landmarks [Wal04]. Länge [Hei38h]. Langen [Cas05]. Langenscheidts [BSP^{+}70, GRE^{+}01]. langsomme [Fis41]. Language [Hei96b, Pag82, Eck12a, Hei69o, Rec10, Bor58, Hei60i]. Langville [Dje07]. Laplacian [FMV07, GV85, Kor05, Maa05]. Laplacians [JZ11]. Large [Hei44c, Hei73b, Hei72d, Hei72l, vW48, HW48]. Larson [Erm89]. Laser [Hak02, HVTK^{+}97, OSWR86, vRFS^{+}00]. Laser/Mantegna [vRFS^{+}00]. Lasertechnologie [Rec85a]. Lasertechnologie/Werner [Rec85a]. last [Hei53e]. Later [Gla00, Hei53b]. Latour [BSMB99]. Lattice [CK95, Dob77, Man91, Mor72c]. Laudatio [Hei63f]. Laudationes [BDR86]. Laudations [BDR86]. Laue [Hei44b, Hei60f, Hei86j, Hei44b, Hei60f, Hei86j, Mor48]. Laureate [Hei65f]. laureates [ABH^{+}55, ABH^{+}87]. Lavoisier [Wag81]. Law [Har73, Ros79, Tau94, Ano34, Gau14, Hei42a, Hei58c, Hei66g, Cop81]. Lawren [Sei90a]. Lawrence [CS94, Car99a, Hen00, Log99, Neu99]. Laws [Sch91a, LM29, Pic08]. Layer [BFH^{+}42b, BFH^{+}42a, DDH41a, DDH41b, DDH42c, DDH89a, DDH89c, DDH89f, DDH89g, Hei41a, HBF^{+}42, Hei89i, HBF^{+}89a, HBF^{+}89b]. layers [HBF^{+}89c]. laying [Hei86m]. leaders [Ols63]. leading [Log96]. leap [Fis10a, Fis12a, Hei46c]. learning [Rin61]. Leben [ARK^{+}99, Büh95, Cas95, Eck12a, Eck12b, Hei80, Hei83a, Hei91a, Hen11, MGG^{+}95, RFS^{+}91, Rec03, Rec09, Rec10, Hei51d, Hei53h, Paw02b]. Lebendigen [BMG^{+}50]. Lebenslauf [Hei86h]. Lebensvorgängen [BMG^{+}50]. Lebensvorgängen/Eichler [BMG^{+}50]. Lebenswerk [PFH20]. Lebenswerkes [Hei49b]. Leccia [Mar73]. Lecourt [Jor78]. Lecture [Hei75s, Hei95, LSK^{+}76, HSD34a, Hei47f, Hei70j, vW41a, Hei31g, Hei89-38, Hei89-37, Hol86, And89]. Lecturees [Hei51k]. Lecturer [Rec84]. Lectures [Hei29c, Hei43p, Hei46a, Hei49j, Hei53l, Hei84-40, Nis32, SM41, Str62a, BDH70, Bor49, Hei52k, Hei52p, HMW^{+}84, SR92, Cal14, Edd07, Edd12, Hei43g, Hei43r, HHvB10, Wal92, Bor53]. Led [KH19]. Lee [Hei57f, Hei58b]. Lee-Model [Hei58b]. left [ES95]. legacies [Wal24b]. Legacy [CS94, Hea11, Smo92b, Smo92c]. legend [Wal24b]. Legenden [Wal93, Wal90b]. Legendenbildung [FS93]. legends [FS93, Wal93, Wal90b]. Lehmann [HEB^{+}80]. Lehrbuch [HEB^{+}80, SPS^{+}64]. Lehrer [Luc82]. Lehrstuhl [AH93]. Leibniz [LMOS15]. Leibnizschen [LMOS15]. leichten [Hei35c]. Leighton [BSP^{+}70]. Leinen [Hei53b]. Leipzig [Ano35a, Ben30, Ben34, Ben35, Clu37, Deh36, Deh37, DS43, EHR^{+}96, Hei31g, KW93, Mos42, Wil36a, Wil36b, Wit83, Ano29, GZ92b, GZ92a, Hei36e, RKW05, Tel93, vWL93, vM05a, Ano08]. Leipzig/Kofler [EHR^{+}96]. Leipziger [RKW05, vM05a, Hei36e]. Leitfaden [EHR^{+}96]. lekcii [BDH70]. Lenard [Sta36, Sta36]. length [DDH40a, DDH40b, DDH89e, DDH89d, Hei38h, Kra95]. Lense [BMG^{+}50]. Leo [Wig96i]. Leon [Hei31c, Hei31c, CS79]. Leopoldina [HH14, Wit83]. Leopoldina-Symposium [HH14]. Lepton [Hei67p]. Leptonenmassen [Hei67p]. Lesart [Kra93]. Leslie [Sei90a, Erm89]. let [Kra87]. Lett [RDM^{+}12a]. Letter [AO44, Ben13, Bet93, DR92, Hei60d, Hei66e, Hei66f, Pow02, Sch22, Wei93, Wig96j, Dem15, Gou47a, Hei56a, Hei79c, HGW^{+}89a, Hei89-30, Pow03, Sch15, Yam02a, Hei48h]. letterature [Cas14b]. Letters [Bro72, Cas05, Cho02, Kan03, KRW05, LRD^{+}91, O’F92, Pei92a, Wel92, BB71a, BB71b, BB05, Hei05b, Hof23, Som36, HHH03]. lettura [Arg09]. letzten [Hei53e, Hei86f]. Leuna [Har43, Har43]. leurs [Hei31c]. Level [Bil03, Lic16, LH24, Mil12, RS11b]. Lew [Erm89]. Lexikon [RFS^{+}91, vRFS^{+}00]. Li [HHHH11, HH17]. Library [Ano54, Hei53b]. Licht [SPS^{+}64, vRFS^{+}00]. Lichte [Hei41d]. Lichtenberg [Wag81]. Lie [DDI03, Scu75]. Liebe [Kan03, HHHH11, Cas05, HHH03]. Lies [Bed03, Bed04]. liess [Kra87]. Life [Ano94, Bac52, CS94, Cas91, Cas95, Cas02c, Cas02d, CD94, Dre92, Eck12a, For22, Fra94, Gar91, Kra86a, Lus10, MGG^{+}95, Mot85, Phi92, Rec10, Roh85, Swe93, WMG^{+}92, Wei85, BDH^{+}89, BR64, Cas92c, Gre05, Hei49b, Hei51d, Hei53h, Hei80, Hei83a, Hei84a, Hei91a, Mat86a, Mil97, Paw02b, RRK^{+}64, Roz67, BDH70]. Lifework [PFH20]. Light [Ble90, BSP^{+}70, Cas02b, CR12, Hei50i, Hei53p, Hei84-46, Wig96j, Hei25b, Hei35c, Hei41d, Hei52o, Hei84-61, KW05]. Like [Ano29, Hei43a, Hei23a]. Lillestol [vRFS^{+}00]. Limit [JTD^{+}98, ML71, Ano11, Isa82, KKM^{+}09, NKD^{+}11, Sch11a, SAR^{+}09, vV78]. limitation [Oza03]. Limitations [Ano34, vBW88]. Limited [LBS^{+}04, BM11, HBB^{+}07, Bad95]. Limits [Cam07b, Hei38f, Hei84-48, Sch11a]. Lindau [Hei93c, Hei75m, Hei93c]. Lindauer [Hei75m]. Lindley [Fer07, For09]. Lindner [MGG^{+}95]. line [Hei22c, SH22b, SH22a]. Linear [CRR93, Hei55f, Hei57f, Hei84q, Hei84r, Hei84-36, Hei84-34, KI64, SPS^{+}64, Unr72, vV78, GbK77, Hei58w, Hei62g, Hei68g, Hei84-29, KSSA78, Nar93]. Linear-algebraic [CRR93]. linearized [BN08]. Linen [Hei53b]. lines [Hei39a]. Linial [LN06]. Linienschärfe [SH22b]. Linienstruktur [Hei22c]. Liouville [Tem91, Van41]. Liouville-space [Tem91]. Lise [Büh98b, Cap57a, Cap57b, Hei69f, Rec91a, Rec91b, Büh98b, Cap57b, Hei69f, Rec91a, Rec91b, Sim89]. Liste [URWB95]. Liste/SSC [URWB95]. Liste/SSC-Physiker [URWB95]. listening [Hol93b]. Literary [Bed09]. Literatur [Str62b, Str62c]. Literature [DW52, Tut05, Cas14b]. Little [Klo93]. liv [BR64, RRK^{+}64]. Liverpool [Row09]. Lives [Bro97, Cas97, Sei90a, Bre97]. Lloyd [Sei90a]. Local [HK00, Kor05, OD03, ZG96]. Localization [JTD^{+}98]. localized [GbK77, KSGB^{+}83]. localized-site [GbK77, KSGB^{+}83]. Lochak [RFS^{+}91]. Löcher [Bes04]. Lochs [RDP^{+}95]. Lochs/Neuer [RDP^{+}95]. Logics [Hei59g]. Logik [Bor58]. London [Bro72, Neu99, Par06, Gou90]. lonen [Hei23a]. Look [Hof22, AMS04, Cor79]. Loopholes [Sch94, Tau94]. loot [Nag16]. Lorentz [JU82]. Lost [Gou46a]. Lösung [MGG^{+}95]. Lösung/Meÿenn [MGG^{+}95]. Lösungen [Hei22b, Hei84w]. Losungsmethoden [Hei50a]. louder [Dje07]. Louis [DC48, Loc07]. love [CG14]. Low [Her69, KGNS10]. Low-temperature [KGNS10]. Ltd [Log00]. Luchner [EHR^{+}96]. Luck [HVTK^{+}97]. Lügen [Bed03, Bed04]. Lukens [Cru23]. Lumière [Hei84-61, Hei53p]. lunaires [VB15]. lunar [VB15]. Lysenko [Jor78].

M [Ben34, CS94, Cru23, Deh37, Hei30b, Hei30c, Hei31g, Hei58l, Kan03, Sei90a, Wig96j, Wil36b, Wil36a, Wil36b]. M. [Hei58l, Hof02, Mos42, Pau34]. Mac [BMG^{+}50]. machine [Hei41g, Hei72a]. machines [Mil97]. Macmillan [Bro72]. MacPherson [Sei90a]. Macroscopic [BKBS93]. Made [Lau58a, Haw11, HvL47, Oor09, Sei90a]. Magazin [Jac11b]. magazine [Hei67j]. Magic [vM02]. magical [AMS04]. Magnet [Mor71, Mor72b, KJ06, Mor72c]. Magneten [DRF90a, DRF90b]. Magnetic [EHR^{+}96, WMG^{+}92, Hei25a]. magnetische [Hei25a]. Magnetisierungskurve [Hei31l]. Magnetism [Fis64, MC97, Sno49]. Magnetismus [LSK^{+}76]. Magnetization [Fal64, Hei31l, HN99]. Magneto [BSP^{+}70, Sno49]. Magneto-Electric [BSP^{+}70]. Magneto-Resistance [Sno49]. Magnetostriction [Sno49, Hei31l]. Magnetostriktion [Hei31l]. magnon [SZ73, SZ74]. Magnons [HWY69, HY72, Dob77]. Magnus [RDP^{+}95]. Magnus-Haus [RDP^{+}95, URWB95]. Mai [Hei30a, Hei64b, Hei89l, Hei58p]. Main [Ano34, Geĭ58]. Mainzer [RFS^{+}91]. Major [AO44]. Majorana [GHH02, Bri65b, GHH02, GR08]. maker [Log93]. Makers [AG02]. Making [Rho86, Rov21, Sei90a, Sta11a, Bel99, Zan02b]. Maksa [Geĭ70a]. Malcolm [Sei90a]. Man [Bro73a, Fac82, Cop81, Brü64, Büh98a, Hei89-40, Hei91a, Kem85, Kra87, Hei50b, Sim79]. mandarins [Rin69, Rin90]. Mandl [KED^{+}60]. Manhattan [Sei90a, BB96, Jon85, Sei90a]. Manifold [Ni04a, Ni04b]. manifolds [Wan98]. Mann [Kra87]. Männer [Sch05]. Mantegna [vRFS^{+}00]. manuscrit [Bar04]. Many [Hei89k, Hei50a, Kra84, SME77, vV78, Hei26a]. Many-Body [Hei89k, vV78, Hei26a]. many-particle [Hei50a]. Maple [EHR^{+}96, vRFS^{+}00]. Mar [Hei89i]. Maran [WMG^{+}92]. March [Cap57a, Cap57b, Hei89i, Cap57b, Kra95, vW41a]. March/Der [Cap57a, Cap57b]. March/The [Cap57b]. Märchen [HVTK^{+}97]. Marelene [CS94]. Maria [Cas05]. Marion [LSK^{+}76]. Mark [Hay90, Klo96, Hei05b]. Marks [Sch00b]. Marley [KED^{+}60]. Martha [Kra86a, Mat86b, Rec85b]. Martin [ARK^{+}99, Sei90a]. más [W^{+}87, WdC02]. Mass [Gam30, Log96, Par06, Tur06, Ber02, El 00, McC24, vW35, vW41b]. Masse [vW41b]. massereichen [RDP^{+}95]. Masses [Hei67p]. massive [Fri97, Fri98]. massless [KKM^{+}09]. matematica [Arg09]. matematizzazione [Fer99]. material [HvL47, HvL47]. Materialforschung [URWB95]. Materialforschung/Wer [URWB95]. Materie [Hei34e, Hei53o, Hei67h, Hei67i, MGG^{+}95, Hei54c]. Materie/Brandmüller [MGG^{+}95]. Materiewellen [Deh36, Hei36b, Ano35b, Haa34]. materja [Inf33]. Math [DPL96, Fri98, Sch02a]. Mathcad [EHR^{+}96]. Mathematica [Bes98]. Mathematical [ARK^{+}99, Hei32a, SPS^{+}64, Hei46b, Hei51h, Hei65p]. Mathematicians [Acz01]. Mathematics [Ber03, BSP^{+}70, Arg09, FF91]. Mathematik [BMG^{+}50, HMSH78]. Mathematik/Dreier [HMSH78]. Mathematik/Mac [BMG^{+}50]. mathematische [Hei46b]. mathematischen [Hei65p, SPS^{+}64]. mathematization [Fer99]. matrice [DC48]. Matrices [KBWJ96, Per88, TR09]. Matrix [Bor72, BHJ62, Bro02a, Bro02b, Cus86, Hei89k, KED^{+}60, Mac77, Mas14, Sta09b, BJLR17, Bor70, DC48, GKZL96, JLP^{+}02, KSSA78, Kle78, KLV80, NRS71, Ras78, Hei89k]. Matrizen [BMG^{+}50]. Matrizen/Heisenberg [BMG^{+}50]. Matrizenmechanik [BHJ62]. Matter [Ano43, ARK^{+}99, Bar07, Gou90, Hei34e, Hei62h, Hei96d, Wit12, Žup65, Car76, Hei36b, Hei53o, Hei54c, Hei66g, Hei67h, Hei67i, Inf33, Inf34, MF69, WH07, Ano35b, Haa34, Han62, Ros79]. Matter/Friedrich [ARK^{+}99]. Matthias [Kra07, Lan08, Lau07]. Maurice [KW87]. Max [Adl03, Ano84b, Bro72, Fis10a, Fis12a, Hei72a, Hei27a, Hei31g, HPE^{+}33, Hei35b, Hei43i, Hei43p, Hei43q, Hei49b, Hei53l, HMW^{+}84, Mar73, Rec81b, Sch05, Sch07, Tre01, vBL96b, BB71a, BB71b, BB05, Fis10a, Fis12a, Geĭ70a, Gre05, Hei27a, HPE^{+}33, Hei35b, Hei38g, Hei43i, Hei43p, Hei44b, Hei47d, Hei47e, Hei49b, Hei53l, Hei60f, Hei62i, Hei62j, Hei70g, Hei71f, Hei71h, Hei86j, Hei86i, Sch05, Sch07]. Max-Auwärter-Preis [Ano84b].

Max-Planck- [Rec81b]. Max-Planck-Gesellschaft [vBL96a, vBL96b]. Max-Planck-Institut [Hei53l, Hei53l, Hei62j]. Maxwell [Bes04, Hei72a, Bes04]. May [Hei30a, Hei64b, Hei89l]. Mayenn [Kra96]. McKay [Sei90a]. Meaning [Bar07, DT18, Hei79a, Ome79, Smo93, Hei71d, Mes04]. Means [Hol84, Hei52m]. Measure [DGS99, GV86]. Measure-valued [DGS99]. Measurement [Cle09, KED^{+}60, RDM^{+}12a, RDM^{+}12b, WZ83, D’A03, SAR^{+}09, Tel61, VL97, Vya99, Wis98]. Measurement-Disturbance [RDM^{+}12a, RDM^{+}12b, Wis98]. Measurements [Cop81, RDM^{+}12a, RDM^{+}12b, BM11, JD06, HBF^{+}42, HBF^{+}89a]. measures [Irv67c, NBH97]. measuring [Oza02]. Mécanique [BH84, BD92, BH28, dB82]. Meccanica [Hei36a]. Mechanic [Gar91]. Mechanical [Geĭ77b, Hei77b, Kae31, Ser77, BPS74, DRB86, Geĭ77a, Hei25c, HBE87, Jes98, Tel61]. Mechanics [Bok10, BHJ62, BHJ67, BH75, BW00a, BW00b, Bru80, CK02b, EHR^{+}96, Fra10, GGK02, GL23, HSD34b, Hei79b, Hen98, Jam74b, Lev95, Los61, Mac77, Mas14, Nis32, OW10, RFS^{+}91, Sta11a, Sta11b, Sta09b, Wey31b, Wey50, dB90, vdW67, Ano35b, Ano20, BD92, Blo76a, Blo76b, Blo85, BJLR17, Bok04, BJ25, BHJ26, BH28, Bor61, Bor63b, BHJ63, BJ67, BH84, BGĬ90, Cam09, CK02a, Cus94, Enz83, ET07, Fri11, Fue61, Gal01, GKZL96, Haa34, HR02a, Hee95, HJ26, Hei26a, Hei26b, Hei26c, Hei26f, Hei26d, Hei27e, Hei27f, Hei27c, Hei27d, Hei30h, Hei30g, Hei30d, Hei31e, Hei32i, Hei32j, Hei33a, Hei34d, Hei58f, Hei60b, Hei65c, Hei77e, Hei77j, Hei83b, Hei83c, Hei84-51, HL84, Hei84-42, Hei85d, Hei93e, Hei10b, Jac01, Kle78, KLV80, LM87]. mechanics [Low97, Nar93, OS07, PC77, Raz11, Sch26, Sch27, Sch66, Sch96b, dB82, vdW68, Hei30b, Mil90, Eng09, Hei28a, Hei28b, Hei30c, Hei31b, Hei36a, Hei43g, Mil12, RS11b, Hei36b, Hei50a]. Mechanics/Neuser [EHR^{+}96]. Mechanics/Spatschek [RFS^{+}91]. Mechanik [GRE^{+}01, vRFS^{+}00, Hei26f, Hei27e, Hei27d, Hei84-42, Hei43g]. mechanischer [Hei25c]. Medal [Ano29]. medarbejdere [BR64, RRK^{+}64]. Media [BSP^{+}70, JHB99]. Media/Deschamps [BSP^{+}70]. Medium [Gro96]. Mediziner [HEB^{+}80]. Mediziner/Hermann [HEB^{+}80]. Medvedev [Jor78]. Meet [PS69]. Meeting [Aas02a, Aas02d, Aas02c, Bar07, CCJ^{+}34, Hei73a, Kra07, Lan08, Lau07, Wig96a, Aas02b, Ano02a, Dör05, Gla00, Hei30a, Hei41b, Hei69b, Kar05a, Paw02a, Ano66]. meetings [Har43, Hei84-55, Hei96c, Hei00a, Hei08c, Hei75m]. Meets [Bok06, Ore98]. Megabit [Ano84a]. mehanicheskih [Geĭ77a]. mehaniki [Geĭ77b]. Mehlin [KED^{+}60]. Mehr [Hei49g]. Mehrfachlinien [SH22a]. Mehrkörperproblem [Hei26a]. Mehrkörperprobleme [Hei27c]. Mehrkörperproblemen [Hei89k]. Mehrteilchen [Hei50a]. Meine [HHHH11, Dür05]. meinem [Sch26]. meinen [Hei53b]. Meitner [Cap57a, Cap57b, DRF90a, DRF90b, MGG^{+}95, Rec91a, Rec91b, BM84, Büh98b, Cap57b, Hei69f, Rec91a, Rec91b, Sim89]. Meixner [Koo89]. members [HR^{+}89]. Membership [Cla23]. Memo [Hei53a]. Memoir [Wig96e]. Memoirs [Sei90a, Erm89, Hei67f]. Memorandum [BBH^{+}62, Hei49g, Ano83a, GHW36, HR53, HGW89b]. Memorial [Hei63e, Hei75q, Wig96a]. memoriam [CDvW77, Her76a, Mey76, Sch82, Wit83]. Memories [Dür05, Hei80, Hei83a, Hei91a, Hei60b, Hei64c, Tel93, Wig96g]. Memory [Rec81a, Sta11a, Geĭ70a, Hei43u, Her76a, Sch82]. Men [Lau58a, Sch05]. Mensch [Hei50b, Sim79]. Menschen [LSK^{+}76]. Mental [Wal08, Hen07]. Mentality [Wal08, Hen07]. Mercer [FW07]. Merely [Ber01]. Merit [Cap57b]. Merite [Cap57a, Cap57b]. Mermin [Wal08]. Meschede [vRFS^{+}00]. Meson [DH62, Hei49i, Hei51l, Hei52n, Hei52g]. Mesonen [Hei49k, Hei49d, Hei51l, Hei84p, Hei43d, Hei52b]. Mesonenerzeugung [Hei52g]. Mesons [Hei43d, Hei52h, Hei84-28, Hei49k, Hei49d, Hei52b, Hei55e, Hei56l, Hei83g, Hei84p]. Mesotron [Hei38a, Hei84i]. Message [GLSC01]. Meßprozess [Tel61]. Messungen [HBF^{+}42, HBF^{+}89a]. mesure [d’E92]. metafora [Cas14b]. Metal [BFH^{+}42b, DDH42b, DDH89b, Hei42b, Hei89g, BFH^{+}42a, Döp42, DDH42a, DDH42c, DDH89a, DDH89c, HBF^{+}42, HBF^{+}89a, HBF^{+}89b, HBF^{+}89c, OD03]. Metall [BFH^{+}42b, DDH42b, DDH89a, DDH89b, HBF^{+}42, Hei89g, HBF^{+}89a, HBF^{+}89c, RDP^{+}95, BFH^{+}42a, HBF^{+}89b, DDH42a, Hei42b]. Metall-Keramik-Multilagen [RDP^{+}95]. Metall-Keramik-Multilagen/Neues [RDP^{+}95]. Metallphysik [SPS^{+}64]. Metallschichtensystem [DDH89c, DDH42c]. metals [HBE87]. metaphor [Cas14b]. Metaphysics [Los61, Hei91c]. metaplectic [GBLW86]. Metastable [JTD^{+}98]. Method [Har73, Hei58b, Hei56e, Hei72b, Hei84e, MC92, OD03, CS94]. Methode [Hei84e]. Methode“ [Hei56e, Hei72b]. Methodik [BSP^{+}70]. Methodik/Happer [BSP^{+}70]. Methodische [BMG^{+}50]. Methodological [Los61, Mon04]. Methodology [MWL^{+}11]. Methods [Hei75b, Hei50a]. Methuen [Log00]. Metric [Hei84u, Hei84-36, Hei58w, Hei72h]. metrics [LN06]. metrology [Ano11, NKD^{+}11]. Meyenn [Bro95, Eis00, GRE^{+}01, HEB^{+}80, Pye81, Rec86, ARK^{+}99, MGG^{+}95, Car97]. Meyer [Dje07]. Michael [Kra07, Log00, Dör05, Lan08, Lau07]. micromechanical [SAR^{+}09]. Microscope [Cor79, KPS^{+}95, Roy78, CL94, Foo94, Tan04, VL97, Vya99, Wil84]. Microscopic [BHX96]. Microscopy [WMG^{+}92]. Mie [HW12a]. Mikroskop [DRF90a, DRF90b, URWB95]. Military [MS15]. Miller [Sei90a]. Mind [Hür22, Han62]. minds [Pic08]. mine [Sch26, Sch66]. Minimally [JD06]. Minimum [HK00]. Minister [GHW36, HGW^{+}89a]. Minutes [Hei89-38]. Miramare [Meh73]. Misconceive [RW92]. Misinterpreted [Pop16c]. Misleading [Roy78]. misrepresentation [Tru72]. Missing [Joh13]. mission [Gou47a, Gou48b]. mistaken [Büh98a, Log96]. Mistakes [Bed03, Bed04]. místicos [W^{+}87, WdC02]. Mitarbeit [KPMN20]. Mitarbeiter [Ano84a]. Mitglied [DRF90a, DRF90b]. Mitglied/Neue [DRF90a, DRF90b]. Mitglieder [HR^{+}89]. Mittel [Hei46e]. Mittelstaedt [And89]. Mitwirkung [Hei65d]. Mladjenovic [ARK^{+}99]. Mobilization [Joh13]. Model [Bra97, Cas79b, CJ72a, CJ72b, Fis64, Gre63, GR08, Hei57f, Hei58b, Hen98, KBWJ96, KI64, ML71, Alb73, BH23b, BR96a, BN08, CA76, DMdA06, DRB86, DLS76, Ech96, GbK77, GWW^{+}09, HWY69, HY72, HN99, Isa82, IM95, KW02, KW03, Mor72a, SL70, Yaf00, Yeh70, ZG96]. modèle [Ech96]. Modelers [Lak96]. Modeling [ARK^{+}99, KGNS10]. Modeling/Origin [ARK^{+}99]. Modelle [BMG^{+}50]. Modellen [BH23b]. Models [BHX96, Joh13, Lak96, Mac77, CL05, Car99b, KF02, KS92, MP92, NBH97, SF71, Wre75, SK73]. Modern [And89, BHN98, Bor58, BW04, Fli32, Han62, HSD34b, HBSA61, Hei89v, Hei96b, Hei04, Hei08d, Hol84, Kle17, Mac58, Rud58, SPS^{+}64, vRFS^{+}00, Ano66, Bra09, Hei31h, Hei31d, Hei32d, Hei36f, Hei37e, Hei41d, Hei43k, Hei48e, Hei48d, Hei52e, Hei52l, Hei52o, Hei53o, Hei58s, Hei59c, HSBA60, Hei60i, Hei61c, Hei61f, Hei62l, Hei63a, Hei65a, Hei70d, Hei72i, Hei74a, Hei74i, Hei74j, Hei84-54, Hei00b, Hei07b, Inf34, HSD34a]. moderna [Hei99, HSBA60, Pol00]. moderne [Ben34, HSD34a, Hei63a, MGG^{+}95, RDP^{+}95, Wil36b]. modernen [Hei31h, Hei31d, Hei32d, Hei36f, Hei37e, Hei41d, Hei43k, Hei48e, Hei48d, Hei53o, Hei60i, Hei61c, Hei84-54, Kle17]. moderner [Hei70d]. modernes [Hei65a]. Modernizing [Joh13]. Modes [HVTK^{+}97]. Modes/Gu [HVTK^{+}97]. modification [Hei24d]. modified [DC87]. modules [LJ08, LWZ10]. Moelwyn-Hughes [BSP^{+}70]. möglich [Hei79b, Hei85d]. Möglichkeit [Hei39e, Hei40a, Hei41f, Hei69k, Hei69l, Hei89h, Hei89-36, Hei89m, Hei89x, vW40]. Möglichkeiten [Hei53g, Hei56j, Hei89w]. Mohsen [RS11b]. Molecular [Ano84b, Joh13]. molecule [Tel30]. Molecules [Hei84c, BH23b, BH24b, Hei24c, Hei32b, Hei33c, Hei84-41, NAZ02]. Molekeln [Hei24c, BH23b, BH24b]. MolekUlen [Hei33c, Hei84-41]. Moloney [HVTK^{+}97]. moment [CG14]. momenta [KF07]. Moments [KBWJ96, Büh98b]. Momentum [Gre62]. mondo [Arg09]. mondo-della-vita [Arg09]. monograph [Lic16]. Monotonicity [Pea76]. Monte [HN99]. Morgan [KED^{+}60]. Morita [Pac88]. Morris [Kra86a, Mat86a, Wei85]. Moseley [Hei72a]. Moskauer [DRF90a, DRF90b]. most [Haw11, Hei40b, Sch13, W^{+}87, WdC02]. motion [Hei22a, Isa82, LL82, Lun84, SHI73]. Motz [Sei90a]. Movability [vBW88]. Movement [Sei90a]. Moyer [CS94]. MR [DPL96, Mil12, RS11a]. Mr. [Fri11, Hei31f, Kae47]. Müller [Cas05]. Multidimensional [EHR^{+}96]. Multilagen/Neues [RDP^{+}95]. Multiparticle [LBS^{+}04, OS07]. multiphoton [IM95]. Multiple [Hei52h, KW76, Hei49k, Hei49d, Hei52b, Hei84p]. multiplet [Hei25d]. multiplets [LH24, SH22a]. Multipletts [LH24]. Multiplettstruktur [Hei25d]. Multiplication [DDH42b, DDH89b, DDH42a, DDH42c, DDH89a, DDH89c]. München [Bun71, Hei62j, Her76a, SM41, Hei58k]. Münchener [BDR86]. mundo [W^{+}87, WdC02]. Munich [BDR86, Cas05, Hei58k, Hei74g, Her76a, SM41, TEW02]. Museum [Ano84b, DRF90a, DRF90b, Hei64b, Hei86o, Fec90, Hei86o, Rec91a, Rec91b, Sim10]. Museum/Heisenberg [Ano84b]. Museum/Heisenberg-Programm [Ano84b]. Museums [Hei64b, Rec91a, Rec91b]. Music [Sim79]. Musik [Sim79]. My [Hei72a, HHHH11, HH17, Dür05, Hei53b]. Mysterious [KH19, Kar05a]. Mystery [Acz01, FR13b, Gla00, Ano20, Bes04, FR98]. mystical [W^{+}87, WdC02]. Myth [Klo96, Wal95].

N [Cru23, Dje07, Hei31f, Wal08]. nabrosku [Hei62n]. nach [Ano84a, GHH02, HVTK^{+}97, Hei25b, Hei53b, Hei56g, KH47, URWB95, vRFS^{+}00, vW41a]. Nachlass [Ano76]. Nachruf [HKvL47, Hei47e, Hei51c, Hei60f, Hei71h]. Nacht [vS14]. Nachtrag [Fis02, Hei41e, Hei89-34, Hei89x]. Nachweis [DDH42a, DDH89a]. Nagging [Paw02a]. nagljadnom [Geĭ77b]. naissance [BD92]. Name [RMRP09, Hei67e, Hei67e]. Named [Ano33]. namely [D’A03]. Namensverleihung [Hei89b]. naming [Hei89b]. Nano [RDP^{+}95]. Nano-Kugellager [RDP^{+}95]. Nano-Kugellager/Bislang [RDP^{+}95]. Nanocrystals [ARK^{+}99]. Nanocrystals/Gershenfeld [ARK^{+}99]. Nanolithography [JTD^{+}98]. nanoscale [Cle09]. narrative [Com56]. Nasledie [Smo92c]. Nation [Ano33]. National [Hay90, Wal89a, Har37a, Har37b, Har81, Hen96, HH11, HW04, RW94, Wal89b, Wal92]. Nationalsozialismus [Mai07]. Natur [EHR^{+}96, Hei69e, WMG^{+}92, Hei79e]. Natur/Cassidy [WMG^{+}92]. Natural [Hei66g, Sil13, Som36, Str62a, Hei35f, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei49l, Hei52e, Hei55b, Hei56f, Hei57c, Hei58e, Hei59a, Hei59i, Hei60a, Hei62b, Hei63b, HG65, Hei66a, Hei67b, Hei68c, Hei70c, Hei71d, HnH72, Hei73j, Hei74d, Hei76e, Hei79d, Sta36, Hei72i, Ros79]. Naturbeschreibung [HVTK^{+}97]. Naturbeschreibung/Demtröder [HVTK^{+}97]. Naturbild [Hei53c, Hei55b, Hei56f, Hei57c, Hei58e, Hei59a, Hei60a, Hei62b, Hei63b, HG65, Hei66a, Hei67b, Hei68c, Hei70c, HnH72, Hei74d, Hei76e, Hei79d]. Nature [ARK^{+}99, Cal14, Edd28, Edd78, Edd07, Edd12, Edd13, Gin75, Hei58v, Hei60g, Hei65l, Hei76a, Hei77a, Pag82, Plo10, RDP^{+}95, Edd00, Geĭ77c, Hei33d, Hei37e, Hei52j, Hei53c, Hei58o, Hei58r, Hei63g, Hei69e, Hei70i, Hei74h, Hei76f, Hei79e, Hei84-58, Meh73, Han62, Joh59, Sch59]. Naturerklärung [Hei33d]. Naturforscher [Hei72a, Wit83, Sta36]. Naturphilosophie [Hei37e]. Naturwissenschaft [AH93, Ben35, Clu37, DS43, Hei34i, Hei34f, Hei34g, Hei34h, Hei82c, Hei84-63, KED^{+}60, LSK^{+}76, Mos42, Str62a, Str62c, AH93, Hei35f, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei48e, Hei48d, Hei48f, Hei49h, Hei49l, Hei59i, Hei61c, Hei61e, Hei65p, Hei71c, Hei72j, Hei72i, Hei73j, Smo93, Tre01]. Naturwissenschaft/Heisenberg [KED^{+}60]. Naturwissenschaft/Lecture [LSK^{+}76]. Naturwissenschaften [Ano35a, Deh37, Som36, Wil36a, Hei71d]. Naturwissenschaftler [vRFS^{+}00]. Naturwissenschaftler/Weizsäcker [vRFS^{+}00]. naturwissenschaftliche [BDR85b, Hei67b, Hei73f]. naturwissenschaftlichen [Hei42d]. nauki [Inf33]. Nazi [Car99a, Hen00, Lan02, Lib68, Log99, Neu99, Wal99, FR08, Bro93, Bro92, Cho02, DR92, FR08, Gla00, GP92b, Irv67b, Joh13, Mac85, Mac93, Mai07, MP01, O’F92, Ros98, Wal95, WHF^{+}96, Zan02a, Klo96]. Nazi-related [Mai07]. Nazis [Bro09, Cha92, Hei48i, Kae48a]. Nazism [Ols63]. názory [Hei69i]. NBC [Hei89-31]. Nearest [Fal64]. Nearest-Neighbor [Fal64]. nebula [HvW48]. need [Hei48l, Hei68b]. needs [Pop16c]. Negative [Sch13]. Neighbor [Fal64, Dob77]. nel [Arg09]. Nelinejnye [Geĭ68]. nelle [Cas14b]. nemeckogo [DF96]. Neon [Hei27b]. Neue [DRF90a, DRF90b, URWB95, Wit83, Rec02b, Ben30, Ben34, Ben35, Clu37]. Neuen [Hei56e, Hei84e, Ano02a, BFH^{+}42a, Hei61a, Hei72b, HBF^{+}89b, Hei89-40, SPS^{+}64]. Neuer [Ano84a, RDP^{+}95, Wal93]. Neuere [Hou30]. Neues [RDP^{+}95]. Neufassung [Dür80]. Neujahrsansprache [Hei89y]. Neumann [Wig96h, Bac22, GW96, Wig96c, WG96]. Neuser [EHR^{+}96]. Neutrinienne [Hei84-61, Hei53p]. Neutrino [GLR14, Hei53p]. Neutrinooszillationen [URWB95]. Neutrinoquelle [URWB95]. Neutrinoquelle/Atomare [URWB95]. Neutron [Bro71, BM84, GR09, Rog13, DK78, DDH42a, DDH42c, Hei37c, DDH42b, DDH89a, DDH89c, DDH89b]. Neutron-Induced [GR09]. Neutronen [DDH40a, DDH40b, DDH89e, DDH89d, DDH89b, Hei52p, Rec85a, DDH42b, Fis41]. Neutronenstrahlung [Hei37c]. Neutronenvermehrung [DDH42a, DDH42c, DDH42b, DDH89a, DDH89c, DDH89b]. Neutronics [GORS09]. Neutrons [DDH42b, DDH89b, DDH40a, DDH40b, DDH89e, DDH89d, Fis41, Hei52p]. Never [Pei93b, Hei48g]. Nevill [Mac78b]. Newcomb [CS94]. News [Stu85, Yam02a, DRF90b, Hei89-31]. Newton [Bes04, Mil12, Büh98b, Fri11, Hei41d, Hei52o]. Newtonian [BKBS93]. Newtons [Büh98b]. Newton’sche [Hei41d]. Nicholas [Hei86n]. Nichols [Sei90a]. nicht [Hei67a, Hei67g, Pop16a]. Nichtlaminare [Hei22b, Hei84w]. Nichtlineare [MGG^{+}95]. nichtlinearen [Hei61b, Hei63d, Hei84k, Hei84r]. nichtlinearer [AH57, Hei53y, HKM55]. nichtrenormierbarer [Hei54f]. Nicolaus [Gin75]. Niels [Adl03, Bor58, Hei86g, Hei10a, Kra07, Lan08, Lau07, Aas02b, Aas02e, BR64, Dör05, Fav05, FK85, Hei35d, Hei63h, Hei10a, Meh87, Pai91, PRW55, RRK^{+}64, Roz67, Sch11c, Yam02a, vD15]. niemals [Hei48g]. niemieckich [Kar06]. Nietzsche [Sei90b]. night [vS14]. Nikolaus [Hei86n]. NLB [Jor78]. NMR [EHR^{+}96]. No [Sch94, Yam02a, Zan02a, DPL96, Hei67g, Hol93b]. Nobel [HSD34a, Rec10, ABH^{+}55, ABH^{+}87, Ano33, Ano20, Gre05, Hei32c, HSD34b, Hei65f, LW34, Läs13]. Nobelpreis [Ano51, Hen11, Rec09, Rec10]. Nobelpreises [HSD34a, Wil36b]. Nobelpreisträger [ABH^{+}55, ABH^{+}87, Hei65f, Läs13]. noble [BDR84j, Hei23a]. noble-gas-like [Hei23a]. noch [Hei51e, NW95, URWB95]. Non [Boh62, Hei55f, Hei57f, Hei68g, Hei84q, Hei84r, Hei84-36, Hei84-34, MWL^{+}11, Hei58w, Hei62g, Hei84-29, Hei84w]. Non-Classical [Boh62]. Non-laminar [Hei84w]. Non-Linear [Hei55f, Hei84q, Hei84r, Hei84-36, Hei84-34, Hei68g, Hei58w, Hei62g, Hei84-29]. Non-Proliferation [MWL^{+}11]. Nonintegrability [DKLN92]. nonintuitive [Hei46c]. nonisospectral [LTG^{+}12]. Nonlaminar [Hei22b]. Nonlinear [BDR84i, DH66, DH84a, Hei60h, Hei63d, Hei65k, Hei67l, Hei84-30, Hei84-32, Wan01, AH57, Bes98, Car71, DHYY65, Hei53y, HKM55, Hei61b, Hei84k, KJ06, Mie80, Mie81, Geĭ68, Hei75c]. Nonlocality [Kas11, OW10, Sei05]. nonrelativistic [BP75]. nonrenormalizable [Hei54f]. nonseparable [KtL79]. Nonviolence [Hei69g]. North [Adl03, BCEP94]. Northwestern [Gou47a]. Norton [Jor78]. Norwegian [Aas02e, Bør12]. Note [Hei48m, May87, Ree19, RDM^{+}12a, Van41, Hei35g, SH22b, Hei31f]. Notes [LSK^{+}76, RS11a, Geĭ36]. notion [Hei48e, Hei48d, Hei74i]. Notiz [Hei31f]. Notwendigkeit [Hei48l]. Nov [Hei89-36, Hei89x]. Nova [Wit83]. November [Hei31g, Hei89-36, Hei89x, Wal20b]. Nowe [Inf33]. Noyau [Hei84b, Pau34, Hei34c]. noyaux [CCJ^{+}34]. Nr [Wit83]. nucléaire [Kar07, Lur95]. Nuclear [ARK^{+}99, BB36, BBF^{+}57a, Bor53, Bør12, Bri65a, BM86, BR96b, BSP^{+}70, CS94, Cas02b, Dah99, Eck90, For22, Gam30, Gou47c, GORS09, GWY05a, HB82b, HVTK^{+}97, Hay90, Hei50i, Hei53m, Hei53n, Hei54b, Hei54d, Hei66j, Hei68e, Hei69h, Hei76b, Hei81a, Hei84-46, Hei86a, Kar05c, KH19, Lib68, MGG^{+}95, MWL^{+}11, MWL^{+}15, MS15, Mil85, Mla98, NBDG40, Rad83, Rad91, Sei90a, She83a, She83b, Sim00, Sim10, Sim12, TV23, Wat01, WMG^{+}92, Wig96m, Ada72, ABH^{+}55, ABH^{+}87, Bad95, Bet91, BDR86, Bro92, CHW91, Eck33, El 00, GWY05b, HB82a, Hau22, Hei32i, Hei32j, Hei43l, Hei52k, Hei53k, Hei53g, Hei56a, Hei56d, Hei56j, Hei57b, Hei68b, Hei89l, Hei89p, Hei89-32, Hof92, Hof23, Hou41, Irv67b, Jun64, Jun72, Jun90, Kae48b, Kar07, KP07, Kra87, Lur95, Maj33]. nuclear [Wal89a, Wal93, WC84, vW35, Ano54, Fel54, Hei89d, Hei89c, Wal05, Wal07, vW41b, Hei54a, Kle54, Sei90a]. Nuclear-Fission [Sim10, Sim12]. Nuclei [BB36, CR12, CCJ^{+}34, Hei35c, Hei38c, Hei38b, Hei65g, Hei65h]. Nucleon [GRE^{+}01]. Nucleon/Sube [GRE^{+}01]. nucleons [Hei56l]. Nucleus [FR13b, Hei75d, BDR84f, BDR84h, BR96a, Geĭ36, Hei32f, Hei32g, Hei33b, Hei34c, Hei35a, Hei36d, Hei37b, Hei37d, Hei39c, Hei43n, Hei47b, HW49, Hei51j, Hei51k, Hei84j, Hei84f, Hei84b, Hei84l, Hou30, Rec02a]. nuklearen [Rad83, Rad91]. Numbering [Hei72a]. Numbers [DH62, vW48]. Numerical [Bec02, Bra97, MC92]. Nuovo [Pol00]. Nur [RFS^{+}91]. Nutzung [HB82a].

O [BBF^{+}45, DDH41a, DDH42a, DDH89a, DDH89c, DDH89f, Hei43f, HBB^{+}89, Bro95, Hei96c, BBF^{+}45, Cav14, DDH41a, DDH42a, DDH42c, DDH89a, DDH89c, DDH89f, Geĭ77a, Geĭ77b, Hei42b, HBB^{+}89, Hei89g]. O-38-metal [DDH42c, DDH89c]. O-38-Metallschichtensystem [DDH89c]. Obituary [Tel76, Wig76, HKvL47, Hei47e, Hei51c, Hei60f, Hei71h]. Object [Hei76b, KH19, JU82]. Objectivity [Car03]. Objects [FW20, MSUU80]. Objekten [GRE^{+}01]. Observability [Lac02a, Lac02b, Wol14]. Observable [Hei43j, Hei44a, Hei43m, Hei62k]. Observations [Ber01, Yeh70]. Obsession [Mil10]. Obtain [Ano34]. Obtaining [CA76, Har73, LB72]. obtenir [Kar07]. occasion [Hei58k, Hei69b, Hei89b, Hei89a, PRW55, Sch82]. occupation [Ano44a]. Occupied [Sch11c]. occurring [Hei38h]. October [CCJ^{+}34, DDH89f, Far01, Hei35d, HBF^{+}89a, IEE06]. octobre [CCJ^{+}34]. O’Dell [BSP^{+}70]. off [D’A03]. öffentlichen [Hei51d, Hei53h]. Öffentlichkeit [DRF90b, DRF90a]. Office [Kra86a, Kra07, Lan08, Sei90a]. official [Hei89b]. offiziellen [Hei89b]. Oguchi [LD66]. ohne [Hei57b]. Okt [DDH89f, HBF^{+}89a]. Oktober [Hei35d]. old [Hei38g, Hei49a, Hei61a]. Olson [HEB^{+}80]. oltre [Hei84-55, Hei00a, Hei08c]. om [Ano44a]. Omnès [EHR^{+}96]. once [Ols63, Hei64c]. once-great [Ols63]. onde [dB58]. ondulatoire [dB82]. One [AK15, Fis64, Kae31, Sta09a, BT78, DMdA06, Fri97, Fri98, Hol93b, Isa82, PDL92, ZG96]. One-Dimensional [Fis64, BT78, DMdA06, Isa82, PDL92, ZG96]. ones [SK09]. Online [CF98]. ont [VB15]. op [vC16]. Open [Bok04, Büh98a, CL05, Hof92, Oor09]. Opening [Hei75s, Sch00b, SM41, Geĭ58]. Operacija [DF96]. Operation [FS93, Hof92, d’E92, DF96, Fra93b, Fra93c, Fra93a, Hof92, Hof93, Hof23, NW95, Fra93a]. Operator [Sch99, Bes98, Fri97, Fri98, OS07, Sny82]. Operators [Kor05, CZ81, Dys51a, Dys51b, KF07]. ÖPG [Ano84b]. Opinion [Gol92, Hei10b, Hei65m, Hei89-32]. Opinions [BDR89b]. Oppenheimer [Ano34, Hol84]. Opposition [Hei47a, Loc07, Sch11b]. Optical [Ano04a, ARK^{+}99, BSP^{+}70, CR12, JTD^{+}98, RFS^{+}91, BH24a, BH63b, Hei24c]. Optics [Aga02, KED^{+}60, PSZM02]. Optics/Mandl [KED^{+}60]. Optik [BMG^{+}50, DRF90a, DRF90b, vRFS^{+}00]. optische [BH24a, BH63b, Hei24c]. Optoelektronische [OSWR86]. Oracles [Sei90a]. orbits [BH63a]. Order [Cap01, EHR^{+}96, Gou90, HK19, Mer38, DRB86, GWW^{+}09, Hei84-52, LTG^{+}12, Ano25, BDR84c, Cap57b, Cru23, Hei89z]. Order/Kafka [EHR^{+}96]. Ordnung [BDR84c, Hei84-52, Hei89z]. Organisation [BDR89b, HR53, Hei54b]. Organization [BDR89b, Hei54b, HR53]. Origin [ARK^{+}99, BR96b, Hei77g, Hei49k, Hei75f, Rad83, Rad91]. Original [Ano34, BDR84a, BDR89a, BDR93, Bro86, Mat87, Paw06]. Originalarbeiten [BDR84a, Bro86, BDR89a, BDR93, Mat87]. Originaltonaufnahmen [Paw06]. Origins [Meh87]. Orthner [Hei39b]. orthogonal [DLW08, VZ09]. orthosymplectic [BDH88]. Oscillator [New35, BDH88, CGM^{+}71, SAR^{+}09]. oscillators [Car71, Tru75]. osnovnye [Geĭ58]. Ostmark [Hei41b]. Otázky [Hei69i]. Other [Hen00, Hei89o, Hof23, Hol86, MGG^{+}95, vW41b]. Otkrytie [Geĭ58]. Otto [GRE^{+}01, Hei69f, KPMN20, KF01, Sim07, Hei54e, Hei68e, Hei68f, Hei69f, KF01, She83a, She83b, Sim07, Sim12, Wal03, Wal06]. Our [Gou48a, Hei43o, Hei61e, Hei65b, Hei69k, Hei69l]. outlook [Hei52j]. ovation [Hei74g]. over-relaxation [BPP11]. Overview [Cas02c, Cas02d, Hak02, Hei84d, Hei53t, Wal08]. Oxford [FBD^{+}22]. oxides [Man91].

P [Ano54, Ano68, Ben34, Bro85, Bro86, Bro97, Cas97, Cru23, Gre10, Hei30b, Hei36c, Hei50a, Kra86a, Mat87, Sch91b, TW91, Wig96j, Wil36b]. P. [Hei30c, Hei36c, Lan02]. Paarläufe [FS06]. Pacifist [Sal19]. package [Flo97]. Packets [Hen93, Lan59]. pact [Cor03]. page [FS93]. Paged [BB96]. PageRank [Dje07]. pages [Bed09, Cas05, Cas11, Eis00, Hei31g, LH93, Neu99, Par06, Wat01]. Pair [Mor71, Mor72b, FS06, Mor72c]. Palermo [Pol00]. Pamjati [Geĭ70a]. Papenfuß [Jac03]. Paper [Hen98, AMS04, Cas78, Hei24a, Pau34]. paperback [Fis02, Kra07]. Papers [Aas02c, Ano75, BDR84e, CS79, HDW75, Kra86a, Kra07, Mat86b, Mat87, Sch58, Sch03, Sei90a, BDR84a, BDR84b, CDvW75, CDvW79, Eck70, FB01, Haw11, Wig96m, Yan13a, BDR89a, BDR93, Bro86]. Paradigms [Bok06]. Paradoxes [Joh13, Hei51i]. Paradoxien [Hei51i]. Paraffin [BFH^{+}42b, Hei41a, Hei89i, HBF^{+}89a, HBF^{+}89b, HBF^{+}89c, BFH^{+}42a, HBF^{+}42, BFH^{+}42a, BFH^{+}42b, HBF^{+}42]. Parents [Kan03, Cas05, HHH03]. Pariser [KF02]. Parlamentarischen [HR^{+}89]. Parliamentary [HR^{+}89]. Parr [KF02]. Part [BDR84a, BDR85a, BDR89a, BDR93, Bro86, Car97, Ess71, Gar91, Mat87, vM96, vM99, vM01, vM05b, BDR85b, Eis00, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96c, Hei96a, Hei01a, Hei06, Wei46, Wei99, Geĭ90, Wig96m]. parte [Hei96c]. Partial [Hei32a]. participation [Hei65d]. Particle [Ari23, ARK^{+}99, Ble90, BH83a, BH83b, HVTK^{+}97, Hei74l, RFS^{+}91, SPS^{+}64, BH82, Cam06, DH65b, Fri97, Fri98, Gau14, Hei50a, Hei76h, Wei98]. Particles [DH66, DH84a, DH84b, Hei29c, Hei39f, Hei43c, Hei55g, Hei56c, Hei57g, Hei58a, Hei58m, Hei58n, Hei60h, Hei62f, Hei62k, Hei63i, Hei63j, Hei64a, Hei65i, Hei65n, Hei65e, Hei68d, Hei68g, Hei75c, Hei76a, Hei77a, Hei84-48, Hei84t, Hei84-57, Hei84v, Hei84z, Hei84-59, Hei84-31, Hei84-33, Hei84-34, Hei84-32, Hei84-38, Hei84-39, Hei84-43, Hei93c, HP93, Pau41, Pau93, Rec81a, BDR84g, BDR85b, DHM^{+}59, DH61b, DH61a, DH65a, Fis10a, Fis12a, Geĭ56, Geĭ77c, Hei38c, Hei38b, Hei38h, Hei43m, Hei43j, Hei44a, Hei50j, Hei51m, Hei51h, Hei51i, Hei53o, Hei54c, Hei55c, Hei56k, Hei59b, Hei61b, HD61, Hei61h, Hei61i, Hei62a, Hei63c, Hei63g, Hei65b, Hei66d, Hei67d, Hei71c, Hei74p, Hei76f, Hei84d, Hei84g, Hei84k, Hei84o]. particles [Hei84m, Hei84-58, Hei84-27, Hei89o, HN99, JD06, KLV80, Hei93b, Wig67, Hei93d]. Partisans [Sal19]. partition [KSSA78, SF71]. partnership [BR87]. Paschen [HKvL47, HKvL47]. Pascual [Sch05, Sch07, Bel85, Sch05]. Pasi [Hei77f]. Päsler [HMSH78]. passage [Hei36d, Hei37b, Hei84l, BDR84f, Hei38c, Hei38b]. passing [Ano77a, Ano77b]. Passion [Jon08]. Passions [Ber91, CS94]. passive [Hei47a]. Past [Hei85b]. Path [Eck92, Gar91, Mas14, STCW94, Hau22]. Paths [HPE^{+}33, Hei43i]. pathways [Inf33]. Patrick [Hei75h, Bab02]. patterns [Tem91]. Paul [Car99a, Kra87, OSWR86, Wig96a, BR87, Erm89, Hen00, KPMN20, Kra87, KW87, Log99, Neu99, Wig96k]. Pauli [ARK^{+}99, Bro95, Car97, CP84, Eis00, GRE^{+}01, HEB^{+}80, Hei31k, Hei59j, Hen86, HvMW79, Kra96, MGG^{+}95, Mie81, OSWR86, Pye81, Rec86, vMHW85, vM93, vM96, vM99, vM01, vM05a, vM05b, BC09, FW07, Hei56o, Hei59k, Hei61g, Mas04, Mil10, Nel72, Ser77, dR99, vM05a]. Pauli-Jung-Dialog [MGG^{+}95]. Paulis [Hei59k]. Paulischen [Hei31k]. pbk [Lan08]. Peace [Sal19, Sei90a, Hei70j, Hei89d, Hei89c, Rec02a]. Peaceful [HB82b, Sal19, Hei53k]. Peacock [ARK^{+}99]. Peculiarities [PdK23]. Peierls [Hei63f]. Pekka [And89]. Penetrating [Hei43c, Hei38d, Hei41c, Hei84n]. people [Hei89o, Wei46, Wei99]. peoples [Hei46e]. perception [Klo96]. performance [Bro09, Wil99]. period [HHH03, Kan03]. periphery [Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06]. Perlmutter [BSP^{+}70]. permutation [KSSA78]. Persico [Hei36a]. Personal [Ber01, Jun58, Sei90a, Bet91, Com56, Jun82, Dür76]. personalities [BBH^{+}62]. Persönliches [Dür76]. Persönlichkeiten [BBH^{+}62]. Perspective [Bru80, Cas00, For22, Jam74b, TW10, Bet91]. perspectives [ET07, Hei59e, SW05, Mac58]. perspektiver [Boh64]. Perturbation [Liu72, CGM^{+}71, Car71]. peste [Hei77f]. Peter [And89, Sei90a]. Phänomenologische [HMSH78]. phänomenologischen [Hei66c]. Pharmazeuten [HEB^{+}80]. Phase [Dah02b, DLS76, AS03, BH23b, Dah02a, DRB86, HBB^{+}07, Sch96b]. Phasenbeziehungen [BH23b]. phenomena [Hei26d, Hei39b, SME77]. phenomenological [Arg09, Geĭ67, Hei66c]. Philadelphia [Wat01]. Philipp [Sta36]. Phillips [CS94]. Philosoph [GZ92b, GZ92a, Mey76, Otr79]. Philosophen [Rec02a]. Philosopher [BHN98, GA02, Cam09, GZ92b, GZ92a, Mey76, Otr79, Rec02a, Bok10, Sta11b, Fra10]. Philosophic [Hei52k, Bor53]. Philosophical [Ano54, BDR85a, Hei53b, Hei79f, DB71, Geĭ58, Hei58g, Hei58h, Hei58i, Hei58u, Hei59k, Hei60e, Hei61g, HBD71, Hei74j, Hei74k, Hei89v, Mil90, BDR85b, BDR86, Hei70h, Hei84-59]. Philosophie [Ano06, BMG^{+}50, Hei59e, HH14, KED^{+}60, LSK^{+}76, SCH^{+}78, Str62b, BD92, Hei58t, Hei59d, Hei65j, Hei72k, Hei78, Hei79g, Hei84-60, Hei87c, Hei90a, Hei90b, Hei00c, Hei03, Hei07c, Hei10a, Hei11, Hor66, Bar04]. Philosophie/Bonhoeffer [BMG^{+}50]. Philosophie/Frisch [KED^{+}60]. philosophische [DB71, Hei59k, HBD71, Hei70h, Hei84-59]. philosophischen [Hei58g, Hei58h, Hei58i]. Philosophy [Alt06, Bab02, Bru80, Cam05, CL83, CW85, CF98, Geĭ90, Han62, Hei58t, Hei11, Jam74b, Lac02b, Mac58, Nic90, Ros60, Rud58, SS16, Shi83, Shi86, UM92, Wüt16, dR99, Ano66, Ano68, BD92, Cas75, ET07, Geĭ70b, GHW09, Hei37e, Hei52e, Hei58s, Hei59d, Hei59c, Hei59e, Hei62l, Hei65j, Hei72k, Hei78, Hei79g, Hei82a, Hei84-60, Hei87c, Hei89-35, Hei90a, Hei90b, Hei96c, Hei97, Hei00c, Hei00b, Hei03, Hei07c, Hei07b, Hei10a, Hor66, Lac02a, Pai91, Boh62]. Philosphers [Acz01]. Philpott [HVTK^{+}97]. Photobiology [Ore98]. Photoeffekt [EHR^{+}96]. Photoeffekt/Kleint [EHR^{+}96]. Photographic [Hei30a]. photographische [Hei30a]. Photon [Eps71, SME77]. Photonenemission [URWB95]. Photonenemission/Schnellere [URWB95]. Photonik [URWB95]. photos [HHHH11]. Photovoltaik [MGG^{+}95]. Photovoltaik/Lindner [MGG^{+}95]. Phys [BMG^{+}50, Fri98, Got04, May87, RDM^{+}12a, Sch02a]. Phys. [DPL96]. Physical [Cal14, CS94, CH84, Edd28, Edd78, Edd07, Edd12, Edd13, GL23, Hal09, Har60, Hei30i, HW04, Hof05, HW07, HW12b, LSK^{+}76, Ome75, Oza03, Pop16c, DB71, Edd00, Hei30d, HPE^{+}33, Hei33d, Hei41b, Hei43i, Hei50f, Hei58f, HBD71, Hei84-49, Hei08b, Mil90, Cas79a, Hei41b, Hei57a, Hei72l, Wal04]. Physicist [Ano29, BHN98, Bok10, Cho02, Fra10, GZ92b, Han62, Hei60g, Hei65l, Joh59, Sch59, Sta11b, Ano20, Ber91, Cam09, GZ92a, Gre05, Hei58r, Hei70i, Hof92, KW87, Meh73, Otr79, Rec02a, Sal01, vL47, vL48a, Mey76, Ber91, CS94]. Physicists [Ano33, BBF^{+}57a, Bro97, Cas97, Dür62, Dür80, Hür22, KH19, PS69, Sal19, TEW02, Wal08, vW02, vWR04, Ada72, Bes04, Bey82, Bre97, Cli65, Dür06, Hen07, Him90, HW12b, Wal93, W^{+}87, WdC02, vS14, HW07]. Physics [Acz01, And89, Ano68, ARK^{+}99, Bar07, Bed09, BB36, BDH70, Ble90, BDR84i, BDR85a, BDR89b, BBSR69, Bør12, BH83a, BH83b, BM86, BSP^{+}70, CL83, Cor71, Cus86, DGH^{+}12, Dir76, Fli32, FW20, Geĭ68, Gil09, Gla00, Gou46b, Gou47a, HVTK^{+}97, Hei32c, HSD34a, Hei41a, Hei43g, Hei53l, Hei53m, Hei53n, Hei54b, Hei56c, Hei58d, Hei58s, Hei58v, Hei59c, HBSA61, Hei62l, Hei65f, Hei69h, Hei70a, Hei70e, Hei71i, Hei71j, Hei74l, Hei75a, Hei76d, Hei82a, Hei84h, Hei84-55, Hei89i, Hei93a, Hei95, Hei96b, Hei00a, Hei00b, Hei07b, Hei08c, Hen96, HH11, Hol84, Jor78, Kle17, KED^{+}60, LSK^{+}76, Meh73, Mil85, Mla98, Pag82, Pan02, RFS^{+}91, Ros60, Sap10, Sch86, Sei90b, She83a, She83b, Sta11a, SM41, SPS^{+}64, TEW02, TW10]. Physics [Wal89b, Wal92, Wal17, Wat01, vRFS^{+}00, Bal14, BJN01, BDH^{+}89, Bey77, BDR86, B^{+}61, Bor49, BB05, Bra09, BH82, Büh98b, BW04, Cas09, Con62, ET07, Fer99, FF91, Ful15, Gam85, Geĭ67, GHW09, Haw11, Hei31h, Hei31d, Hei32d, Hei35b, Hei36i, Hei36f, Hei37e, Hei38e, Hei39b, Hei39d, Hei40b, Hei41d, Hei43k, Hei49e, Hei51f, Hei52a, Hei52d, Hei52e, Hei52l, Hei52o, Hei53e, Hei53c, Hei53o, Hei55b, Hei56i, Hei56n, Hei56f, Hei56p, Hei57c, Hei58c, Hei58o, Hei58p, Hei58e, Hei58h, Hei58i, Hei58u, Hei59b, Hei59a, HSBA60, Hei60a, Hei60e, Hei60i, Hei61f, Hei62a, Hei62b, Hei63b, Hei63k, Hei65a, HG65, Hei66a, Hei66c, Hei67l, Hei68c, Hei69c, Hei69d, Hei70c, Hei71a, HnH72, Hei72e, Hei73e, Hei73c]. physics [Hei74f, Hei74d, Hei74j, Hei75i, Hei75k, Hei76e, Hei77d, Hei79d, Hei79f, Hei81b, Hei81c, Hei83d, Hei83e, Hei84-54, Hei84-56, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei89v, HBS96, Hei96a, Hei01a, Hei04, Hei06, Hei10a, Kae48b, Kem85, KHFA67, Meh75, Men36, Pai91, PRW55, Rec92, Rec02b, Rig87, Str11, Wal91, Wal08, WP85, WH07, Wit12, Hei04, Alt06, Ano44a, BDR84c, BDR84d, BDR85b, BDR89b, Bor58, Bun71, Cas75, CCJ^{+}34, Fie80, Geĭ90, Han62, Hei43n, Hei47b, HW49, Hei58t, Hei59d, Hei59e, Hei65j, Hei71b, Hei72k, Hei75s, Hei78, Hei84-60, Hei97, Hei00c, Hei07c, Hei11, LW34, Mac58, Rec96, Sch07, Sch11b, Wal05, Wal07, Ano54, Wig96m, DT18, Ess71, Fel54, Hei54a, Hei32a, Kle54, Wal08, Boh62]. Physics [Hei72a, Huf71, Rud58, Sei90a]. physics/Bigelow [HVTK^{+}97]. Physics/Churchman [KED^{+}60]. Physics/Fraser [vRFS^{+}00]. Physics/Grosse [ARK^{+}99]. Physics/Hüfner [RFS^{+}91]. Physics/Kaiser [BSP^{+}70]. Physics/Klatt [BSP^{+}70]. Physics/Marion [LSK^{+}76]. Physics/Peacock [ARK^{+}99]. Physics/Schroeder [vRFS^{+}00]. Physics/Sharman [SPS^{+}64]. Physik [Ano51, Ano06, BMG^{+}50, BDR85a, BDR84c, BDR84d, BDR85b, BDR89b, B^{+}61, Büh98b, DRF90a, DRF90b, EHR^{+}96, GRE^{+}01, HEB^{+}80, HVTK^{+}97, Hei35b, Hei36i, Hei39b, Hei41a, Hei43g, Hei43n, Hei47b, HW49, Hei53e, Hei53l, Hei56c, Hei59b, Hei62a, Hei62j, Hei71b, Hei84-54, Hei89i, Hei10a, HH14, KED^{+}60, MGG^{+}95, RFS^{+}91, Rec96, Rec02b, SM41, SCH^{+}78, Str62b, SPS^{+}64, Tre78, vRFS^{+}00, Fie80, Hei31h, Hei31d, Hei32d, Hei36f, Hei37e, Hei38e, Hei41d, Hei53c, Hei53o, Hei55b, Hei56f, Hei57c, Hei58e, Hei58t, Hei59a, Hei59d, Hei59e, Hei60a, Hei60i, Hei62b, Hei63b, Hei65f, HG65, Hei65j, Hei66a, Hei66c, Hei68c, Hei70a, Hei70c, Hei71a, HnH72, Hei72k, Hei73e, Hei74d, Hei76e, Hei78, Hei79d, Hei84-60, Hei93a]. Physik [Hei00c, Hei07c, Hei11, Kle17, Men36, Rec81b, Rec92, Sch07, Sch11b, Wal05, Wal07]. Physik-affären [Sch11b]. Physik-Lehrbuch [HEB^{+}80]. Physik/Brüche [SPS^{+}64]. Physik/Greub [SPS^{+}64]. Physik/Hermann [SCH^{+}78]. Physik/Kassel [EHR^{+}96]. Physik/Kudo [RFS^{+}91]. Physik/Vilenkin/Shellard [MGG^{+}95]. Physikalische [BSP^{+}70, Cas79a, DB71, HW07, HBD71, Hei84-49, Hei41b, Hei72l]. Physikalischen [Ano84a, DRF90a, DRF90b, RDP^{+}95, Ben30, Hei30d, HPE^{+}33, Hei33d, Hei41b, Hei43i, Hei58f, Hei08b, Pau31]. physikalisches [HVTK^{+}97]. Physikdaten [SCH^{+}78]. Physikdaten/Behrens [SCH^{+}78]. Physikdaten/Sexl [SCH^{+}78]. Physiker [Bes04, Bey82, Dür62, Dür80, GZ92b, GZ92a, Him90, Rec02a, URWB95, vRFS^{+}00, vS14, vW02, vWR04, HW07, Lor15, Mey76, vL47, vL48a]. Physikerstreit [Sal01]. Physikertagung [SPS^{+}64]. Physik“ [Hei43k]. physique [CCJ^{+}34, Rhe91]. physiques [Hei57a, Mil58]. picture [Fri97, Fri98, Nar93, Sny82, Kas11]. Pictures [Dir64, Per64, ST65, dG14, Her76c, Her84, Her01, Tel93]. Pierre [Mar73]. Pile [HW48]. pilot [Hei42b, Hei89g]. Pinzette [DRF90a, DRF90b]. Pioneer [CS94, Jac01]. Pionier [Jac01]. Piper [Bun71]. places [Hei89o]. Planck [Adl03, Fis10a, Fis12a, Hei31g, HPE^{+}33, Hei35b, Hei38g, Hei43i, Hei43p, Hei43q, Hei47e, Hei53l, Hei58l, Hei62j, HMW^{+}84, Hei86i, Hei89a, Rec81b, Tre01, vBL96a, vBL96b, Cas75, CG14, Fis10a, Fis12a, Hei31g, Hei38g, Hei45a, Hei47d, Hei47e, Hei49b, Hei58l, Hei58g, Hei58h, Hei58i, Hei58u, Hei60e, Hei74k, Hei86i, Hei89a, Rhe91]. Planck-Feier [Hei89a]. Planck-Institut [Hei43p, HMW^{+}84]. Plancks [Hei49b]. Plancksche [Hei45a, Hei58g, Hei58h, Hei58i]. Plane [Hei57d]. Plänen [Hei56d]. planka [Geĭ58]. plans [Hei56d, Hei57d, Kra87]. plant [Hei42b, Hei56d, Hei89g]. Plaschko [MGG^{+}95]. plasma [Hei63k]. Plasmaphysik [RFS^{+}91, Hei63k]. Plasmarohren [Ano84a]. Plasmazustand [RDP^{+}95]. Plato [Fer99, Hei53o]. Platone [Fer99]. Platons [Hei53o]. Play [Aas00, Cas22, Fra05, MB11a]. Plays [Ros00]. Plenarsaal [Hei53q]. plenipotentiaries [Mai07]. Plenum [Klo96]. plight [Hei50d]. plunder [Gim90]. Pluralism [Bok08]. Plus [EHR^{+}96]. Pobedy [Smo92a]. Poenaru [HVTK^{+}97]. Poet [Ano33]. Poincaré [Adl03]. Point [New35]. Poland [Got04, Ber04]. Polarisation [Hei25b, Hei25a]. polarization [Hei25b, Hei25a]. Polarized [OSWR86]. Pole [Ano04b]. poles [Ter82]. Policies [Sei90a, CHW91, WC84]. Policy [Car02a, Car02b, Eck90, BBH^{+}62, Hei72l]. Polish [Inf33]. política [Hei96c]. Political [BDR89b, Hei73b, Sei90a, Hei61e, Hei72d, Hei80, Hei83a, Hei91a, Sch05, Sch07]. Politics [Hab69, Jor78, SW05, Sim12, TW10, Bey77, BDR89b, BB05, Car99b, Fav05, Hei69g, Hei74m, Hei96c, Sch07, Wal08, Sei90a]. Politik [BDR89b, Sch07, Fav05, BBH^{+}62, Hei69g]. Politische [BDR89b, Hei80, Hei83a, Hei91a, Sch05, Sch07]. politischen [Hei61e]. polity [Pai91]. Pollaczek [Koo89]. polya [Hei62n]. polymer [Fed04]. Polymers [EHR^{+}96, HVTK^{+}97]. Polymers/Ambegaokar [HVTK^{+}97]. Polymers/Ruder [EHR^{+}96]. Polynomial [FT94]. polynomials [BMP87, BM94, DLW08, Koo89, Per88, VZ09]. polynuclear [OD03]. Pomerans [Han62, Sch59, Huf71, Joh59]. Pople [KF02]. popular [BDR85a, BDR85b, BDR86]. Porazhenija [Smo92a]. Poröses [RDP^{+}95]. Portrait [CW85]. Portraits [LMOS15, Far01]. Porträts [LMOS15]. Portrays [Aas00]. Portuguese [Hei96c]. Position [Oza02, SAR^{+}09]. Positive [Wan01, Pac88]. Positivism [Hei91c, Spi98, Lan72, Hei31g]. Positivismo [Spi98]. Positivismus [Lan72, Hei31g]. Positron [HE06, Hei34b, Hei34a, HE35, Hei52s, Hei52s]. Positrons [Hei34a, HE35]. Positrons“ [Hei34b]. possibilities [Hei53g, Hei56j, Mon04, Hei89w]. possibility [Hei39e, Hei40a, Hei41f, Hei69k, Hei69l, Hei89h, Hei89-36, Hei89m, Hei89x, Hei89-40, vW40]. possible [Hei85d, Per39]. Post [Hei56b]. postdoctoral [Hei01b]. postgraduate [Lic16]. Postscript [Pau93]. postwar [Gim90]. potency [Car76]. potential [NRS71, TR09]. Pour [Cap57a, Cap57b, Kar07]. Power [Gou47c, Hay90, MWL^{+}15, Sch91a, Sei90a, Bet91, Wal89a]. Powers [Cas93a, Hol93a, LH93, Pei93b, Swe93, Wil94, Ber93, Kra93]. Pp [Bro72, Cas93a, Hen86, Hen11, Klo96, Kra07, Sta11b, Bok10, Bun71, Cru23, Fra94, Hay90, Hei53b, Hol93a, Jor78, Kra86a, Lan02, Lan08, Lib68, Log00, Pol00, Tur06, Wil94, Ano94]. Practice [Sta09a, Sta11a, Set10]. Prag [Hei89t]. Pragmatism [Spi98]. Pragmatismo [Spi98]. Praktiker [EHR^{+}96]. Praktiker/Mathcad [EHR^{+}96]. praktische [vW1 ]. Präp [DDH41a, DDH41b, DDH89f, DDH89g]. Präparat [DDH40a, DDH89d, Hah40, Hei41a, Hei89i]. Präsidenten [Ano71, Hei50g, Hei53q]. precision [Wit12]. predicting [OD03]. Prediction [GLR14]. Preface [Hei93b, Hei52r]. Preis [Ano35a, Ano84b, Ben30, Ben34, Ben35, Clu37, Deh36, Deh37, DS43, DRF90a, DRF90b, Mos42, Wil36a, Wil36b, Wit83]. Preis/Verwundbarkeit [DRF90a, DRF90b]. Preliminary [BFH^{+}42b, HBF^{+}89c]. preparation [DDH40a, DDH41a, DDH41b, DDH89d, DDH89f, DDH89g, Hei41a, Hei89i, Hah40]. prepare [HW84b]. prescriptive [Kav24]. prescriptive-dynamical [Kav24]. Present [Hei29c, Hei50i, Hei53u, Hei63i, Hei65n, Hei68g, Hei74l, Hei79c, Hei79b, Hei84-34, Hei84-38, Hei84-46, MS15, CHW91, Geĭ56, Hei38f, Hei49e, Hei52d, Hei65b, Hei69e, Hei84-27, WC84, Hei43o]. present-day [Hei49e, Hei52d]. presentation [Hei89-27]. presented [CDvW75, CDvW79]. President [Ano71, Hei50g, Hei53q]. Press [Bed09, Bok10, Cas11, Hay90, Jor78, Lan02, Neu99, Par06, Sta11b, Tur06]. Preußischen [Hei86c]. Prevent [LH93]. Preview [Gou47d]. Primacy [Eck90]. primary [Hen96, HH11]. Primas [MGG^{+}95]. primer [SR92, McC24]. Princeton [Ano84a]. principe [Rhe91]. principes [Hei57a, Mil58]. principielle [Hei35e]. principio [Bla86]. Principle [Ari23, Bac52, Bel92, Cul05, Har73, Jau29, Loe63, PC14, Pan02, RS94, Sei05, Tya63, dB90, Ano12, Bel85, BCL^{+}69, D’A03, DJR93, Hab83, Hei35e, Hei52l, JLP^{+}02, Lap91, LM29, LH01, MRV24, Mes04, NAZ02, Oza02, Rob69, Wil99, Wit12, Wol14, Ano35c, Ano83b, Ano00, App01, AP22, Bla86, Bro73b, Com31, Cop81, Cow49a, Cow49b, Dah02a, Dah02b, Dar31, EMGM94, Fli32, Fli33, Gam58, Hei25b, Hei31k, Hei75f, Hen93, HK00, HS81, Ing70, JHB99, Lar30, Man11, Mar31, McC94, Moo93a, Moo93b, New35, Noy31, OW10, Pol95, PC77, Rob80, Rot93, San81, Sch00a, Sch30, Sch94, SR49, STCW94, Tut05, Van41, Žup65, Rob82]. Principles [Cha92, CH84, Hei30c, Hei30i, KED^{+}60, Los61, RFS^{+}91, WMG^{+}92, Hei27f, Hei30b, Hei30d, Hei50f, Hei57a, Hei58f, Hei08b, KLV80, War90, Con29, Meh87]. Prinzipielle [Hei36f]. Prinzipien [Ben30, BMG^{+}50, Hei30d, Hei58f, Hei08b, Pau31]. Priroda [Geĭ77c]. Prisma [Pol00]. Privatarchiv [HHHH11]. private [HHHH11]. Prize [Ano29, Hei32c, HSD34a, Rec10, Ano33, Läs13]. Prizes [LW34]. prób [Kar06]. probabiliste [dB82]. Probabilistic [RFS^{+}91, dB90, dB82]. Probability [Plo10, HVTK^{+}97, Hei38i, Hei84-45]. Probably [Bro97, Cas97, Bre97]. probe [sMDN^{+}11]. Problem [Bor58, Gro96, Hei24d, HJ26, Hei48k, Hei75b, Maa05, DC87, GV85, GV86, Hei26a, Hei42c, Hei48b, Hei52g, KJ06, Lap91]. Problem-Selection [Hei75b]. Probleme [Hei58h, Hei58i, Hei70h, Hei84-54, Hei84-59, Hei27a, Hei31h, Hei32d, Hei64e]. problemi [Hei60e]. Problems [BDR84i, Bor53, DH32, Geĭ68, Hei75s, Hei76d, Hei84h, Nis32, CL05, Dun86, Geĭ58, GZ95, Hei27c, Hei38e, Hei49e, Hei50a, Hei52d, Hei52k, Hei58h, Hei58i, Hei58u, Hei60e, Hei67l, Hei70h, Hei74k, Hei75i, Hei77d, Hei79f, Hei84-54, Hei84-56, Hei84-59, Hei27a, Hei31h, Hei32d, Hei64e]. problemy [Geĭ58, Geĭ68]. proceedings [AK15, NBH97, BCEP94]. process [Tel61]. processes [Hei32h, Hei49k, Hei49d, Hei84p, Tho80]. Production [HEB^{+}80, Hei49i, Hei51l, Hei52h, Hei84-28, Hei41f, Hei43l, Hei49d, Hei52b, Hei55e, Hei83g, Hei84p, Hei89-36, Hei89l, Hei89x, vW41b, vW41b]. Production/Olson [HEB^{+}80]. Produktion [Hei51l]. proekta [DF96]. Prof [Ano73, Eic50, Hei36i, Hei47f]. Prof. [Ano73, Hei36i, Hei47f, Hei89t]. Profanität [Hei50b]. profanity [Hei50b]. Professional [Cas79b]. Professor [Ano29, Ano71, ARK^{+}99, GZ95, Hei53q, Hei56c, Hei65f, Hei67h, Hei67i, Hei72m, Hei86g, Hei89-37, Eic50, Gou47a, Hei67h, Hei67i, Hei70j, Hei72m, Hei75n, Hei86g, Hei89-37]. professors [Wei46, Wei99, LW34]. Profound [Bet91]. Profs. [Ano33]. progetto [Ces18]. Program [Ano93a, Bet93, Cus86, Hei93d, Sei90a, Flo97, Hei89-38, KSSA78, Rec96]. Programm [Ano84b, URWB95, Ano83a, RDP^{+}95]. Programm/1 [Ano84a]. Programme [URWB95]. Programme/Neue [URWB95]. Progress [Gou49, Hei49f, Hei58a, Hei71k, Hei84-37, Gou47a, Hei34j, Hei70b, Hei74c, Hei84d, Hei37c, Hei62h]. prohibition [Hei89-32]. Project [Car99a, Hen00, Lan02, Log99, Neu99, PdK23, Wal99, Bal07, Ces18, DF96, Nag16, Pop18, Ros98, Bet00, BB96, Cas02b, Cas17, Gou47c, Kan02, Sim12, Pei93b, Sei90a]. Projects [MWL^{+}15, Hei65d]. Projekt- [URWB95]. Proletarian [Jor78]. Proliferation [MWL^{+}11]. prolongation [DC87]. Promotionen [Cas03, Hei01b, HR02b, RW02]. Proof [HN55, DDH42a]. Propaganda [Wal92]. Propagation [BSP^{+}70]. Properties [ARK^{+}99, Hei84-48, MF69, BPS74, CCJ^{+}34, Flo97, HWY69, HY72, Hei84g, LL86, May87, Wre75]. property [DL08]. Proposed [HS81, Rob80, Hei42b, Hei89g, Rob82]. propriétés [CCJ^{+}34]. Prospects [Rec81a, Boh64, Boh67, Hei63k]. protagonisti [Hei84-55, Hei00a, Hei08c]. protagonists [Hei84-55, Hei00a, Hei08c]. Protest [BS47, Lor15]. Protestant [BBH^{+}62]. protocols [Hof93, Wal93, Hof23]. Protokoll [Hei89-38]. Protokolle [Hof93, Hof23, Wal93]. proton [Hei53i]. Proved [Acz01]. Proyecto [Bal07]. Prüfungsprotokolle [Hei01b, HR02b, RW02, Cas03]. Prussian [Hei86c]. PS [Hei68b, Hei68b]. Pseudoparticle [Woo77]. Psychoanalysis [CL83]. Psychophysical [Cop81]. Pu [DK78]. Public [Cas11, Del13, Hei73h, HGW^{+}89a, Hen10, Lal15, Sta11b, BJN01, Car10, Hei51d, Hei53h, Set11]. Publication [GR08]. Publisher [RDM^{+}12a]. Publishers [Ano68]. Publishing [Hei53b, Log00]. Pullman [Hei51e]. Pullmann [Hei51e]. Pumping [BSP^{+}70]. pupils [RKW05]. Pure [Hei56p]. Pursuit [Mil10]. Put [Kae31]. puts [Dur00]. Pythagoras [Gre10, Cre10].

QCD [Wal83]. Quadratic [Gid71]. quadric [GS12]. Quandary [Kae31]. Quanta [BH84, BH28, Hei57a, Inf33, Inf34, DB71, HBD71, Mil58]. quantal [CGM^{+}71]. Quanten [DB71, HBD71]. Quantendynamik [HP29]. Quantenei [Rec01]. Quantenelektrodynamik [Hei79c, Hei79b, Hei85b]. Quantenlogik [Hei59g]. Quantenmechanik [Ano35b, BHJ26, BHJ63, Deh36, GRE^{+}01, Haa34, HJ26, Hei27c, Hei77e, Hei79b, Hei85d, Sch26, BJ25, Bor61, Bor63b, Fue61, Hei26b, Hei26c, Hei26d, Hei30h, Hei30g, Hei31b, Hei31e, Hei34d, Hei60b, Hei65c, HL84, Hei10b, Jac01, Wey31a, Hei28b, Hei36b]. Quantenmechanik/Thomas/Weise [GRE^{+}01].

Quantenmechanik“ [Hei27f]. quantenmechanische [Tel61]. Quantenphysik [Ful15, GRE^{+}01]. Quantenrätsel [Bes04]. Quantensprung [Fis10a, Fis12a, Hei46c]. quantentheoretische [Hei25c, Hei26f, Hei84-42, Hei27d]. quantentheoretischen [Hei27e]. Quantentheorie [AH57, Ben30, BH24b, Hei22c, Hei24d, Hei25d, Hei28d, Hei29a, HP30, Hei38i, Hei43h, Hei46b, Hei50j, Hei50h, Hei51m, Hei54f, HKM55, Hei56e, Hei56m, Hei62a, HH84, Hei84-47, Hei84-45, Hei86i, Hei93c, Hou30, LSK^{+}76, Hei29b, Hei30d, Hei38f, Hei47f, Hei51a, Hei56h, Hei58f, HB63, Hei69o, Hei79g, Hei87c, Hei90a, Hei90b, Hei03, Hei08b, LBBH69, Hei36c, Pau31]. Quantentheorie/Berichte [LSK^{+}76]. Quantenvektoren [Hei27b]. Quantenwelt [Lan72, Arr06]. quantique [BD92, Rhe91]. Quantiques [Hei31c]. Quantisierung [Hei53y]. Quantitative [BDR84j, Hei23a]. quantitativer [HVTK^{+}97]. Quantitatives [BDR84j, Hei23a]. quantitie [Hei43m]. Quantities [Hei62k, KW02]. Quantity [Ber35]. Quantization [Hei55f, Hei57f, Hei84-29, dG14, GKZL96, Hei53y, KtL79, Tru75]. Quantum [Aga02, Ano75, Arr06, AG88, Bar07, Bed09, Bel99, BBE92, Bes04, BHX96, Boh62, Bok08, Bok10, BHJ67, BH75, BGĬ90, BW00a, BW00b, Bro02a, Bro02b, Bru80, Cas92b, CK02b, CH84, Cus94, DH32, DH62, DHYY65, EHR^{+}96, ES95, Eng09, ET07, Fin96, Fra10, FBD^{+}22, GGK02, Gar91, Gil09, GL23, HR02a, HVTK^{+}97, Hee95, Hei28a, Hei28b, Hei30b, Hei30c, Hei30i, Hei31b, HSD34b, Hei50h, Hei50i, Hei57g, Hei59g, HB63, Hei65k, Hei75j, HDW75, Hei77b, Hei79c, Hei79b, Hei84-53, Hei84-30, Hei84-31, Hei84-46, Hei93c, Hei93e, Hei96d, Hür22, Jad14, Jam74b, Jau29, Joh13, Jon08, Kas11, KED^{+}60, Log00, Los61, sMDN^{+}11, ML71, Mil12, Nis32, OW10, Pag82, Plo10, RG72, RS11b, Rov21, Sap10, Sch96b, Sch02c, Sch58, Sch94, Sch03]. Quantum [Sei05, Sei90b, Sta11a, Sta11b, Tho80, Vya99, Wey31b, Wey50, WZ83, Wüt16, vdW67, Alb73, Ano35b, Ano11, Ano20, AH57, BD92, Blo76a, Blo76b, Blo85, Bok04, BH24b, BJ25, BHJ26, Bor61, Bor63b, BHJ63, BJ67, BPS74, BP75, Cam09, Car71, Cas76, Cas09, CGST91, CK02a, Cli65, Con62, CG14, D’A03, DT18, Dys51a, Dys51b, DLS76, Enz83, FO94, Fed04, FR08, Fis10a, Fis12a, Fri13, Fri11, Fue61, Ful15, Gal01, Gam85, Geĭ77a, GZ95, GHW09, Gre05, Haa34, Haw11, Hei22c, Hei24d, Hei25d, Hei25c, HJ26, Hei26a, Hei26f, Hei26d, Hei27e, Hei27b, Hei27f, Hei27c, Hei28d, Hei29a, Hei29b, HP29, Hei30h, Hei30g, Hei30d, HP30, Hei31e, Hei33a, Hei34d, Hei35e, Hei36b, Hei36c, Hei38f, Hei38i, Hei43h, Hei46b]. quantum [Hei46c, Hei50f, Hei50j, Hei51m, Hei51a, Hei53j, Hei54f, Hei55d, HKM55, Hei56e, Hei56h, Hei56m, Hei57a, Hei58f, Hei58u, Hei60b, Hei62a, Hei65c, Hei69o, Hei73d, Hei75l, Hei77e, Hei77j, Hei79f, Hei83b, Hei83c, Hei84-51, HH84, Hei84-47, Hei84-42, Hei84-45, Hei85b, Hei85d, Hei86i, Hei91b, Hei07a, Hei08b, HBE87, Hou30, Jac01, Jes98, JLP^{+}02, KF07, KHFA67, LM87, LBBH69, Lan72, LL86, LM29, LFK99, LH01, Low97, MSUU80, May87, Mil97, Mil90, MB11a, NKD^{+}11, Nar93, OS07, PC77, Raz11, Rec01, SW72, Sch26, Sch66, Set10, SME77, Str11, Tel61, VL97, Whi96, dSL02, vdW68, BV09, CDvW75, CDvW77, CDvW79, Geĭ77b, Hei26b, Hei26c, Hei27d, Hei31c, Hei47f, Hei64d, Hei79g, HL84, Hei87c]. Quantum [Hei90a, Hei90b, HBS96, Hei03, Hei10b, Lev95, Wal08, W^{+}87, WdC02, Lic16, Sta11a]. Quantum- [Wüt16]. quantum-classical [JLP^{+}02]. Quantum-Electrodynamics [Hei84-53, Hei53j]. quantum-mechanical [Geĭ77a, HBE87, Tel61]. Quantum-Theoretical [Hei77b, Plo10, Hei77j, Geĭ77b]. quantum-to-classical [LH01]. Quantummechanik [Hei26a]. quasi [BT96]. quatre [VB15]. Quest [DT18, Hay90, KH19, Com56, Wal89a]. question [Hei25b, Hei51m, Hei89-32, Hou41]. questioners [Cli65]. Questions [Hei52l, Hei36f, Hei58g, Paw02a, Hei69i]. Qui [Dem15, VB15].

R [Bun71, CS94, Deh37, Erm89, Hei39b, Hei63f, Sei90a]. R. [Hei39b]. rêveurs [VB15]. raccord [dB58]. Race [Dah99, Gou46a, Sei90a, Wat01, Mac86, Mac87, Sei90a]. Rad [HvL47]. radially [PL91]. Radiation [EH38, HEB^{+}80, Hei31j, Hei39f, Hei39h, Hei43o, Hei46a, Hei51l, Hei53u, Hei75s, Hei76d, Hei84h, HMW^{+}84, BBB^{+}84, GE38, GE39a, GE39b, Hei31a, Hei32e, Hei36g, Hei36j, Hei37c, Hei37f, Hei38d, Hei38a, Hei43p, Hei75i, Hei77d, Hei84i, KH25]. Radiations [RCE30]. Radical [Hee75]. radicals [Bro09]. Radio [BMG^{+}50, Hei89y]. Radio-Technisches [BMG^{+}50]. Radioactive [Hei43d, MA40, SPS^{+}64, RCE30]. Radioactivity [GR09]. radioaktive [Hei43d]. radiological [Bro92]. Radium [Kae48b]. radius [JLP^{+}02]. Radon [He04]. Rahmen [Hei46b, Hei53d]. Rahmenplan [Ano84a]. raised [Boh67]. Rajnath [Tiw14]. Raleigh [BCEP94]. rally [Hei53q]. Randomly [Rus64]. Rang [HVTK^{+}97]. Rankings [Dje07]. Räntsch [BMG^{+}50]. rapport [Hei62d, Hei62e, Ano44a]. Rapports [CCJ^{+}34, Hei62c]. rasche [DDH42b, DDH89b]. Raster [DRF90a, DRF90b]. Raster-Tunnel-Mikroskop [DRF90a]. Rasterkraftmikroskops [URWB95]. Rasterkraftmikroskops/Magnus [URWB95]. Rasterkraftmikroskops/Magnus-Haus [URWB95]. Rate [Žup65]. Rates [HR^{+}89]. Rath [GRE^{+}01]. Rätselkabinett [HVTK^{+}97]. Raums [Hei56m]. Ray [Hei75s, Hei89s, Hei37a, SH22b]. Rayner [CS94]. Rayner-Canham [CS94]. Rays [Hei39g, Hei84y, RFS^{+}91, BDR84g, EH84, Hei31i, Hei32h, Hei33c, Hei39j, Hei41c, Hei43r, Hei51f, Hei53t, Hei53l, Hei84n, Hei84-41, HW84a, Hei84-44]. Razavy [RS11b]. reaction [El 00]. reactions [Hei89w, Hou41]. Reactor [GORS09, KH19, Hei57d, Kar05b, McC24, Wal05, Wal07]. readers [Mil12, Hei60d]. readings [WH07]. Reaktionen [Hei89w]. Reaktor [Kar05b]. Reaktorforschung [Wal05, Wal07]. Reaktorstation [Hei57d, Hei56d]. Real [DT18, Arg09, FR98, Hei31g]. reale [Hei31g]. Realität [Hau22]. realitè [Rhe91]. Reality [Bor58, Cru23, Hei96b, HK19, Mes04, BDR84c, Hau22, Hei60i, Hei82c, Hei84-52, Hei89z, Ano25]. Realization [LC71, AFW94, Liu72]. Realizations [BP75]. realm [Kae48b]. realtà [Arg09]. reason [BLW19]. Reasoning [HVTK^{+}97]. Rebhan [vRFS^{+}00]. Reborn [Gil09]. Receive [Ano29]. Recensioni [Cir88]. Recently [Aas02c]. Reception [Hei86a, Hei89-28]. Rechenberg [Cas05, Cru23, Hen11, Hof11, Kra86a, TW91, Bro85, Bro86, Eck12a, Eck12b, Jac11a, Jac11b, Mat87, Sch91b]. Rechnen [BMG^{+}50]. Recollections [Kra86a, Mat86a, Mot85, Roh85, Wei85, Hei84a, Pei71]. Reconsidered [Hol22]. Reconsidering [Sta11a, BV09]. record [Ano99]. recordings [Ber96, Paw06]. recovery [Hei89n]. Redaktion [Ano84a]. Rede [Hei72f, Hei72m, Hei89-28, Hei89-27]. Redivivus [Ros79]. Reductionism [Bok08]. redundant [Ter82]. reestablishment [Hei56g]. Reexamining [Bok08]. refinements [Wit12]. Refining [MRV24]. Reflections [CS94, Shi83, Shi86, DB71, Wal93, Spi98]. Reflexiones [Spi98]. reformulation [Oza03]. Regeln [Hei24d]. Regime [MP01, Sch11b]. Regimmotstånd [Sch11b]. Regnant [Sch86]. regular [ES95]. regularization [Sch02a, Sch02b]. reibende [Hei22b, Hei84w]. Reich [BDR89b, GHW36, Hau22, Hei67a, HW11, Bal13, Bal14, Bed03, Bed04, Bey77, Bey82, BLW19, BDR89b, Cas92a, Cas93c, Hau22, Hei47a, Hei67a, Hei68i, HGW^{+}89a, Hof05, HW07, HW11, HW12b, Jor78, Pop18]. Reichsminister [GHW36, HGW^{+}89a]. reinterpretation [Hei25c, Kav24]. Reise [Hei75p]. Rejection [Ser77]. Related [Sei90a, Mai07]. relatif [Per39]. Relating [Fal64, Aas02b, Aas02d]. Relation [AG88, Bok08, DGH^{+}12, NRS71, AS03, Bok04, Chi01, Jau93, Mar08, Oza03, Sch26, Sch66, Tri02, Wis98]. Relations [Aga02, Fed18, Geĭ77b, HN55, Hei76c, Hei77b, Hei77g, Jam74a, KF02, BH23b, DD13, FT94, Geĭ77a, Hei25c, Hei31d, Jes98, SK09, vD15, CM02, Hei53e]. Relationship [RDM^{+}12a, RDM^{+}12b, Hei75m, Sch13]. Relativistic [Pau41, Fri97, Fri98, Hei50h, Low97, SH22b]. relativistische [SH22b]. relativistischen [Hei50h]. Relativitätstheorie [Bes04, KPMN20, vRFS^{+}00]. Relativity [Bes04, Cop81, HVTK^{+}97, Fin96, Cop81]. relaxation [BPP11]. Release [Aas02b, Aas02d, EHR^{+}96, Hei49f, Hou41]. Released [Aas02c, Paw02a]. relevant [Mai07, RDP^{+}95]. religiao [Hei96c]. religion [Hei91c, Hei96c, Tre01]. religiose [Hei73f, BDR85b]. religious [BDR85b, Hei73f, Hei74n]. Remark [Hei24a]. Remarks [Con29, Hei58n, Hei63j, Hei77g, Hei84t, Hei84-33, Ome79, Wig96d, Hei50g, Bor61, Hei31a, Hei34b, Hei34a, Hei35a, Hei48b, Hei52b, Hei53p, Hei56e, Hei75f, Hei84-61]. Remarques [Hei53p, Hei84-61]. Remembering [Wig96k]. remembrances [LL77, Rec91c]. Reminiscences [Blo76b, Hei85e, KW87]. Removal [LD66]. Renewal [Hab83]. reopen [Bro93, GP92b]. reparations [Gim90]. Reparatur [Ano84a]. replaced [SK09]. Replies [Kae47]. Reply [HH75, Lau58b, Mor48, Rob82, Hei36i]. Report [Hei41e, Hei62d, Hei62e, Hei68g, Hei84-34, Eic53, Hei61h, Hei61i, Hei64b, Hei67p, Hei89x, Hei89-34, HBF^{+}89c, KHFA67, Ano44a, BBF^{+}45, Döp42, Hah40, Har43, Hei30a, Hei40a, Hei41a, Hei84-48, Hei84g, Hei89h, Hei89i, HBB^{+}89, Hei02, vW41a]. Reports [DW52, HSD34b, Hei49f, Hei62c, KRW05, Dir28, Hei01b, Hei05b, CCJ^{+}34].

Representation [Bor72, Hei58v, WB03, BT78, Bor70, KSSA78, SL70, d’E92]. Representations [Bil03, FO94, FO99, ES95, EHUY92, Low97, MMSV98, Sch99, Tor91]. Repressed [Rad83, Rad91]. Repression [Wal06, Wal03]. Republic [Hei57d, Hei65d, Hei86a, Sch07, Hau22, Sch05]. Research [Eic50, Eic53, Gou47a, Gou49, GWY05a, Hei30a, Hei47g, Hei47h, Hei49f, Hei50c, Hei50g, HR53, Hei53q, Hei58w, Hei60h, Hei84-35, Hei84-36, Hei84-32, Hei89-38, Lib68, TW10, BDR89b, CR13, GWY05b, Hau22, Hei48l, Hei49c, Hei49g, Hei51e, Hei53f, Hei56n, Hei56j, Hei56p, Hei63k, Hei64e, Hei65d, Hei68b, Hei89p, Irv67b, Irv67c, KP07, Mac93, Mai07, Rec96, Rot93, Wal05, Wal07, Ano29, Hei54b, Hei73a, Hei75r, Läs13]. researcher [Lic16, Rec02a]. researchers [Hei36e, Hof23, Jun57]. reshaping [Mil97]. Resistance [Pow76, Sno49]. Resistive [Gro96]. resolution [Gau14]. Resolved [RFS^{+}91, SAR^{+}09]. Resolved-sideband [SAR^{+}09]. Resonance [WMG^{+}92, Hea11, Hei25a, Hei26a, Hei27c, LL82]. Resonanz [Hei26a, Hei27c]. Resonanzfluoreszenz [Hei25a]. respected [Sch13]. response [Tan04, vV78]. Responsibility [Hei70e, Wal06, KF01, Tel71, Wal03]. Rest [Ano93b]. Results [Ano34, BFH^{+}42b, CZ81, CRR93, DMdA06, Dob77, HBF^{+}89c]. rész [Hei83h]. rethinking [Pfe77]. rethinks [Car03]. Retires [Cor71]. retrospective [Fie80]. Rev [RDM^{+}12a]. Revelations [Ber93]. Reverse [Cop81]. Review [And89, Ano54, Ano68, Ano94, Ano25, Bed09, Ber93, BDR84e, Bok10, Bor58, Bor49, Bor53, Bro72, Bro85, Bro86, Bro95, Bro97, Bun71, Car97, Car99a, Cas74, Cas93a, Cas97, Cas02a, Cas03, Cas05, Cas11, Cru23, DT18, Dre92, Ess71, Fay07, Fel54, Fer07, For09, Fra94, Fra10, Gre10, Hab83, Han62, Hay90, Hei72a, Hei25a, Hei27b, Hei27a, Hei28b, Hei31b, Hei31c, Hei32a, Hei35b, Hei36b, Hei36a, Hei36c, Hei39a, Hei39b, Hei43g, Hei43h, Hei43i, Hei50a, Hei50b, Hei51l, Hei58a, Hei71k, Hei75g, Hei84-37, Hen86, Hen00, Hen10, Hen11, Hol93a, Hol00a, Huf71, Joh59, Kea09, Kle54, Kra86a, Kra96, Kra07, Lan02, Lan08, Lau07, Lev95, Lib68, Log99, Log00, Lus10, Mac78b, Mac58, Mar94, Mar73, Mil58, Mot85, Neu99]. Review [Par06, Pau31, Phi92, Pol00, Pye81, Rec84, Rec85b, Rec86, Roh85, Ros79, Rud58, Sap10, Sch59, Sch91b, Sta11a, Str62a, TW91, Tre01, Tre78, Tur06, Wat01, Wei85, Wig67, Wil94, All06, BDR84b, Ols63, Pei93b, Wal99, Eck12a, Hei53u, Paw02b, Wal04]. Reviews [BDR86, Dje07, Gar91, Lal15, Sei90a, Sta11b, Swe93, Wal08, Jor78, HPE^{+}33, HvLM^{+}35, HJFS36, HSBR36, HGT^{+}39, HSBB43, HF46, HHKG50, KHBF43, MPB^{+}24, Som36]. Revised [Ano34, Hei31b, Dür80]. revision [Lur95]. Revisited [HS83, Wil84, Ano98, Bro77, Wal19, vV78]. Revisiting [BL08, Luc07]. Revolution [Cas92b, Han62, Kae48b, Mac58, O’F92, Rov21, Rud58, Bel99, Gre05, Hei58s, Hei59c, Hei62l, Hei00b, Hei07b, Lud01]. Reweighting [dSL02]. rewriting [Lur95]. Rey [Hei89s]. Reynolds [vW48]. Reynoldschen [vW48]. Rezension [Eck12a, Eck12b, Eis00, Ess71]. Rezensionen [Som36]. Rhetoric [CS94]. Rhodes [Sei90a]. Richard [Bro97, Cas97, Hei86l, Sei90a, Hei86l]. richtet [DRF90a, DRF90b]. Richtigkeitskriterien [Hei73e]. rigged [WB03]. right [ES95]. ring [GBLW86]. rings [Hei72f]. Rise [Dje07, Mac77, Rad83, Rad91, She83a, She83b]. Risikostudie [DRF90a, DRF90b]. rivoluzione [Lud01]. Road [Sei90a, Bet91]. Robert [AO44, Hei72a, Gre10]. Roger [Sei90a]. Rohr [EHR^{+}96]. Rol [Geĭ67]. Role [Eck90, Hei74l, Hei75b, dR99, AS03, Bro93, Hei31d, Hei38a, Hei51d, Hei53h, Hei66c, Hei84i, Mai07, Zan02a, Geĭ67]. Roles [BW00a, BW00b]. Rolle [Hei31d, Hei38a, Hei51d, Hei53h, Hei66c, Hei84i, Mai07]. Romanian [Hei77f]. romoting [Hei31g]. Ronald [CS94]. Röntgen [MGG^{+}95, Som36]. Röntgendubletts [SH22b]. Röntgens [Som36]. Röntgenspiegel [RDP^{+}95]. Rontgenstrahlen [Hei33c, Hei31i, Hei84-41]. Roosevelt [Dem15]. Rosbaud [Kra87, Kra87]. Rose [Hen00, Lan02, Log99, Neu99, Car99a]. Rosenfeld [CS79]. Rotation [HEB^{+}80]. Rotation/Haas [HEB^{+}80]. Rothberg [HVTK^{+}97]. Routledge [CF98]. Row [Ano68, Bun71]. Royal [Cla23]. RTM [URWB95]. RTM-induzierter [URWB95]. Ru [vW1 ]. Rückblick [Fie80, Paw02b]. Ruder [EHR^{+}96]. Rudjer [Hei59f]. Rühle [WMG^{+}92]. Rule [Sno49]. rules [Hei24d]. Rumscheidt [Cru23]. run [Hei57b]. Rundfunk [Hei89y]. Runge [BMG^{+}50]. Runge/Willers [BMG^{+}50]. runs [FS06]. Rupp [vD15]. Russell [Bro72, Hab83, Mar73, FW07]. Russian [Ano33, Ano77a, BDH70, DF96, Geĭ36, GE38, GE39a, GE39b, Geĭ56, Geĭ58, Geĭ67, Geĭ68, Geĭ70a, Geĭ70b, Geĭ77a, Geĭ77b, Geĭ77c, Geĭ87, Geĭ90, Hei62n, Smo92a, Smo92c]. Rutherford [Lak96]. Rüttgers [RDP^{+}95].

S [Ano35a, Ano08, Ben30, Ben34, Ben35, Büh95, Cal14, Clu37, Deh36, Deh37, DS43, FS93, HVTK^{+}97, Hei27b, Hei53b, Kra86a, Mat86a, Mos42, Rec03, Str62b, Str62c, Wei85, Wil36a, Wil36b, Ano29, Hei58l, Sal19]. S. [Hei27b]. S.J. [Bab02]. Saale [HH14, Wit83]. sabotaged [Tel01]. Saboteur [Bro93]. sabotiert [Tel01]. saggio [Cap01]. Said [Hei49f, Hol93b]. Sallhofer [Lak96]. Sam [vC16, vCH18]. Same [FR13a, Büh98a]. samenstelling [Hei39c, Hei84j]. Samuel [Kae47]. Sanctifies [Hol84]. Sander [Ano06]. sanfte [RDP^{+}95]. Savant [Bro93]. savants [Luc16]. saved [Sim07]. savoir [Bar04]. says [Tiw14]. scalar [Tem91]. Scale [Hei44c, HW48, Hei89-33, KW76]. scaling [Ano11, NKD^{+}11]. Scan [Yam02a]. scattered [Ols63]. Scattering [Alb73, Hei33c, Hei84-41, KH25, Wil84, BT78]. Schaaf [Hof02]. Schäfke [SPS^{+}64]. Scharf [Wit83]. Schatten [Läs13]. Schauer [Hei36g, Hei36j, Hei39j, Hei43c, BDR84k, Hei43a]. scheitern [Kra87]. Schichten [DDH42a, DDH89a]. Schichtenanordnungen [BFH^{+}42a, DDH41a, DDH41b, DDH89f, DDH89g, Hei41a, HBF^{+}42, Hei89i, HBF^{+}89a, HBF^{+}89b]. Schichtenkugel [BFH^{+}42b, HBF^{+}89c]. Schicksal [ARK^{+}99, Jun56, Jun64, Jun72, Jun90]. Schicksal/Kärger [ARK^{+}99]. Schiemann [Paw08]. Schimmel [BSP^{+}70]. Schlitten [RDP^{+}95]. Schlußwort [Hei50g, Hei53q]. Schmidt [HMSH78]. Schmidt-Rohr [EHR^{+}96]. Schnellere [URWB95]. Schocken [Jor78]. scholar [Hei68f]. Scholars [Ano34, Hol93b]. scholarship [Wei46, Wei99]. Schöne [Hei79e]. Schönen [BDR85b, Hei71d, Smo93]. Schönheit [Hah10, HHvB10]. Schools [Hei89-30]. Schöpfer [Ful15]. Schreckens [Hei89j]. Schriften [BDR85a, BDR85b, BDR86]. Schritte [Rec02b, Hei71l, Hei73g, Hei76g, Hei77h, Hei84-62, Hei89-29]. Schröder [Tre01]. Schrödinger [Adl03, Ano35b, Ano51, ARK^{+}99, Ben34, Deh36, Haa34, HF46, Hei61d, Hei63e, LW34, Lak96, Sal01, Wil36b, Ano35b, BC09, BD92, Cla23, Dir64, DD13, FR13a, Fri97, Fri98, Ful15, Haa34, Han79, Han62, Hei63e, JD06, KJ06, Meh87, Mil97, Per64, Plo10, Sal01, Shi83, Shi86, ST65, dG14]. Schrödingers [KPMN20, RFS^{+}91]. Schroeder [vRFS^{+}00]. Schroedinger [Ano33]. Schroeer [Sei90a]. Schule [Hei89-30]. Schüler [KPMN20, RKW05]. Schumacher [HEB^{+}80]. Schuster [Lib68]. Schwankungserscheinungen [Hei26d]. Schwarzen [Bes04, RDP^{+}95]. schwere [Hei38a, Hei84i]. schwerem [DDH40b, DDH89e]. schweren [vW41b]. Schwerpunkte [RDP^{+}95]. schwersten [URWB95]. schwieriger [Luc82]. Schwierigkeiten [Hei93c]. Schwinger [CM02]. Science [AG02, Ano36, Ano94, Bab02, Ber01, Bor53, Bri29, CS94, Car02a, Car02b, Cas91, Cas92a, Cas11, CD94, CW85, Cre10, Del13, Dje07, Dre92, Dur00, Fay07, Fra94, GGK02, Gar91, Gim90, Gre10, Hab69, Hei52m, Hei53q, Hei73h, Hei74l, Hei74m, Hei75b, Hei76c, Hei76i, Hei83j, HGW^{+}89a, Hei08d, Hen10, IEE06, Jon08, Jor78, Kae31, Klo96, Kra86a, Kra07, Lac02b, Lal15, Lan08, LMOS15, Mac58, MGG^{+}95, Mat86b, Mer38, Mor47, MTV22, Nic90, Par06, Phi92, Rec10, RMRP09, RW94, Sch11c, Smo93, Sta11b, Sta37, Swe93, Tre01, Tur06, UM92, Wal08, WMG^{+}92, Wüt16, Ano02b, Bal13, Bet91, BDR89b, BR87, Car99b, Car10, Cas92c, Cas93c, Cor03, DR92, Fav05, FW07, FR04, FS06, FR98, Gou47f, Gra00, Gre05, Hei34i]. science [Hei34f, Hei34g, Hei34h, Hei35f, Hei36e, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei48e, Hei48d, Hei49l, Hei50d, Hei51d, Hei52k, Hei53h, Hei58s, Hei59c, Hei59i, Hei61c, Hei62l, Hei65p, Hei70b, Hei70d, Hei71c, Hei72i, Hei72l, Hei73j, Hei74a, Hei74h, Hei74i, Hei75m, Hei77i, Hei82c, Hei84-63, Hei00b, Hei07b, Hol86, Hol98, Hol05, Inf33, Inf34, Lac02a, Lig05, Lin07, Lin08, Oor09, Pic08, PFH20, Pop16b, Rot93, SW05, Sch96a, Wal95, BDR89b, GHW36, Han62, Hei34j, Hei46e, Hei49h, Hei61e, Hei72j, Hei73h, Kae31, Rür79, Som36, Fer07, Jor78, Rud58, For09, Hab83]. Science/Goetzberger/Voß/Knobloch [MGG^{+}95]. Sciences [CS94, Hei31g, Hei79a, Str62a, Hal09, Hee95, Hei53r, Hei71d, Men84, Hei86c, RS11a, Ome79]. Scientific [Ano29, BDR84b, BDR84e, Bro86, Bro95, CS94, Car97, Eck90, Eck12a, Eis00, Gin75, GW96, Hei72d, Hei73b, Hei74n, Hen86, Jon78, Lar30, LS93b, Mil10, Rec84, Rod19, UM92, WG96, vMHW85, vM93, vM96, vM99, vM01, vM05b, Ano66, BDR84a, BDR85b, BDR89a, BDR93, Haw11, Hei42d, Hei43f, Hei48l, Hei52j, HR53, Hei55a, Hei56b, Hei64e, Hei67b, Hei69k, Hei69l, Hei71n, Hei74o, Lud01, Mac93, Mat87, Ols63, Rec10, Sch13, Wig96m, Hei73f, HvMW79, PHvMW01, PvM93, Rec86]. scientifica [Lud01]. Scientist [CS94, GA02, Gou47a, Gou49, Hei70e, Jor78, LH93, Ols63, Car03, Eck01, KF01, Pei92b, Sta36, Tel71, CS94]. Scientists [Acz01, Ano12, Bad95, Ber92, Ber01, BS47, Bey77, Bur62, DG99, Gol95, Gou47b, Hol93b, Jun58, Kae31, MP01, Sei90a, vL48b, BBF^{+}57b, Cas14b, Cor03, Erm89, Hal09, Hei57e, Hei68a, HW11, Jun56, Jun64, Jun72, Jun82, Jun90, Log96, Luc16, OS98, Str11, Bey82, Jor78]. scienza [Pol00]. scienzato [GA02]. scienziati [Cas14b]. Scope [Lic16, Mil12, RS11b]. scoperta [Hei60e]. SCRAPS [Hei65l]. Scribbles [Luc04]. Search [Dje07, Sim00, vRFS^{+}00, Kra95, Mes04]. Sechs [DS43]. Second [Hei53x, Hei93b, Dob77, DRB86, Sim07]. second-neighbor [Dob77]. second-order [DRB86]. Secret [Ber93, Ces18, Hol93a, Kar05c, LH93, Pei93b, Swe93, Wal04, Wil94, Ber96, Cas93a, FB00, Gou48b, Hof92, Kar07, Nag16, Pow93a, Pow93b, Pow94, Pow00]. secrète [Kar07]. section [Fis41]. sections [DK78]. Sector [Nel72]. Security [MWL^{+}11]. Seems [GLR14]. seen [RKW05, Roz67]. Segrè [Wal08]. segreti [Ces18]. sehr [Hei36d, Hei37b, Hei38c, Hei38b, Hei84l]. sei [Hei67g]. sein [Ano02a, vWL93]. seine [Eic53, Hei38a, Hei84i, KPMN20, MGG^{+}95]. seinem [GHH02, Hei43r, Hei44b, Hei05b, KRW05]. seiner [RKW05, RDP^{+}95, SCH^{+}78, Tre78]. seines [Sch82]. Seiten [Ben30, DS43, Eis00, Hei31g, Wit83]. Selbstbiographie [Hei82b]. Selbstenergie [Hei30e]. Selbstreflexionen [Wal93]. selbstvergessene [Fis02]. Selbstzeugnissen [Her76c, Her84, Her01]. Selected [CS79, Sch58, Sch03, Sei90a, Yan13a, Bor63a, Hei87a]. Selection [Hei75b]. Selections [CS94]. Self [Per88, Gra00, Hei30e, Her76c, Her84, Her01, Wal93]. self-deception [Gra00]. Self-dual [Per88]. self-energy [Hei30e]. Self-reflections [Wal93]. Sellerio [Pol00]. Sellevag [vRFS^{+}00]. seltsamen [DH61a, Hei61h, Hei61i]. Semiclassical [KtL79, GKZL96]. Semiconductor [ARK^{+}99, SPS^{+}64]. semidirect [Low97]. Semiempirical [OD03]. semigroups [WB03]. Semilinear [Maa05]. Seminar [Hei52f]. semisimple [Scu75]. Sense [Hei79a, Rov21, Ome79]. sensitivity [BM11]. Separated [NBDG40]. separation [HJ43]. Sept [Hei89-33]. September [Har43, Meh73, Aas02c, Hei89-33]. septième [CCJ^{+}34]. Serber [McC24]. Serie [Pol00]. Series [BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93, Bro86, Kra86a, Mat87, BDR84e]. Service [Ces18]. Serving [Bal13, Bal14]. servizi [Ces18]. Set [Ano34]. Seth [Sta11a]. Sets [RS94]. settling [Enz83]. Seventh [CCJ^{+}34, Far01]. Sexl [SCH^{+}78]. sfondo [Hei99, Pol00]. shadow [Läs13]. shape [HvW48]. Share [Ano33]. Sharman [SPS^{+}64]. sharpness [SH22b]. Shellard [MGG^{+}95]. Shepherd [Dje07]. Shepherd-Barr [Dje07]. Shield [Sei90a]. Shifts [BSP^{+}70]. shock [Hei52g]. Shockley [Dje07]. shook [Gam85, Haw11]. Short [Ano33, Unr72, Hei61h, Hei61i, Hei67p, Hei89u]. Short-Story [Ano33]. Should [vW57, Got02b, Hei89-40]. Shower [DeG74, Hei36g, Hei36j].

Showers [Hei39g, Hei43a, Hei49i, Hei52q, Hei53s, Hei84y, BDR84k, Hei37a, Hei39j, Hei43c]. showing [Ano11, NKD^{+}11]. Shurkin [Dje07]. Si [Hei68h]. sideband [SAR^{+}09]. Sides [FR13a]. Significance [Hei69j, Hei53r]. Silence [Ano93b]. Silizium [RDP^{+}95]. similar [BDR84j]. similitudes [GS12]. Simon [CS94, Lib68]. Simple [Hei84c, AFW94, Hei32b, Yeh70]. simply [DDI03]. simulation [HN99, sMDN^{+}11]. Simulations [HVTK^{+}97]. Simulations/Poenaru [HVTK^{+}97]. since [Meh75]. Singh [Tiw14]. Single [Hür22, HJ43]. singular [CGM^{+}71]. singularities [NRS71]. Singularity [DKLN92, Wre75]. Sir [Adl03, CS94, Mar94]. sisteme [Geĭ67]. site [GbK77, KSGB^{+}83]. Situation [Hei29c, Hei63i, Hei65n, Hei68g, Hei84-34, Hei84-38, Hei93a, Sim07, Hei84-27]. Sitzung [Hei41b]. Six [Sei90a]. Sixty [FR13b, Mil90]. Sixty-two [Mil90]. size [Hei53i]. skæbne [Jun57]. Sketch [CK95, Ano61b]. Skizze [Ano61b]. Skurrile [Arr06]. Slept [Bro97, Cas97, Bre97]. slit [Wis98]. slow [Fis41]. Small [Mac85, HN99]. Smallest [Ano43, Fis10a, Fis12a, Hei53o, Hei71c, Kra95]. Smith [Sei90a]. Smoking [Dah22]. so-called [Hei89j]. sobre [Hei96c, Spi98]. Social [Mer38, Hee95]. Socialism [Hay90, Har37a, Har37b, Har81, Hen96, HH11, HW04, RW94, Wal89a, Wal89b, Wal92]. Socialist [CS94]. Society [BDR89b, GGK02, Hei30a, Hei31g, Hei41b, Hei76c, Rür79, Hei75m, Tel71, Cla23, HW04, Hof05, HW07, HW12b, Mai07, vBL96b]. Sociology [UM92]. soderzhanii [Geĭ77b]. Software [ARK^{+}99, EHR^{+}96, GRE^{+}01, vRFS^{+}00]. sogenannte [Hei89j]. Sokalratic [BSMB99]. sole [Jun57]. Solid [EHR^{+}96, RFS^{+}91]. Solid-State [EHR^{+}96]. Solids [HEB^{+}80]. Soliton [HVTK^{+}97]. solitude [Wal91]. soll [Hei89-40]. sollte [Got02b]. Solution [Bra97, BN08, LB72, Nel72, Wan01, Bes98, CGM^{+}71, DGS99, Hei50a, LFK99, dB58]. Solutions [HK00, Hei22b, Hei84w, IM95, JZ11, KS92, Mie80, Mie81, NRS71, OS07, SMMM87]. Solvability [Kor05]. Solvay [BV09, CCJ^{+}34, Far01, Meh75, Sta11a, CCJ^{+}34]. Some [CZ81, Hei74p, Hei84-43, Ome79, Yeh70, Hei27b, Jes98, Ols63, CL05]. Sommer [Hei51k, Hei85d]. Sommerfeld [Hei39a, Hei43g, Hei43r, Hei48a, Hei51c, Lak96, LSK^{+}76, Sta11a, Ano84b, DHYY65, Hei43r, Hei51b, Hei51c, Hei68j, Hei69j, Hei86m, PS69, Set10, vM02]. Sommerfeld-Ausstellung [Ano84b]. Sommerfeld-Gesamtausgabe [Hei68j]. Sommerfeld/Hermann [LSK^{+}76]. Sommerfelds [Hei86m]. Sommerfeldschen [vM02]. Sonnen [Jun56, Jun64, Jun72, Jun90]. Sonnenenergie [MGG^{+}95]. sono [Arg09]. sootnoshenij [Geĭ77a]. Sorge [Hei48f]. sostojanie [Geĭ56]. Soul [Fer07, For09, Wal08, Bal13, Bal14, Lin07, Lin08]. Sources [Bar04, KHFA67, vdW67, vdW68, Hen96, HH11]. sous [CCJ^{+}34]. Soviet [Jor78]. Sovremennoe [Geĭ56]. Space [Dah02b, Gre62, Hei84u, Hei84-36, AS03, CDvW75, CDvW77, CDvW79, Dah02a, GV86, Hei56m, Hei56o, Hei58w, Hei72h, Mie80, Sch02a, Sch02b, Sch96b, Tem91]. space-time [Mie80]. spacelike [Mie80]. spaces [MMSV98, WB03]. Spaltningen [Fav05].

Spaltung [Hah43]. Spanish [FR08, WdC02, Bla86, Bal07, FR04, HBS96, SHD91, Spi98, W^{+}87]. späten [Hei53b]. Spatschek [RFS^{+}91]. Speak [Bur62]. speaks [Dje07]. specific [KH47]. spectra [Hei26e, Hei27b]. Spectral [KBWJ96, DMdA06, Hei39a]. spectroscopic [Mas04]. Spectroscopy [HVTK^{+}97, LBS^{+}04, URWB95, WMG^{+}92, RFS^{+}91]. Spectroscopy/Callaghan [WMG^{+}92]. Spectroscopy/Finkelstein [HVTK^{+}97]. spectrum [DK78, Gau14, JKS13, vW48]. Speech [Hei72f, Hei69b, Hei89u, Hei58k, Hei58l, Hei72m, Hei86b, Hei89b, Hei89a, Hei89q, Hei89-28, Hei89-27]. Speeches [BDR86]. Speicherchip [Ano84a]. Speicherringe [Hei72f]. Spektra [Hei26e]. Spektrallinien [Hei39a]. Spektren [Hei27b]. Spektrum [Büh95, Rec03, vW48]. Spencer [Sei90a]. speziellen [SPS^{+}64]. spezifische [KH47]. spezifischen [URWB95]. Sphere [BFH^{+}42b, Del13, Set11, Car10, Isa82, Cas11, Sta11b, Hen10, Lal15]. Spherical [Gel84, Gre80]. spherically [DPL94, DPL96]. Spiegel [Hei67g, Hei75m, RKW05, Hei67g, Hei67j]. Spiegel-Gesprach [Hei67g]. Spiess [EHR^{+}96]. Spin [BHX96, CL05, Eps71, Fal64, Fis64, KBWJ96, KA71, LB72, LC71, ML71, Mor71, Mor72b, SK73, Sor02, Zel87, BT96, BG96, Bec02, BPP11, DKLN92, DM99, DC87, DL08, Dur00, Flo97, Fri97, Fri98, JKS13, LFK99, Liu72, sMDN^{+}11, Man91, MC92, Mor72c, OD03, PL91, PDL92, SL70, Tru72]. Spin- [BHX96, JKS13, Man91]. spin-12 [Flo97]. Spin-density [CL05]. spin-Heisenberg [Bec02]. Spin-Pair [Mor71, Mor72b, Mor72c]. Spin-Wave [LC71]. Spinor [DH66, DH84a, Hei60h, Hei63d, Hei65k, Hei68g, Hei75c, Hei84q, Hei84r, Hei84-30, Hei84-36, Hei84-34, Hei84-32, DHYY65, Hei58w, Hei61b, Hei62g, Hei84k, Mie81]. Spinortheorie [Hei61b, Hei63d, Hei84k, Hei84r]. SPINSGA [Flo97]. spiral [HvW48]. Spiralnebel [HvW48]. splitting [Hah43]. Spookiest [Acz01]. Sprache [Eck12a, Eck12b, Hei69o, Hen11, Hof11, Jac11a, Jac11b, Rec09, Rec10, Hei60i]. Sprengstoff [Mül40]. Springer [Cru23, Eis00, Hen86, Hen11, Kra86a]. Springer-Verlag [Kra86a]. Sproull [SPS^{+}64]. spumed [Hei48i]. Spur [Hei10a]. Spurion [Hei63j, DH65a]. Spurion’-Theory [Hei63j]. spurned [Kae48a]. Spy [Sei90a]. square [Man91]. SSC-Physiker [URWB95]. Stabilität [Hei23b, Hei24c, Hei24e]. Stability [Hei84x, Hei85a, Hei23b, Hei24e, Hei51g, Hei52i, Hei24c]. Stabilizers [GS12]. Stable [RS94]. Stadt [Hei58k]. Stærkere [Jun57]. Stage [Dje07, Ano02b, Gla00]. staircase [Fis10a, Fis12a]. Stallion [Sei90a]. Stand [Hei43o, Hei53t, Hei53u, Hei55c, Hei61b, Hei65b, Hei84k, Hei84m, Hei63k]. Standing [JTD^{+}98, Pfe69]. standsrates [Hei64b]. Stanley [vRFS^{+}00]. Stannard [Hab83]. Starke [RDP^{+}95]. stärkster [RDP^{+}95]. State [Bra97, EHR^{+}96, Hei43o, Hei84u, Nel72, RFS^{+}91, BT78, BK88, Geĭ56, Hei55c, Hei61b, Hei65b, Hei72h, Hei84k, Hei84m, KSGB^{+}83, Sch02a, Sch02b, Tru72, Wre75, Hei63k]. Statement [Gou47a, Hei89-31]. statements [Hei38i, Hei84-45]. States [BB36, Hei65i, Hei84z, LBS^{+}04, Bec02, CM02, Fri13, Hei50h, Hei56o, Kle78, Ras78, SZ73, SZ74, Sch86]. Stationäre [Hei50h]. Stationary [Hei65i, Hei84z, BB36, Hei50h]. Statistical [Fed18, Hei58q, Sla72, Bac22, Bor61, Bor63b, Hei48n, Hei48j, HF46]. statistics [Hei31c]. Statistiques [Hei31c]. statistischen [Bor61, Bor63b, Hei48n]. Stehle [EHR^{+}96]. Steinmetz [CS94]. stellt [RDP^{+}95]. Stellung [Bor58]. Stellungnahme [Hei65m, Hei89-32]. Stellungnahmen [BDR89b]. step [MRV24]. Steps [Ano36, Rec02b, Geĭ87, Hei71l, Hei73g, Hei76g, Hei77f, Hei77h, Hei84-62, Hei89-29]. Sternstunden [Büh98b]. Stiftung [Ano71, BDR89b, DRF90b, Hei56g, Hei89u, Hei89-28, Hei71n]. Stiftung/Jürgen [DRF90a, DRF90b]. still [Kae48b, NW95]. Stipendiaten [Hei69b]. Stipendium [Ano84b]. Stitch [OSWR86]. stochastic [Tho80]. Stockholm [HSD34a, Wil36b]. storage [Hei72f]. Storck [SCH^{+}78]. Stores [Ree19]. Störfall [Pop16b, RFS^{+}91]. storia [Lud01]. Story [Acz01, Ano33, Jon08, Kra86b, MP01, Mor47, Tut05, Gam85]. Stoßwellenproblem [Hei52g]. Strahlenbelastung [Rec85a]. Strahlenschutzrecht [SPS^{+}64]. Strahlung [BBB^{+}84, Hei34e, Hei39f, Hei39j, Hei39h, Hei43p, Hei43r, Hei53l, HMW^{+}84, Hei84-44, KH25, BDR84g, EH38, EH84, Hei36g, Hei43o, Hei53t, Hei53u]. Strahlungen [Hei51f]. Strahlungsfeld [Hei31j]. Strahlungstheorie [Hei31a]. strained [Sch13]. Strange [DH65b, DH66, DH84a, DH84b, Hei63j, Hei84-33, Log00, DH61b, DH61a, DH65a, HD61, Hei61h, Hei61i]. Strange-particle [DH65b]. Strategic [Hei85a]. Street [Ano68, FW07]. Streifzüge [EHR^{+}96]. Streuung [Hei31i, Hei33c, Hei84-41, KH25]. Strichartz [FMV07]. Strings [MGG^{+}95]. Strobl [HVTK^{+}97]. Strong [EHR^{+}96, Pac88, San81]. stronger [SK09]. strongly [DGS99]. Structure [CCJ^{+}34, GRE^{+}01, Hei84c, Hei84b, Hei96d, Pau34, Pol95, Ros79, BK04, DKLN92, DC87, DHYY65, Hei22c, Hei25d, Hei32f, Hei32g, Hei32b, Hei33b, Hei34c, Hei35c, Hei39a, Hei65g, Hei65h, Hei66g, LH24, Mor72a, Tem91, CCJ^{+}34, Hei75d]. structures [CDvW75, CDvW77, CDvW79, Hei70b]. Struggle [Fer07, For09, Wal08, Bal13, Bal14, Hei35b, Lin07, Lin08]. Struktur [Hei35c]. Stück [Fra05]. Student [Pei93a]. students [BJN01]. Studien [KP07, vBL96a, vBL96b]. Studienförderpreis [DRF90a, DRF90b]. Studiengang [DRF90a, DRF90b]. Studies [Ano29, CS94, Nic90, SS16, UM92, Hei37f, HW84a, KP07, vWL93, vBL96b, BFH^{+}42a, HBF^{+}89b]. Studium [vWL93]. Study [BHX96, Car99a, GORS09, Lan02, Log99, Neu99, Wal99, dR99, Arg09, Fis10a, Fis12a, MC92, Ras78, Ros98]. Stufe [LH24]. stuff [Haw11]. Stuttgart [Hei53b, Kra86a, Str62b, Str62c, SPS^{+}64, AH93]. Style [Cas79b]. sub [JZ11]. sub-Laplacians [JZ11]. Sube [GRE^{+}01]. Subelliptic [Wan01, GV85]. subgroups [Scu75]. Subsemigroups [WB03]. Subspaces [RS94, GS12]. substances [RCE30]. Success [Hol84]. Such [Hen00]. Suche [vRFS^{+}00]. sulla [Hei28c, HSBA60]. Suman [Sta11a]. Summary [Hei52n]. Summer [Hei51k, Hei85d]. Sundberg [Wig96j]. Sündermann [HEB^{+}80]. Suns [Jun58, Jun64, Jun72, Jun90, Jun56, Jun57, Jun82, Lau58a]. Super [Hei68b]. Super-PS [Hei68b, Hei68b]. superalgebras [BDH88]. superconducting [HvL47]. Superconductive [Cop81]. Superconductivity [Hei29c, Hei46f, BDR84l, Hei47j, Hei47k, Hei47l, Hei48k]. Superconductor [Hei47c, EMGM94, KH47]. superconductors [Hei48c]. supergenerator [Tem91]. supermagnetic [GWW^{+}09]. supersymmetries [GH10]. Supersymmetry [Shi02]. Supplement [Hei89-34, Fis02]. supplemented [Hof23]. Support [TW10]. Supported [Cho02]. supraleitendem [DRF90a, DRF90b, HvL47]. Supraleiter [KH47, Hei47c, Hei48c]. Supraleitung [BDR84l, Hei46f, Hei47j, Hei47k, Hei47l, Hei48k]. Surface [BG96, WMG^{+}92, Dob77, Isa82]. Surface-embeddability [BG96]. Surfaces [OSWR86, TR09]. Surfaces/Paul [OSWR86]. Survey [Hei65n, Hei84-38, GHH02, PC77]. Survived [FTF01]. Surviving [Mac93]. Swastika [Luc11, Mac93]. Swedish [Sch11b, Sch15]. switching [HN99]. Switzerland [Cru23]. Sword [Sei90a]. Sylves [Sei90a]. symbols [KF07]. Symmetric [KS92, Bec02, DPL94, DPL96, Flo97, JD06, PL91]. Symmetric-group [KS92]. Symmetrie [Hei68d]. Symmetries [Sud06, JKS13]. Symmetry [Hei68d, vRFS^{+}00, LFK99]. symplectic [BP75]. Symposium [BW04, Gin75, Hei04, HH14, IEE06, Meh73, Rec81a, And89, Sch00b, Wig96d]. Synchrotron [HEB^{+}80, Hei53i]. Synthesis [HVTK^{+}97, Fin96]. System [DDH42a, DDH89a, Hei66c, Mai07, DDH42c, DDH89c, Geĭ67, PL91]. Systeme [MGG^{+}95]. Systeme/Bunde/Havlin [MGG^{+}95]. Systems [Fis64, Hei89k, RFS^{+}91, BT96, BPS74, BP75, Flo97, Hei26e, Hei66c, KtL79, KGNS10, sMDN^{+}11, Nar93, Pug80, Wan98, RS11a]. Szilard [Wig96i].

T [Lev95, Sei90a]. Tabelle [Fis02, HH14]. table [Fis02]. Tagebuchaufzeichnungen [HHHH11]. Tagebücher [Hah45]. Tagung [Hei30a]. Tagungen [Hei75m]. Tailfingen [Läs13]. tajnych [Kar06]. take [Wei98]. talentierter [KPMN20]. Talk [Ano60, Hei56l]. Talks [BDR84e, Ess71, BDR84b, HHvB10]. Tamm [Hei56e, Hei72b, Hei84e, Hei58b, Hei72b, Hei84e]. Tamper [RMRP09]. tapes [Hof92]. Tartaglia [Gro96]. Taschenbuchausgabe [Fis02]. Taschenwörterbuch [BSP^{+}70]. Task [Gou48a, Hei72f]. Tätigkeit [Eic53]. Tätigkeitsbereich [Hei56g]. taught [CG14]. tausend [Jun56, Jun64, Jun72, Jun90]. Taylor [SPS^{+}64]. teacher [Luc82, Ost11]. teachers [Hei36e]. teaching [Hei52o]. Team [Ano93a, CP93]. Team-B [CP93]. Technical [Hei47g, Hei47h, Hei84-35, Hei34j, Hei39e, Hei40a, Hei46d, Hei63a, Hei89h, Hei89w, Hei89m, Cas79a, Rür79]. technically [Hei67b]. Technik [AH93, BDR89b, DRF90a, DRF90b, EHR^{+}96, Hei61e, Hei70f, WMG^{+}92, SCH^{+}78]. Technik/Maran [WMG^{+}92]. Technikgeschichte [PFH20]. technique [dSL02]. Techniques [HEB^{+}80]. technisch [Hei67b]. technische [Hei63a]. Technischen [Cas79a, Hei46d, Rür79, Hei39e, Hei66b, Hei89w, Hei89m]. technischer [Hei34j, Hei40a, Hei89h]. Technisches [BMG^{+}50]. Technological [MGG^{+}95]. Technology [BDR89b, Hei66b, Kra86a, Kra07, Lan08, MTV22, RW94, Sch11c, Sei90a, Gim90, Hei53k, Hei53g, Hei61e, Hei70f, Hei74h, Hei74m, Hei89d, Hei89c, Mil97, PFH20, Rad83, Rad91]. Teil [ARK^{+}99, BDR85b, BDR89a, BDR93, BSP^{+}70, Bun71, Eis00, Ess71, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06, OSWR86, vM96, vM99, vM01, vM05b]. Teilchen [Fis10a, Fis12a, Hei61h, Hei61i, Hei71c, DH61a, Hei43c]. Telephone [Ano04b]. Tell [Ber95, Hol00a]. Teller [Dür05, Dür05, MF69, Sei90a, Wig96b]. tells [FS06]. Temper [Sch86]. Temperature [Her69, Mor71, Mor72b, KS92, KGNS10, Mor72c]. Ten [Jon08, Str62a]. tendency [Hei70d]. Tendenz [Hei70d]. tension [CBS20]. tensors [Tem91]. Tentative [dB58]. tentatives [Kar07]. tenu [CCJ^{+}34]. teoreticheskoj [Geĭ67]. teoretisk [Ano44a]. teorii [Geĭ36, GE38, GE39a, GE39b, Geĭ56, Geĭ58, Hei62n]. teorij [Geĭ67]. Teorija [Geĭ70b]. Term [Gal01, LH24]. Terms [Yam02a, Kav24, ZG96]. Termstruktur [LH24]. territory [Hei57d]. terror [Hei89j]. Tesla [DRF90a, DRF90b]. Test [Bad95, HS81, Rob80, URWB95, Cap01, Hei42b, Hei89g, HR02b, RW02, Rob82]. Testing [Sei90a, Sei90a]. Tests [Kar05c, Hei89-32, Hei89-33]. Teubner [OSWR86]. textbook [RS11b]. Texte [Hei87a]. Thales [Lak96]. Thank [Hei67g]. theater [FR98]. Their [LSK^{+}76, Hei31c, Hei39b, Läs13, NW95, Pac88, SH22a, vBL96b]. them [Hei67g, Ols63, Pic08]. Themen [RDP^{+}95]. Theodosius [Adl03]. Theology [Los61, Moo93a, Moo93b, Hab83, Oor09]. Theorem [Van41]. theorems [Yeh70]. Theoretic [Hee75, Ras78]. Theoretical [Ano44a, ARK^{+}99, Cus86, Geĭ77b, Hei43g, Hei77b, Meh73, Plo10, Sch86, SM41, Bor49, Geĭ67, Hei25c, Hei26f, Hei27e, Hei27d, Hei34c, Hei38e, Hei42e, Hei43k, Hei66c, Hei69c, Hei77j, Hei83b, Hei84-56, Hei84-42, Hei89n, LL82, EH38, EH84, Hei32e, Hei37f, Hei41e, Hei44c, Hei84b, HW84a, Hei89-33, Hei89-34]. Theoretische [MGG^{+}95, RFS^{+}91, vRFS^{+}00, Hei43g, SM41, EH38, EH84, Hei32e, Hei37f, Hei44c, HW84a, Hei89-33]. theoretischen [Hei38e, Hei42e, Hei43k, Hei66c, Hei89n]. Theoretischer [Hei41e, Hei89-34]. Theorie [BDR84h, DHM^{+}59, DH61a, Hei24a, Hei28e, Hei30f, Hei31l, Hei34b, Hei34a, Hei35a, HE35, Hei36j, Hei37d, Hei38h, Hei39j, Hei43m, Hei43j, Hei43a, Hei44a, Hei47j, Hei47k, Hei47l, Hei48n, Hei51i, Hei52b, Hei55c, Hei58a, Hei58m, Hei61h, Hei61i, Hei62f, Hei65e, Hei67p, Hei70h, Hei84d, Hei84f, Hei84s, Hei84m, Hei84-59, Kar05b, SPS^{+}64, vW35, Hei84-61, Hei83d, Hei53p, Hei57a, Hei67k, Mil58, dB58, BDR84k, Hei50a, Hei51j, Hei51k, Hei52p, Hei53s, Hei84-48, KH47]. Theorie“ [Hei48e, Hei48d]. Theorien [Ano35b, Deh36, Haa34, Hei51l, Hei66c, Hei73e]. Theories [Bok06, Bri65b, Hei51l, Hei65n, Hei84-38, Kae31, Pau41, Rec81a, TV23, Ano35b, Eck33, Geĭ67, Haa34, Hei66c, Hei73e, KED^{+}60]. théoriques [Hei34c, Hei84b]. theorist [Log96]. Theory [Acz01, Ano34, Boh62, Bre23, Bro02a, Bro02b, BM84, CK95, CH84, Cop81, DH32, DH61b, DH66, DH84a, DH84b, GKZL96, GR09, Hak02, HVTK^{+}97, Hei28b, Hei29c, Hei30i, Hei31b, HSD34a, Hei39g, Hei43b, Hei46f, Hei50i, Hei52q, Hei53p, Hei55g, Hei57g, Hei58a, Hei58m, Hei58n, Hei58q, Hei60h, HD61, Hei62f, Hei62h, Hei62k, Hei63d, HB63, Hei63i, Hei63j, Hei64a, Hei65k, Hei65e, Hei66j, Hei67p, Hei68g, Hei75j, Hei75c, Hei84q, Hei84r, Hei84s, Hei84t, Hei84-57, Hei84v, Hei84-30, Hei84-31, Hei84-33, Hei84-36, Hei84-34, Hei84-32, Hei84-39, Hei84y, Hei84-43, Hei84-46, Hei86a, Hei89-35, Hei93b, Hei93c, HP93, Hei93d, Hei96d, HE06, KED^{+}60, LC71, MC97, MS15, Pau93, RFS^{+}91, Rus64, Sno49, Squ67, Sta11a, Wey31b, Wey50]. Theory [WZ83, Wig67, Alb73, AH57, BT78, BV09, Bes04, BDR84l, BDR84h, BH24b, Bro77, CL05, Car71, Cas76, CDvW75, CDvW77, CDvW79, Cli65, DHM^{+}59, DH61a, DHYY65, DH65a, Gam85, Geĭ36, GE38, GE39a, GE39b, Geĭ56, Geĭ58, GZ95, Hei22c, Hei24d, Hei24a, Hei25d, Hei28d, Hei28e, Hei29a, Hei29b, Hei30f, Hei31a, Hei31l, Hei34b, Hei34a, Hei35a, HE35, Hei35e, Hei36j, Hei37d, Hei38f, Hei38h, Hei38i, Hei39j, Hei41d, Hei43m, Hei43j, Hei43h, Hei44a, Hei46b, Hei47f, Hei47j, Hei47k, Hei47l, Hei48e, Hei48d, Hei48n, Hei48j, Hei50f, Hei50j, Hei50h, Hei51m, Hei51a, Hei51h, Hei51i, Hei52b, Hei54f, Hei55d, Hei55c, HKM55, Hei56e, Hei56h, Hei56m, Hei57a, Hei58g, Hei58u, Hei58w, Hei61b, Hei61h]. theory [Hei61i, Hei62a, Hei62g, Hei62n, Hei63c, Hei64d, Hei66d, Hei67d, Hei67k, Hei67m, Hei69o, Hei70h, Hei73d, Hei74i, Hei74k, Hei74p, Hei75l, Hei79g, Hei84d, Hei84-48, Hei84f, Hei84k, Hei84o, HH84, Hei84m, Hei84-59, Hei84-27, Hei84-47, Hei84-61, Hei84-45, Hei86i, Hei87c, Hei90a, Hei90b, Hei91b, Hei03, Hei07a, Hei08b, HW12a, HBE87, Hou30, Ing32, Kar05b, KH47, LBBH69, Maj33, MSUU80, McC24, NRS71, SW72, Set10, Tiw14, Tru72, dB58, vV78, vW35, BDR84k, Geĭ70b, Hei43a, Hei51j, Hei51k, Hei52p, Hei53s, OD03, Hei36c, Hei50a]. Theory/Brückner [RFS^{+}91]. Theory/Marley [KED^{+}60]. there [Ols63, Hei48g]. Thermal [Pug80, vRFS^{+}00, DDH40a, DDH40b, DDH89e, DDH89d, Jau93]. thermischer [DDH40a, DDH40b, DDH89e, DDH89d]. thermo [EHUY92]. thermodynamic [KW02, KKM^{+}09]. Thermodynamical [Hei48k]. Thermodynamics [HF46]. Thermodynamik [HMSH78]. Thermodynamik/Schmidt [HMSH78]. Thermodynamische [Hei48k]. thermonuclear [Hei89w]. thermonuklearen [Hei89w]. Thesis [HH75, Hei75h]. Things [Hen00]. thinkers [Sch13]. Thinking [Plo10, Pfe77]. Third [Hei67a, Jor78, Bed03, Bed04, Bey77, Bey82, BLW19, BDR89b, Cas93c, Hau22, Hei47a, Hei68i, Hof05, HW07, HW11, HW12b, Pop18]. Thirty [Gam85, Wig96l]. Thomas [Ber93, CS94, Cas93a, Hol93a, LH93, Pei93b, Swe93, Wil94, GR08, Sil13]. Thomson [Lak96]. Thought [Cam07b, BCL^{+}69, Hei65a, Hei70b, Hol93b, Hol18, Rob69]. Thoughts [Hei83i, Hei37e, Hei75p]. Thousand [Jun58, Jun64, Jun72, Jun90, Lau58a, Jun56, Jun57, Jun82]. threat [Mas04]. Three [Ano33, HSD34b, O’F92, Hei39b, Sch05]. Tidal [HEB^{+}80]. tier [DDH42a]. Tiger [BM84]. til [Ano44a]. Tillich [Los61]. Time [Cop81, Mac86, Mac87, Mor71, Mor72b, RFS^{+}91, B^{+}61, Büh98a, BPS74, BP75, CGM^{+}71, Cas05, CDvW75, CDvW77, CDvW79, CP84, Hei51i, Hei60b, Hei61e, Hei64c, Hof23, KW76, Mie80, Mor72c, Sei90a]. time-dependent [CGM^{+}71]. Time-Resolved [RFS^{+}91]. Times [Bro97, Cas97, LH93, Lus10, MGG^{+}95, BB05, Bre97, Hei34f, Hei34g, Hei34h, Hei52e, Luc82, Pai91]. Timothy [Mac78b]. Titus [Sei90a]. Today [WF19, Hei69j, Hei72j, Hei58p]. Tode [Hei68f]. todo [Hei96c]. Toeplitz [DMdA06]. together [Hei41b]. Tokio [RDP^{+}95]. Tokio/Auseinandergelebt [RDP^{+}95]. Tokyo [Kra86a]. told [BR64, RRK^{+}64]. tomorrow [Hei70f]. Tonbänder [Hof92, WR92]. tonnes [BBF^{+}45, Hei42b, HBB^{+}89, Hei89g]. Too [Hür22]. topics [Bor49]. Topological [MGG^{+}95]. torno [Bla86]. Toroidal [FO99]. tour [vW41a]. Town [Ano60, Mac85]. track [Hei10a]. Trackdown [Ols63]. Tracking [GP92a, KH19]. trade [D’A03]. trade-off [D’A03]. Tradition [Hei73h, Hei75b, Hei76i, Hei83j, HMSH78, Men84, Hei77i]. Tragedy [Jon08]. Trajectories [Ari23, BH23a]. Trans [Bro72]. transcriptions [Fra93a]. Transcripts [Ber92, Dah22, Hol22, Fra93b, Fra93c, Fra93a, Mar94]. Transfer [Hei47c, Mai07]. Transform [He04, BK88, Gel84]. Transformation [Cam05, Cam07b, Men84, Nel72, LTG^{+}12, Hei39h, Hei84-44]. transformations [Hah43]. transforms [Zak96]. Transistoren [URWB95]. Transistors [SPS^{+}64]. Transistors/Beiträge [SPS^{+}64]. Transition [Hol84, OD03]. transitions [DRB86, DLS76]. transitive [DDI03]. Translated [Bun71, Kra86a, Cru23, Han62, Jor78]. Translation [BJLR17, Hei53b]. Translations [Ano75, DDI03, HDW75]. translator [Sch59]. transmission [DL08]. transmutation [Per39]. Transparent [Hei65o]. Transparente [Hei65o]. Transuranium [CR13, Sim00]. transverse [DKLN92]. Traum [Irv67a]. Travelling [Rod19]. treason [BLW19]. treasury [FF91]. Treatment [KA71, Unr72, Hei89k]. Treaty [Bad95]. Treffen [Ano02a, Kar05a, Paw02a]. Trenneffektes [HJ43]. triangular [MRV24]. triatomic [Hei24c]. Tribal [Fer92]. Tribute [Rec91a, Rec91b]. Trieste [Meh73, BDH70]. trillion [CP93]. Trinity [Sei90a]. triumphs [Smo92a, Smo92d]. True [MP01]. Truth [Klo96, RMRP09, Arg09, BDR85b, Hei73f, Hei74n, Wal95]. try [Hei67a]. trying [Hol00a]. Tschechoslowakischen [Hei89y]. tseloe [Geĭ90]. TU [RDP^{+}95]. Tübingen [Hei69b, Hei69b]. tun [Hei89-40]. Tunazima [Hei31f, Hei31f]. Tunnel [DRF90a, DRF90b]. Turbulence [Hei58q, Hei23b, Hei24e, Hei48b, Hei48n, Hei48j, Hei51g, vW48]. Turbulenz [Hei23b, Hei24e, vW48, Hei48n]. Turbulenzproblem [Hei48b]. turn [BJLR17]. tusind [Jun57]. Tutzing [CDvW79]. Twentieth [Cas97, Hei75k, Lig05, Bro72, TW10]. twentieth-century [Lig05]. Twenty [CW85]. Twenty-five [CW85]. twin [Wal24b, Wis98]. twin-slit [Wis98]. Twisted [Bil03, Abd97, LJ08, LWZ10]. Two [Bes04, Bor49, FR13a, Hei29c, Hei49j, Hei84-40, KH19, KA71, LB72, Mor71, Mor72b, Mor72c, SZ73, SZ74, Dür62, Dür80, Dür06, Fis15, Fra05, Hei26e, Sch13, SL70, Woo77, Ano34, LL77, RS11a, Mil90]. two-dimensional [Woo77]. Two-magnon [SZ73, SZ74]. Two-Spin [KA71, SL70]. Two-Spin- [LB72]. Two-Time [Mor71, Mor72b, Mor72c]. type [DDI03, HR02a, JZ11].

U [BFH^{+}42a, BBF^{+}45, GRE^{+}01, HBB^{+}89, HBF^{+}89b, vW40, Ano08, ARK^{+}99, Haa34, HEB^{+}80, HVTK^{+}97, MGG^{+}95, OSWR86, Otr79, RFS^{+}91, SCH^{+}78, Wil36b, vRFS^{+}00, vW1 , Bec02, BBF^{+}45, DK78, HBB^{+}89, Tur46a, Tur46b, vW40]. U-metall [BFH^{+}42a, HBF^{+}89b]. U. [Hei75g]. u.a [Car97, Hen86, HvMW79, Kra96, PvM93, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05b]. u.a. [BDR85b, BDR86, BDR89b, Eis00]. ua. [PHvMW01]. überall [EHR^{+}96]. Überblick [Hei58a, Hei84d]. Überfahrt [Hei24b]. Übergabe [Hei89b]. Überlegungen [Hei32e]. Übermikroskopie [BMG^{+}50]. Übermikroskopie/Krieger [BMG^{+}50]. übermorgen [Hei70f]. Übersetzung [Hei53b]. Übersicht [Hei43o, Hei51l, Hei53t, Hei53u]. Übersichtsartikel [BDR84b, BDR84e]. Übertragung [Hei47c]. Übungsaufgaben [LSK^{+}76]. Udviklingen [Hei35e]. Ultimate [Mes04]. ultra [Hei37f]. Ultrastrahlung [Hei32h, Hei37f, Hei51l, HW84a]. umbral [DLW08]. Umdenken [Pfe77]. Umdeutung [Hei25c, Wüt16, Kav24]. Umgang [Döp42]. umgearb [Hei31b]. Umkreis [Ess71, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96a, Hei01a, Hei06]. Umsatzzuwächse [RDP^{+}95]. umstritten [DRF90a, DRF90b]. umstritten/Internationale [DRF90a, DRF90b]. Umwelt [RDP^{+}95]. Umwelt- [RDP^{+}95]. unanschauliche [Hei46c]. Unbestimmtheitsrelationen [Hei31d]. Unbounded [Maa05, Kae48b]. unbridled [Hei74c]. Uncertain [Fra94, Hol93a, Yam02a, BB05, Pei92b]. Uncertainties [Ano83b, RFS^{+}91, Swe93, dB82]. Uncertainty [Aga02, Ari23, AG88, Bed09, Bes04, Cas92b, Cas92c, Cop81, DGH^{+}12, EMGM94, Fed18, Gar91, Hab83, Har73, Hei77g, Hen93, HK00, Ing70, Kae31, Lap90, Lin07, Lin08, Loe63, Man11, Meh87, MB11b, Moo93a, Moo93b, NBH97, San81, Sap10, Sch30, Sei05, Smo92d, Sta11a, Tan04, Tut05, Tya63, Van41, WMG^{+}92, WF19, Zan02b, Žup65, dB90, AS03, Ano12, BCL^{+}69, CBS20, Cas09, Chi01, CG14, DJR93, DD13, FR04, FR08, Hei31d, Jad14, Jau93, Jes98, Lap91, LM29, Lic16, Mar08, Mil90, NAZ02, Oza02, Oza03, Phi92, Rob69, SAR^{+}09, SK09, Tiw14, Tri02, Wil99, Wit12, Zan02a, vD15, Ano35c, Ano83b, Ano00, App01, AP22, BKBS93, Bla86, Bro73b, CS94, Com31, Con29, Cow49a, Cow49b, Dar31, Dre92, Fer07]. Uncertainty [Fli32, Fli33, Fra94, Gam58, Hei75f, JHB99, Lar30, Mar31, McC94, New35, Noy31, OW10, PC77, Rot93, Sch94, Smo92a, SR49, STCW94, Swe93, For09, Kea09, Ano94, Cas91, CD94, Gar91, Phi92]. Unclassified [DW52]. Uncovered [Hür22]. underdetermination [Mas04]. undergraduate [DJR93, RS11b]. undergraduates [Mil12]. undergrowth [Gra00]. Underlie [Sei05]. Understand [Wal22]. Understanding [AMS04, Hei46e, Hei52m, Hei69c]. Unendlichen [vRFS^{+}00]. Unendlichen/Rebhan [vRFS^{+}00]. Unerkannte [vRFS^{+}00]. Unfälle [Döp42]. Unfinished [DT18]. ung [Ano35b]. Ungleichheit [LSK^{+}76]. unificación [SHD91]. Unification [SHD91]. Unified [Hei62h, Hei64a, Hei65n, Hei84-57, Hei84v, Hei84-38, Hei84-43, Hei93b, Hei93d, Rec81a, Wig67, Hei62n, Hei63c, Hei66d, Hei67d, Hei67m, Hei74p, Hei84o]. union [Hei52c]. unitaire [Hei67k]. unitary [Hei67k, Low97]. United [Sch86]. unity [Hei42d, Hei49c, Hei52j, Hei53f, Hei69e, Lan72]. universal [Hei38h, Hei69k, Hei69l]. Universe [Bar07, Fac82, Kae31, LSK^{+}76, Pol95]. Universe/Grau [LSK^{+}76]. universelle [Hei38h]. universeller [Hei69k, Hei69l]. Universität [AH93, Hei89-27, RDP^{+}95, Rür79, SM41, SPS^{+}64]. Universität/Kommentar [SPS^{+}64]. Universitets [Ano44a]. universities [Har37a, Har37b, Har81, Hei68a, Rin61]. University [Bok10, Cas79a, Cas11, Hay90, Jor78, Kra86a, Lan02, Lan08, Neu99, Par06, Rür79, Sta11b, SM41, Tur06, Hei72i, Gou47a, Hei36e, Hei89-27]. Unlikely [Acz01]. Unmechanischer [Ser77]. unpolitical [Hei80, Hei83a, Hei91a]. Unpolitischen [Hei80, Hei83a, Hei91a]. Unpublished [O’F92, HHHH11]. Unravelling [FR13b]. Unschärfe [Bes04]. Unschärfeprinzip [Sch30]. Unschärferelation [Hei75f]. unserem [Hei69k, Hei69l]. unserer [B^{+}61, Hei43o, Hei61e, Hei65b, SPS^{+}64]. Untergang [EHR^{+}96]. Untergang/Verhulst [EHR^{+}96]. Untersuchungen [Hei37f, HW84a, BFH^{+}42a, Hei59g, HBF^{+}89b]. Untold [Kra86b]. unusual [MF69]. unveröffentlichten [HHHH11]. unzuständig [Hei50e]. upgrade [RDP^{+}95]. upon [Hei64c]. Uran [DDH89a, Fis41, Mül40, Ano42, DDH42a]. Uran-Metall [DDH89a, DDH42a]. Uranbombe [Hei93f, WR92]. Uranbrenners [HW48, HW84b]. Uranisotop [vW41b]. Uranium [Bet00, For22, HW48, Hei93f, Kan02, KH19, MWL^{+}15, NBDG40, PdK23, Ree19, Ano42, BDJ57, Ber96, BFH^{+}42a, DF96, Döp42, DDH42a, DDH89a, Fis41, Hah43, Hei39e, Hei40a, Hei41g, Hei42c, Hei42e, HW84b, Hei89h, Hei89m, Hei89n, HBF^{+}89b, KF01, Mül40, Per39, Pop18, vW41b, Sim07, Nag16]. Uranmaschine [Hei41g]. Uranmetall [Döp42]. uranovogo [DF96]. Uranproblem [Hei42c]. Uranprojekt [Nag16, Pop18, KPMN20]. Urans [Hah43]. Uranspaltung [BDJ57, GRE^{+}01, Hei40a, Hei89h, Hei89m, Hei89n, KF01, Sim07, Hei39e, Hei42e]. Uranuntersuchungen [vW1 ]. Ursprung [Rad83, Rad91]. usability [Mül40]. Use [HB82b, Sti47, vW41b, HB82a, Hei46d]. uses [HVTK^{+}97]. using [Flo97, JD06, MC92]. utilization [Hei89w].

V [EHR^{+}96, Mat86a, Pye81, BDH70, BDR89b, Geĭ67, Geĭ68, Hei24a]. Valediction [Hei69a]. Valentini [Sta11a]. value [Dun86, DHYY65]. valued [DGS99]. values [Hei61a]. Vanish [Gre62]. Vannevar [Gol92]. variables [BC09, Hei43j, Hei44a]. variation [VL97, Vya99]. Variational [GbK77, KLV80, KSGB^{+}83, KS92]. variety [Hei65p]. Various [DH32]. VB [TR09]. Vechernie [BDH70]. Vectors [Nel72, Hei27b]. Vedas [Tiw14]. Vektoren [BMG^{+}50]. venner [BR64, RRK^{+}64]. Ventures [Cas22]. verabschiedet [Ano84a]. Veranderung [Hei82c]. Verantwortlichkeit [Tel71]. Verantwortung [GRE^{+}01, KF01, Wal03]. verb [Haa34]. Verbots [Hei89-32]. verdrängte [Rad83, Rad91]. Verdrängung [Wal03]. Vergleichende [Hei51l]. Verhalten [Hei48c]. Verhältnis [Sch26]. Verhältnisbestimmung [Tre01]. Verhältnisse [Hei93a]. Verhulst [EHR^{+}96]. verification [DJR93, NAZ02]. verknüpften [Hei34e]. Verknüpfung [Ano06]. Verlag [Ano35a, Ano08, Ben30, Ben34, Büh95, Bun71, Clu37, Deh37, DS43, Kra86a, Mos42, Str62b, Str62c, Wil36a, Wil36b]. Verlagsanstalt [Hei53b]. Verlagsbuchhandlung [Cas05]. Verlagsgesellschaft [Deh36, Hei31g]. verlängert [Ano84b]. verlängert/Ausschreibung [Ano84b]. Verleihung [Hei89-27]. verm [Haa34]. Vernera [Ano77a, Smo92a, Smo92c]. verschiedener [HJ43]. verschränktes [Paw02b]. Verschwinden [GHH02]. Version [GRE^{+}01]. Verständigung [Hei46e]. Versuch [BBF^{+}45, Hei42b, HBB^{+}89, Hei89g, Tre01]. Versuche [Hei41a, Hei89i, DDH41a, DDH41b, DDH89f, DDH89g]. Versuches [Hei89e, Hei43e]. versuchte [Hei67a]. versus [CP84, dR99]. Vertrieb [ARK^{+}99]. Verwandlung [Hei39h, Hei84-44]. Verwendbarkeit [Mül40]. Verwendung [vW41b]. Verwundbarkeit [DRF90a, DRF90b]. Very [Hei84-28, Hei38c, Hei38b, Hei55e, Hei83g]. Vexation [Fay07, Par06, Tur06, Hol05]. Vibrations [SPS^{+}64, Jau93]. Victor [Ber91, CS94, Rec86]. Victory [Hol05, Fay07, Par06, Tur06]. videnskab [Fav05]. Vielfacherzeugung [Hei52b]. Vielfachprozessen [Hei49k, Hei49d, Hei84p]. Vienna [Hei41b]. Vietnam [Tut05]. View [Got02a, Hei42d, Hei55b, Hei56f, Hei57c, Hei58e, Hei59a, Hei60a, Hei62b, Hei63b, HG65, Hei66a, Hei68c, Hei70c, HnH72, Hei74h, Hei74d, Hei76e, Hei79d]. Viewed [Dah02b, Dah02a]. viewpoints [Bet91]. views [Cam07a, Hei59k, Hei61g, HBD71]. VII [GbK77]. viii [Kra07, Lan08, GORS09]. világképe [Hei58p]. Vilenkin [MGG^{+}95]. Vincent [Sei90a]. vingt [Hei68h]. violating [BCL^{+}69, Rob69]. Violation [RDM^{+}12b, RDM^{+}12a]. Virasoro [Bil03, FO99, LJ08, LWZ10]. Virasoro-Toroidal [FO99]. virke [BR64, RRK^{+}64]. virus [Irv67c]. Vision [Bri29]. Visionen [HVTK^{+}97]. Visit [Wal20b, Aas02e, Hei89r]. Visual [Hei77j]. viszonya [Sch66]. vita [Arg09]. vitae [Hei86d, Hei86h]. Vogel [EHR^{+}96]. Voice [CS94]. Vol [Bro95, Kra07, Mac58, OSWR86, Pye81, Wig96m]. Völkern [Hei46e]. Volksbildung [HGW^{+}89a]. vollständig [Hof23]. vols [Hen11]. Volume [BDR84c, BDR84d, BDR85b, BDR86, BDR89b, Car97, CDvW79, Eis00, Hei43g, Kra96, PvM93, Rec86, Sta09a, UM92, vMHW85, vM93, vM96, vM99, vM01, vM05b, CDvW77, FK85, HvMW79]. Vom [BMG^{+}50, SPS^{+}64, Hau22, HGW89b, HH14, WR92, vW41a, Rec02a, vM02]. Vor [Hei64b, ABH^{+}55, ABH^{+}87, Ano02a, Hei74c, Hei86c, Hof93, RDP^{+}95]. vor/Wir [RDP^{+}95]. Voraussetzungen [Hei59b]. Vorbereitung [HW48, HW84b]. Vorfeld [Hei57b]. Vorlaufiger [HBF^{+}89c, BFH^{+}42b].

Vorlesung [Hei47f, Hei51k, Hei52p]. Vorlesungen [Hei43g]. Vorsicht [RDP^{+}95]. Vorsitzenden [Hei64b]. Vorstellungen [Hei53o]. Vorstoß [vRFS^{+}00]. vortex [Hei22a, Rec02a]. Vortrag [Hei31g, WR92, Hei89-37]. Vorträge [Ano35a, Ben35, BDR84b, BDR84e, Deh37, DS43, Hei43p, Hei53l, HMW^{+}84, HHvB10, Mos42, SM41, Str62c, Wil36a, Wil36b, HSD34a, Hei43r]. Vortragsreise [vW41a]. Vortragsveranstaltung [Hei89-38]. Vorwort [Hei39i, Hei43s, Hei53x, Hei53v, Hei53w, Hei59h, Hei60j, Hei62m, Hei63l, Hei64f, Hei64g, Hei65q, Hei66h, Hei66i, Hei67n, Hei67o, Hei68j, Hei68k, Hei69m, Hei69n, Hei71m, Hei72n, Hei73i, Hei74q, Hei75t, Hei89-39]. Voß [MGG^{+}95].

W [Ano54, Ano94, Ben30, Ben34, Ben35, Bor53, Bro85, Bro86, CS94, Fra94, Hei51k, Hei53q, Kle54, Kra86a, Kva76, Mat87, Mil58, Ome79, Sch91b, Sei90a, Str62a, Str62b, Str62c, Wil36a, Wil36b]. W. [Bor49, Fel54, Hei51k, Pau31, Paw06, Ros79, TW91]. Wachstum [HMSH78]. Waddington [Adl03]. Waerden [Hei48h]. Waffenschmiede [Wal05, Wal07]. während [Hei73a, vM05a]. Wahrheit [BDR85b, Hei73f]. Wahrscheinlichkeitsaussagen [Hei38i, Hei84-45]. walked [Hol18]. Walker [Hay90, Klo96]. Walther [Hei64h]. Wandel [BMG^{+}50]. wanderer [Fis15]. Wandlungen [Ano35a, Ben35, Clu37, Deh37, KED^{+}60, Mos42, Som36, Str62c, Wil36a, DS43, Hei34i, Hei34f, Hei34g, Hei34h, Hei35f, Hei36h, Hei43t, Hei44d, Hei45b, Hei47i, Hei49l, Hei59i, Hei73j, Str62a]. want [Büh98a]. War [Ber93, Cas17, FTF01, Gou46b, Gou49, Hei81a, Hol93a, Jon78, KH19, Kra86b, LH93, Pei93b, Pow93a, Sim12, Swe93, Tho97, Wil94, Boh64, Boh67, Bro93, Cas93a, Cor03, Hei56b, Hei64c, Hei89-40, Mai07, Pow93b, Pow94, Pow00, vL47, Sal19, Sim07, Tut05]. war-relevant [Mai07]. Wardenburg [Tho97]. Warfare [Fer92]. Wärme [KH47]. warn [ABH^{+}55, ABH^{+}87]. warnen [ABH^{+}55, ABH^{+}87]. Wartime [Dah99, GWY05a, GWY05b, Sch22, Wat01, vL48a, vL48b, SW05, Sch15]. Warum [Pop16a]. Was [Ano93b, Bro93, Gil09, Hei08a, Hol93a, Pei93b, Sei90a, Brü64, Hei48g, Hei48l, Hol93b, Brü64, Hei76h, Hei89-40]. Washington [Cas11]. Wasn’t [Dah22]. Wasser [DDH41b, DDH89e, DDH89g, DDH40b]. Wasserstoffmolekülion [Tel30]. Wastes [SPS^{+}64]. Water [Bør12, Dah99, SPS^{+}64, Wat01, DDH40b, DDH41b, DDH89e, DDH89g]. Water/Taylor [SPS^{+}64]. Watts [WMG^{+}92]. Wave [FMV07, Hei50i, Hei55f, Hei84-46, Hen93, Lan59, LC71, RFS^{+}91, Wei98, dB90, AH57, Cam06, Fed04, HP29, HP30, Hei38i, Hei43h, Hei46b, Hei47f, Hei50a, Hei50h, Hei52g, Hei54f, HKM55, Hei56e, Hei56m, Hei84-29, Hei84-45, Hei10b, Sch27, Tru72, dB58, dB82]. wave-particle [Cam06]. waveform [MRV24]. wavefunction [sMDN^{+}11]. wavefunctions [Ter82]. Wavelets [Tor91]. Wavepacket [CK02b, CK02a]. Waves [JTD^{+}98, PSZM02, SPS^{+}64, Ano35b, Haa34, Hei36b, Liu72]. Waves/Heimendahl [SPS^{+}64]. way [Hei46e]. WE-Heraeus-Stiftung [DRF90a]. WE-Heraeus-Stiftung/Jürgen-Geiger-Preis/Verwundbarkeit [DRF90a]. Weak [RDM^{+}12a, RDM^{+}12b]. Weapon [Hei76b, Hau22, Kar07]. Weapons [For22, GWY05a, ABH^{+}55, ABH^{+}87, Bad95, Bet91, Boh64, Boh67, Bro92, GWY05b, Hei57b, Hei89-32]. Weart [Sei90a]. Weaver [Sei90a]. Weber [WMG^{+}92, Cop81]. Wechselwirkung [Hei56k, Hei84-48]. Weg [Hau22, URWB95]. Wege [HPE^{+}33, Hei43i]. wegzuschaffen [Hei89-40]. Weimar [Bro09]. Weinberg [MGG^{+}95, Wig96j]. Weird [Str11]. Weisbuch [RFS^{+}91]. Weise [GRE^{+}01]. Weiss [MGG^{+}95, Hei48l, Sno49]. ‘Weiße [Sta37]. Weisskopf [CS94, HEB^{+}80, Ano75, Ber91, HDW75, Mat86a, Pye81, Rec86]. weitere [Hof23]. weiterhin [DRF90a, DRF90b]. Weizäcker [Ano02a, Ano02a]. Weizsäcker [Cas14a, GRE^{+}01, Hei72c, Hei02, HH14, KF01, Rec02a, vRFS^{+}00, vS14, Cas15, Hei72c, Hei02, KF01, Rec02a, vWL93, vS14]. Welch-Stipendium [Ano84b]. welcome [Hei69b, Hei89u, Hei64b, Hei89f]. Well [Ros00]. Wellen [Hei56e]. Wellen-felder [Hei56e]. Wellenfelder [HP30, Hei47f, HP29, Hei38i, Hei46b, Hei50h, Hei56m, Hei84-45, Hei43h]. Wellengleichungen [AH57, HKM55, Hei54f]. Wellenmechanik [Hei10b, Hei50a, Sch27]. Welt [Hei89-40, Hei65o, Hei67b]. Weltanschauung [Hei35b]. Weltbild [Hei84-63, Hei84-49]. Weltbildes [HVTK^{+}97, Hei42d]. welten [Fis15]. Welterbe [PFH20]. Weltformel [HHvB10, Hah10]. Weltgeltung [Pfe69]. Weltkrieg [Sim07]. Weltkrieges [Hei89-40]. Weltperspektiven [Hei59e]. Wenn [Büh98a]. Wenske [MGG^{+}95]. went [Ols63]. Wentzel [Hei43h]. Wer [URWB95, Hei48l]. werden [RDP^{+}95]. Were [Sei90a, Hei68h]. Werer [Fra94]. Werk [Büh95, MGG^{+}95, Pfe69, Pfe77, Cas95, Hei55a, Rec03]. Werk/Weinberg [MGG^{+}95]. Werke [BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR85b, BDR86, BDR89a, BDR89b, BDR93, Bro85, Bro86, Kra86a, Mat87, Sch91b, TW91, BDR84e]. Werke/Collected [Kra86a, Sch91b]. Werner [Adl03, Ano29, Ano35a, Ano68, Ano71, Ano94, Ano06, Bar04, Bes04, Bor58, Bro72, Bro85, Bro86, Büh95, Büh98a, Bun71, Cap57a, Cap57b, CS94, Cas91, Cas14a, CDvW77, CD94, Clu37, Deh37, DS43, Dre92, DB71, Dür76, Dür05, Eck12a, Eck12b, EHR^{+}96, Ess71, Ful15, Gar91, GZ92b, GRE^{+}01, Hah10, Han62, Hei54a, Hei80, Hei83a, Hei91a, Hen11, Her76a, Huf71, Jac03, Jac11b, Joh59, Kra86a, Kra07, KF01, Lan08, Lau07, Mar73, MGG^{+}95, Mat86a, Mat87, Mey76, Mos42, Mot85, Pfe77, Phi92, Rec85a, Rec91a, Rec91b, Rec91c, Rec01, Rec02b, Roh85, Rud58, Sch82, Sch05, Sch07, Sch91b, SCH^{+}78, SPS^{+}64, Swe93, Tre01, Tre78, WR92, WMG^{+}92, Wei85, Wit83, vM02, vRFS^{+}00, vW02, vWR04, Aas02e, Alt06, Ano44b, Ano77a]. Werner [Ano77b, Ano20, Ano25, Anoxx, Arg09, Bes04, Bla86, Ble90, BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR84e, BDR85b, BDR86, BDR89a, BDR89b, BDR93, B^{+}61, BR87, Büh98a, BW04, Cap57b, Cap01, Cas74, Cas76, CR84, CB84, Cas92c, Cas93c, Cas95, Cas01, Cas02c, Cas02d, Cas05, Cas15, Ces18, Cir88, Cle06, Cru23, Dör05, DB71, Dür76, D^{+}92, Dür05, Eck01, Eck12a, Erm89, Fav05, Fis02, Fis10b, Fis12b, Fis15, Fue61, Ful15, Gar91, Gem87, GA02, GZ92b, GZ92a, Gou76, Hah06, Hah10, Hei35g, Hei36e, Hei48m, HBD71, Hei74e, Hei80, Hei83a, Hei84a, Hei86k, Hei89b, Hei91a, Hei01b, Hei02, HR02b, Hei04, Hei05b, Hei08a, Hen98, Her76a, Her76b, Her76c, Her84, Her01, Hof11, Hol00b, Hor66, Jac01, Jam76, Jor71, Kan02, Kan03, KW93]. Werner [KRW05, Kra95, KF01, Lac02a, Lac02b, Loc07, Lus10, Mac58, Mas14, Meh87, Mey76, Mil85, MB11b, MP77, O’F92, Otr79, PLS02, Paw08, Pfe77, Pol00, Pop18, Rec81a, Rec81b, Rec84, Rec91a, Rec91b, Rec91c, Rec92, Rec93, Rec96, Rec01, RW02, Rec02b, RKW05, Rec09, Rec10, RH02, Sch82, Sch05, Sch07, Sch59, Sil13, Smo92a, Smo92b, Smo92c, Smo92d, Spi98, Tel76, Wal92, Wal20b, Wal22, Wig67, Wig76, Yam02a, Yam02b, Yam02c, Yan13b, vM02, vWvdW77, vW02, vWR04, Ano08, Cas03, Jac11a, Mac78b, Mat86b, Rec85b, Rec03]. Werner-Heisenberg-Gymnasium [Hei89b, Hei89b]. Werte [Hei61a]. wes [Haa34]. Wesen [BMG^{+}50]. West [Rec85a, Car99b]. West-Berlin [Rec85a]. Western [AFG90]. Westphal [SPS^{+}64]. Weyl [Hei28b, Hei31b, BK88, CM02, CP84, GBLW86, Gel84, JD06, KF07, Mie81, RS94, Tor91, Tru75, VZ09, WB03, Zak96]. wheel [HvL47]. Wheeler [GZ95]. where [Ols63]. which [Gre62]. ‘White [Sta37]. Whitehead [Adl03]. Whittaker [LWZ10]. Who [Ano84b, GLSC01, Hei08a, Lau58a, Ano20, Dem15, Gre05, VB15, Dem15, Hei48l]. Whole [Ess71, BDR85b, Geĭ90, Hei69d, Hei72e, Hei73c, Hei81b, Hei81c, Hei83e, Hei84-50, Hei85c, Hei86e, Hei87b, Hei88, Hei96c, Hei96a, Hei01a, Hei06, Sim07]. wichtig [Hei48l]. Wider [KP07]. Wie [RDP^{+}95, NW95]. Wiederbegrundung [Hei56g]. Wiemers [Ano08, EHR^{+}96]. Wien [Hei41b]. Wieviele [RFS^{+}91]. Wigner [Bri65b, Eck33, Wig96m, Wig96j]. Wild [LMOS15]. wilde [LMOS15]. Wilfried [Tre01]. Wilhelm [Hah40, Hei25a, Hei31g, Hei41a, Hei50b, Hei71b, Hei89i, Läs13, Mai07, MGG^{+}95, Rec81b, Rec96, SW05, Sim07, Wal05, Wal07, vBL96a, vBL96b, Hei25a, Hei50b]. Wilhelm-Gesellschaft [Hei31g]. Wilhelm-Instituts [Hah40]. Wilhelm/Max [vBL96b]. Will [Com31, Mar31, Büh98a]. Willem [Bor58]. Willers [BMG^{+}50]. William [Dje07, Sei90a]. Williams [Sei90a]. Windows [GRE^{+}01]. winners [Läs13]. Wins [Ano29]. Winter [Hei47f, Hei52p]. Wir [RDP^{+}95, Hei67g]. Wirbel [Rec02a]. Wirbelbewegung [Hei22a]. wird [Ano84b, Büh98a, DRF90a, DRF90b, Hei48l, RDP^{+}95, URWB95]. Wirken [Eck12a, Eck12b, Hen11, Rec09, Rec10]. Wirklichkeit [Hei60i, Hei84-52, URWB95, BDR84c, Hei89z]. Wirklichkeitsbegriffs [Hei82c]. Wirkung [Pfe77]. Wirkungsgrades [HJ43]. Wirkungsquantum [Hei45a]. wishes [Rec02a]. Wiśniewski [Hei24a]. Wissenschaft [BDR89b, Cas93c, Hei36e, HA^{+}48, Hei51d, Hei53h, Hei53q, Hei75m, HGW^{+}89a, Hen11, HMSH78, MGG^{+}95, Rec09, Bor58, FS06, GHW36, Hei34j, Hei46e, Hei50d, Hei70b, Hei70d, Hei77i, Rür79, Sta37]. Wissenschaft/Cassidy [MGG^{+}95]. Wissenschaft/Päsler [HMSH78]. Wissenschaften [Hei31g, Hei86c, Rec85a]. Wissenschaft“ [Rec10].

Wissenschaftler [Hei68a, HW11, Bey82]. Wissenschaftlers [GRE^{+}01, Tel71, KF01]. Wissenschaftlers/Dürr [GRE^{+}01]. Wissenschaftliche [BDR84a, BDR84b, BDR89a, BDR93, Bro86, EHR^{+}96, Mat87, Eck12a, Eck12b, Hei43f, Hen11, Rec09, Rec10, BDR84e]. wissenschaftlichen [Hei53d, HR53, Hei64e, Hei71n]. Wissenschaftlicher [ARK^{+}99, Car97, Eis00, GRE^{+}01, HEB^{+}80, Hen86, HvMW79, MGG^{+}95, OSWR86, Pye81, vMHW85, vM93, vM96, vM99, vM01, vM05b, Hei48l, Hei69k, Hei69l, PHvMW01, PvM93, Kra96]. wissenschaftliches [Hei55a]. Wissenschafts [PFH20]. Wissenschafts- [PFH20]. Wissenschaftsakademie [LMOS15]. Wissenschaftsgeschichte [Pop16b]. wissenschaftspolitische [Hei72l]. Wissenschaftsrat [Ano84a, URWB95]. Wissenstransfer [Mai07]. within [GBLW86, Hei57d]. without [Hei57b]. witnesses [Sch01]. Wizard [Jon78]. WKB [Kle78]. Woher [Got02b]. Wohnhäusern [Rec85a]. Wolf [KED^{+}60]. Wolfgang [ARK^{+}99, Car97, Eis00, HEB^{+}80, Kra96, MGG^{+}95, Pye81, Rec86, Bro95, Eis00, Hei59j, Hei59k, Hei61g, Hen86, HvMW79, vMHW85, vM93, vM96, vM99, vM01, vM05b, Car97, Pye81]. won [Ano20]. words [Dje07, Hei53q, Hei63e, Hei75q, Hei69f, Hei86l, Hei86m]. Work [Cas95, Cas02c, Cas02d, Eic50, GW96, Rec10, WG96, Eck12a, Hah40, Hei31f, Hei34b, Hei36e, Hei41g, Hei43f, Hei46d, Hei49b, Hei55a, Hei69j, Hei74o, Hou30, Pfe77, Roz67, Sch05, vL47, Hei64a, Pfe69]. Worked [Wal22]. workers [RRK^{+}64]. Working [KH19, Hei42c]. Works [BDR84a, BDR84b, BDR85a, BDR84c, BDR84d, BDR84e, BDR85b, BDR86, BDR89b, Bro86, Kra86a, Mat87, TW91, Yam02b, Yam02c, BDR89a, BDR93, BR64, Bor63a, Hei87a, PLS02, RRK^{+}64, Wig96m, Bro85, Sch91b]. Workshop [DRF90a, DRF90b, NBH97]. World [Arr06, Cal14, CK95, Edd28, Edd78, Edd07, Edd12, Edd13, Hür22, Mac58, PFH20, RS11a, Sei90a, Sim07, Arg09, Edd00, FF91, Gre05, Hah10, Haw11, Hei31g, Hei42d, Hei59e, Hei65o, Hei67b, Hei74h, Hei84-63, Hei84-49, Hei89-40, HHvB10, Inf34, Lan72, W^{+}87, WdC02, Cas17, KH19, Kra86b, LH93, Sim12, AFG90]. worlds [Fis15]. worldview [Hei58p]. Worte [Hei86l, Hei86m]. WÖrterbuch [BMG^{+}50, MGG^{+}95]. WÖrterbuch/Dingler [BMG^{+}50]. wreath [Hei86m]. wreath-laying [Hei86m]. Write [Ano33]. writer [Gou47a]. Writing [Cas22]. Writings [Kra86a, BDR85a, BDR85b, BDR86, CB84, Cas01, W^{+}87, WdC02, Rec85b, Mat86b]. Wrong [Mil12, RS11a, Fri11]. Wunner [EHR^{+}96]. Wünschen [Rec02a]. wußte [Brü64].

x [Hay90, Hei31i, Hei33c, Hei84-41, SH22b]. X-ray [SH22b]. X-rays [Hei31i, Hei33c, Hei84-41]. XI [BDR84h, Bro72, KSGB^{+}83]. xii [Bok10, Fra94, Sta11b, Wil94]. xiv [Par06, Tur06]. xvi [Cas11, Wat01]. xx [Lan02, Neu99]. xxi [Hen11]. xxiv [Sta11b]. xxix [Hen86]. XXZ [KKM^{+}09].

Yale [Jor78]. Year [Cas15, FR13b, PC77, Hei89y]. Years [And89, ARK^{+}99, CW85, Gla00, HB82b, Jac03, Mla98, PLS02, Rog13, Wig96l, Yam02b, Yam02c, AK15, Boh64, Boh67, CR13, Cas79a, Con62, Gam85, GHH02, HB82a, Hei38g, Hei42a, Hei49a, Hei51a, Hei53e, Hei53b, Hei56g, Kae48b, Mil90, vM05a]. years“ [Hei53b]. yedinoy [Hei62n]. yield [HJ43]. York [Ano54, Ano68, Ano94, Bed09, Bun71, Cas11, Eis00, Fra94, Hay90, Hei53b, Hen86, Hol93a, Kra86a, Lib68, Pei93b, Wil94]. youngest [Hei39d]. youth [Rec10]. Yukawa [Hei50i, Hei62e, Hei84-46].

z [Kar06]. Zagreb [Hei89-27, Hei89-27]. zahlen [vW48]. Zahlzeichen [BMG^{+}50]. Zamechanija [Geĭ36]. Zamechaniya [Hei62n]. Zee [vRFS^{+}00]. Zeeman [Bre23, Hei22c, Hei24d, Hei25d, HJ26, Hen98, SH22a]. Zeeman-Effekte [Hei24d]. Zeemaneffekte [Hei22c, Hei25d, HJ26]. Zeemankomponenten [SH22a]. Zehn [Str62c]. Zehnder [Som36, Som36]. Zeichnen [GRE^{+}01]. Zeilinger [Fis10a, Fis12a]. Zeit [BMG^{+}50, Cas05, Hei60b, HHH03, Kan03, SCH^{+}78, SPS^{+}64, B^{+}61, Hei34i, Hei34f, Hei34g, Hei34h, Hei61e, Luc82, Tre78]. Zeit/Lense [BMG^{+}50]. Zeit/Sproull [SPS^{+}64]. Zeit/Storck [SCH^{+}78]. Zeitalter [Hei66b, Hei69k, Hei69l]. Zeitbegriffs [Hei51i]. zeitgenössische [Hof23]. Zeitschrift [DRF90a, DRF90b]. Zeitzeugen [Sch01, Cas02a]. Zen-Ereignisse [Ano84a]. Zentralverbandes [RDP^{+}95]. Zentrifuge [HJ43]. Zentrifugenanordnungen [HJ43]. Zentrum [DRF90a, DRF90b]. Zerfall [Hei43d]. Zero [Bil03, New35]. Zero-Point [New35]. Zeros [SF71]. Zerstrahlung [Hei52s]. Zertrümmerungsprozesse [Hei32h]. Zhizn [BDH70]. Zhores [Jor78]. ZILSH [OD03]. Zollernburg [Läs13, Läs13]. Zones [Hei76b]. zugleich [Büh98a]. Zukunft [HMSH78]. zukunftige [Hei79c, Hei79b, Hei85b]. Zukunftsauftrag [SPS^{+}64]. Zum [Hei31f, Hei68f, Rec01, vM02, Ano83a, Ano84a, DB71, Fis10a, Fis12a, Fue61, Hei35d, Hei41e, Hei48a, Hei48b, Hei48k, Hei54e, Hei62i, Hei64h, HBD71, Hei72c, Hei86j, Hei89x, Hei89-34, Hei89-37, Jac01, Paw02a, PFH20, Rec02a, Rec02b, RH02, RDP^{+}95, URWB95, Hei31k, Hei36i, Hei43u, Hei52r, Hei71n, Hei86n, Sch30]. Zur [Hei24a, BBH^{+}62, Bor61, Bor63b, EH38, EH84, Fis02, GRE^{+}01, HJ43, Hau22, Hei31a, Hei31f, Hei31g, Hei32e, HPE^{+}33, Hei34b, Hei34a, Hei35a, Hei37f, Hei43i, Hei46d, Hei46e, HW48, Hei52b, Hei53x, Hei56e, Hei59g, Hei64a, Hei70d, Hei72b, Hei75p, Hei84e, Hei84f, HW84b, HW84a, Hei89w, Hei89-32, KP07, KPMN20, KED^{+}60, KF01, LSK^{+}76, Rec02b, Rür79, Sch27, SM41, SPS^{+}64, AH57, BDR84l, BH24b, BJ25, BHJ26, BHJ62, BHJ63, DHM^{+}59, DH61a, Hei22c, Hei25d, Hei28d, Hei28e, HP29, HP30, Hei31l, Hei33d, Hei36j, Hei39j, Hei41g, Hei46f, Hei47j, Hei47k, Hei47l, Hei48n, Hei50j, Hei51m, Hei52s, Hei53y, Hei54f, HKM55, Hei61h, Hei61i, Hei67p, Hei69o, Hei84-47, Hei86o, Hou41, vW35]. Zusammenarbeit [Dür05, Hei56i, URWB95, vM05a]. Zusammenarbeit/Projekt [URWB95]. Zusammenarbeit/Projekt- [URWB95]. Zustände [Hei50h]. Zwang [Ser77]. Zwei [Ano35a, Ben35, Bes04, Wil36a, Dür62, Dür80, Fis15, Fra05, Hei26e, LL77]. Zweige [Hei39d]. Zweiten [Sim07]. zwischen [Fis15, Hei53e, Hei75m, HW07].

References

Aaserud:2000:CPP

[Aas00] Finn Aaserud. “Copenhagen” play portrays Bohr and Heisenberg. AIP History Newsletter, 32(1):1–2, ???? 2000. CODEN ???? ISSN 1048-1338.

Aaserud:2002:BHM

[Aas02a] Finn Aaserud. The Bohr–Heisenberg meeting from a distance.

In Anonymous [Ano02b]. URL http://web.gc.cuny.edu/ashp/

nml/artsci/aaserud.htm. Three videocassetes recorded at the Graduate Center of the City University of New York, Baird Auditorium, March 2, 2002. Apparently not published in print.

Aaserud:2002:NBA

[Aas02b] Finn Aaserud. Niels Bohr Archive: Release of documents relating to 1941 Bohr–Heisenberg meeting. Web document, February 6,

2002. URL http://www.nba.nbi.dk/release.html.

Aaserud:2002:RRP

[Aas02c] Finn Aaserud. Recently released papers on the Bohr–Heisenberg meeting in Copenhagen in September 1941. Web document, March

8, 2002. URL http://web.gc.cuny.edu/sciart/copenhagen/
bhpapers.htm.

Aaserud:2002:RDR

[Aas02d] Finn Aaserud. Release of documents relating to 1941 Bohr– Heisenberg meeting. Naturens Verden, 84(8–9):1–40, ???? 2002. CODEN NAVDAL. ISSN 0028-0895. URL http:/

/www.nba.nbi.dk/papers/introduction.htm;   http://www.
nba.nbi.dk/release.html.

Aaserud:2002:WHB

[Aas02e] Finn Aaserud. Werner Heisenberg’s besøk hos Niels Bohr i 1941: Drama og dokumentasjon. (Norwegian) [Werner Heisenberg’s visit with Niels Bohr in 1941: drama and documentation]. Forskningspolitikk, 3(??):14–15, 17, ???? 2002. CODEN ???? ISSN 0333-0273.

Abdesselam:1997:THA

[Abd97] Boucif Abdesselam. The twisted Heisenberg algebra U_{h,w}(H(4)). Journal of Mathematical Physics, 38(12):6045–6060, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1955:ANW

[ABH^{+}55] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. Keesing’s Archiv der Gegenwart, 25(??):5251H–??, ???? 1955. The original typescript, entitled “Mainauer Kundgebung” [Mainau rally], signed by authors, is reproduced in [ABH^{+}87].

Adler:1987:ANW

[ABH^{+}87] Kurt Adler, Max Born, W. Heisenberg, et al. Achtzehn Nobelpreisträger warnen vor Atomwaffen. (German) [Eighteen Nobel laureates warn against nuclear weapons]. In ????, editor, 30 Jahre Göttinger Erklärung. Nachdenken über die Rolle des Wissenschaftlers in der Gesellschaft. (German) [30 years Göttingen declaration. Reflection on the role of the scientist in the company], page 42. ????, Munich, West Germany, 1987. Reproduction of original signed typescript of [ABH^{+}55].

Aczel:2001:EGM

[Acz01] Amir D. Aczel. Entanglement: The Greatest Mystery in Physics: The Unlikely Story of How Scientists, Mathematicians, and Philosphers Proved Einstein’s Spookiest Theory. Four Walls Eight Windows, New York, NY, USA, 2001. ISBN 1-56858-232-3. xviii + 284 pp. LCCN QC174.12 .A29 2001. URL http://www.loc.gov/

catdir/enhancements/fy0831/2002069338-d.html.

Adams:1972:FGN

[Ada72] J. B. Adams, F.R.S. Four generations of nuclear physicists. Notes and Records of the Royal Society of London, 27(1):75–94, August 1972. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Adler:2003:HPM

[Adl03] Mortimer J. Adler, editor. Henri Poincaré, Max Planck, Alfred North Whitehead, Albert Einstein, Sir Arthur Eddington, Niels Bohr, G. H. Hardy, Werner Heisenberg, Erwin Schrödinger, Theodosius Dobzhansky, C. H. Waddington, volume 56 of Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, 2003. ISBN 0-85229-531-6. xvi + 749 pp. LCCN AC1 .G72 1990.

Adler:1990:GBW

[AFG90] Mortimer Jerome Adler, Clifton Fadiman, and Philip W. Goetz, editors. Great books of the Western world. Encyclopaedia Britannica, Chicago, IL, USA, second edition, 1990. ISBN 0-85229-531-6 (set). 749 pp. LCCN AC1 .G72 1990.

Atakishiyev:1994:SDR

[AFW94] Natig M. Atakishiyev, Alejandro Frank, and Kurt Bernardo Wolf. A simple difference realization of the Heisenberg q-algebra. Journal of Mathematical Physics, 35(7):3253–3260, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i7/
p3253_s1.

Arthurs:1988:QCG

[AG88] E. Arthurs and M. S. Goodman. Quantum correlations: a generalized Heisenberg uncertainty relation. Physical Review Letters, 60(24):2447–2449, June 1988. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

prl.aps.org/abstract/PRL/v60/i24/p2447_1.

Allaby:2002:MS

[AG02] Michael Allaby and Derek Gjertsen. Makers of Science. Makers of science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-521680-6 (set). 96 (vol. 1) pp. LCCN Q141 .A44 2002. URL http://www.gbv.de/dms/

goettingen/360423353.pdf;   http://www.loc.gov/catdir/
enhancements/fy0611/2001048396-d.html.

Agarwal:2002:HUR

[Aga02] Girish Saran Agarwal. Heisenberg’s uncertainty relations and quantum optics. In Papenfuss et al. [PLS02], pages 146– 153. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Ascoli:1957:QNW

[AH57] R. Ascoli and W. Heisenberg. Zur Quantentheorie nichtlinearer Wellengleichungen. IV. Elektrodynamik. (German) [On the quantum theory of nonlinear wave equations. IV. Electrodynamics].

Zeitschrift für Naturforschung, 12a:177–187, March 1957. CO-DEN ZNTFA2. ISSN 0372-9516.

Albrecht:1993:NTG

[AH93] Helmuth Albrecht and Armin Hermann, editors. Naturwissenschaft und Technik in der Geschichte: 25 Jahre Lehrstuhl für Geschichte der Naturwissenschaft und Technik am Historischen Institut der Universität Stuttgart. Verlag für Geschichte der Naturwissenschaften und der Technik, Stuttgart, Germany, 1993. ISBN 3-928186-15-9. 401 pp. LCCN Q183.4.G293 U556 1993.

Aaserud:2015:OHY

[AK15] Finn Aaserud and Helge Kragh, editors. One hundred years of the Bohr atom: proceedings from a conference, volume 1 of Scientia Danica. Series M, Mathematica et physica. Det Kongelige Danske Videnskabernes Selskab, Copenhagen, Denmark, 2015. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/
2015/06/Bohr-Book.pdf.

Albeverio:1973:STM

[Alb73] Sergio Albeverio. Scattering theory in a model of quantum fields. I. Journal of Mathematical Physics, 14(12):1800–1816, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v14/i12/p1800_s1.

Allen:2006:HBG

[All06] Michael Thad Allen. Hitler’s Bombe: Die geheime Geschichte der deutschen Kernwaffenversuche (review). Technology and Culture, 47(3):671–672, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/201438.

Altavilla:2006:FFW

[Alt06] Costanza Altavilla. Fisica e filosofia in Werner Heisenberg. (Italian) [Physics and Philosophy of Werner Heisenberg], volume 4 of Transazioni. Guida, Napoli, Italia, 2006. ISBN 88-6042-391-0. 368 pp. LCCN QC16.H35 L48 2006. 23.00 EUR. URL http://www.

loc.gov/catdir/toc/casalini05/07041675.pdf.

Aitchison:2004:UHM

[AMS04] Ian J. R. Aitchison, David A. MacManus, and Thomas M. Snyder. Understanding Heisenberg’s “magical” paper of July 1925: a new look at the calculational details. American Journal of Physics, 72 (11):1370–1379, November 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v72/i11/p1370_s1.

Andersen:1989:BRP

[And89] Per H. Andersen. Book review: Pekka Lahti and Peter Mittelstaedt (editors), Symposium on the Foundations of Modern Physics 1987: The Copenhagen Interpretation 60 Years after the Como Lecture. Physics Today, 42(9):92, 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v42/i9/p92/s1.

Anonymous:1929:PWS

[Ano29] Anonymous. Professor, 27, wins scientific honor: Dr. [Werner] Heisenberg of Leipzig get[s] medal and $2,500 from Research Corporation here. Award to Dr. [Bergen] Davis, also. Columbia physicist to receive like prize for studies in exchange of energy. New York Times, ??(??):25, April 23, 1929. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
results/13F38E18F585DFDACE3.

Anonymous:1933:NPG

[Ano33] Anonymous. Nobel Prize goes to Bunin, Russian; short-story write and poet, exile in France, is first of his nation to get award; three physicists named; Profs. Dirac and Schroedinger share for 1933 — Heisenberg gets the 1932 grant. New York Times, ?? (??):23, November 10, 1933. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/100827281.

Anonymous:1934:ETD

[Ano34] Anonymous. Electron theory of Dirac revised: California scholars dispose of main difficulties but obtain original results. Extend Heisenberg law. Two limitations to experimental knowledge set up by Oppenheimer and Furry. Electron theory of Dirac revised. New York Times, ??(??):n1, February 18, 1934. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1935:BWG

[Ano35a] Anonymous. Buchbesprechung: Wandlungen in den Grundlagen der Naturwissenschaften. Zwei Vorträge von Werner Heisenberg. Leipzig 1935. Verlag von S. Hirzel. 45 S. Preis geh. RM 2. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 41(7):458–459, 1935. CODEN ZEAPAA. ISSN 0005-9021.

Anonymous:1935:MQE

[Ano35b] Anonymous. Materiewellen und Quantenmechanik: eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. (German) [Matter waves and quantum mechanics: an introduction the foundations of the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Nature, 135(3403):88, January 19, 1935. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v135/n3403/pdf/135088d0.pdf.

Anonymous:1935:UP

[Ano35c] Anonymous. The Uncertainty Principle. Nature, 135(3407):261, February 16, 1935. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v135/n3407/pdf/135261a0.pdf. See [Dar31].

Anonymous:1936:GSG

[Ano36] Anonymous. German science goose-steps. New York Times, ?? (??):20, March 12, 1936. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/101908004/.

Anonymous:1942:EUG

[Ano42] Anonymous. Energiegewinnung aus Uran. (German) [Energy generation by uranium]. Report AMPG-I-34-105, Archiv der MPG, Berlin, Germany, February 26–28, 1942. 144 pp. Secret report [Ergebnisse der vom Heereswaffenamt Veranlaßten Forschungsarbeiten zur Nutzbarmachung von Atomkernenergien == Results of the research work initiated by the Army Ordnance Office on the utilization of nuclear energy] by unattributed Uranverein scientists to General Emil Leeb, head of German Army Ordinance. Cassidy [Cas09, page 321] says that Heisenberg probably did not write this report, but some of the wording and ideas are taken from Heisenberg’s known writings. Popp [Pop21] suggests that Kurt Diebner may have been the editor, but that it was, and still is, German government tradition to omit mention of authors.

Anonymous:1943:SBB

[Ano43] Anonymous. The smallest building blocks of matter. Krakauer Zeitung, ??(302):??, December 18, 1943. News report of Heisenberg lecture in Cracow, Poland. See [Ber04] for the circumstances and audience of this lecture.

Anonym:1944:ROB

[Ano44a] Anonym. Rapport om Begivenhederne under Besættelsen af Universitets Institut for teoretisk Fysik fra d. 6. December 1943 til d. 3. Februar 1944. (Danish) [Report of events during the occupation of the Institute for Theoretical Physics from 6 December 1943 to 3 February 1944]. Report ??, Institute for teoretisk Fysik, Københavns Universitet, Copenhagen, Denmark, 1944.

Anonymous:1944:WH

[Ano44b] Anonymous. Werner Heisenberg. Das Reich, ??(??):1, May 14, 1944.

Anonymous:1951:NPH

[Ano51] Anonymous. Der Nobelpreis für Physik: Heisenberg–Schrödinger 1932 Dirac 1933/Chadwick 1935/1936 Hess–Anderson. Physikalische Blätter, 7(9):404–409, September 1951. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19510070904/
abstract.

Anonymous:1954:BRH

[Ano54] Anonymous. Book review: Heisenberg, W., Nuclear physics. New York: Philosophical Library, 1953. 225 P. $4.75. Science Education, 38(4):323, 1954. CODEN SEDUAV. ISSN 1098-237X.

Anonymous:1960:TTC

[Ano60] Anonymous. The talk of the town: Connections. The New Yorker, 36(35):33–??, October 15, 1960. ISSN 0028-792X. URL http://

archives.newyorker.com/?i=1960-10-15#folio=032.

Anonymous:1961:EEG

[Ano61a] Anonymous, editor. Erkenntnis und Erziehung. (German) [Knowledge and education]. Österreichischer Bundesverlag für Unterricht, Wissenschaft und Kunst, Wien, Austria, 1961. 184 pp. LCCN LB7 .E7.

Anonymous:1961:HAS

[Ano61b] Anonymous. Heisenberg (autobiographischen Skizze). (German) [Heisenberg (autobiographical sketch)]. In ????, editor, Stifterverband für die Deutsche Wissenschaft. Jahrbuch 1961. (German) [Donors’ Association for German Science. Yearbook 1961], page 42. Stifterverband für die Deutsche Wissenschaft, Essen, West Germany, 1961.

Anonymous:1966:FMS

[Ano66] Anonymous. Frontiers of modern scientific philosophy and humanism: The Athens Meeting, 1964. Elsevier, Amsterdam, The Netherlands, 1966. ix + 91 pp. LCCN Q175 .F9 1966.

Anonymous:1968:BRH

[Ano68] Anonymous. Book review: Heisenberg, Werner. Physics and philosophy. New York 10016 (49 East 33rd Street): Harper & Row, Publishers. 218 P. $1.40. Science Education, 52(1):90, 1968. CO-DEN SEDUAV. ISSN 0036-8326 (print), 1098-237X (electronic).

Anonymous:1971:IPA

[Ano71] Anonymous. Interview mit dem Präsidenten der Alexander von Humboldt-Stiftung Professor Dr. Werner Heisenberg. (German) [Interview with the President of the Alexander von Humboldt Foundation, Professor Dr. Werner Heisenberg]. Alexander von Humboldt-Stiftung. Mitteilungen, 20(??):2–5, ???? 1971.

Anonymous:1973:IPH

[Ano73] Anonymous. Interview mit Prof. Heisenberg. (German) [Interview with Prof. Heisenberg]. In ????, editor, Maximilians-Gymnasium München. Bericht über das Schuljahr 1972/73 und Stimmen aus dem Max-Gymnasium. (German) [Maximilians Gymnasium in Munich. Report on the school year 1972/73 and voices from the Max Gymnasium], pages 14–17. Maximilians-Gymnasium, Munich, West Germany, 1973. LCCN ????

Anonymous:1975:DGC

[Ano75] Anonymous. Quantum electrodynamics: Translations of four papers by Heisenberg, Dirac, and Weisskopf. Technical Report RL-75-073, Science and Technology Facilities Council ePublication Archive, Chilbolton, Daresbury, and Rutherford Appleton Lab-

oratories, UK, 1975. URL http://epubs.cclrc.ac.uk/work-
details?w=21197.

Anonymous:1976:NHG

[Ano76] Anonymous. Nachlass Heisenberg. (German) [Heisenberg estate]. Max Planck-Gesellschaft, Archiv, Berlin, III. Abteilung, Repertorium 93., 1976.

Anonymous:1977:KKV

[Ano77a] Anonymous. K konchine Vernera K. Geĭzenberga. (Russian) [On the passing of Werner K. Heisenberg]. Uspekhi Fizicheskikh Nauk, 121(4):656–677, April 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/en/

articles/1977/4/d/.

Anonymous:1977:PWK

[Ano77b] Anonymous. On the passing of Werner K. Heisenberg. Soviet Physics. Uspekhi, 20(4):335, April 30, 1977. CODEN SOPUAP.

ISSN 0038-5670. URL http://ufn.ru/en/articles/1977/4/d/.

Anonymous:1983:MHP

[Ano83a] Anonymous. Memorandum zum Heisenberg-Programm. Physikalische Blätter, 39(12):411, December 1983. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19830391212/
abstract.

Anonymous:1983:UAU

[Ano83b] Anonymous. Uncertainties about Uncertainty Principle. Nature, 302(5907), March 31, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v302/n5907/pdf/302377a0.pdf.

Anonymous:1984:ANM

[Ano84a] Anonymous. Aktuelles: Neuer Mitarbeiter in der Redaktion der Physikalischen Blätter/Bisher über 1000 Bewerber im Heisenberg-Programm/1-Megabit-Speicherchip in Experimentalversion/Reparatur von Ionenlaser-Plasmarohren in England/ Forschungsbudget 1984/Zen-Ereignisse /Wissenschaftsrat verabschiedet Empfehlungen zum 14. Rahmenplan für den Hochschulbau/Faltings geht nach Princeton. Physikalische Blätter, 40 (7):235–237, July 1984. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19840400723/abstract.

Anonymous:1984:SAI

[Ano84b] Anonymous. Sommerfeld-Ausstellung im Deutschen Museum/ Heisenberg-Programm wird verlängert/Ausschreibung für den Max-Auwärter-Preis 1985/Welch-Stipendium 1986 (Ausschreibung)/European Who’s Who in Molecular Beams/Bericht über die Jahrestagung 1984 der ÖPG. Physikalische Blätter, 40(12): 392–393, December 1984. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19840401213/abstract.

Anonymous:1993:CGB

[Ano93a] Anonymous. Corrections: [German bomb program and Team B]. Bulletin of the Atomic Scientists, 49(4):6, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet93, CP93].

Anonymous:1993:RWS

[Ano93b] Anonymous. The rest was silence. New York Times, ??(??): BR3, March 14, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109207251/.

Anonymous:1994:BRB

[Ano94] Anonymous. Book review: Uncertainty, the Life and Science of Werner Heisenberg, David C. Cassidy, W. H. Freeman and Co., New York, 1992, 669 pp. Systems Research, 11(2):78–80, 1994. CODEN SYREER. ISSN 0731-7239 (print), 1099-1735 (electronic).

Anonymous:1998:HR

[Ano98] Anonymous. Heisenberg revisited. Nature, 396(6710):427–428, December 3, 1998. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v396/n6710/full/396427a0.html.

Anonymous:1999:HBH

[Ano99] Anonymous. Heisenberg, the bomb and the historical record. Nature, 400(6742):308, July 22, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v400/n6742/full/400308b0.html.

Anonymous:2000:UPG

[Ano00] Anonymous. An Uncertainty Principle for geometry. Nature, 404 (6779):705–706, April 13, 2000. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v404/n6779/full/404705a0.html.

Anonymous:2002:AVB

[Ano02a] Anonymous. Angst vor der Bombe: Carl Friedrich von Weizäcker über die neuen Dokumente und sein Treffen mit Bohr und Heisenberg. (German) [Fear of the bomb: Carl Friedrich von Weizäcker on the new documents and his meeting with Bohr and Heisenberg]. Süddeutsche Zeitung, ??(??):??, February 8, 2002.

Anonymous:2002:CIS

[Ano02b] Anonymous, editor. The Copenhagen interpretation: science and history on stage. Graduate Center of the City University of New York and Smithsonian Institute, New York, NY, USA, 2002. Three videocassetes recorded at the Graduate Center of the City University of New York, Baird Auditorium, March 2, 2002. Apparently not published in print.

Anonymous:2004:CHO

[Ano04a] Anonymous. Cheating Heisenberg with optical combs. Science, 306(5704):1997, December 17, 2004. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/306/5704/1997.2.full.pdf.

Anonymous:2004:HHT

[Ano04b] Anonymous. How to hit a telephone pole. Physics in Perspective (PIP), 6(4):486, December 2004. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Anonymous:2005:HF

[Ano05] Anonymous. Heisenberg force. In Van Nostrand’s Scientific Encyclopedia. Wiley, New York, NY, USA, 2005. ISBN 0-471-74398-4.

Anonymous:2006:BWH

[Ano06] Anonymous. Buchbesprechung: Werner Heisenberg: Die Verknüpfung von Physik und Philosophie. Von K. Sander (Hrsg.). Physik in unserer Zeit, 37(4):197, 2006. CODEN PHUZAH. ISSN 1521-3943.

Anonymous:2008:BHL

[Ano08] Anonymous. Werner Heisenberg in Leipzig, von Ch. Kleint u. G. Wiemers. Akademie-Verlag 1993. 264 S., Brosch ISBN 3-05-501585-1. Physik Journal, 7(??):??, September 18, 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (elec-

tronic). URL http://www.pro-physik.de/details/rezension/
1109813/Werner_Heisenberg_in_Leipzig.html. See [KW93].

Anonymous:2011:IBQ

[Ano11] Anonymous. Interaction-based quantum metrology showing scaling beyond the Heisenberg limit. Nature, 471(7339):486–489, March 23, 2011. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v471/n7339/full/nature09778.html.

Anonymous:2012:SCD

[Ano12] Anonymous. Scientists cast doubt on Heisenberg’s uncertainty principle. ScienceDaily, ??(??):??, September 7, 2012. URL http:

//www.sciencedaily.com/releases/2012/09/120907125154.

htm. News story about physics research [RDM^{+}12b].

Anonymous:2020:MAW

[Ano20] Anonymous. The mystery about Werner Heisenberg, the physicist who won the Nobel for discovering quantum mechanics.

Web site, May 10, 2020. URL https://www.explica.co/the-
mystery-about-werner-heisenberg-the-physicist-who-won-
the-nobel-for-discovering-quantum-mechanics/.

Anonymous:2025:RWH

[Ano25] Anonymous. Review of Werner Heisenberg’s Reality and its Order — corrigendum. Philosophy of Science, 92(1):237, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL https://www.cambridge.org/
core/journals/philosophy-of-science/article/review-of-
werner-heisenbergs-reality-and-its-order-corrigendum/
21410DA022A813360485CE8DC863C4DE. See [Cru23].

Anonymous:20xx:WKH

[Anoxx] Anonymous. Werner Karl Heisenberg (1901–1976). Infoamérica

Web site., 20xx. URL http://www.infoamerica.org/teoria/
heisenberg1.htm.

Alvarez:1944:LMR

[AO44] Luis Alvarez and J. R. Oppenheimer. Letter to Major Robert Furman. Web document, June 5, 1944. URL http://www.lanl.

gov/history/road/pdf/A%20German%20Atomic%20Bomb.pdf.

Atkinson:2022:HCH

[AP22] David Atkinson and Jeanne Peijnenburg. How certain is Heisenberg’s Uncertainty Principle? HOPOS: Journal of the International Society for the History of Philosophy of Science, 12(1):1–21, Spring 2022. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Appell:2001:NUP

[App01] David Appell. The new Uncertainty Principle. Scientific American, 284(1):18–19, January 2001. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v284/n1/pdf/
scientificamerican0101-18.pdf.

Argentieri:2009:CSE

[Arg09] Niccolò Argentieri. Ci sono elettroni nel mondo-della-vita?: esperienza, matematica, realtà: una lettura fenomenologica dell’epistemologia di Werner Heisenberg. (Italian) [Electrons in the real world: experience, mathematics, truth: a phenomenological study of the epistemology of Werner Heisenberg], volume 7 of Tascabili Bonanno. Filosofia. Bonanno, Acireale, Italia, 2009. ISBN 88-7796-549-5. 126 pp. LCCN QC16.H35 A76

2009. 10.00 EUR. URL http://www.loc.gov/catdir/toc/
casalini08/09777733.pdf.

Aristarhov:2023:HUP

[Ari23] Serj Aristarhov. Heisenberg’s uncertainty principle and particle trajectories. Foundations of Physics, 53(1):??, February 2023. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10701-022-00646-x.

Atmanspacher:1999:BSM

[ARK^{+}99] Harald Atmanspacher, Helmut Rechenberg, Horst Kant, Rolf Hempelmann, Reinhold Blümel, Gerhard Börner, Houjun Mo, Philippe Blanchard, Marius Grundmann, Dietrich Stauffer, and Christian Uebing. Bücher und Software: Meÿenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Band IV/Teil II: 1953–1954/Mladjenovic: The Defining Years in Nuclear Physics. 1932–1960s/Frenkel: Professor Friedrich Houtermans Arbeiten, Leben, Schicksal/Kärger and Heitjans: Diffusion in Condensed Matter/Friedrich: Theoretical Atomic Physics/Peacock: Cosmological Physics/Grosse u Martin: Particle Physics and the Schrödinger Equation/ Gaponenko: Optical Properties of Semiconductor Nanocrystals/ Gershenfeld: The Nature of Mathematical Modeling/Origin 6.0: Vertrieb. Physikalische Blätter, 55(9):79–83, September 1999. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19990550919/abstract.

ArroyoCamejo:2006:SQG

[Arr06] Silvia Arroyo Camejo. Skurrile Quantenwelt. (German) [Crazy Quantum World]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-29720-0. xiii + 246 pp. LCCN QC174.12 .A77 2006. EUR 29.95,

SFR 51.00. URL http://www.springer.com/physics/quantum+
physics/book/978-3-540-29720-8.

Anastopoulos:2003:RPS

[AS03] Charis Anastopoulos and Ntina Savvidou. The role of phase space geometry in Heisenberg’s uncertainty relation. Annals of Physics, 308(1):329–353, November 2003. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

Bopp:1961:WHP

[B^{+}61] F. (Fritz) Bopp et al., editors. Werner Heisenberg und die Physik unserer Zeit. (German) [Werner Heisenberg and the physics of his time]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961. 310 pp. LCCN QC3 .B59.

Babich:2002:HPS

[Bab02] Babette E. Babich. Hermeneutic Philosophy of Science, Van Gogh’s Eyes, and God: Essays in Honor of Patrick A. Heelan, S.J., volume 225 of Boston Studies in the Philosophy and History of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 90-481-5926-1, 94-017-1767-2 (e-book). ISSN 0068-0346. xvii + 484 pp. LCCN

Q174 .B67 vol. 225. URL https://link.springer.com/book/
10.1007/978-94-017-1767-0.

Bachem:1952:HIP

[Bac52] A. Bachem. Heisenberg’s indeterminacy principle and life. Philosophy of Science, 19(4):261–272, October 1952. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185393.

Bacciagaluppi:2022:SIB

[Bac22] Guido Bacciagaluppi. The statistical interpretation: Born, Heisenberg, and von Neumann, 1926–27. In Freire et al. [FBD^{+}22], page ?? ISBN 0-19-884449-2 (hardcover). LCCN

QC173.98 .O94 2022. URL https://global.oup.com/academic/
product/the-oxford-handbook-of-the-history-of-quantum-
interpretations-9780198844495.

Badash:1995:SDN

[Bad95] Lawrence Badash. Scientists and the development of nuclear weapons: from fission to the Limited Test Ban Treaty, 1939–1963. The control of nature. Humanities Press, Atlantic Highlands, NJ, USA, 1995. ISBN 0-391-03873-7, 0-391-03874-5 (paperback). ix + 129 pp. LCCN U264 .B34 1995.

Boya:2007:HPA

[Bal07] Luis Joaquín Boya Balet. Heisenberg y el Proyecto Atómico Alemán. (Spanish) [Heisenberg and the German atomic bomb project]. Revista Iberoamericana de Física, 3(1):78–82, 2007. ISSN

1888-2188.  URL https://dialnet.unirioja.es/revista/
12154/A/2007.

Ball:2013:SRS

[Bal13] Philip Ball. Serving the Reich: the struggle for the soul of science under Hitler. The Bodley Head Ltd, London, UK, 2013. ISBN 1-84792-248-1. ix + 303 pp. LCCN ????

Ball:2014:SRS

[Bal14] Philip Ball. Serving the Reich: the struggle for the soul of physics under Hitler. University of Chicago Press, Chicago, IL, USA, 2014. ISBN 0-226-20457-X (hardcover), 0-226-20460-X (e-book). ix + 303 pp. LCCN QC773.3.G3 B35 2014.

Barberousse:2004:AOB

[Bar04] Anouk Barberousse. Analyses d’ouvrages: Philosophie : Le manuscrit de 1942. ^{≪}Sources du savoir ^{≫}par Werner Heisenberg; Catherine Chevalley. Revue d’Histoire des Sciences, 57(1):214– 216, January 2004. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23634141.

Barad:2007:MUH

[Bar07] Karen Michelle Barad, editor. Meeting the Universe Halfway: Quantum Physics and the Entanglement of Matter and Meaning. Duke University Press, Durham, NC, USA, 2007. ISBN 0-8223-3901-3 (hardcover), 0-8223-3917-X (paperback). xiii + 524 pp. LCCN QC6 .B328 2007.

Bethe:1936:NPS

[BB36] H. A. Bethe and R. F. Bacher. Nuclear physics A. Stationary states of nuclei. Reviews of Modern Physics, 8(2):82–229, April 1, 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.8.82;   http://rmp.aps.org/
abstract/RMP/v8/i2/p82_1.

Born:1969:AEH

[BB69] Max Born and Hedwig Born, editors. Albert Einstein — Hedwig und Max Born. Briefwechsel 1916–55. (German) [Albert Einstein — Hedwig and Max Born. Correspondence 1916–55]. Nymphenburger Verlagshandlung, Munich, Germany, 1969. 329 pp. Preface by Bertrand Russell and foreword by Werner Heisenberg.

Born:1971:BELa

[BB71a] Max Born and Hedwig Born, editors. The Born–Einstein letters: correspondence between Albert Einstein and Max and Hedwig Born from 1916–1955. Macmillan Publishing Company, New York, NY, USA, 1971. ISBN 0-333-11267-9. xi + 240 + 5 pp. LCCN QC16.E5 A4513 1971b. Preface by Bertrand Russell and foreword by Werner Heisenberg. Translated by Irene Born.

Born:1971:BELb

[BB71b] Max Born and Hedwig Born, editors. The Born–Einstein letters; correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1971. ISBN 0-8027-0326-7. x + 240 pp. LCCN QC16.E5 A4513.

Boorse:1996:MPB

[BB96] Henry A. Boorse and Albert A. Bartlett. Manhattan Project: Book faulted, Heisenberg paged. Physics Today, 49(1):83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v49/i1/
p83/s1.

Born:2005:BEL

[BB05] Max Born and Hedwig Born, editors. The Born–Einstein letters: friendship, politics, and physics in uncertain times: correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Macmillan Publishing Company, New York, NY, USA, 2005. ISBN 1-4039-4496-2. xxxii + 235 pp. LCCN QC16.E5 A4 2005. URL http://www.

loc.gov/catdir/bios/hol059/2004061027.html;   http://
www.loc.gov/catdir/description/hol054/2004061027.html;
 http://www.loc.gov/catdir/enhancements/fy0618/2004061027-

t.html. Note on the new edition by Gustav Born. New preface by Diana Buchwald and Kip S. Thorne. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. Translated by Irene Born.

Biermann:1984:KSG

[BBB^{+}84] L. Biermann, P. Budini, J. Buschmann, M. Deutschmann, E. Freese, and K. Gottstein. Kosmische strahlung. (German) [Cosmic radiation]. In Blum et al. [BDR84e], pages 477–508. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_39/.

Bennett:1992:QC

[BBE92] Charles H. Bennett, Gilles Brassard, and Artur K. Ekert. Quantum cryptography. Scientific American, 267(4):50–?? (Intl. ed. 26–??), October 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bopp:1945:BVD

[BBF^{+}45] Fritz Bopp, Walther Bothe, Erich Fischer, Erwin Fünfer, Werner Heisenberg, O. Ritter, and Karl Wirtz. Bericht über einen Versuch mit 1.5 to D_{2}O und U und 40 cm Kohlernückstreumantel (B7). (German) [Report on an experiment with 1.5 tonnes D_{2}O and U, and 40 cm graphite backscatter cloak (B7)]. Kernphysikalische Forschungsberichte G-300, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, January 3, 1945. 7 pp.

Bopp:1957:DGN

[BBF^{+}57a] Fritz Bopp, Max Born, Rudolf Fleischmann, Walter Gerlach, Otto Hahn, Otto Haxel, Werner Heisenberg, Hans Kopfermann, Max von Laue, Heinz Maier-Leibnitz, Josef Mattauch, Friedrich-Adolf Paneth, Wolfgang Paul, Wolfgang Riezler, Fritz Strassmann, Wilhelm Walcher, Carl Friedrich von Weizsäcker, and Karl Wirtz. Declaration of the German nuclear physicists. Bulletin of the Atomic Scientists, 13(6):228, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The declaration calls for the Federal Republic of Germany (‘West Germany’) to “renounce explicitly and voluntarily the possession of atomic weapons of any kind”, and states that none of the authors is ‘ready in any way to take part in the production, the tests, or the application of atomic weapons”. See followup [vW57].

Bopp:1957:EAA

[BBF^{+}57b] Fritz Bopp, Max Born, Rudolf Fleischmann, Walter Gerlach, Otto Hahn, Otto Haxel, Werner Heisenberg, Hans Kopfermann, Max von Laue, Heinz Maier-Leibnitz, Josef Mattauch, Friedrich-Adolf Paneth, Wolfgang Paul, Wolfgang Riezler, Fritz Strassmann, Wilhelm Walcher, Carl Friedrich von Weizsäcker, and Karl Wirtz. Erklärung von Achtzehn Atomforschern. (German) [Declaration of eighteen atomic scientists]. Mitteilungen aus der Max-Planck-Gesellschaft zur Förderung der Wissenschaften, 2(??):62–64, ???? 1957.

Becker:1962:MAE

[BBH^{+}62] Helmut Becker, Joachim Beckmann, Werner Heisenberg, Günter Howe, Georg Picht, Ludwig Reiser, and Carl Friedrich von Weizsäcker. Das Memorandum der acht evangelischen Persönlichkeiten zur Bonner Politik. (German) [The memorandum of eight Protestant personalities to Bonn policy]. Süddeutsche Zeitung, 49(??):4, February 26, 1962.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989.

US$11.95. URL http://www.loc.gov/catdir/description/
dover031/89012033.html.

Bernstein:1995:BAF

[BC95] Jeremy Bernstein and David Cassidy. Bomb apologetics: Farm Hall, August 1945. Physics Today, 48(8):32–36, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v48/i8/p32_s1.

Bacciagaluppi:2009:HSP

[BC09] Guido Bacciagaluppi and Elise Crull. Heisenberg (and Schrödinger, and Pauli) on hidden variables. Studies in History and Philosophy of Modern Physics, 40(4):374–382, December 2009. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S1355219809000501.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12–17, 1993, volume 73 of Proceedings in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Billette:1969:CTE

[BCL^{+}69] J. J. Billette, C. Campillo, C. Lee, R. D. McConnell, G. Pariseau, and G. Fischer. Concerning “A thought experiment violating Heisenberg’s uncertainty principle”. Canadian Journal of Physics = Journal canadien de physique, 47(21):2415–2416, ???? 1969. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). See [Rob69].

Bitbol:1992:ESP

[BD92] Michel Bitbol and Olivier Darrigol, editors. Erwin Schrödinger: philosophy and the birth of quantum mechanics = philosophie et naissance de la mécanique quantique. Editions Frontières, Gif-sur-Yvette, France, 1992. ISBN 2-86332-116-1. xiii + 460 pp. LCCN QC16.S265 E78 1992. 263.75F.

Bethe:1970:ZVF

[BDH70] H. A. Bethe, P. A. M. Dirac, and V. Heisenberg. Zhizn’ v fizike. Vechernie lekcii v Trieste. (Russian) [Life in physics: Evening lectures at Trieste]. Uspekhi Fizicheskikh Nauk, 102(10):279–312, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-

6652 (electronic). URL http://ufn.ru/ru/articles/1970/10/

f/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Beckers:1988:HOS

[BDH88] J. Beckers, D. Dehin, and V. Hussin. On the Heisenberg and orthosymplectic superalgebras of the harmonic oscillator. Journal of Mathematical Physics, 29(7):1705–1711, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/
p1705_s1.

Bethe:1989:LP

[BDH^{+}89] Hans A. (Hans Albrecht) Bethe, P. A. M. Dirac, W. Heisenberg, E. P. Wigner, O. Klein, L. D. Landau, and E. M. Lifshitz, editors. From a life of physics. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1989. ISBN 9971-5-0937-7. ix + 92 pp. LCCN QC71 .F74 1989. URL http://

adsabs.harvard.edu/abs/1989liph.book.....B.

Bagge:1957:UBC

[BDJ57] Erich Bagge, Kurt Diebner, and Kenneth Jay. Von der Uranspaltung bis Calder Hall. (German) [From the fission of uranium to Calder Hall], volume 41 of Rowohlts deutsche Enzyklopädie Sachgebiet Physik. Rowohlt, Hamburg, West Germany, 1957. 165 pp. LCCN ????

Blum:1984:WHGa

[BDR84a] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke: Abt. A: Wissenschaftliche Originalarbeiten. 1. 2.. (German) [Werner Heisenberg: Collected Works: Series A: Original scientific papers. Part 1. Part 2]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-13400-X (A part 1), 0-387-13847-1 (A part 2). xi + 633 (A part 1), x + 717 (A part 2) pp. LCCN QC3 .H33 1984. URL http://www.springer.

com/series/183.

Blum:1984:WHGb

[BDR84b] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abt. B: Wissenschaftliche Übersichtsartikel, Vorträge und Bücher. (German) [Werner Heisenberg: Collected Works: Series B: Scientific review papers, talks, and books]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-540-13020-9, 0-387-13020-9. x + 937 pp. LCCN QC3 .H33 1984.

Blum:1984:WHGd

[BDR84c] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 1. Physik und Erkenntnis 1927–1955. Ordnung der Wirklichkeit. (German) [Werner Heisenberg: Collected Works: Series C, Volume 1. Physics and knowledge] 1927–1955. Order of reality. Piper, Munich, West Germany, 1984. ISBN 3-492-02925-6. x + 453 pp. LCCN ????

Blum:1984:WHGe

[BDR84d] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 2. Physik und Erkenntnis 1956–1968. (German) [Werner Heisenberg: Collected Works: Series C, Volume 2. Physics and knowledge 1956–1968]. Piper, München, West Germany, 1984. ISBN 3-492-02926-4. 440 pp. LCCN ????

Blum:1984:WHGf

[BDR84e] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke: Collected Works: Series B: Scientific Review Papers, Talks, and Books: Wissenschaftliche Übersichtsartikel, Vorträge und Bücher, volume 1 of Gesammelte Werke / Collected Works. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). xx

+ 936 pp. LCCN ???? URL http://www.springerlink.com/
content/978-3-642-61742-3.

Blum:1984:DEK

[BDR84f] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Durchgang energiereicher Korpuskeln durch den Atomkern. (German) [Passage of energetic corpuscles through the atomic nucleus]. In Blum et al. [BDR84e], page 261. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_22?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Blum:1984:EKS

[BDR84g] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Elementarteilchen und kosmische Strahlung. (German) [Elementary particles and cosmic rays]. In Blum et al. [BDR84e], pages 333– 334. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-61742-3_25?coverImageUrl=/static/0.
363/sites/link/images/abstract_cover_placeholder.png.

Blum:1984:FTA

[BDR84h] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Fortschritte in der Theorie des Atomkerns. (15. XI. 37). (German) [Advances in the theory of the atomic nucleus. (15 XI. 37)]. In Blum et al. [BDR84e], page 260. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_21?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Blum:1984:NPP

[BDR84i] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Nonlinear problems in physics. In Blum et al. [BDR84e], pages 863– 874. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_63/.

Blum:1984:QDE

[BDR84j] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Quantitatives über die Deformierbarkeit edelgasähnlicher Ionen. (German) [Quantitative on the deformability of noble gas similar ions]. In Blum et al. [BDR84e], pages 27–28. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_4?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Blum:1984:TS

[BDR84k] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Theorie der Schauer. (German) [Theory of showers]. In Blum et al. [BDR84e], pages 247–248. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_17?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Blum:1984:ES

[BDR84l] Walter Blum, Hans-Peter Dürr, and Helmut Rechenberg. Zur Elektronentheorie der Supraleitung. (German) [Electron theory of superconductivity]. In Blum et al. [BDR84e], page 413. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-61742-3_32?coverImageUrl=/static/0.
363/sites/link/images/abstract_cover_placeholder.png.

Blum:1984:WHGc

[BDR85a] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abt. C: Allgemeinverständliche Schriften. 1: Physik und Erkenntnis 1927–1955. 2: Physik und Erkenntnis 1956–1968. 3: Physik und Erkenntnis 1969–1976. (German) [Werner Heisenberg: Collected Works: Series C: Philosophical and popular writings. Part 1. Physics and Epistemology 1927–1955. Part 2: Physics and Epistemology 1956–1968. Part 3: Physics and Epistemology 1969–1976]. Piper-Verlag, München, Germany, 1984–1985. ISBN 3-492-02925-6 (C part 1), 3-492-02926-4 (C part 2), 3-492-02927-2 (C part 3). 453 (C part 1), viii + 440 (C part 2), 542 (C part 3) pp. LCCN ????

Blum:1985:WHG

[BDR85b] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 3. Allgemeinverständliche Schriften. Physik und Erkenntnis 1969–1976: der Teil und das Ganze, die Bedeutung des Schönen, naturwissenschaftliche und religiöse Wahrheit, Elementarteilchen u.a.. (German) [Werner Heisenberg: Collected Works: Series C, Volume 3. Philosophical and popular writings. Physics and knowledge 1969–1976: The part and the whole, the importance of beauty, scientific and religious truth, elementary particles]. Piper, München, Germany; Zürich, Switzerland, 1985. ISBN 3-492-02927-2. viii + 542 pp. LCCN ????

Blum:1986:WHG

[BDR86] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 4. Allgemeinverständliche Schriften. Biographisches und Kernphysik: Autobiographisches, Laudationes, Buchbesprechungen, Kernphysik, Münchener Festrede u.a. (German) [Werner Heisenberg: Collected Works: Series C, Volume 4: Philosophical and popular writings. Biographical and nuclear physics. Autobiographical, Laudations, Book Reviews, Nuclear Physics, Munich Festival Speeches, a.o.]. Piper, München, West Germany, 1986. ISBN 3-492-02928-0. x + 505 pp. LCCN ????

Blum:1989:WHGa

[BDR89a] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung A. Teil II. Wissenschaftliche Originalarbeiten. (German) [Collected works. Series A. Part II. Original scientific papers]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-13847-1, 3-540-13847-1. x + 717 pp. Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Blum:1989:WHGb

[BDR89b] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke. Abteilung C, Band 5. Wissenschaft und Politik: Physik im Dritten Reich, Organisation der Forschung, Internationale Wissenschaft, A. v. Humboldt-Stiftung, Physik, Technik und Gesellschaft, Politische Stellungnahmen u.a.. (German) [Werner Heisenberg: Collected Works: Series C, Volume 5. Science and politics. Physics in the Third Reich. Organization of research, international science, A. v. Humboldt Foundation, Physics, Technology and Society, Political Opinions, a.o.]. Piper-Verlag, München, Germany, 1989. ISBN 3-492-02929-9. xiii + 652 pp. LCCN ???? URL http://d-nb.

info/900168226/04.

Blum:1993:WHG

[BDR93] W. (Walter) Blum, H.-P. (Hans-Peter) Dürr, and Helmut Rechenberg, editors. Werner Heisenberg: Gesammelte Werke.

Abteilung A. Teil III. Wissenschaftliche Originalarbeiten. (German) [Werner Heisenberg: Collected works. Series A. Part III. Original scientific papers]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-13848-X. x + 700 pp. Annotated by Rolf Hagedorn, Heinz Koppe, H.-P. Dürr and H. Rechenberg.

Beccaria:2002:NAC

[Bec02] Matteo Beccaria. Numerical algorithm for the calculation of the ground states in the U_{q}SU(2) symmetric spin-Heisenberg chain. Applied Numerical Mathematics: Transactions of IMACS, 41(2): 269–281, May 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bedurftig:2003:HAB

[Bed03] Friedemann Bedürftig. Als Hitler die Atombombe baute — Lügen und Irrtümer über das Dritte Reich. (German) [When Hitler Built the Atomic Bomb — Lies and Mistakes About the Third Reich]. Piper. Piper, München, Germany, 2003. ISBN 3-492-04443-3. 267 pp. LCCN DD256.47 .B42 2003.

Bedurftig:2004:HAB

[Bed04] Friedemann Bedürftig. Als Hitler die Atombombe baute — Lügen und Irrtümer über das Dritte Reich. (German) [When Hitler Built the Atomic Bomb — Lies and Mistakes About the Third Reich], volume 4146 of Piper. Piper, München, Germany, 2004. ISBN 3-492-24146-8 (paperback). 267 pp. LCCN DD256.47 .B42 2003.

Bederson:2009:BRD

[Bed09] Benjamin Bederson. Book review: David C. Cassidy, Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb. New York: Bellevue Literary Press, 2009, 456 pages. $27.00 (cloth). Physics in Perspective (PIP), 11(3):351–353, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Beijerinck:2013:CB

[Bei13] Herman C. W. Beijerinck. Crossing borders. Europhysics News, 44(4):21, July/August 2013. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic).

Beller:1985:PJI

[Bel85] Mara Beller. Pascual Jordan’s influence on the discovery of Heisenberg’s indeterminacy principle. Archive for History of Exact Sciences, 33(4):337–349, December 1985. CODEN AH-ESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&
issn=0003-9519&volume=33&issue=4&spage=337.

Beller:1992:GBC

[Bel92] Mara Beller. The genesis of Bohr’s complementarity principle and the Bohr–Heisenberg dialogue. In The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4 [UM92], chapter 18, pages 273–293. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346.

LCCN Q174 .S354 1992. URL https://link.springer.com/
chapter/10.1007/978-94-011-2688-5_18.

Beller:1999:QDM

[Bel99] Mara Beller. Quantum dialogue: the making of a revolution. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1999. ISBN 0-226-04181-6 (hardcover), 0-226-04182-4. xv + 365 + 8 pp. LCCN QC174.13 .B45 1999. URL http:

//www.loc.gov/catdir/description/uchi052/99035499.htm;
 http://www.loc.gov/catdir/toc/uchi052/99035499.htm.

Bennewitz:1930:NBP

[Ben30] ?. Bennewitz. Neue Bücher: Die physikalischen Prinzipien der Quantentheorie. Von W. Heisenberg. 114 Seiten. Verlag S. Hirzel, Leipzig 1930. Preis geh. RM. 7, geb. RM. 8,50. Angewandte Chemie, 43(41):914, 1930. CODEN ANCEAD. ISSN 1521-3757.

Bennewitz:1934:NBM

[Ben34] ?. Bennewitz. Neue Bücher: Die moderne Atomtheorie. W. Heisenberg; E. Schrödinger; P. A. M. Dirac. Verlag S. Hirzel, Leipzig 1934. Preis geh. RM. 2,50. Angewandte Chemie, 47(37):655, 1934. CODEN ANCEAD. ISSN 1521-3757.

Bennewitz:1935:NBW

[Ben35] ?. Bennewitz. Neue Bücher: Wandlungen in den Grundlagen der Naturwissenschaft. Von W. Heisenberg. Zwei Vorträge. S. Hirzel, Leipzig 1935. Preis br. RM. 2. Angewandte Chemie, 48(26):406, 1935. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Benedek:2013:LEA

[Ben13] Giorgio Benedek. Letter to the editor: About “Crossing borders”. Europhysics News, 44(6):32, November/December

2013. URL http://www.europhysicsnews.org/articles/epn/
pdf/2013/06/epn2013-44-6.pdf. Brief comment on [Bei13,

Luc07].

Beresford:1935:IQ

[Ber35] John Beresford. The irrational quantity: pq − qp = \sqrt{−1}h̄/(2π). Manchester Guardian, ??(??):18, February 14, 1935. ISSN 0307-

756X. URL https://theguardian.newspapers.com/image/
258883415/.

Bernstein:1991:CP

[Ber91] Jeremy Bernstein. The charms of a physicist: The Joy of Insight. Passions of a Physicist. by Victor Weisskopf. New York Review of Books, 38(7):47–50, April 11, 1991. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1991/apr/11/the-charms-of-a-physicist/.

Bernstein:1992:FHT

[Ber92] Jeremy Bernstein. The Farm Hall transcripts: The German scientists and the bomb. New York Review of Books, 39(49):47– 53, August 13, 1992. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1992/aug/13/the-farm-hall-transcripts-the-german-scientists-
an/.

Bernstein:1993:RFH

[Ber93] Jeremy Bernstein. Revelations from Farm Hall: Book review: Heisenberg’s War: The Secret History of the German Bomb, by Thomas Powers. Science, 259(5103):1923–1926, March 26, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/259/5103/
1923.extract.

Bernstein:1995:WDH

[Ber95] Jeremy Bernstein. What did Heisenberg tell Bohr about the bomb? Scientific American, 272(5):92–97 (Intl. ed. 72–??), May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v272/n5/pdf/scientificamerican0595-
92.pdf.

Bernstein:1996:HUC

[Ber96] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Bernstein:2001:MPO

[Ber01] Jeremy Bernstein. The Merely Personal: Observations on Science and Scientists. Ivan R. Dee, Chicago, IL, USA, 2001. ISBN 1-56663-344-3. 245 pp. LCCN Q171 .B5375 2001.

Bernstein:2002:HCM

[Ber02] Jeremy Bernstein. Heisenberg and the critical mass. American Journal of Physics, 70(9):911–916, September 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/70/911/1.

Bernstein:2003:DWH

[Ber03] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bernstein:2004:HP

[Ber04] Jeremy Bernstein. Heisenberg in Poland. American Journal of Physics, 72(3):300–304, March 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.

aapt.org/resource/1/ajpias/v72/i3/p300_s1. See comment

and Bernstein’s reply [Got04].

Beskrovnyi:1998:AMA

[Bes98] V. N. Beskrovnyi. Applying Mathematica to the analytical solution of the nonlinear Heisenberg operator equations. Computer Physics Communications, 111(1–3):76–86, June 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465598000290.

Beste:2004:GPQ

[Bes04] Dieter Beste, editor. Grosse Physiker: Quantenrätsel — Werner Heisenberg und die Unschärfe; Relativitätstheorie — Albert Einstein und die Schwarzen Löcher; Zwei Klassiker —- Isaac Newton und James Clerk Maxwell. (German) [Great physicists: Quantum mystery — Werner Heisenberg and Uncertainty; Relativity theory — Albert Einstein and black holes; Two classical authors — Isaac Newton and James Clerk Maxwell]. Spektrum der Wissenschaft Dossier. Spektrum-der-Wiss.-Verl.-Ges., Heidelberg, Germany, 2004. ISBN 3-936278-75-X. 82 pp. LCCN ????

Bethe:1991:RAP

[Bet91] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1993:LBG

[Bet93] Hans A. Bethe. Letter: Bethe on the German bomb program. Bulletin of the Atomic Scientists, 49(1):53–54, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See corrections [Ano93a].

Bethe:2000:GUP

[Bet00] Hans A. Bethe. The German Uranium Project. Physics Today, 53(7):34–36, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v53/i7/p34_s1.

Beyerchen:1977:SUH

[Bey77] Alan D. Beyerchen. Scientists under Hitler: politics and the physics community in the Third Reich. Yale University Press, New Haven, CT, USA, 1977. ISBN 0-300-01830-4. xii + 287 + 5 pp. LCCN QC9.G3 B48.

Beyerchen:1982:WHP

[Bey82] Alan D. Beyerchen. Wissenschaftler unter Hitler: Physiker im Dritten Reich. (German) [Scientists under Hitler: physicists in the Third Reich], volume 34098 of Ullstein-Buch; Ullstein-Buch. Ullstein, Frankfurt/M, West Germany, 1982. ISBN 3-548-34098-9. 379 pp. LCCN ????

Bopp:1942:UNS

[BFH^{+}42a] Fritz Bopp, Erich Fischer, Werner Heisenberg, Carl-Friedrich von Weizsäcker, and Karl Wirtz. Untersuchungen mit neuen Schichtenanordnungen aus U-metall und Paraffin. (German) [Studies with a new layer arrangement of uranium metal and paraffin]. Kernphysikalische Forschungsberichte G-127, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, March 1942. 4 pp.

Bopp:1942:VBE

[BFH^{+}42b] Fritz Bopp, Erich Fischer, Werner Heisenberg, Carl-Friedrich von Weizsäcker, and Karl Wirtz. Vorläufiger Bericht über Ergebnisse an einer Schichtenkugel aus 38-Metall und Paraffin. (German) [Preliminary results with a layer sphere of 38 metal and paraffin]. Technical Report G-126, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, January 6, 1942. 3 pp.

Balakrishnan:1996:SEA

[BG96] Radha Balakrishnan and Partha Guha. Surface-embeddability approach to the dynamics of the inhomogeneous Heisenberg spin chain. Journal of Mathematical Physics, 37(8):3651–3661, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Born:1990:QMI

[BGĬ90] M. Born, V. Geĭzenberg, and P. Ĭordan. Quantum mechanics. II (a fragment). In Einstein collection, 1986–1990 (Russian), pages 146–157. Nauka, Moscow, Russia, 1990. Translated from the German by A. N. Temchin.

Born:1923:EIA

[BH23a] Max Born and Werner Heisenberg. Die Elektronenbahnen im angeregten Heliumatom. (German) [The electron trajectories in the excited helium atom]. Zeitschrift für Physik, 16(1):229–243, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01327396;   http://www.springerlink.
com/content/t75836n61671070k/.

Born:1923:PBM

[BH23b] Max Born and Werner Heisenberg. Über Phasenbeziehungen bei den Bohrschen Modellen von Atomen und Molekeln. (German) [On phase relations in the Bohr model of atoms and molecules]. Zeitschrift für Physik, 14(1):44–55, December 1923. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://www.springerlink.com/content/v71005762t157n89/.

Born:1924:EDI

[BH24a] Max Born and Werner Heisenberg. Über den Einfluß der Deformierbarkeit der Ionen auf optische und chemische Konstanten. I. (German) [On the influence of the deformability of ions on optical and chemical constants. I]. Zeitschrift für Physik, 23 (1):388–410, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.springerlink.

com/content/h0588704m33v5l60/.

Born:1924:QMG

[BH24b] Max Born and Werner Heisenberg. Zur Quantentheorie der Molekeln. (German) [Toward a quantum theory of molecules]. Annalen der Physik (1900), 379(9):1–31, ???? 1924. CODEN ???? ISSN 1521-3889.

Born:1928:MQF

[BH28] Max Born and Werner Heisenberg. La mécanique des quanta. (French) [The mechanics of quanta]. In ????, editor, Electrons et Photons: Rapports et Discussions du Cinquième Conseil de Physique Tenu à Bruxelles du 24 au 29 Octobre 1927 sous les Auspices de l’Institut International de Physique Solvay, pages 143– 184. Gauthier-Villars et cie, Paris, France, 1928. LCCN ????

Born:1963:EIA

[BH63a] Max Born and Werner Heisenberg. Die Elektronenbahnen im angeregten Heliumatom. (German) [Electron orbits in the excited helium atom]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 48, pages 23–37. LCCN ???? Reprinted from Z. Physik 16, 229–243 (1923).

Born:1963:EDI

[BH63b] Max Born and Werner Heisenberg. Über den Einfluß der Deformierbarkeit der Ionen auf optische und chemische Konstanten I. (German) [On the influence of the deformability of ions on optical and chemical constants I]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 49, pages 38–60. LCCN ???? Reprinted from Z. Physik 23, 388–410 (1924).

Born:1975:QM

[BH75] Max Born and Werner Heisenberg. 5. quantum mechanics. In Mehra [Meh75], chapter 6, pages 146–149. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Brown:1982:BEP

[BH82] Laurie M. Brown and Lillian Hoddeson. The birth of elementaryparticle physics. Physics Today, 35(4):36–43, April 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v35/i4/p36_s1. This is a much-abridged version of [BH83a].

Brown:1983:BEP

[BH83a] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. In The Birth of Particle Physics [BH83b], pages 3–36. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Brown:1983:BPP

[BH83b] Laurie M. Brown and Lillian Hoddeson, editors. The Birth of Particle Physics. Cambridge University Press, Cambridge, UK, 1983. ISBN 0-521-24005-0 (hardcover), 0-521-33837-9 (paperback). LCCN QC793 .B57 1983.

Born:1984:MQ

[BH84] Max Born and Werner Heisenberg. La mécanique des quanta. (French) [The mechanics of quanta ]. In Blum et al. [BDR84e], pages 58–99. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book).

LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-3-642-61742-3_7/.

Born:1926:QIG

[BHJ26] Max Born, Werner Heisenberg, and Pascual Jordan. Zur Quantenmechanik II. (German) [Toward quantum mechanics. II]. Zeitschrift für Physik, 35(8–9):557–615, August 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01379806;
http://www.springerlink.com/content/r46w151t550pl624/.

English translation in [BHJ67].

Born:1962:BMG

[BHJ62] Max Born, Werner Heisenberg, and Pascual Jordan, editors. Zur Begründung der Matrizenmechanik. (German) [On the Foundations of Matrix Mechanics], volume 2 of Dokumente der Naturwissenschaft. Abteilung Physik. E. Battenberg, Stuttgart, West Germany, 1962. 135 pp. LCCN QC174.3 .Z8.

Born:1963:QIG

[BHJ63] Max Born, Werner Heisenberg, and Pascal Jordan. Zur Quantenmechanik II. (German) [On quantum mechanics II]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 54, pages 155–213. LCCN ???? Reprinted from Z. Physik 85, 557–615 (1926).

Born:1967:QMc

[BHJ67] Max Born, Werner Heisenberg, and Pascual Jordan. On quantum mechanics II. In van der Waerden [vdW67], pages 321–385. LCCN QC174.12 S655. English translation of [BHJ26].

Baird:1998:HHC

[BHN98] Davis Baird, R. I. G. Hughes, and Alfred Nordmann. Heinrich Hertz: Classical Physicist, Modern Philosopher, volume 198 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346. xi + 324 pp. LCCN ???? URL https://link.

springer.com/book/10.1007/978-94-015-8855-3.

Bishop:1996:MSQ

[BHX96] R. F. Bishop, R. G. Hale, and Y. Xian. A microscopic study of the quantum critical behavior of the spin-1/2 anisotropic Heisenberg models. International Journal of Quantum Chemistry, 57(5):919–927, March 5, 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/abstract?ID=60430.

Billig:2003:RTH

[Bil03] Yuly Billig. Representations of the twisted Heisenberg–Virasoro algebra at level zero. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 46(??):529–537, ???? 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Born:1925:QGT

[BJ25] Max Born and Pascual Jordan. Zur Quantenmechanik. (German) [Toward quantum mechanics]. Zeitschrift für Physik, 34(1): 858–888, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.

com/article/10.1007/BF01328531. English translation of pages

858–882 in [BJ67].

Born:1967:QMb

[BJ67] Max Born and Pascual Jordan. On quantum mechanics. In van der Waerden [vdW67], pages 277–306. LCCN QC174.12 S655. English translation of pages 858–882 of [BJ25].

Blum:2017:THH

[BJLR17] Alexander Blum, Martin Jähnert, Christoph Lehner, and Jürgen Renn. Translation as heuristics: Heisenberg’s turn to matrix mechanics. Studies in History and Philosophy of Modern Physics, 60 (??):3–22, November 2017. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219816301770.

Bergstrom:2001:PCS

[BJN01] L. Bergström, K. E. Johansson, and Ch. Nilsson. The physics of Copenhagen for students and the general public. Physics Education, 36(5):388–393, September 2001. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.

iop.org/0031-9120/36/i=5/a=303.

Basu:1988:HWG

[BK88] Debabrata Basu and T. K. Kar. The Heisenberg–Weyl group in the coherent state basis and the Bargmann transform. Journal of Mathematical Physics, 29(8):1854–1859, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/
p1854_s1.

Bedford:2004:HIF

[BK04] Donald Bedford and Peter Krumm. Heisenberg indeterminacy and the fine structure constant. American Journal of Physics, 72(7): 969–970, July 2004. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?AJP/
72/969/1.

Barone:1993:NCH

[BKBS93] S. R. Barone, E. E. Kunhardt, J. Bentson, and A. Syljuasen. Newtonian chaos + Heisenberg Uncertainty = macroscopic indeterminacy. American Journal of Physics, 61(5):423–437, May 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/61/423/1.

Bernstein:2003:HBE

[BL03] Jeremy Bernstein and Peter D. Lax. Heisenberg & the bomb: An exchange. New York Review of Books, 50(17):??, November 6, 2003. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2003/nov/06/heisenberg-the-bomb-an-exchange/. Reply by

Thomas Powers.

Boya:2008:ARF

[BL08] Luis J. Boya and Amand A. Lucas. About “Revisiting Farm Hall”. Europhysics News, 39(2):13–15, March/April 2008. CO-DEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (elec-

tronic). URL http://www.europhysicsnews.org/articles/
epn/pdf/2008/02/epn2008-39-2.pdf. See [Luc07].

BlancoCendon:1986:TAP

[Bla86] Fernando Blanco Cendon. En torno al principio de indeterminación de Werner Karl Heisenberg. (Spanish) [On the Uncertainty Principle of Werner Karl Heisenberg], volume 9 of Publicaciones de los Institutos Pontificios de Filosofía y Teología. Serie III, Cuadernos de filosofía. Universidad pontificia de Santo Tomás de Manila, Instituto de filosofía de Madrid, Madrid, Spain, 1986. ISBN 84-600-4226-X. 133 pp. LCCN QC174.17.H4 B56 1986.

Bleuler:1990:WHI

[Ble90] K. Bleuler. Werner Heisenberg’s ideas on particle physics in the light of recent achievements. Zeitschrift für Naturforschung, 45(9– 10):1051–1058, ???? 1990. CODEN ZNTFA2. ISSN 0372-9516.

Bloch:1976:HED

[Blo76a] Felix Bloch. Heisenberg and the early days of quantum mechanics. Physics Today, 29(12):23–27, December 1, 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3024633.

Bloch:1976:RHE

[Blo76b] Felix Bloch. Reminiscences of Heisenberg and the early days of quantum mechanics. Physics Today, 29(12):23–27, December 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/

v29/i12/p23/s1. Reprinted in [WP85, pages 319–323].

Bloch:1985:HED

[Blo85] Felix Bloch. Heisenberg and the early days of quantum mechanics. In Weart and Phillips [WP85], pages 319–323. ISBN 0-88318-468-0 (paperback). LCCN QC7 .H694 1985. Reprint of [Blo76b].

Bjorkman:2019:ICT

[BLW19] Maria Björkman, Patrik Lundell, and Sven Widmalm, editors. Intellectual collaboration with the Third Reich: treason or reason? Routledge studies in Second World War history. Routledge, Abingdon, Oxon, UK, 2019. ISBN 0-8153-9474-8 (hardcover), 1-351-18508-X (mobi), 1-351-18509-8 (e-pub), 1-351-18510-1 (adobe), 1-351-18511-X (e-book). ???? pp. LCCN DD256.6 .H58 2019.

Brown:1984:LTM

[BM84] Laurie M. Brown and Donald F. Moyer. Lady or tiger? — The Meitner–Hupfeld effect and Heisenberg neutron theory. American Journal of Physics, 52(2):130–136, February 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Brown:1986:HNP

[BM86] Laurie M. Brown and Arthur I. Miller. Heisenberg and nuclear physics. Physics Today, 39(6):116–??, June 1986. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v39/i6/p116/s1.

Baskerville:1993:BHG

[BM93] W. K. Baskerville and S. Majid. The braided Heisenberg group. Journal of Mathematical Physics, 34(8):3588–3606, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v34/i8/p3588_s1.

Brambilla:1994:CBC

[BM94] Nora Brambilla and Emilio Montaldi. On the connection between continuous Hahn polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 35(1):389–392, January 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i1/
p389_s1.

Braun:2011:HLS

[BM11] Daniel Braun and John Martin. Heisenberg-limited sensitivity with decoherence-enhanced measurements. Nature Communications, 2(3):223, March 1, 2011. CODEN NCAOBW. ISSN

2041-1723 (electronic). URL http://www.nature.com/ncomms/
journal/v2/n3/full/ncomms1220.html.

Bilharz:1950:BRC

[BMG^{+}50] H. Bilharz, E. Moritz, H. Görtler, W. Lorey, K. Philipp, G. Möllenstedt, E. G. Ramberg, and U. Jetter. Bücher: Runge: Carl Runge/Willers: Zahlzeichen und Rechnen im Wandel der Zeit/Lense: Vom Wesen der Mathematik/Mac Duffee, Jung: Vektoren, Determinanten und Matrizen/Heisenberg: Die Physik der Atomkerne/Räntsch: Die Optik in Der Feinmesstechnik/Borries: Die Übermikroskopie/Krieger: Radio-Technisches WÖrterbuch/ Dingler: Methodische Philosophie/Bonhoeffer: Phys. Chem. Modelle von Lebensvorgängen/Eichler: Prinzipien des Lebendigen. Physikalische Blätter, 6(3):134–138, March 1950. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19500060307/abstract.

Bender:1987:CHP

[BMP87] Carl M. Bender, Lawrence R. Mead, and Stephen S. Pinsky. Continuous Hahn polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 28(3):509–513, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/
p509_s1.

Bruning:2008:SLM

[BN08] E. Brüning and S. Nagamachi. Solution of a linearized model of Heisenberg’s fundamental equation II. Journal of Mathematical Physics, 49(5):052304, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v49/i5/p052304_s1.

Bohm:1962:CNC

[Boh62] David Bohm. Classical and non-classical concepts in the quantum theory : an answer to Heisenberg’s Physics and Philosophy. British Journal for the Philosophy of Science, XII(48):265–280, February 1962. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://bjps.oxfordjournals.org/
content/XII/48/265.full.pdf+html;   http://www.jstor.
org/stable/685396.

Bohr:1964:KAOa

[Boh64] Aage Bohr. Krigens år og atomvåbnenes perspektiver. (Danish) [The war years and the prospects of atomic weapons]. In Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers] [RRK^{+}64], pages 191–214. LCCN ????

URL http://www.nba.nbi.dk/papers/rozental.htm. Contains

comments on the Bohr–Heisenberg meeting in the summer of 1941.

Bohr:1967:WYP

[Boh67] Aage Bohr. The war years and the prospects raised by the atomic weapons. In Rozental [Roz67], page ?? LCCN QC16.B63 N53.

Bokulich:2004:OCD

[Bok04] Alisa Bokulich. Open or closed? Dirac, Heisenberg, and the relation between classical and quantum mechanics. Studies in History and Philosophy of Modern Physics, 35(3):377–396, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219804000139.

Bokulich:2006:HMK

[Bok06] Alisa Bokulich. Heisenberg meets Kuhn: Closed theories and paradigms. Philosophy of Science, 73(1):90–107, January 2006. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (elec-

tronic). URL http://www.jstor.org/stable/10.1086/510176.

Bokulich:2008:RQC

[Bok08] Alisa Bokulich. Reexamining the Quantum–Classical Relation: Beyond Reductionism and Pluralism. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-85720-1. x + 195 pp. LCCN QA805 .B685 2008. URL http://www.loc.gov/

catdir/enhancements/fy0834/2008018902-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0834/2008018902-d.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008018902-t.html.

Bokulich:2010:BRK

[Bok10] Alisa Bokulich. Book review: Kristian Camilleri, Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher (2009) Kristian Camilleri, Cambridge University Press, Cambridge, ISBN-13 978-0-521-88484-6 pp. xii + 199. Studies in History and Philosophy of Modern Physics, 41(4):368–370, November 2010. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981000047X. See [Cam09].

Born:1949:TTT

[Bor49] Max Born. Two topics in theoretical physics (review of two lectures by W. Heisenberg). Nature, 164(4161):165, July 30, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v164/
n4161/pdf/164165a0.pdf.

Born:1953:BRB

[Bor53] Max Born. Book review: Philosophic Problems of Nuclear Science. Eight Lectures, by W. Heisenberg and F. C. Hayes. The Philosophical Quarterly, 3(10):88–90, January 1953. CODEN ???? ISSN 0031-8094 (print), 1467-9213 (electronic). URL http://

www.jstor.org/stable/pdfplus/2216712.pdf.

Borgers:1958:BRB

[Bor58] Alfons Borgers. Book review: The Causality Problem in Atomic Physics, by Niels Bohr; Language and Reality in Modern Physics, by Werner Heisenberg; Die Stellung der Logik im Gebaude der Heutigen Wissenschaft, by Evert Willem Beth. Journal of Symbolic Logic, 23(1):66, March 1958. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://www.jstor.

org/stable/2964516.

Born:1961:BSD

[Bor61] Max Born. Bemerkungen zur statistischen Deutung der Quantenmechanik. (German) [Remarks on the statistical interpretation of quantum mechanics]. In W. Heisenberg und die Physik unserer Zeit. (German) [W. Heisenberg and the physics of our time], pages 103–118. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1961.

Born:1963:AAGb

[Bor63a] Max Born. Ausgewählte Abhandlungen. (German) [Selected works], volume 2. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1963. viii + 706 pp. LCCN ???? Mit einem Verzeichnis der wissenschaftlichen Schriften (German) [With a table of contents of scientific writings].

Born:1963:BSD

[Bor63b] Max Born. Bemerkungen zur statistischen Deutung der Quantenmechanik. (German) [Comments on the statistical interpretation of quantum mechanics]. In Ausgewählte Abhandlungen. (German) [Selected works] [Bor63a], chapter 70, pages 454–468. LCCN ???? Reprinted from Werner Heisenberg und die Physik unserer Zeit. Vieweg, Braunschweig 1961, pp. 103–118.

Borie:1970:MHR

[Bor70] Edith Borie. S matrix in the Heisenberg representation. Journal of Mathematical Physics, 11(12):3487–3496, December 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i12/p3487_s1. See errata [Bor72].

Borie:1972:EMH

[Bor72] Edith Borie. Errata: “S Matrix in the Heisenberg Representation”. Journal of Mathematical Physics, 13(10):1664, October 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v13/i10/p1664_s1. See [Bor70].

Borresen:2012:FNP

[Bør12] Hans Christofer Børresen. Flawed nuclear physics and atomic intelligence in the campaign to deny Norwegian heavy water to Germany, 1942–1944. Physics in Perspective (PIP), 14(4):471–497, December 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-012-0094-9.

Burdet:1975:RCE

[BP75] G. Burdet and M. Perrin. Realizations of the central extension of the inhomogeneous symplectic algebra as time dependent invariance algebras of nonrelativistic quantum systems. Journal of Mathematical Physics, 16(8):1692–1703, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i8/
p1692_s1.

Bernaschi:2011:BGC

[BPP11] M. Bernaschi, G. Parisi, and L. Parisi. Benchmarking GPU and CPU codes for Heisenberg spin glass over-relaxation. Computer Physics Communications, 182(6):1265–1271, June 2011. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465511000749.

Burdet:1974:ETD

[BPS74] G. Burdet, M. Perrin, and P. Sorba. On the explicit-time-dependent invariance properties of quantum mechanical systems. Journal of Mathematical Physics, 15(12):2253–2256, December 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v15/i12/p2253_s1.

Bohr:1964:NBH

[BR64] Niels Bohr and S. (Stefan) Rozental, editors. Niels Bohr: hans liv og virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: his life and works told by a circle of friends and colleagues]. J. H. Schultz, København, Danmark, 1964. 341 pp. LCCN QC16.B63 N5. English translation in [Roz67].

Brown:1987:PDW

[BR87] Laurie M. Brown and Helmut Rechenberg. Paul Dirac and Werner Heisenberg—a partnership in science. In Kurşunoğlu and Wigner [KW87], pages 117–162. ISBN 0-521-34013-6. LCCN QC16.D57 R46 1987. URL http://adsabs.

harvard.edu/abs/1987ragp.book.....D;   http://www.loc.
gov/catdir/description/cam023/86033409.html;   http://
www.loc.gov/catdir/toc/cam028/86033409.html.

Brown:1996:HMN

[BR96a] Laurie M. Brown and Helmut Rechenberg. Heisenberg’s model of the nucleus (1932–33). In The Origin of the Concept of Nuclear Forces [BR96b], chapter 2.2, pages 33–35. ISBN 0-7503-0373-5. LCCN QC793.3.B5 B76 1996. URL http://catdir.loc.gov/

catdir/enhancements/fy0701/96031758-d.html.

Brown:1996:OCN

[BR96b] Laurie M. Brown and Helmut Rechenberg. The Origin of the Concept of Nuclear Forces. Institute of Physics Publishers, Bristol, UK, 1996. ISBN 0-7503-0373-5. xii + 392 pp. LCCN

QC793.3.B5 B76 1996. URL http://catdir.loc.gov/catdir/
enhancements/fy0701/96031758-d.html.

Bracken:1997:NSG

[Bra97] P. Bracken. Numerical solution for the ground-state energy of the anisotropic Heisenberg model. International Journal of Quantum Chemistry, 62(1):13–21, ???? 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=42485;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=42485&PLACEBO=IE.pdf.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Breit:1923:HTA

[Bre23] G. Breit. The Heisenberg theory of the Anomalous Zeeman Effect. Nature, 112(2811):396, September 15, 1923. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v112/
n2811/pdf/112396a0.pdf.

Brennan:1997:HPS

[Bre97] Richard P. Brennan. Heisenberg probably slept here: the lives, times, and ideas of the great physicists of the 20th Century. Wiley, New York, NY, USA, 1997. ISBN 0-471-15709-0 (cloth). xi + 274 pp. LCCN QC15 .B74 1997; 97.E02533.

Bridgman:1929:NVS

[Bri29] Percy Williams Bridgman. The new vision of science. Harper’s Monthly Magazine, 158(??):443–451, March 1929. ISSN 0361-

7815. URL http://harpers.org/archive/1929/03/the-new-
vision-of-science/.

Brink:1965:NF

[Bri65a] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon, New York, NY, USA, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brink:1965:THW

[Bri65b] David Maurice Brink. The theories of Heisenberg, Wigner and Majorana. In Nuclear Forces [Bri65a], chapter 2, pages 13–25. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965.

Blum:2006:PH

[BRL06] Barbara Blum, Helmut Rechenberg, and Anna Ludovico. Per Heisenberg, volume 158 (or 11??) of Scienze storiche, filosofiche, pedagogiche e psicologiche. Aracne, Roma, Italy, 2006. ISBN 88-548-0636-6. 91 pp. LCCN ???? URL http://digital.casalini.

it/8854806366.

Bromberg:1971:INB

[Bro71] Joan Bromberg. The impact of the neutron: Bohr and Heisenberg. Historical Studies in the Physical Sciences, 3(??):307–341, ???? 1971. CODEN HSPSAS. ISSN 0073-2672. URL http://www.

jstor.org/stable/27757321.

Bromberg:1972:BRT

[Bro72] Joan Bromberg. Book review: Twentieth Century The Born– Einstein Letters. Correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Trans. by Irene Born. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. London: Macmillan, 1971. Pp. xi + 240. £3.85. British Journal for the History of Science, 6 (2):222–223, December 1972. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4025309.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:HPU

[Bro73b] Jacob Bronowski. Heisenberg’s Principle of Uncertainty. In The Ascent of Man [Bro73a], pages 364–368. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Brown:1977:HTF

[Bro77] H. A. Brown. Heisenberg theory of ferromagnetism: constant coupling revisited. American Journal of Physics, 45(8):707–710, August 1977. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/45/707/1.

Brown:1985:BRBb

[Bro85] Laurie M. Brown. Book review: Gesammelte Werke; Collected Works by Werner Heisenberg; W. Blum; H.-P. Durr; H. Rechenberg. Isis, 76(3):451–452, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232933.

Brown:1986:BRBa

[Bro86] Laurie M. Brown. Book review: Gesammelte Werke — Collected Works. Series A/Part 1: Wissenschaftliche Originalarbeiten — Original Scientific Papers by Werner Heisenberg; W. Blum; H.- P. Durr; H. Rechenberg. Isis, 77(3):527–528, September 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/231627.

Brooks:1992:HNW

[Bro92] Geoffrey Brooks. Hitler’s nuclear weapons: the development and attempted deployment of radiological armaments by Nazi Germany. L. Cooper, London, UK, 1992. ISBN 0-85052-344-3. 212 pp. LCCN UA710 .B8442 1992.

Broad:1993:NFR

[Bro93] William J. Broad. New findings reopen debate on Heisenberg’s role in German war effort.: Was Heisenberg saboteur or savant of Nazi drive for A-bomb? New York Times, ??(??):C1– C2, September 1, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109030191/.

Brown:1995:BRK

[Bro95] Laurie M. Brown. Book review: Karl von Meyenn, Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg a. o.; Vol. III: 1940–1949. Physics Today, 48(4):86, April 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v48/i4/
p86/s1.

Brown:1997:BRB

[Bro97] Neil Brown. Book review: Heisenberg Probably Slept Here: The Lives, Times, and Ideas of the Great Physicists of the 20th Century, by Richard P. Brennan. Physics Education, 32 (5):??, September 1997. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://iopscience.iop.

org/0031-9120/32/5/029.

Brodsky:2002:HMFa

[Bro02a] Stanley J. Brodsky. The Heisenberg matrix formulation of quantum field theory. Fortschritte der Physik = Progress of Physics, 50 (5–7):503–511, 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Brodsky:2002:HMFb

[Bro02b] Stanley J. Brodsky. The Heisenberg matrix formulation of quantum field theory. In Papenfuss et al. [PLS02], pages 74– 82. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Brown:2009:WRN

[Bro09] Timothy Scott Brown. Weimar radicals: Nazis and communists between authenticity and performance, volume 28 of Monographs in German history. Berghahn Books, New York, NY, USA, 2009. ISBN 1-84545-564-9 (hardcover). xii + 213 pp. LCCN HN460.R3 B77 2009.

Bruche:1964:WWM

[Brü64] E. Brüche. Was wußte man 1943/44 in Deutschland von der Atombombe? (German) [What was known in 1943/44 in Germany about the atomic bomb?]. Physikalische Blätter, 20(5):220–225, May 1964. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19640200504/.

Brush:1976:IIF

[Bru76] Stephen G. Brush. Irreversibility and indeterminism: Fourier to Heisenberg. Journal of the History of Ideas, 37(4):603–630, October/December 1976. CODEN JHIDAT. ISSN 0022-5037 (print),

1086-3222 (electronic). URL https://www.jstor.org/stable/
2709027.

Brush:1980:CCP

[Bru80] Stephen G. Brush. The chimerical cat: Philosophy of quantum mechanics in historical perspective. Social Studies of Science, 10 (4):393–447, November 1, 1980. CODEN SSSCDH. ISSN 0306-3127 (print), 1460-3659 (electronic). URL http://www.jstor.

org/stable/284918;   https://journals.sagepub.com/doi/
pdf/10.1177/030631278001000401.

Bethe:1947:GSA

[BS47] H. A. Bethe and H. S. Sack. German scientists in army employment: II. A protest. Bulletin of the Atomic Scientists, 3(2):65, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bricmont:1999:SDC

[BSMB99] Jean Bricmont, Alan Sokal, N. David Mermin, and Mara Beller. Sokalratic debate continues, fueled by Latour and Copenhagen interpretations. Physics Today, 52(8):15–17, 1999. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v52/i8/p15/s1.

Bruche:1970:BHT

[BSP^{+}70] E. Brüche, K. Strubecker, M. Päsler, G. Höhler, D. Lilienthal, J. Eggert, W. A. P. Luck, E. Brüche, R. Torge, and W. Kranzer. Bücher: Heisenberg: Der Teil und das Ganze/ Hewitt: Abstract Harmonic Analysis II/O’Dell: The Electrodynamics of Magneto-Electric Media/Deschamps: Mathematics Applied to Physics/Kaiser and Perlmutter: Coral Gables Conf. on fundamental interactions at high energy/Kowalski: Nuclear Electronics/Moelwyn-Hughes: Physikalische Chemie/Fletscher: The Chemical Physics of Ice/Schimmel: Elektronenmikroskopische Methodik/Happer: Light Propagation and Light Shifts in Optical Pumping Experiments/Leighton: Exercises in introductory Physics/Klatt: Langenscheidts Taschenwörterbuch Englisch Deutsch 1970. Physikalische Blätter, 26(7):331–335, July 1970.

CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19700260712/abstract.

Babbitt:1978:GSR

[BT78] Donald Babbitt and Lawrence Thomas. Ground state representation of the infinite one-dimensional Heisenberg ferromagnet. III. Scattering theory. Journal of Mathematical Physics, 19(8):1699– 1704, August 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i8/p1699_s1.

Bergvelt:1987:DDA

[Bt87] M. J. Bergvelt and A. P. E. ten Kroode. Differential-difference AKNS equations and homogeneous Heisenberg algebras. Journal of Mathematical Physics, 28(2):302–306, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i2/
p302_s1.

Bagarello:1996:HDS

[BT96] F. Bagarello and C. Trapani. The Heisenberg dynamics of spin systems: a quasi ^{∗}-algebras approach. Journal of Mathematical Physics, 37(9):4219–4234, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buhrke:1995:BWH

[Büh95] Thomas Bührke. Bücher: Werner Heisenberg — Leben und Werk. Von David C. Cassidy. 19 Abb., 808 S., geb. DM 68,-. Spektrum Akademischer Verlag, Heidelberg 1995. ISBN 3-86025-315-8. Physik in unserer Zeit, 26(6):285, 1995. CODEN PHUZAH. ISSN 0031-9252.

Buhrke:1998:LWM

[Büh98a] Thomas Bührke. “Wenn man beide Augen zugleich aufmachen will, dann wird man irre.” Werner Heisenberg (1901–1976). (German) [“If you want to open both eyes at the same time, then you will be mistaken.” Werner Heisenberg (1901–1976)]. In Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner] [Büh98b], pages 184–206. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). LCCN ????

Buhrke:1998:NAS

[Büh98b] Thomas Bührke. Newtons Apfel: Sternstunden der Physik; von Galilei bis Lise Meitner. (German) [Newton’s apple: great moments of physics; from Galileo to Lise Meitner], volume 1202 of Beck’sche Reihe. Verlag C. H. Beck, München, Germany, 1998. ISBN 3-406-44402-4, 3-406-42002-8 (paperback). 258 pp. LCCN ????

Bunge:1971:BRT

[Bun71] Mario Bunge. Book review: Der Teil Und Das Ganze, by Werner Heisenberg, 334 pp. R. Piper Verlag, München, 1969. [Translated as Physics and Beyond: Encounters and Conversations, 272 pp. Harper and Row, New York. 1971]. Physics Today, 24(1):63–64, January 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v24/i1/p63/s1.

Bunge:1977:IHI

[Bun77] Mario Bunge. The interpretation of Heisenberg’s inequalities. In Pfeiffer [Pfe77], pages 146–156. ISBN 3-492-02275-8. LCCN QC16.H35 D46. DM28.00.

Burkhardt:1962:GSS

[Bur62] Gerd Burkhardt. German scientists speak up. Bulletin of the Atomic Scientists, 18(6):45–46, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacciagaluppi:2009:QTC

[BV09] Guido Bacciagaluppi and Antony Valentini, editors. Quantum theory at the crossroads: reconsidering the 1927 Solvay conference. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-81421-9 (hardcover). LCCN QC173.96 .B33 2009. URL http://

www.loc.gov/catdir/toc/ecip0817/2008019585.html.

Born:2000:BRQ

[BW00a] Gustav Born and Alvin M. Weinberg. Copenhagen — roles in quantum mechanics. Physics Today, 53(7):74–75, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v53/i7/p74/s1.

Born:2000:CRQ

[BW00b] Gustav Born and Alvin M. Weinberg. Copenhagen — roles in quantum mechanics. Physics Today, 53(7):74–75, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v53/i7/p74/s1.

Buschhorn:2004:FPH

[BW04] Gerd W. Buschhorn and Julius Wess, editors. Fundamental physics — Heisenberg and beyond: Werner Heisenberg Centennial Symposium “Developments in Modern Physics”. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20201-3. LCCN QC173.96 .W47 2004. URL

http://www.springer.com/978-3-540-20201-1.

Cleveland:1976:OHH

[CA76] Charles L. Cleveland and Rodrigo Medina A. Obtaining a Heisenberg Hamiltonian from the Hubbard model. American Journal of Physics, 44(1):44–46, January 1976. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/?AJP/44/44/1.

Callaway:2014:AEN

[Cal14] H. G. Callaway. Arthur S. Eddington: The Nature of the Physical World: Gifford Lectures of 1927. Cambridge Scholars Publishing, Newcastle upon Tyne, England, annotated edition, 2014. ISBN 1-4438-6386-6, 1-4438-6703-9 (e-book). 431 pp. LCCN

Q143.E4 .A784 2014eb. URL http://alltitles.ebrary.com/
Doc?id=10934892.

Camilleri:2005:HTK

[Cam05] Kristian Camilleri. Heisenberg and the transformation of Kantian philosophy. International Studies in the Philosophy of Science, 19(3):271–287, 2005. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/02698590500462273.

Camilleri:2006:HWP

[Cam06] Kristian Camilleri. Heisenberg and the wave-particle duality. Studies in History and Philosophy of Modern Physics, 37(2):298–315, June 2006. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219806000025.

Camilleri:2007:BHD

[Cam07a] Kristian Camilleri. Bohr, Heisenberg and the divergent views of complementarity. Studies in History and Philosophy of Modern Physics, 38(3):514–528, September 2007. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S135521980600092X.

Camilleri:2007:ILC

[Cam07b] Kristian Camilleri. Indeterminacy and the limits of classical concepts: The transformation of Heisenberg’s thought. Perspectives on Science, 15(2):178–201, Summer 2007. CO-DEN PRSIEU. ISSN 1063-6145 (print), 1530-9274 (elec-

tronic). URL http://muse.jhu.edu/journals/perspectives_
on_science/v015/15.2camilleri.html;   http://muse.jhu.
edu/journals/perspectives_on_science/v015/15.2camilleri.
pdf;   http://www.mitpressjournals.org/doi/abs/10.1162/
posc.2007.15.2.178.

Camilleri:2009:HIQ

[Cam09] Kristian Camilleri. Heisenberg and the interpretation of quantum mechanics: the physicist as philosopher. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-88484-5 (hardcover). xii + 199 pp. LCCN QC16.H395 .C34 2009. URL http://www.

loc.gov/catdir/toc/ecip0825/2008034113.html.

Cap:1957:AMPa

[Cap57a] F. Cap. Arthur March/Der Pour le Merite für Lise Meitner und Werner Heisenberg. Physikalische Blätter, 13(7):317–318, July 1957. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19570130705/abstract.

Cap:1957:AMPb

[Cap57b] F. Cap. Arthur March/Der Pour le Mérite für Lise Meitner und Werner Heisenberg. (German) [Arthur March/the Order of Merit for Lise Meitner and Werner Heisenberg]. Physik Journal, 13(7): 317–318, July 1957. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Cappelletti:2001:DAC

[Cap01] Valentina Cappelletti. Dall’ordine alle cose: saggio su Werner Heisenberg. (Italian) [From Order to Chaos: test on Werner Heisenberg], volume 16, 471 of Lo spoglio dell’occidente; Di fronte e attraverso. Jaca Book, Milano, Italia, 2001. ISBN 88-16-40571-6. 234 pp. LCCN QC174.12 .C348 2001. 16.53 EUR.

Carhart:1971:CPT

[Car71] R. A. Carhart. Canonical perturbation theory for nonlinear quantum oscillators and fields. Journal of Mathematical Physics, 12 (8):1748–1762, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v12/i8/p1748_s1.

Carella:1976:HCM

[Car76] M. J. Carella. Heisenberg concept of matter as potency. Diogenes, 96(??):25–37, ???? 1976. CODEN ???? ISSN 1467-7695.

Carson:1997:BRBb

[Car97]  Cathryn Carson.  Book review: Wolfgang Pauli: Wis-

senschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Volume 4, Part 1: 1950–1952. [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, et al. Volume 4, Part 1: 1950–1952.] by Wolfgang Pauli; Karl von Meyenn. Isis, 88 (4):726–727, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237871.

Carson:1999:BRB

[Car99a] Cathryn Carson. Book review: Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture by Paul Lawrence Rose. Isis, 90(4):835–836, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237723.

Carson:1999:NMS

[Car99b] Cathryn Carson. New models for science in politics: Heisenberg in West Germany. Historical Studies in the Physical and Biological Sciences, 30(1):115–171, ???? 1999. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

Carson:2002:HFSa

[Car02a] Cathryn Carson. Heisenberg and the framework of science policy. Fortschritte der Physik = Progress of Physics, 50(5–7):432– 436, 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Carson:2002:HFSb

[Car02b] Cathryn Carson. Heisenberg and the framework of science policy. In Papenfuss et al. [PLS02], pages 3–7. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Carson:2003:OSH

[Car03] C. Carson. Objectivity and the scientist: Heisenberg rethinks. Science in Context, 16(1–2):243–269, March 2003. CODEN SC-COEW. ISSN 0269-8897 (print), 1474-0664 (electronic).

Carson:2010:HAA

[Car10] Cathryn Carson. Heisenberg in the atomic age: science and the public sphere. Publications of the German Historical Institute. German Historical Institute, Washington, DC, USA, 2010. ISBN 0-521-82170-3 (hardcover). xvi + 541 + 8 pp. LCCN QC16.H45 C37 2010.

Casimir:1974:BRW

[Cas74] H. B. G. Casimir. Book review: Werner Heisenberg, Across the Frontiers. Physics Today, 27(8):48–??, August 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v27/i8/p48/s2.

Casertano:1975:FFE

[Cas75] Giovanni Casertano. Fisica e filosofia: Einstein–Infeld, de Broglie, Heisenberg, Planck. (Italian) [Physics and philosophy: Einstein– Infeld, de Broglie, Heisenberg, Planck]. Il tripode, Napoli, 1975. 150 pp. LCCN QC6 .F55.

Cassidy:1976:WHC

[Cas76] David Charles Cassidy. Werner Heisenberg and the crisis in quantum theory, 1920–1925. Thesis (Ph.D.), Purdue University, West Lafayette, IN, USA, 1976. xiv + 520 pp.

Cassidy:1978:HFP

[Cas78] David C. Cassidy. Heisenberg’s first paper. Physics Today, 31(7):23–28, July 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v31/i7/p23/s1.

Cassidy:1979:GHH

[Cas79a] David C. Cassidy. Gustav Hertz, Hans Geiger und das Physikalische Institut der Technischen Hochschule Berlin in den Jahren 1933 bis 1945. (German) [Gustav Hertz, Hans Geiger and the Physical Institute of the Technical University of Berlin in the years 1933 to 1945]. In Rürup [Rür79], pages 373–387. ISBN 0-387-09672-8, 3-540-09672-8. LCCN T173 .B51938 1979.

Cassidy:1979:HFC

[Cas79b] David C. Cassidy. Heisenberg’s first core model of the atom: The formation of a professional style. Historical Studies in the Physical Sciences, 10(??):187–224, ???? 1979. CODEN HSPSAS. ISSN

0073-2672. URL http://www.jstor.org/stable/27757390.

Cassidy:1991:BLS

[Cas91] D. C. Cassidy. Uncertainty — the Life and Science of Werner Heisenberg. Current Contents, 49(??):5–13, ???? 1991.

Cassidy:1992:HGS

[Cas92a] D. C. Cassidy. Heisenberg, German science, and the 3rd Reich. Social Research, 59(3):643–661, ???? 1992.

Cassidy:1992:HUQ

[Cas92b] David C. Cassidy. Heisenberg, uncertainty and the quantum revolution. Scientific American, 266(5):106–112 (Intl. ed. 64–??), May 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v266/n5/pdf/scientificamerican0592-
106.pdf.

Cassidy:1992:ULS

[Cas92c] David C. Cassidy. Uncertainty: the life and science of Werner Heisenberg. W. H. Freeman and Company, New York, NY, USA, 1992. ISBN 0-7167-2243-7. xii + 669 + 16 pp. LCCN QC16.H35 C37 1992; QC 16 .H35C37 1992. US$29.95.

Cassidy:1993:ACB

[Cas93a] David C. Cassidy. Atomic conspiracies: Book review: Heisenberg’s war: the secret history of the German bomb, by Thomas Powers, Knopf/Cape: 1993. Pp. 610. $27.50, £20. Nature, 363(6427):311– 312, May 27, 1993. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v363/n6427/pdf/363311a0.pdf.

Cassidy:1993:GBN

[Cas93b] David C. Cassidy. Germany and the bomb: New evidence. Scientific American, 268(2):120, February 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v268/n2/pdf/
scientificamerican0293-120.pdf.

Cassidy:1993:WHD

[Cas93c] David C. Cassidy. Werner Heisenberg — Die deutsche Wissenschaft und das Dritte Reich. (German) [Werner Heisenberg — German science and the Third Reich]. In Albrecht and Hermann [AH93], pages 65–?? ISBN 3-928186-15-9. LCCN Q183.4.G293 U556 1993.

Cassidy:1995:WHL

[Cas95] David C. Cassidy. Werner Heisenberg: Leben und Werk. (German) [Werner Heisenberg: Life and Work]. Spektrum, Akademische Verlag, Heidelberg / Berlin, Germany, 1995. ISBN 3-86025-315-8. 786 pp. LCCN ???? URL

http://www.dandelon.com/intelligentSEARCH.nsf/alldocs/
A5B4016FF90055E0C12570FC00275999/.

Cassidy:1997:BRBb

[Cas97] David C. Cassidy. Book review: Heisenberg Probably Slept Here: The Lives, Times, and Ideas of the Great Physicists of the Twentieth Century by Richard P. Brennan. Isis, 88(4):726, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237870.

Cassidy:2000:HPC

[Cas00] David C. Cassidy. A historical perspective on Copenhagen. Physics Today, 53(7):28–32, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v53/i7/p28/s1.

Cassidy:2001:WHB

[Cas01] David C. Cassidy. Werner Heisenberg: a bibliography of his writings. Whittier Publications, Inc., Island Park, NY, USA, second edition, 2001. ISBN 1-57604-115-8 (paperback). xvii + 226 pp. LCCN Z8395.513 .C38 2001.

Cassidy:2002:BRB

[Cas02a] David C. Cassidy. Book review: Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. Isis, 93(4):720–721, December 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/10.1086/376027.

Cassidy:2002:NLC

[Cas02b] David C. Cassidy. New light on “Copenhagen” and the German Nuclear Project. Physics in Perspective (PIP), 4(4):447–455, December 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s000160200004.

Cassidy:2002:WHOa

[Cas02c] David C. Cassidy. Werner Heisenberg: An overview of his life and work. Fortschritte der Physik = Progress of Physics, 50(5–7): 437–439, 2002. CODEN FPYKA6. ISSN 1521-3978.

Cassidy:2002:WHOb

[Cas02d] David C. Cassidy. Werner Heisenberg: An overview of his life and work. In Papenfuss et al. [PLS02], pages 8– 10. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Cassidy:2003:BRB

[Cas03] David C. Cassidy. Book review: Werner Heisenberg: Gutachtenund Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). Isis, 94(2):390–391, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/379455.

Cassidy:2005:BRW

[Cas05] D. C. Cassidy. Book review: Werner Heisenberg, Liebe Eltern! Briefe aus kritischer Zeit 1918 bis 1945. [(German) [Dear parents! Letters from a critical time 1918 to 1945]]. Edited by Anna Maria Hirsch-Heisenberg. With a foreword by Helmut Rechenberg. Munich, Germany: Langen Müller in der F. A. Herbig Verlagsbuchhandlung, 2003, 400 pages. EUR26.00 (cloth). Physics in Perspective (PIP), 7(2):253–256, June 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Cassidy:2011:BRC

[Cas11] David C. Cassidy. Book review: Cathryn Carson, Heisenberg in the Atomic Age: Science and the Public Sphere. Washington, DC: German Historical Institute and New York: Cambridge University Press, 2010, xvi + 541 pages. $80.00 (cloth). Physics in Perspective (PIP), 13(2):250–255, June 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Cassidy:2014:CFW

[Cas14a] David C. Cassidy. Carl Friedrich von Weizsäcker und Werner Heisenberg. In Hentschel and Hoffmann [HH14], pages 485–502. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Castellari:2014:FMF

[Cas14b] Marco Castellari, editor. Formula e metafora: figure di scienziati nelle letterature e culture contemporanee. (Italian) [Formula and metaphor: figures of scientists in the literature and contemporary cultures], volume 8 of Di/segni. Ledizioni, Milano, Italy, 2014. ISBN 88-6705-207-1. LCCN ????

Cassidy:2015:WHC

[Cas15] David C. Cassidy. Werner Heisenberg and Carl Friedrich Freiherr von Weizsäcker: A fifty-year friendship. Physics in Perspective (PIP), 17(1):33–54, March 2015. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-014-0152-6.

Cassidy:2017:FHG

[Cas17] David C. In Cassidy. Farm Hall and the German Atomic Project of World War II: a dramatic history. Springer International Publishing, Cham, Switzerland, 2017. ISBN 3-319-59577-6 (print), 3-319-59578-4 (e-book). xiv + 125 + 32 pp. LCCN PS3553.A87 F37 2017.

Cassidy:2022:DFH

[Cas22] David C. Cassidy. The drama of Farm Hall: a historian ventures into play writing. Berichte zur Wissenschaftsgeschichte, 45(1–2): 245–260, June 2022. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Cavecchi:2014:BHD

[Cav14] Mariacristina Cavecchi. Bohr e Heisenberg, O dell’indeterminazione. (Italian) [Bohr and Heisenberg, of the indetermination]. In Castellari [Cas14b], pages 363–376. ISBN 88-6705-207-1. LCCN ????

Cassidy:1984:WHB

[CB84] David C. Cassidy and Martha Baker. Werner Heisenberg: a bibliography of his writings, volume 9 of Berkeley papers in history of science. Office for History of Science and Technology, University of California, Berkeley, Berkeley, CA, USA, 1984. ISBN 0-918102-10-3 (paperback). ISSN 0145-0379. vi + 153 pp. LCCN Z8395.513

.C38 1984; QC16.H395. URL http://www.aip.org/history/
heisenberg/bibliography/contents.htm. See also [Rec93].

Capozziello:2020:ACH

[CBS20] Salvatore Capozziello, Micol Benetti, and Alessandro D. A. M. Spallicci. Addressing the cosmological H_{0} tension by the Heisenberg uncertainty. Foundations of Physics, 50(9):893–899, September 2020. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-020-00356-2.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Chevalley:1994:AOB

[CD94] Catherine Chevalley and Olivier Darrigol. Analyses d’ouvrages: Uncertainty: The Life and Science of Werner Heisenberg par David C. Cassidy. Revue d’Histoire des Sciences, 47(3–4):501– 503, juillet–décembre 1994. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/23633505.

Castell:1975:QTS

[CDvW75] Lutz Castell, Michael Drieschner, and Carl Friedrich Freiherr von Weizsäcker, editors. Quantum theory and the structures of time and space: papers presented at a conference held in Feldafing, July 1974, Einzelarbeiten aus dem Max-Planck-Institut zur Erforschung der Lebensbedingungen der wissenschaftlichtechnischen Welt. Carl Hanser Verlag, München, West Germany, 1975. ISBN 3-446-12075-0. LCCN QC173.96 .C66 1974. DM18.20.

Castell:1977:QTS

[CDvW77] Lutz Castell, Michael Drieschner, and Carl Friedrich Freiherr von Weizsäcker, editors. Quantum theory and the structures of time and space, volume 2; In memoriam Werner Heisenberg, Einzelarbeiten aus dem Max-Planck-Institut zur Erforschung der Lebensbedingungen der wissenschaftlich-technischen Welt. Carl Hanser Verlag, München, West Germany, 1977. ISBN 3-446-12459-4.

LCCN QC173.96 .C66 1976. URL http://catalog.hathitrust.
org/api/volumes/oclc/3619411.html.

Castell:1979:QTS

[CDvW79] L. Castell, M. Drieschner, and Carl Friedrich Freiherr von Weizsäcker, editors. Quantum theory and the structures of time and space: papers presented at a conference held in Tutzing, July 1978. Volume 3. Carl Hanser Verlag, München, West Germany, 1979. ISBN 3-446-12911-1. LCCN ????

Cesareo:2018:PSS

[Ces18] R. Cesareo. Il progetto dei servizi segreti americani di assassinare Werner Heisenberg. (Italian) [The US Secret Service’s project to assassinate Werner Heisenberg]. Physis: Rivista Internazionale di Storia della Scienza. Nuova Serie, 53(1–2):??, ???? 2018. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Craig:1998:REP

[CF98] Edward Craig and Luciano Floridi, editors. Routledge Encyclopedia of Philosophy Online. Routledge, London, UK, version 2.0 edition, 1998. ISBN 0-415-16916-X (CD-ROM), 0-415-19608-6 (user guide), 0-415-16917-8 (10 vol set + CD-ROM). LCCN B51. URL

http://www.rep.routledge.com/.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Camiz:1971:EST

[CGM^{+}71] P. Camiz, A. Gerardi, C. Marchioro, E. Presutti, and E. Scacciatelli. Exact solution of a time-dependent quantal harmonic oscillator with a singular perturbation. Journal of Mathematical Physics, 12(10):2040–2043, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i10/p2040_s1.

Celeghini:1991:QHG

[CGST91] Enrico Celeghini, Riccardo Giachetti, Emanuele Sorace, and Marco Tarlini. The quantum Heisenberg group H(1)_{q}. Journal of Mathematical Physics, 32(5):1155–1158, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i5/
p1155_s1.

Compton:1984:PPQ

[CH84] Arthur H. Compton and Werner Heisenberg. The physical principles of the quantum theory. In Blum et al. [BDR84e], pages 117– 166. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_10/.

Charles:1992:HPK

[Cha92] Dan Charles. Heisenberg’s principles kept bomb from Nazis. New Scientist, 135(1837):4–??, September 5, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http://

www.newscientist.com/article/mg13518370.300-heisenbergs-
principles-kept-bomb-from-nazis.html. See rebuttals [Pei92a,

Wel92].

Chisolm:2001:GHU

[Chi01] Eric D. Chisolm. Generalizing the Heisenberg uncertainty relation. American Journal of Physics, 69(3):368–371, March 2001. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/69/368/1.

Cho:2002:LAP

[Cho02] Adrian Cho. Letters aver physicist supported Nazi bomb. Science, 295(5558):1211, February 15, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/295/5558/1211.1.summary. See re-

buttal [Got02a].

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Ciri:1988:RGG

[Cir88] Giuseppe Ciri. Recensioni: Giuseppe Gembillo, Werner Heisenberg. La filosofia di un fisico. Studi Kantiani, 1(??):121–122, ???? 1988. ISSN 1123-4938 (print), 1724-1812 (electronic). URL http:

//www.jstor.org/stable/24344698.

Chen:1972:EEI

[CJ72a] H. H. Chen and R. I. Joseph. Erratum: “Exchange Interaction Model of Ferromagnetism”. Journal of Mathematical Physics, 13 (11):1842, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i11/p1842_s1. See [CJ72b].

Chen:1972:EIM

[CJ72b] H. H. Chen and R. K. Joseph. Exchange interaction model of ferromagnetism. Journal of Mathematical Physics, 13(5):725–739, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i5/p725_s1. See erratum [CJ72a].

Carazza:1995:HLW

[CK95] Bruno Carazza and Helge Kragh. Heisenberg’s lattice world — the 1930 theory sketch. American Journal of Physics, 63(7):595–605, July 1995. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://ajp.aapt.org/resource/1/ajpias/
v63/i7/p595_s1.

Chiao:2002:HICa

[CK02a] R. Y. Chiao and P. G. Kwiat. Heisenberg’s introduction of the ‘collapse of the wavepacket’ into quantum mechanics. Fortschritte der Physik = Progress of Physics, 50(5–7):614–623, ???? 2002. CO-DEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Chiao:2002:HICb

[CK02b] Raymond Y. Chiao and Paul G. Kwiat. Heisenberg’s introduction of the collapse of the wavepacket into quantum mechanics. In Papenfuss et al. [PLS02], pages 185–194. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Cohen:1983:PPP

[CL83] R. S. Cohen and L. Laudan, editors. Physics, Philosophy and Psychoanalysis: Essays in Honour of Adolf Grünbaum, volume 76 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN 0068-0346. xviii + 339 pp. LCCN ???? URL https://link.springer.

com/book/10.1007/978-94-009-7055-7.

Clauser:1994:HMD

[CL94] John F. Clauser and Shifang Li. “Heisenberg microscope” decoherence atom interferometry. Physical Review A (Atomic, Molecular, and Optical Physics), 50(3):2430–2433, September 1994. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://pra.aps.org/abstract/PRA/v50/i3/p2430_
1.

Capelle:2005:SDF

[CL05] K. Capelle and Valter L. Líbero. Spin-density functional theory: Some open problems and application to inhomogeneous Heisenberg models. International Journal of Quantum Chemistry, 105 (6):679–686, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clary:2023:FMR

[Cla23] David C. Clary. Foreign membership of the Royal Society: Schrödinger and Heisenberg? Notes and Records of the Royal Society of London, 77(3):513–536, September ??, 2023. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Cleveland:2006:HW

[Cle06] Cutler J. Cleveland. Heisenberg, Werner. The Encyclopedia

of Earth, 2006. URL http://www.eoearth.org/view/article/
153467/.

Clerk:2009:MFH

[Cle09] Aashish Clerk. Measurement: Facing Heisenberg at the nanoscale. Nature Nanotechnology, 4(12):796–798, December 2009. CO-DEN NNAABX. ISSN 1748-3387 (print), 1748-3395 (elec-

tronic). URL http://www.nature.com/nnano/journal/v4/n12/
full/nnano.2009.368.html.

Cline:1965:QPQ

[Cli65] Barbara Lovett Cline. The questioners: physicists and the quantum theory. Crowell, New York, 1965. vii + 274 pp.

LCCN QC15 .C4. URL http://catalog.hathitrust.org/api/
volumes/oclc/372589.html.

Clusius:1937:NBW

[Clu37] K. Clusius. Neue Bücher: Wandlungen in den Grundlagen der Naturwissenschaft. Von Werner Heisenberg. 2. erweiterte Auflage. Verlag S. Hirzel, Leipzig 1936. Preis kart. RM. 2,50. Angewandte Chemie, 50(1):60, 1937. CODEN ANCEAD. ISSN 1521-3757.

Chaturvedi:2002:SCI

[CM02] S. Chaturvedi and N. Mukunda. The Schwinger SU(3) construction. II. Relations between Heisenberg–Weyl and SU(3) coherent states. Journal of Mathematical Physics, 43(11):5278–5309, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Compton:1931:UPF

[Com31] Arthur H. Compton. The Uncertainty Principle and free will. Science, 74(1911):172, August 14, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/74/1911/172.1.full.pdf. Comment

on [Dar31]. See further comments in [Mar31].

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Condon:1929:RUP

[Con29] E. U. Condon. Remarks on Uncertainty Principles. Science, 69 (1796):573–574, May 31, 1929. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/69/1796/573.full.pdf.

Condon:1962:YQP

[Con62] Edward U. Condon. 60 years of quantum physics. Physics Today, 15(10):37–49, October 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/PHTOAD/v15/i10. Delayed 1951 Presidential ad-

dress at the 1500th regular meeting of the American Philosophical Society of Washington, 2 December 1962, at the Natural History Museum Auditorium of the Smithsonian Institution, on the 60th anniversary of Planck’s constant, h. Reprinted in [WP85, pages 310–318].

Cope:1981:RUD

[Cop81] F. W. Cope. On the relativity and uncertainty of distance, time, and energy measurements by man. 1. Derivation of the Weber Psychophysical Law from the Heisenberg Uncertainty Principle applied to a superconductive biological detector. 2. The reverse derivation. 3. A human theory of Relativity. Physiological Chemistry and Physics, 13(4):305–311, ???? 1981. CODEN PLCHB4. ISSN 0031-9325.

Correspondent:1971:PHR

[Cor71] A. Correspondent. Physics: Heisenberg retires. Nature, 229(5280): 80, January 8, 1971. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v229/n5280/pdf/229080b0.pdf.

Corben:1979:ALT

[Cor79] H. C. Corben. Another look through the Heisenberg microscope. American Journal of Physics, 47(12):1036–1037, December 1979. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://ajp.aapt.org/resource/1/ajpias/
v47/i12/p1036_s1.

Cornwell:2003:HSS

[Cor03] John Cornwell. Hitler’s scientists: science, war, and the Devil’s pact. Viking Press, New York, NY, USA, 2003. ISBN 0-670-03075-9. xvi + 535 pp. LCCN Q127.G3 C67 2003.

Cowley:1949:UPa

[Cow49a] W. L. Cowley. Uncertainty Principle. Nature, 163(4143):492, March 26, 1949. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v163/n4143/pdf/163492a0.pdf.

Cowley:1949:UPb

[Cow49b] W. L. Cowley. The Uncertainty Principle. Nature, 164(4162): 246, August 6, 1949. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v164/n4162/pdf/164246a0.pdf.

Cowling:1984:BVT

[CP84] Michael G. Cowling and John F. Price. Bandwidth versus time concentration: the Heisenberg–Pauli–Weyl inequality. SIAM Journal on Mathematical Analysis, 15(1):151–165, January 1984. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Cahn:1993:TBT

[CP93] Anne Hessing Cahn and John Prados. Team-B: the trillion dollar experiment. Bulletin of the Atomic Scientists, 49(3):22–31, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-13610752.

html. See corrections [Ano93a].

Cassidy:1984:BDW

[CR84] David C. Cassidy and Helmut Rechenberg. Biographical data: Werner Heisenberg (1901–1976). In Blum et al. [BDR84a], pages 1–16. ISBN 0-387-13400-X (A part 1), 0-387-13847-1 (A part 2). LCCN QC3 .H33 1984. URL http:

//link.springer.com/chapter/10.1007/978-3-642-61742-3_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
642-61742-3/1.pdf.

Clark:2012:LAN

[CR12] Charles W. Clark and Joseph Reader. Light, atoms and nuclei: The optical discovery of deuterium. Optics and Photonics News, 23(5):36–41, May 2012. CODEN OPPHEL. ISSN 1047-6938 (print), 1541-3721 (electronic). URL https://www.osa-opn.org/

home/articles/volume_23/issue_5/features/light,_atoms_
and_nuclei_the_optical_discovery_of/. See also [?].

Caciuffo:2013:YRI

[CR13] Roberto Caciuffo and Joseph Richter. 50 years of research at the Institute for Transuranium Elements in Karlsruhe, volume 25925 of EUR. Publications Office of the European Union, Luxembourg, 2013. ISBN 92-79-29459-8. viii + 199 pp. LCCN ???? URL

http://bookshop.europa.eu/en/50-years-of-research-at-
the-institute-for-transuranium-elements-in-karlsruhe-

pbLCNA25925/. See also [KH19] for an account of one of the German uranium cubes found in 2013 in the USA, and [MWL^{+}11, MWL^{+}15] for nuclear forensics of the reactor cubes.

Crease:2010:GEB

[Cre10] Robert P. Crease. The Great Equations: Breakthroughs in Science from Pythagoras to Heisenberg. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33793-6. 315 pp. LCCN Q175 .C884 2010.

Choi:1993:LAR

[CRR93] Man Duen Choi, Jeffrey S. Rosenthal, and Peter Rosenthal. Linear-algebraic results associated with antiferromagnetic Heisenberg chains. SIAM Journal on Matrix Analysis and Applications, 14(3):830–852, July 1993. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Crull:2023:RKK

[Cru23] Elise Crull. Review of Konrad Kleinknecht’s Werner Heisenberg: Reality and Its Order — Konrad Kleinknecht, ed., Werner Heisenberg: Reality and Its Order, with introduction by H. Rechenberg and commentary by E. P. Fischer, translated by M. B. Rumscheidt, N. Lukens, and I. Heisenberg. Cham, Switzerland: Springer (2019), 148 pp., $129.99. Philosophy of Science, 90(4):1010–1013, October 2023. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL

https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-konrad-kleinknechts-werner-
heisenberg-reality-and-its-order-konrad-kleinknecht-ed-
werner-heisenberg-reality-and-its-order-with-introduction-
by-h-rechenberg-and-commentary-by-e-p-fischer-translated-

by-m-b-rumscheidt-n-lukens-and-i-heisenberg-cham-switzerland-
springer-2019-148-pp-12999/0C9AC6F396502302FB501A1DE763F381.

Cohen:1979:SPL

[CS79] R. S. (Robert Sonné) Cohen and John J. Stachel, editors. Selected Papers of Léon Rosenfeld, volume 21 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1979. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (ebook). ISSN 0068-0346. xxxiv + 929 pp. LCCN Q174 .B67

vol. 21 QC7. URL https://link.springer.com/book/10.1007/
978-94-009-9349-5.

Carson:1994:RBS

[CS94] Cathryn Carson and Silvan S. Schweber. Recent biographical studies in the physical sciences: David C. Cassidy: Uncertainty: The Life and Science of Werner Heisenberg. Ronald R. Kline: Steinmetz: Engineer and Socialist. Albert E. Moyer: A Scientist’s Voice in American Culture: Simon Newcomb and the Rhetoric of Scientific Method. Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham: Harriet Brooks: Pioneer Nuclear Scientist. John M. Thomas and David Phillips: Selections and Reflections: The Legacy of Sir Lawrence Bragg. Victor Weisskopf: The Joy of Insight: Passions of a Physicist. Isis, 85(2):284–292, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236495.

Culetu:2005:HP

[Cul05] Hristu Culetu. Λ and the Heisenberg principle. Foundations of Physics, 35(9):1511–1519, September 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/s10701-005-6479-8.

Cushing:1986:IHM

[Cus86] James T. Cushing. The importance of Heisenberg’s S-matrix program for the theoretical high-energy physics of the 1950’s. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(2):110–149, June 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Cushing:1994:QMH

[Cus94] James T. Cushing. Quantum mechanics: historical contingency and the Copenhagen hegemony. Science and its conceptual foundations. University of Chicago Press, Chicago, IL, USA, 1994. ISBN 0-226-13202-1, 0-226-13204-8 (paperback). xvi + 317 pp. LCCN QC173.98 .C87 1994. URL http://www.loc.

gov/catdir/description/uchi051/94008427.html;   http:/
/www.loc.gov/catdir/toc/uchi051/94008427.html.

Cohen:1985:PTF

[CW85] Robert S. Cohen and Marx W. Wartofsky, editors. A Portrait of Twenty-five Years: Boston Colloquium for the Philosophy of Science 1960–1985, volume 89 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1985. ISBN 90-277-1971-3, 94-009-5345-3 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-009-5345-1.

Cecchini:1981:SRC

[CZ81] C. Cecchini and A. Zappa. Some results on the center of an algebra of operators on VN(G) for the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 33 (??):1469–1486, ???? 1981. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Durr:1992:WH

[D^{+}92] H.-P. (Hans-Peter) Dürr et al., editors. Werner Heisenberg. Hanser, München, Germany, 1992. ISBN 3-446-17217-3. 123 pp. LCCN QC16.H35 W47 1992.

DAriano:2003:HPN

[D’A03] G. M. D’Ariano. On the Heisenberg principle, namely on the information-disturbance trade-off in a quantum measurement. Fortschritte der Physik = Progress of Physics, 51(4–5):318–330, May 2003. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Dahl:1999:HWW

[Dah99] Per F. Dahl. Heavy Water and the Wartime Race for Nuclear Energy. Institute of Physics Publishers, Bristol, UK, 1999. ISBN 0-7503-0633-5 (hardcover). xvi + 399 pp. LCCN QD169.W3

D25 1999. URL http://www.loc.gov/catdir/enhancements/
fy0661/99033672-d.html.

Dahl:2002:BHCa

[Dah02a] J. P. Dahl. The Bohr–Heisenberg Correspondence Principle viewed from phase space. Fortschritte der Physik = Progress of Physics, 50(5–7):630–635, ???? 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Dahl:2002:BHCb

[Dah02b] Jens P. Dahl. The Bohr-Heisenberg Correspondence Principle viewed from phase space. In Papenfuss et al. [PLS02], pages 201– 206. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Dahn:2022:FHT

[Dah22] Ryan Dahn. The Farm Hall transcripts: The smoking gun that wasn’t. Berichte zur Wissenschaftsgeschichte, 45(1–2):202–218, June 2022. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Darwin:1931:UP

[Dar31] Charles Galton Darwin. The Uncertainty Principle. Science, 73 (1903):653–660, June 19, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/73/1903/653.full.pdf. See comments [Com31,

Mar31, Noy31, Fli32, Fli33, Ano35c, New35] and later papers [Cow49a, Cow49b, SR49, Gam58, Žup65].

deBroglie:1958:TRE

[dB58] Louis de Broglie. Tentative de raccord entre l’équation de Heisenberg et l’équation de l’onde u en théorie de la double solution. (French) [Attempt at connecting the Heisenberg equation to the wave equation u in the double-solution theory]. Comptes rendus de l’Académie des sciences, Paris, 246(??):2077–??, ???? 1958. ISSN 0001-4036 (print), 2419-6304 (electronic).

Durr:1971:QFP

[DB71] Hans-Peter Dürr and Erich Bagge, editors. Quanten und Felder: Physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg. (German) [Quanta and fields: physical and philosophical reflections on the 70th birthday of Werner Heisenberg]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 3-528-08317-4. 366 pp.

LCCN ???? URL http://www.springer.com/physics/book/
978-3-528-08317-5.

deBroglie:1982:IHI

[dB82] Louis de Broglie. Les incertitudes d’Heisenberg et l’interprétation probabiliste de la mécanique ondulatoire. (French) [Heisenberg uncertainties and the probabilistic interpretation of wave mechanics]. Gauthier-Villars et cie, Paris, France, 1982. ISBN 2-04-015411-6. xlii + 304 pp. With a preface and with notes by Georges Lochak.

deBroglie:1990:HUP

[dB90] Louis de Broglie. Heisenberg’s Uncertainty Principle and the Probabilistic Interpretation of Wave Mechanics. Number 40 in Fundamental Theories of Physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0929-4. xliv + 302 pp. LCCN QC174.2 .B653513 1990. English translation by Alwyn van der Merwe.

Destouches:1948:CEC

[DC48] Jean-Louis Destouches and Michel Cazin. Corrélation entre la complémentarité de Louis de Broglie et la matrice caractéristique de Heisenberg. (French) [Correlation between the complementarity of Louis de Broglie and the characteristic matrix of Heisenberg]. Comptes rendus de l’Académie des sciences, Paris, 226(??):566– 567, ???? 1948. ISSN 0001-4036 (print), 2419-6304 (electronic).

De:1987:IPP

[DC87] Archan K. De and A. Roy Chowdhury. On the inverse problem and prolongation structure for the modified anisotropic Heisenberg spin chain. Journal of Mathematical Physics, 28(2):319– 322, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p319_s1.

Dou:2013:GHS

[DD13] Yan-Ni Dou and Hong-Ke Du. Generalizations of the Heisenberg and Schrödinger uncertainty relations. Journal of Mathematical Physics, 54(10):103508, October 2013. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dopel:1940:BDTa

[DDH40a] Robert Döpel, Klara Döpel, and Werner Heisenberg. Bestimmung der Diffusionslänge thermischer Neutronen in Präparat 38. (German) [Determination of the diffusion length of thermal neutrons in preparation 38]. Kernphysikalische Forschungsberichte G-22, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, December 5, 1940. 5 pp.

Dopel:1940:BDTb

[DDH40b] Robert Döpel, Klara Döpel, and Werner Heisenberg. Bestimmung der Diffusionslänge thermischer Neutronen in schwerem Wasser. (German) [Determination of the diffusion length of thermal neutrons in heavy water]. Kernphysikalische Forschungsberichte G-23, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, August 7, 1940. 8 pp.

Dopel:1941:VSD

[DDH41a] Robert Döpel, Klara Döpel, and Werner Heisenberg. Versuche mit einer Schichtenanordnungen von D_{2}O und Präp[arat] 38. (German) [Experiments with a layer arrangement of D_{2}O and preparation 38]. Kernphysikalische Forschungsberichte G-75, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, October 28, 1941. 18 pp.

Dopel:1941:VSW

[DDH41b] Robert Döpel, Klara Döpel, and Werner Heisenberg. Versuche mit Schichtenanordnungen von Wasser und Präp[arat] 38. (German) [Experiments with a layer arrangement of water and preparation 38]. Kernphysikalische Forschungsberichte G-74, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, April 28, 1941. 13 pp.

Dopel:1942:ENE

[DDH42a] Robert Döpel, Klara Döpel, and Werner Heisenberg. Der experimentelle Nachweis der effektiven Neutronenvermehrung in einem Kugel-Schichten-System aus D_{2}O und Uran-Metall. (German) [The experimental proof of the effective neutron multiplication in a ball-tier system of D_{2}O and uranium metal]. Kernphysikalische Forschungsberichte G-136, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, July 1942. 8 pp.

Dopel:1942:NMD

[DDH42b] Robert Döpel, Klara Döpel, and Werner Heisenberg. Die Neutronenvermehrung in 38-Metall durch rasche Neutronen. (German) [Neutron multiplication in 38 metal by fast neutrons]. Technical Report G-137, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, February 26–28, 1942. 7 pp.

Dopel:1942:NED

[DDH42c] Robert Döpel, Klara Döpel, and Werner Heisenberg. Die Neutronenvermehrung in einem D2O-38-Metallschichtensystem. (German) [The neutron multiplication in a D_{2}O-38-metal layer system]. Kernphysikalische Forschungsberichte G-373, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, March 1942. 3 pp.

Dopel:1989:ENE

[DDH89a] R. Döpel, K. Döpel, and Werner Heisenberg. Der experimentelle Nachweis der effektiven Neutronenvermehrung in einem Kugel-Schichten System aus D_{2}O und Uran-Metall (Jul. 1942). (German) [Neutron multiplication in a ball-layer system composed of D_{2}O and uranium metal (July 1942)]. In Blum et al. [BDR89a], pages 536–544. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:NMD

[DDH89b] R. Döpel, K. Döpel, and Werner Heisenberg. Die Neutronenvermehrung in 38-Metall durch rasche Neutronen (26.-28. Feb. 1942). (German) [Neutron multiplication in 38 metal by fast neutrons (26–28 February 1942)]. In Blum et al. [BDR89a], pages 529–535. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:NED

[DDH89c] R. Döpel, K. Döpel, and Werner Heisenberg. Die Neutronenvermehrung in einem D_{2}O-38-Metallschichtensystem (26.–28. Feb. 1942). (German) [Neutron multiplication in a D_{2}O-38-metal layer system (26 to 28 February 1942)]. In Blum et al. [BDR89a], pages 526–528. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:BDTb

[DDH89d] Robert Döpel, Klara Döpel, and Werner Heisenberg. Bestimmung der Diffusionslänge thermischer Neutronen in Präparat 38 (5 Dec. 1940). (German) [Determination of the diffusion length of thermal neutrons in preparation 38 (5 December 1940)]. In Blum et al. [BDR89a], pages 427–431. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:BDTa

[DDH89e] Robert Döpel, Klara Döpel, and Werner Heisenberg. Bestimmung der Diffusionslänge thermischer Neutronen in schwerem Wasser (7 Aug 1940). (German) [Determination of the diffusion length of thermal neutrons in heavy water (7 Aug 1940)]. In Blum et al. [BDR89a], pages 419–426. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:VSD

[DDH89f] Robert Döpel, Klara Döpel, and Werner Heisenberg. Versuche mit einer Schichtenanordnungen von D_{2}O und Präp[arat] 38 (28. Okt. 1941). (German) [Experiments with a layer arrangement of D_{2}O and preparation 38 (28 October 1941)]. In Blum et al. [BDR89a], pages 481–498. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Dopel:1989:VSW

[DDH89g] Robert Döpel, Klara Döpel, and Werner Heisenberg. Versuche mit Schichtenanordnungen von Wasser und Präp[arat] 38 (28 Apr. 1941). (German) [Experiments with a layer arrangement of water and preparation 38 (28 April 1941)]. In Blum et al. [BDR89a], pages 463–475. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

DeCat:2003:TST

[DDI03] Tine De Cat, Karel Dekimpe, and Paul Igodt. Translations in simply transitive affine actions of Heisenberg type Lie groups. Linear Algebra and its Applications, 359(1–3):101– 111, January 15, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.elsevier.

nl/gej-ng/10/30/19/208/25/34/abstract.html;   http://
www.sciencedirect.com/science?_ob=GatewayURL&_origin=
SOCJLA&_urlversion=4&_method=citationSearch&_version=
1&_piikey=S0024379502004354&_volkey=00243795%23359%23101&
_refkey=DeCat%232003%23101%23111&md5=549eb30683727b5fb34d937aa95c2f99.

dEspagna:1992:RHO

[d’E92] B. d’Espagna. La représentation de Heisenberg et l’opération de mesure. In Bitbol and Darrigol [BD92], pages 421–424. ISBN 2-86332-116-1. LCCN QC16.S265 E78 1992. 263.75F.

DeGrand:1974:HES

[DeG74] Thomas A. DeGrand. Heisenberg’s explosion shower and the Froissart bound. American Journal of Physics, 42(10):862–865, October 1974. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/42/862/1.

Dehlinger:1936:BMQ

[Deh36] U. Dehlinger. Buchbesprechung: Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von De Broglie, Schrödinger, Heisenberg und Dirac. 4. und 5. Aufl. Von Arthur Haas. Leipzig 1934, Akademische Verlagsgesellschaft. 299 S., 7 Abb. Preis geb. RM 7.80. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 42(1):43, 1936. CODEN ZEA-PAA. ISSN 0005-9021.

Dehlinger:1937:BWG

[Deh37] U. Dehlinger. Buchbesprechung: Wandlungen in den Grundlagen der Naturwissenschaften. Drei Vorträge von Werner Heisenberg.

2. erweiterte Aufl. Verlag von S. Hirzel. Leipzig 1936. 63 S. Preis kart. R. M. 2.50. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 43(10):849, 1937. CODEN ZEAPAA. ISSN 0005-9021.

Deltete:2013:HAA

[Del13] Robert J. Deltete. Heisenberg in the atomic age: Science and the public sphere. Annals of Science, 70(3):419–422, 2013. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “The letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CO-DEN RQSCAN. ISSN 0035-2160. URL https://www.unamur.

be/sciences/philosoc/revueqs/textes-en-ligne/rqs_186_
1_demortier.

Djakov:1996:OEI

[DF96] B. B. D’jakov and V. Ja. Frenkel’. Operacija “Epsilon” ili konec nemeckogo uranovogo proekta. (Russian) [Operation Epsilon and the end of the German uranium project], volume 1677 of Preprint / Fiziko-Techniceskij Institut im. A. F. Ioffe, Rossijskaja Akademija Nauk. FTI, Sankt-Peterburg, Russia, 1996. 60 pp. LCCN ????

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp.

LCCN Q141 .D52 1994. URL http://www.loc.gov/catdir/
enhancements/fy0637/99488304-d.html;   http://www.loc.
gov/catdir/enhancements/fy0637/99488304-t.html;   http:
//www.netLibrary.com/urlapi.asp?action=summary&v=1&bookid=
12306.

deGosson:2014:BJQ

[dG14] Maurice A. de Gosson. Born–Jordan quantization and the equivalence of the Schrödinger and Heisenberg pictures. Foundations of Physics, 44(10):1096–1106, October 2014. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10701-014-9831-z.pdf.

DiDomenico:2012:HUR

[DGH^{+}12] Antonio Di Domenico, Andreas Gabriel, Beatrix C. Hiesmayr, Florian Hipp, Marcus Huber, Gerd Krizek, Karoline Mühlbacher, Sasa Radic, Christoph Spengler, and Lukas Theussl. Heisenberg’s uncertainty relation and Bell inequalities in high energy physics. Foundations of Physics, 42(6):778–802, June 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-011-9575-y.

Ding:1999:MVS

[DGS99] Shijin Ding, Boling Guo, and Fengqiu Su. Measure-valued solution to the strongly degenerate compressible Heisenberg chain equations. Journal of Mathematical Physics, 40(3):1153–1162, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dirac:1932:VPQ

[DH32] P. A. M. Dirac and W. Heisenberg. Various Problems in Quantum Theory. Kei Mekai, Tokyo, Japan, 1932. ???? pp. LCCN ???? Based on notes of lectures held in Tokyo on September 2–7, 1929. Translated by Y. Nishina, M. Kotani, and T. Inui.

Durr:1961:TSTa

[DH61a] H. P. Dürr and W. Heisenberg. Zur Theorie der ‘seltsamen’ Teilchen. (German) [On the theory of ‘strange’ particles]. Zeitschrift für Naturforschung, 16a(8):726–747, ???? 1961. CO-DEN ZNTFA2. ISSN 0372-9516. URL http://labs.adsabs.

harvard.edu/ui/abs/1961ZNatA..16..726D.

Durr:1961:TSP

[DH61b] H.-P. Dürr and Werner Heisenberg. Theory of the strange particles. In P. Falk-Vairant E. Cremieu-Alcan and O. Lebey., editors, The Aix-en-Provence International Conference on Elementary Particles, Volume 1: Parallel Sessions. Proceedings of the conference held September 14–20, 1961, pages 163–?? C. E. N. Saclay, Gif-sur-Yvette, France, 1961. LCCN ???? URL http://

labs.adsabs.harvard.edu/ui/abs/1961elpa.conf..163D.

Durr:1962:QNM

[DH62] H. P. Dürr and W. Heisenberg. The quantum numbers of the ωmeson. Il Nuovo Cimento (10), 23(5):807–816, March 1, 1962. CO-DEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Durr:1965:STS

[DH65a] H.-P. Dürr and W. Heisenberg. On the “spurion” theory of strange particles. Il Nuovo Cimento (10), 37(4):1446–1486, June 16, 1965. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Durr:1965:SPC

[DH65b] H.-P. Dürr and W. Heisenberg. Strange-particle coupling constants. Il Nuovo Cimento (10), 37(4):1487–1510, June 16, 1965.

CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Durr:1966:SPN

[DH66] H. P. Dürr and Werner Heisenberg. Strange particles in the nonlinear spinor theory. In Ya. Smorodinskii, V. G. Grishin, A. A. Komar, V. A. Nikitin, L. D. Solovyev, and S. M. Korenchenko, editors, XII International Conference on High Energy Physics, Dubna, August 5–15, 1964. Volume 2, pages 245–247. Atomizdat, Moscow, USSR, 1966. URL ????.

Durr:1984:SPN

[DH84a] H. P. Dürr and W. Heisenberg. Strange particles in the nonlinear spinor theory. In Blum et al. [BDR84e], pages 634– 636. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_58/.

Durr:1984:TSP

[DH84b] H.-P. Durr and W. Heisenberg. Theory of the strange particles. In Blum et al. [BDR84e], pages 592–597. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_52/.

Durr:1959:TEG

[DHM^{+}59] H.-P. Dürr, W. Heisenberg, H. Mitter, S. Schlieder, and K. Yamazaki. Zur Theorie der Elementarteilchen. (German) [On the theory of elementary particles]. Zeitschrift für Naturforschung, 14a:441–485, 1959. CODEN ZNTFA2. ISSN 0372-9516.

Durr:1965:QEN

[DHYY65] H.-P. Dürr, W. Heisenberg, H. Yamamoto, and K. Yamazaki. Quantum electrodynamics in the nonlinear spinor theory and the value of Sommerfeld’s fine-structure constant. Il Nuovo Cimento (10), 38(3):1220–1242, August 1, 1965. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). See remarks [Hei64a].

Dirac:1928:DFR

[Dir28] P. A. M. Dirac. Discussion following reports of Bohr and Heisenberg. In ????, editor, Electrons and photons (5th Conseil de physique de l’Institut International de Physique, Solvay, 24–29 October 1927, Bruxelles), pages 258–263. Gauthier-Villars et cie, Paris, France, 1928. LCCN ????

Dirac:1964:ESH

[Dir64] P. A. M. Dirac. Equivalence of the Schrödinger and Heisenberg pictures. Nature, 204(4960):772, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v204/
n4960/pdf/204772a0.pdf.

Dirac:1976:HIP

[Dir76] P. A. M. Dirac. Heisenberg’s influence on physics. Commentarii — Pontificia Academia Scientiarum, 3(14):1–15, ???? 1976. CODEN CPOSAF. ISSN 0554-6648.

Djerassi:2007:BRW

[Dje07] Carl Djerassi. Book reviews: When acting speaks louder than words: Science on stage: From ‘Doctor Faustus’ to ‘Copenhagen’, by Kirsten Shepherd-Barr. Google’s PageRank and Beyond: The Science of Search Engine Rankings, by Amy N. Langville and Carl D. Meyer. Broken Genius The Rise and Fall of William Shockley, Creator of the Electronic Age, by Joel N. Shurkin. Physics Today, 60(2):63–64, 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v60/i2/p63/s1.

DeYoung:1993:EVH

[DJR93] P. A. DeYoung, P. L. Jolivette, and N. Rouze. Experimental verification of the Heisenberg uncertainty principle — an advanced undergraduate laboratory. American Journal of Physics, 61(6): 560–563, June 1993. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?AJP/
61/560/1.

Davis:1978:FCS

[DK78] M. C. Davis and G. F. Knoll. Fission cross-sections of ^{235}U and 239Pu averaged over ^{252}Cf neutron spectrum. Annals of Nuclear Energy, 5(11–12):583–588, January 1978. CODEN ANENDJ. ISSN 0306-4549 (print), 1873-2100 (electronic).

Daniel:1992:SSA

[DKLN92] M. Daniel, M. D. Kruskal, M. Lakshmanan, and K. Nakamura. Singularity structure analysis of the continuum Heisenberg spin chain with anisotropy and transverse field: Nonintegrability and chaos. Journal of Mathematical Physics, 33(2):771–776, February 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i2/p771_s1.

Ding:2008:TPD

[DL08] Qing Ding and Wei Lin. The transmission property of the discrete Heisenberg ferromagnetic spin chain. Journal of Mathematical Physics, 49(9):093501, September 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v49/i9/p093501_s1.

delaPena:1980:CIG

[dlP80] Luis de la Peña. Conceptually interesting generalized Heisenberg inequality. American Journal of Physics, 48(9):775–776, September 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/48/775/1.

Dyson:1976:PTQ

[DLS76] Freeman J. Dyson, Elliott H. Lieb, and Barry Simon. Phase transitions in the quantum Heisenberg model. Physical Review Letters, 37(3):120–123, 1976. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Dattoli:2008:HAU

[DLW08] G. Dattoli, D. Levi, and P. Winternitz. Heisenberg algebra, umbral calculus and orthogonal polynomials. Journal of Mathematical Physics, 49(5):053509, May 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v49/i5/p053509_s1.

Daniel:1999:GEC

[DM99] M. Daniel and K. Manivannan. On the geometric equivalence of certain discrete integrable Heisenberg ferromagnetic spin chains. Journal of Mathematical Physics, 40(6):2560–2567, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v40/i6/p2560_s1.

DeChiara:2006:EED

[DMCF06] Gabriele De Chiara, Simone Montangero, Pasquale Calabrese, and Rosario Fazio. Entanglement entropy dynamics of Heisenberg chains. Journal of Statistical Mechanics: Theory and Experiment, 3:14, 2006. CODEN JSMTC6. ISSN 1742-5468.

Damak:2006:TAS

[DMdA06] Mondher Damak, Marius Mantoiu, and Rafael Tiedra de Aldecoa. Toeplitz algebras and spectral results for the one-dimensional Heisenberg model. Journal of Mathematical Physics, 47(8):082107, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v47/i8/p082107_s1.

Dobson:1977:ERS

[Dob77] J. Dobson. Exact results for second-neighbor surface magnons in an FCC lattice. Journal of Mathematical Physics, 18(1):116– 117, January 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v18/i1/p116_s1.

Dopel:1942:BUB

[Döp42] Robert Döpel. Bericht über Unfälle beim Umgang mit Uranmetall. (German) [Report on accidents when handling uranium metal]. Kernphysikalische Forschungsberichte G-135, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, July 9, 1942.

Dorries:2005:MFC

[Dör05] Matthias Dörries, editor. Michael Frayn’s Copenhagen in debate: historical essays and documents on the 1941 meeting between Niels Bohr and Werner Heisenberg, volume 20 (Berkeley); 33 (Uppsala) of Berkeley papers in history of science; Uppsala studies in history of science. Office for History of Science and Technology, University of California, Berkeley, Berkeley, CA, USA, 2005. ISBN 0-9672617-2-4 (paperback). vii + 195 pp. LCCN PR6056.R3 C6636 2005; PR6056.R3 C6436 2005; PR6056.R3 C6765 2005.

Daniel:1994:IIS

[DPL94] M. Daniel, K. Porsezian, and M. Lakshmanan. On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions. Journal of Mathematical Physics, 35 (12):6498–6510, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v35/i12/p6498_s1. See erra-

tum [DPL96].

Daniel:1996:EII

[DPL96] M. Daniel, K. Porsezian, and M. Lakshmanan. Erratum: “On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions”, [J. Math. Phys. 35 (1994), no. 12, 6498–6510; MR 95h:82022]. Journal of Mathematical Physics, 37(9):4768, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [DPL94].

Dietz:1992:LNS

[DR92] Chris Dietz and Ernest A. Ryavec. Letter: “Nazi science” debate continued. Bulletin of the Atomic Scientists, 48(9):46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GP92b].

deRegt:1999:PVH

[dR99] Henk W. de Regt. Pauli versus Heisenberg: a case study of the heuristic role of philosophy. Foundations of Science, 4(4):405–426, ???? 1999. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-

8471 (electronic). URL http://link.springer.com/article/
10.1023/A:1009643801616.

Drumheller:1986:IMM

[DRB86] John E. Drumheller, David Raffaelle, and Mark Baldwin. An improved mechanical model to demonstrate the first- and secondorder phase transitions of the easy-axis Heisenberg antiferromagnet. American Journal of Physics, 54(12):1130–1133, December 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://link.aip.org/link/?AJP/54/1130/1.

Dresden:1992:BRD

[Dre92] Max Dresden. Book review: David C. Cassidy, Uncertainty: The Life and Science of Werner Heisenberg. Physics Today, 45(6):79– ??, June 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v45/i6/p79/s1.

Dreisigacker:1990:ABFa

[DRF90a] E. Dreisigacker, F. Rieß, and G. Fischer. Aktuelles: Berliner Forschungsreaktor weiterhin umstritten/Internationale Fachausstellung Optik 1990 in Berlin/CERN: Finnland wird Mitglied/ Neue Zeitschrift der Moskauer Physikalischen Gesellschaft/ Bergakademie Freiberg richtet Studiengang Physik ein/Gesamtdeutsche DFG?/Deutsches Museum ehrt Meitner und Heisenberg/DVT-Preis Technik und Öffentlichkeit /Das Raster-Tunnel-Mikroskop als Pinzette für Atome/Europäisches Astronautenzentrum in Köln/20,7 Tesla mit supraleitendem Magneten erreicht!/Zentrum für hochauflösende Elektronenmikroskopie im Forschungszentrum Jülich eingeweiht/Studienförderpreis der WE-Heraeus-Stiftung/ Jürgen-Geiger-Preis/Verwundbarkeit der Industrielandschaft Europa/Risikostudie Kernkraftwerke/Computer Workshop IV ein Bericht. Physikalische Blätter, 46(6):189–198, June 1990. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

URL http://onlinelibrary.wiley.com/doi/10.1002/phbl.
19900460614/abstract.

Dreisigacker:1990:ABFb

[DRF90b] E. Dreisigacker, F. Rieß, and G. Fischer. Aktuelles: Berliner Forschungsreaktor weiterhin umstritten/Internationale Fachausstellung Optik 1990 in Berlin/CERN: Finnland wird Mitglied/ Neue Zeitschrift der Moskauer Physikalischen Gesellschaft/ Bergakademie Freiberg richtet Studiengang Physik ein/Gesamtdeutsche DFG?/Deutsches Museum ehrt Meitner und Heisenberg/DVT Preis Technik und Öffentlichkeit /Das Raster Tunnel Mikroskop als Pinzette für Atome/Europäisches Astronautenzentrum in Köln/20,7 Tesla mit supraleitendem Magneten erreicht!/Zentrum für hochauflösende Elektronenmikroskopie im Forschungszentrum Jülich eingeweiht/Studienförderpreis der WE Heraeus Stiftung/ Jürgen Geiger Preis/Verwundbarkeit der Industrielandschaft Europa/Risikostudie Kernkraftwerke/Computer Workshop IV — ein Bericht. (German) [News: …]. Physik Journal, 46(6):189–198, June 1990. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Dehlinger:1943:WGN

[DS43] U. Dehlinger and G. V. Schulz. Wandlungen in den Grundlagen der Naturwissenschaft. Sechs Vorträge von Werner Heisenberg. 3., erweiterte Auflage. 95 Seiten. Verlag von S. Hirzel, Leipzig 1942. Preis kart. RM 3.50. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 49(8):406, 1943. CODEN ZEAPAA. ISSN 0005-9021.

deSouza:2002:RHT

[dSL02] Adauto J. F. de Souza and M. L. Lyra. Reweighting histogram technique for the quantum Heisenberg ferromagnet. Computer Physics Communications, 146(1):16–23, June 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465502004307.

Dale-Trotter:2018:BRQ

[DT18] Iain Dale-Trotter. Book review: The quantum heretics: What is Real: the Unfinished Quest for the Meaning of Quantum Physics by Adam Becker. Physics World, 31(12):??, December 2018. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL https://physicsworld.com/a/the-quantum-heretics/.

Dunkl:1986:BVP

[Dun86] Charles F. Dunkl. Boundary value problems for harmonic functions on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):478–512, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Durrenmatt:1962:PKZ

[Dür62] Friedrich Dürrenmatt. Die Physiker: Eine Komödie in zwei Akten. (German) [The Physicists: A comedy in two acts]. Verlag Die Arche, Zürich, Switzerland, 1962. 71 pp. LCCN ????

Durr:1976:PAW

[Dür76] Hans Peter Dürr. Persönliches: Abschied von Werner Heisenberg. (German) [Personal: Farewell to Werner Heisenberg]. Physikalische Blätter, 32(3):97–104, March 1976. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19760320301/
abstract.

Durrenmatt:1980:PKZ

[Dür80] Friedrich Dürrenmatt. Die Physiker: Eine Komödie in zwei Akten. Neufassung 1980. (German) [The Physicists: A comedy in two acts. Revised 1980]. Diogenes, Zürich, Switzerland, 1980. ISBN 3-257-20837-5. 93 pp. LCCN ????

Durrani:2000:HSH

[Dur00] Matin Durrani. History of science: Heisenberg hoax puts author in a spin. Physics World, 13(7):7, July 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://

stacks.iop.org/2058-7058/13/i=7/a=7.

Durr:2005:ETW

[Dür05] Hans-Peter Dürr. Von Edward Teller zu Werner Heisenberg — Erinnerungen an meine Zusammenarbeit mit Heisenberg. (German) [From Edward Teller to Werner Heisenberg — memories of my collaboration with Heisenberg]. In Kleint et al. [KRW05], pages 321–331. ISBN 3-7776-1402-5. ISSN 0365-6470. LCCN

QC16.H35 W474 2005. URL http://www.loc.gov/catdir/toc/
fy0613/2006402966.html.

Durrenmatt:2006:PCT

[Dür06] Friedrich Dürrenmatt. The physicists: a comedy in two acts. Grove Press, New York, NY, USA, 2006. ISBN 0-8021-4427-6. 76 pp. LCCN PT2607.U493 P4913 2006. Translated by Joel Agee from the German original, Die Physiker (The Physicists), published by Verlag der Arche, Zürich 1962 and revised version by Diogenes Verlag AG, Zürich 1980.

David:1952:GRA

[DW52] L. R. David and L. A. Warheit. German reports on atomic energy: a bibliography of unclassified literature. Report TID-3030, Technical Information Service, Oak Ridge, TN, USA, 1952.

Dyson:1951:HOQa

[Dys51a] F. J. Dyson. Heisenberg operators in quantum electrodynamics. I. Physical Review (2), 82:428–439, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Dyson:1951:HOQb

[Dys51b] F. J. Dyson. Heisenberg operators in quantum electrodynamics. II. Physical Review (2), 83:608–627, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Einstein:1982:BGC

[EB82] Albert Einstein and Max Born. Briefwechsel: 1916–1955. (German) [Correspondence: 1916–1955]. Edition Erbrich, Frankfurt am Main, Germany, 1982. ISBN 3-88682-005-X. 330 pp. LCCN QC16.E5 A25 1982. With contributions by Hedwig Born, with an introduction by Bertrand Russell, and with a foreword by Werner Heisenberg. Reprint of the 1969 edition.

Echerbault:1996:MHF

[Ech96] Mireille Echerbault. Sur le modèle d’Heisenberg. (French) [on the Heisenberg model]. Lecture Notes in Mathematics, 1626:162– 177, 1996. CODEN LNMAA2. ISBN 3-540-61336-6 (print), 3-540-68463-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0094647/.

Eckart:1933:CNT

[Eck33] Carl Eckart. A comparison of the nuclear theories of Heisenberg and Wigner. I. Physical Review (2), 44(2):109, July 1933. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v44/i2/p109_1.

Eckart:1970:CEP

[Eck70] Carl Eckart. Carl Eckart papers, 1921–1973 (bulk 1935–1970). ????, ????, 1970. 8 (4002413) pp. LCCN Oversize 0113K; 0810D.

URL http://hdl.loc.gov/loc.mss/eadmss.ms001014.

Eckert:1990:PDF

[Eck90] Michael Eckert. Primacy doomed to failure: Heisenberg’s role as scientific adviser for nuclear policy in the FRG. Historical Studies in the Physical and Biological Sciences, 21(1):29–58, ???? 1990. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355

(electronic). URL http://www.jstor.org/stable/27757654.

Eckert:1992:HP

[Eck92] Michael Eckert. Heisenberg’s path. Science, 255(5047):1001– 1002, February 21, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/255/5047/1001.full.pdf.

Eckert:2001:WHC

[Eck01] Michael Eckert. Werner Heisenberg: controversial scientist. Physics World, 14(12):35–40, December 2001. CODEN PH-WOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL

http://stacks.iop.org/2058-7058/14/i=12/a=31.

Eckert:2012:RBH

[Eck12a] Michael Eckert. Rezension: Werner Heisenberg — Die Sprache der Atome. Leben und Wirken — Eine wissenschaftliche Biographie von Helmut Rechenberg. (German) [Review: Werner Heisenberg — The language of the atoms. Life and work — A Scientific Biography, by Helmut Rechenberg]. Berichte zur Wissenschaftsgeschichte, 35(1):73–74, March 2012. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Eckert:2012:RWH

[Eck12b] Michael Eckert. Rezension: Werner Heisenberg Die Sprache der Atome. Leben und Wirken: Eine wissenschaftliche Biographie von Helmut Rechenberg. Berichte zur Wissenschaftsgeschichte, 35(1): 73–74, 2012. CODEN BEWID8. ISSN 1522-2365.

Eddington:1928:NPW

[Edd28] Sir Arthur Stanley Eddington. The Nature of the Physical World, volume 1927 of Gifford lectures. Cambridge University Press, Cambridge, UK, 1928. xix + 361 + 1 pp. LCCN Q175 .E3. URL

https://archive.org/details/natureofphysical00eddi.

Eddington:1978:NPW

[Edd78] Arthur Stanley Eddington. The Nature of the Physical World, volume 15 of Ann Arbor Paperbacks. University of Michigan Press, Ann Arbor, MI, USA, 1978. ISBN 0-472-06015-5. xix + 361 pp. LCCN Q175 .E3 1978. Fifth printing.

Eddington:2000:NPW

[Edd00] Sir Arthur Stanley Eddington. The nature of the physical world. Kessinger Publishing’s rare reprints. Kessinger Publishing, LLC, Whitefish, MT, USA, 2000. ISBN 1-4179-0718-5. xvii + 361 pp. LCCN Q175 .E3 2000.

Eddington:2007:NPW

[Edd07] Sir Arthur Stanley Eddington. The Nature of the Physical World: Gifford Lectures (1927). Henry Foundation and The Johns Hopkins University Press, Baltimore, MD, USA, 2007. ISBN [none]. 361 pp. LCCN Q175 .E3 2007. URL http://henry.pha.jhu.

edu/eddington.html.

Eddington:2012:NPW

[Edd12] Sir Arthur Stanley Eddington. The Nature of the Physical World: Gifford Lectures (1927). Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-66385-7 (paperback). 361 pp. LCCN Q175 .E3 2012.

Eddington:2013:NPW

[Edd13] Sir Arthur Stanley Eddington. The Nature of the Physical World. Literary Licensing, LLC, ????, 2013. ISBN 1-4940-9894-6. URL

https://books.google.com/books?id=2zyUngEACAAJ.

Euler:1938:TGD

[EH38] Hans Euler and Werner Heisenberg. Theoretische Gesichtspunkte zur Deutung der kosmischen Strahlung. (German) [Theoretical aspects in the interpretation of cosmic radiation]. Ergebnisse der Exakten Naturwissenschaften, 17(2):1–69, ???? 1938. CODEN EENAA3. ISSN 0367-0325.

Euler:1984:TGD

[EH84] H. Euler and W. Heisenberg. Theoretische Gesichtspunkte zur Deutung der kosmischen Strahlung. (German) [Theoretical aspects of the interpretation of cosmic rays]. In Blum et al. [BDR84e], pages 262–330. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_23/.

Eckhardt:1996:BWS

[EHR^{+}96] B. Eckhardt, G. Hellbardt, H. Rechenberg, I. Heber, R. Haag, M. Carrier, S. Hüfner, M. Pillin, R. R. Ernst, H. Friedrich, C. Kleint, G. Baumann, and A. Ruider. Buchbesprechungen — Wissenschaftliche Software: Stehle: Order, Chaos, Order/Kafka: Gegen den Untergang/Verhulst: Der Glanz von Kopenhagen/ Luchner: Physik ist überall: Streifzüge durch Natur, Alltag, Technik und Forschung/Omnès: The Interpretation of Quantum Mechanics/Neuser: Natur und Begriff/Vogel: Gerthsen Physik/ Kassel: Quantum Groups/Schmidt-Rohr, Spiess: Multidimensional Solid-State NMR and Polymers/Ruder, Wunner, Herold, Geyer: Atoms in Strong Magnetic Fields/Herrmann: Der Photoeffekt/Kleint, Wiemers: Werner Heisenberg in Leipzig/Kofler: Maple V, Release 2: Einführung und Leitfaden für den Praktiker/ Mathcad Plus 6.0. Physikalische Blätter, 52(1):55–60, January 1996. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19960520116/abstract.

Evans:1992:HIR

[EHUY92] T. S. Evans, I. Hardman, H. Umezawa, and Y. Yamanaka. Heisenberg and interaction representations in thermo field dynamics. Journal of Mathematical Physics, 33(1):370–378, January 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i1/p370_s1.

Eickemeyer:1950:IPH

[Eic50] H. Eickemeyer. Interview mit Prof. Heisenberg/Die Arbeit des Deutschen Forschungsrates. (German) [Interview with Professor Heisenberg/the work of the German Research Council]. Physikalische Blätter, 6(4):170–174, April 1950. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19500060405/abstract.

Eickemeyer:1953:ADF

[Eic53] H. Eickemeyer, editor. Abschlußbericht des Deutschen Forschungsrates über seine Tätigkeit. (German) [Final report of the German Research Council about its activities]. Oldenbourg, Munich, West Germany, 1953. ???? pp. LCCN ???? 117–121.

Eisner:2000:RBP

[Eis00] Werner Eisner. Rezension: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Herausgegeben von Karl von Meyenn (Band IV, Teil I: 1950–1952; Teil II: 1953– 1954). Berlin/Heidelberg/New York: Springer 1996 und 1999; 968 und 1100 Seiten. (German) [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, and others. Edited by Karl of Meyenn (Volume IV, part I: 1950–1952, Part II: 1953–1954). Berlin / Heidelberg / New York: Springer 1996 and 1999, 968 and 1100 pages]. Berichte zur Wissenschaftsgeschichte, 23(1):59–62, ???? 2000. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

ElNaschie:2000:HCM

[El 00] M. S. El Naschie. Heisenberg’s critical mass calculations for an explosive nuclear reaction. Chaos, Solitons & Fractals, 11(6):987– 997, May 2000. CODEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Elion:1994:DDH

[EMGM94] W. J. Elion, M. Matters, U. Geigenmüller, and J. E. Mooij. Direct demonstration of Heisenberg’s Uncertainty Principle in a superconductor. Nature, 371(6498):594–595, October 13, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v371/
n6498/pdf/371594a0.pdf.

Engler:2009:KCB

[Eng09] Fynn Ole Engler. K. Camilleri: Heisenberg and the Interpretation of Quantum Mechanics. Physik Journal, 8(10):47, ???? 2009. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Enz:1983:HAQ

[Enz83] Charles P. Enz. Heisenberg’s applications of quantum mechanics (1926–33) or the settling of the new land. Helvetica Physica Acta, 56(5):993–1001, ???? 1983. CODEN HPACAK. ISSN 0018-0238. Extended version of the Introduction to Group 4 of the Collected Works of Werner Heisenberg.

Epstein:1971:PSF

[Eps71] Kenneth J. Epstein. Photon spin formalism. Journal of Mathematical Physics, 12(1):125–130, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v12/i1/p125_s1.

Ermenc:1989:ABS

[Erm89] Joseph J. Ermenc, editor. Atomic bomb scientists: memoirs, 1939–1945: interviews with Werner Karl Heisenberg, Paul Harteck, Lew Kowarski, Leslie R. Groves, Aristid von Grosse, C. E. Larson. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1989. ISBN 0-88736-267-2. 385 pp. LCCN QC774.A2 A86 1989. US$120.00.

Ellinas:1995:QHG

[ES95] Demosthenes Ellinas and Jan Sobczyk. Quantum Heisenberg group and algebra: contraction, left and right regular representations. Journal of Mathematical Physics, 36(3):1404–1412, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Essler:1971:RTG

[Ess71] Wilhelm K. Essler. Rezension: Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik von Werner Heisenberg. (German) [Book review: The Part and the Whole: Talks on the Field of Atomic Physics by Werner Heisenberg]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 2(2):333–336, September 1971. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://

www.jstor.org/stable/25170219.

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp. LCCN QC174.12 .Q346 2007. URL http://www.loc.gov/

catdir/enhancements/fy0824/2006934045-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0824/2006934045-d.
html;   http://www.loc.gov/catdir/toc/fy0708/2006934045.
html.

College:1982:MU

[Fac82] Faculty of Lynchburg College, editor. Man and the Universe, volume 1 (series 2) of Classical selections on great issues. University Press of America, Washington, DC, USA, 1982. ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). xvi + 370 pp. LCCN Q125 .M316 1982.

Falk:1964:IRN

[Fal64] Harold Falk. Inequalities relating the nearest-neighbor spin correlation and the magnetization for the Heisenberg Hamiltonian. Journal of Mathematical Physics, 5(10):1478–1480, October 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i10/p1478_s1.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137– 138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Favrholdt:2005:SNB

[Fav05] David Favrholdt. Spaltningen: Niels Bohr og Werner Heisenberg i videnskab og politik. (Danish) [Fission: Niels Bohr and Werner Heisenberg in science and politics]. Lindhardt og Ringhof, Copenhagen, Denmark, 2005. ISBN 87-595-2450-2. 282 pp. LCCN QC16.B63 F387 2005.

Faye:2007:BRG

[Fay07] Jan Faye. Book review: Gerald Holton: Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Isis, 98(3):660– 661, September 2007. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
10.1086/524275.

Frayn:2000:CSI

[FB00] Michael Frayn and David Burke. Celia’s secret: an investigation. Faber, London, UK, 2000. ISBN 0-571-20530-5. viii + 110 pp. LCCN PR6056.R3 C45 2000.

Frayn:2001:CPI

[FB01] Michael Frayn and David Burke. The Copenhagen papers: an intrigue. Metropolitan Books, New York, NY, USA, 2001. ISBN 0-8050-6752-3. 129 pp. LCCN PR6056.R3

C45 2001. URL http://www.loc.gov/catdir/enhancements/
fy0667/00051965-b.html;   http://www.loc.gov/catdir/enhancements/
fy0667/00051965-d.html.

Freire:2022:OHH

[FBD^{+}22] Olival Freire, Guido Bacciagaluppi, Olivier Darrigol, Thiago Hartz, Christian Joas, Alexei Kojevnikov, and Osvaldo Pessoa, editors. The Oxford Handbook of the History of Quantum Interpretations. Oxford handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2022. ISBN 0-19-884449-2 (hardcover). xiv + 1296 pp. LCCN

QC173.98 .O94 2022. URL https://global.oup.com/academic/
product/the-oxford-handbook-of-the-history-of-quantum-
interpretations-9780198844495.

Fechter:1990:AMH

[Fec90] Egidius Fechter. The Atomkeller-Museum in Haigerloch. Town of Haigerloch, Haigerloch, Germany, 1990. 20 pp.

Federbush:2004:PEQ

[Fed04] Paul Federbush. A polymer expansion for the quantum Heisenberg ferromagnet wave function. Journal of Mathematical Physics, 45 (1):44–50, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v45/i1/p44_s1.

Fedullo:2018:HUR

[Fed18] Aniello Fedullo. Heisenberg uncertainty relations as statistical invariants. Foundations of Physics, 48(11):1546–1556, November 2018. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Feld:1954:BRW

[Fel54] Bernard T. Feld. Book review: W. Heisenberg, Nuclear Physics. Physics Today, 7(8):18–??, August 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v7/i8/p18/s2.

Ferguson:1992:TW

[Fer92] R. Brian Ferguson. Tribal warfare. Scientific American, 266(1): 108–?? (Intl. ed. 90–??), January 1992. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Ferrari:1999:MDF

[Fer99] F. Ferrari. La matematizzazione della fisica: Galileo e Heisenberg epigoni di Platone? (Italian) [The mathematization of physics: Heisenberg as a follower of Plato and Galileo?]. Quaderni di Storia della fisica, 4:39–56, 1999. CODEN ???? ISSN 1594-9974 (print),

1827-6164 (electronic). URL http://en.sif.it/journals/qsf/
econtents.

Ferreira:2007:BRD

[Fer07] Pedro G. Ferreira. Book review: David Lindley, Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for The Soul of Science. Physics Today, 60(8):57–??, August 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v60/i8/p57/s1.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Fierz:1980:PDJ

[Fie80] M. Fierz. Physik in den dreissiger Jahren — ein Rückblick. (German) [Physics in the 1930s — a retrospective]. Physikalische Blätter, 36(6):133–136, June 1980. CODEN PHBLAG.

ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19800360603/
abstract.

Finkelstein:1996:QRS

[Fin96] David Ritz Finkelstein. Quantum relativity: a synthesis of the ideas of Einstein and Heisenberg. Texts and monographs in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-57084-5, 0-387-57084-5. xii + 577 pp. LCCN ????

Fischer:1941:BAU

[Fis41] E. Fischer. Bestimmung des Absorptionsquerschnittes von Uran für langsomme Neutronen. (German) [Determination of the absorption cross section of uranium for slow neutrons]. Report, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, June 9, 1941. ?? pp. See addendum [Hei41e].

Fisher:1964:MOD

[Fis64] Michael E. Fisher. Magnetism in one-dimensional systems — the Heisenberg model for infinite spin. American Journal of Physics, 32(5):343–346, May 1964. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/?AJP/32/343/1.

Fischer:2002:WHS

[Fis02] Ernst Peter Fischer. Werner Heisenberg: das selbstvergessene Genie; mit einem Nachtrag zur Taschenbuchausgabe; mit einer Tabelle. (German) [Werner Heisenberg: the forgotten genius; with a supplement to the paperback and a table], volume 3701 of Serie Piper. Piper, München, Germany, ungekürzte taschenbuchausgabe edition, 2002. ISBN 3-492-23701-0. 287 pp. LCCN ???? EUR 9.90.

Fischer:2010:HQE

[Fis10a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2010:WH

[Fis10b] Ernst Peter Fischer. Werner Heisenberg (1901–1976). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis10a], pages 171–187. ISBN 3-7766-2643-7. LCCN ????

Fischer:2012:HQE

[Fis12a] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fischer:2012:WH

[Fis12b] Ernst Peter Fischer. Werner Heisenberg (1901–1976). In Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger] [Fis12a], pages 171–187. ISBN 3-596-19406-7. LCCN ????

Fischer:2015:WHW

[Fis15] Ernst Peter Fischer. Werner Heisenberg: ein wanderer zwischen zwei welten. (German) [Werner Heisenberg: a wanderer between two worlds]. Springer Spektrum, Berlin, Germany, 2015. ISBN 3-662-43441-5, 3-662-43442-3 (e-book). xii + 374 + 43 pp. LCCN

QC16.H35. URL http://www.springerlink.com/content/978-
3-662-43442-0.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Flint:1932:UPM

[Fli32] H. T. Flint. The Uncertainty Principle in modern physics. Nature, 129(3264):746–747, May 21, 1932. CODEN NATUAS.

ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

www.nature.com/nature/journal/v129/n3264/pdf/129746a0.

pdf. See [Dar31].

Flint:1933:UP

[Fli33] H. T. Flint. The Uncertainty Principle. Nature, 132(3329):282, August 19, 1933. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v132/n3329/pdf/132282b0.pdf. See [Dar31, Fli32].

Flocke:1997:SPP

[Flo97] Norbert Flocke. SPINSGA: a program package for the evaluation of properties of spin-12 Heisenberg systems using the symmetric group approach. Computer Physics Communications, 106(1–2):114–124, October 2, 1997. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465597000738.

Furioli:2007:SIW

[FMV07] Giulia Furioli, Camillo Melzi, and Alessandro Veneruso. Strichartz inequalities for the wave equation with the full Laplacian on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 59(??):1301–1322, ???? 2007. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabbri:1994:RQH

[FO94] Marc A. Fabbri and Frank Okoh. Representations of quantum Heisenberg algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 46(??):920–929, ???? 1994. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fabbri:1999:RVH

[FO99] Marc A. Fabbri and Frank Okoh. Representations of Virasoro– Heisenberg algebras and Virasoro-toroidal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 51(3):523–545, June 1999. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Foot:1994:HM

[Foo94] Christopher Foot. A Heisenberg microscope. Nature, 371(6500): 744–745, October 27, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v371/n6500/pdf/371744a0.pdf.

Forman:2009:BRD

[For09] Paul Forman. Book review: David Lindley: Uncertainty: Einstein, Heisenberg, Bohr, and the Struggle for the Soul of Science. Isis, 100(1):180–181, March 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/599679.

Forstner:2022:LLI

[For22] Christian Forstner. Laboratory life instead of nuclear weapons: a new perspective on the German Uranium Club. Physics in Perspective (PIP), 24(4):181–207, December 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL https://

link.springer.com/article/10.1007/s00016-022-00294-8.

Flynn:1998:BRM

[FR98] Michael Flynn and Linda Rothstein. Bulletin: The real mystery science theater. Bulletin of the Atomic Scientists, 54(6):9–10, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fernandez-Ranada:2004:CIC

[FR04] Antonio Fernández-Rañada. Ciencia, incertidumbre y conciencia Heisenberg. (Spanish) [Heisenberg science, uncertainty, and conscience], volume 18 of Científicos para la historia. Nivola, Madrid, Spain, 2004. ISBN 84-95599-83-x. 276 pp. LCCN ????

Fernandez-Ranada:2008:HIC

[FR08] Antonio Fernández-Rañada. Heisenberg de la incertidumbre cuántica a la bomba atómica nazi. (Spanish) [Heisenberg of the quantum uncertainty to the Nazi atomic bomb], volume 1 of Científicos para la historia. Serie mayor. Nivola, Madrid, Spain, second edition, 2008. ISBN 84-96566-81-1. 334 pp. LCCN ????

Fernandez:2013:HST

[FR13a] Bernard Fernandez and Georges Ripka. Heisenberg and Schrödinger, two sides of the same coin. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13b], page 139. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13b] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Frank:1993:OET

[Fra93a] Sir Charles Frank, editor. Opération Epsilon: les transcriptions de Farm Hall. (French) [Operation Epsilon: the Farm Hall transcripts]. Flammarion, Paris, France, 1993. ISBN 2-08-211218-7. 382 pp. LCCN ????

Frank:1993:OEFa

[Fra93b] Sir Charles Frank, editor. Operation Epsilon: the Farm Hall transcripts. Institute of Physics Publishers, Bristol, UK, 1993. ISBN 0-7503-0274-7. ix + 313 pp. LCCN Q141 .O56 1993. US$15.00.

Frank:1993:OEFb

[Fra93c] Sir Charles Frank, OBE, FRS, editor. Operation Epsilon: the Farm Hall transcripts. University of California Press, Berkeley, CA, USA, 1993. ISBN 0-520-08499-3. ix + 313 pp. LCCN Q141 .O56 1993b.

Frank:1994:BRU

[Fra94] Sir Charles Frank, F.R.S. Book review: An uncertain life. David C. Cassidy, Uncertainty: The Life and Science of Werer Heisenberg. W. H. Freeman and Co., New York, 1991. $29.95, pp. xii + 669. ISBN 0-7167-2243-7. Notes and Records of the Royal Society of London, 48(1):147–151, January 1994. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Frayn:1998:C

[Fra98] Michael Frayn. Copenhagen. Methuen drama. Methuen Drama, London, UK, 1998. ISBN 0-413-72490-5. 116 pp. LCCN PR6056.R3 C64 1998.

Frayn:2000:C

[Fra00] Michael Frayn. Copenhagen. Anchor Books, New York, NY, USA, 2000. ISBN 0-385-72079-3 (paperback). 132 pp.

LCCN PR6056.R3 C64 2000. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;   http://www.loc.gov/catdir/

description/random044/00055814.html;   http://www.pbs.
org/hollywoodpresents/copenhagen/.

Frayn:2005:KSZ

[Fra05] Michael Frayn. Kopenhagen: Stück in zwei Akten. (German) [Copenhagen: play in two acts]. Wallstein, Göttingen, Germany, fourth edition, 2005. ISBN 3-89244-635-0. LCCN ???? German translation by Matthias Dörries of [Fra00].

Frappier:2010:BRK

[Fra10] Mélanie Frappier. Book review: Kristian Camilleri: Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher. Isis, 101(1):228–229, March 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/653885.

Frick:1997:NHO

[Fri97] Rudolf A. Frick. A new Hamiltonian operator for a massive relativistic particle with spin one in a generalized Heisenberg/ Schrödinger picture. Journal of Mathematical Physics, 38(7): 3457–3465, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [Fri98].

Frick:1998:ENH

[Fri98] Rudolf A. Frick. Erratum: “A new Hamiltonian operator for a massive relativistic particle with spin one in a generalized Heisenberg/Schrödinger picture” (J. Math. Phys. 38, 3457–3465 (1997)). Journal of Mathematical Physics, 39(4):2455, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Fri97].

Fritzsch:2011:YWM

[Fri11] Harald Fritzsch. You are wrong, Mr. Einstein!: Newton, Einstein, Heisenberg, and Feynman discussing quantum mechanics. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4324-99-X. xxi + 178 pp. LCCN QC174.12 .F755 2011.

Friederich:2013:IHI

[Fri13] Simon Friederich. Interpreting Heisenberg interpreting quantum states. Philosophia Naturalis, 50(1):85–114, ???? 2013. CODEN ???? ISSN 0031-8027.

Frank-Steiner:1993:LBF

[FS93] Vincent C. Frank-Steiner. Legendenbildung begann in Farm Hall: Zu D. Hoffmann: ,,Operation Epsilon”, Dezember-Heft 1992, S. 989. (German) [Creation of legends began in Farm Hall: From D. Hoffman, “Operation Epsilon”, December issue 1992, page 989]. Physikalische Blätter, 49(4):268–269, April 1993. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/

phbl.19930490404/abstract. See [Hof92]. Review of [Hof93], with response by H. Rechenberg.

Fischer:2006:EPF

[FS06] Ernst Peter Fischer and Klaus Sander. Ernst Peter Fischer erzählt, Paarläufe der Wissenschaft. (German) [Ernst Peter Fischer tells the pair runs of science]. Suppoš, Köln, Germany, 2006. ISBN 3-932513-68-1. LCCN ???? URL http://deposit.ddb.

de/cgi-bin/dokserv?id=2758637\%26prov=M\%26dok\_var=1\
%26dok\_ext=htm. 4 CDs.

Fleury:1994:PRH

[FT94] Norbert Fleury and Alexander Turbiner. Polynomial relations in the Heisenberg algebra. Journal of Mathematical Physics, 35 (11):6144–6149, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v35/i11/p6144_s1.

Folberth:2001:HCS

[FTF01] Otto G. Folberth, Michael Thorwart, and Egidius Fechter. Haigerloch cave survived the war. Physics Today, 54(5):93–94, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v54/i4/
p93/s1.

Fues:1961:WHG

[Fue61] E. Fues. Werner Heisenberg zum 60. Geburtstag: Die Erstgeburt der Quantenmechanik. (German) [Werner Heisenberg’s 60th birthday: the first-born of quantum mechanics]. Physikalische Blätter, 17(12):560–569, December 1961. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19610171203/
abstract.

Fulde:2015:SQW

[Ful15] Peter Fulde. Schöpfer der Quantenphysik: Werner Heisenberg und Erwin Schrödinger. (German) [Creator of quantum physics: Werner Heisenberg and Erwin Schrödinger]. In Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy] [LMOS15], pages 436–460. ISBN 3-11-037598-2 (hardcover). LCCN AS182.B38 B477 2015.

Feldman:2007:MSE

[FW07] Burton Feldman and Katherine Williams. 112 Mercer Street: Einstein, Russell, Gödel, Pauli, and the end of innocence in science. Arcade Publishing, New York, NY, USA, 2007. ISBN 1-55970-704-6. xx + 243 pp. LCCN Q141 .F345 2007. URL http://www.

loc.gov/catdir/toc/ecip078/2007001194.html.

Forstner:2020:BHP

[FW20] Christian Forstner and Mark Walker, editors. Biographies in the History of Physics: Actors, Objects, Institutions. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0. vi + 324 + 38 pp. LCCN QC16.E5 .B564 2020. URL http://www.

springerlink.com/content/978-3-030-48509-2.

Gembillo:2002:WHS

[GA02] Giuseppe Gembillo and Costanza Altavilla, editors. Werner Heisenberg: scienzato e filosofo. (Italian) [Werner Heisenberg: Scientist and Philosopher], volume 14 of Interazioni. A. Siciliano, Messina, Italy, 2002. 291 pp. LCCN QC16.H35 W473 2002.

Galles:2001:BCT

[Gal01] Carlos D. Galles. Born coined the term [quantum mechanics]. Physics Today, 54(5):94, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v54/i4/p94_s1.

Gamow:1930:MDC

[Gam30] George Gamow. Mass defect curve and nuclear constitution. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 126(803):632–644, March 3, 1930. ISSN 0950-1207 (print), 2053-9150 (electronic).

URL http://www.jstor.org/stable/pdfplus/95297.pdf.

Gamow:1958:PU

[Gam58] George Gamow. The Principle of Uncertainty. Scientific American, 198(1):51–57, January 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v198/n1/pdf/
scientificamerican0158-51.pdf.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty years that shook physics: the story of quantum theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12

.G35 1985. US$4.95. URL http://www.loc.gov/catdir/
description/dover032/85006797.html.

Garfield:1991:BRD

[Gar91] E. Garfield. Book reviews: David Cassidy’s biography of Werner Heisenberg — the quantum mechanic. part 1. the path to uncertainty [Uncertainty The Life and Science of Werner Heisenberg by David C. Cassidy]. Current Contents, 47(??):186–193, Novem-

ber 25, 1991. URL http://www.garfield.library.upenn.edu/
essays/v14p186y1991.pdf.

Gauderis:2014:ENP

[Gau14] Tjerk Gauderis. To envision a new particle or change an existing law? Hypothesis formation and anomaly resolution for the curious case of the β decay spectrum. Studies in History and Philosophy of Modern Physics, 45(??):27–45, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219813000920.

Garciaa-bach:1977:VLS

[GbK77] M. A. Garcíaa-bach and D. J. Klein. Variational localized-site cluster expansions. VII. Higher ansätze and the linear Heisenberg model. International Journal of Quantum Chemistry, 12(2):273– 289, August 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garcia-Bulle:1986:MGW

[GBLW86] Mauricio García-Bullé, Wolfgang Lassner, and Kurt Bernardo Wolf. The metaplectic group within the Heisenberg–Weyl ring. Journal of Mathematical Physics, 27(1):29–36, January 1986.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p29_s1.

Geizenberg:1938:KTK

[GE38] V. Geĭzenberg and G. Euler. K teorii kosmicheskovo islychenia. (Russian) [On the theory of cosmic radiation]. Uspekhi Fizicheskikh Nauk, 21(2):130–161, February 1938. CODEN EENAA3. ISSN 0367-0325. Russian translation of [EH38].

Geizenberg:1939:KTKa

[GE39a] V. Geĭzenberg and G. Euler. K teorii kosmicheskogo izluchenija. (Russian) [The theory of cosmic radiation]. Uspekhi Fizicheskikh Nauk, 21(2):130–161, February 1939. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ ru/articles/1939/3/c/. Russian translation of [EH38].

Geizenberg:1939:KTKb

[GE39b] V. Geĭzenberg and G. Euler. K teorii kosmicheskogo izluchenija. (Russian) [The theory of cosmic radiation]. Uspekhi Fizicheskikh Nauk, 21(3):261–300, March 1939. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1939/3/c/.

Geizenberg:1936:ZKT

[Geĭ36] V. Geĭzenberg. Zamechanija k teorii atomnogo jadra. (Russian) [Notes on the theory of the atomic nucleus]. Uspekhi Fizicheskikh Nauk, 16(1):1–7, January 1936. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1936/1/a/.

Geizenberg:1956:SST

[Geĭ56] V. Geĭzenberg. Sovremennoe sostojanie teorii jelementarnyh chastitsa. (Russian) [The present state of the theory of elementary particles]. Uspekhi Fizicheskikh Nauk, 60(11):413–424, November 1956. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1956/11/c/.

Geizenberg:1958:OPO

[Geĭ58] V. Geĭzenberg. Otkrytie planka i osnovnye filosofskie problemy atomnoj teorii. (Russian) [Opening bracket and the main philosophical problems of atomic theory]. Uspekhi Fizicheskikh Nauk, 66(10):163–175, October 1958. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1958/10/b/.

Geizenberg:1967:RFT

[Geĭ67] V. Geĭzenberg. Rol’ fenomenologicheskih teorij v sisteme teoreticheskoj fiziki. (Russian) [Role in the system of phenomenological theories of theoretical physics]. Uspekhi Fizicheskikh Nauk, 91(4):731–733, April 1967. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1967/4/g/.

Geizenberg:1968:NPV

[Geĭ68] V. Geĭzenberg. Nelinejnye problemy v fizike. (Russian) [Nonlinear problems in physics]. Uspekhi Fizicheskikh Nauk, 94(1):155–166, January 1968. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-

6652 (electronic). URL http://ufn.ru/ru/articles/1968/1/g/

.

Geizenberg:1970:PMB

[Geĭ70a] V. Geĭzenberg. Pamjati Maksa Borna. (Russian) [In memory of Max Born]. Uspekhi Fizicheskikh Nauk, 102(9):148–152, September 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1970/9/g/.

Geizenberg:1970:TKF

[Geĭ70b] V. Geĭzenberg. Teorija, kritika i filosofija. (Russian) [Theory, criticism, and philosophy]. Uspekhi Fizicheskikh Nauk, 102 (10):298–312, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1970/10/k/.

Geizenberg:1977:KIK

[Geĭ77a] V. Geĭzenberg. O kvantovoteoreticheskom istolkovanii kinematicheskih i mehanicheskih sootnoshenij. (Russian) [On the interpretation of quantum-mechanical and kinematic relations]. Uspekhi Fizicheskikh Nauk, 122(8):574–586, ???? 1977. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1977/8/c/.

Geizenberg:1977:NSK

[Geĭ77b] V. Geĭzenberg. O nagljadnom soderzhanii kvantovoteoreticheskoj kinematiki i mehaniki. (Russian) [Quantum-theoretical interpretation of kinetic and mechanical relations]. Uspekhi Fizicheskikh Nauk, 122(8):651–671, August 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ ru/articles/1977/8/h/. Russian translation of [Hei27e].

Geizenberg:1977:PJC

[Geĭ77c] V. Geĭzenberg. Priroda jelementarnyh chastits. (Russian) [The nature of elementary particles]. Uspekhi Fizicheskikh Nauk, 121(4):657–677, April 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/ articles/1977/4/e/. Russian translation of [Hei76a].

Geizenberg:1987:RSB

[Geĭ87] V. Geĭzenberg. ???? (Russian) [Steps beyond the horizon]. Mir, Moscow, USSR, 1987. ???? pp.

Geizenberg:1990:FFC

[Geĭ90] V. Geĭzenberg. Fizika i filosofiya. Chast i tseloe. (Russian) [Physics and Philosophy. Part and whole]. Nauka, Moscow, Russia, 1990. ISBN 5-02-012452-4. 400 pp. Translated from the German by I. A. Akchurin, È. P. Andreev and V. V. Bibikhin, With a chapter by A. V. Akhutin.

Geller:1984:SHW

[Gel84] Daryl Geller. Spherical harmonics, the Weyl transform and the Fourier transform on the Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(??):615– 684, ???? 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gembillo:1987:WHF

[Gem87] Giuseppe Gembillo. Werner Heisenberg : la filosofia di un fisico, volume 8 of Filosofia e storiografia. Giannini, Napoli, Italia, 1987. ISBN ???? ISSN 1123-4938. xiv + 195 pp. LCCN QC16.H395.

Galison:2002:QMS

[GGK02] Peter Galison, Michael Gordin, and David Kaiser, editors. Quantum Mechanics: Science and Society. Routledge, London, UK, 2002. ISBN 1-136-70972-X. 433 (est.) pp. LCCN ????

Geloun:2010:GHA

[GH10] Joseph Ben Geloun and Mahouton Norbert Hounkonnou. q-graded Heisenberg algebras and deformed supersymmetries. Journal of Mathematical Physics, 51(2):023502, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v51/i2/
p023502_s1.

Giacomuzzi:2002:EMB

[GHH02] Salvatore Matteo Giacomuzzi, Gerhard Holzmüller, and Gerhard Huemer. Ettore Majorana (1906–1938): Eine Bestandsaufnahme 64 Jahre nach seinem Verschwinden. (German) [Ettore Majorana (1906–1938): A survey 64 years after his disappearance]. Berichte zur Wissenschaftsgeschichte, 25(2):137–148, June 2002. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Geiger:1936:MRE

[GHW36] Hans Geiger, Werner Heisenberg, and Max Wein. [Memorandum to Der Reichsminister für Erziehung und Wissenschaft [The Reich Minister for Education and Science]], ???? 1936. Letter opposing the Nazi discrediting of theoretical physics, cited in [BDR84e, page 7].

Greenberger:2009:CQP

[GHW09] Daniel M. Greenberger, Klaus Hentschel, and Friedel Weinert, editors. Compendium of quantum physics: concepts, experiments, history and philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-70626-7, 3-540-70622-4. xvi + 901 pp. LCCN C174.12 .C66 2009.

Gidas:1971:QFI

[Gid71] Basilis Gidas. Quadratic fermion interaction Hamiltonian. Journal of Mathematical Physics, 12(7):1414–1419, July 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/
p1414_s1.

Gilder:2009:AEW

[Gil09] Louisa Gilder. The Age of Entanglement: When Quantum Physics Was Reborn. Vintage Books, New York, NY, USA, 2009. ISBN 1-4000-9526-3. xviii + 443 pp. LCCN QC174.12 .G528 2009.

Gimbel:1990:STR

[Gim90] John Gimbel. Science, technology, and reparations: exploitation and plunder in postwar Germany. Stanford University Press, Stanford, CA, USA, 1990. ISBN 0-8047-1761-3. xv + 280 pp.

LCCN Q149.U5 G53 1990. US$35.00. URL http://www.loc.

gov/catdir/description/cam024/89037835.html;   http://
www.loc.gov/catdir/toc/cam027/89037835.html.

Gingerich:1975:NSD

[Gin75] Owen Gingerich, editor. The Nature of Scientific Discovery: a Symposium Commemorating the 500th Anniversary of the Birth of Nicolaus Copernicus, volume 5 of Smithsonian international symposia series. Smithsonian Institution Press, Washington, DC, USA, 1975. ISBN 0-87474-148-3. LCCN QB36.C8 N37.

Greenberg:1996:HMM

[GKZL96] William R. Greenberg, Abraham Klein, Ivaylo Zlatev, and Ching-Teh Li. From Heisenberg matrix mechanics to semiclassical quantization: Theory and first applications. Physical Review A (Atomic, Molecular, and Optical Physics), 54(3):1820–1837, September 1996. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://pra.aps.org/

abstract/PRA/v54/i3/p1820_1.

Golub:2023:HPF

[GL23] Robert Golub and Steve Keith Lamoreaux. The Historical and Physical Foundations of Quantum Mechanics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2023. ISBN 0-19-186123-5, 0-19-255536-7, 0-19-882218-9 (hardcover), 0-19-882219-7 (paperback). xiii + 747 pp. LCCN QC174.12 .G65 2023.

Glanz:2000:PFA

[Gla00] James Glanz. Of physics, friendship and the atomic bomb: 60 years later, a mystery abides of physics, friendship and Nazi Germany’s atomic bomb efforts mystery of a 1941 meeting becomes drama on the stage. New York Times, ??(??):F1–F2, March 21, 2000. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/91737705/.

Guerra:2014:WEC

[GLR14] Francesco Guerra, Matteo Leone, and Nadia Robotti. When energy conservation seems to fail: The prediction of the neutrino. Science & Education (Springer), 23(6):1339–1359, June 2014. CO-DEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Gottstein:2001:HMB

[GLSC01] Klaus Gottstein, Harry J. Lipkin, Donald C. Sachs, and David C. Cassidy. Heisenberg’s message to Bohr: Who knows? Physics Today, 54(4):14, 92–93, April 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v54/i4/p14/s1.

Goldberg:1992:ICO

[Gol92] Stanley Goldberg. Inventing a climate of opinion: Vannevar Bush and the decision to build the bomb. Isis, 83(3):429–452, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233904.

Goldberg:1995:GSC

[Gol95] Stanley Goldberg. Groves and the scientists: Compartmentalization and the building of the bomb. Physics Today, 48(8):38– 43, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v48/i8/p38_s1.

Grasso:2009:NSH

[GORS09] Giacomo Grasso, Carlo Oppici, Federico Rocchi, and Marco Sumini. A neutronics study of the 1945 Haigerloch B-VIII nuclear reactor. Physics in Perspective (PIP), 11(3):318–335, September 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/

10.1007/s00016-008-0396-0. See also [KH19, CR13, MWL^{+}11].

Gottstein:1983:CGE

[Got83] Klaus Gottstein. Commentary: The Göttingen Eighteen. Bulletin of the Atomic Scientists, 39(8):62–63, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottstein:2002:HBA

[Got02a] Klaus Gottstein. Heisenberg and Bohr — another view. Science, 295(5563):2211, March 22, 2002. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/295/5563/2211.2.full. Rebuttal to

[Cho02].

Gottstein:2002:WSH

[Got02b] Klaus Gottstein. Woher sollte Heisenberg das gewusst haben? (German) [How should Heisenberg know that?]. Physik Journal, 1 (6):29–??, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print),

1619-6597 (electronic).  URL http://www.pro-physik.de/
details/physikjournalIssue/1089827/Issue_6_2002.html.

Gottstein:2004:CHP

[Got04] Klaus Gottstein. Comment on “Heisenberg in Poland” by Jeremy Bernstein [Am. J. Phys. 72 (3), 300–304 (2004)]. American Journal of Physics, 72(9):1143–1144, September 2004. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/72/1143/1. See [Ber04].

Goudsmit:1946:HGL

[Gou46a] Samuel A. Goudsmit. How Germany lost the race. Bulletin of the Atomic Scientists, 1(7):4–5, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1946:WPG

[Gou46b] Samuel A. Goudsmit. War physics in Germany. Review of Scientific Instruments, 17(1):49–52, ???? 1946. CODEN RSINAK. ISSN 0034-6748 (print), 1089-7623 (electronic).

Goudsmit:1947:GAR

[Gou47a] Samuel A. Goudsmit. German atom research: Scientist disputes statement on development of bomb: the writer of the following letter, Professor of Physics at Northwestern University, headed the American mission investigating Germany’s progress in atomic research. New York Times, ??(??):E8, November 9, 1947. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/108045246/.

Goudsmit:1947:GSA

[Gou47b] Samuel A. Goudsmit. German scientists in army employment: I. The case analyzed. Bulletin of the Atomic Scientists, 3(2):64, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1947:HGN

[Gou47c] Samuel A. Goudsmit. Heisenberg on the German Nuclear Power Project. Bulletin of the Atomic Scientists, 3(11):343, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Hei47g].

Goudsmit:1947:PA

[Gou47d] Samuel A. Goudsmit. [Preview of Alsos]. Life Magazine, page ??, October 20, 1947.

Goudsmit:1947:A

[Gou47e] Samuel Abraham Goudsmit. Alsos. H. Schuman, New York, NY, USA, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947.

Goudsmit:1947:AFG

[Gou47f] Samuel Abraham Goudsmit. Alsos: the failure in German science. Sigma Books, London, UK, 1947. xiv + 259 pp. LCCN D810.S2 G6 1947a.

Goudsmit:1948:OTG

[Gou48a] Samuel A. Goudsmit. Our task in Germany. Bulletin of the Atomic Scientists, 4(4):106, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1948:LSA

[Gou48b] Samuel Abraham Goudsmit. L’Allemagne et le secret atomique (la mission Alsos). (French) [Germany and the atomic secret (Alsos mission)]. A. Fayard, Paris, France, 1948. 250 + 1 pp. LCCN D810.S2 G615.

Goudsmit:1949:GWR

[Gou49] Samuel A. Goudsmit. German war research: Scientist’s claim of progress with atom is disputed. New York Times, ??(??): E8, January 9, 1949. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105811389/. The writer of the

following letter, co-discovered of the “spin of the Electron”, was detailed to the War Department as chief of the ALSOS mission investigating Germany’s progress in Atomic research. He is now a senior physicist at the Brookhaven National Laboratory.

Goudsmit:1976:WH

[Gou76] Samuel A. Goudsmit. Werner Heisenberg (1901–1976). Yearbook of the American Philosophical Society, pages 74–80,

1976. URL http://www.aip.org/history/nbl/collections/
goudsmit/colls/box46/box46f148.html.

Goudsmit:1983:A

[Gou83] Samuel Abraham Goudsmit. Alsos. History of modern physics, 1800–1950. Tomash Publishers, Los Angeles, CA, USA, 1983. ISBN 0-938228-09-9. xxx + 259 pp. LCCN D810.S2 G6 1983.

Goudaroulis:1990:MOC

[Gou90] Yorgos Goudaroulis. A matter of order: A controversy between Heisenberg and London. In Greek Studies in the Philosophy and History of Science [Nic90], chapter 18, pages 279–291. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346.

LCCN ???? URL https://link.springer.com/chapter/10.
1007/978-94-009-2015-6_18.

Goudsmit:1996:A

[Gou96] Samuel Abraham Goudsmit. Alsos. History of modern physics and astronomy. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-415-5. xxxviii + 259 pp. LCCN D810.S2 G6 1996.

Goldberg:1992:ATG

[GP92a] Stanley Goldberg and Thomas Powers. Alsos: Tracking the German bomb. Bulletin of the Atomic Scientists, 48(7):35, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:DFR

[GP92b] Stanley Goldberg and Thomas Powers. Declassified files reopen “Nazi bomb” debate. Bulletin of the Atomic Scientists, 48(7):32– 35, 38–40, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DR92].

Goldberg:1992:GFH

[GP92c] Stanley Goldberg and Thomas Powers. The “guests” at Farm Hall. Bulletin of the Atomic Scientists, 48(7):36–37, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guerra:2008:EMF

[GR08] Francesco Guerra and Nadia Robotti. Ettore Majorana’s forgotten publication on the Thomas–Fermi model. Physics in Perspective (PIP), 10(1):56–76, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0340-8;
http://www.springerlink.com/content/53l26g8353l2421g/.

Guerra:2009:EFD

[GR09] Francesco Guerra and Nadia Robotti. Enrico Fermi’s discovery of neutron-induced artificial radioactivity: The influence of his theory of beta decay. Physics in Perspective (PIP), 11(4):379– 404, December 2009. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://www.springerlink.com/
content/65w213132524072n/.

Gratzer:2000:USD

[Gra00] W. B. (Walter Bruno) Gratzer. The undergrowth of science: delusion, self-deception, and human frailty. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-850707-0. ix + 328 pp. LCCN Q125 .G785 2000. URL http://www.

loc.gov/catdir/enhancements/fy0639/00036309-d.html;
http://www.loc.gov/catdir/enhancements/fy0639/00036309-
t.html;   http://www.loc.gov/catdir/enhancements/fy0725/
00036309-b.html.

Greenberg:1962:HFW

[Gre62] O. W. Greenberg. Heisenberg fields which vanish on domains of momentum space. Journal of Mathematical Physics, 3(5):859–866, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v3/i5/p859_s1.

Greenberg:1963:BKF

[Gre63] Newton I. Greenberg. Binary kernel formulation of a Heisenberg model of ferromagnetism. Journal of Mathematical Physics, 4(3): 405–409, March 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v4/i3/p405_s1.

Greiner:1980:SHH

[Gre80] Peter C. Greiner. Spherical harmonics on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 23(??):383–396, ???? 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Gornitz:2001:BSM

[GRE^{+}01] Thomas Görnitz, Helmut Rechenberg, Berthold-Georg Englert, Thomas Walcher, Katja Bammel, Thomas W. Beneke, Wolfgang W. Schwippert, and Filip Floecel. Bücher und Software:

Meyenn: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg U. A./Kraus: Von der Uranspaltung zur Göttinger Erklärung. Otto Hahn, Werner Heisenberg und Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers/Dürr: Bohmsche Mechanik als Grundlage der Quantenmechanik/Thomas/Weise: The Structure of the Nucleon/ Sube: Langenscheidts Fachwörterbuch Physik. Deutsch-Englisch/ ConceptDraw 1.6 Zeichnen mit intelligenten Objekten /Rath: Quantenphysik für Windows Version 1.5. Physikalische Blätter, 57 (11):82–86, November 2001. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.20010571123/abstract.

Greenspan:2005:ECW

[Gre05] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Gregory:2010:BRR

[Gre10] Andrew Gregory. Book review: Robert P. Crease: The Great Equations: Breakthroughs in Science from Pythagoras to Heisenberg. Isis, 101(3):626–627, September 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/657178.

Groetsch:1996:TIP

[Gro96] C. W. Groetsch. Tartaglia’s inverse problem in a resistive medium. American Mathematical Monthly, 103(7):546–551, August/September 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/2974665.

Gulde:2012:SSU

[GS12] Michael Gulde and Markus Stroppel. Stabilizers of subspaces under similitudes of the Klein quadric, and automorphisms of Heisenberg algebras. Linear Algebra and its Applications, 437(4):1132–1161, August 15, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0024379512002492.

Gaveau:1985:DPS

[GV85] Bernard Gaveau and Jacques Vauthier. The Dirichlet problem for the subelliptic Laplacian on the Heisenberg group. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 37(??):760–766, ???? 1985. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gaveau:1986:DPH

[GV86] Bernard Gaveau and Jacques Vauthier. The Dirichlet problem on the Heisenberg group. III. harmonic measure of a certain halfspace. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(??):666–671, ???? 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldstine:1996:SWJ

[GW96] H. H. Goldstine and E. P. Wigner. The scientific work of John von Neumann. In Wigner [Wig96m], pages 123–126. ISBN 3-540-56972-3. Reprinted from [?].

Guo:2009:IHO

[GWW^{+}09] Jia-Feng Guo, Shi-Kun Wang, Ke Wu, Zhao-Wen Yan, and Wei-Zhong Zhao. Integrable higher order deformations of Heisenberg supermagnetic model. Journal of Mathematical Physics, 50(11): 113502, November 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v50/i11/p113502_s1.

Grunden:2005:WNWa

[GWY05a] Walter E. Grunden, Mark Walker, and Masakatsu Yamazaki. Wartime nuclear weapons research in Germany and Japan. Osiris (Series 2), 20(??):107–130, ???? 2005. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic). URL http://www.

jstor.org/stable/3655253.

Grunden:2005:WNWb

[GWY05b] Walter E. Grunden, Mark Walker, and Masakatsu Yamazaki. Wartime nuclear weapons research in Germany and Japan. In Sachse and Walker [SW05], page ?? ISBN 0-226-73329-7, 0-226-73328-9. ISSN 0369-7827 (print), 1933-8287 (electronic). LCCN Q49.K14174 P65 2005; AS182 .K357 2005.

Geyer:1992:WHP

[GZ92a] B. Geyer and C. Zylka. Werner Heisenberg als Physiker und Philosoph in Leipzig. (German) [Werner Heisenberg as a physicist and philosopher in Leipzig]. Physikalische Blätter, 48(2):118–119, February 1992. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19920480219/abstract.

Geyer:1992:DIW

[GZ92b] Bodo Geyer and Christian Zylka. Dokumentation und Information: Werner Heisenberg als Physiker und Philosoph in Leipzig. (German) [Werner Heisenberg: Physicist and philosopher in Leipzig]. Berichte zur Wissenschaftsgeschichte, 15(1):56–58, ???? 1992. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Greenberger:1995:FPQ

[GZ95] Daniel M. Greenberger and Anton Zeilinger, editors. Fundamental problems in quantum theory: a conference held in honor of Professor John A. Wheeler, volume 755 of Annals of the New York Academy of Sciences. New York Academy of Sciences, New York, NY, USA, 1995. CODEN ANYAA9. ISBN 0-89766-921-5, 0-89766-922-3 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). LCCN Q11 .N5 vol. 755; QC173.96.

Heisenberg:1948:ADW

[HA^{+}48] Werner Heisenberg, H. J. Abs, et al. Arbeit durch wissenschaft. aufruf! Die Welt, 143(??):3, ???? 1948.

Haas:1934:MQE

[Haa34] Arthur Haas. Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. 4. u. 5. verb. u. abermals wes. verm. Aufl. (German) [Matter waves and quantum mechanics. An introduction based on the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Akad. Verlagsges., Leipzig, Germany, 1934. viii + 299 + 7 pp.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Habgood:1983:UPT

[Hab83] John Habgood. Uncertainty principle of theology: Book review: Science and the Renewal of Belief, by Russell Stannard. Nature, 302(5906):362, March 24, 1983. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v302/n5906/pdf/302362a0.pdf.

Hahn:1940:BAK

[Hah40] Otto Hahn. Bericht über die Arbeiten des Kaiser Wilhelm-Instituts für Chemie über “Präparat 38”. (German) [Report on the work of the Kaiser Wilhelm Institute for Chemistry on “Preparation 38”]. Report FA 002 514, Archives of the Deutsches Museum (ADM), Munich, Germany, December 10, 1940.

Hahn:1943:KAS

[Hah43] Otto Hahn. Künstliche atomumwandlungen und die spaltung des urans. (German) [Artificial atom transformations and the splitting of uranium]. Report FA 002 737, Archives of the Deutsches Museum (ADM), Munich, Germany, May 6, 1943.

Hahn:1945:FHT

[Hah45] Otto Hahn. Farm Hall Tagebücher. (German) [Farm Hall diaries]. Unpublished typescript., 1945. Archive of the Max-Planck-Gesellschaft, Berlin-Dahlem, Germany.

Hahn:2006:CKA

[Hah06] Ralf Hahn. C. Kleint et al. (Hrsg.): Werner Heisenberg 1901– 1976. Physik Journal, 5(12):60, ???? 2006. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Hahn:2010:CDW

[Hah10] Matthias Hahn. CD/DVD: Werner Heisenberg — Die Schönheit der Weltformel. (German) [CD / DVD: Werner Heisenberg — the beauty of the world formula]. Physik Journal, 9(7):62, ???? 2010. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
physikjournalIssue/1089649/Issue_7_2010.html.

Hahn:2011:CCB

[Hah11] Matthias Hahn. C. Carson: Heisenberg in the Atomic Age. Physik Journal, 10(2):46, ???? 2011. CODEN PJHOB2.

ISSN 1617-9439 (print), 1619-6597 (electronic). URL http:/

/www.pro-physik.de/details/physikjournalIssue/1089637/
Issue_2_2011.html.

Haken:2002:HEL

[Hak02] Hermann Haken. Heisenberg’s equations in laser theory. A historical overview. Fortschritte der Physik = Progress of Physics, 50 (5–7):642–645, 2002. CODEN FPYKA6. ISSN 1521-3978.

Hall:2009:GSP

[Hal09] Derek Hall, editor. Great scientists: Physical sciences. Facts at your fingertips. Brown Bear Books, Redding, CT, USA, 2009. ISBN 1-933834-46-3. 64 pp. LCCN QC15 .P468 2009.

Hanson:1962:RBP

[Han62] Norwood Russell Hanson. Review: The Physicist’s Conception of Nature by Werner Heisenberg (translated by Arnold J. Pomerans), Physics and Philosophy: The Revolution in Modern Science by Werner Heisenberg; Mind and Matter by Erwin Schrödinger. The Journal of Philosophy, 59(12):320–322, June 7, 1962. CODEN ???? ISSN 0022-362X (print), 1939-8549 (electronic). URL http:

//www.jstor.org/stable/2022902.

Hanle:1979:IBH

[Han79] Paul A. Hanle. Indeterminacy before Heisenberg: The case of Franz Exner and Erwin Schrödinger. Historical Studies in the Physical Sciences, 10(??):225–269, ???? 1979. CODEN HSP-

SAS. ISSN 0073-2672. URL http://www.jstor.org/stable/
27757391.

Hartshorne:1937:GUNa

[Har37a] Edward Yarnall Hartshorne, Jr. The German universities and National Socialism. Harvard University Press, Cambridge, MA, USA, 1937. 179 + 27 + 181–184 pp. LCCN LA728 .H36 1937a.

Hartshorne:1937:GUNb

[Har37b] Edward Yarnall Hartshorne, Jr. The German universities and National Socialism. G. Allen and Unwin, London, UK, 1937. 184 pp. LCCN LA728 .H36.

Harteck:1943:BBB

[Har43] Paul Harteck. Bericht über die Besprechungen in Berlin am 27.9.43 und in Leuna am 28.9.43. (German) [Report on the meetings in Berlin on 27 September 1943 and in Leuna on 28 September 1943]. Report FA 002 796, Archives of the Deutsches Museum (ADM), Munich, Germany, October 1943.

Harteck:1960:PCB

[Har60] Paul Harteck. Physical chemists in Berlin: 1919–1933. Journal of Chemical Education, 37(9):462, 1960. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL http://

pubs.acs.org/doi/abs/10.1021/ed037p462.

Harney:1973:MOF

[Har73] R. C. Harney. A method for obtaining force law information by applying the Heisenberg uncertainty principle. American Journal of Physics, 41(1):67–70, January 1973. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/41/67/1.

Hartshorne:1981:GUN

[Har81] Edward Yarnall Hartshorne, Jr. The German universities and National Socialism. AMS Press, New York, NY, USA, 1981. ISBN 0-404-16943-0. 184 pp. LCCN LA728 .H36 1981.

Haupt:2022:DWB

[Hau22] Heinz Dieter Haupt. Deutschlands Weg zur Bombe: Chimäre oder Realität?: vom Dritten Reich bis zur Bundesrepublik: die Geschichte alternativer Kernwaffenentwicklungen in Deutschland im Kontext der internationalen Forschung. (German) [Germany’s path to the bomb: chimera or reality?: from the Third Reich to the Federal Republic: the history of alternative nuclear weapon developments in Germany in the context of international research]. Literareon. Literareon, im Utzverlag, München, Germany, 2022. ISBN 3-8316-2328-7 (hardbound). 810 pp. LCCN U264.5.G3 H385 2022.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Hayes:1990:BRB

[Hay90] Peter Hayes. Book review: German National Socialism and the Quest for Nuclear Power, 1939-1949. Mark Walker. Cambridge University Press, New York, 1989. x, 290 pp., illus. $29.95. Science, 249(4968):570–??, August 3, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/249/4968/570.citation.

Heisenberg:1963:KDQ

[HB63] Werner Heisenberg and Niels Bohr, editors. Die Kopenhagener Deutung der Quantentheorie. (German) [The Copenhagen Interpretation of Quantum Theory]. E. Battenberg, Stuttgart, West Germany, 1963. 67 pp. LCCN QC174.1 .K65.

Heisenberg:1972:ADA

[HB^{+}72] Werner Heisenberg, A. Butenandt, et al. An alle Deutschen [Anzeige]. (German) [To all Germans [advertisement]]. Frankfurter Allgemeine Zeitung, 105(??):14, May 6, 1972.

Hanle:1982:JNK

[HB82a] Wilhelm Hanle and Erich Bagge. 40 Jahre Nutzung der Kernenergie. (German) [Forty years of use of nuclear energy]. Atomkernenergie, 40(??):3–7, ???? 1982. CODEN ATKEDU. ISSN 0171-5747.

Hanle:1982:PUN

[HB82b] Wilhelm Hanle and Erich Bagge. Peaceful use of nuclear energy during 40 years. Atomkernenergie, 40(??):1–3, ???? 1982. CODEN ATKEDU. ISSN 0171-5747.

Heisenberg:1989:BBVb

[HBB^{+}89] Werner Heisenberg, F. Bopp, W. Bothe, E. Fischer, O. Ritter, and K. Wirtz. Bericht über einen Versuch mit 1.5 to D_{2}O und U und 40 cm Kohlerückstreumantel (B7) (3 Jan. 1945). (German) [Report on an experiment with 1.5 tonnes D_{2}O and U, and 40 cm graphite backscatter cloak (B7 (3 Jan. 1945))]. In Blum et al. [BDR89a], pages 595–601. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Higgins:2007:EFH

[HBB^{+}07] B. L. Higgins, D. W. Berry, S. D. Bartlett, H. M. Wiseman, and G. J. Pryde. Entanglement-free Heisenberg-limited phase estimation. Nature, 450(7168):393–396, November 15, 2007.

CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v450/
n7168/full/nature06257.html.

Heisenberg:1971:QFP

[HBD71] Werner Heisenberg, Erich Bagge, and H. Peter Dürr, editors. Quanten und Felder; physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg. (German) [Quanta and fields: physical and philosophical views. On the 70th Birthday of Werner Heisenberg]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1971. 366 pp. LCCN QC174.45 .Q35.

Hoddeson:1987:DQM

[HBE87] Lillian Hoddeson, Gordon Baym, and Michael Eckert. The development of the quantum-mechanical electron theory of metals: 1928–1933. Reviews of Modern Physics, 59(1):287–327, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.287;   http://rmp.aps.org/abstract/
RMP/v59/i1/p287_1.

Heisenberg:1942:MSM

[HBF^{+}42] Werner Heisenberg, Fritz Bopp, Erich Fischer, Carl-Friedrich von Weizsäcker, and Karl Wirtz. Messungen an Schichtenanordnungen aus 38-Metall und Paraffin. (German) [Measurements on arrays of 38-layer metal and paraffin]. Kernphysikalische Forschungsberichte G-162, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, October 30, 1942. 17 pp.

Heisenberg:1989:MSM

[HBF^{+}89a] Werner Heisenberg, F. Bopp, E. Fischer, C. F. von Weizsäcker, and K. Wirtz. Messungen an Schichtenanordnungen aus 38-Metall und Paraffin (30 Okt. 1942). (German) [Measurements on arrays of 38-layer metal and paraffin (30 October 1942)]. In Blum et al. [BDR89a], pages 553–569. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:UNS

[HBF^{+}89b] Werner Heisenberg, F. Bopp, E. Fischer, C. F. von Weizsäcker, and K. Wirtz. Untersuchungen mit neuen Schichtenanordnungen aus U-metall und Paraffin (26.-28. Feb. 1942). (German) [Studies with a new layer arrangement of uranium metal and paraffin (26– 28 February 1942)]. In Blum et al. [BDR89a], pages 522–525. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:VBB

[HBF^{+}89c] Werner Heisenberg, F. Bopp, E. Fischer, C. F. von Weizsäcker, and K. Wirtz. Vorlaufiger Bericht über Ergebnisse an einer Schichtenkugel aus 38-Metall und Paraffin [Bill] (6 Jan. 1942). (German) [Preliminary report on results of layers of 38-ball metal and paraffin [bill] (6 January 1942)]. In Blum et al. [BDR89a], pages 514–516. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1996:FCS

[HBS96] Werner Heisenberg, Niels Bohr, and Erwin Schrödinger. Física cuántica: (Spanish) [Quantum physics]. Círculo de Lectores, Barcelona, Spain, 1996. ISBN ???? ???? pp. LCCN ????

Heisenberg:1961:MPb

[HBSA61] Werner Heisenberg, Max Born, Erwin Schrödinger, and Pierre Auger. On Modern Physics. Clarkson N. Potter. Inc., New York, NY, USA, 1961. 108 pp. LCCN QC71 .O5. URL http://alsos.

wlu.edu/information.aspx?id=997. English translation by M.

Goodman and J. W. Binns.

Heisenberg:1961:TSP

[HD61] Werner Heisenberg and H.-P. Dürr. Theory of the strange particles. In E. Cremieu-Alcan et al., editors, Conference Internationale d ?Aix-en-Provence sur les Particules Elementaires The Aix-en-Provence — International Conference on Elementary Particles. Vol. 1, pages 163–168. C.E.N. Saclay, Gif-sur-Yvette, France, 1961. LCCN ????

Heisenberg:1975:QET

[HDW75] W. Heisenberg, Paul A. M. Dirac, and V. Weisskopf. Quantum electrodynamics: Translations of four papers by Heisenberg, Dirac, and Weisskopf. Technical Report RL-75-073, Science and Technology Facilities Council ePublication Archive, Chilbolton, Daresbury, and Rutherford Appleton Laboratories,

UK, 1975. URL http://epubs.stfc.ac.uk/work-details?
w=21197;   http://purl.org/oai/oai:epubs.cclrc.ac.uk:
work/21197.

Heisenberg:1935:FDT

[HE35] Werner Heisenberg and Hans Euler. Folgerungen aus der Diracschen Theorie des Positrons. (German) [Consequences of Dirac’s theory of the positron]. Zeitschrift für Physik, 98(11–12):714–732, November 1935. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01343663;   http://www.springerlink.
com/content/p710710041x81t32/. English translation in [HE06].

He:2004:IFR

[He04] Jianxun He. An inversion formula of the Radon transform on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 47(3):389–397, September 2004. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Heisenberg:2006:CDT

[HE06] Werner Heisenberg and Hans Euler. Consequences of Dirac’s theory of the positron. ArXiv Physics e-prints, May 4, 2006. URL

http://arxiv.org/pdf/physics/0605038. English translation

by W. Korolevski and H. Kleinert.

Healy:1977:GKH

[Hea77] W. P. Healy. A generalization of the Kramers–Heisenberg dispersion formula. Physical Review A (Atomic, Molecular, and Optical Physics), 16(4):1568–1574, October 1977. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL

http://pra.aps.org/abstract/PRA/v16/i4/p1568_1.

Healy:2011:HCL

[Hea11] Eamonn F. Healy. Heisenberg’s chemical legacy: resonance and the chemical bond. Foundations of Chemistry, 13(1):39–49, April 2011. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10698-011-9104-2.

Haensel:1980:BFK

[HEB^{+}80] R. Haensel, Karl Ehrlich, K. Bethge, H. Soffel, W. von Witsch, and H. Rechenberg. Bücher Filme: Kunz: Synchrotron Radiation Techniques and Applications/Lehmann: Interaction of Radiation with Solids and Elementary Defect Production/Olson u.

Schumacher: Collective Ion Acceleration/Brosche u Sündermann: Tidal Friction and the Earth’s Rotation/Haas u. Koch: Physik-Lehrbuch für Pharmazeuten und Mediziner/Hermann, Meyenn u. Weisskopf: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a., Bd. I: 1919–1929. Physikalische Blätter, 36(10):318–320, October 1980. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19800361013/abstract.

Heelan:1975:HRT

[Hee75] Patrick A. Heelan. Heisenberg and radical theoretic change. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):113–136, ???? 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL

http://www.jstor.org/stable/25170347.

Heelan:1995:QMS

[Hee95] Patrick A. Heelan. Quantum mechanics and the social sciences: After hermeneutics. Science & Education (Springer), 4(2):127– 136, April 1995. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (electronic).

Heisenberg:1922:ADK

[Hei22a] Werner Heisenberg. Die absoluten Dimensionen der Kármánschen Wirbelbewegung. (German) [The absolute dimension of the Kármán vortex motion]. Zeitschrift für Physik, 23(1):363–366, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/

10.1007/BF01327603. English translation in [vdW67, pages 223– 252].

Heisenberg:1922:NLD

[Hei22b] Werner Heisenberg. Nichtlaminare Lösungen der Differentialgleichungen für reibende Flussigkeiten. (German) [Nonlaminar solutions of differential equations for grinding fluids]. In Th. von Kármán and T. Levi-Civita, editors, Vortrage aus dem Gebiete der Hydro- und Aerodynamik (Innsbruck 1922), pages 139–142. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1922. LCCN ????

Heisenberg:1922:QLA

[Hei22c] Werner Heisenberg. Zur Quantentheorie der Linienstruktur und der anomalen Zeemaneffekte. (German) [On the quantum theory of line structure and the anomalous Zeeman effect]. Zeitschrift für Physik, 8(1):273–297, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01329602;   http://www.
springerlink.com/content/p2g7004401m14n38/.

Heisenberg:1923:QDE

[Hei23a] Werner Heisenberg. Quantitatives über die Deformierbarkeit edelgasahnlicher lonen. (German) [Quantitative on the deformability of noble-gas-like ions]. Verh. Dtsch. Phys. Ges. (3), 5(??):7–8, ???? 1923.

Heisenberg:1923:STF

[Hei23b] Werner Heisenberg. Über Stabilität und Turbulenz von Flüssigkeitsströmmen. (German) [On stability and turbulence of fluid flow]. Dissertation, Ludwig-Maximilians-Universität München, Munich, Germany, September 1923. 59 pp.

Heisenberg:1924:BAF

[Hei24a] Werner Heisenberg. Bemerkung zu einer Arbeit von F. v. Wiśniewski: ,,Zur Theorie des Heliums“. (German) [Remark on a paper by F. v. Wiśniewski: “On the theory of helium”]. Zeitschrift für Physik, 25(1):175–176, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01327520;   http:/
/www.springerlink.com/content/g1815p05283257m4/.

Heisenberg:1924:KUG

[Hei24b] Werner Heisenberg. Der Kampf um die Überfahrt. (German) [The battle for the crossing]. Die Spur in ein deutsches Jugendland, 3 (3):37–39, ???? 1924.

Heisenberg:1924:EDI

[Hei24c] Werner Heisenberg. Über den Einfluß der Deformierbarkeit der Ionen auf optische und chemische Konstanten. II: Stabilität und Bildungswärme dreiatomiger Molekeln und Ionen. (German) [On the influence of the deformability of the ions on optical and chemical constants. II: Stability and heat of formation of triatomic molecules and ions]. Zeitschrift für Physik, 26(1):196–204, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01327329;   http://www.
springerlink.com/content/rw1356k1t240k063/.

Heisenberg:1924:AFR

[Hei24d] Werner Heisenberg. Über eine Abänderung der formalen Regeln der Quantentheorie beim Problem der anomalen Zeeman-Effekte. (German) [On a modification of the formal rules of quantum theory in the problem of the anomalous Zeeman effect]. Zeitschrift für Physik, 26(1):291–307, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01327336;   http:/
/www.springerlink.com/content/v2803t507860nw30/.

Heisenberg:1924:STF

[Hei24e]  Werner Heisenberg.  Über Stabilität und Turbulenz von

Flüssigkeitsströmmen. (German) [On stability and turbulence of fluid flow]. Annalen der Physik (1900), 379(15):577–627, ????

1924. ISSN 1521-3889. URL http://adsabs.harvard.edu/abs/

1924AnP...379..577H. An English translation is available in [Hei51g].

Heisenberg:1925:BWH

[Hei25a] Werner Heisenberg. Besprechung: Wilhelm Hanle. Über magnetische Beeinflussung der Polarisation der Resonanzfluoreszenz. (German) [Book review: Wilhelm Hanle. On magnetic influence on the polarization of the resonance fluorescence]. Physikalische Berichte, 6(??):806–807, ???? 1925.

Heisenberg:1925:AKF

[Hei25b] Werner Heisenberg. Über eine Anwendung des Korrespondenzprinzips auf die Frage nach der Polarisation des Fluoreszenzlichtes. (German) [An application of the Correspondence Principle to the question of the polarization of fluorescence light]. Zeitschrift für Physik, 31(1):617–626, February/April 1925. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF02980618;   http://www.springerlink.
com/content/r8tl178265164143/.

Heisenberg:1925:QUK

[Hei25c] Werner Heisenberg. Über quantentheoretische Umdeutung kinematischer und mechanischer Beziehungen. (German) [On quantum theoretical reinterpretation of kinematic and mechanical relations]. Zeitschrift für Physik, 33(1):879–893, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01328377;   http://www.springerlink.
com/content/w22847j178u03029/. English translation in [vdW67,

pages 261–276].

Heisenberg:1925:QMA

[Hei25d] Werner Heisenberg. Zur Quantentheorie der Multiplettstruktur und der anomalen Zeemaneffekte. (German) [On the quantum theory of multiplet structure and the anomalous Zeeman effect]. Zeitschrift für Physik, 32(1):841–860, December 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01331720;
http://www.springerlink.com/content/r85rk473x733t55v/.

Heisenberg:1926:MRQ

[Hei26a] Werner Heisenberg. Mehrkörperproblem und Resonanz in der Quantummechanik. (German) [Many-body problem and resonance in quantum mechanics]. Zeitschrift für Physik, 38(6–7): 411–426, June 1926. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01397160;   http://www.springerlink.
com/content/mth33737521280v1/.

Heisenberg:1926:QGQa

[Hei26b]  Werner Heisenberg.  Quantenmechanik. (German) [Quan-

tum mechanics]. Naturwissenschaften, 14(45):989–994, ???? 1926. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
t1770183l1922681.

Heisenberg:1926:QGQb

[Hei26c] Werner Heisenberg. Quantenmechanik. (German) [Quantum mechanics]. Zeitschrift für Angewandte Chemie, 39(??):1177, ???? 1926. ISSN 0932-2132. Abstract of [Hei26b].

Heisenberg:1926:SQG

[Hei26d] Werner Heisenberg. Schwankungserscheinungen und Quantenmechanik. (German) [Fluctuation phenomena and quantum mechanics]. Zeitschrift für Physik, 40(7):501–506, July 1926. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01440827;
http://www.springerlink.com/content/n102340512v9254q/.

Heisenberg:1926:SAZ

[Hei26e] Werner Heisenberg. Über die Spektra von Atomsystemen mit zwei Elektronen. (German) [On the spectra of atomic systems with two electrons]. Zeitschrift für Physik, 39(7–8):499– 518, July 1926. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01322090;   http://www.springerlink.
com/content/r1m57104030g5565/.

Heisenberg:1926:QKM

[Hei26f] Werner Heisenberg. Über quantentheoretische Kinematik und Mechanik. (German) [On quantum theoretical kinematics and mechanics]. Mathematische Annalen, 95(1):683–705, 1926. CODEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Heisenberg:1927:BMB

[Hei27a] Werner Heisenberg. Besprechung: Max Born. Probleme der Atomdynamik. (German) [Book review: Max Born. Problems of Atomic Dynamics]. Fysisk Tidsskrift, 25(??):96–97, ???? 1927.

Heisenberg:1927:BGB

[Hei27b] Werner Heisenberg. Besprechung: S. Goudsmit und E. Back. Die Kopplung der Quantenvektoren bei Neon, Argon und einigen Spektren der Kohlenstoffgruppe. (German) [Book review: S. Goudsmit und E. Back. Coupling of the quantum vectors in neon, argon, and some spectra of the carbon group]. Physikalische Berichte, 8(??): 470, ???? 1927.

Heisenberg:1927:MRQ

[Hei27c] Werner Heisenberg. Mehrkörperprobleme und Resonanz in der Quantenmechanik. II. (German) [Few-body problems and resonance in quantum mechanics. II]. Zeitschrift für Physik, 41 (3):239–267, August 1927. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01391241;   http://www.
springerlink.com/content/ln0385427144375w/.

Heisenberg:1927:QMG

[Hei27d] Werner Heisenberg. Quantentheoretische Mechanik. (German) [Quantum theoretical mechanics]. Deutsche Mathematiker-Vereinigung. Jahresbericht, 36(??):24–25, ???? 1927.

Heisenberg:1927:AIQ

[Hei27e] Werner Heisenberg. Über den anschaulichen Inhalt der quantentheoretischen Kinematik und Mechanik. (German) [On the ideological content of quantum theoretical kinematics and mechanics]. Zeitschrift für Physik, 43(3–4):172–198, March 1927. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01397280;
http://www.springerlink.com/content/t8173612621026q5/.

English translation in [WZ83, pages 62–84]. This is the paper in which Heisenberg introduces the famous Uncertainty Principle (Unschärferelation or Ungenauigkeit (impreciseness)).

Heisenberg:1927:GGQ

[Hei27f] Werner Heisenberg. Über die Grundprinzipien der ,,Quantenmechanik“. (German) [On the basic principles of “quantum mechanics”]. Forschungen und Fortschritte, 3(??):83, ???? 1927.

Heisenberg:1928:BBG

[Hei28a] Werner Heisenberg. Besprechung: Birthwistle, G., The new Quantum Mechanics. Naturwissenschaften, 16(??):318, ???? 1928. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1928:BHW

[Hei28b] Werner Heisenberg. Besprechung: Hermann Weyl, Gruppentheorie und Quantenmechanik. (German) [Book review: Hermann Weyl, Group Theory and Quantum Mechanics]. Deutsche Literaturzeitung, 49(??):2473–2474, ???? 1928.

Heisenberg:1928:DSC

[Hei28c] Werner Heisenberg. Discussione sulla communicazione Bohr. (Italian) [Discussions of Bohr’s communication]. In ????, editor, Atti del Congresso Internazionale dei Fisici, 11–20 Settembre 1927-V, Como–Pavia–Roma, pages 593–594, 597. Nicola Zanichelli, Bologna, Italy, 1928. LCCN ????

Heisenberg:1928:QFG

[Hei28d] Werner Heisenberg. Zur Quantentheorie des Ferromagnetismus. (German) [On the quantum theory of ferromagnetism]. In P. Debye, editor, Probleme der modernen Physik: Arnold Sommerfeld zum 60. Geburtstage gewidmet von semen Schülern. (German) [Problems of modern physics: Arnold Sommerfeld’s 60th Birthday, dedicated by his pupils], pages 114–122. S. Hirzel, Leipzig, Germany, 1928. LCCN ????

Heisenberg:1928:TFG

[Hei28e] Werner Heisenberg. Zur Theorie des Ferromagnetismus. (German) [On the theory of ferromagnetism]. Zeitschrift für Physik, 49(9–10):619–636, September 1928. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01328601;   http://www.
springerlink.com/content/w77836623761874g/.

Heisenberg:1929:EQG

[Hei29a] Werner Heisenberg. Die Entwicklung der Quantentheorie 1918– 1928. (German) [The development of the quantum theory 1918– 1928]. Naturwissenschaften, 17(26):490–496, ???? 1929. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/p1179238357h7776.

Heisenberg:1929:HEQ

[Hei29b] Werner Heisenberg. Die historische Entwicklung der Quantentheorie. (German) [The historical development of quantum theory]. Fr. Oldecops Erben, Oschatz, Germany, 1929. ???? pp.

Heisenberg:1929:TLP

[Hei29c] Werner Heisenberg. Two Lectures: 1. The Present Situation in the Theory of Elementary Particles. 2. The Electron Theory of Superconductivity. Cambridge University Press, Cambridge, UK, 1929. 52 pp. LCCN ????

Heisenberg:1930:BGE

[Hei30a] E. Heisenberg. Bericht über die Gründung und die erste Tagung der Deutschen Gesellschaft für photographische Forschung (23. bis 25. Mai 1930). (German) [Report on the establishment and first meeting of the German Society for Photographic Research (23 to May 25, 1930)]. Naturwissenschaften, 18(52):1130–1131, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/u5158064h268j564.

Heisenberg:1930:BBP

[Hei30b] Werner Heisenberg. Besprechung: The principles of Quantum mechanics. Von P. A. M. Dirac. Metallwirtschaft, 9(??):988, ???? 1930.

Heisenberg:1930:BPM

[Hei30c] Werner Heisenberg. Buchbesprechung: P. A. M. Dirac The Principles of Quantum Mechanics. Metallwirtschaft, 9(??):988–??, ???? 1930.

Heisenberg:1930:PPQa

[Hei30d] Werner Heisenberg. Die physikalischen Prinzipien der Quantentheorie. (German) [The physical principles of quantum mechanics]. Hirzel, Leipzig, Germany, 1930. viii + 115 pp. LCCN ????

Heisenberg:1930:SEG

[Hei30e] Werner Heisenberg. Die Selbstenergie des Elektrons. (German) [The self-energy of the electron]. Zeitschrift für Physik, 65(1–2):4–13, November 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01397404;   http://www.
springerlink.com/content/t61j113n618m2j11/.

Heisenberg:1930:FTF

[Hei30f] Werner Heisenberg. Fortschritte in der Theorie des Ferromagnetismus. (German) [Advances in the theory of ferromagnetism]. Metallwirtschaft, 9(??):843–844, ???? 1930.

Heisenberg:1930:KQGb

[Hei30g] Werner Heisenberg. Kausalgesetz und Quantenmechanik. (German) [Causality and quantum mechanics]. Z. angew. Phys., 43 (??):853–854, ???? 1930. Abstract of [Hei31e].

Heisenberg:1930:KQGa

[Hei30h] Werner Heisenberg. Kausalität und Quantenmechanik. (German) [Causality and quantum mechanics]. Naturwissenschaften, 18(??): 1068–??, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1930:PPQb

[Hei30i] Werner Heisenberg. The Physical Principles of the Quantum Theory. University of Chicago Press, Chicago, IL, USA, 1930. ???? pp. LCCN ???? English translation of [Hei30d].

Heisenberg:1931:BSG

[Hei31a] Werner Heisenberg. Bemerkungen zur Strahlungstheorie. (German) [Remarks on radiation theory]. Annalen der Physik (1900), 401(3):338–346, March 1931. ISSN 1521-3889. URL http:/

/onlinelibrary.wiley.com/doi/10.1002/andp.19314010305/
abstract.

Heisenberg:1931:BHW

[Hei31b] Werner Heisenberg. Besprechung: Hermann Weyl, Gruppentheorie und Quantenmechanik, 2. umgearb. Aufl. (German) [Book review: Hermann Weyl, Group Theory and Quantum Mechanics, 2nd revised edition]. Deutsche Literaturzeitung, 52(??):424–425, ???? 1931.

Heisenberg:1931:BLB

[Hei31c] Werner Heisenberg. Besprechung: Leon Brillouin, Les Statistiques Quantiques et leurs applications. (German) [Book review: Leon Brillouin, Quantum statistics and their applications]. Physikalische Zeitschrift, 32(??):150–151, ???? 1931. CODEN PHZTAO. ISSN 0369-982X.

Heisenberg:1931:RUM

[Hei31d] Werner Heisenberg. Die Rolle der Unbestimmtheitsrelationen in der modernen Physik. (German) [The role of the uncertainty relations in the modern physics]. Monatshefte für Mathematik und Physik, 38(1):365–372, 1931. CODEN MMPHA8. ISSN 0026-9255 (print), 1436-5081 (electronic).

Heisenberg:1931:KQG

[Hei31e] Werner Heisenberg. Kausalgesetz und Quantenmechanik. (German) [Causality and quantum mechanics]. Erkenntnis, 2(1):

172–182, ???? 1931. CODEN ????  ISSN ????  URL

http://www.springerlink.com/content/r1v3282l62551554/
fulltext.pdf.

Heisenberg:1931:NAH

[Hei31f] Werner Heisenberg. Notiz zur Arbeit des Herrn N. Tunazima: Zum Ferromagnetismus. (German) [Note on the work of Mr. N. Tunazima: On the ferromagnetism]. Zeitschrift für Physik, 68(9–10):720, September 1931. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01390593;   http://www.
springerlink.com/content/k34635h0n1n05026/.

Heisenberg:1931:PRA

[Hei31g] Werner Heisenberg. Positivismus und reale Außenwelt. Von M. Planck. Vortrag, gehalten am 12. November 1930 im Harnack-Haus der Kaiser Wilhelm-Gesellschaft zur Förderung der Wissenschaften. 35 Seiten. Akademische Verlagsgesellschaft, Leipzig 1931. (German) [Positivism and the real world outside. On Max Planck. Lecture, held on 12 November 1930 at the Harnack House the Kaiser Wilhelm Society for the F romoting the Sciences. 35 pages. academic Verlagsgesellschaft, Leipzig, 1931]. Angewandte Chemie, 44(10):189, March 1931. CODEN ANCEAD. ISSN 1521-

3757. URL http://onlinelibrary.wiley.com/doi/10.1002/
ange.19310441010/abstract.

Heisenberg:1931:PMP

[Hei31h] Werner Heisenberg. Probleme der modernen Physik. (German) [Problems of modern physics]. Berliner Tageblatt, 60(607):??, December 25, 1931. Morgen-Ausgabe, 1 Beiblatt [Morning Edition, 1 Supplement].

Heisenberg:1931:ISR

[Hei31i]  Werner Heisenberg.  Über die inkohärente Streuung von

Röntgenstrahlen. (German) [On the incoherent diffusion of Xrays]. Physikalische Zeitschrift, 32(19):737–740, October 1, 1931. CODEN PHZTAO. ISSN 0369-982X.

Heisenberg:1931:EES

[Hei31j] Werner Heisenberg. Über Energieschwankungen in einem Strahlungsfeld. (German) [Energy fluctuations in a radiation field]. Sächs. Akad. Wiss. (Leipzig), Berichte d. math.-phys. Kl., 83(??):3–9, ???? 1931.

Heisenberg:1931:PAG

[Hei31k] Werner Heisenberg. Zum Paulischen Ausschließungsprinzip. (German) [On Pauli’s Exclusion Principle]. Annalen der Physik (1900), 402(7):888–904, 1931. ISSN 1521-3889. URL http:/

/onlinelibrary.wiley.com/doi/10.1002/andp.19314020710/
abstract.

Heisenberg:1931:TMM

[Hei31l] Werner Heisenberg. Zur Theorie der Magnetostriktion und der Magnetisierungskurve. (German) [On the theory of magnetostriction and magnetization curve]. Zeitschrift für Physik, 69(5–6): 287–297, May 1931. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01391350;   http://www.springerlink.
com/content/t88384n177787623/.

Heisenberg:1932:BRB

[Hei32a] Werner Heisenberg. Book review: Bateman, Partial Differential Equation of Mathematical Physics. Physikalische Zeitschrift, 33 (??):454, ???? 1932. CODEN PHZTAO. ISSN 0369-982X.

Heisenberg:1932:CDS

[Hei32b] Werner Heisenberg. Contribution to discussion on the structure of simple molecules. In ????, editor, Chemistry at the Centenary (1931) Meeting of the British Association for the Advancement of Science, pages 247–248. W. Heifer & Sons, Cambridge, UK, 1932. LCCN ????

Heisenberg:1932:NPP

[Hei32c] Werner Heisenberg. The Nobel Prize in Physics 1932. No-

belprize.org, 1932. URL http://www.nobelprize.org/nobel_
prizes/physics/laureates/1932/. The Nobel Prize in Physics

1932 was awarded to Werner Heisenberg “for the creation of quantum mechanics, the application of which has, inter alia, led to the discovery of the allotropic forms of hydrogen”.

Heisenberg:1932:PMP

[Hei32d] Werner Heisenberg. Probleme der modernen Physik. (German) [Problems of modern physics]. Deutsche Optische Wochenschrift, 18(2):27–28, ???? 1932. Reprint of [Hei31h].

Heisenberg:1932:TUH

[Hei32e] Werner Heisenberg. Theoretische Überlegungen zur Höhenstrahlung. (German) [Theoretical considerations of cosmic radiation]. Annalen der Physik (1900), 405(4):430–452, 1932. ISSN 1521-

3889. URL http://onlinelibrary.wiley.com/doi/10.1002/
andp.19324050404/abstract.

Heisenberg:1932:BAG

[Hei32f] Werner Heisenberg. Über den Bau der Atomkerne. I. (German) [On the structure of the nucleus. I]. Zeitschrift für Physik, 77(1–2):1–11, January 1932. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01342433;   http://www.
springerlink.com/content/gt251k5258372318/. See also par-

tial English translation [Hei65g].

Heisenberg:1932:BAI

[Hei32g] Werner Heisenberg. Über den Bau der Atomkerne. II. (German) [On the structure of the nucleus. II]. Zeitschrift für Physik, 78(3–4):156–164, March 1932. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01337585;   http://www.
springerlink.com/content/m77927k66q803736/.

Heisenberg:1932:DUH

[Hei32h] Werner Heisenberg. Über die durch Ultrastrahlung hervorgerufenen Zertrümmerungsprozesse. (German) [On the cause of fragmentation processes from cosmic rays]. Naturwissenschaften, 20 (21):365–366, ???? 1932. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/j3134692417w4647.

Heisenberg:1932:KGNa

[Hei32i] Werner Heisenberg. Über Kernmechanik. (German) [On nuclear mechanics]. Göttinger Nachrichten, ??(??):??, ???? 1932. Sonderausgabe [special edition].

Heisenberg:1932:KGNb

[Hei32j] Werner Heisenberg. Über Kernmechanik. (German) [On nuclear mechanics]. Physikalische Blätter, 18(??):557–558, ???? 1932. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL https://www.aip.org/history/heisenberg/
bibliography/1960-64.htm. Reprint of [Hei32i].

Heisenberg:1933:DQM

[Hei33a] Werner Heisenberg. The development of quantum mechanics. Nobel Lectures Web site, December 11, 1933. URL https://www.

nobelprize.org/prizes/physics/1932/heisenberg/lecture/

.

Heisenberg:1933:BAI

[Hei33b] Werner Heisenberg. Über den Bau der Atomkerne. III. (German) [On the structure of the nucleus. III]. Zeitschrift für Physik, 80(9–10):587–596, September 1933. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01335696;   http://www.
springerlink.com/content/j05335316h41783t/. Partial En-

glish translation in [?].

Heisenberg:1933:SRM

[Hei33c] Werner Heisenberg. Über die Streuung von Rontgenstrahlen an MolekUlen und Kristallen. (German) [On the scattering of X-rays of molecules and crystals]. Ergeb. Tech. Röntgenkd., 3(??):26–31, ???? 1933.

Heisenberg:1933:GPN

[Hei33d] Werner Heisenberg. Zur Geschichte der physikalischen Naturerklärung. (German) [The history of the physical nature of explanation]. Ber. Sächs. Akad. Wiss. (Leipzig) math. phys. Kl., 85(??):29–40, ???? 1933. English translation in [Hei52k, pages 27–40].

Heisenberg:1934:BDT

[Hei34a]  Werner Heisenberg.  Bemerkungen zur Diracschen Theo-

rie des Positrons. (German) [Remarks on the Dirac theory of the positron]. Zeitschrift für Physik, 90(3–4):209–231, March 1934. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01333516;   http://www.springerlink.
com/content/vr14wp0778r73025/. See correction [Hei34b].

Heisenberg:1934:BAB

[Hei34b] Werner Heisenberg. Berichtigung zu der Arbeit: ,,Bemerkungen zur Diracschen Theorie des Positrons“. (German) [Correction on the work: “Remarks on the Dirac theory of the positron”]. Zeitschrift für Physik, 92(9–10):692, September 1934. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01340782;
http://www.springerlink.com/content/l0710nm9813n7622/.

See [Hei34a].

Heisenberg:1934:CTG

[Hei34c] Werner Heisenberg. Considérations théoriques générales sur la structure du noyau. (French) [General theoretical considerations of the structure of the nucleus]. In Cockcroft et al. [CCJ^{+}34], pages 289–335. LCCN ???? Publiés par la commission administrative de l’institut.

Heisenberg:1934:EQG

[Hei34d] Werner Heisenberg. Die Entwicklung der Quantenmechanik. (German) [The development of quantum mechanics]. In Die moderne Atomtheorie: Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Lectures on modern atomic theory: the acceptance of the Nobel Prize in 1933 in Stockholm], pages 1–18. S. Hirzel, Leipzig, 1934. LCCN ????

Heisenberg:1934:EMS

[Hei34e] Werner Heisenberg. Über die mit der Entstehung von Materie aus Strahlung verknüpften Ladungsschwankungen. (German) [Charge fluctuations connected with the creation of matter]. Sächs. Akad. Wiss. (Leipzig), Berichte d. math.-phys. Kl., 86(??):317–322, ???? 1934.

Heisenberg:1934:WGEb

[Hei34f] Werner Heisenberg. Wandlungen der Grundlagen der exakten Naturwissenschaft in jüngster Zeit. (German) [Changes in the fundamentals of the exact science in recent times]. Angewandte Chemie, 47(45):763, 1934. CODEN ANCEAD. ISSN 1521-3757.

Heisenberg:1934:WGEc

[Hei34g] Werner Heisenberg. Wandlungen der Grundlagen der exakten Naturwissenschaft in jüngster Zeit. (German) [Changes in the fundamentals of the exact science in recent times]. Naturwissenschaften, 22(40):669–675, ???? 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/l54t407881124377.

Heisenberg:1934:WGEd

[Hei34h] Werner Heisenberg. Wandlungen der Grundlagen der exakten Naturwissenschaft in jüngster Zeit. (German) [Changes in the fundamentals of the exact science in recent times]. Naturwissenschaften, 22(41):700, ???? 1934. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/qt45k7u7146n008m.

Heisenberg:1934:WGEa

[Hei34i] Werner Heisenberg. Wandlungen der Grundlagen der exakten Naturwissenschaft in jüngster Zeit. (German) [Recent changes in the foundations of exact science]. Angewandte Chemie, 47(41): 697–702, 1934. CODEN ANCEAD. ISSN 1521-3757. English translation in [Hei52k, 697–702].

Heisenberg:1934:WTF

[Hei34j] Werner Heisenberg. Wissenschaft und technischer Fortschritt. (German) [Science and technical progress]. Stahl und Eisen, 54 (??):749–752, ???? 1934.

Heisenberg:1935:BTA

[Hei35a] Werner Heisenberg. Bemerkungen zur Theorie des Atomkerns. (German) [Remarks on the theory of the atomic nucleus]. In ????, editor, Pieter Zeeman 1865–25 Mei-1935 Verhandelingen op 25 Mei 1935 Aangeboden aan Prof Dr P Zeeman. (Dutch) [Pieter Zeeman 1865–25 May 1935 Discourses on 25 May 1935. Presented to Prof. Dr. P Zeeman], pages 108–116. Martinus Nijhoff, The Hague, The Netherlands, 1935. LCCN ????

Heisenberg:1935:BMP

[Hei35b] Werner Heisenberg. Besprechung: Max Planck, Die Physik im Kampf um die Weltanschauung. (German) [Book review: Max Planck, The physics in the struggle of belief]. Naturwissenschaften, 23(??):321, ???? 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1935:SLA

[Hei35c] Werner Heisenberg. Die Struktur der leichten Atomkerne. (German) [The structure of the light atomic nuclei]. Zeitschrift für Physik, 96(7–8):473–484, July 1935. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01337702;   http://www.
springerlink.com/content/v1145371h3464r31/.

Heisenberg:1935:NBF

[Hei35d] Werner Heisenberg. Niels Bohr zum fünfzigsten Geburtstage am 7. Oktober 1935. (German) [Niels Bohr’s fiftieth birthday, 7 October 1935]. Naturwissenschaften, 23(40):679, October 4, 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/g341u926u7452621.

Heisenberg:1935:UAK

[Hei35e] Werner Heisenberg. Udviklingen af Kvanteteoriens principielle Grundlag efter 1925. (Danish) [Development of quantum theory’s principle foundations after 1925]. Fysik Tidsskrift, 33(??):96–101, ???? 1935.

Heisenberg:1935:WGN

[Hei35f] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1935. 45 pp. LCCN ????

Heisenberg:1935:WHA

[Hei35g] Werner Heisenberg. Werner Heisenberg [autobiographical note]. In ????, editor, Les Prix Nobel en 1933. (French) [The Nobel Prizes in 1933], page 85. P. A. Norstedt & Söner, Stockholm, Sweden, 1935. LCCN ????

Heisenberg:1936:BEP

[Hei36a] Werner Heisenberg. Besprechung: Enrico Persico, Fondamenti della Meccanica Atomica. (German) [Book review: Enrico Persico, Fundamentals of Atomic Mechanics]. Naturwissenschaften, 24(??):639, ???? 1936. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1936:BEF

[Hei36b] Werner Heisenberg. Besprechung: Erwin Fues, Beugungsversuche mit Materiewellen. Einführung in die Quantenmechanik. (German) [Book review. Erwin Fues, Diffraction experiments with matter waves. Introduction to quantum mechanics]. Deutsche Literaturzeitung, 57(??):684–686, ???? 1936.

Heisenberg:1936:BPJ

[Hei36c] Werner Heisenberg. Besprechung: P. Jordan, Anschauliche Quantentheorie. (German) [Book review: P. Jordan, Illustrative quantum theory]. Naturwissenschaften, 24(??):558–559, ???? 1936. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1936:DSE

[Hei36d] Werner Heisenberg. Der Durchgang sehr energiereicher Korpuskeln durch den Atomkern. (German) [The passage of highenergy corpuscles through the nucleus]. Ber. Sächs, Akad. Wiss., 89(??):369–384, ???? 1936.

Heisenberg:1936:LHA

[Hei36e] Werner Heisenberg. Leipziger Hochschullehrer bei der Arbeit. Werner Heisenberg der Forscher an der Grenze der Wissenschaft [Interview]. (German) [Leipzig University teachers at work. Werner Heisenberg, the researchers at the frontier of science]. Neue Leipziger Zeitung, 26(??):4, January 26, 1936.

Heisenberg:1936:PFM

[Hei36f] Werner Heisenberg. Prinzipielle Fragen der modernen Physik. (German) [Fundamental questions of modern physics]. In ????, editor, Neuere Fortschritte in den exakten Wissenschaften. Fünf Wiener Vorträge. Dritter Zyklus. (German) [Recent advances in the exact sciences. Five Vienna talks. Third cycle], pages 91–102. Franz Deuticke, Leipzig, Germany, 1936. LCCN ???? English translation in [Hei52l].

Heisenberg:1936:SKS

[Hei36g] Werner Heisenberg. Über die ‘Schauer’ in der Kosmischen Strahlung. (German) [On the ‘showers’ in the cosmic radiation]. Forsch. Fortscher., 12(??):341–342, ???? 1936. English translation in [Hei37a].

Heisenberg:1936:WGN

[Hei36h] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1936. 63 pp. LCCN ????

Heisenberg:1936:ADJ

[Hei36i] Werner Heisenberg. Zum Artikel: Deutsche und jüdische Physik. Entgegnung von Prof. Dr. Heisenberg. (German) [On the article: German and Jewish physics. Reply by Prof. Dr. Heisenberg]. Völkischer Beobachter (Northern edition (A)), 59(6):6, February 28, 1936. This is a reply to [Men36]. See also the American news story [Ano36].

Heisenberg:1936:TSH

[Hei36j] Werner Heisenberg. Zur Theorie der ‘Schauer’ in der Höhenstrahlung. (German) [On the theory of the ‘showers’ in cosmic radiation]. Zeitschrift für Physik, 101(9–10):533–540, September 1936. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/

article/10.1007/BF01349603;   http://www.springerlink.
com/content/k54465232750n11r/.

Heisenberg:1937:CCR

[Hei37a] Werner Heisenberg. Concerning cosmic-ray showers. Research and Progress, 3(??):32–34, ???? 1937. English translation of [Hei36g].

Heisenberg:1937:DSE

[Hei37b] Werner Heisenberg. Der Durchgang sehr energiereicher Korpuskeln durch den Atomkern. (German) [The passage of highenergy corpuscles through the nucleus]. Naturwissenschaften, 25 (46):749–750, ???? 1937. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/w31h75005681021m.

Heisenberg:1937:FGN

[Hei37c] Werner Heisenberg. Fortschritte auf dem Gebiet der Neutronenstrahlung. (German) [Progress in the field of neutron radiation]. Deutsche Allgemeine Zeitung (Reichs-Ausgabe), 107–108 (??):6, March 6, 1937. Unterhaltungsblatt [Entertainment Journal].

Heisenberg:1937:FTA

[Hei37d] Werner Heisenberg. Fortschritte in der Theorie des Atomkerns (Abstrakt). (German) [Advances in the theory of the atomic nucleus (abstract)]. Jahresbericht des Physikalischen Vereins zu Frankfurt am Main für das Rechnungsjahr 1937/38, ??(??):27–??, ???? 1937.

Heisenberg:1937:GAN

[Hei37e] Werner Heisenberg. Gedanken der antiken Naturphilosophie in der modernen Physik. (German) [Thoughts of the ancient philosophy of nature in modern physics]. Die Antike. Z. Kunst. u. Kultur des kl. Altertums, 13(??):118–124, ???? 1937. English translation in [Hei52e].

Heisenberg:1937:TUU

[Hei37f] Werner Heisenberg. Theoretische Untersuchungen zur Ultrastrahlung. (German) [Theoretical studies on ultra radiation]. Verh. Dtsch. Phys. Ges. (3), 18(??):50–??, ???? 1937.

Heisenberg:1938:SEM

[Hei38a] Werner Heisenberg. Das schwere Elektron (Mesotron) und seine Rolle in der Höhenstrahlung. (German) [The heavy electron (mesotron) and its role in the cosmic radiation]. Angewandte Chemie, 52(??):41–42, ???? 1938. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic). Physikalisches Colloquium Hamburg 1 December 1938.

Heisenberg:1938:DSEb

[Hei38b] Werner Heisenberg. Der Durchgang sehr energiereicher Korpuskeln durch den Atomkern (Abstrakt). (German) [Passage of very high energy particles through atomic nuclei (abstract)]. Verh. Dtsch. Phys. Ges. (3), 19(??):2, ???? 1938.

Heisenberg:1938:DSEa

[Hei38c] Werner Heisenberg. Der Durchgang sehr energiereicher Korpuskeln durch den Atomkern. (German) [Passage of very high energy particles through atomic nuclei]. Il Nuovo Cimento (8), 15(1):31–34, January 1938. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://www.springerlink.

com/content/m7h603q384211473/.

Heisenberg:1938:ADK

[Hei38d] Werner Heisenberg. Die Absorption der durchdringenden Komponente der Höhenstrahlung. (German) [Absorption of the penetrating component of cosmic radiation]. Annalen der Physik (1900), 425(7):594–599, 1938. ISSN 1521-3889.

Heisenberg:1938:GAT

[Hei38e] Werner Heisenberg. Die gegenwärtigen Aufgaben der theoretischen Physik. (German) [The current problems of theoretical physics]. Scientia (3), 32(??):61–69, ???? 1938.

Heisenberg:1938:GAB

[Hei38f] Werner Heisenberg. Die Grenzen der Anwendbarkeit der bisherigen Quantentheorie. (German) [The limits of applicability of the present quantum theory]. Zeitschrift für Physik, 110(3–4):251– 266, March 1938. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01342872;   http://www.springerlink.
com/content/uh55r4vm27168125/.

Heisenberg:1938:MPA

[Hei38g] Werner Heisenberg. Max Planck. Achtzig Jahre alt. (German) [Max Planck. Eighty years old]. Kölnische Zeitung, 201(??):1–2, April 23, 1938.

Heisenberg:1938:TEA

[Hei38h] Werner Heisenberg. Über die in der Theorie der Elementarteilchen auftretende universelle Länge. (German) [On the theory of elementary particles in the occurring universal length]. Annalen der Physik (1900), 424(1–2):20–33, 1938. ISSN 1521-3889.

Heisenberg:1938:WQW

[Hei38i] Werner Heisenberg. Wahrscheinlichkeitsaussagen in der Quantentheorie der Wellenfelder. (German) [Probability statements in the quantum theory of wave fields]. Actual. Sci. & Ind. (Paris), 743/I (??):45–51, ???? 1938.

Heisenberg:1939:BAS

[Hei39a] Werner Heisenberg. Besprechung: Arnold Sommerfeld, Atombau und Spektrallinien. 2. Band. (German) [Book review: Arnold Sommerfeld, Atomic structure and spectral lines]. Naturwissenschaften, 27(??):741–742, ???? 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1939:BRO

[Hei39b] Werner Heisenberg. Besprechung: R. Orthner, Die drei Grundphänomene der Physik und ihre Deutung. (German) [Book review: R. Orthner, Three basic phenomena of physics and their interpretation]. Z. techn. Phys, 20(??):326, ???? 1939.

Heisenberg:1939:AHS

[Hei39c] Werner Heisenberg. De atoomkern in hare samenstelling. (Dutch) [The nucleus in its composition]. Nederlands Tijdschrift voor Natuurkunde, 6(??):89–98, ???? 1939.

Heisenberg:1939:EJZ

[Hei39d] Werner Heisenberg. Die Entwicklung der jüngsten Zweige der Atomphysik in Deutschland. (German) [The development of the youngest branches of atomic physics in Germany]. Forschungen und Fortschritte, 15(??):241–244, ???? 1939. English translation in [Hei40b].

Heisenberg:1939:MTE

[Hei39e] Werner Heisenberg. Die Möglichkeit der technischen Energiegewinnung aus der Uranspaltung. (German) [The possibility of technical acquisition of energy from uranium fission]. Kernphysikalische Forschungsberichte G-39, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, December 6, 1939. 19 pp.

Heisenberg:1939:EKS

[Hei39f] Werner Heisenberg. Elementarteilchen und kosmische Strahlung (Abstrakt). (German) [Elementary particles and cosmic radiation (abstract)]. Verhandlungen der Deutschen Physikalischen Gesellschaft. 3. Reihe, 20(??):9, ???? 1939. CODEN VDPEAZ. ISSN 0372-5448.

Heisenberg:1939:TESb

[Hei39g] Werner Heisenberg. On the theory of explosion showers in cosmic rays (abstract). Reviews of Modern Physics, 11(3–4):241, July 1939. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.11.241;   http://rmp.aps.org/abstract/
RMP/v11/i3-4/p241_1.

Heisenberg:1939:VKS

[Hei39h] Werner Heisenberg. Verwandlung der kosmischen Strahlung in der Atmosphäre (Abstrakt). (German) [Transformation of cosmic radiation in the atmosphere (abstract)]. Verhandlungen der Deutschen Physikalischen Gesellschaft. 3. Reihe, 20(??):21–22, ???? 1939. CODEN VDPEAZ. ISSN 0372-5448.

Heisenberg:1939:VGF

[Hei39i] Werner Heisenberg. Vorwort. (German) [Foreword]. In Louis de Broglie, editor, Licht und Materie, pages 7–8. H. Groverts-Verlag, Hamburg, Germany, 1939. LCCN ????

Heisenberg:1939:TESa

[Hei39j] Werner Heisenberg. Zur Theorie der explosionsartigen Schauer in der kosmischen Strahlung. II. (German) [On the theory of explosion showers in cosmic rays. II]. Zeitschrift für Physik, 113(1–2):61–86, January 1939. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01371656;   http://www.
springerlink.com/content/l4811k547275n054/.

Heisenberg:1940:BMT

[Hei40a] Werner Heisenberg. Bericht über die Möglichkeit technischer Energiegewinnung aus der Uranspaltung (II). (German) [Report on the technical possibility of energy from fission of uranium (II)]. Kernphysikalische Forschungsberichte G-40, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, February 29, 1940. 22 pp.

Heisenberg:1940:DMR

[Hei40b] Werner Heisenberg. The development of the most recent branches of atomic physics in Germany. Research and Progress, 6(??):33–42, ???? 1940. English translation of [Hei39d].

Heisenberg:1941:BVS

[Hei41a] Werner Heisenberg. Bericht über die Versuche mit Schichtenanordnungen von Präparat 38 und Paraffin am Kaiser Wilhelm Institut für Physik in Berlin-Dahlem. (German) [Report on experiments with a layer arrangement of preparation 38 and paraffin at the Kaiser Wilhelm Institute for Physics in Berlin-Dahlem]. Kernphysikalische Forschungsberichte G-93, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, May 1941.

Heisenberg:1941:CPG

[Hei41b] Werner Heisenberg. Chemisch-Physikalische Gesellschaft in Wien (gemeinsam mit dem Gauverein Ostmark der Deutschen physikalischen Gesellschaft) Sitzung am 30. April 1941. (German) [Chemical-Physical Society in Vienna (together with the German Ostmark Gauverein physical Society) meeting on 30 April 1941]. Angewandte Chemie, 54(29–30):348, 1941. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1941:DKH

[Hei41c] Werner Heisenberg. Die durchdringende Komponente der Höhenstrahlung. (German) [The penetrating component of cosmic rays]. Angewandte Chemie, 54(??):348, ???? 1941. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1941:GNF

[Hei41d] Werner Heisenberg. Die Goethe’sche und die Newton’sche Farbenlehre im Lichte der modernen Physik. (German) [Goethe’s and Newton’s theory of color in the light of modern physics]. Geist der Zeit (N.F.), 19(??):261–275, ???? 1941. English translation in [Hei52o].

Heisenberg:1941:TNB

[Hei41e] Werner Heisenberg. Theoretischer Nachtrag (zum Bericht von E. Fischer). (German) [Theoretical appendix to E. Fischer’s report]. Kernphysikalische forschungsberichte, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, June 26, 1941. 4 pp. See [Fis41].

Heisenberg:1941:MEH

[Hei41f] Werner Heisenberg. Über die Möglichkeit der Energieerzeugung mit Hilfe des Isotops 238. (German) [On the possibility of energy production with the help of isotope 238]. Kernphysikalische Forschungsberichte G-92, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, November 19, 1941. 15 pp.

Heisenberg:1941:DAU

[Hei41g] Werner Heisenberg. Zur Durchführung der Arbeiten an der Uranmaschine. (German) [On carrying out work on the uranium machine]. Report 34, Archives of the Max Planck Society, Berlin, Germany (AMPG). I. Abteilung, Berlin, Germany, November 27, 1941.

Heisenberg:1942:JEG

[Hei42a] Werner Heisenberg. 100 Jahre Energiegesetz. (German) [100 years of the energy law]. Deutsche Allgemeine Zeitung, 536(??):1–2, November 9, 1942.

Heisenberg:1942:BGH

[Hei42b] Werner Heisenberg. Bemerkungen zu dem geplanten halbtechnischen Versuch mit 1,5 to D2O und 3 to 38-Metall. (German) [Comments on the proposed pilot plant test with 1.5 tonnes D_{2}O and 3 tonnes 38-metal]. Kernphysikalische Forschungsberichte G-161, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, July 31, 1942. 8 pp.

Heisenberg:1942:AUG

[Hei42c] Werner Heisenberg. Die Arbeiten am Uranproblem. (German) [Working on the uranium problem]. Report 34, Archives of the Max Planck Society, Berlin, Germany (AMPG). I. Abteilung, Berlin, Germany, June 4, 1942. ???? pp.

Heisenberg:1942:ENW

[Hei42d] Werner Heisenberg. Die Einheit des naturwissenschaftlichen Weltbildes. (German) [The unity of the scientific world view]. Barth, Leipzig, Germany, 1942. 31 pp. LCCN ???? English translation in [Hei52j].

Heisenberg:1942:TGE

[Hei42e] Werner Heisenberg. Die theoretischen Grundlagen [für?] der Energiegewinnung aus der Uranspaltung. (German) [The theoretical basis of the energy from the fission of uranium]. Report G-323, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, February 26, 1942. 5 pp. Reprinted in [BDR89a, page 517]. Also preserved as Deutsches Museum Archiv, report FA-002-761.

Heisenberg:1943:TES

[Hei43a] Werner Heisenberg. 12. Theorie der explosionsartigen Schauer. (German) [12. Theory of explosion-like showers]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages 115–118. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:KGC

[Hei43b] Werner Heisenberg. 2. Die Kaskadentheorie. (German) [2. Cascade theory]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages 10–24. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:SDT

[Hei43c] Werner Heisenberg. 5. Schauer mit durchdringenden Teilchen. (German) [5. Showers with penetrating particles]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages 54–60. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:RZM

[Hei43d] Werner Heisenberg. 8. Der radioaktive Zerfall der Mesonen. (German) [8. The radioactive decay of mesons]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages 78–84. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:AGV

[Hei43e] Werner Heisenberg. Auswertung des Gottower Versuches. (German) [Evaluation of the Gottow experiment]. Report G-219, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, 1943. 12 pp.

Heisenberg:1943:BGWa

[Hei43f] Werner Heisenberg. B. O. Grönbloms wissenschaftliche Arbeiten. (German) [B. O. Grönblom’s scientific work]. Acta Academiae Aboenis. Math. et Phys., 14(4):1–18, ???? 1943.

Heisenberg:1943:BAS

[Hei43g] Werner Heisenberg. Besprechung: Arnold Sommerfeld, Vorlesungen über theoretische Physik. Bd. I. Mechanik. (German) [Book review: Arnold Sommerfeld, Lectures on Theoretical Physics. Volume 1. Mechanics]. Naturwissenschaften, 31(??):350–351, ???? 1943. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1943:BGWb

[Hei43h] Werner Heisenberg. Besprechung: Gregor Wentzel, Einführung in die Quantentheorie der Wellenfelder. (German) [Book review: Gregor Wentzel, Introduction to the quantum theory of wave fields]. Naturwissenschaften, 31(??):251–252, ???? 1943. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1943:BMP

[Hei43i] Werner Heisenberg. Besprechung: Max Planck: Wege zur physikalischen Erkenntnis. Bd. II. (German) [Book review: Max Planck, Paths to physical knowledge]. Leipziger Neueste Nachrichten, Ausg. A, ??(??):??, June 25, 1943.

Heisenberg:1943:BGTb

[Hei43j] Werner Heisenberg. Die “beobachtbaren Größen” in der Theorie der Elementarteilchen. II. (German) [“Observable variables” in the theory of elementary particles. II]. Zeitschrift für Physik, 120(11–12):673–702, November 1943. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01336936;   http://www.
springerlink.com/content/r744562625813k35/.

Heisenberg:1943:BMT

[Hei43k] Werner Heisenberg. Die Bewertung der ,,modernen theoretischen Physik“. (German) [The evaluation of “modern theoretical physics”]. Zeitschrift für die gesamte Naturwissenschaft, 9(??): 201–212, October 1943. According to [Ros98, pages 269–270] and [Wal89b, page 74], this article was written three years earlier, in 1940, but was delayed until Heisenberg had received Heinrich Himmler’s letter of protection. It is Heisenberg’s response to the “deutsche Physik” campaign to suppress Jews in science, and Jewish contributions to physics. See also [Cas09, page 338] for more about this article.

Heisenberg:1943:EAG

[Hei43l] Werner Heisenberg. Die Energiegewinnung aus der Atomkernspaltung. (German) [The energy production from nuclear fission]. Kernphysikalische Forschungsberichte G-217, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, May 6, 1943. 6 pp.

Heisenberg:1943:BGTa

[Hei43m] Werner Heisenberg. Die ,,beobachtbaren Größen“in der Theorie der Elementarteilchen. (German) [‘Observable quantities’ in the theory of elementary particles]. Zeitschrift für Physik, 120(7–10):513–538, July 1943. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01329800;   http:/
/www.springerlink.com/content/jm3516j05614784j/. English

translation in [Hei62k]. Reprinted in [BDR84a, pages 611–636].

Heisenberg:1943:PAG

[Hei43n] Werner Heisenberg. Die Physik der Atomkerne. (German) [Physics of the atomic nucleus]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1943. 180 pp. LCCN ????

Heisenberg:1943:EUJ

[Hei43o] Werner Heisenberg. Einleitung. 1. Übersicht über den jetzigen Stand unserer Kenntnisse von der kosmischen Strahlung. (German) [Introduction. 1. Present state of our knowledge of cosmic radiation]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages 1–10. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:KSV

[Hei43p] Werner Heisenberg, editor. Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1943. vi + 173 pp. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:MP

[Hei43q] Werner Heisenberg. [Max Planck]. Völkischer Beobachter, ??(??): ??, March 1943. Incompletely cited in [Cas09, page 339, 450].

Heisenberg:1943:VKS

[Hei43r] Werner Heisenberg. Vorträge über kosmische Strahlung: Arnold Sommerfeld zu seinem 75. Geburtstag am 5. Dezember 1943 gewidmet. (German) [Lectures on cosmic rays: dedicated to Arnold Sommerfeld on his 75th birthday on 5 December 1943]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1943. vi + 173 pp.

Heisenberg:1943:VGF

[Hei43s] Werner Heisenberg. Vorwort. (German) [Foreword]. In Kosmische Strahlung: Vorträge gehalten im Max Planck-Institut, Berlin-Dahlem. (German) [Cosmic radiation: Lectures held in the Max Planck Institute, Berlin-Dahlem] [Hei43p], pages v–vi. LCCN ???? Lectures in honor of Arnold Sommerfeld. With contributions by Erich Bagge, Fritz Bopp, Siegfried Flügge, Werner Heisenberg, A. Klemm, J. Meixner, G. Molière, H. Volz, Carl Friedrich von Weizsäcker, and Karl Wirtz. English translation in [Hei46a].

Heisenberg:1943:WGN

[Hei43t] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1943. 93 pp. LCCN ????

Heisenberg:1943:ADH

[Hei43u] Werner Heisenberg. Zum Andenken an David Hilbert. (German) [In memory of David Hilbert]. Physikalische Zeitschrift, 44(??): 277–278, ???? 1943. CODEN PHZTAO. ISSN 0369-982X.

Heisenberg:1944:BGT

[Hei44a] Werner Heisenberg. Die ,,beobachtbaren Größen“in der Theorie der Elementarteilchen. III. (German) [The “observable variables” in the theory of elementary particles. III]. Zeitschrift für Physik, 123(1–2):93–112, March 1944. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01375146;   http://www.
springerlink.com/content/t33g3p1272652730/. Part IV was

not published in a journal, but was finally printed in [BDR89a, pages 687–698].

Heisenberg:1944:MLS

[Hei44b] Werner Heisenberg. Max von Laue. Zu seinem 65. Geburtstag. (German) [Max von Laue. On his 65th birthday]. Deutsche Allgemeine Zeitung (Reichsausgabe), 273(??):2, October 6, 1944.

Heisenberg:1944:TAD

[Hei44c] Werner Heisenberg. Theoretische Auswertung der Dahlemer Grosßversuche. (German) [Theoretical evaluation of the Dahlem large scale experiments]. Report G-220, Uranverein, Kaiser Wilhelm Institut für Physik in Berlin-Dahlem, Berlin, Germany, September 1944. 7 pp.

Heisenberg:1944:WGN

[Hei44d] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1944. 95 pp. LCCN ????

Heisenberg:1945:PWG

[Hei45a] Werner Heisenberg. Das Plancksche Wirkungsquantum. (German) [Planck’s constant]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1945. 19 pp. LCCN ????

Heisenberg:1945:WGN

[Hei45b] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1945. 95 pp. LCCN ????

Heisenberg:1946:CRL

[Hei46a] Werner Heisenberg, editor. Cosmic Radiation. 15 Lectures. Dover, New York, NY, USA, 1946. 192 pp. LCCN ???? English translation by T. H. Johnson of [Hei43p], in commemoration of the 75th birthday of Arnold Sommerfeld.

Heisenberg:1946:MRQ

[Hei46b] Werner Heisenberg. Der mathematische Rahmen der Quantentheorie der Wellenfelder. (German) [The mathematical framework of the quantum theory of wave fields]. Zeitschrift für Naturforschung, 1:608–622, 1946. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1946:UQG

[Hei46c] Werner Heisenberg. Der unanschauliche Quantensprung. (German) [The nonintuitive quantum leap]. Neue Physikalische Blätter, 2(??):4–6, ???? 1946.

Heisenberg:1946:ATA

[Hei46d] Werner Heisenberg. Über die Arbeiten zur Technischen Ausnutzung der Atomkernenergie in Deutschland. (German) [On the work toward the technical use of atomic energy in Germany]. Naturwissenschaften, 33(11):325–329, ???? 1946. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//www.springerlink.com/content/j0m0u5j553402765. English

translation in [Hei47g].

Heisenberg:1946:WMV

[Hei46e] Werner Heisenberg. Wissenschaft als Mittel zur Verständigung unter den Völkern. (German) [Science as a way of understanding among peoples]. Deutsche Beiträge, 1(??):164–174, ???? 1946. English translation in [Hei52m].

Heisenberg:1946:ESG

[Hei46f] Werner Heisenberg. Zur Elektronentheorie der Supraleitung. (German) [On the electron theory of superconductivity]. Neue Physikalische Blätter, 2(??):220, ???? 1946.

Heisenberg:1947:APO

[Hei47a] Werner Heisenberg. Die aktive und die passive Opposition im Dritten Reich. (German) [Active and passive opposition in the Third Reich]. Unpublished paper in Heisenberg Archives in Munich, cited in [Cas09, page 220, 437]., November 12, 1947.

Heisenberg:1947:PAG

[Hei47b] Werner Heisenberg. Die Physik der Atomkerne. (German) [Physics of the atomic nucleus]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1947. vi + 180 pp. LCCN ???? Elaborated by W. Westphal and Dr. Jorges. The third edition is reproduced in the Collected Works.

Heisenberg:1947:UEK

[Hei47c] Werner Heisenberg. Die Übertragung elektromagnetischer Kräfte im Supraleiter. (German) [The transfer of electromagnetic forces in the superconductor]. Nachrichten der Akademie der Wissenschaften in Göttingen, Math.-Phys. Klasse, ??(??):23–26, ???? 1947.

Heisenberg:1947:MP

[Hei47d] Werner Heisenberg. Max Planck. Göttinger Universitäts-Zeitung, 2(10):6, ???? 1947.

Heisenberg:1947:MPN

[Hei47e] Werner Heisenberg. Max Planck [Nachruf]. (German) [Max Planck [obituary]]. Göttinger Universitäts-Zeitung, 2(21):1–2, ???? 1947.

Heisenberg:1947:QWV

[Hei47f] Werner Heisenberg. Quantentheorie der Wellenfelder: Vorlesung von Prof. Heisenberg, Winter 1946/47. (German) [Quantum theory of wave fields: lecture of Prof. Heisenberg, Winter 1946/47]. Max-Planck-Institut für Physik, Göttingen, West Germany, 1947. ???? pp. LCCN ???? Elaborated by K. Wildermuth.

Heisenberg:1947:RGTa

[Hei47g] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Nature, 160(4059):211–215, August 16, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v160/

n4059/pdf/160211a0.pdf. Abridged translation of German original [Hei46d]. See also [Hei53m, Hei53n], and response [Gou47c].

Heisenberg:1947:RGTb

[Hei47h] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Chemical and Engineering News, 1947 (25):2620–2621, September 15, 1947. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605X (electronic). URL http://pubs.

acs.org/doi/pdf/10.1021/cen-v025n037.p2620.

Heisenberg:1947:WGN

[Hei47i] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, Germany, 1947. 98 pp. LCCN ????

Heisenberg:1947:TSGa

[Hei47j] Werner Heisenberg. Zur Theorie der Supraleitung. (German) [On the theory of superconductivity]. Zeitschrift für Naturforschung, 2a(4):185–201, ???? 1947. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1947:TSGb

[Hei47k] Werner Heisenberg. Zur Theorie der Supraleitung. (German) [On the theory of superconductivity]. Forsch. Fortschr., 21–23(??): 243–244, ???? 1947.

Heisenberg:1947:TSGc

[Hei47l] Werner Heisenberg. Zur Theorie der Supraleitung. (German) [On the theory of superconductivity]. Angewandte Chemie, 59(2):59, February 1947. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1948:ASD

[Hei48a] Werner Heisenberg. Arnold Sommerfeld zum 5. Dezember 1948. (German) [Arnold Sommerfeld to 5th December 1948]. Zeitschrift für Naturforschung, 3a(??):429, ???? 1948. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1948:BTG

[Hei48b] Werner Heisenberg. Bemerkungen zum Turbulenzproblem. (German) [Remarks on the turbulence problem]. Zeitschrift für Naturforschung, 3a(??):434–437, ???? 1948. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1948:EVS

[Hei48c] Werner Heisenberg. Das elektrodynamische Verhalten der Supraleiter. (German) [The electrodynamic behavior of superconductors]. Zeitschrift für Naturforschung, 3a(??):65–75, ???? 1948. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1948:BGA

[Hei48d] Werner Heisenberg. Der Begriff ,,Abgeschlossene Theorie“in der modernen Naturwissenschaft. (German) [The notion of a “closed theory” in modern science]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):331–336, November 1948. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). English translation in [Hei74i].

Heisenberg:1948:BAT

[Hei48e] Werner Heisenberg. Der Begriff ,,Abgeschlossene Theorie“in der modernen Naturwissenschaft. (German) [The notion of a “closed theory” in modern science]. Dialectica: International Review of Philosophy of Knowledge, 2(3–4):331–336, November 1948. CO-DEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). English translation in [Hei74i].

Heisenberg:1948:SN

[Hei48f] Werner Heisenberg. Die Sorge um die Naturwissenschaft. Göttinger Universitäts-Zeitung, 3(??):7, January 16, 1948.

Heisenberg:1948:GND

[Hei48g] Werner Heisenberg. Es gab niemals eine deutsche Atombombe. (German) [There never was a German atomic bomb]. Schwäbische Donauzeitung, ??(??):3, August 2, 1948.

Heisenberg:1948:LBV

[Hei48h] Werner Heisenberg. [Letter to B. L. van der Waerden]. Cassidy [Cas09, page 314] writes of this letter “Heisenberg himself seems to suggest another purpose for his visit [to Niels Bohr in Copenhagen] … Heisenberg implied that he was attempting to stave off an Allied crash program on nuclear fission research.”, April 28, 1948.

Heisenberg:1948:NSI

[Hei48i] Werner Heisenberg. Nazis spumed idea of an atomic bomb (interview). New York Times, ??(??):10, December 28, 1948. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Heisenberg:1948:TSI

[Hei48j] Werner Heisenberg. On the theory of statistical and isotropic turbulence. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 195:402–406, 1948. CODEN PRLAAZ. ISSN 0080-4630.

Heisenberg:1948:TBP

[Hei48k] Werner Heisenberg. Thermodynamische Betrachtungen zum Problem der Supraleitung. (German) [Thermodynamical considerations in the problem of superconductivity]. Annalen der Physik (1900), 438(1):289–296, 1948. ISSN 1521-3889.

Heisenberg:1948:WWW

[Hei48l] Werner Heisenberg. Wer weiß, was wichtig wird. Die Notwendigkeit wissenschaftlicher Forschung. (German) [Who knows what is important? The need for scientific research]. Die Welt, 143(??):3, December 9, 1948.

Heisenberg:1948:WHA

[Hei48m] Werner Heisenberg. Werner Heisenberg (autobiographical note). In ????, editor, Der Geist weht wo er will. Kalender für das Jahr 1949. Zero-Verlag, Ulm, West Germany, 1948. 10. Woche [10th week].

Heisenberg:1948:STT

[Hei48n] Werner Heisenberg. Zur statistischen Theorie der Turbulenz. (German) [On the statistical theory of turbulence]. Zeitschrift für Physik, 124(7–12):628–657, July 1948. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01668899;   http:/
/www.springerlink.com/content/p225013t42156018/. English

translation in [Hei58q].

Heisenberg:1949:AEJ

[Hei49a] Werner Heisenberg. Albert Einstein 70 Jahre alt. (German) [Albert Einstein 70 years old]. Welt am Sonntag, 11(??):5, March 13, 1949.

Heisenberg:1949:ALM

[Hei49b] Werner Heisenberg. Die Auswirkungen des Lebenswerkes Max Plancks. (German) [The impact of the life work of Max Planck]. Angewandte Chemie, 61(4):115–117, 1949. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1949:EFG

[Hei49c] Werner Heisenberg. Die Einheit der Forschung. (German) [The unity of research]. Deutsche Zeitung und Wirtschaftszeitung, 103 (??):4, December 24, 1949. Reprinted in [Hei53f].

Heisenberg:1949:EMVb

[Hei49d] Werner Heisenberg. Die Erzeugung von Mesonen in Vielfachprozessen. (German) [The production of mesons in multiple processes]. Il Nuovo Cimento (9), 6 (supplement)(??):493–497, 503– 508, ???? 1949. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Heisenberg:1949:GGA

[Hei49e] Werner Heisenberg. Die gegenwartigen Grundprobleme der Atomphysik. (German) [Fundamental problems of present-day atomic physics]. In Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science] [Hei49l], pages 89–101. LCCN ????

Heisenberg:1949:GAR

[Hei49f] Werner Heisenberg. German atom research: Appraisal of progress said to await release of reports. New York Times, ??(??):8E– 8E, January 30, 1949. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/105774675/. Section 4.

Heisenberg:1949:MFF

[Hei49g] Werner Heisenberg. Mehr Fürsorge für die Forschung. Eine Denkschrift des Deutschen Forschungsrates. (German) [More care for research. A memorandum by the German Research Council]. Deutsche Zeitung und Wirtschaftszeitung, 75(??):5–10, September 17, 1949. Reprinted in [Hei53d].

Heisenberg:1949:NHB

[Hei49h] Werner Heisenberg. Naturwissenschaft und humanistische Bildung. (German) [Science and education in the humanities]. In H. Scharold, editor, 100 Jahre Maximilians-Gymnasium, em Beitrag zur Geschichte des Gymnasiums in Bayern. (German) [Centennial of the Maximilians-Gymnasium, a contribution to the history of Bavarian high schools], pages 153–164. ????, Munich, West Germany, 1949. LCCN ????

Heisenberg:1949:PMS

[Hei49i] Werner Heisenberg. Production of meson showers. Nature, 164 (4158):65–66, July 9, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v164/n4158/pdf/164065c0.pdf.

Heisenberg:1949:TL

[Hei49j] Werner Heisenberg. Two Lectures. Cambridge University Press, Cambridge, UK, 1949. 52 pp. LCCN ????

Heisenberg:1949:EMVa

[Hei49k] Werner Heisenberg. Über die Entstehung von Mesonen in Vielfachprozessen. (German) [On the origin of mesons in multiple processes]. Zeitschrift für Physik, 126(6):569–582, July 1949. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01330108;
http://www.springerlink.com/content/u8614153126336l7/.

Heisenberg:1949:WGN

[Hei49l] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Leipzig, East Germany, 1949. 112 pp. LCCN ????

Heisenberg:1950:BBL

[Hei50a] Werner Heisenberg. Besprechung: Theorie und Losungsmethoden des Mehrteilchen problems der Wellenmechanik, von P. Gombás. (German) [Book review: Theory and solution methods of manyparticle problems of wave mechanics]. Umschau, 50(??):550, ???? 1950.

Heisenberg:1950:BWK

[Hei50b] Werner Heisenberg. Besprechung: Wilhelm Kamlah: Der Mensch in der Profanität. (German) [Book review: Wilhelm Kamlah: Man in profanity]. Naturwissenschaften, 37(??):311, ???? 1950. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1950:DFI

[Hei50c] Werner Heisenberg. Der deutsche Forschungsrat (Interview). (German) [The German Research Council (interview)]. Physikalische Blätter, 6(4):170–172, April 1950. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19500060405/
abstract.

Heisenberg:1950:DWG

[Hei50d] Werner Heisenberg. Die Not der deutschen Wissenschaft. (German) [The plight of German science]. Die Tat, 230(??):30, August 15, 1950.

Heisenberg:1950:AUI

[Hei50e] Werner Heisenberg. Für die Atombombe unzuständig (Interview). (German) [Jurisdiction for the atomic bomb [????] (interview)]. Die Welt, 28(??):2, February 2, 1950.

Heisenberg:1950:PPQ

[Hei50f] Werner Heisenberg. The physical principles of the quantum theory. Dover, New York, NY, USA, 1950. ISBN 0-486-60113-7. 183 pp. LCCN QC174.1 .H4 1950. Translated into English by Carl Eckart and Frank C. Hoyt.

Heisenberg:1950:SPD

[Hei50g] Werner Heisenberg. Schlußwort des Präsidenten des Deutschen Forschungsrates. (German) [Closing remarks by the President of the German Research Council]. In ????, editor, Vom Geist der Deutschen Wissenschaft. Ansprachen und Vortrage. (German) [The Spirit of German Science. Speeches and Lectures], pages 45–47. Franz Steiner Verlag, Wiesbaden, West Germany, 1950. Reprinted in [Hei53q].

Heisenberg:1950:SZR

[Hei50h] Werner Heisenberg. Stationäre Zustände in der relativistischen Quantentheorie der Wellenfelder. (German) [Stationary states in the relativistic quantum theory of wave fields]. Zeitschrift für Naturforschung, 5a:367–373, 1950. CODEN ZNTFA2. ISSN 0372-

Heisenberg:1950:YTN

[Hei50i] Werner Heisenberg. The Yukawa theory of nuclear forces in the light of present quantum theory of wave fields. Progress of Theoretical Physics, 5(4):523–525, July/August 1950. CODEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Heisenberg:1950:QEG

[Hei50j] Werner Heisenberg. Zur Quantentheorie der Elementarteilchen. (German) [On the quantum theory of elementary particles]. Zeitschrift für Naturforschung, 5a:251–259, 1950. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1951:JQG

[Hei51a] Werner Heisenberg. 50 Jahre Quantentheorie. (German) [50 years of quantum theory]. Naturwissenschaften, 38(3):49–55, ???? 1951. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/k1t4655274v37550.

Heisenberg:1951:AS

[Hei51b] Werner Heisenberg. Arnold Sommerfeld. Naturwissenschaften, 38 (??):337–338, ???? 1951. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1951:ASN

[Hei51c] Werner Heisenberg. Arnold Sommerfeld [Nachruf]. (German) [Arnold Sommerfeld [obituary]]. Naturwissenschaften, 38(15):337– 338, ???? 1951. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/r02343hv28210677.

Heisenberg:1951:RWI

[Hei51d] Werner Heisenberg. Die Rolle der Wissenschaft im öffentlichen Leben. (German) [The role of science in public life]. Deutsche Universitätszeitung, 6(12):15–16, June 29, 1951.

Heisenberg:1951:DFN

[Hei51e] Werner Heisenberg. Hat Deutschlands Forschung noch eine Chance? (Interview with A. Pullmann). (German) [Does Germany’s research have a chance? (Interview with A. Pullman)]. Göttinger Tageblatt, 265(??):??, November 10, 1951.

Heisenberg:1951:KSA

[Hei51f] Werner Heisenberg. Kosmische Strahlungen und Atomphysik. (German) [Cosmic rays and atomic physics]. In ????, editor, Jahresbericht 1951 der Max-Planck-Gesellschaft zur Forderung der Wissenschaften e.V.. (German) [Annual Report for 1951 of the Max Planck Society for the Needs of Sciences e.V.], pages 229– 236. Max-Planck-Gesellschaft, Göttingen, West Germany, 1951.

Heisenberg:1951:STF

[Hei51g] Werner Heisenberg. On stability and turbulence of fluid flows. Techical Memo 1291, National Commission for Aeronautics, Washington, DC, USA, 1951. 60 pp. English translation of Heisenberg’s doctoral dissertation [Hei24e].

Heisenberg:1951:MFT

[Hei51h] Werner Heisenberg. On the mathematical frame of the theory of elementary particles. Communications on Pure and Applied Mathematics (New York), 4(1):15–22, 1951. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Heisenberg:1951:PZT

[Hei51i] Werner Heisenberg. Paradoxien des Zeitbegriffs in der Theorie der Elementarteilchen. (German) [Paradoxes of the time concept in the theory of elementary particles]. In Festschrift zur Feier des zweihundertjährigen Bestehens der Akademie der Wissenschaften in Göttingen. I. Math.-Phys. Kl. (German) [Festschrift in honor of the bicentennial anniversary of the Academy of Sciences in Goettingen. I. Math Phys. kl], pages 50–64. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1951.

Heisenberg:1951:TAG

[Hei51j] Werner Heisenberg. Theorie des Atomkerns. (German) [Theory of the atomic nucleus]. Max-Planck-Institut für Physik, Göttingen, West Germany, 1951. 162 pp. LCCN ????

Heisenberg:1951:TAV

[Hei51k] Werner Heisenberg. Theorie des Atomkerns. Vorlesung von W. Heisenberg im Sommer 1950. (German) [Theory of the atomic nucleus. Lecturees of W. Heisenberg, Summer 1950]. Max-Planck-Institut für Physik, Göttingen, West Germany, 1951. ???? pp. LCCN ???? Elaborated by W. Macke.

Heisenberg:1951:VUT

[Hei51l] Werner Heisenberg. Vergleichende Übersicht über Theorien der π-Mesonen Produktion in der Ultrastrahlung. (German) [Comparative review of theories of π-meson production in cosmic radiation]. In ????, editor, Diskussionskonferenz über Probleme der Kernphysik und Ultrastrahlung, 1–3 Juli 1951. (German) [Talk conference on problems of nuclear physics and cosmic rays, 1– 3 July 1951], pages E1–E6. Universität Heidelberg, Heidelberg, West Germany, 1951. LCCN ????

Heisenberg:1951:FKQ

[Hei51m] Werner Heisenberg. Zur Frage der Kausalität in der Quantentheorie der Elementarteilchen. (German) [On the question of causality in the quantum theory of elementary particles]. Zeitschrift für Naturforschung, 6a:281–284, 1951. CODEN ZNTFA2. ISSN 0372-9516.

Heisenberg:1952:AKG

[Hei52a] Werner Heisenberg. Atomphysik und Kausalgesetz. (German) [Atomic physics and causality]. Merkur, 6(??):701–711, ???? 1952. English translation in [Hei58c].

Heisenberg:1952:BTV

[Hei52b] Werner Heisenberg. Bemerkungen zur Theorie der Vielfacherzeugung von Mesonen. (German) [Remarks on the theory of multiple production of mesons]. Naturwissenschaften, 39(3):69, ???? 1952. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/
t83u7v30822q1515.

Heisenberg:1952:EAG

[Hei52c] Werner Heisenberg. Europas Atomunion. (German) [Europe’s atomic union]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 7(??):989–990, ???? 1952. CODEN UNIVA8. ISSN 0041-9079.

Heisenberg:1952:FPP

[Hei52d] Werner Heisenberg. Fundamental problems of present-day atomic physics. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 95–108. LCCN QC71 .H3513 1952.

Heisenberg:1952:INP

[Hei52e] Werner Heisenberg. Ideas of the natural philosophy of ancient times in modern physics. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 53–59. LCCN QC71 .H3513 1952. English translation of [Hei37e].

Heisenberg:1952:KFB

[Hei52f] Werner Heisenberg. Kolloquium in Freiburg i. Br. (German) [Seminar in Freiburg i. Br.]. Angewandte Chemie, 64(8):225–226, 1952. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1952:MSG

[Hei52g] Werner Heisenberg. Mesonenerzeugung als Stoßwellenproblem. (German) [Meson generation as a shock wave problem]. Zeitschrift für Physik, 133(1–2):65–79, September 1952. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://

link.springer.com/article/10.1007/BF01948683;   http:/
/www.springerlink.com/content/q1434v1k36w64418/.

Heisenberg:1952:MPM

[Hei52h] Werner Heisenberg. Multiple production of mesons. In O. Koefod-Hansen et al., editors, Report on the International Physics Conference, June 3–17, 1952, pages 39–41. Institute of Theoretical Physics, Copenhagen, DK, 1952. LCCN ????

Heisenberg:1952:SLF

[Hei52i] Werner Heisenberg. On the stability of laminar flow. In Proceedings of the International Congress of Mathematicians, Cambridge, Massachusetts, 1950, vol. 2, pages 292–296. American Mathematical Society, Providence, RI, USA, 1952.

Heisenberg:1952:USO

[Hei52j] Werner Heisenberg. On the unity of the scientific outlook on nature. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 77–94. LCCN ???? English translation of [Hei42d].

Heisenberg:1952:PPN

[Hei52k] Werner Heisenberg. Philosophic problems of nuclear science: Eight lectures. Faber and Faber, London, UK, 1952. 126 pp. LCCN QC71 .H3513 1952.

Heisenberg:1952:QPM

[Hei52l] Werner Heisenberg. Questions of principle in modern physics. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 41–52. LCCN QC71 .H3513 1952. English translation of [Hei36f].

Heisenberg:1952:SMI

[Hei52m] Werner Heisenberg. Science as means of international understanding. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 109–120. LCCN QC71 .H3513 1952. English translation of [Hei46e].

Heisenberg:1952:SHM

[Hei52n] Werner Heisenberg. Summary of heavy meson data. In O. Koefod-Hansen et al., editors, Report on the International Physics Conference, June 3–17, 1952, pages 35–37. Institute of Theoretical Physics, Copenhagen, DK, 1952. LCCN ????

Heisenberg:1952:TGN

[Hei52o] Werner Heisenberg. The teaching of Goethe and Newton on colour in the light of modern physics. In Philosophic problems of nuclear science: Eight lectures [Hei52k], pages 60–76. LCCN QC71 .H3513 1952. English translation of [Hei41d].

Heisenberg:1952:TNV

[Hei52p] Werner Heisenberg. Theorie der Neutronen, Vorlesung gehalten im Winter 1950/51. (German) [Theory of neutrons, lectures held in Winter 1950/51]. Max-Planck-Institut für Physik, Göttingen, West Germany, 1952. 123 pp.

Heisenberg:1952:TES

[Hei52q] Werner Heisenberg. Theory of explosion showers. Technical Translation TT-320, National Research Council of Canada, Ottawa, ON, Canada, 1952. English translation of [Hei53s].

Heisenberg:1952:GGR

[Hei52r] Werner Heisenberg. Zum Geleit. (German) [Preface]. In Das Atom. (German) [The Atom], page 5. “Musterschmidt” Wiss. Verlag, Göttingen, West Germany, 1952. LCCN ????

Heisenberg:1952:ZEP

[Hei52s] Werner Heisenberg. Zur Zerstrahlung von Elektron und Positron. (German) [On the annihilation of the electron and positron]. Physikalische Blätter, 8(4):191, April 1952. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19520080410/
abstract.

Heisenberg:1953:AGM

[Hei53a] Werner Heisenberg. Aktennotiz. (German) [Memo]. In Eickemeyer [Eic53], pages 59–61. LCCN ???? On the discussion of 19 January 1951 concerning the desperate situation of university institutes.

Heisenberg:1953:MSJ

[Hei53b] Werner Heisenberg. Aus meinen späten Jahren, von Albert Einstein. (Übersetzung von Dr. Hildegard Blomeyer, nach der amerikanischen Ausgabe ,,Out of my later years“, Philosophical Library, New York 1950). Deutsche Verlagsanstalt, Stuttgart. 1952. 280 S., Leinen DM 13.80. (German) [Out of my later years, by Albert Einstein. (Translation by Dr. Hildegard Blomeyer, according to the American edition “Out of my later years”, Philosophical Library, New York, 1950). German Publishing house, Stuttgart. 1952, 280 pp, Linen DM 13.80]. Angewandte Chemie, 65(1):43, 1953. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Heisenberg:1953:NHP

[Hei53c] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The image of nature of contemporary physics]. In Jahresbericht 1953 der Max-Planck-Gesellschaft zur Forderung der Wissenschaften e.V.. (German) [Annual Report for 1953 of the Max Planck Society for the Needs of Sciences e.V.], pages 32–54. Max-Planck-Gesellschaft, Göttingen, West Germany, 1953. English translation in [Hei58o].

Heisenberg:1953:DDF

[Hei53d] Werner Heisenberg. Denkschrift des Deutschen Forschungsrates über die Betreuung der wissenschaftlichen Forschung im Rahmen der Bundesregierung. In Eickemeyer [Eic53], pages 117–121. LCCN ???? Reprint of [Hei49g].

Heisenberg:1953:BZP

[Hei53e] Werner Heisenberg. Die Beziehungen zwischen Physik und Chemie in den letzten 75 Jahren. (German) [Relations between physics and chemistry in the last 75 years]. Naturwiss. Rundschau, 6(??):1–7, ???? 1953.

Heisenberg:1953:EFG

[Hei53f] Werner Heisenberg. Die Einheit der Forschung. (German) [The unity of research]. In Eickemeyer [Eic53], pages 213–215. LCCN ???? Reprint of [Hei49c].

Heisenberg:1953:MAD

[Hei53g] Werner Heisenberg. Die Möglichkeiten der Atomtechnik in Deutschland. (German) [The possibilities of nuclear technology in Germany]. Ubersee-Club, Hamburg, West Germany, 1953. 12 pp. LCCN ????

Heisenberg:1953:RWI

[Hei53h] Werner Heisenberg. Die Rolle der Wissenschaft im öffentlichen Leben. (German) [The role of science in public life]. In Eickemeyer [Eic53], pages 219–221. LCCN ???? 117–121.

Heisenberg:1953:DSP

[Hei53i] Werner Heisenberg. Discussion on the size of the proton synchrotron. In The Theory and Design of an Alternating Gradient Proton Synchrotron, pages 179–180. Institute of Physics, Geneva, Switzerland, 1953. LCCN ????

Heisenberg:1953:DHQ

[Hei53j] Werner Heisenberg. Doubts and hopes in quantum-electrodynamics. Physica, 19:897–908, 1953. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/1953Phy....19..897H.

Heisenberg:1953:FAA

[Hei53k] Werner Heisenberg. Friedliche Atomtechnik. Die Anwendung der Atomenergie in Deutschland. (German) [Peaceful nuclear technology. The application of atomic energy in Germany]. Deutsche Universitätszeitung, 8(6):10–13, March 23, 1953.

Heisenberg:1953:KSV

[Hei53l] Werner Heisenberg, editor. Kosmische Strahlung: Vorträge gehalten am Max-Planck-Institut für Physik in Göttingen. (German)

[Cosmic rays: Lectures held at the Max-Planck-Institut for Physics in Göttingen]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1953. viii + 620 pp. LCCN ???? With contributions by L. Biermann, P. Budini, J. Buschmann, M. Deutschmann, E. Freese, K. Gottstein, R. Hagedorn, Werner Heisenberg, Fritz G. Houtermans, H. Jahn, G. Lüders, R. Lüst, W. Macke, H. M. Mayer, P. Meyer, G. Molière, R. Oehme, K. Ott, W. Paul, F. Sauter, A. Schlüter, K. Symanzik, M. Teucher, Carl Friedrich von Weizsäcker, Karl Wirtz, and B. Zumino.

Heisenberg:1953:NPa

[Hei53m] Werner Heisenberg. Nuclear Physics. Methuen and Co. Ltd., London, UK, 1953. ???? pp. LCCN ???? English translation of [HW49].

Heisenberg:1953:NPb

[Hei53n] Werner Heisenberg. Nuclear Physics. Philosophical Library, New York, NY, USA, 1953. ISBN 0-8065-3033-2. ix + 225 pp. LCCN

QC173 .H3854. URL http://alsos.wlu.edu/information.

aspx?id=95. English translation of [HW49]. See also [Hei47g].

Heisenberg:1953:PVK

[Hei53o] Werner Heisenberg. Platons Vorstellungen von den kleinsten Bausteinen der Materie und die Elementarteilchen der modernen Physik. (German) [Plato’s ideas about the smallest building blocks of matter and the fundamental particles of modern physics]. In ????, editor, Im Umkreis der Kunst. Eine Festschift für Emil Pretorius. (German) [Within the art. A Festschift for Emil Pretorius], pages 137–140. Insel-Verlag, Wiesbaden, West Germany, 1953. LCCN ????

Heisenberg:1953:RTN

[Hei53p] Werner Heisenberg. Remarques sur la théorie neutrinienne de la lumière. (French) [Remarks on the neutrino theory of light]. In ????, editor, Louis de Broglie, Physicien et Penseur. (French) [Louis de Broglie, Physicist and Thinker], pages 283–286. Albin Michel, Paris, France, 1953. LCCN ????

Heisenberg:1953:SPD

[Hei53q] Werner Heisenberg. Schlußwort des Präsidenten des Deutschen Forschungsrates, Professor Dr. W. Heisenberg auf der Kundgebung der deutschen Wissenschaft am 4. August 1950 im Plenarsaal des Bundestages in Bonn. (German) [Closing words of the President of German Research Council, Professor W. Heisenberg at the rally of German Science on 4 August 1950 in the Chamber of Bundestag in Bonn]. In Eickemeyer [Eic53], pages 217–219. LCCN ???? Reprint of [Hei50g].

Heisenberg:1953:SBE

[Hei53r] Werner Heisenberg. The significance of beauty in the exact sciences. Report at the Bavarian Fine Arts Academy., November 17, 1953.

Heisenberg:1953:TEG

[Hei53s] Werner Heisenberg. Theorie der Explosionsschauer. (German) [Theory of explosion showers]. In Kosmische Strahlung: Vorträge gehalten am Max-Planck-Institut für Physik in Göttingen. (German) [Cosmic rays: Lectures held at the Max-Planck-Institut for Physics in Göttingen] [Hei53l], pages 148–164. LCCN ???? English translation in [Hei52q].

Heisenberg:1953:BBH

[Hei53t] Werner Heisenberg. Übersicht über den heutigen Stand der Kenntnisse von der kosmischen Strahlung. (German) [Overview of the current knowledge of cosmic rays]. In Kosmische Strahlung: Vorträge gehalten am Max-Planck-Institut für Physik in Göttingen. (German) [Cosmic rays: Lectures held at the Max-Planck-Institut for Physics in Göttingen] [Hei53l], pages 1–8. LCCN ???? With contributions by L. Biermann, P. Budini, J. Buschmann, M. Deutschmann, E. Freese, K. Gottstein, R. Hagedorn, Werner Heisenberg, Fritz G. Houtermans, H. Jahn, G. Lüders, R. Lüst, W. Macke, H. M. Mayer, P. Meyer, G. Molière, R. Oehme, K. Ott, W. Paul, F. Sauter, A. Schlüter, K. Symanzik, M. Teucher, Carl Friedrich von Weizsäcker, Karl Wirtz, and B. Zumino.

Heisenberg:1953:UHS

[Hei53u] Werner Heisenberg. Übersicht über den heutigen Stand der Kenntnisse von der kosmischen Strahlung. (German) [Review of the present knowledge of cosmic radiation]. In Kosmische Strahlung: Vorträge gehalten am Max-Planck-Institut für Physik in Göttingen. (German) [Cosmic rays: Lectures held at the Max-Planck-Institut for Physics in Göttingen] [Hei53l], pages 1–8. LCCN ???? With contributions by L. Biermann, P. Budini, J. Buschmann, M. Deutschmann, E. Freese, K. Gottstein, R.

Hagedorn, Werner Heisenberg, Fritz G. Houtermans, H. Jahn, G. Lüders, R. Lüst, W. Macke, H. M. Mayer, P. Meyer, G. Molière, R. Oehme, K. Ott, W. Paul, F. Sauter, A. Schlüter, K. Symanzik, M. Teucher, Carl Friedrich von Weizsäcker, Karl Wirtz, and B. Zumino.

Heisenberg:1953:VGFa

[Hei53v] Werner Heisenberg. Vorwort. (German) [Foreword]. In Eickemeyer [Eic53], pages 9–10. LCCN ???? 117–121.

Heisenberg:1953:VGFb

[Hei53w] Werner Heisenberg. Vorwort. (German) [Foreword]. In A. Hildesheimer, editor, Die Welt der ungewohnten Dimensionen. (German) [The world of unusual dimensions], page ?? A. W. Sijthoff, Leiden, The Netherlands, 1953.

Heisenberg:1953:VAG

[Hei53x] Werner Heisenberg. Vorwort zur 2. Auflage. (German) [Foreword to the second edition]. In Kosmische Strahlung: Vorträge gehalten am Max-Planck-Institut für Physik in Göttingen. (German) [Cosmic rays: Lectures held at the Max-Planck-Institut for Physics in Göttingen] [Hei53l], page iv. LCCN ???? With contributions by L. Biermann, P. Budini, J. Buschmann, M. Deutschmann, E. Freese, K. Gottstein, R. Hagedorn, Werner Heisenberg, Fritz G. Houtermans, H. Jahn, G. Lüders, R. Lüst, W. Macke, H. M. Mayer, P. Meyer, G. Molière, R. Oehme, K. Ott, W. Paul, F. Sauter, A. Schlüter, K. Symanzik, M. Teucher, Carl Friedrich von Weizsäcker, Karl Wirtz, and B. Zumino.

Heisenberg:1953:QNG

[Hei53y] Werner Heisenberg. Zur Quantisierung nichtlinearer Gleichungen. (German) [On the quantization of nonlinear equations]. Nachr. Akad. Wiss. Göttingen. Math.-Phys. Kl. Math.-Phys.-Chem. Abt., 153:111–127, 1953.

Heiberg:1954:BBP

[Hei54a] Edvard Heiberg. Books: Nuclear Physics, by Werner Heisenberg. Bulletin of the Atomic Scientists, 10(6):233, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heisenberg:1954:EOK

[Hei54b] Werner Heisenberg. Die Europäische Organisation für kernphysikalische Forschung. (German) [The European Organization for Nuclear Physics Research]. Mitteilungen aus der Max-Planck-Gesellschaft, 3(??):137–140, ???? 1954.

Heisenberg:1954:EMG

[Hei54c] Werner Heisenberg. Elementarteile der Materie. (German) [Elementary particles of matter]. In ????, editor, Vom Atom zum Weltsystem. Eine Vortragsreihe über die Ergebnisse der Atomkernforschung. (German) [From the atom to the world system. A lecture series on the results of nuclear research], pages 45–58. Alfred Kröner, Stuttgart, West Germany, 1954. LCCN ????

Heisenberg:1954:NF

[Hei54d] Werner Heisenberg. Nuclear forces. American Journal of Physics, 22(6):415, June 1954. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://labs.adsabs.harvard.
edu/ui/abs/1954AmJPh..22..415H.

Heisenberg:1954:OHG

[Hei54e] Werner Heisenberg. Otto Hahn zum 75. Geburtstag. (German) [Otto Hahn’s 75 birthday]. Frankfurter Allgemeine Zeitung, 55 (??):3, March 6, 1954.

Heisenberg:1954:QNW

[Hei54f] Werner Heisenberg. Zur Quantentheorie nichtrenormierbarer Wellengleichungen. (German) [On the quantum theory of nonrenormalizable wave equations]. Zeitschrift für Naturforschung, 9a:292–303, 1954. CODEN ZNTFA2. ISSN 0372-9516. URL http:

//labs.adsabs.harvard.edu/ui/abs/1954ZNatA...9..292H.

Heisenberg:1955:AEW

[Hei55a] Werner Heisenberg. Albert Einsteins wissenschaftliches Werk. (German) [Albert Einstein’s scientific work]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 10(??):897–902, ???? 1955. CODEN UNIVA8. ISSN 0041-9079.

Heisenberg:1955:NHP

[Hei55b] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie Physik. Rowohlt, Reinbek bei Hamburg, Germany, 1955. 149 pp. LCCN ????

Heisenberg:1955:GST

[Hei55c] Werner Heisenberg. Der gegenwärtige Stand der Theorie der Elementarteilchen. (German) [The current state of the theory of elementary particles]. Naturwissenschaften, 42(24):637–641, ???? 1955. CODEN NATWAY. ISSN 0028-1042 (print), 1432-

1904 (electronic). URL http://labs.adsabs.harvard.edu/
ui/abs/1955NW.....42..637H;   http://www.springerlink.
com/content/l3701963116383x3.

Heisenberg:1955:DIQ

[Hei55d] Werner Heisenberg. The development of the interpretation of the quantum theory. In Pauli et al. [PRW55], pages 12–29. LCCN

QC71 .P3 1955. URL http://labs.adsabs.harvard.edu/ui/

abs/1955nbdp.book...12H. English translation of [Hei56h].

Heisenberg:1955:PMV

[Hei55e] Werner Heisenberg. The production of mesons in very high energy collisions. Il Nuovo Cimento (10), 2(Supplement 1):96–103, January 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-

6121 (electronic). URL http://link.springer.com/article/

10.1007/BF02746079. Reproduced from the magnetophone-tape after the lecture of the author.

Heisenberg:1955:QNL

[Hei55f] Werner Heisenberg. Quantization of non-linear wave equations. In E. H. Bellamy and R. G. Moorhouse, editors, Proceedings of the 1954 Glasgow Conference on Nuclear and Meson Physics, pages 293–295. Pergamon, New York, NY, USA, 1955. LCCN ????

Heisenberg:1955:TEP

[Hei55g] Werner Heisenberg. Theory of elementary particles. Science, 122 (3182):1215–1220, December 23, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://labs.

adsabs.harvard.edu/ui/abs/1955Sci...122.1215H;   http:
//www.sciencemag.org/content/122/3182/1215.full.pdf.

Heisenberg:1956:ABG

[Hei56a] Werner Heisenberg. An ‘Die Atomwirtschaft’ (Brief). (German) [On ‘The nuclear industry’ (letter)]. Die Atomwirtschaft, 1(??):3, ???? 1956.

Heisenberg:1956:AED

[Hei56b] Werner Heisenberg. Atomic energy development in Germany. post war scientific developments. Atomics, 7(??):136–138, 146, ????

Heisenberg:1956:FPH

[Hei56c] Werner Heisenberg. Aus dem Festvortrag von Professor Heisenberg über ,,Die Physik der Elementarteilchen“. (German) [From Heisenberg’s address on “The physics of elementary particles”]. Mitteilungen der Max-Planck-Gesellschaft, 4(??):194–195, ???? 1956.

Heisenberg:1956:BPD

[Hei56d] Werner Heisenberg. Bemerkungen zu den Plänen für eine deutsche Reaktorstation. (German) [Comments on the plans for a German nuclear plant]. In B. Rajewski and G. Schreiber, editors, Aus der deutschen Forschung der letzten Dezennien: Dr. Ernst Telschow zum 65. Geburtstag gewidmet. 31. Oktober 1954. (German) [On German research in the last decades: dedicated to Dr. Telschow’s 65th birthday, 31st October 1954], pages 121–123. Georg Thieme Verlag, Stuttgart, West Germany, 1956.

Heisenberg:1956:BNT

[Hei56e] Werner Heisenberg. Bemerkungen zur ,,Neuen Tamm–Dancoff-Methode“in der Quantentheorie der Wellen-felder. (German) [Remarks on the “New Tamm–Dancoff method” in the quantum theory of wave fields]. Nachr. Akad. Wiss. Göttingen. Math.-Phys. Kl. IIa, 1956:27–36, 1956. ISSN 0065-5295.

Heisenberg:1956:NHP

[Hei56f] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie / Hrsg.: Ernesto Grassi. Rowohlt, Reinbek bei Hamburg, Germany, 1956. 149 pp. LCCN ????

Heisenberg:1956:AHS

[Hei56g] Werner Heisenberg. Die Alexander von Humboldt-Stiftung: Berliner Jahre und Tätigkeitsbereich nach der Wiederbegrundung im Jahre 1953. (German) [The Alexander von Humboldt Foundation: Berlin years and activity on the reestablishment in 1953]. Mitteilungen aus der Max-Planck-Gesellschaft, 1(??):40–46, ???? 1956.

Heisenberg:1956:EDQ

[Hei56h] Werner Heisenberg. Die Entwicklung der Deutung der Quantentheorie. (German) [The development of the interpretation of the quantum theory]. Physikalische Blätter, 12(7):289–304, July 1956.

CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/

phbl.19560120701/abstract. English translation in [Hei55d].

Heisenberg:1956:EZG

[Hei56i] Werner Heisenberg. Die europaische Zusammenarbeit auf dem Gebiet der Atomphysik. (German) [European cooperation in the field of atomic physics]. In G. Erler and H. Kruse, editors, Deutsches Atomenergierecht. Erganzbare Sammlung geltender Vorschriften auf dem Gebiet der Atomenergie. (German) [German Atomic Energy Law. ???? collection of existing legislation in the field of Atomic,], pages 1955–1962. Verlag Otto Schwartz, Göttingen, West Germany, 1956.

Heisenberg:1956:MAA

[Hei56j] Werner Heisenberg. Die Möglichkeiten der angewandten Atomforschung in Deutschland. (German) [The possibilities of applied nuclear research in Germany]. Landtagsamt, Munich, West Germany, 1956. 15 pp.

Heisenberg:1956:WEG

[Hei56k] Werner Heisenberg. Die Wechselwirkung der Elementarteilchen. (German) [The interaction of elementary particles]. Forschung und Fortschritte, 30(??):193–194, ???? 1956.

Heisenberg:1956:DDB

[Hei56l] Werner Heisenberg. Discussion of D. I. Blokhintsev’s talk “on the generation of mesons in collisions of high energy nucleons”. In A. Citron et al., editors, CERN Symposium on High Energy Accelerators and Pion Physics, Geneva, 11–23 June 1956. Proceedings. Volume 2. Pion Physics, pages 157–158. CERN, Geneva, Switzerland, 1956. LCCN ????

Heisenberg:1956:EHR

[Hei56m] Werner Heisenberg. Erweiterungen des Hilbert-Raums in der Quantentheorie der Wellenfelder. (German) [Extensions of Hilbert space in the quantum theory of wave fields]. Zeitschrift für Physik, 144(1–3):1–8, February 1956. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://www.springerlink.

com/content/t1r654687012m058/.

Heisenberg:1956:GAF

[Hei56n] Werner Heisenberg. Grundlagenforschung und angewandte Forschung in der Atomphysik. (German) [Basic research and applied research in atomic physics]. In ????, editor, Wissenschaft und Wirtschaft: Festschrift der Metallgesellschaft. (German) [Science and industry: the commemorative of the Metal Society], pages 153–164. Metallgesellschaft, Frankfurt am Main, West Germany, 1956. English translation in [Hei56p].

Heisenberg:1956:HSI

[Hei56o] Werner Heisenberg. Hilbert space II and the “ghost” states of Pauli and Källén. Il Nuovo Cimento (10), 4(Supplemento):743– 747, 1956. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Heisenberg:1956:PAR

[Hei56p] Werner Heisenberg. Pure and applied research in atomic physics. In ????, editor, Science and Economic Order, page ?? Metallgesellschaft, Frankfurt am Main, West Germany, 1956. English translation in [Hei56n].

Heisenberg:1957:PPT

[Hei57a] W. Heisenberg. Les principes physiques de la théorie des quanta. (French) [Physical principles of quantum theory]. Gauthier-Villars et cie, Paris, France, 1957. x + 124 + 2 pp. Préface de Louis de Broglie.

Heisenberg:1957:KVA

[Hei57b] Werner Heisenberg. 1000 Kilometer Vorfeld ohne Atomwaffen. (German) [1000 km run without nuclear weapons]. Frankfurter Allgemeine Zeitung, 109(??):3, May 11, 1957.

Heisenberg:1957:NHP

[Hei57c] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1957. 149 pp. LCCN ???? Edited by Ernesto Grassi.

Heisenberg:1957:PFR

[Hei57d] Werner Heisenberg. Die Plane für eine Reaktorstation im Gebiet der Bundesrepublik. (German) [The plans for a reactor facility within the territory of the Federal Republic [of Germany]].

In H. Zimmermann, editor, Zur Ökonomik und Technik der Atomzeit. (German) [The economics and technology of the Atomic Age], pages 1–7. List-Gesellschaft, Tübingen, West Germany, 1957.

Heisenberg:1957:EAG

[Hei57e] Werner Heisenberg. Erklärung von 18 Atomforschern. (German) [Declaration of 18 atomic scientists]. Mitteilungen aus der Max-Planck-Gesellschaft, 2(??):62–64, ???? 1957.

Heisenberg:1957:LMQ

[Hei57f] Werner Heisenberg. Lee model and quantization of non linear field equations. Nuclear Physics, 4:532–563, August/November 1957. CODEN NUPHA7. ISSN ???? URL http://labs.adsabs.

harvard.edu/ui/abs/1957NucPh...4..532H.

Heisenberg:1957:QTF

[Hei57g] Werner Heisenberg. Quantum theory of fields and elementary particles. Reviews of Modern Physics, 29(3):269–278, July 1957. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.29.269;   http://rmp.aps.org/abstract/
RMP/v29/i3/p269_1.

Heisenberg:1958:AUF

[Hei58a] Werner Heisenberg. Allgemeiner Überblick und Fortschritte in der Theorie der Elementarteilchen. (German) [General review and progress of the theory of elementary particles]. Naturwissenschaften, 45(3):2–??, January 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). Mitteilungen der GDNÄ.

Heisenberg:1958:ATD

[Hei58b] Werner Heisenberg. Application of the Tamm–Dancoff-method to the Lee-model. Nuclear Physics, 5(1):195–201, January/ March 1958. CODEN NUPHA7. ISSN ???? URL

http://labs.adsabs.harvard.edu/ui/abs/1958NucPh...5..
195H;   http://www.sciencedirect.com/science/article/
pii/0029558258900208.

Heisenberg:1958:APC

[Hei58c] Werner Heisenberg. Atomic physics and causal law. In The physicist’s conception of nature [Hei58r], pages 32–50. LCCN Q175 .H393. English translation of [Hei52a].

Heisenberg:1958:AGA

[Hei58d] Werner Heisenberg. Atomphysik. (German) [Atomic physics]. In ????, editor, Forschungsberichte 2. Aus dem Schwerpunktsprogramm der Deutschen Forschungsgemeinschaft. (German) [Research reports 2 from the priority program of the German Research Foundation], pages 51–54. Franz Steiner Verlag, Wiesbaden, West Germany, 1958.

Heisenberg:1958:NHP

[Hei58e] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1958. 149 pp. LCCN ????

Heisenberg:1958:PPQ

[Hei58f] Werner Heisenberg. Die physikalischen Prinzipien der Quantentheorie. (German) [The physical principles of quantum mechanics]. Bibliographisches Institut, Mannheim, Germany, 1958. viii + 117 pp. LCCN ????

Heisenberg:1958:PEPa

[Hei58g] Werner Heisenberg. Die Plancksche Entdeckung und die philosophischen Grundfragen der Atomlehre. (German) [Planck’s discovery and the philosophical basic questions of atomic theory]. Naturwissenschaften, 45(10):227–234, January 1958. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/wupx14qt11u1631q. En-

glish translation in [Hei74b].

Heisenberg:1958:PEPb

[Hei58h] Werner Heisenberg. Die Plancksche Entdeckung und die philosophischen Probleme der Atomphysik. (German) [Planck’s discovery, and the philosophical problems of atomic physics]. Universitas: Zeitschrift für interdisziplinäre Wissenschaft, 14(??): 135–148, ???? 1958. CODEN UNIVA8. ISSN 0041-9079. Reprint of [Hei58g].

Heisenberg:1958:PEPc

[Hei58i] Werner Heisenberg. Die Plancksche Entdeckung und die philosophischen Probleme der Atomphysik. (German) [Planck’s discovery, and the philosophical problems of atomic physics]. In Louis Leprince-Ringuet et al., editors, L’Homme et Atome: textes des conférences et des entretiens organisés par les Rencontres internationales de Genève, 1958. (French) [Man and the atom: texts of lectures and discussions organized by the International Meetings of Geneva, 1958], pages 37–53. Editions de la Baconnière, ????, 1958. LCCN ???? English translation in [Hei58u].

Heisenberg:1958:F

[Hei58j] Werner Heisenberg. Festrede. In ????, editor, ????, page 16. Münchner Zeitungsverlag, Munich, West Germany, 1958. LCCN ???? The 800th anniversary of Munich.

Heisenberg:1958:FAJ

[Hei58k] Werner Heisenberg. Festrede [anläßlich der 800-Jahrfeier der Stadt München] (German) [Speech [on the occasion of the 800th anniversary of Munich]]. Münchener Zeitungsverlag, Munich, West Germany, 1958. 15 pp. LCCN ????

Heisenberg:1958:FMP

[Hei58l] Werner Heisenberg. Festrede (on M. Planck?s 100th birthday). (German) [Speech (on M. Planck’s 100th birthday)]. Stimmen aus dem Max-Gymnasium, 6(??):6–17, September 1958.

Heisenberg:1958:FTE

[Hei58m] Werner Heisenberg. Fortschritte in der Theorie der Elementarteilchen. (German) [Advances in the theory of elementary particles]. Wissenschaft und Fortschritt, 8(??):243–245, ???? 1958.

Heisenberg:1958:GRC

[Hei58n] Werner Heisenberg. General remarks concerning the theory of elementary particles. In ????, editor, International Conference on Mesons and Recently Discovered Particles e 43^{o} Congresso Nazionale di Fisica: Communicazioni, page ix. Ciclografia, Padova, Italy, 1958. LCCN ????

Heisenberg:1958:INC

[Hei58o] Werner Heisenberg. The idea of nature in contemporary physics. In The physicist’s conception of nature [Hei58r], pages 7–31. LCCN Q175 .H393. English translation of [Hei53c].

Heisenberg:1958:MFV

[Hei58p] Werner Heisenberg. A mai fizika világképe. (Hungarian) [Today’s physics worldview]. Studium könyvek. Gondolat, Budapest, Hungary, 1958. 134 pp. LCCN ???? Translation from German by Zoltán Morlin.

Heisenberg:1958:STT

[Hei58q] Werner Heisenberg. On the statistical theory of turbulence. Technical Memorandum 1431, National Advisory Committee for Aeronautics, Washington, DC, USA, 1958. English translation of [Hei48n].

Heisenberg:1958:PCN

[Hei58r] Werner Heisenberg. The physicist’s conception of nature. Harcourt, Brace, New York, NY, USA, 1958. 192 pp. LCCN Q175 .H393. Translated by Arnold J. Pomerans from the German original [Hei58e].

Heisenberg:1958:PPR

[Hei58s] Werner Heisenberg. Physics and philosophy: the revolution in modern science, volume 19 of World perspectives. Harper, New York, NY, USA, 1958. xv + 206 pp. LCCN QC6 .H473. Planned and edited by Ruth Nanda Anshen. Introduction by F. S. C. Northrop. German translation in [Hei58t].

Heisenberg:1958:PPG

[Hei58t] Werner Heisenberg. Physik und Philosophie. (German) [Physics and Philosophy]. S. Hirzel, Stuttgart, West Germany, 1958. 201 pp. LCCN ???? German translation of [Hei58s].

Heisenberg:1958:PQT

[Hei58u] Werner Heisenberg. Planck’s quantum theory and the philosophical problems of atomic physics. Universitas (English edition), 3 (??):105–118, ???? 1958. English translation of [Hei58h].

Heisenberg:1958:RNC

[Hei58v] Werner Heisenberg. The representation of nature in contemporary physics. Dædalus, 87(3):95–108, Summer 1958. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026454.

Heisenberg:1958:RNL

[Hei58w] Werner Heisenberg. Research on the non-linear spinor theory with indefinite metric in Hilbert space. In B. Ferretti, editor, 1958 International Conference on High-Energy Physics at CERN. Conference held June 30–July 5, 1958, in Geneva, Switzerland, pages 119–126. CERN, Geneva, Switzerland, 1958.

LCCN ???? URL http://labs.adsabs.harvard.edu/ui/abs/
1958hep..conf..119H.

Heisenberg:1959:NHP

[Hei59a] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie Sachgebiet Physik. Rowohlt, Reinbek bei Hamburg, Germany, 1959. 149 pp. LCCN ????

Heisenberg:1959:GVP

[Hei59b] Werner Heisenberg. Grundlegende Voraussetzungen in der Physik der Elementarteilchen. (German) [Basic conditions in the physics of elementary particles]. In ????, editor, Martin Heidegger zum siebzigsten Geburtstag. Eine Festschrift. (German) [Martin Heidegger’ 70th Birthday: A celebration], pages 291–297. Neske, Pfullingen, West Germany, 1959. LCCN ????

Heisenberg:1959:PPR

[Hei59c] Werner Heisenberg. Physics and philosophy: the revolution in modern science, volume 15 of World perspectives. George Allen & Unwin, London, UK, 1959. 176 pp. LCCN QC6 .H473.

Heisenberg:1959:PPG

[Hei59d] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Hirzel, Stuttgart, West Germany, 1959. 201 pp. LCCN ????

Heisenberg:1959:PPW

[Hei59e] Werner Heisenberg. Physik und Philosophie: Weltperspektiven. (German) [Physics and philosophy: world perspectives], volume 249 of Ullstein Buch. Ullstein Taschenbücher-Verlag, Frankfurt, West Germany, 1959. 196 pp. LCCN ????

Heisenberg:1959:RB

[Hei59f] Werner Heisenberg. Rudjer Bošković. In ????, editor, Actes du Symposium International R.J. Bošković 1968 [1958??] (Belgrad, Zagreb, Lubljana), pages 29–30. ????, ????, 1959. LCCN ????

Heisenberg:1959:UQG

[Hei59g] Werner Heisenberg. Untersuchungen zur Quantenlogik. (German) [Investigations on quantum logics]. Bayerische Akademie der Wissenschaften, Math.-Naturwiss. Kl. Sitzungsberichte, pages 38–39, October 9, 1959. Report on a paper of P. Mittelstaedt.

Heisenberg:1959:VGF

[Hei59h] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1958 bis zum 30. September 1959. (German) [Report of the Alexander von Humboldt Foundation about their activity from 1 October 1958 to 30th September 1959], page 3. Humboldt-Stiftung, Bonn, West Germany, 1959. LCCN ????

Heisenberg:1959:WGN

[Hei59i] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Stuttgart, West Germany, 1959. 183 pp. LCCN ????

Heisenberg:1959:WP

[Hei59j] Werner Heisenberg. Wolfgang Pauli 25.4.1900–15.12.1958. In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1959. (German) [Bavarian Academy of Science, Yearbook 1959], pages 183–187. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1959. LCCN ????

Heisenberg:1959:WPP

[Hei59k] Werner Heisenberg. Wolfgang Paulis philosophische Auffassungen. (German) [Wolfgang Pauli’s philosophical views]. Naturwissenschaften, 46(24):661–663, ???? 1959. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:/

/www.springerlink.com/content/n0472548308qm7m1. English

translation in [Hei61g].

Heisenberg:1960:NHP

[Hei60a] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie Sachgebiet Physik. Rowohlt, Reinbek bei Hamburg, Germany, 1960. 149 pp. LCCN ????

Heisenberg:1960:EZE

[Hei60b] Werner Heisenberg. Erinnerungen an die Zeit der Entwicklung der Quantenmechanik. (German) [Memories of the time of the evolution of quantum mechanics]. In Theoretical physics in the twentieth century (Pauli memorial volume), pages 40–47. Interscience Publishers, New York, NY, USA, 1960.

Heisenberg:1960:FIH

[Hei60c] Werner Heisenberg. Feier im Hörsaal des Instituts. (German) [Ceremony in the auditorium of the Institute]. Mitteilungen aus der Max-Planck–Gesellschaft, 6(??):325–338, ???? 1960.

Heisenberg:1960:CRL

[Hei60d] Werner Heisenberg. For the ‘Courier? Readers (letter). CERN Courier, 8(??):10, ???? 1960.

Heisenberg:1960:SPP

[Hei60e] Werner Heisenberg. La scoperta di Planck e i problemi filosofici della fisica atomica. (Italian) [The discovery of Planck and philosophical problems of atomic physics]. In Discussione sulla fisica moderna. (Italian) [Discussions on modern physics] [HSBA60], pages 3–33.

Heisenberg:1960:MLN

[Hei60f] Werner Heisenberg. Max von Laue [Nachruf]. (German) [Max von Laue [obituary]]. Mitteilungen aus der Max-Planck-Gesellschaft, 4(??):194–196, ???? 1960.

Heisenberg:1960:PCN

[Hei60g] Werner Heisenberg. The physicist’s conception of nature. American Journal of Physics, 28(1):82–83, January 1960. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/1960AmJPh..
28...82H.

Heisenberg:1960:RRN

[Hei60h] Werner Heisenberg. Recent research on the nonlinear spinor theory of elementary particles. In E. C. G. Sudarshan, J. H. Tinlot, and A. C. Melissinos, editors, Proceedings of the 1960 Annual International Conference on High Energy Physics at Rochester, pages 851–858. University of Rochester, Rochester, NY, USA, 1960. LCCN ????

Heisenberg:1960:SWM

[Hei60i] Werner Heisenberg. Sprache und Wirklichkeit in der modernen Physik. (German) [Language and reality in modern physics]. In ????, editor, Gestalt und Gedanke. Vol. 6: Wort und Wirklichkeit. Jahrbuch der Bayerischen Akademie der Schönen Künste. (German) [Shape and idea. Vol 6: Word and reality. Annual of the Bavarian Academy of Fine Arts], pages 32–62. Oldenbourg, Munich, West Germany, 1960. LCCN ????

Heisenberg:1960:VGF

[Hei60j] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1959 bis zum 30. September 1960. (German) [Report of the Alexander von Humboldt Foundation on its activities from 1 October 1959 to 30th September 1960], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1960. LCCN ????

Heisenberg:1961:AWN

[Hei61a] Werner Heisenberg. Alte Werte in neuen Formen. (German) [Old values in new forms]. Magnum, 35(??):39, April 1961.

Heisenberg:1961:DSN

[Hei61b] Werner Heisenberg. Der derzeitige Stand der nichtlinearen Spinortheorie der Elementarteilchen. (German) [The current state of the nonlinear spinor theory of elementary particles]. Acta Physica Austriaca, 14:328–339, 1961. CODEN APASAP. ISSN 0001-6713.

Heisenberg:1961:AMN

[Hei61c] Werner Heisenberg. Die Abstraktion in der modernen Naturwissenschaft. (German) [Abstraction in modern science]. In ????, editor, Reden und Gedenkworte. Orden Pour le Mérite für Wissenschaft und Künste. Vol. 4. (German) [Speeches and commemorative words. Pour le Mérite for Arts and Science. Vol 4], pages 141–164. Lambert Schneider, Heidelberg, West Germany, 1961. LCCN ???? English translation in [Hei74a].

Heisenberg:1961:ESG

[Hei61d] Werner Heisenberg. Erwin Schrödinger 12.8.1887–4.1.1961. (German). In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1961. (German) [Bavarian Academy of Science, Yearbook 1961], pages 194–196. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1961.

Heisenberg:1961:NTI

[Hei61e] Werner Heisenberg. Naturwissenschaft und Technik im politischen Geschehen unserer Zeit. (German) [Science and technology in the political events of our time]. In ????, editor, Dauer im Wandel. Festschrift zum 70. Geburtstag von Carl J. Burckhardt.

(German) [Period of change. Festschrift for the 70th birthday of Carl J. Burckhardt], pages 194–197. Georg D. W. Callwey, Munich, West Germany, 1961.

Heisenberg:1961:MPa

[Hei61f] Werner Heisenberg. On modern physics. Orion Press, London, UK, 1961. 108 pp. LCCN QC71 .O5 1961. English translation from the Italian by M. Goodman and J. W. Binns.

Heisenberg:1961:WPP

[Hei61g] Werner Heisenberg. Wolfgang Pauli’s philosophical views. Main Currents in Modern Thought, 17(??):51–54, ???? 1961. English translation of [Hei59k].

Heisenberg:1961:TSTa

[Hei61h] Werner Heisenberg. Zur Theorie der ‘seltsamen’ Teilchen (Kurzbericht). (German) [On the theory of strange particles (short report)]. Sitz. ber. Akad. Wiss., math.-naturwiss. Kl, ??(??):11– 12, March 3, 1961.

Heisenberg:1961:TSTb

[Hei61i] Werner Heisenberg. Zur Theorie der ‘seltsamen’ Teilchen (Kurzbericht). (German) [On the theory of strange particles (short report)]. Sitz. ber. Akad. Wiss., math.-naturwiss. Kl, ??(??):30– 32, October 6, 1961.

Heisenberg:1962:BIQ

[Hei62a] Werner Heisenberg. Bohrs Interpretation der Quantentheorie und die Physik der Elementarteilchen. (German) [Bohr’s interpretation of quantum theory and the physics of elementary particles]. Fysisk Tidsskrift, 60(??):47–53, ???? 1962.

Heisenberg:1962:NHP

[Hei62b] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, tenth edition, 1962. 149 pp. LCCN ????

Heisenberg:1962:DRH

[Hei62c] Werner Heisenberg. Discussion des rapports de Heitler et Feynman. (French) [Discussion of the reports of Heitler and Feynman]. In R. Stoops, editor, La Théorie Quantique des Champs. Rapports et Discussions du Douzième Conseil de Physique à Bruxelles, 9-14 Octobre, 1961, pages 92–97, 169–170, 173–176, 201–203, 243–247. Interscience Publishers, New York, NY, USA, 1962. LCCN ????

Heisenberg:1962:DDRa

[Hei62d] Werner Heisenberg. Discussion du rapport de Källén. (French) [Discussion of the report of Källén]. In R. Stoops, editor, La Théorie Quantique des Champs. Rapports et Discussions du Douzième Conseil de Physique à Bruxelles, 9-14 Octobre, 1961, pages 169–170, 173–176. Interscience Publishers, New York, NY, USA, 1962. LCCN ????

Heisenberg:1962:DDRc

[Hei62e] Werner Heisenberg. Discussion du rapport de Yukawa. (French) [Discussion of the report of Yukawa]. In R. Stoops, editor, La Théorie Quantique des Champs. Rapports et Discussions du Douzième Conseil de Physique à Bruxelles, 9-14 Octobre, 1961, pages 243–247. Interscience Publishers, New York, NY, USA, 1962. LCCN ????

Heisenberg:1962:ETE

[Hei62f] Werner Heisenberg. Einführung in die Theorie der Elementarteilchen. (German) [Introduction to the Theory of Elementary Particles]. Max-Planck-Institut für Physik und Astrophysik, Munich, West Germany, 1962. viii + 259 pp. LCCN ???? Elaborated by H. Rechenberg and K. Lagally.

Heisenberg:1962:ENL

[Hei62g] Werner Heisenberg. Electrodynamics in the non-linear spinor theory. In J. Prentki, editor, 1962 International Conference on High-Energy Physics at CERN. Conference held July 4–11, 1962, in Geneva, Switzerland, pages 675–680. CERN, Geneva, Switzerland,

1962. LCCN ???? URL http://labs.adsabs.harvard.edu/ui/
abs/1962hep..conf..675H.

Heisenberg:1962:FEF

[Hei62h] Werner Heisenberg. Fortschritte in der einheitlichen Feldtheorie der Elementarteilchen. (German) [Progress in the unified field theory of matter]. Naturwissenschaftliche Rundschau, 15(??):392– 393, ???? 1962. CODEN NARSAC. ISSN 0028-1050.

Heisenberg:1962:MBA

[Hei62i] Werner Heisenberg. Max Born zum achtzigsten Geburtstag. (German) [Max Born on his eightieth birthday]. Naturwissenschaften, 49(23):529, January 1962. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/p138q381372tp36r.

Heisenberg:1962:MPI

[Hei62j] Werner Heisenberg. Max-Planck-Institut für Physik und Astrophysik in München. In ????, editor, Max-Planck-Gesellschaft zur Forderung der Wissenschaften e.V. Jahrbuch 1961. Part II. (German) [Max Planck Society for the Promotion of Sciences E.v. Yearbook 1961. Part II], pages 632–643. Max-Planck-Gesellschaft, Göttingen, West Germany, 1962. LCCN ????

Heisenberg:1962:OQT

[Hei62k] Werner Heisenberg. The “Observable quantities” in the theory of elementary particles. Report UCRL-trans-808, Lawrence Radiation Laboratory, Berkeley, CA, USA, 1962. English translation of [Hei43m].

Heisenberg:1962:PPR

[Hei62l] Werner Heisenberg. Physics and philosophy; the revolution in modern science, volume TB 549 of Harper torchbooks: The science library. Harper & Row, New York, NY, USA, 1962. ISBN 0-06-130549-9. 213 pp. LCCN QC6 .H473 1962. Introduction by F. S. C. Northrop.

Heisenberg:1962:VGF

[Hei62m] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1960 bis zum 30. September 1961. (German) [Report of the Alexander von Humboldt Foundation about their activity from 1 October 1960 to the 30th September 1961], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1962.

Heisenberg:1962:ZKE

[Hei62n] Werner Heisenberg. Zamechaniya k eynshteynovskomu nabrosku yedinoy teorii polya. (Russian) [Einstein’s draft of a unified field theory]. In ????, editor, Eynshteyn i razvitiye fizikomathematicheskoy mysli. Sbornik statey, pages 63–68. Academy of Sciences, Moscow, USSR, 1962. LCCN ????

Heisenberg:1963:AMT

[Hei63a] Werner Heisenberg. Anpassung an die moderne technische Entwicklung. (German) [Adapting to the modern technical development]. In ????, editor, Eine Freundesgabe der Wissenschaft für Ernst Hellmut Vits zur Vollendung seines 60. Lebensjahres am 19. Dezember 1963. (German) [A friends gift of science for Ernst Hellmut Vits on his 60th birthday, 19 December 1963], pages 77– 83. Fritz Knapp Verlag, Frankfurt am Main, 1963. LCCN ????

Heisenberg:1963:NHP

[Hei63b] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of rowohlts deutsche enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1963. 149 pp. LCCN ????

Heisenberg:1963:EEF

[Hei63c] Werner Heisenberg. Die Entwicklung der einheitlichen Feldtheorie der Elementarteilchen. (German) [The development of the unified field theory of elementary particles]. Naturwissenschaften, 50(1): 3–7, ???? 1963. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/t8727t31283wn35j.

Heisenberg:1963:ENS

[Hei63d]  Werner Heisenberg.  Elektrodynamik in der nichtlinearen

Spinortheorie [Abstrakt]. (German) [Electrodynamics in the nonlinear spinor theory (abstract)]. In E. Brüche and H. Franke, editors, Physikertagung Stuttgart: Hauptvorträge der Jahrestagung 1962 des Verbandes Deutscher Physikalischer Gesellschaften. (German) [Physicists Conference Stuttgart: Keynotes of the 1962 Annual Meeting of the German Physical Society], page 93. Physik Verlag, Mosbach/Baden, West Germany, 1963. LCCN ????

Heisenberg:1963:GFE

[Hei63e] Werner Heisenberg. Gedenkworte für Erwin Schrödinger. (German) [Memorial words for erwin schrödinger]. In ????, editor, Orden Pour le Mérite für Wissenschaft und Künste. Reden und Gedenkworte. Vol. 5. (German) [Pour le Mérite for Science and Art. Speeches and commemorative words. Vol 5], pages 27–35. Lambert Schneider, Heidelberg, West Germany, 1963. LCCN ????

Heisenberg:1963:LRP

[Hei63f] Werner Heisenberg. Laudatio (für R. E. Peierls). Phys. Verh. (2), 3(??):354–357, ???? 1963.

Heisenberg:1963:NEP

[Hei63g] Werner Heisenberg. The nature of elementary particles. Pontificia Academia Scientiarum. Commentarii, 1(??):1–4, ???? 1963.

Heisenberg:1963:NB

[Hei63h] Werner Heisenberg. Niels Bohr 7.10.1885–18.11.1962. In ????, editor, Bayerische Akademie der Wissenschaften. Jahrbuch 1963. (German), pages 204–207. Bayerische Akademie der Wissenschaften, Munich, West Germany, 1963.

Heisenberg:1963:PST

[Hei63i] Werner Heisenberg. The present situation in the theory of elementary particles. In ????, editor, Commemoration of the Fiftieth Anniversary of Niels Bohr’s First Papers on Atomic Constitution Held in Copenhagen on 8–15 July 1963. Part 2: Session on Elementary Particles, pages 1–21. Institute for Theoretical Physics, Copenhagen, DK, 1963. URL ????.

Heisenberg:1963:RST

[Hei63j] Werner Heisenberg. Remarks on the ‘spurion’-theory of strange particles. In ????, editor, Rendiconti della Conferenza Internazionale di Siena sulle Particelle Elementari — Proceedings of the Sienna Conference on Elementary Particles. Volume 1, pages 445–447. Società Italiana di Fisica, Bologna, Italy, 1963. LCCN ????

Heisenberg:1963:SAF

[Hei63k] Werner Heisenberg. Stand und Aussichten der Forschung auf dem Gebiet der Plasmaphysik. (German) [State and prospects of research in the field of plasma physics]. Die Atomwirtschaft, 8(??): 290–291, ???? 1963.

Heisenberg:1963:VGF

[Hei63l] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1961 bis zum 30. September 1962. (German) [Report of the Alexander von Humboldt Foundation, on their activity from 1 October 1961 to 30 September 1962], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1963.

Heisenberg:1964:AEF

[Hei64a] Werner Heisenberg. Arbeiten zur einheitlichen Feldtheorie der Elementarteilchen. (German) [Work on the unified field theory of elementary particles]. Bayerische Akademie der Wissenschaften, Math.-Naturwiss. Kl. Sitzungsberichte, pages 30–32, November 6, 1964. Remarks on [DHYY65].

Heisenberg:1964:BDV

[Hei64b] Werner Heisenberg. Begrüßung durch den Vorsitzenden des Vor standsrates; Bericht über den Festakt am 7. Mai 1964 im Kongreßsaal des Deutschen Museums: Ehrungen. (German) [Welcome by the Chairman of the Board of Council, report on the ceremony on 7 May 1964 in the Congress Hall of the German Museum: Honours]. In ????, editor, Verwaltungsbericht über das 60. Geschäftsjahr 1963 und Bericht über die Jahresversammlung 1964 im Deutschen Museum. (German) [Management report for the 60th fiscal year 1963 and report on the 1964 Annual Meeting at the German Museum], pages 26, 51–54, 54–56. Deutsches Museum, Munich, West Germany, 1964.

Heisenberg:1964:WEE

[Hei64c] Werner Heisenberg. Es war einmal. Erinnerungen an den 31 Juli 1914. (German) [Once upon a time. Memories of 31 July 1914]. Epoca, 2(7):31, ???? 1964.

Heisenberg:1964:KOD

[Hei64d] Werner Heisenberg. Kvanteteorien og dens fortolkning. (Danish) [Quantum theory and its interpretation]. In Bohr and Rozental [BR64], pages 90–103. LCCN QC16.B63 N5. English translation in [Roz67, pages 94–108].

Heisenberg:1964:PFW

[Hei64e] Werner Heisenberg. Probleme bei der Forderung der wissenschaftlichen Forschung. (German) [Problems with the demand of the scientific research focuses]. Die neue Gesellschaft, 11(??): 184–194, ???? 1964.

Heisenberg:1964:VGFa

[Hei64f] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1962 bis zum 30. September 1963. (German) [Report of the Alexander von Humboldt Foundation on its activities from 1 October 1962 to 30th September 1963], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1964. LCCN ????

Heisenberg:1964:VGFb

[Hei64g] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Über den Stand der Wissenschaft in Deutschland im internationalen Vergleich. Auszüge aus Berichten von Forschungsstipendiaten der Alexander von Humboldt-Stiftung 1962/63. (German) [On the state of science in Germany by international standards. Extracts from reports research fellows of the Alexander von Humboldt Foundation 1962/63], pages 5–6. Humboldt-Stiftung, Bad Godesberg, West Germany, 1964. LCCN ????

Heisenberg:1964:WGG

[Hei64h] Werner Heisenberg. Walther Gerlach zum 75. Geburtstag. (German) [Walther Gerlach’ 75th birthday]. Süddeutsche Zeitung, 198 (??):16, July 31, 1964.

Heisenberg:1965:AMD

[Hei65a] Werner Heisenberg. Atomphysik und modernes Denken. (German) [Atomic physics and modern thought]. In ????, editor, Aufruf und Aufbruch. Festgabe für Günter Jacob. (German) [Call and leave. Festschrift for Günter Jacob], pages 118–138. Evangelische Verlagsanstalt, Berlin, West Germany, 1965.

Heisenberg:1965:HSU

[Hei65b] Werner Heisenberg. Der heutige Stand unserer Kenntnis von den Elementarteilchen. (German) [The present state of our knowledge of elementary particles]. In ????, editor, Wissenschaft und Menschheit. (German) [Science and Humanity], pages 207–217. Urania Verlag, Leipzig and Jena, East Germany; Berlin, West Germany, 1965. LCCN ????

Heisenberg:1965:EQG

[Hei65c] Werner Heisenberg. Die Entwicklung der Quantenmechanik. (German) [The development of quantum mechanics]. In ????, editor, Nobel Lectures Including Presentation Speeches and Laureates: Biographies: Physics 1922–1941, pages 290–301. Elsevier, Amsterdam, The Netherlands, 1965. LCCN ???? Reprint of [Hei34d].

Heisenberg:1965:MBG

[Hei65d] Werner Heisenberg. Die Mitwirkung der Bundesrepublik an den großen internationalen Forschungsprojekten. (German) [The participation of the Federal Republic [of Germany] in the great international research projects]. In B. Martin, editor, Jahrbuch der auswärtigen Kulturbeziehungen 1965. (German) [Yearbook of Foreign Cultural Relations 1965], pages 51–56. Akademischer Verlag, Bonn, West Germany, 1965. LCCN ????

Heisenberg:1965:TEG

[Hei65e] Werner Heisenberg. Die Theorie der Elementarteilchen. (German) [The theory of elementary particles]. Naturwissenschaftliche Rundschau, 18(??):70, ???? 1965. CODEN NARSAC. ISSN 0028-1050.

Heisenberg:1965:GNP

[Hei65f] Werner Heisenberg. Gespräch mit Nobelpreisträger Professor Dr. Werner Heisenberg Physik. (German) [Interview with Nobel Laureate Professor Dr. Werner Heisenberg, physics]. In ????, editor, Wohin führt die Wissenschaft? Neun Gespräche mit deutschen Gelehrten. Aus einer Sendereihe von Radio Bremen. Bremer Beitrage. (German) [Where is the science? Nine interviews with German scholars. From a series of Radio Bremen. Bremen contributions], pages 39–54. Verlag B. C. Heye, Bremen, West Germany, 1965.

Heisenberg:1965:SANa

[Hei65g] Werner Heisenberg. On the structure of atomic nuclei, I. In Nuclear Forces [Bri65a], pages 144–154. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Partial English translation of [Hei32f].

Heisenberg:1965:SANb

[Hei65h] Werner Heisenberg. On the structure of atomic nuclei, III. In Nuclear Forces [Bri65a], pages 155–160. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Partial English translation of [Hei33b].

Heisenberg:1965:PCS

[Hei65i] Werner Heisenberg. Particles as collective stationary states. Progress of Theoretical Physics, Supplement:298–303, 1965. CO-DEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/1965PThPS.

E65..298H. Supplement of the Progress of Theoretical Physics Commemoration Issue for the 30th Anniversary of the Meson Theory by Dr. H. Yukawa, 1965.

Heisenberg:1965:PPG

[Hei65j] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. ????, Frankfurt, West Germany, 1965. 197 pp. LCCN ????

Heisenberg:1965:QEN

[Hei65k] Werner Heisenberg. Quantum electrodynamics in the nonlinear spinor theory. In H. Rechenberg, editor, Proceedings of the Seminar on Unified Theories of Elementary Particles, July 1965, pages 281–299. MPI für Physik und Astrophysik, Munich, West Germany, 1965. URL ????.

Heisenberg:1965:SPC

[Hei65l] Werner Heisenberg. SCRAPS: The physicist’s concept of nature. The Physics Teacher, 3(7):336, October 1965. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Heisenberg:1965:SGO

[Hei65m] Werner Heisenberg. Stellungnahme. (German) [Opinion]. In S. Wiesenthal, editor, Verjährung. 200 Persönlichkeiten des öffentlichen Lebens sagen nein: Eine Dokumentation. (German) [Statute of Limitations. 200 personalities of the public life saying no: A documentary ], page 72. Europäische Verlagsanstalt, Frankfurt am Main, 1965. LCCN ????

Heisenberg:1965:SPS

[Hei65n] Werner Heisenberg. Survey on the present situation concerning unified theories of elementary particles. In H. Rechenberg, editor, Proceedings of the Seminar on Unified Theories of Elementary Particles, July 1965, pages 1–15. Max-Planck-Institut für Physik und Astrophysik, Munich, West Germany, 1965. LCCN ???? URL

????.

Heisenberg:1965:TWG

[Hei65o] Werner Heisenberg. Transparente Welt. (German) [Transparent world]. Huber, Bern, Switzerland, 1965. xviii + 360 pp. LCCN ????

Heisenberg:1965:BFM

[Hei65p] Werner Heisenberg. Über den Formenreichtum in der mathematischen Naturwissenschaft. (German) [On the variety of forms in mathematical science]. In ????, editor, Transparente Welt. Festschrift zum 60. Geburtstag von Jean Gebser. (German) [Transparent world. Festschrift for the 60th birthday of Jean Gebser], pages 1–5. Verlag Hans Huber, Bern, Switzerland; Stuttgart, West Germany, 1965. LCCN ????

Heisenberg:1965:VGF

[Hei65q] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1963 bis zum 30. September 1964. (German) [Report of the Alexander von Humboldt Foundation about their activity from 1 October 1963 to the 30th September 1964], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1965. LCCN ????

Heisenberg:1966:NHP

[Hei66a] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie Physik. Rowohlt-Taschenbuch-Verlag, Hamburg, West Germany, 1966. 150 pp. LCCN ????

Heisenberg:1966:KIT

[Hei66b] Werner Heisenberg. Die Künste im technischen Zeitalter. (German) [The Arts in the Age of Technology]. Akademie für das Graphische Gewerbe, Munich, West Germany, 1966. 96 pp. LCCN ????

Heisenberg:1966:RPT

[Hei66c] Werner Heisenberg. Die Rolle der phänomenologischen Theorien im System der theoretischen Physik. (German) [The role of phenomenological theories in systems of theoretical physics]. In A. De-Shalit, H. Feshbach, and L. van Hove, editors, Preludes in Theoretical Physics in honor of V. F. Weisskopf, pages 166– 169. North-Holland Publishing Co., Amsterdam, The Netherlands,

1966. LCCN ???? URL http://labs.adsabs.harvard.edu/ui/
abs/1966ptp..book..166H.

Heisenberg:1966:IUF

[Hei66d] Werner Heisenberg. Introduction to the unified field theory of elementary particles. Wiley, New York, NY, USA, 1966. ix + 177 pp. LCCN ????

Heisenberg:1966:LDIa

[Hei66e] Werner Heisenberg. Letter to David Irving. Report AMPG III-093-1743 01 088, Archiv der MPG, Berlin, Germany, June 10,

Heisenberg:1966:LDIb

[Hei66f] Werner Heisenberg. Letter to David Irving. Report AMPG III-093-1743 01 061, Archiv der MPG, Berlin, Germany, July 6, 1966.

Heisenberg:1966:NLS

[Hei66g] Werner Heisenberg. Natural law and the structure of matter. In Frontiers of modern scientific philosophy and humanism: The Athens Meeting, 1964 [Ano66], pages 26–42. LCCN Q175 .F9 1966.

Heisenberg:1966:VGFa

[Hei66h] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1964 bis zum 30. September 1965. (German) [Report of the Alexander von Humboldt Foundation on its activity from 1 October 1964 to 30th September 1965], pages 3–4. Humboldt-Stiftung, Bonn, West Germany, 1966. LCCN ????

Heisenberg:1966:VGFb

[Hei66i] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Ausbildung und Laufbahn von Hochschullehrern im Ausland. (German) [Education and career of university teachers abroad], page 7. Akademischer Verlag, Bonn, West Germany, 1966. LCCN ????

Heisenberg:1966:NT

[Hei66j] Werner Karl Heisenberg. Nuclear theory. In Henry A. Boorse and Lloyd Motz, editors, The World of the Atom, volume 1, pages 1607–?? Basic Books, New York, NY, USA, 1966. LCCN

QC173 B71 v.1. URL http://labs.adsabs.harvard.edu/ui/

abs/1966woat.book.1607H. With a foreword by I. I. Rabi.

Heisenberg:1967:DRV

[Hei67a] Werner Heisenberg. Das Dritte Reich versuchte nicht, die Atombombe zu bauen. (German) [The Third Reich did not try to build the atomic bomb]. Frankfurter Allgemeine Zeitung, Beilage ‘Bilder und Zeiten’, 286(??):4, December 9, 1967.

Heisenberg:1967:NGT

[Hei67b] Werner Heisenberg. Das Naturbild Goethes und die technisch naturwissenschaftliche Welt. (German) [The natural image of Goethe and the technically scientific world]. In A. B. Wachsmuth, editor, Goethe Neue Folge des Jahrbuchs der Goethe-Gesellschaft. (German) [Goethe New Series of the Yearbook of the Goethe Society], pages 27–42. Hermann Böhlaus Nachf., Weimar, East Germany, 1967. LCCN ???? English translation in [Hei74h].

Heisenberg:1967:EJD

[Hei67c] Werner Heisenberg. Ein Experiment aus dem Jahre 1942 in Deutschland. (German) [An experiment in 1942 in Germany]. Die Atomwirtschaft, 12(??):581, ???? 1967.

Heisenberg:1967:EEF

[Hei67d] Werner Heisenberg. Einführung in die einheitliche Feldtheorie der Elementarteilchen. (German) [Introduction to the unified field theory of elementary particles]. Hirzel, Stuttgart, West Germany, 1967. xi + 179 pp. LCCN ????

Heisenberg:1967:EGN

[Hei67e] Werner Heisenberg. Einstein ist ein guter Name. (German) [Einstein is a good name]. Süddeutsche Zeitung, 27(??):13, January 27, 1967.

Heisenberg:1967:EGM

[Hei67f] Werner Heisenberg. Erinnerungen. (German) [Memoirs]. In ????, editor, The Yoshio Fujioka Commemorative Issue, pages 359–360. Institute for Optical Research and University of Education, Tokyo, Japan, 1967. LCCN ????

Heisenberg:1967:GSD

[Hei67g] Werner Heisenberg. Gott sei Dank, wir konnten sie nicht bauen (Spiegel-Gesprach). (German) [Thank God we could no build them (Spiegel conversation)]. Der Spiegel, 28(??):79–83, July 3, 1967.

Heisenberg:1967:HMGa

[Hei67h] Werner Heisenberg. Harmonie der Materie. Ein Gesprach mit Professor Werner Heisenberg. (German) [Harmony of matter. A conversation with Professor Werner Heisenberg]. Neue Stadt, 10(2): 7–9, February 1967.

Heisenberg:1967:HMGb

[Hei67i] Werner Heisenberg. Harmonie der Materie. Ein Gesprach mit Professor Werner Heisenberg. (German) [Harmony of matter. A conversation with Professor Werner Heisenberg]. Neue Stadt, 10(3): 14–15, March 1967.

Heisenberg:1967:ISM

[Hei67j] Werner Heisenberg. Interview with the Spiegel magazine. Der Spiegel, 28(??):??, July 3, 1967. URL http://alsos.wlu.edu/

information.aspx?id=1997.

Heisenberg:1967:TDC

[Hei67k] Werner Heisenberg. La théorie du champ unitaire. (French) [The theory of the unitary field]. In ????, editor, Science et Synthèse, pages 38–47. Galimard, Paris, France, 1967. LCCN ???? Translation from English original in [Hei67m].

Heisenberg:1967:NPP

[Hei67l] Werner Heisenberg. Nonlinear problems in physics. Physics Today, 20(5):27–33, May 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v20/i5/p27/s1.

Heisenberg:1967:UFT

[Hei67m] Werner Heisenberg. The unified field theory. In ????, editor, Science and Synthesis. An International Colloquium organized by UNESCO on the Tenth Anniversary of the Death of Einstein and Teilhard de Chardin, pages 12–17. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. LCCN ???? French translation in [Hei67k].

Heisenberg:1967:VGFa

[Hei67n] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Gästehaus der Alexander von Humboldt-Stiftung errichtet mit den Mitteln der Stiftung Volkswagenwerk. (German) [Guesthouse of the Alexander von Humboldt Foundation built with funds from the Volkswagen Foundation], page 3. Bundesdruckerei, Bonn, West Germany, 1967. LCCN ????

Heisenberg:1967:VGFb

[Hei67o] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit vom 1. Oktober 1965 bis zum 30. September 1966, pages 7–8. Humboldt-Stiftung, Bonn, West Germany, 1967. LCCN ????

Heisenberg:1967:TLK

[Hei67p] Werner Heisenberg. Zur Theorie der Leptonenmassen (Kurzbericht). (German) [On the theory of lepton masses (short report)].

Bayerische Akademie der Wissenschaften, Math.-Naturwiss. Kl. Sitzungsberichte, pages 7–9, January 13, 1967.

Heisenberg:1968:AWD

[Hei68a] Werner Heisenberg. Ausländische Wissenschaftler an deutschen Hochschulen. (German) [Foreign scientists at German universities]. Süddeutsche Zeitung, 185(??):2, August 2, 1968.

Heisenberg:1968:BEK

[Hei68b] Werner Heisenberg. Braucht Europas Kernforschung das ‘Super-PS’? (Interview). (German) [Does Europe need nuclear research, the ‘super-PS’?]. Die Welt, 149(??):2, June 29, 1968.

Heisenberg:1968:NHP

[Hei68c] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1968. 149 pp. LCCN ????

Heisenberg:1968:KES

[Hei68d] Werner Heisenberg. Kosmologie, Elementarteilchen, Symmetrie. (German) [Cosmology, elementary particles, symmetry]. Naturwissenschaftliche Rundschau, 21(??):389–390, ???? 1968. CODEN NARSAC. ISSN 0028-1050.

Heisenberg:1968:OHD

[Hei68e] Werner Heisenberg. Otto Hahn, discoverer of nuclear fission, dies. Physics Today, 21(10):101–??, October 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v21/i10/p101/s1.

Heisenberg:1968:OHT

[Hei68f] Werner Heisenberg. Otto Hahn. Zum Tode des Gelehrten. (German) [Otto Hahn. On the death of the scholar]. Süddeutsche Zeitung, 186(??):89–90, August 3–4, 1968.

Heisenberg:1968:RPS

[Hei68g] Werner Heisenberg. Report on the present situation in the nonlinear spinor theory of elementary particles. In ????, editor, Fundamental Problems in Elementary Particle Physics: Proceedings of the Fourteenth Conference on Physics at the University of Brussels, October 1967, pages 129–144. Interscience Publishers, New York, NY, USA, 1968. LCCN ????

Heisenberg:1968:SJV

[Hei68h] Werner Heisenberg. ‘Si j’avais vingt ans’ (interview). (French) [‘If I were 20 (interview)]. Le Nouvel Observateur, ??(??):22–25, April 30, 1968.

Heisenberg:1968:TRA

[Hei68i] Werner Heisenberg. The Third Reich and the atomic bomb. Bulletin of the Atomic Scientists, 24(6):34–35, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://alsos.wlu.edu/information.aspx?id=157. Translated

by Margaret Seckel from the German original.

Heisenberg:1968:VFS

[Hei68j] Werner Heisenberg. Vorwort für die Sommerfeld-Gesamtausgabe. (German) [Foreword to the Sommerfeld Complete Edition]. In F. Sauter, editor, Arnold Sommerfeld. Gesammelte Schriften. (German) [Arnold Sommerfeld. collected writings], page 3. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1968.

Heisenberg:1968:VGF

[Hei68k] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über die Tätigkeit im Jahre 1967. (German) [Report of the Alexander von Humboldt Foundation of the activity in 1967], page 7. Humboldt-Stiftung, Bad Godesberg, West Germany, 1968. LCCN ????

Heisenberg:1969:AGV

[Hei69a] Werner Heisenberg. Abschiedsworte. (German) [Valediction]. In ????, editor, Zur Erinnerung an Paul Scherrer-Sonderegger. 3. Februar 1880 bis 25. September 1969. (German) [In memory of Paul Scherrer Sonderegger. Third February 1880 to 25 September 1969], pages 9–11. ????, Zurich, Switzerland, 1969. LCCN ????

Heisenberg:1969:ABA

[Hei69b] Werner Heisenberg. Auszüge aus der Begrüßungsansprache anläßlich der Einführungstagung der Humboldt-Stipendiaten in Tübingen, 1968. (German) [Excerpts from the welcome speech on the occasion of the introductory meeting of the Humboldt Fellows in Tübingen, 1968]. Alexander von Humboldt-Stiftung. Mitteilungen, 17(??):32–35, ???? 1969.

Heisenberg:1969:CUT

[Hei69c] Werner Heisenberg. The concept of ‘understanding’ in theoretical physics. In H. Mark and S. Fernbach, editors, Properties of Matter Under Unusual Conditions, pages 7–10. Interscience Publishers, New York, NY, USA, 1969. LCCN ????

Heisenberg:1969:TGG

[Hei69d] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. R. Piper, München, West Germany, 1969. ISBN 3-492-01791-6. 333 pp. LCCN ????

Heisenberg:1969:ENA

[Hei69e] Werner Heisenberg. Die Einheit der Natur bei Alexander von Humboldt und in der Gegenwart. (German) [The unity of nature at the Alexander von Humboldt Foundation and in the present]. Alexander von Humboldt-Stiftung. Mitteilungen, 18(??): 17–25, ???? 1969.

Heisenberg:1969:GOH

[Hei69f] Werner Heisenberg. Gedenkworte für Otto Hahn und Lise Meitner. (German) [Words of commemoration for Otto Hahn and Lise Meitner]. In ????, editor, Orden Pour le Mérite für Wissenschaften und Künste. Reden und Gedenkworte. (German) [Pour le Mérite for Arts and Sciences. Speeches and commemorative words], pages 111–119. Verlag Lambert Schneider, Heidelberg, West Germany, 1969. LCCN ????

Heisenberg:1969:GPG

[Hei69g] Werner Heisenberg. Gewaltlosigkeit in der Politik. (German) [Nonviolence in politics]. In ????, editor, Mahatma Gandhi 1869/1969, pages 67–68. Inter Nationes, Bonn, West Germany, 1969. LCCN ????

Heisenberg:1969:NP

[Hei69h] Werner Heisenberg. Nuclear Physics. Greenwood Press, Westport, CT, USA, 1969. ISBN 0-8371-2089-6. viii + 224 pp. LCCN QC173 .H3854. Reprint of [Hei53m]. See also [Hei47g].

Heisenberg:1969:ONI

[Hei69i] Werner Heisenberg. Otázky a názory (interview). (Czech) [Questions and comments]. Csekoslovensky Casopis pro Fysikul, A19 (??):93–95, ???? 1969.

Heisenberg:1969:SSW

[Hei69j] Werner Heisenberg. Significance of Sommerfeld’s work today. In F. Bopp and H. Kleinpoppen, editors, Physics of the One- and Two-Electron Atoms, Proceedings of the Arnold Sommerfeld Centennial Memorial Meeting and of the International Symposium on the Physics of the One- and Two-Electron Atoms, 10–14 September, 1968, Munich, pages 44–52. North-Holland Publishing Co., Amsterdam, The Netherlands, 1969. LCCN ???? URL http://

labs.adsabs.harvard.edu/ui/abs/1969pote.conf...44H.

Heisenberg:1969:BMU

[Hei69k] Werner Heisenberg. Über die Möglichkeit universeller wissenschaftlicher Bildung in unserem Zeitalter. (German) [On the possibility of universal scientific education in our age]. In H. Pfeiffer, editor, Alexander von Humboldt: Werk und Weltgeltung. (German) [Alexander von Humboldt: business and international standing], pages 9–13. R. Piper Verlag, Munich, West Germany, 1969. LCCN ????

Heisenberg:1969:MUW

[Hei69l] Werner Heisenberg. Über die Möglichkeit universeller wissenschaftlicher Bildung in unserem Zeitalter. (German) [On the possibility of universal scientific education in our age]. In Pfeiffer [Pfe69], page ???? LCCN Q143.H9 A68.

Heisenberg:1969:VGFa

[Hei69m] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Bericht der Alexander von Humboldt-Stiftung über ihre Tätigkeit im Jahre 1968. (German) [Report of the Alexander von Humboldt Foundation on its activities in 1968], page 7. Humboldt-Stiftung, Bonn, West Germany, 1969. LCCN ????

Heisenberg:1969:VGFb

[Hei69n] Werner Heisenberg. Vorwort. (German) [Foreword]. In Born and Born [BB69], pages 9–13. Preface by Bertrand Russell and foreword by Werner Heisenberg.

Heisenberg:1969:SQG

[Hei69o] Werner Heisenberg. Zur Sprache in der Quantentheorie. (German) [On the language of quantum theory]. Physikalische Blätter, 25 (3):112–113, March 1969. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Heisenberg:1970:APG

[Hei70a] Werner Heisenberg. Abschluß der Physik. (German) [Completion of physics]. Süddeutsche Zeitung, ??(??):44, October 6, 1970. Reprinted in [Hei71a]. English translation in [Hei74f].

Heisenberg:1970:ADI

[Hei70b]  Werner Heisenberg.  Änderungen der Denkstrukturen im

Fortschritt der Wissenschaft. (German) [Changes in the structures of thought in the progress of science]. Zeitschrift fur die Einheit der Wissenschaften im Studium generale, 23(9):808–816, ???? 1970. ISSN 0039-4149.

Heisenberg:1970:NHP

[Hei70c] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of rowohlts deutsche enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, fifteenth edition, 1970. ISBN 3-499-55008-3 (kart.). 149 pp. LCCN ????

Heisenberg:1970:TAM

[Hei70d] Werner Heisenberg. Die Tendenz zur Abstraktion in moderner Kunst und Wissenschaft. (German) [The tendency toward abstraction in modern art and science]. In K. Gaiser, editor, Das Altertum und jedes neue Gute. Festschrift für Wolfgang Schadewaldt zum 15. März 1970. (German) [Antiquity and each new good. Festschrift for Wolfgang Schadewaldt 15 March 1970], pages 485–494. W. Kohihammer, Stuttgart, West Germany, 1970. LCCN ????

Heisenberg:1970:GAA

[Hei70e] Werner Heisenberg. Genesis of the atomic age: Physics and beyond: Encounters and conversations: The responsibility of the scientist (1945–1950). Bulletin of the Atomic Scientists, 26(9):33– 36, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from [Hei71j].

Heisenberg:1970:ITU

[Hei70f] Werner Heisenberg. Industrie und Technik von übermorgen (Interview). (German) [Industry and technology of tomorrow (interview)]. Süddeutsche Zeitung, Sonderbeilage “Bayern baut an der Zukunft”, 25(??):8, ???? 1970.

Heisenberg:1970:MB

[Hei70g] Werner Heisenberg. Max Born. Physics Bulletin, 21(??):137, ???? 1970. CODEN PHSBB4. ISSN 0031-9112.

Heisenberg:1970:PPT

[Hei70h] Werner Heisenberg. Philosophische Probleme in der Theorie der Elementarteilchen. (German) [Philosophical problems in the theory of elementary]. In ????, editor, Sächsische Akademie der Wissenschaften. Jahrbuch 1966–1968. (German) [Saxony Academy of Sciences, Yearbook 1966–1968], pages 136–137. ????, ????, 1970. LCCN ????

Heisenberg:1970:PCN

[Hei70i] Werner Heisenberg. The physicist’s conception of nature. Greenwood Press, Westport, CT, USA, 1970. ISBN 0-8371-3107-3. 192 pp. LCCN Q175 .H393 1970.

Heisenberg:1970:PHL

[Hei70j] Werner Heisenberg. Professor Heisenberg’s lecture (atoms and peace). In ????, editor, Dansk Ingeniørforening. Award of the Niels Bohr Gold Medal to Werner Karl Heisenberg. 7th October 1970, pages 20–33. Dansk Ingeniørforening, Copenhagen, DK, 1970. LCCN ????

Heisenberg:1971:APG

[Hei71a] Werner Heisenberg. Abschluß der Physik. (German) [Completion of physics]. In Schritte über Grenzen. (German) [Steps across borders] [Hei71l], pages 306–313. LCCN ???? Reprint of [Hei70a].

Heisenberg:1971:KWI

[Hei71b] Werner Heisenberg. Das Kaiser-Wilhelm-Institut für Physik. Geschichte eines Instituts. (German) [The Kaiser Wilhelm Institute for Physics. History of an institution]. In ????, editor, Max-Planck-Gesellschaft zur Forderung der Wissenschaften e.V. Jahrbuch 1971. (German) [Max Planck Society for the Promotion of E.v. Sciences Yearbook 1971], pages 46–89. Max-Planck-Gesellschaft, Munich, West Germany, 1971. LCCN ????

Heisenberg:1971:BKT

[Hei71c] Werner Heisenberg. Der Begriff der kleinsten Teilchen in der Entwicklung der Naturwissenschaft. (German) [The concept of the smallest particles in the development in science]. In ????, editor, Meyers Enzyklopadisches Lexikon. Vol. 2, pages 870–879. Bibliographisches Institut, Mannheim, West Germany, 1971. LCCN ????

Heisenberg:1971:BSE

[Hei71d] Werner Heisenberg. Die Bedeutung des Schönen in den exakten Naturwissenschaften. (German) [The meaning of beauty in the exact natural sciences]. In ????, editor, Ensemble 2, pages 228– 243. R. Oldenbourg, Munich, West Germany, 1971. LCCN ????

Heisenberg:1971:EC

[Hei71e] Werner Heisenberg. Encounters and conversations. Physics and beyond. Allen and Unwin, London, UK, 1971. viii + 256 pp. LCCN ???? English translation by Arnold J. Pomerans.

Heisenberg:1971:CMB

[Hei71f] Werner Heisenberg. In commemoration of Max Born. Universitas (English edition), 12(??):289–294, ???? 1971. English translation of [Hei71h].

Heisenberg:1971:I

[Hei71g] Werner Heisenberg. Introduction. In Born and Born [BB71a], page ?? ISBN 0-333-11267-9. LCCN QC16.E5 A4513 1971b. Preface by Bertrand Russell and foreword by Werner Heisenberg. Translated by Irene Born.

Heisenberg:1971:MBD

[Hei71h] Werner Heisenberg. Max Born 11 Dezember 1882–5. Januar 1970 [Nachruf]. (German) [[obituary]]. In ????, editor, Jahrbuch der Akademie der Wissenschaften in Göttingen für das Jahr 1970. (German) [Yearbook of the Academy of Sciences in Göttingen for 1970], pages 59–64. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1971. LCCN ???? English translation in [Hei71f].

Heisenberg:1971:PBEa

[Hei71i] Werner Heisenberg. Physics and beyond: encounters and conversations, volume 23 of World perspectives. G. Allen and Unwin, London, UK, 1971. ISBN 0-04-925020-5 (paperback), 0-04-925008-6. viii + 256 pp. LCCN QC173 .H38613x 1971b.

Heisenberg:1971:PBEb

[Hei71j] Werner Heisenberg. Physics and beyond: encounters and conversations, volume 42 of World perspectives. Harper & Row, New York, NY, USA, 1971. xviii + 247 pp. LCCN QC173 .H38613 1971. Translated from the German by Arnold J. Pomerans.

Heisenberg:1971:RRP

[Hei71k] Werner Heisenberg. Review of recent progress. In H. P. Dürr, editor, Proceedings of the Symposium on Basic Questions in Elementary Particle Physics, June 8–18, 1971, pages 1–10. Max-Planck-Institut für Physik under Astrophysik, Munich, West Germany, 1971. LCCN ????

Heisenberg:1971:SGG

[Hei71l] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1971. 313 pp. LCCN ????

Heisenberg:1971:VGF

[Hei71m] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Alexander von Humboldt-Stiftung. Jahresbericht 1970. (German) [Alexander von Humboldt Foundation. Annual report 1970], pages 5–6. Humboldt-Stiftung, Bonn, West Germany, 1971. LCCN ????

Heisenberg:1971:WAH

[Hei71n] Werner Heisenberg. Zum wissenschaftlichen Austausch der Humboldt-Stiftung. (German) [For a scientific exchange of the Humboldt Foundation]. Alexander von Humboldt-Stiftung. Mitteilungen, 21(??):24–25, ???? 1971.

Heilbron:1972:BRP

[Hei72a] John L. Heilbron. Book review: Physics in My Generation by Max Born; Bahnbrecher des Atomzeitalters. Grosse Naturforscher von Maxwell bis Heisenberg by Friedrich Herneck; Moseley and the Numbering of the Elements by Bernard Jaffe; The Big Machine by Robert Jungk. Isis, 63(1):111–112, March 1972. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/229207;   http:
//www.jstor.org/stable/pdfplus/229207.pdf.

Heisenberg:1972:BNT

[Hei72b] Werner Heisenberg. Bemerkungen zur ,,neuen Tamm–Dancoff-Methode“. (German) [Comments on the new “Tamm–Dancoff method”]. In Problems of theoretical physics (in memory of I. E. Tamm) (Russian), pages 34–36, 483, 490. Nauka, Moscow, Russia,

Heisenberg:1972:CFW

[Hei72c] Werner Heisenberg. C. F. von Weizsäcker zum 60. Geburtstag (28 Juni 1972). (German) [C. F. von Weizsäcker’s 60th birthday (28 June 1972)]. Physikalische Blätter, 28(7):319–321, July 1972. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19720280706/abstract.

Heisenberg:1972:CLA

[Hei72d] Werner Heisenberg. Construction of large accelerators: Scientific and political aspects. Science, 179(??):643–647, ???? 1972. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). English translation of [Hei72l].

Heisenberg:1972:TGG

[Hei72e] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. R. Piper, München, West Germany, fourth edition, 1972. ISBN 3-492-01791-6. 333 pp. LCCN ????

Heisenberg:1972:ASG

[Hei72f] Werner Heisenberg. Die Aufgabe der Speicherringe: Rede bei der Einweihung der Speicherringe in CERN am 16.10.1971. (German) [The task of the storage rings: Speech at the inauguration of the storage rings in CERN on 16.10.1971]. Physikalische Blätter, 28 (3):107–111, March 1972. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19720280304/abstract.

Heisenberg:1972:EKB

[Hei72g] Werner Heisenberg. Die Evolution ist kein Betriebsunfall. (German) [Evolution is not an accident]. In A. Reif, editor, Interviews, pages 94–108. Hoffmann & Campe, Hamburg, West Germany, 1972. LCCN ????

Heisenberg:1972:IMS

[Hei72h] Werner Heisenberg. Indefinite metric in state space. In Salam and Wigner [SW72], pages 129–136. ISBN 0-521-08600-0.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.uk/
cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http://www.
loc.gov/catdir/enhancements/fy1001/72075298-d.html;

http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Heisenberg:1972:NHHb

[Hei72i] Werner Heisenberg. Naturwissenschaft in der heutigen Hochschule. (German) [Natural science in the modern university]. ????, Munich, West Germany, 1972. 18 pp. LCCN ????

Heisenberg:1972:NHHa

[Hei72j] Werner Heisenberg. Naturwissenschaft in der heutigen Hochschule. (German) [Science in today’s college]. Alexander von Humboldt-Stiftung. Mitteilungen, 23(??):8–15, ???? 1972.

Heisenberg:1972:PPG

[Hei72k] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Hirzel, Stuttgart, West Germany, 1972. 201 pp. LCCN ????

Heisenberg:1972:PWG

[Hei72l] Werner Heisenberg. Physikalische und wissenschaftspolitische Gesichtspunkte beim Bau von Großbeschleunigern. (German) [Physical and science policy considerations in the construction of large accelerators]. In ????, editor, Festgabe für Otto Mittelstaedt. Die wissen schaftliche Redaktion No. 7. (German) [Festschrift for Otto Mittelstaedt. The scientific editorial, number 7], pages 62– 74. Bibliographisches Institut, Mannheim, West Germany, 1972. LCCN ???? English translation in [Hei73b].

Heisenberg:1972:RPH

[Hei72m] Werner Heisenberg. Rede von Professor Heisenberg. (German) [Speech by Professor Heisenberg]. In ????, editor, Werner-Heisenberg-Gymnasium. Jahresbericht 1971/72. (German) [Werner Heisenberg High School. Annual report 1971/72], pages 41–44. Werner-Heisenberg-Gymnasium, Garching, West Germany, 1972. LCCN ????

Heisenberg:1972:VGF

[Hei72n] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Alexander von Humboldt-Stiftung. Jahresbericht 1971. (German) [Alexander von Humboldt Foundation. Annual Report 1971 ], pages 5–6. Humboldt-Stiftung, Bonn, West Germany, 1972. LCCN ????

Heisenberg:1973:BFW

[Hei73a] Werner Heisenberg. Begrüßungsworte an die Forschungsstipendiaten während der Jahrestagung 1973. (German) [Greeting to the Research Fellows during the annual meeting 1973]. Alexander von Humboldt-Stiftung. Mitteilungen, 26(??):14–16, ???? 1973.

Heisenberg:1973:CLA

[Hei73b] Werner Heisenberg. Construction of large accelerators: Scientific and political aspects. Science, 179(4074):643–647, February 16, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/179/4074/
643.full.pdf.

Heisenberg:1973:TGG

[Hei73c] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1973. 287 pp. LCCN ????

Heisenberg:1973:DCH

[Hei73d] Werner Heisenberg. Development of concepts in the history of quantum theory. In Mehra [Meh73], chapter 11, pages 264–275. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972.

URL http://www.springer.com/us/book/9789027703453.

Heisenberg:1973:RAT

[Hei73e] Werner Heisenberg. Die Richtigkeitskriterien der abgeschlossenen Theorien in der Physik. (German) [The correctness criteria of complete theories in physics]. In E. Scheibe and G. Süßmann, editors, Einheit und Vielheit. Festschrift für Carl Friedrich von Weizsäcker zum 60. Geburtstag. (German) [Unity and multiplicity. Festschrift for Carl Friedrich von Weizsäcker’s 60th birthday], pages 140–144. Vandenhoeck & Ruprecht, Göttingen, West Germany, 1973. LCCN ???? English translation in [Hei83d].

Heisenberg:1973:NRW

[Hei73f] Werner Heisenberg. Naturwissenschaftliche und religiose Wahrheit. (German) [Scientific and religious truth]. Mitteilungen aus der Max-Planck-Gesellschaft, 2(??):73–90, ???? 1973. English translation in [Hei74n].

Heisenberg:1973:SGG

[Hei73g] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1973. 351 pp. LCCN ????

Heisenberg:1973:TSS

[Hei73h] Werner Heisenberg. Tradition in science. Science and public affairs. Bulletin of the Atomic Scientists, 29(10):4–10, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heisenberg:1973:VGF

[Hei73i] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Alexander von Humboldt-Stiftung. Jahresbericht 1972. (German) [Alexander von Humboldt Foundation. Annual report 1972], pages 5–7. Humboldt-Stiftung, Bonn, West Germany, 1973. LCCN ????

Heisenberg:1973:WGN

[Hei73j] Werner Heisenberg. Wandlungen in den Grundlagen der Naturwissenschaft. (German) [Changes in the fundamentals of natural science]. Hirzel, Stuttgart, West Germany, 1973. ISBN 3-7776-0247-7. 183 pp. LCCN ????

Heisenberg:1974:AMS

[Hei74a] Werner Heisenberg. Abstraction in modern science. In Across the Frontiers [Hei74b], pages 70–88. ISBN 0-06-011824-5. LCCN

???? URL http://alsos.wlu.edu/information.aspx?id=941.

English translation of [Hei61c].

Heisenberg:1974:AF

[Hei74b] Werner Heisenberg. Across the Frontiers, volume 48 of World perspectives. Harper & Row, New York, NY, USA, 1974. ISBN 0-06-011824-5. xxii + 229 pp. LCCN ???? URL http://alsos.

wlu.edu/information.aspx?id=941.

Heisenberg:1974:AVH

[Hei74c] Werner Heisenberg. Angst vor dem hemmungslosen Fortschritt (Interview). (German) [Fear of unbridled progress (interview)]. Südwestpresse, ??(??):15, June 29, 1974.

Heisenberg:1974:NHP

[Hei74d] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics]. MI-Compact-Studium. Verlag Moderne Industrie, München, West Germany, 1974. LCCN ????

Heisenberg:1974:DDW

[Hei74e] Werner Heisenberg. Double dialogue with Werner Heisenberg. THEORIA to theory, 8(??):11–33, ???? 1974.

Heisenberg:1974:EP

[Hei74f] Werner Heisenberg. The end of physics? In Across the Frontiers [Hei74b], pages 184–191. ISBN 0-06-011824-5. LCCN ????

URL http://alsos.wlu.edu/information.aspx?id=941. En-

glish translation of [Hei70a].

Heisenberg:1974:FOA

[Hei74g] Werner Heisenberg. Festival ovation for the 800th anniversary of the City of Munich. In Across the Frontiers [Hei74b], pages 47–64. ISBN 0-06-011824-5. LCCN ???? URL http://alsos.wlu.edu/

information.aspx?id=941.

Heisenberg:1974:GVN

[Hei74h] Werner Heisenberg. Goethe’s view of nature and the world of science and technology. In Across the Frontiers [Hei74b], pages 122– 141. ISBN 0-06-011824-5. LCCN ???? URL http://alsos.wlu. edu/information.aspx?id=941. English translation of [Hei67b].

Heisenberg:1974:NCT

[Hei74i] Werner Heisenberg. The notion of a “closed theory” in modern science. In Across the Frontiers [Hei74b], pages 39–46. ISBN 0-06-011824-5. LCCN ???? URL http://alsos.wlu.edu/

information.aspx?id=941.

Heisenberg:1974:PBM

[Hei74j] Werner Heisenberg. The philosophical background of modern physics. Encyclopedia Moderna, 28(9):133–141, ???? 1974.

Heisenberg:1974:PDP

[Hei74k] Werner Heisenberg. Planck’s discovery and the philosophical problems of atomic theory. In Across the Frontiers [Hei74b], pages 8–29. ISBN 0-06-011824-5. LCCN ???? URL http://alsos.

wlu.edu/information.aspx?id=941.

Heisenberg:1974:REP

[Hei74l] Werner Heisenberg. The role of elementary particle physics in the present development of science. Address Delivered to the Royal Swedish Academy of Sciences by Professor Werner Heisenberg on April 24, 1974, 1974. Dokumenta 14 (Royal Swedish Academy Stockholm).

Heisenberg:1974:STC

[Hei74m] Werner Heisenberg. Science and technology in contemporary politics. In Across the Frontiers [Hei74b], pages 64–69. ISBN 0-06-011824-5. LCCN ???? URL http://alsos.wlu.edu/

information.aspx?id=941.

Heisenberg:1974:SRT

[Hei74n] Werner Heisenberg. Scientific and religious truth. In Across the Frontiers [Hei74b], pages 213–229. ISBN 0-06-011824-5. LCCN

???? URL http://alsos.wlu.edu/information.aspx?id=941.

English translation of [Hei73f].

Heisenberg:1974:SWA

[Hei74o] Werner Heisenberg. The scientific work of Albert Einstein. In Across the Frontiers [Hei74b], pages 1–7. ISBN 0-06-011824-5.

LCCN ???? URL http://alsos.wlu.edu/information.aspx?
id=941.

Heisenberg:1974:UFT

[Hei74p] Werner Heisenberg. The unified field theory of elementary particles: Some recent advances. Naturwissenschaften, 61(1):1– 5, ???? 1974. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/g456025757286218.

Heisenberg:1974:VGF

[Hei74q] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Alexander von Humboldt-Stiftung. Jahresbericht 1973. (German) [Alexander von Humboldt Foundation. Annual report 1973], pages 7–8. Humboldt-Stiftung, Bonn, West Germany, 1974. LCCN ????

Heisenberg:1975:DCC

[Hei75a] W. Heisenberg. The development of concepts in 20th-Century physics. Voprosy Filosofii, ??(1):79–88, ???? 1975. ISSN 0042-8744.

Heisenberg:1975:GTR

[Hei75b] W. Heisenberg. The great tradition & its role in problem-selection, methods and concepts in science — end of an epoch. Encounter, 44(3):52–58, ???? 1975. ISSN 0013-7073.

Heisenberg:1975:NST

[Hei75c] Werner Heisenberg. 5.4. Nonlinear spinor theory of elementary particles. In Mehra [Meh75], chapter 12, pages 319–320. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Heisenberg:1975:SN

[Hei75d] Werner Heisenberg. 7. Structure of the nucleus. In Mehra [Meh75], chapter 8, pages 222–226. ISBN 90-277-0635-2. LCCN QC1.S792 M43.

Heisenberg:1975:BGA

[Hei75e] Werner Heisenberg. Begegnungen und Gesprache mit Albert Einstein. (German) [Encounters and conversations with Albert Einstein]. In Tradition in der Wissenschaft. (German) [Tradition in science] [Hei77i], pages 111–125. ISBN 3-492-00454-7. LCCN ???? English translation in [Hei83f].

Heisenberg:1975:BEU

[Hei75f] Werner Heisenberg. Bemerkungen über die Entstehung der Unschärferelation. (German) [Remarks on the origin of Uncertainty Principle]. Physikalische Blätter, 31(5):193–196, May 1975. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/

phbl.19750310501/abstract. English translation in [Hei77g].

Heisenberg:1975:BRU

[Hei75g] Werner Heisenberg. Book review: U. Hoyer, Die Geschichte der Bohrschen Atomtheorie. Physikalische Blätter, 31(1):45, January 1975. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19750310111/abstract.

Heisenberg:1975:CPH

[Hei75h] Werner Heisenberg. Comments to [Patrick A.] Heelans thesis. Journal for General Philosophy of Science = Zeitschrift für allgemeine Wissenschaftstheorie, 6(1):137–138, ???? 1975. CODEN JGPSE4, ZAWTA2. ISSN 0044-2216 (print), 0925-4560 (print),

1572-8587 (electronic). URL http://www.springerlink.com/
content/j86613m83m724j82/.

Heisenberg:1975:CRF

[Hei75i] Werner Heisenberg. Cosmic radiation and fundamental problems in physics. In ????, editor, International Conference on Cosmic Rays. Conference Papers, volume 2, pages 3461–3474. Max-Planck-Institut für Extraterrestrische Physik, Munich, West Germany, 1975. LCCN ???? Reprinted in [Hei76d].

Heisenberg:1975:DCHb

[Hei75j] Werner Heisenberg. Development of concepts in history of quantum theory. Ceskoslovensky Casopis pro Fyziku Sekce A, 25:397– ??, ???? 1975. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Heisenberg:1975:DCP

[Hei75k] Werner Heisenberg. The development of concepts in physics of the twentieth century. In Académie Royale de Belgique, editor, Connaissance Scientifique et Philosophie. Colloque Organisé le 16 et 17 Mai 1973 (Publications du Deuxième Centenaire No. 4), pages 161–177. Palais des Academies, Brussels, Belgium, 1975. LCCN ????

Heisenberg:1975:DCHa

[Hei75l] Werner Heisenberg. Development of concepts in the history of quantum theory. American Journal of Physics, 43(5):389–394, March 1975. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://labs.adsabs.harvard.edu/ui/abs/

1975AmJPh..43..389H. Based on a lecture at Harvard University.

Heisenberg:1975:BZG

[Hei75m] Werner Heisenberg. Die Beziehungen zwischen Gesellschaft und Wissenschaft im Spiegel der Lindauer Tagungen. (German) [The relationship between society and science in the Lindau Meetings]. In A. Dées de Steno, editor, Nobel führte sie zusammen. Begegnungen in Lindau. (German) [Nobel brought them together. Meetings at Lindau], pages 84–99. Belser Presse, Stuttgart, West Germany, 1975. LCCN ????

Heisenberg:1975:DPH

[Hei75n] Werner Heisenberg. Discussion with Professor Heisenberg. In Gingerich [Gin75], pages 556–573. ISBN 0-87474-148-3. LCCN QB36.C8 N37.

Heisenberg:1975:F

[Hei75o] Werner Heisenberg. Foreword. In Mehra [Meh75], pages v–vii. ISBN 90-277-0635-2. LCCN ????

Heisenberg:1975:GRK

[Hei75p] Werner Heisenberg. Gedanken zur ‘Reise der Kunst ins Innere’. (German) [Thoughts on the ‘journey of the art inside’]. In V. Duff and G. v. Molnár, editors, Versuche zu Goethe. Festschrift für Erich Heller zum 65. Geburtstag am 27.3.1976. (German) [Tests of Goethe. Festschrift for Erich Heller’s 65th Birthday on 03/27/1976], pages 321–325. Lothar Stiehm Verlag, Heidelberg, West Germany, 1975. LCCN ???? English translation in [Hei83i].

Heisenberg:1975:GFH

[Hei75q] Werner Heisenberg. Gedenkworte für Hans Kienle. (German) [Memorial words for Hans Kienle]. In ????, editor, Orden Pour le Mérite für Wissenschaften und Künste. Reden und Gedenkworte. (German) [Pour le Mérite for Arts and Sciences. Speeches and commemorative words], volume 12, pages 129–134. Lambert Schneider, Heidelberg, West Germany, 1975. LCCN ????

Heisenberg:1975:GFG

[Hei75r] Werner Heisenberg. Grußwort an die Forschungsstipendiaten. (German) [Greetings to the Research Fellows]. Alexander von Humboldt-Stiftung. Mitteilungen, 31(??):1, ???? 1975.

Heisenberg:1975:ILG

[Hei75s] Werner Heisenberg. Inaugural lecture given at the opening of the 14th International Cosmic Ray Conference: Cosmic Radiation and Fundamental Problems in Physics. In ????, editor, Proceedings from the 14th International Cosmic Ray Conference, held in München, Germany, 15–29 August, 1975. Volume 11 (Invited Lectures and Rapporteur Papers)., pages 3461–?? ????, ????, 1975.

ISBN ???? LCCN ???? URL http://labs.adsabs.harvard.
edu/ui/abs/1975ICRC...11.3461H.

Heisenberg:1975:VGF

[Hei75t] Werner Heisenberg. Vorwort. (German) [Foreword]. In ????, editor, Alexander von Humboldt-Stiftung. Jahresbericht 1974. (German) [Alexander von Humboldt Foundation. Annual report 1974], pages 9–11. Humboldt-Stiftung, Bonn, West Germany, 1975. LCCN ????

Heisenberg:1976:NEPb

[Hei76a] W. Heisenberg. Nature of elementary particles. Naturwissenschaften, 63(1):1–7, ???? 1976. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Heisenberg:1976:NWF

[Hei76b] W. Heisenberg. Nuclear weapon free zones as object of international arms-control diplomacy. Europa Archiv, 31(13):445–452, ???? 1976. ISSN 0014-2476.

Heisenberg:1976:RBS

[Hei76c] W. Heisenberg. Relations between society and science. Universitas-Stuttgart, 18(2):97–104, ???? 1976. ISSN 0041-9079.

Heisenberg:1976:CRF

[Hei76d] Werner Heisenberg. Cosmic radiation and fundamental problems in physics. Naturwissenschaften, 63(2):63–67, February 1976. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904

(electronic). URL http://www.springerlink.com/content/

r680225007732133. Reprint of [Hei75i]. Lecture given at the 14th International Cosmic Ray Conference, August 18, 1975, in Munich (due to the illness of the lecturer, the text was read by Prof. R. Lüst).

Heisenberg:1976:NHP

[Hei76e] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1976. ISBN 3-499-55008-3. 149 pp. LCCN ????

Heisenberg:1976:NEPa

[Hei76f] Werner Heisenberg. The nature of elementary particles. Physics Today, 29(3):32–39, March 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/ link/phtoad/v29/i3/p32/s1. English translation of [Hei76h].

Heisenberg:1976:SGG

[Hei76g] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1976. ISBN 3-492-01995-1. 351 pp. LCCN ????

Heisenberg:1976:WEG

[Hei76h] Werner Heisenberg. Was ist ein elementarteilchen? (German) [What is an elementary particle?]. Naturwissenschaften, 63(1): 1–7, ???? 1976. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/

content/r517163867623308. English translation in [Hei76f].

Heisenberg:1976:TS

[Hei76i] Werner K. Heisenberg. Tradition in science. Ceskoslovensky Casopis pro Fyziku Sekce A, 26(4):403–410, ???? 1976. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Heisenberg:1977:NEP

[Hei77a] W. Heisenberg. Nature of elementary particles. Uspekhi Fizicheskikh Nauk, 121(4):657–668, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Heisenberg:1977:QTI

[Hei77b] W. Heisenberg. Quantum-theoretical interpretation of kinetic and mechanical relations. Uspekhi Fizicheskikh Nauk, 122(4):574–586, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Heisenberg:1977:BGA

[Hei77c] Werner Heisenberg. Begegnungen und Gesprache mit Albert Einstein. (German) [Encounters and conversations with Albert Einstein]. In Tradition in der Wissenschaft. (German) [Tradition in science] [Hei77i], pages 111–125. ISBN 3-492-00454-7. LCCN ???? English translation in [Hei83f].

Heisenberg:1977:CRF

[Hei77d] Werner Heisenberg. Cosmic radiation and fundamental problems in physics. Uspekhi Fizicheskikh Nauk, 121(4):669–677, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://labs.adsabs.harvard.edu/ui/abs/
1977UsFiN.121..669H.

Heisenberg:1977:AQG

[Hei77e] Werner Heisenberg. Die Anfange der Quantenmechanik in Göttingen. (German) [The beginning of quantum mechanics in Göttingen]. In Tradition in der Wissenschaft. (German) [Tradition in science] [Hei77i], pages 43–60. ISBN 3-492-00454-7. LCCN ????

Heisenberg:1977:PPG

[Hei77f] Werner Heisenberg. Pasi peste granite. (Romanian) [Steps Abroad]. Ed. Politica, Bucureşti, Romania, 1977. xxxv + 323 pp. LCCN ????

Heisenberg:1977:ROR

[Hei77g] Werner Heisenberg. Remarks on the origin of the relations of uncertainty. In Price and Chissick [PC77], pages 3–6. ISBN 0-471-99414-6. LCCN QC174.125 .U5. English translation of [Hei75f].

Heisenberg:1977:SGG

[Hei77h] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1977. ISBN 3-492-01995-1. 351 pp. LCCN ????

Heisenberg:1977:TWG

[Hei77i] Werner Heisenberg. Tradition in der Wissenschaft. (German) [Tradition in science]. Piper, Munich, West Germany, 1977. ISBN 3-492-00454-7. 145 pp. LCCN ????

Heisenberg:1977:VCQ

[Hei77j] Werner Heisenberg. Visual contents of quantum-theoretical kinematics and mechanics. Uspekhi Fizicheskikh Nauk, 122(4):651– 671, ???? 1977. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Heisenberg:1978:PPG

[Hei78] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Hirzel, Stuttgart, West Germany, 1978. ISBN 3-7776-0339-2. xxx + 201 pp. LCCN ????

Heisenberg:1979:SMB

[Hei79a] W. Heisenberg. Sense and meaning of beauty in exact sciences. Voprosy Filosofii, ??(12):49–60, ???? 1979. ISSN 0042-8744.

Heisenberg:1979:AHBb

[Hei79b] Werner Heisenberg. Anlage zu [414] (Bisherige und zukunftige Quantenelektrodynamik). Ist eine deterministische Ergänzung der Quantenmechanik möglich?. (German) [Addendum to [414] (present and future quantum electrodynamics). Can quantum mechanics be completed deterministically?]. In Hermann et al. [HvMW79], pages 409–418. ISBN 0-387-08962-4, 3-540-08962-4. LCCN QC16.P37 A34; QC16.P37 W64. URL http://www.

springer.com/physics/book/978-3-540-08962-9. Addendum

to letter of Heisenberg to Pauli of 25 April 1935.

Heisenberg:1979:AHBa

[Hei79c] Werner Heisenberg. Anlage zu Heisenbergs Briefe [407] (Bisherige und zukunftige Quantenelektrodynamik). (German) [Addendum to Heisenberg’s letter [407] (on present and future quantum electrodynamics)]. In Hermann et al. [HvMW79], pages 386– 392. ISBN 0-387-08962-4, 3-540-08962-4. LCCN QC16.P37 A34;

QC16.P37 W64. URL http://www.springer.com/physics/

book/978-3-540-08962-9. Addendum to letter of Heisenberg to Pauli of 25 April 1935.

Heisenberg:1979:NHP

[Hei79d] Werner Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1979. ISBN 3-499-55008-3. 149 pp. LCCN ????

Heisenberg:1979:SNG

[Hei79e] Werner Heisenberg. Das Schöne in der Natur. (German) [The beauty in nature]. In Simon [Sim79], page ?? ISBN 3-7013-0578-1. LCCN ML3800 .M33 1979.

Heisenberg:1979:PPQa

[Hei79f] Werner Heisenberg. Philosophical problems of quantum physics. Ox Bow Press, Woodbridge, CT, USA, 1979. ISBN 0-918024-14-5, 0-918024-15-3 (paperback). 126 pp. LCCN QC6.2 .H44 1979. Reprint of [Hei52k].

Heisenberg:1979:QPG

[Hei79g] Werner Heisenberg. Quantentheorie und Philosophie. (German) [Quantum theory and philosophy]. Reclam-Verlag, Stuttgart, West Germany, 1979. ISBN 3-15-009948-X. 125 pp. LCCN ????

Heisenberg:1980:PLU

[Hei80] Elisabeth Heisenberg. Das politische Leben eines Unpolitischen: Erinnerungen an Werner Heisenberg. (German) [The political life of an unpolitical: memories of Werner Heisenberg]. Piper, München, West Germany, 1980. ISBN 3-492-02552-8. 201 + 10 pp. LCCN ????

Heisenberg:1981:DNW

[Hei81a] W. Heisenberg. The dangers of nuclear war. Journal of Strategic Studies, 4(3):326–328, ???? 1981. ISSN 0140-2390 (print), 1743-937x (electronic).

Heisenberg:1981:TGGa

[Hei81b] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1981. 288 pp. LCCN ????

Heisenberg:1981:TGGb

[Hei81c] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Piper, Munich, West Germany, 1981. ISBN 3-492-02212-X. 333 pp. LCCN ????

Heisenberg:1982:PP

[Hei82a] Werner Heisenberg. Physics and philosophy. In Faculty of Lynchburg College [Fac82], page ?? ISBN 0-8191-2295-5, 0-8191-2252-1 (paperback). LCCN Q125 .M316 1982.

Heisenberg:1982:SGA

[Hei82b] Werner Heisenberg. Selbstbiographie. (German) [Autobiography]. Nuova Acta Leopoldina (NF), 55(248):11–16, ???? 1982.

Heisenberg:1982:VWE

[Hei82c] Werner Heisenberg. Veranderung des Wirklichkeitsbegriffs der exakten Naturwissenschaft (Abstrakt). (German) [Alteration of the concept of reality of the exact science (abstract)]. In ????, editor, Die Mittwochs-Gesellschaft: Protokolle aus dem geistigen Deutschland 1932 bis 1944. (German) [The Wednesday Society: Minutes from the intellectual Germany from 1932 to 1944], pages 332–333. Severin und Siedler, Berlin, West Germany, 1982. LCCN ????

Heisenberg:1983:PLU

[Hei83a] Elisabeth Heisenberg. Das politische Leben eines Unpolitischen: Erinnerungen an Werner Heisenberg. (German) [The political life of an unpolitical: memories of Werner Heisenberg], volume 279 of Serie Piper. Piper, München, West Germany, second edition, 1983. ISBN 3-492-00579-9. 201 pp. LCCN ????

Heisenberg:1983:ACQ

[Hei83b] Werner Heisenberg. The actual content of quantum theoretical kinematics and mechanics. Zhurnal Physik, 43(??):172–198, De-

cember 1983. URL http://labs.adsabs.harvard.edu/ui/abs/

1983ZhPhy..43..172H. English translation of [Hei27e].

Heisenberg:1983:BQM

[Hei83c] Werner Heisenberg. The beginnings of quantum mechanics in Göttingen. In Tradition in Science [Hei83j], pages 37–55. ISBN 0-8164-2488-8. LCCN ???? All German chapters translated by Peter Heath.

Heisenberg:1983:CCC

[Hei83d] Werner Heisenberg. The correctness criteria for ‘closed theories’ in physics. In ????, editor, Tradition in Science, pages 123–129. Seabury Press, New York, NY, USA, 1983. LCCN ???? English translation of [Hei73e].

Heisenberg:1983:TGG

[Hei83e] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1983. ISBN 3-423-00903-9. 287 pp. LCCN ????

Heisenberg:1983:ECA

[Hei83f] Werner Heisenberg. Encounters and conversations with Albert Einstein. In Tradition in Science [Hei83j], pages 107–122. ISBN 0-8164-2488-8. LCCN Q175 .H3932 1981. English translation of [Hei77c].

Heisenberg:1983:PMV

[Hei83g] Werner Heisenberg. The production of mesons in very high energy collisions. In Anonymous, editor, Celebrazioni del Trentennale della Scuola Internazionale di Fisica ≪Enrico Fermi≫, pages 96– 103. Editrice Compositori e Società Italiana di Fisica, Bologna, Italia, 1983. ISBN ???? LCCN ???? Posthumous republication of lecture given in 1954 at the II Course of the International School of Physics at Villa Monastero, Varenna, Italy. Also available in [Hei55e].

Heisenberg:1983:REB

[Hei83h] Werner Heisenberg. A rész és az egész: beszélgetések az atomfizikáról. Gondolat, Budapest, third edition, 1983. ISBN 963-281-292-1. 330 pp. LCCN ????

Heisenberg:1983:TAJ

[Hei83i] Werner Heisenberg. Thoughts on ‘The artist’s journey into the interior. In Tradition in Science [Hei83j], pages 130–135. ISBN 0-8164-2488-8. LCCN ???? English translation of [Hei75p].

Heisenberg:1983:TS

[Hei83j] Werner Heisenberg. Tradition in Science. Seabury Press, New York, NY, USA, 1983. ISBN 0-8164-2488-8. 141 pp. LCCN Q175 .H3932 1981. All German chapters translated by Peter Heath.

Heisenberg:1984:IER

[Hei84a] Elisabeth Heisenberg. Inner exile: recollections of a life with Werner Heisenberg. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1984. ISBN 0-8176-3146-1. xvii + 170 + 20 pp. LCCN QC16.H35 H4413 1984.

Heisenberg:1984:CTG

[Hei84b] M. W. Heisenberg. Considérations théoriques générales sur la structure du noyau. (French) [Theoretical general considerations on the structure of the nucleus]. In Blum et al. [BDR84e], pages 179–225. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book).

LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-3-642-61742-3_14/.

Heisenberg:1984:CDS

[Hei84c] Professor W. Heisenberg. Contribution to discussion on the structure of simple molecules. In Blum et al. [BDR84e], pages 171– 172. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-61742-3_12?coverImageUrl=/static/0.
363/sites/link/images/abstract_cover_placeholder.png.

Heisenberg:1984:AUF

[Hei84d] W. Heisenberg. Allgemeiner Überblick und Fortschritte in der Theorie der Elementarteilchen. (German) [General overview and progress in the theory of elementary particles]. In Blum et al. [BDR84e], page 571. ISBN 3-642-61743-3 (print), 3-642-61742-

5 (e-book). LCCN ???? URL http://link.springer.com/
accesspage/chapter/10.1007/978-3-642-61742-3_49?coverImageUrl=
/static/0.363/sites/link/images/abstract_cover_placeholder.
png.

Heisenberg:1984:BNT

[Hei84e] W. Heisenberg. Bemerkungen zur Neuen Tamm–Dancoff-Methode. (German) [Comments on the new Tamm–Dancoff method]. In Blum et al. [BDR84e], pages 901–903. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/chapter/10.1007/978-3-642-61742-3_
66/.

Heisenberg:1984:BTA

[Hei84f] W. Heisenberg. Bemerkungen zur Theorie des Atomkerns. (German) [Comments on the theory of the atomic nucleus]. In Blum et al. [BDR84e], pages 238–246. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_16/.

Heisenberg:1984:BAEb

[Hei84g] W. Heisenberg. Bericht über die allgemeinen Eigenschaften der Elementarteilchen. (German) [Report on the general properties of elementary particles]. In Blum et al. [BDR84e], pages 346–

358. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_29/.

Heisenberg:1984:CRF

[Hei84h] W. Heisenberg. Cosmic radiation and fundamental problems in physics. In Blum et al. [BDR84e], pages 928–932. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_70/.

Heisenberg:1984:SEM

[Hei84i] W. Heisenberg. Das schwere Elektron (Mesotron) und seine Rolle in der Höhenstrahlung. (German) [The heavy electron (mesotron) and its role in cosmic radiation]. In Blum et al. [BDR84e], pages 331–333. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book).

LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-3-642-61742-3_24/.

Heisenberg:1984:AHS

[Hei84j] W. Heisenberg. De atoomkern en hare samenstelling. (Dutch) [The atomic nucleus and its composition]. In Blum et al. [BDR84e], pages 335–344. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_27/.

Heisenberg:1984:DSN

[Hei84k] W. Heisenberg. Der derzeitige Stand der nichtlinearen Spinortheorie der Elementarteilchen. (German) [The current state of nonlinear spinor theory of elementary particles]. In Blum et al. [BDR84e], pages 580–591. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_51/.

Heisenberg:1984:DSE

[Hei84l] W. Heisenberg. Der Durchgang sehr energiereicher Korpuskeln durch den Atomkern. (German) [The passage of highly energetic corpuscles through the atomic nucleus]. In Blum et al. [BDR84e], pages 256–259. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_20/.

Heisenberg:1984:GST

[Hei84m] W. Heisenberg. Der gegenwärtige Stand der Theorie der Elementarteilchen. (German) [The current state of the theory of elementary particles]. In Blum et al. [BDR84e], pages 547– 551. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_45/.

Heisenberg:1984:DKH

[Hei84n] W. Heisenberg. Die durchdringende Komponente der Höhenstrahlung. (German) [The penetrating component of the cosmic rays]. In Blum et al. [BDR84e], page 359. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_30?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:EEF

[Hei84o] W. Heisenberg. Die Entwicklung der einheitlichen Feldtheorie der Elementarteilchen. (German) [The development of the unified field theory of elementary particles]. In Blum et al. [BDR84e], pages 605–609. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_55/.

Heisenberg:1984:EMV

[Hei84p] W. Heisenberg. Die Erzeugung von Mesonen in Vielfachprozessen. (German) [The production of mesons in multiple processes]. In Blum et al. [BDR84e], pages 457–467. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_36/.

Heisenberg:1984:ENL

[Hei84q] W. Heisenberg. Electrodynamics in the non-linear spinor theory. In Blum et al. [BDR84e], pages 598–603. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_53/.

Heisenberg:1984:ENS

[Hei84r] W. Heisenberg. Elektrodynamik in der nichtlinearen Spinortheorie. (German) [electrodynamics in the non-linear spinor theory]. In Blum et al. [BDR84e], page 604. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_54?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:FTF

[Hei84s] W. Heisenberg. Fortschritte in der Theorie des Ferromagnetismus. (German) [Advances in the theory of ferromagnetism]. In Blum et al. [BDR84e], pages 167–170. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_11/.

Heisenberg:1984:GRC

[Hei84t] W. Heisenberg. General remarks concerning the theory of elementary particles. In Blum et al. [BDR84e], page 562. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-

642-61742-3_47?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:IMS

[Hei84u] W. Heisenberg. Indefinite metric in state space. In Blum et al. [BDR84e], pages 904–911. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_67/.

Heisenberg:1984:IUFb

[Hei84v] W. Heisenberg. Introduction to the unified field theory of elementary particles. In Blum et al. [BDR84e], pages 677–861. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_62/.

Heisenberg:1984:NLD

[Hei84w] W. Heisenberg. Nichtlaminare Lösungen der Differentialgleichungen für reibende Flüssigkeiten. (German) [Non-laminar solutions of differential equations for grinding fluids]. In Blum et al. [BDR84e], pages 23–27. ISBN 3-642-61743-3 (print), 3-642-61742-

5 (e-book). LCCN ???? URL http://link.springer.com/
chapter/10.1007/978-3-642-61742-3_3/.

Heisenberg:1984:SLF

[Hei84x] W. Heisenberg. On the stability of laminar flow. In Blum et al. [BDR84e], pages 471–475. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_38/.

Heisenberg:1984:TES

[Hei84y] W. Heisenberg. On the theory of explosion showers in cosmic rays. In Blum et al. [BDR84e], page 345. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_28?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:PCS

[Hei84z] W. Heisenberg. Particles as collective stationary states. In Blum et al. [BDR84e], pages 637–642. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_59/.

Heisenberg:1984:PST

[Hei84-27] W. Heisenberg. The present situation in the theory of elementary particles. In Blum et al. [BDR84e], pages 610–630. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_56/.

Heisenberg:1984:PMV

[Hei84-28] W. Heisenberg. The production of mesons in very high energy collisions. In Blum et al. [BDR84e], pages 527–534. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_43/.

Heisenberg:1984:QNL

[Hei84-29] W. Heisenberg. Quantization of non-linear wave equations. In Blum et al. [BDR84e], pages 524–526. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_42/.

Heisenberg:1984:QEN

[Hei84-30] W. Heisenberg. Quantum electrodynamics in the nonlinear spinor theory. In Blum et al. [BDR84e], pages 658–676. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_61/.

Heisenberg:1984:QTF

[Hei84-31] W. Heisenberg. Quantum theory of fields and elementary particles. In Blum et al. [BDR84e], pages 552–561. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_46/.

Heisenberg:1984:RRN

[Hei84-32] W. Heisenberg. Recent research on the nonlinear spinor theory of elementary particles. In Blum et al. [BDR84e], pages 572– 579. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_50/.

Heisenberg:1984:RFT

[Hei84-33] W. Heisenberg. Remarks on the ^{≪}Spurion^{≫}-theory of strange particles. In Blum et al. [BDR84e], pages 631–633. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/chapter/10.1007/978-3-642-61742-3_
57/.

Heisenberg:1984:RPS

[Hei84-34] W. Heisenberg. Report on the present situation in the non-linear spinor theory of elementary particles. In Blum et al. [BDR84e], pages 875–890. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_64/.

Heisenberg:1984:RGT

[Hei84-35] W. Heisenberg. Research in Germany on the technical application of atomic energy. In Blum et al. [BDR84e], pages 414– 418. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_33/.

Heisenberg:1984:RNL

[Hei84-36] W. Heisenberg. Research on the non-linear spinor theory with indefinite metric in Hilbert space. In Blum et al. [BDR84e], pages 563–570. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_48/.

Heisenberg:1984:RRP

[Hei84-37] W. Heisenberg. Review of recent progress. In Blum et al. [BDR84e], pages 891–900. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_65/.

Heisenberg:1984:SPS

[Hei84-38] W. Heisenberg. Survey on the present situation concerning unified theories of elementary particles. In Blum et al. [BDR84e], pages 643–657. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_60/.

Heisenberg:1984:TEP

[Hei84-39] W. Heisenberg. Theory of elementary particles. In Blum et al. [BDR84e], pages 535–546. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_44/.

Heisenberg:1984:TL

[Hei84-40] W. Heisenberg. Two lectures. In Blum et al. [BDR84e], pages 443– 455. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_35/.

Heisenberg:1984:SRM

[Hei84-41] W. Heisenberg. Über die Streuung von Röntgenstrahlen an Molekülen und Kristallen. (German) [On the scattering of X-rays by molecules and crystals]. In Blum et al. [BDR84e], pages 173– 178. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_13/.

Heisenberg:1984:QKM

[Hei84-42] W. Heisenberg. Über quantentheoretische Kinematik und Mechanik. (German) [On quantum theoretical kinematics and mechanics]. In Blum et al. [BDR84e], pages 29–51. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/chapter/10.1007/978-3-642-61742-3_
5/.

Heisenberg:1984:UFT

[Hei84-43] W. Heisenberg. The unified field theory of elementary particles: Some recent advances. In Blum et al. [BDR84e], pages 912– 916. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_68/.

Heisenberg:1984:VKS

[Hei84-44] W. Heisenberg. Verwandlung der kosmischen Strahlung in der Atmosphäre. (German) [Transformation of cosmic rays in the atmosphere]. In Blum et al. [BDR84e], page 334. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_26?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:WQW

[Hei84-45] W. Heisenberg. Wahrscheinlichkeitsaussagen in der Quantentheorie der Wellenfelder. (German) [Probability statements in the quantum theory of wave fields]. In Blum et al. [BDR84e], pages 249–255. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book).

LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-3-642-61742-3_19/.

Heisenberg:1984:YTN

[Hei84-46] W. Heisenberg. The Yukawa theory of nuclear forces in the light of present quantum theory of wave fields. In Blum et al. [BDR84e], pages 468–470. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-

book). LCCN ???? URL http://link.springer.com/chapter/
10.1007/978-3-642-61742-3_37/.

Heisenberg:1984:QF

[Hei84-47] W. Heisenberg. Zur Quantentheorie des Ferromagnetismus. (German) [On the quantum theory of ferromagnetism]. In Blum et al. [BDR84e], pages 100–108. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_8/.

Heisenberg:1984:BAEa

[Hei84-48] Werner Heisenberg. Bericht über die allgemeinen Eigenschaften der Elementarteilchen. II. Die Wechselwirkung der Elementarteilchen. III. Die Grenzen der bisherighen Theorie. (German) [Report on the general properties of elementary particles. II. The interaction of elementary particles. III. The limits of current theory]. In Blum et al. [BDR84b], pages 346–358. ISBN 3-540-13020-9, 0-387-13020-9. LCCN QC3 .H33 1984. Parts II and III were to be presented at the 8th Solvay Conference in Brussels, 22–29 October 1939. The conference was canceled because of the outbreak of World War II in September 1939, and the report was not published until the Collected Works were prepared. Wolfgang Pauli prepared Part I of this joint work; it was published in revised form in [Pau41].

Heisenberg:1984:PWG

[Hei84-49] Werner Heisenberg. Das physikalische Weltbild (1942). (German) [The physical world]. In Blum et al. [BDR84c], pages 193–201. ISBN 3-492-02925-6. LCCN ????

Heisenberg:1984:TGG

[Hei84-50] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1984. ISBN 3-423-00903-9. 287 pp. LCCN ????

Heisenberg:1984:DQM

[Hei84-51] Werner Heisenberg. The development of quantum mechanics. In Blum et al. [BDR84e], pages 226–237. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_15/.

Heisenberg:1984:OWG

[Hei84-52] Werner Heisenberg. [Die] Ordnung der Wirklichkeit (1942). (German) [The order of reality]. In Blum et al. [BDR84c], pages 217– 306. ISBN 3-492-02925-6. LCCN ????

Heisenberg:1984:DHQ

[Hei84-53] Werner Heisenberg. Doubts and hopes in quantum-electrodynamics. In Blum et al. [BDR84e], pages 509–520. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.

springer.com/chapter/10.1007/978-3-642-61742-3_40/.

Heisenberg:1984:EPM

[Hei84-54] Werner Heisenberg. Erkenntnistheoretische Probleme in der modernen Physik (1939). (German) [Epistemological problems in modern physics (1939)]. In Blum et al. [BDR84c], pages 22–28. ISBN 3-492-02925-6. LCCN ????

Heisenberg:1984:FOI

[Hei84-55] Werner Heisenberg. Fisica e oltre: incontri con i protagonisti: 1920–1965. (Italian) [Physics and Beyond: meetings with the protagonists: 1920–1965]. Prove: saggi scientifici. Boringhieri, Torino, Italia, 1984. 261 pp. LCCN ????

Heisenberg:1984:FPT

[Hei84-56] Werner Heisenberg. Fundamental problems of theoretical physics (1959). In Blum et al. [BDR84d], pages 256–262. ISBN 3-492-02926-4. LCCN ????

Heisenberg:1984:IUFa

[Hei84-57] Werner Heisenberg. Introduction to the unified field theory of elementary particles. In Blum et al. [BDR84b], pages 677–861. ISBN 3-540-13020-9, 0-387-13020-9. LCCN QC3 .H33 1984. Reprint of [Hei66d].

Heisenberg:1984:NEP

[Hei84-58] Werner Heisenberg. The nature of elementary particles. In Blum et al. [BDR84e], pages 917–927. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_69/.

Heisenberg:1984:PPT

[Hei84-59] Werner Heisenberg. Philosophische Probleme in der Theorie der Elementarteilchen (1967). (German) [Philosophical problems in the theory of elementary particles (1967)]. In Blum et al. [BDR84d], pages 410–422. ISBN 3-492-02926-4. LCCN ????

Heisenberg:1984:PPG

[Hei84-60] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Ullstein, Frankfurt/M, West Germany, 1984. ISBN 3-548-35132-8. 196 pp. LCCN ????

Heisenberg:1984:RTN

[Hei84-61] Werner Heisenberg. Remarques sur la théorie neutrinienne de la lumière. (French) [Remarks on the theory of light]. In Blum et al. [BDR84e], pages 521–523. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_41/.

Heisenberg:1984:SGG

[Hei84-62] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1984. ISBN 3-492-00636-1. 318 pp. LCCN ????

Heisenberg:1984:BWN

[Hei84-63] Werner Heisenberg. Über das Weltbild der Naturwissenschaft (1942). (German) [On the world of science]. In Blum et al. [BDR84c], pages 207–215. ISBN 3-492-02925-6. LCCN ????

Heisenberg:1985:SCS

[Hei85a] W. Heisenberg. Strategic and crisis stability. Bulletin of Peace Proposals, 16(3):225–231, ???? 1985. ISSN 0007-5035.

Heisenberg:1985:BZQ

[Hei85b] Werner Heisenberg. Bisherige und zukünftige Quantenelektrodynamik (25 April 1935). (German) [Past and future quantum electrodynamics (April 25 1935)]. In von Meyenn et al. [vMHW85], pages 386–392. ISBN 0-387-13609-6 (New York), 3-540-13609-6

(Berlin). LCCN ???? URL http://www.springerlink.com/
content/978-3-540-13609-5.

Heisenberg:1985:TGG

[Hei85c] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1985. ISBN 3-423-00903-9. 287 pp. LCCN ????

Heisenberg:1985:DEQ

[Hei85d] Werner Heisenberg. Ist eine deterministische Erganzung der Quantenmechanik möglich? (Sommer 1935). (German) [Is a complement of deterministic quantum mechanics possible? (summer 1935)]. In von Meyenn et al. [vMHW85], pages 409–418. ISBN 0-387-13609-6 (New York), 3-540-13609-6 (Berlin). LCCN

???? URL http://www.springerlink.com/content/978-3-
540-13609-5.

Heisenberg:1985:R

[Hei85e] Werner Heisenberg. Reminiscences from 1926 and 1927. In French and Kennedy [FK85], pages 163–171. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Heisenberg:1986:RAD

[Hei86a] W. Heisenberg. The reception of American deterrence theory in the Federal Republic of Germany and the German nuclear debate of the 1950s. Journal of Strategic Studies, 9(4):70–83, ???? 1986. ISSN 0140-2390 (print), 1743-937X (electronic).

Heisenberg:1986:ABK

[Hei86b] Werner Heisenberg. Ansprache bei der Beisetzung von K.F. Bonhoeffer (1957). (German) [Speech at the funeral of K. F. Bonhoeffer (1957)]. In Blum et al. [BDR86], pages 97–98. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:AVP

[Hei86c] Werner Heisenberg. Antrittsrede vor der Preußischen Akademie der Wissenschaften (1943). (German) [Inaugural address to the Prussian Academy of Sciences (1943)]. In Blum et al. [BDR86], page 16. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:CVF

[Hei86d] Werner Heisenberg. Curriculum vitae (finished 1975). In Blum et al. [BDR86], pages 35–36. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:TGG

[Hei86e] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Piper, Munich, West Germany, 1986. ISBN 3-423-00903-9. 336 pp. LCCN ????

Heisenberg:1986:EEA

[Hei86f] Werner Heisenberg. Erinnerung an die Entwicklung der Atomphysik in den letzten 50 Jahren (1968). In Blum et al. [BDR86], pages 22–34. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:GPN

[Hei86g] Werner Heisenberg. Glückwunschschreiben an Professor Niels Bohr (1961). (German) [Congratulations to Professor Niels Bohr (1961)]. In Blum et al. [BDR86], pages 129–130. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:LGC

[Hei86h] Werner Heisenberg. Lebenslauf (1943). (German) [Curriculum vitae]. In Blum et al. [BDR86], pages 14–15. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:MPQ

[Hei86i] Werner Heisenberg. Max Planck und die Quantentheorie (1943). (German) [Max Planck and quantum theory (1943)]. In Blum et al. [BDR86], pages 65–67. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:ML

[Hei86j] Werner Heisenberg. Max von Laue zum 9.10.1949. (German) [Max von Laue on 1949.10.9]. In Blum et al. [BDR86], pages 81–82. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:WH

[Hei86k] Werner Heisenberg. Werner Heisenberg (1958). In Blum et al. [BDR86], page 18. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:WBR

[Hei86l] Werner Heisenberg. Worte an der Bahre Richard Beckers (1955). (German) [Words at the grave of Richard Becker (1955)]. In Blum et al. [BDR86], pages 89–90. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:WKG

[Hei86m] Werner Heisenberg. Worte bei der Kranzniederlegung am Grabe Arnold Sommerfelds (1951). (German) [Words at the wreathlaying ceremony at the grave of Arnold Sommerfeld (1951)]. In Blum et al. [BDR86], page 83. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:GNK

[Hei86n] Werner Heisenberg. Zum 500 Geburtstag von Nikolaus Kopernikus (1972). (German) [On the 500th birthday of Nicholas Copernicus (1972)]. In Blum et al. [BDR86], page 201. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1986:EEI

[Hei86o] Werner Heisenberg. Zur Einweihung der Einsteinbüste im Deutschen Museum (1972). (German) [On the inauguration of the bust of Einstein in the German Museum (1972)]. In Blum et al. [BDR86], pages 195–197. ISBN 3-492-02928-0. LCCN ????

Heisenberg:1987:ATG

[Hei87a] Werner Heisenberg. Ausgewählte Texte. (German) [Selected works]. Goldmann, Munich, West Germany, 1987. ISBN 3-442-11021-1. 127 pp. LCCN ????

Heisenberg:1987:TGG

[Hei87b] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1987. ISBN 3-423-00903-9. 287 pp. LCCN ????

Heisenberg:1987:QPG

[Hei87c] Werner Heisenberg. Quantentheorie und Philosophie. (German) [Quantum theory and philosophy], volume 9948 of Universal-Bibliothek [Universal Library]. Reclam-Verlag, Stuttgart, West Germany, 1987. ISBN 3-15-009948-X. 126 pp. Vorlesungen und Aufsätze. [Lectures and essays], With an afterword by Jürgen Busche.

Heisenberg:1988:TGG

[Hei88] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Deutsche Taschenbuchverlag, Munich, West Germany, 1988. ISBN 3-423-00903-9. 287 pp. LCCN ????

Heisenberg:1989:AAP

[Hei89a] Werner Heisenberg. Ansprache anläßlich der Planck-Feier (1948). (German) [Speech on the occasion of the celebration of Planck (1948)]. In Blum et al. [BDR89b], pages 592–596. ISBN 3-492-

02929-9. LCCN ???? URL http://d-nb.info/900168226/04.

Heisenberg:1989:AAO

[Hei89b] Werner Heisenberg. Ansprache aus Anlaß der offiziellen Namensverleihung und Übergabe des Erweiterungsbaus an das Werner-Heisenberg-Gymnasium Göppingen (1975). (German) [Speech on the occasion of the official naming and handing-over ceremony of the extension at the Werner-Heisenberg-Gymnasium Göppingen (1975)]. In Blum et al. [BDR89b], pages 448–452. ISBN

3-492-02929-9. LCCN ???? URL http://d-nb.info/900168226/
04.

Heisenberg:1989:AIF

[Hei89c] Werner Heisenberg. Atomtechnik im Frieden (1951). (German) [Nuclear technology in peace (1951)]. In Blum et al. [BDR89b], pages 128–130. ISBN 3-492-02929-9. LCCN ???? URL http://

d-nb.info/900168226/04.

Heisenberg:1989:AFG

[Hei89d] Werner Heisenberg. Atomtechnik in Frieden (1950). (German) [Nuclear technology in peace (1950)]. In Blum et al. [BDR89b], pages 125–127. ISBN 3-492-02929-9. LCCN ???? URL http://

d-nb.info/900168226/04.

Heisenberg:1989:AGV

[Hei89e] Werner Heisenberg. Auswertung des Gottower Versuches (1943/ 1944). (German) [Evaluation of Gottower experiment (1943/ 1944)]. In Blum et al. [BDR89a], pages 576–587. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:BGA

[Hei89f] Werner Heisenberg. Begrüßungsansprache (Göttinger Akademie, 1951). (German) [Welcome address (Göttingen Academy, 1951)]. In Blum et al. [BDR89b], pages 412–413. ISBN 3-492-02929-9.

LCCN ???? URL http://d-nb.info/900168226/04.

Heisenberg:1989:BGH

[Hei89g] Werner Heisenberg. Bemerkungen zu dem geplanten halbtechnischen Versuch mit 1,5 to D2O und 3 to 38-Metall (31 Jul. 1942). (German) [Comments on the proposed pilot plant test with 1.5 tonnes D_{2}O and 3 tonnes 38-metal (31 July 1942)]. In Blum et al. [BDR89a], pages 545–552. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:BBM

[Hei89h] Werner Heisenberg. Bericht über die Möglichkeit technischer Energiegewinnung aus der Uranspaltung (II) (29 Feb 1940). (German) [Report on the technical possibility of energy from fission of uranium (II) (29 February 1940)]. In Blum et al. [BDR89a], pages 397–418. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:BBVa

[Hei89i] Werner Heisenberg. Bericht über die Versuche mit Schichtenanordnungen von Präparat 38 und Paraffin am Kaiser Wilhelm Institut für Physik in Berlin-Dahlem (Mar 41). (German) [Report on experiments with layer arrangements of preparation 38 and paraffin at the Kaiser Wilhelm Institute for Physics in Berlin Dahlem (March 41)]. In Blum et al. [BDR89a], pages 432– 462. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:SGS

[Hei89j] Werner Heisenberg. Das sogenannte Gleichgewicht des Schreckens (1969). (German) [The so-called balance of terror]. In Blum et al. [BDR89b], pages 557–559. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Heisenberg:1989:BMH

[Hei89k] Werner Heisenberg. Die Behandlung von Mehrkörperproblemen mit Hilfe der η-Matrix. (German) [The treatment of many-body systems by the η-matrix]. In Blum et al. [BDR89a], pages 687–698. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Unpublished paper, probably from 1944, subtitled ‘Die beobachtbaren Größen in der Theorie der Elementarteilchen. IV’ [The “observable variables” in the theory of elementary particles. IV].

Heisenberg:1989:EAM

[Hei89l] Werner Heisenberg. Die Energiegewinnung aus der Atomkernspaltung (5 Mai 1943). (German) [The energy production from nuclear fission (5 may 1943)]. In Blum et al. [BDR89a], pages 570–575. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:MTE

[Hei89m] Werner Heisenberg. Die Möglichkeit der technischen Energiegewunnung der Uranspaltung (6 Dec 1939). (German) [The possibility of technical energy extraction from uranium fission (6 December 1939)]. In Blum et al. [BDR89a], pages 378–396. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:TGF

[Hei89n] Werner Heisenberg. Die theoretischen Grundlagen für die Energie gewinnung aus der Uranspaltung (26 Feb. 1942). (German) [The theoretical basis for energy recovery from the fission of uranium (26 February 1942)]. In Blum et al. [BDR89a], pages 517– 521. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:EEO

[Hei89o] Werner Heisenberg. Encounters with Einstein: And other essays on people, places, and particles. Princeton Science Library. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-02433-2. vi + 141 pp. Reprint of the 1983 original, With an epilogue by Hans-Peter Dürr.

Heisenberg:1989:ECA

[Hei89p] Werner Heisenberg. European cooperation for the advancement of nuclear research (1953). In Blum et al. [BDR89b], pages 292– 293. ISBN 3-492-02929-9. LCCN ???? URL http://d-nb.info/

900168226/04.

Heisenberg:1989:FGA

[Hei89q] Werner Heisenberg. Festrede (Göttinger Akademie, 1951). (German) [Speech (Göttingen Academy, 1951)]. In Blum et al. [BDR89b], pages 414–419. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Heisenberg:1989:ICV

[Hei89r] Werner Heisenberg. Interview (on the Copenhagen visit in 1941; 1965). In Blum et al. [BDR89b], page 44. ISBN 3-492-02929-9.

LCCN ???? URL http://d-nb.info/900168226/04.

Heisenberg:1989:IHR

[Hei89s] Werner Heisenberg. Interview von Heman Ray (1949). (German) [Interview of Heman Rey]. In Blum et al. [BDR89b], pages 537– 538. ISBN 3-492-02929-9. LCCN ???? URL http://d-nb.info/

900168226/04.

Heisenberg:1989:IPD

[Hei89t] Werner Heisenberg. Interview with Prof. Dr. Werner Heisenberg (Prag; 1968). In Blum et al. [BDR89b], pages 618–620. ISBN 3-

492-02929-9. LCCN ???? URL http://d-nb.info/900168226/
04.

Heisenberg:1989:KBF

[Hei89u] Werner Heisenberg. Kurze Begrüßungsansprache bei der Feierstunde (Humboldt-Stiftung, 1974). (German) [Short welcome speech at the ceremony (Humboldt Foundation, 1974)]. In Blum et al. [BDR89b], pages 368–369. ISBN 3-492-02929-9. LCCN ????

URL http://d-nb.info/900168226/04.

Heisenberg:1989:MPP

[Hei89v] Werner Heisenberg. Modern physics and its philosophical consequences (1957). In Blum et al. [BDR89b], pages 598–607. ISBN

3-492-02929-9. LCCN ???? URL http://d-nb.info/900168226/
04.

Heisenberg:1989:MTA

[Hei89w] Werner Heisenberg. Möglichkeiten zur technischen Ausnutzung der thermonuklearen Reaktionen (1957). (German) [Possibilities for utilization of the technical thermonuclear reactions (1957)]. In Blum et al. [BDR89b], pages 192–194. ISBN 3-492-02929-9. LCCN

???? URL http://d-nb.info/900168226/04.

Heisenberg:1989:NBB

[Hei89x] Werner Heisenberg. Nachtrag zum Bericht: über die Möglichkeit der Energieerzeugung mit Hilfe des Isotops 238 (Nov. 1941). (German) [Addendum to report: On the possibility of energy production with the help of isotope 238 (November 1941)]. In Blum et al. [BDR89a], pages 510–513. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:NFT

[Hei89y] Werner Heisenberg. Neujahrsansprache für den Tschechoslowakischen Rundfunk (1968). (German) [New Year’s address on the Czechoslovak Radio (1968)]. In Blum et al. [BDR89b], page 554. ISBN 3-492-02929-9. LCCN ???? URL http://d-nb.info/

900168226/04.

Heisenberg:1989:OWG

[Hei89z] Werner Heisenberg. Ordnung der Wirklichkeit. (German) [Order of reality]. Piper, Munich, West Germany, 1989. ISBN 3-492-10945-4. 175 pp. LCCN ????

Heisenberg:1989:RVE

[Hei89-27] Werner Heisenberg. Rede bei der Verleihung des Ehrendoktorgrades der Universität Zagreb (1969). (German) [Speech at the presentation of the honorary doctorate at the University of Zagreb (1969)]. In Blum et al. [BDR89b], pages 427–432. ISBN 3-492-

02929-9. LCCN ???? URL http://d-nb.info/900168226/04.

Original publication (in Croatian) in Encyclopedis Moderna 11, 57–61.

Heisenberg:1989:RFE

[Hei89-28] Werner Heisenberg. Rede für den Empfang bei der Alexander von Humboldt-Stiftung (1974). (German) [Speech for the reception at the Alexander von Humboldt Foundation (1974)]. In Blum et al. [BDR89b], pages 370–374. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Heisenberg:1989:SGG

[Hei89-29] Werner Heisenberg. Schritte über Grenzen. (German) [Steps across borders]. Piper, Munich, West Germany, 1989. ISBN 3-492-10336-7. 318 pp. LCCN ????

Heisenberg:1989:SBB

[Hei89-30] Werner Heisenberg. Schule und Bildung (Brief; 1953). (German) [Schools and education (letter; 1953)]. In Blum et al. [BDR89b], page 420. ISBN 3-492-02929-9. LCCN ???? URL http://d-nb.

info/900168226/04.

Heisenberg:1989:SNN

[Hei89-31] Werner Heisenberg. Statement for NBC News (1971). In Blum et al. [BDR89b], page 494. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Heisenberg:1989:SFV

[Hei89-32] Werner Heisenberg. Stellungnahme zur Frage des Verbots der Kernwaffenversuche (1960). (German) [Opinion on the question of the prohibition of nuclear weapons tests (1960)]. In Blum et al. [BDR89b], page 545. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Heisenberg:1989:TAD

[Hei89-33] Werner Heisenberg. Theoretische Auswertung der Dahlemer Großversuche (Sept. 1944). (German) [Theoretical analysis of the Dahlem-scale tests (September 1944)]. In Blum et al. [BDR89a], pages 588–594. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:TNB

[Hei89-34] Werner Heisenberg. Theoretischer Nachtrag (zum Bericht von E Fischer) (26 Jun 1941). (German) [Theoretical supplement (to the report by E. Fischer) (26 June 1941)]. In Blum et al. [BDR89a], pages 477–480. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:TCP

[Hei89-35] Werner Heisenberg. Theory, criticism and a philosophy. In Bethe et al. [BDH^{+}89], pages 31–46. ISBN 9971-5-0937-7. LCCN QC71

.F74 1989. URL http://adsabs.harvard.edu/abs/1989liph.
book.....B.

Heisenberg:1989:BME

[Hei89-36] Werner Heisenberg. Über die Möglichkeit der Energieerzeugung mit Hilfe des Isotops 238 (19 Nov. 1941). (German) [On the possibility of energy production with the help of isotope 238 (19 November 1941)]. In Blum et al. [BDR89a], pages 499–509. ISBN 0-387-13847-1, 3-540-13847-1. LCCN ???? Wissenschaftliche Originalarbeiten. [Original scientific papers], Edited and with a preface by W. Blum, H.-P. Dürr and H. Rechenberg.

Heisenberg:1989:VGP

[Hei89-37] Werner Heisenberg. Vortrag zum 60. Geburtstag von Professor Bopp (1970). (German) [Lecture on the 60th birthday of Professor Bopp (1970)]. In Blum et al. [BDR89b], pages 621–626. ISBN 3-

492-02929-9. LCCN ???? URL http://d-nb.info/900168226/
04.

Heisenberg:1989:VDF

[Hei89-38] Werner Heisenberg. Vortragsveranstaltung ,,Deutscher Forschungsrat“(Protokoll, 1949). (German) [Lecture program “German Research Council” (minutes, 1949)]. In Blum et al. [BDR89b], pages 72–86. ISBN

3-492-02929-9. LCCN ???? URL http://d-nb.info/900168226/
04.

Heisenberg:1989:VGF

[Hei89-39] Werner Heisenberg. Vorwort. (German) [Foreword]. In Blum et al.

[BDR89b], pages 35–36. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04. For a collection of essays on

the Uranium Project, 1948.

Heisenberg:1989:WSM

[Hei89-40] Werner Heisenberg. Was soll man tun, um die Gefahr eines neuen Weltkrieges in Europa und der Welt endgultig wegzuschaffen? (1970). (German) [What should be done to avoid the possibility of another world war in Europe and the world be finally carried away?]. In Blum et al. [BDR89b], page 560. ISBN 3-492-02929-9.

LCCN ???? URL http://d-nb.info/900168226/04. Original

publication (in Polish) in Express Wieczorny No. 109 (9 May), 3.

Heisenberg:1990:QPGa

[Hei90a] Werner Heisenberg. Quantentheorie und Philosophie. (German)

[Quantum theory and philosophy], volume 9948 of Universal-Bibliothek [Universal Library]. Reclam-Verlag, Stuttgart, West Germany, 1990. ISBN 3-15-009948-X. 125 pp. Vorlesungen und Aufsätze. [Lectures and essays], With an afterword by Jürgen Busche.

Heisenberg:1990:QPGb

[Hei90b] Werner Heisenberg. Quantentheorie und Philosophie. (German) [Quantum theory and philosophy]. Hirzel, Stuttgart, West Germany, 1990. ISBN 3-7776-0457-7. xxxiii + 201 pp. With contributions from Günther Rasche und Bartel L. van der Waerden.

Heisenberg:1991:PLU

[Hei91a] Elisabeth Heisenberg. Das politische Leben eines Unpolitischen: Erinnerungen an Werner Heisenberg. (German) [The political life of an unpolitical man: memories of Werner Heisenberg], volume 1084 of Serie Piper. Piper, München, Germany, third edition, 1991. ISBN 3-492-11084-3. 201 pp. LCCN ????

Heisenberg:1991:CIQ

[Hei91b] Werner Heisenberg. Copenhagen interpretation of quantum theory. In Ferris and Fadiman [FF91], pages 86–96. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Heisenberg:1991:PMR

[Hei91c] Werner Heisenberg. Positivism, metaphysics, and religion. In Ferris and Fadiman [FF91], pages 821–827. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Heisenberg:1993:VDP

[Hei93a] Werner Heisenberg. Die Verhältnisse in der deutschen Physik. (German) [The situation of physics in Germany]. In Blum et al. [BDR93], pages 614–617. ISBN 3-540-13848-X. LCCN ???? Undated draft of memorandum (probably 1938). The referenced appendices are missing.

Heisenberg:1993:FPS

[Hei93b] Werner Heisenberg. Facsimile of the “Preface to the Second Edition” of the Book Introduction to the Unified Field Theory of Elementary Particles. In Blum et al. [BDR93], pages 586–588. ISBN 3-540-13848-X. Unpublished manuscript (about 1973). First edition [Hei66d], reprinted in [Hei84-57]. The second edition was incomplete at the time of Heisenberg’s death.

Heisenberg:1993:FSQ

[Hei93c] Werner Heisenberg. Fortschritte und Schwierigkeiten in der Quantentheorie der Elementarteilchen (Lindau 1953). (German) [Advances and difficulties in the quantum theory of elementary particles (Lindau 1953)]. In Blum et al. [BDR93], pages 309–319. ISBN 3-540-13848-X. LCCN ???? Lecture presented 2 July 1953 at the Nobel Prize Winners’ Meeting in Lindau, West Germany.

Heisenberg:1993:PEB

[Hei93d] Werner Heisenberg. Program for the 2nd edition of the book Introduction to the Unified Field Theory of Elementary Particles. In Blum et al. [BDR93], pages 584–585. ISBN 3-540-13848-X. Unpublished manuscript (about 1973). The second edition was incomplete at the time of Heisenberg’s death.

Heisenberg:1993:QM

[Hei93e] Werner Heisenberg. Quantum mechanics (1930s). In Blum et al. [BDR93], pages 593–596. ISBN 3-540-13848-X. LCCN ???? Unpublished lecture (1930s).

Heisenberg:1993:UGU

[Hei93f] Werner Heisenberg. Über die Uranbombe (1945). (German) [On the uranium bomb]. In Blum et al. [BDR93], pages 618–636. ISBN 3-540-13848-X. LCCN ???? Transcript of original tape of colloquium at Farm Hall, 14 August 1945.

Heisenberg:1995:LBP

[Hei95] Werner Heisenberg. A lecture on bomb physics: February 1942. Physics Today, 48(8):27–30, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/www.physicstoday.org/resource/1/phtoad/v48/i8/p27_s1.

Translated by William Sweet, with an introduction by David Cassidy.

Heisenberg:1996:TGG

[Hei96a] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Piper, Munich, Germany, 1996. ISBN 3-492-22297-8. 287 pp. LCCN ????

Heisenberg:1996:LRM

[Hei96b] Werner Heisenberg. Language and reality in modern physics. In Schirmacher [Sch96a], page ?? ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996.

Heisenberg:1996:PTE

[Hei96c] Werner Heisenberg. A parte e o todo: encontros e conversas sobre física, filosofia, religiao e política. (Portuguese) [The part and the whole: meetings and conversations about fiica, philosophy, religion and politics]. Contraponto, Rio de Janeiro, Brazil, 1996. ISBN 85-85910-13-5. 286 pp. LCCN ????

Heisenberg:1996:QTS

[Hei96d] Werner Heisenberg. Quantum theory and the structure of matter. In Schirmacher [Sch96a], page ?? ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). LCCN QC9.G3 G47 1996.

Heisenberg:1997:FFC

[Hei97] Werner Heisenberg. Fizika i filozofija. (Croatian) [Physics and philosophy]. Kruzak, Zagreb, Croatia, 1997. ISBN 953-96477-3-8. 178 pp. LCCN ????

Heisenberg:1999:SFD

[Hei99] Werner Heisenberg. Lo sfondo filosofico della fisica moderna, volume 19 of Serie Filosofia della scienza. Sellerio, Palermo, Italia, 1999. ISBN 88-389-1450-8. 210 pp. LCCN ???? Edited by Giuseppe Gembillo e Enrico Antonio Giannetto.

Heisenberg:2000:FOI

[Hei00a] Werner Heisenberg. Fisica e oltre: incontri con i protagonisti: 1920–1965. (Italian) [Physics and Beyond: meetings with the protagonists: 1920–1965]. Saggi. Scienze. Bollati Boringhieri, Torino, Italia, 2000. ISBN 88-339-0127-0. 261 pp. LCCN ????

Heisenberg:2000:PPR

[Hei00b] Werner Heisenberg. Physics and philosophy: the revolution in modern science. Penguin classics. Penguin, London, UK, 2000. ISBN 0-14-118215-6. xvii + 143 pp. LCCN ???? introduction by Paul Davies.

Heisenberg:2000:PPG

[Hei00c] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Hirzel, Stuttgart, Germany, sixth edition, 2000.

ISBN 3-7776-1024-0. 288 pp. LCCN QC6. DM 36.00, sfr 36.00, S 263.00.

Heisenberg:2001:TGG

[Hei01a] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Piper, Munich, Germany, 2001. ISBN 3-492-22297-8. 287 pp. LCCN ????

Heisenberg:2001:WHG

[Hei01b] Werner Heisenberg. Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen . (German) [Werner Heisenberg: reports and examination reports for doctorates and postdoctoral]. ERS-Verlag, Berlin, Germany, 2001. ISBN 3-928577-42-5. 266 pp. LCCN ????

Heisenberg:2002:GWH

[Hei02] Werner Heisenberg. Gutachten von Werner Heisenberg [für Carl Friedrich Freiherr von Weizsäcker]. (German) [Report of Werner Heisenberg [for Carl Friedrich Freiherr von Weizsäcker]]. In Rechenberg and Wiemers [RW02], pages 218–221. ISBN 3-928577-43-3. LCCN QC16.H35 W475 2002.

Heisenberg:2003:QPG

[Hei03] Werner Heisenberg. Quantentheorie und Philosophie. (German) [Quantum theory and philosophy], volume 9948 of Universal-Bibliothek [Universal Library]. Reclam-Verlag, Stuttgart, Germany, 2003. ISBN 3-15-009948-X. 125 pp. Vorlesungen und Aufsätze. [Lectures and essays], With an afterword by Jürgen Busche.

Heisenberg:2004:FPH

[Hei04] Werner Heisenberg. Fundamental physics, Heisenberg and beyond: Werner Heisenberg Centennial Symposium ‘Developments in Modern Physics’. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20201-3. ix + 188 pp. LCCN ????

Heisenberg:2005:I

[Hei05a] Werner Heisenberg. Introduction. In Born and Born [BB05], pages xxxiv–xxxvii. ISBN 1-4039-4496-2. LCCN QC16.E5 A4 2005.

URL http://www.loc.gov/catdir/bios/hol059/2004061027.
html;   http://www.loc.gov/catdir/description/hol054/

2004061027.html;   http://www.loc.gov/catdir/enhancements/

fy0618/2004061027-t.html. Note on the new edition by Gustav Born. New preface by Diana Buchwald and Kip S. Thorne. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. Translated by Irene Born.

Heisenberg:2005:WHB

[Hei05b] Werner Heisenberg. Werner Heisenberg 1901–1976; Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976, contributions, reports, letters: Festschrift to mark his 100th birthday]. Verlag der Sächsischen Akademie der Wissenschaften, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. 424 pp. LCCN ????

Heisenberg:2006:TGG

[Hei06] Werner Heisenberg. Der Teil und das Ganze: Gespräche im Umkreis der Atomphysik. (German) [The part and the whole: Discussions in the periphery of atomic physics]. Piper, Munich, Germany, 2006. ISBN 3-492-04890-0. 287 pp. LCCN ????

Heisenberg:2007:CIQ

[Hei07a] Werner Heisenberg. The Copenhagen interpretation of quantum theory. In Whitfield and Hicks [WH07], page ?? ISBN 1-880323-91-5. LCCN QC7.5 .W53 2007. URL http://store.greatbooks. org/adu-wtm.html. Foreword by Alan Lightman.

Heisenberg:2007:PPR

[Hei07b] Werner Heisenberg. Physics and philosophy: the revolution in modern science. Harper Perennial modern classics. HarperPerennial, New York, NY, USA, 2007. ISBN 0-06-120919-8 (paperback). xxi + 201, 24 pp. LCCN QC6 .H335 2007. URL http://

www.loc.gov/catdir/toc/fy0803/2008270999.html. Introduc-

tion by David Lindley.

Heisenberg:2007:PPG

[Hei07c] Werner Heisenberg. Physik und Philosophie. (German) [Physics and philosophy]. Hirzel, Stuttgart, Germany, seventh edition, 2007. ISBN 3-7776-1024-0. 288 pp. LCCN ???? EUR

18.40. URL http://www.gbv.de/dms/hebis-darmstadt/toc/

184174120.pdf. With contributions by Günther Rasche and Bartel Leendert van der Waerden.

Heisenberg:2008:WWW

[Hei08a] Jochen Heisenberg. Who was Werner Heisenberg? http://wernerheisenberg.unh.edu/, September 25, 2008.

Heisenberg:2008:PPQ

[Hei08b] Werner Heisenberg. Die physikalischen Prinzipien der Quantentheorie. (German) [The physical principles of quantum theory]. Hirzel, Stuttgart, Germany, 2008. ISBN 3-7776-1616-8. xx + 117 pp. LCCN ????

Heisenberg:2008:FOI

[Hei08c] Werner Heisenberg. Fisica e oltre: incontri con i protagonisti: 1920–1965. (Italian) [Physics and Beyond: meetings with the protagonists: 1920–1965], volume 554 of Universale Bollati Boringhieri. Bollati Boringhieri, Torino, Italia, 2008. ISBN 88-339-1865-3. 271 pp. LCCN ????

Heisenberg:2008:AMS

[Hei08d] Werner Karl Heisenberg. Abstraction in modern science. In Nishina Memorial Foundation, editor, Nishina Memorial Lectures: Creators of Modern Physics, volume 746 of Lecture Notes in Physics, pages 1–16. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 4-431-77056-9. LCCN ????

Heisenberg:2010:SNB

[Hei10a] Werner Heisenberg. In der Spur von Niels Bohrs Physik und Philosophie. (German) [On the track of Niels Bohr’ physics and philosophy]. In Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die ,,fröhliche Wissenschaft“(Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)] [Rec10], pages 223–297. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). LCCN QC16.H35 R434 2010. URL http://

ebooks.ciando.com/book/index.cfm/bok_id/46829;   http:
//www.ciando.com/img/books/width167/3540692223_k.jpg;
 http://www.ciando.com/pictures/bib/3540692223bib_t_1.
jpg.

Heisenberg:2010:QWA

[Hei10b] Werner Heisenberg. Quantenmechanik, Wellenmechanik und Anschauung. (German) [Quantum mechanics, wave mechanics and opinion]. In Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die ,,fröhliche Wissenschaft“(Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)] [Rec10], pages 449–525. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). LCCN QC16.H35 R434 2010. URL http://

ebooks.ciando.com/book/index.cfm/bok_id/46829;   http:
//www.ciando.com/img/books/width167/3540692223_k.jpg;
 http://www.ciando.com/pictures/bib/3540692223bib_t_1.
jpg.

Heisenberg:2011:PP

[Hei11] Werner Heisenberg. Physik und Philosophie. (German) [Physics and Philosophy]. Hirzel, Stuttgart, Germany, eighth edition, 2011. ISBN 3-7776-2153-6. 288 pp. LCCN ????

Hendry:1986:BRW

[Hen86] John Hendry. Book review: Wolfgang Pauli. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Berlin, Heidelberg and New York: Springer, 1985. Pp. xxix + 783. ISBN 3-540-13609-6. DM 298.00. British Journal for the History of Science, 19(3): 348, November 1986. CODEN BJHSAT. ISSN 0007-0874 (print),

1474-001X (electronic). URL http://www.jstor.org/stable/
4026519.

Henderson:1993:DWP

[Hen93] Giles Henderson. 2-D wave packets and the Heisenberg Uncertainty Principle. Journal of Chemical Education, 70(12):972–976, December 1993. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic).

Hentschel:1996:PNS

[Hen96] Klaus Hentschel, editor. Physics and National Socialism: an anthology of primary sources, volume 18 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 3-7643-5312-0 (Basel), 0-8176-5312-0 (Boston). ci + 406 + civ pp. LCCN Q127.G3 P48 1996.

Hentschel:1998:HHM

[Hen98] Klaus Hentschel. Heinrich Hertz’s mechanics: A model for Werner Heisenberg’s April 1925 paper on the anomalous Zeeman effect. In Heinrich Hertz: Classical Physicist, Modern Philosopher [BHN98], chapter 12, pages 183–223. ISBN 90-481-4881-2, 94-015-8855-4 (e-book). ISSN 0068-0346. LCCN

???? URL https://link.springer.com/chapter/10.1007/
978-94-015-8855-3_12.

Hentschel:2000:ERH

[Hen00] K. Hentschel. Essay review: Heisenberg, German culture, and other such horrifying things. Paul Lawrence Rose, Heisenberg and the Nazi Atomic Bomb Project, 1939–1945, ISBN 0-520-21077-8. Annals of Science, 57(3):301–306, 2000. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Hentschel:2007:MAM

[Hen07] Klaus Hentschel. The mental aftermath: the mentality of German physicists 1945–1949. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920566-3 (hardcover). 205 pp. LCCN QC9.G3 H46 2007. URL http://www.loc.gov/

catdir/toc/fy0805/2007280005.html.

Hentschel:2010:ERH

[Hen10] Klaus Hentschel. Essay review: Heisenberg contextualized: Cathryn Carson: Heisenberg in the Atomic Age: Science and the Public Sphere. Isis, 101(4):843–845, December 2010. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/10.1086/657517.

Hentschel:2011:BRH

[Hen11] Klaus Hentschel. Book review: Helmut Rechenberg, Werner Heisenberg — Die Sprache der Atome. Leben und Wirken — Eine wissenschaftliche Biographie. Die ‘Fröhliche Wissenschaft’ (Jugend bis Nobelpreis). 2 vols. Heidelberg: Springer, 2010. Pp. xxi + 603. ISBN 978-3-540-69221-8. £168.00 (hardback). British Journal for the History of Science, 44(4):612–614, 2011. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Herbert:1969:KIH

[Her69] D. C. Herbert. Kinematic interaction for the Heisenberg antiferromagnet at low temperature. Journal of Mathematical Physics, 10 (12):2255–2263, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v10/i12/p2255_s1.

Hermann:1976:MWH

[Her76a] Armin Hermann. In memoriam Werner Heisenberg: Akademische Gedenkfeier in München. (German) [In memory of Werner Heisenberg: Academic commemoration in Munich]. Physikalische Blätter, 32(10):433–435, October 1976. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19760321001/
abstract.

Hermann:1976:WH

[Her76b] Armin Hermann. Werner Heisenberg, 1901–1976. Inter Nationes, Bonn-Bad Godesberg, West Germany, 1976. 146 pp. LCCN QC16.H35 H4813.

Hermann:1976:WHS

[Her76c] Armin Hermann. Werner Heisenberg in Selbstzeugnissen und Bilddokumenten. (German) [Werner Heisenberg: in self certifications and pictures], volume 240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, 1976. 152 pp. LCCN QC16.H35 .H47; QC16.H35 H47.

Hermann:1984:WHS

[Her84] Armin Hermann. Werner Heisenberg mit Selbstzeugnissen und Bilddokumenten. (German) [Werner Heisenberg: in self certifications and pictures], volume 240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, 1984. ISBN 3-499-50240-2. 152 pp. LCCN QC16.H37 H47 1984.

Hermann:2001:WHS

[Her01] Armin Hermann. Werner Heisenberg: mit Selbstzeugnissen und Bilddokumenten dargestellt. (German) [Werner Heisenberg: in self certifications and pictures], volume 50240 of Rowohlts Monographien. Rowohlt, Reinbek bei Hamburg, Germany, seventh edition, 2001. ISBN 3-499-50240-2. 151 pp. LCCN QC16.H35 .H47.

Heisenberg:1946:BES

[HF46] Werner Heisenberg and F. Firbas. Besprechungen: Erwin Schrödinger, Statistical Thermodynamics. (German) [Book Reviews]. Naturwissenschaften, 33(10):319, ???? 1946. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL

http://www.springerlink.com/content/wvk1r381221v8g10.

Heisenberg:1965:NHP

[HG65] Werner Heisenberg and Ernesto Grassi. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche enzyklopädie. Rowohlt, Reinbek bei Hamburg, Germany, 1965. 149 pp. LCCN ????

Heisenberg:1939:BGR

[HGT^{+}39] W. Heisenberg, Walter Gross, Rudolf Tschesche, Max Hilzheimer, and P. Schulze. Besprechungen. (German) [Reviews]. Naturwissenschaften, 27(44):741–744, ???? 1939. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/p0332413k8281150.

Heisenberg:1989:HRF

[HGW^{+}89a] Werner Heisenberg, H. Geiger, M. Wien, et al. An den Herrn Reichsminister für Erziehung, Wissenschaft und Volksbildung (Brief). (German) [To the Reich Minister of Education, Science and Public Education (letter)]. In Blum et al. [BDR89b], pages 12–13. ISBN 3-492-02929-9. LCCN ???? URL http://d-nb.

info/900168226/04.

Heisenberg:1989:MAG

[HGW89b] Werner Heisenberg, W. Gerlach, and K. Wirtz. Memorandum vom 7. August 1945. (German) [Memorandum of 7 August 1945]. In Blum et al. [BDR89b], pages 26–27. ISBN 3-492-02929-9. LCCN

???? URL http://d-nb.info/900168226/04.

Heisenberg:1975:CHT

[HH75] Werner Heisenberg and Patrick A. Heelan. Comments to Heelan’s thesis [with reply]. Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 6(1):137–138, ???? 1975. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587

(electronic). URL http://www.jstor.org/stable/25170348.

Heisenberg:1984:EQ

[HH84] W. Heisenberg and G. Hertz. Die Entwicklung der Quantentheorie 1918–1928. (German) [The development of quantum theory 1918–1928]. In Blum et al. [BDR84e], pages 109–115. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_9/.

Hentschel:2011:PNS

[HH11] Klaus Hentschel and Ann Hentschel, editors. Physics and National Socialism: an anthology of primary sources. Modern BirkhÃäuser classics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-0348-0202-1 (paperback), 3-0348-0203-X (e-book). ci + 406 pp. LCCN Q127.G3 P48 2011.

Hentschel:2014:CFW

[HH14] Klaus Hentschel and Dieter Hoffmann, editors. Carl Friedrich von Weizsäcker: Physik — Philosophie — Friedensforschung: Leopoldina-Symposium, vom 20. bis 22. Juni 2012 in Halle (Saale): mit 110 Abbildungen und 1 Tabelle, volume 63 of Acta Historica Leopoldina. Deutsche Akademie der Naturforscher Leopoldina — Nationale Akademie der Wissenschafte, Halle (Saale), Germany, 2014. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Heisenberg:2017:MDL

[HH17] Werner Heisenberg and Elisabeth Heisenberg. My Dear Li: Correspondence, 1937–1946. Yale University Press, New Haven, CT, USA, 2017. ISBN 0-300-19693-8, 0-300-22501-6 (e-book). xvi + 310 (est.) pp. LCCN QC16.H45 A4 2016. Edited by Anna Maria Hirsch-Heisenberg and translated by Irene Heisenberg Werner and Elisabeth Heisenberg.

Heisenberg:2003:LEB

[HHH03] Werner Heisenberg and Anna M. Hirsch-Heisenberg. Liebe Eltern! Briefe aus kritischer Zeit; 1918 bis 1945. (German) [Dear parents! Letters from a critical period, 1918 to 1945]. Langen Müller, Munich, Germany, 2003. ISBN 3-7844-2900-9. 400 pp. LCCN ????

Heisenberg:2011:MLL

[HHHH11] Werner Heisenberg, Elisabeth Heisenberg, and Anna Maria Hirsch-Heisenberg. “Meine liebe Li!”: der Briefwechsel, 1937– 1946: mit Fotos und bislang unveröffentlichten Tagebuchaufzeichnungen aus dem Privatarchiv. (German) [“My dear Li”: correspondence, 1937–1946: with photos and an unpublished diary from the private archive]. Residenz, St. Pölten, Austria, 2011. ISBN 3-7017-3247-7. 349 pp. LCCN QC16.H45 A4 2011.

Heisenberg:1950:BGR

[HHKG50] W. Heisenberg, K. H. Hellwege, R. Kallenbach, and J. Goubeau. Besprechungen. (German) [Reviews]. Naturwissenschaften, 37 (13):311–312, ???? 1950. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/t708073pp644270q.

Heisenberg:2010:SWV

[HHvB10] Werner Heisenberg, Elisabeth Heisenberg, and Gero von Boehm. Die Schönheit der Weltformel Vorträge und Gespräche. (German) [The beauty of the world formula: Lectures and talks]. Quartino, Munich, Germany, 2010. ISBN 3-86750-062-2. ???? pp. LCCN ???? 1 DVD + 1 MP3-CD (530 minutes).

Himmelheber:1990:DPB

[Him90] Max Himmelheber. Die deutschen Physiker und die Bombe. (German) [The German physicists and the bomb]. Frankfurter Allgemeine Zeitung, 155(??):8, July 7, 1990. Letter to the editor.

Heisenberg:1926:AQP

[HJ26] W. Heisenberg and P. Jordan. Anwendung der Quantenmechanik auf das Problem der anomalen Zeemaneffekte. (German) [Application of quantum mechanics to the problem of the anomalous Zeeman effect]. Zeitschrift für Physik, 37(4–5):263– 277, April 1926. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01397100;   http://www.springerlink.
com/content/j387501075121133/.

Harteck:1943:BTA

[HJ43] Paul Harteck and Johannes Jensen. Berechnung des Trenneffektes und der Ausbeute verschiedener Zentrifugenanordnungen zur Erho hung des Wirkungsgrades einer einzelnen Zentrifuge. (German) [Calculation of the separation effect and the yield of different centrifuge arrangements to increase the efficiency of a single centrifuge]. Report FA 002 592, Archives of the Deutsches Museum (ADM), Munich, Germany, February 1943.

Heisenberg:1936:BGRa

[HJFS36] W. Heisenberg, P. Jordan, M. H. Fischer, and O. Schlüter. Besprechungen. (German) [Reviews]. Naturwissenschaften, 24(35): 558–560, ???? 1936. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/h478417q5487xn54.

Hoffman:2000:HLM

[HK00] David K. Hoffman and Donald J. Kouri. Hierarchy of local minimum solutions of Heisenberg’s Uncertainty Principle. Physical Review Letters, 85(25):5263–5267, December 18, 2000. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic),

1092-0145. URL http://prl.aps.org/abstract/PRL/v85/i25/
p5263_1.

Heisenberg:2019:RO

[HK19] Werner Heisenberg and Konrad Kleinknecht, editors. Reality and Its Order. Springer International Publishing, Cham, Switzerland, 2019. ISBN 3-030-25695-2 (print), 3-030-25696-0 (ebook), 3-030-25697-9, 3-030-25698-7. ix + 148 + 2 pp. LCCN

QC5.53; QC6.9. URL http://ezaccess.libraries.psu.edu/
login?url=https://doi.org/10.1007/978-3-030-25696-8.

Heisenberg:1955:QNW

[HKM55] W. Heisenberg, F. Kortel, and H. Mitter. Zur Quantentheorie nichtlinearer Wellengleichungen. III. (German) [On the quantum theory of nonlinear wave equations. III]. Zeitschrift für Naturforschung, 10a:425–446, 1955. CODEN ZNTFA2. ISSN

0372-9516. URL http://labs.adsabs.harvard.edu/ui/abs/
1955ZNatA..10..425H.

Heisenberg:1947:FPN

[HKvL47] Werner Heisenberg, H. Kopfermann, and Max von Laue. Friedrich Paschen [Nachruf]. (German) [Friedrich Paschen [obituary]]. Annalen der Physik (1900), 436(4–5):137–138, 1947. ISSN 1521-3889.

Heisenberg:1984:Q

[HL84] Werner Heisenberg and O. Loewi. Quantenmechanik. (German) [Quantum mechanics]. In Blum et al. [BDR84e], pages 52– 57. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_6/.

Hermann:1978:BFH

[HMSH78] A. Hermann, W. Maaß, K. Strubecker, and G. Hellbardt. Bücher Filme: Heisenberg: Tradition in der Wissenschaft/ Päsler: Phänomenologische Thermodynamik/Schmidt: Intensivkurs Mathematik/Dreier: Zukunft durch kontrolliertes Wachstum. Physikalische Blätter, 34(2):93–95, February 1978. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19780340208/abstract.

Heisenberg:1984:KS

[HMW^{+}84] Werner Heisenberg, G. Molière, K. Wirtz, A. Klemm, H. Vol, C. F. v. Wei säcker, F. Bopp, S. Flügge, and J. Meixner. Kosmische strahlung: Vorträge gehalten im max planck-institut berlin– dahlem. (German) [Cosmic radiation: lectures held in the max planck-institut berlin–dahlem]. In Blum et al. [BDR84e], pages 360–412. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book).

LCCN ???? URL http://link.springer.com/chapter/10.

1007/978-3-642-61742-3_31/. Arnold Sommerfeld zu seinem 75. Geburtstag am 5. Dezember 1943 gewidmet. (German) [Dedicated to Arnold Sommerfeld on his 75th birthday on 5 December 1943].

Haar:1955:PHR

[HN55] D. Ter Haar and W. M. Nicol. Proof of the Heisenberg relations. Nature, 175(4467):1046, June 11, 1955. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v175/
n4467/pdf/1751046a0.pdf.

Hinzke:1999:MCS

[HN99] D. Hinzke and U. Nowak. Monte Carlo simulation of magnetization switching in a Heisenberg model for small ferromagnetic particles. Computer Physics Communications, 121–122: 334–337, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599003483.

Heisenberg:1972:NHP

[HnH72] Werner Heisenberg and Werner Carl Associated name Heisenberg. Das Naturbild der heutigen Physik. (German) [The natural view of contemporary physics], volume 8 of Rowohlts deutsche Enzyklopädie. Rowohlt-Taschenbuch-Verlag, Reinbek bei Hamburg, West Germany, 1972. ISBN 3-499-55008-3. 149 pp. LCCN ????

Hoffmann:1992:FHT

[Hof92] Dieter Hoffmann. Farm-Hall-Tonbänder: Operation Epsilon: Die Geheimdienstakten über die Internierung der deutschen Atomphysiker im englischen Farm Hall sind geöffnet. (German) [Farm Hall tapes: Operation Epsilon: The secret files about the internment of German nuclear physicist in English Farmhall are open]. Physikalische Blätter, 48(12):989–993, December 1992. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19920481205/abstract. See comment [FS93].

Hoffmann:1993:OEF

[Hof93] Dieter Hoffmann. Operation Epsilon: die Farm-Hall-Protokolle oder die Angst der Alliierten vor der deutschen Atombombe. (German) [Operation Epsilon: The Farm Hall protocols or Fear of the Allies against the German Atomic Bomb]. Rowohlt, Berlin, Germany, 1993. ISBN 3-87134-082-0. 379 pp. LCCN ????

Hoffmann:2002:MSB

[Hof02] Dieter Hoffmann. M. Schaaf: Heisenberg, Hitler und die Bombe. Physik Journal, 1(9):67, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http:/

/www.pro-physik.de/details/physikjournalIssue/1089823/
Issue_9_2002.html.

Hoffmann:2005:BAA

[Hof05] Dieter Hoffmann. Between autonomy and accommodation: The German Physical Society during the Third Reich. Physics in Perspective (PIP), 7(3):293–329, September 2005. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-004-0235-x.

Hoffmann:2011:HRB

[Hof11] Dieter Hoffmann. H. Rechenberg: Werner Heisenberg — Sprache der Atome. Physik Journal, 10(2):46, ???? 2011. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL

http://www.pro-physik.de/details/physikjournalIssue/1089637/
Issue_2_2011.html.

Hoffmann:2022:FHA

[Hof22] Dieter Hoffmann. Farm Hall — another look. Berichte zur Wissenschaftsgeschichte, 45(1–2):200–201, June 2022. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Hoffmann:2023:OEF

[Hof23] Dieter Hoffmann. Operation Epsilon: die Farm-Hall-Protokolle erstmals vollständig, ergänzt um zeitgenössische Briefe und weitere Dokumente der 1945 in England internierten deutschen Atomforscher. (German) [Operation Epsilon: the Farm Hall Protocols complete for the first time, supplemented by contemporary letters and other documents from the German nuclear researchers interned in England in 1945]. GNT-Verlag, Berlin, Germany, 2023. ISBN 3-86225-111-X (hardcover), 3-86225-508-5, 3-87134-082-0, 978-38622-5-1. 588 pp. LCCN QC773.A1.

Holton:1984:SSM

[Hol84] Gerald Holton. Success sanctifies the means: Heisenberg, Oppenheimer, and the transition to modern physics. In Mendelsohn [Men84], pages 155–173. ISBN 0-521-26724-2. LCCN Q126.8 .T73

1984. URL http://www.loc.gov/catdir/description/cam022/
84007832.html;   http://www.loc.gov/catdir/toc/cam028/
84007832.html.

Holton:1986:ASB

[Hol86] Gerald James Holton. The advancement of science, and its burdens: the Jefferson Lecture and other essays. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-25244-X, 0-521-27243-2 (paperback). xi + 351 pp. LCCN Q173 .H734

1986. URL http://www.loc.gov/catdir/description/cam032/
86009722.html.

Hollinger:1993:BRH

[Hol93a] David A. Hollinger. Book review: How uncertain was he?: Heisenberg’s War: The Secret History of the German Bomb. By Thomas Powers. Illustrated 608 pp. New York: Alfred A. Knopf. New York Times, ??(??):BR3–BR4, March 14, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/109209420/.

Hollinger:1993:EHS

[Hol93b] David A. Hollinger. Eavesdropping on Hitler’s scientists: What 10 German scholars, interned by the Allies in 1945, said when they thought no one was listening. New York Times, ??(??):A32, October 31, 1993. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://search.proquest.com/
hnpnewyorktimes/docview/109047172/. Review of [Fra93b].

Holton:1998:ASB

[Hol98] Gerald James Holton. The advancement of science, and its burdens: with a new introduction. Harvard University Press, Cambridge, MA, USA, 1998. ISBN 0-674-00530-9 (paperback). xlix + 352 pp. LCCN QK596.C8 Y46.

Holton:2000:BRW

[Hol00a] Gerald Holton. Book review: What is Copenhagen trying to tell us? Los Angeles Times, ??(??):??, December 31, 2000.

URL https://www.latimes.com/archives/la-xpm-2000-dec-
31-bk-6553-story.html.

Holton:2000:WHA

[Hol00b] Gerald Holton. Werner Heisenberg and Albert Einstein. Physics Today, 53(7):38–??, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v53/i7/p38/s1.

Holton:2005:VVS

[Hol05] Gerald James Holton. Victory and vexation in science: Einstein, Bohr, Heisenberg, and others. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01519-3. xi + 229 pp. LCCN Q180.A3 H65 2005.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (e-book). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Holton:2022:FHT

[Hol22] Gerald Holton. Farm Hall transcripts reconsidered. Berichte zur Wissenschaftsgeschichte, 45(1–2):261–264, June 2022. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Horz:1966:WHP

[Hor66] Herbert Horz. Werner Heisenberg und die Philosophie. (German) [Werner Heisenberg and philosophy]. Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1966. 299 + 8 pp. LCCN ????

Houtermans:1930:NAQ

[Hou30] Fritz G. Houtermans. Neuere Arbeiten über Quantentheorie des Atomskerns. (German) [New work on the quantum theory of the atomic nucleus]. Ergebnisse der Exakten Naturwissenschaften, 9 (??):123–221, ???? 1930. CODEN EENAA3. ISSN 0367-0325.

Houtermans:1941:FAK

[Hou41] Fritz G. Houtermans. Zur Frage der Auslösung von Kern-Kettenreaktionen. (German) [On the question of the release of nuclear chain reactions]. Report ????, Mitteilung aus dem Laboratorium Manfred von Ardenne, Berlin-Lichterfelde-Ost, Germany, August 1941.

Heisenberg:1929:QWG

[HP29] Werner Heisenberg and Wolfgang Pauli. Zur Quantendynamik der Wellenfelder. (German) [On the quantum dynamics of wave fields]. Zeitschrift für Physik, 56(1–2):1–61, January 1929. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01340129;
http://www.springerlink.com/content/mr542891u38632q0/.

Heisenberg:1930:QWI

[HP30] W. Heisenberg and W. Pauli. Zur Quantentheorie der Wellenfelder. II. (German) [On the quantum dynamics of wave fields. II]. Zeitschrift für Physik, 59(3–4):168–190, March 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL

http://link.springer.com/article/10.1007/BF01341423;
http://www.springerlink.com/content/m516326152712757/.

Heisenberg:1993:IGT

[HP93] Werner Heisenberg and Wolfgang Pauli. On the isospin group in the theory of the elementary particles. In Blum et al. [BDR93], pages 337–351. ISBN 3-540-13848-X. LCCN ???? Preprint of March 1958. See disagreement [Pau93] in which Pauli broke off collaborative work with Heisenberg in this area. Pauli died on 15 December 1958.

Heisenberg:1933:BMP

[HPE^{+}33] W. Heisenberg, J. Picht, P. P. Ewald, K. Fajans, H. Cassel, et al. Besprechung: Max Planck, Wege zur physikalischen Erkenntnis. (German) [Reviews: Max Planck, Paths to physical knowledge].

Naturwissenschaften, 21(33):608–611, ???? 1933. CODEN NAT-WAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//www.springerlink.com/content/v458v1565v5716jg.

Heisenberg:1953:MDB

[HR53] Werner Heisenberg and F. H. Rein. Memorandum des DFR über die Organisation der wissenschaftlichen Forschung in Deutschland. (German) [Memorandum of DFR [German Research Council] on the organization of scientific research in Germany]. In Eickemeyer [Eic53], pages 122–126. LCCN ???? 117–121.

Heisenberg:1989:MPR

[HR^{+}89] Werner Heisenberg, E. Regener, et al. An die Mitglieder des Parlamentarischen Rates, Bonn (1948). (German) [To the members of the Parliamentary Council in Bonn (1948)]. In Blum et al. [BDR89b], page 71. ISBN 3-492-02929-9. LCCN ???? URL

http://d-nb.info/900168226/04.

Hall:2002:QMH

[HR02a] M. J. W. Hall and M. Reginatto. Quantum mechanics from a Heisenberg-type equality. Fortschritte der Physik = Progress of Physics, 50(5–7):646–651, ???? 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Heisenberg:2002:WHG

[HR02b] Werner Heisenberg and Helmut Rechenberg, editors. Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). (German) [Werner Heisenberg: Appraisal and test certificates for graduations and Habilitations (1929–1942)], volume 29 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS-Verl., Berlin, Germany, second edition, 2002. ISBN 3-928577-43-3. 266 pp. LCCN ???? EUR 27.50.

Hurst:1970:CIH

[HS70] C. A. Hurst and S. Sherman. Correlation inequalities for Heisenberg ferromagnets. Journal of Mathematical Physics, 11(8):2473– 2480, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v11/i8/p2473_s1.

Home:1981:CPN

[HS81] D. Home and S. Sengupta. Comments on a proposed new test of Heisenberg’s Principle. Journal of Physics A (Mathematical and General), 14(2):539–541, February 1981. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Home:1983:HGE

[HS83] D. Home and S. Sengupta. Heisenberg Gedanken experiment revisited. American Journal of Physics, 51(6):567–569, June 1983. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Heisenberg:1960:DSF

[HSBA60] Werner Heisenberg, Erwin Schrödinger, Max Born, and Pierre Auger. Discussione sulla fisica moderna. (Italian) [Discussions on modern physics], volume 1952 of Quattro conferenze organizzate dalle “Rencontres internationales de Genève”. Biblioteca di cultura scientifica, 59. Paolo Boringhieri, Torino, 1960. v + 131 + 1 pp.

Heisenberg:1943:BGR

[HSBB43] Werner Heisenberg, ?. Suhrmann, Günther Briegleb, and ?. Burgeff. Besprechungen. (German) [Reviews]. Naturwissenschaften, 31(29–30):350–352, ???? 1943. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

link.springer.com/article/10.1007/BF01475430;   http:/
/www.springerlink.com/content/r22468135227u5u5.

Heisenberg:1936:BGRb

[HSBR36] W. Heisenberg, E. Sauter, A. Butenandt, and F. Ruttner. Besprechungen. (German) [Reviews]. Naturwissenschaften, 24(40): 639–640, ???? 1936. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/p839112r61026t82.

Heisenberg:1934:MAE

[HSD34a] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Die moderne Atomtheorie: die bei der entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. (German) [Modern Atomic Theory: the 1933 Nobel Prize in Physics lecture in Stockholm]. Verlag von S. Hirzel, Leipzig, Germany, 1934. 2 + 45 + 1 pp. LCCN QC173 .H38.

Heisenberg:1934:MQM

[HSD34b] Werner Heisenberg, Erwin Schrödinger, and P. A. M. (Paul Adrien Maurice) Dirac. Modern Quantum Mechanics: Three Nobel Reports. Technico-Theoretical State Press, Leningrad and Moscow, USSR, 1934. ???? pp.

Hufbauer:1971:BRB

[Huf71] Karl Hufbauer. Book review: Physics and Beyond. Encounters and Conversations by Werner Heisenberg; Arnold J. Pomerans. Isis, 62(4):558–560, Winter 1971. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/229861.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Heisenberg:1947:BRS

[HvL47] Werner Heisenberg and Max von Laue. Das Barlowsche Rad aus supraleitendem Material. (German) [The Barlow wheel made of superconducting material]. Zeitschrift für Physik, 124(7–12):514– 518, July 1947. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01668888;   http://www.springerlink.
com/content/v2r2424412k347x9/.

Heisenberg:1935:BGR

[HvLM^{+}35] W. Heisenberg, M. v. Laue, A. Meissner, F. Trendelenburg, O. v. Auwers, et al. Besprechungen. (German) [Reviews]. Naturwissenschaften, 23(20):321–324, ???? 1935. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/w28hk15267444q42.

Hermann:1979:WPW

[HvMW79] Armin Hermann, Karl von Meyenn, and Victor F. (Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band I: 1919– 1929. (German) [Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume I: 1919–1929], volume 2 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-08962-4, 3-540-08962-4. xlvii + 577 pp. LCCN QC16.P37 A34; QC16.P37 W64. URL http://www.springer.

com/physics/book/978-3-540-08962-9. Edited by Armin Her-

mann, K. von Meyenn and Victor F. Weisskopf, With a foreword by Victor F. Weisskopf, With an introduction by Armin Hermann.

Hardy-Vennen:1997:BGG

[HVTK^{+}97] A. Hardy-Vennen, N. Treitz, L. Kienle, A. Weis, P. Mittelstaedt, K. Kremer, W. Selke, F. Rieß, P. Kienle, and K. Fesser. Buchbesprechung: Gilmore: Die geheimnisvollen Visionen des Herrn S.: Ein physikalisches Märchen nach Charles Dickens/ Krätz: Das Rätselkabinett des Doktor Krätz/Rang: Begriffliche Fundamente quantitativer Naturbeschreibung/Demtröder: Laser Spectroscopy/Finkelstein: Quantum Relativity. A Synthesis of the Ideas of Einstein and Heisenberg/Strobl: The Physics of Polymers/Ambegaokar: Reasoning About Luck: probability and ist uses in physics/Bigelow, Moloney, Philpott u. Rothberg: Nuclear and Particle Physics Simulations/Poenaru: Nuclear Decay Modes/Gu: Soliton Theory and Its Applications/ Kaufmann: Die Entstehung des Weltbildes der Physik (Filmbesprechung). Physikalische Blätter, 53(3):254–260, March 1997. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19970530314/abstract.

Heisenberg:1948:GSG

[HvW48] W. Heisenberg and C. F. von Weizsäcker. Die Gestalt der Spiralnebel. (German) [The shape of the spiral nebula]. Zeitschrift für Physik, 125(4–6):290–292, April 1948. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01454899;   http://www.
springerlink.com/content/tjn1227865720317/.

Heisenberg:1948:GVK

[HW48] Werner Heisenberg and Karl Wirtz. Großversuche zur Vorbereitung der Konstruktion eines Uranbrenners. (German) [Large scale experiments for the construction of a uranium pile]. In Walther Bothe and Siegfried Flügge, editors, FIAT Review of German Science 1939–1945: Nuclear Physics and Cosmic Rays, pages 143– 165. Dieterichsche Verlagsbuchhandlung, Wiesbaden, West Germany, 1948. LCCN ????

Heisenberg:1949:PAG

[HW49] Werner Heisenberg and Wilhelm Westphal. Die Physik der Atomkerne. (German) [Physics of the atomic nucleus]. Friedrich Vieweg und Sohn, Braunschweig, Germany, third extended edition, 1949. viii + 192 pp. LCCN ???? English translation in [Hei53m, Hei53n].

Heisenberg:1984:TUU

[HW84a] W. Heisenberg and E. Wilhelmy. Theoretische Untersuchungen zur Ultrastrahlung. (German) [Theoretical studies on cosmic rays]. In Blum et al. [BDR84e], page 248. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http:

//link.springer.com/accesspage/chapter/10.1007/978-3-
642-61742-3_18?coverImageUrl=/static/0.363/sites/link/
images/abstract_cover_placeholder.png.

Heisenberg:1984:GVK

[HW84b] W. Heisenberg and K. Wirtz. Grossversuche zur Vorbereitung der Konstruktion eines Uranbrenners. (German) [Gross attempts to prepare for the construction of a uranium burner]. In Blum et al. [BDR84e], pages 419–441. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN ???? URL http://link.springer.

com/chapter/10.1007/978-3-642-61742-3_34/.

Hoffmann:2004:GPS

[HW04] Dieter Hoffmann and Mark Walker. The German Physical Society under National Socialism. Physics Today, 57(12):52–58, December 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://scitation.aip.org/content/aip/
magazine/physicstoday/article/57/12/10.1063/1.1878335.

Hoffmann:2007:PZA

[HW07] Dieter Hoffmann and Mark Walker, editors. Physiker zwischen Autonomie und Anpassung: Die Deutsche Physikalische Gesellschaft im Dritten Reich. (German) [Physicists between autonomy and adaptation: The German Physical Society in the Third Reich]. Wiley-VCH, Weinheim, Germany, 2007. ISBN 3-527-40585-2 (hardcover). xii + 675 pp. LCCN QC9.G3 P59 2007.

Hoffmann:2011:FWI

[HW11] Dieter Hoffmann and Mark Walker, editors. “Fremde” Wissenschaftler im Dritten Reich: die Debye-Affäre im Kontext. (German) [“Foreign” scientists in the Third Reich: the Debye affair in context]. Wallstein Verlag, Göttingen, Germany, 2011. ISBN 3-8353-0625-1 (paperback). LCCN Q127.G3 F73 2011. URL

http://d-nb.info/1008461261/04.

Hergert:2012:MTB

[HW12a] W. (Wolfram) Hergert and Thomas Wriedt, editors. The Mie theory: basics and applications, volume 169 of Springer series in optical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-28737-9, 3-642-28738-7 (e-book). ISSN 0342-4111. LCCN QC427.8.M5 M54 2012; QC350 .S76 v. 169.

Hoffmann:2012:GPS

[HW12b] Dieter Hoffmann and Mark Walker, editors. The German Physical Society in the Third Reich: physicists between autonomy and accommodation. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-00684-8 (hardcover). xxiii + 458 pp. LCCN Q49 .G225 2012.

Heisenberg:1969:MMA

[HWY69] W. Heisenberg, H. Wagner, and K. Yamazaki. Magnons in a model with antiferromagnetic properties. Il Nuovo Cimento (10), 59A(3):377–391, February 1969. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/1969NCimA..59..377H.

Heisenberg:1972:MMA

[HY72] W. Heisenberg and K. Yamazaki. Magnons in a model with antiferromagnetic properties. II. Il Nuovo Cimento B, 11B(1):125–137, September 11, 1972. CODEN NCIBAW. ISSN ???? URL http:

//labs.adsabs.harvard.edu/ui/abs/1972NCimB..11..125H.

IEEE:2006:AIS

[IEE06] IEEE, editor. 47th Annual IEEE Symposium on Foundations of Computer Science: FOCS 2006: 21–24 October, 2006, Berkeley, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76

.S974 2006. URL http://ieeexplore.ieee.org/servlet/opac?

punumber=4031329. IEEE Computer Society Order Number P2720.

Iwasawa:1995:ESH

[IM95] H. Iwasawa and K. Matsuo. Exact solutions of Heisenberg equations for multiphoton Jaynes–Cummings model. Journal of Mathematical Physics, 36(6):2720–2726, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Infeld:1933:NDN

[Inf33] Leopold Infeld. Nowe drogi nauki: kwanty i materja (Polish) [New pathways of science: quanta and matter], volume 2 of Z Dziedziny Nauki i Techniki. nakł. Mathesis Polskiej, Warszawa, Poland, 1933. x + 284 pp. LCCN ????

Infeld:1934:WMS

[Inf34] Leopold Infeld. The world of modern science: matter and quanta. G. P. Putnam’s Sons, New York, NY, USA, 1934. 287 pp. LCCN ???? Translation by Louis Infield of [Inf33]. Introduction by Albert Einstein.

Inglis:1932:HTF

[Ing32] D. R. Inglis. The Heisenberg theory of ferromagnetism. Physical Review (2), 42(3):442–443, November 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v42/i3/p442_1.

Ingarden:1970:GHU

[Ing70] R. S. Ingarden. On a generalisation of Heisenberg Uncertainty Principle. Bulletin de l’Académie Polonaise des Sciences — Série des Sciences Mathématiques Astronomiques et Physiques, 18(1): 39–??, ???? 1970.

Irving:1967:TDA

[Irv67a] David John Cawdell Irving. Der Traum von der deutschen Atombombe. (German) [The dream of the German atomic bomb]. Sigbert Mohn Verlag, Gütersloh, Germany, 1967. 328 pp. LCCN QC773.A1 I715 1967.

Irving:1967:GAB

[Irv67b] David John Cawdell Irving. The German atomic bomb; the history of nuclear research in Nazi Germany. Simon and Schuster, New York, NY, USA, 1967. 329 pp. LCCN QC773.A1 I69.

Irving:1967:VHG

[Irv67c] David John Cawdell Irving. The virus house: [Germany’s atomic research and Allied counter-measures]. William Kimber, London, UK, 1967. 288 + 12 pp. LCCN QC773.A1 I7.

Isaacson:1982:CLC

[Isa82] D. Isaacson. The continuum limit of a classical 3-component, onedimensional Heisenberg model is Brownian motion on the surface of a sphere. Journal of Mathematical Physics, 23(1):138–140, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i1/p138_s1.

Jacobi:2001:WHG

[Jac01] Manfred Jacobi. Werner Heisenberg zum 100. Geburtstag: Pionier der Quantenmechanik. (German) [Werner Heisenberg’s 100th birthday: Pioneer of quantum mechanics]. Physik in unserer Zeit, 32(6):274–278, 2001. CODEN PHUZAH. ISSN 1521-3943.

Jacobi:2003:BYW

[Jac03] Manfred Jacobi. Buchbesprechung: 100 Years Werner Heisenberg von D. Papenfuß. Physik in unserer Zeit, 34(3):141, 2003. CODEN PHUZAH. ISSN 0031-9252.

Jacobi:2011:BHS

[Jac11a] Manfred Jacobi. Werner Heisenberg — Die Sprache der Atome, von H. Rechenberg. Physik in unserer Zeit, 42(2):100, March 2011. CODEN PHUZAH. ISSN 0031-9252. URL http:

//onlinelibrary.wiley.com/doi/10.1002/piuz.201190013/
abstract.

Jacobi:2011:BMW

[Jac11b] Manfred Jacobi. Bücher: Magazin: Werner Heisenberg, Die Sprache der Atome. Von H. Rechenberg. Physik in unserer Zeit, 42(2):100, 2011. CODEN PHUZAH. ISSN 1521-3943.

Jadczyk:2014:QFH

[Jad14] Arkadiusz Jadczyk. Quantum fractals: from Heisenberg’s uncertainty to Barnsley’s fractality. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2014. ISBN 981-4569-86-0 (hardcover), 1-322-03070-7 (e-book). xi + 345 pp. LCCN QC20.7.F73 J33 2014.

Jammer:1974:IR

[Jam74a] Max Jammer. The indeterminacy relations. In The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective [Jam74b], chapter 3, pages 55–84. ISBN 0-471-43958-4. LCCN QC173.98 .J35.

Jammer:1974:PQM

[Jam74b] Max Jammer. The Philosophy of Quantum Mechanics: the Interpretations of Quantum Mechanics in Historical Perspective. Wiley, New York, NY, USA, 1974. ISBN 0-471-43958-4. xi + 536 pp. LCCN QC173.98 .J35.

Jammer:1976:WH

[Jam76] Max Jammer. Werner Heisenberg (1901–1976). Zeitschrift für allgemeine Wissenschaftstheorie / Journal for General Philosophy of Science, 7(1):1–10, ???? 1976. CODEN ZAWTA2. ISSN 0044-2216 (print), 1572-8587 (electronic). URL http://www.jstor.

org/stable/25170378.

Jauncey:1929:HIP

[Jau29] G. E. M. Jauncey. Heisenberg’s indetermination principle and the quantum. Nature, 124(3115):57, July 13, 1929. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v124/
n3115/pdf/124057a0.pdf.

Jauch:1993:HUR

[Jau93] W. Jauch. Heisenberg’s uncertainty relation and thermal vibrations in crystals. American Journal of Physics, 61(10):929–932, October 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/?AJP/61/
929/1.

Janzing:2006:MDH

[JD06] Dominik Janzing and Thomas Decker. Minimally disturbing Heisenberg–Weyl symmetric measurements using hard-core collisions of Schrödinger particles. Journal of Mathematical Physics, 47(8):082102, August 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v47/i8/p082102_s1.

Jesudason:1998:SDC

[Jes98] C. G. Jesudason. On some difficulties concerning Heisenberg uncertainty and quantum mechanical relations. Physics Essays, 11 (1):6–33, ???? 1998. CODEN PHESEM. ISSN 0836-1398.

Jasny:1999:MUP

[JHB99] Barbara Jasny, R. Brooks Hanson, and Floyd E. Bloom. A media Uncertainty Principle. Science, 283(5407):1453, March 5, 1999. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/283/5407/
1453.full.

Joel:2013:ISS

[JKS13] Kira Joel, Davida Kollmar, and Lea F. Santos. An introduction to the spectrum, symmetries, and dynamics of spin-1/2 Heisenberg chains. American Journal of Physics, 81(6):450–457, June 2013. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/81/450/1.

Jia:2002:QCC

[JLP^{+}02] Y. W. Jia, Q. H. Liu, X. H. Peng, Wang-Xin, and K. C. Shen. The quantum-classical correspondence of matrix elements of hydrogen atom’s inverse radius in Heisenberg principle. Acta Physica Sinica, 51(2):201–204, ???? 2002.

Johnstone:1959:BRBb

[Joh59] Henry W. Johnstone, Jr. Book review: The Physicist’s Conception of Nature by Werner Heisenberg; Arnold J. Pomerans. Philosophy of Science, 26(4):376, October 1959. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185379.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http:

//www.jstor.org/stable/10.1525/hsns.2013.43.4.391.

Jones:1978:WWB

[Jon78] R. V. (Reginald Victor) Jones. The Wizard War: British Scientific Intelligence, 1939–1945. Coward, McCann and Geoghegan, New York, NY, USA, 1978. ISBN 0-698-10896-5. xx + 556 + 16 pp.

LCCN D810.C88 J66 1978. URL https://en.wikipedia.org/
wiki/Reginald_Victor_Jones.

Jones:1985:MAA

[Jon85] Vincent C. Jones. Manhattan, the Army and the atomic bomb. United States Army in World War II. Special studies. Center of Military History, U.S. Army, Washington, DC, USA, 1985. xx + 660 + 4 pp. LCCN QC773.3.U5 J65 1985.

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Jordan:1971:WHJ

[Jor71] Pascual Jordan. Werner Heisenberg 70 Jahre. Physikalische Blätter, 27(12):559–562, December 1971. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19710271206/
abstract.

Joravsky:1978:SCB

[Jor78] David Joravsky. The scientist as conformist: Book reviews: Soviet Science, by Zhores A. Medvedev, Norton, 262 pp., $10.95. Scientists under Hitler: Politics and the Physics Community in the Third Reich, by Alan D. Beyerchen, Yale University Press, 287 pp., $18.50. Proletarian Science? The Case of Lysenko, by Dominique Lecourt, translated by Ben Brewster, NLB, distributed by Schocken, 170 pp., $11.50. New York Review of Books, 25(15): ??, October 12, 1978. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1978/oct/12/the-scientist-as-conformist/.

Johnson:1998:LMA

[JTD^{+}98] K. S. Johnson, J. H. Thywissen, N. H. Dekker, K. K. Berggren, A. P. Chu, R. Younkin, and M. Prentiss. Localization of metastable atom beams with optical standing waves: Nanolithography at the Heisenberg limit. Science, 280(5369):1583–1586, June 5, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/280/
5369/1583.full.pdf.

Jacquot:1982:LIE

[JU82] J. L. Jacquot and M. Umezawa. Lorentz invariance of the extended object. Journal of Mathematical Physics, 23(9):1693–1696, September 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v23/i9/p1693_s1.

Jungk:1956:HTS

[Jun56] Robert Jungk. Heller als tausend Sonnen. Das Schicksal der Atomforscher. (German) [Brighter than a thousand suns. The fate of the atomic scientists]. Schertz und Goverts, Stuttgart, West Germany, 1956. 368 pp. LCCN ????

Jungk:1957:SET

[Jun57] Robert Jungk. Stærkere end tusind sole: Atomforskernes skæbne. (Danish) [Brighter than a thousand suns: the fate of the atomic researchers]. Jespersen og Pio, Copenhagen, Denmark, 1957. 308 pp. LCCN ???? Translation from German by Inger Gudmundsen.

Jungk:1958:BTT

[Jun58] Robert Jungk. Brighter Than a Thousand Suns: a Personal History of the Atomic Scientists. Harcourt Brace, New York, NY, USA, 1958. 369 pp. LCCN QC773 .J813. Translated by James Cleugh.

Jungk:1964:HTS

[Jun64] Robert Jungk. Heller als tausend Sonnen: das Schicksal der Atomforscher. (German) [Brighter than a Thousand Suns: the fate of the nuclear scientists], volume 600/601 of rororo Taschenbuch. Rowohlt Taschenbuch Verlag, Reinbek bei Hamburg, West Germany, 1964. 346 pp. LCCN ????

Jungk:1972:HTS

[Jun72] Robert Jungk. Heller als tausend Sonnen: das Schicksal der Atomforscher. (German) [Brighter than a Thousand Suns: the fate of the nuclear scientists], volume 6629 of Rororo-Taschenbuch. Rowohlt, Reinbek bei Hamburg, West Germany, 1972. ISBN 3-499-16629-1 (kart.). 346 pp. LCCN ????

Jungk:1982:BTT

[Jun82] Robert Jungk. Brighter than a thousand suns: a personal history of the atomic scientists. Penguin, New York, NY, USA, 1982. ISBN 0-14-020667-1 (paperback). 329 pp.

Jungk:1990:HTS

[Jun90] Robert Jungk. Heller als tausend Sonnen: das Schicksal der Atomforscher. (German) [Brighter than a Thousand Suns: the fate of the nuclear scientists], volume 108 of Heyne-Bücher 19, Heyne-Sachbuch. Heyne, München, Germany, expanded edition, 1990. ISBN 3-453-04019-8. 411 pp. LCCN ????

Jin:2011:FSK

[JZ11] Yongyang Jin and Genkai Zhang. Fundamental solutions of Kohn sub-Laplacians on anisotropic Heisenberg groups and H-type groups. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 54(1):126–140, March 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Kuiper:1971:GFT

[KA71] L. K. Kuiper and Jeffrey J. Adels. Green’s function treatment of two-spin Heisenberg ferromagnet. American Journal of Physics, 39(3):253–256, March 1971. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/?AJP/39/253/1.

Kaempffert:1931:HEU

[Kae31] Waldemar Kaempffert. How to explain the universe? Science in a quandary: One after another the theories put out by the scientists have been exploded and now science in its uncertainty has been forced to become idealistic and to drop the idea of a mechanical universe. New York Times, ???(??):120, January 11, 1931. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/99177884/13F390B7A1F783DA798.

Kaempffert:1947:MKR

[Kae47] Waldemar Kaempffert. Mr. Kaempffert replies [to Samuel A. Goudsmit]. New York Times, ??(??):E8, November 9, 1947. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/108047021/. See [Gou47a].

Kaempffert:1948:NSI

[Kae48a] Waldemar Kaempffert. Nazis spurned idea of an atomic bomb. New York Times, ??(??):10, December 28, 1948. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. Interview with Werner Heisenberg.

Kaempffert:1948:RRB

[Kae48b] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/108348269/.

Kant:2002:WHG

[Kan02] Horst Kant. Werner Heisenberg and the German Uranium Project, volume 203 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wissenschaftsgeschichte, Berlin, Germany, 2002. 40 pp. LCCN ????

Kant:2003:MHH

[Kan03] Horst Kant. A. M. Hirsch-Heisenberg: Werner Heisenberg — Liebe Eltern! Briefe aus kritischer Zeit 1918–1945. (German) [A. M. Hirsch-Heisenberg: Werner Heisenberg — Dear Parents! Letters from the critical period 1918–1945]. Physik Journal, 2(10):60, ???? 2003. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-

physik.de/details/physikjournalIssue/1089799/Issue_10_

2003.html. See [HHH03].

Karlsch:2005:GTK

[Kar05a] Rainer Karlsch. Ein geheimnisvolles Treffen in Kopenhagen. (German) [A mysterious meeting in Copenhagen]. In Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests] [Kar05c], pages 78–81. ISBN 3-421-05809-1. LCCN

QC773.3.G3 K37 2005. URL http://www.loc.gov/catdir/toc/
fy054/2005433177.html.

Karlsch:2005:HTR

[Kar05b] Rainer Karlsch. Heisenbergs Theorie für Reaktor und Bombe. (German) [Heisenberg’s theory for reactor and bomb]. In Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests] [Kar05c], pages 39–40. ISBN 3-421-05809-1. LCCN

QC773.3.G3 K37 2005. URL http://www.loc.gov/catdir/toc/
fy054/2005433177.html.

Karlsch:2005:HBG

[Kar05c] Rainer Karlsch. Hitlers Bombe: die geheime Geschichte der deutschen Kernwaffenversuche. (German) [Hitler’s bomb: The Secret History of German Nuclear Tests]. Deutsche Verlags-Anstalt, München, Germany, 2005. ISBN 3-421-05809-1. 415 pp. LCCN

QC773.3.G3 K37 2005. URL http://www.loc.gov/catdir/toc/
fy054/2005433177.html.

Karlsch:2006:ABH

[Kar06] Rainer Karlsch. Atomowa bomba Hitlera: historia tajnych niemieckich prób z bronia jadrowa. Wojny — konflikty. Wydawn. Dolnoślaskie, Wrocław, Poland, 2006. ISBN 83-7384-512-7. 366 pp. LCCN ???? Polish translation by Jerzy Pasieka of [Kar05c].

Karlsch:2007:BHH

[Kar07] Rainer Karlsch. La bombe de Hitler: histoire secrète des tentatives allemandes pour obtenir l’arme nucléaire. (French) [Hitler’s bomb: the secret history of German attempts to get the nuclear weapon]. Le Grand livre du mois, Paris, France, 2007. ISBN 2-286-03945-3 (hardcover). 522 pp. LCCN ???? URL http://catalogue.bnf.

fr/ark:/12148/cb41169004k.

Kastner:2011:QNE

[Kas11] Ruth E. Kastner. Quantum nonlocality: Not eliminated by the Heisenberg Picture. Foundations of Physics, 41(7):1137–1142, July 2011. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/s10701-011-9536-5.

Katscher:1996:WDH

[Kat96] Friedrich Katscher. What did Heisenberg know? Nature, 380 (6571):196, March 21, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v380/n6571/pdf/380196a0.pdf.

Kaveh:2024:RHU

[Kav24] Shahin Kaveh. A reinterpretation of Heisenberg’s Umdeutung in prescriptive-dynamical terms. Studies in History and Philosophy of Science Part A, 107(??):43–53, October 2024. CO-DEN SHPSB5. ISSN 0039-3681 (print), 1879-2510 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0039368124001092.

Karwowski:1996:EMH

[KBWJ96] Jacek Karwowski, Dorota Bielínska-Waż, and Jacek Jurkowski. Eigenvalues of model Hamiltonian matrices from spectral density distribution moments: The Heisenberg spin Hamiltonian. International Journal of Quantum Chemistry, 60(1):185–193, October 5, 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X

(electronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=60582.

Kean:2009:RBU

[Kea09] Sam Kean. Review: Beyond Uncertainty by David C. Cassidy. New Scientist, ??(2695):??, February 11, 2009. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http:

//www.newscientist.com/article/mg20126952.500-review-
beyond-uncertainty-by-david-c-cassidy.html.

Kockel:1960:BHW

[KED^{+}60] B. Kockel, W. H. Estphal, F. Dessauer, K. Strubecker, H. Siedentopf, F. Borgnis, H. Falkenhagen, W. Finkelnburg, and R. Vieweg. Bücher: Heisenberg: Wandlungen in den Grundlagen der Naturwissenschaft/Heisenberg: Physik und Philosophie/Frisch: Beiträge zur Physik und Chemie des 20. Jahrhunderts/Bellman: Introduction to Matrix Analysis/Mehlin: Astronomy/Born und Wolf: Principles of Optics/Mandl: Introduction to Quantum Field Theory/Marley, Morgan: Health Physics/Churchman: Measurement Definitions and Theories. Physikalische Blätter, 16(6):344– 347, June 1960. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-

3722 (electronic). URL http://onlinelibrary.wiley.com/doi/
10.1002/phbl.19600160610/abstract.

Kemmer:1985:HMH

[Kem85] Nicholas Kemmer. Heisenberg: the man and his physics. Nature, 313(6005):826, February 28, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v313/n6005/pdf/313826a0.pdf.

Kraus:2001:UGE

[KF01] Elisabeth Kraus and Carl Joachim Friedrich. Von der Uranspaltung zur Göttinger Erklärung: Otto Hahn, Werner Heisenberg, Carl Friedrich von Weizsäcker und die Verantwortung des Wissenschaftlers. (German) [From uranium fission to the Göttingen Declaration: Otto Hahn, Werner Heisenberg, Carl Friedrich von Weizsäcker and the responsibility of the scientist]. Königshausen and Neumann, Würzburg, Germany, 2001. ISBN 3-8260-1987-3. 449 pp. LCCN ????

Karwowski:2002:RBP

[KF02] Jacek Karwowski and Norbert Flocke. Relations between Pariser– Parr–Pople and Heisenberg models. International Journal of Quantum Chemistry, 90(3):1091–1098, ???? 2002. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Krivoruchenko:2007:WSH

[KF07] M. I. Krivoruchenko and Amand Faessler. Weyl’s symbols of Heisenberg operators of canonical coordinates and momenta as quantum characteristics. Journal of Mathematical Physics, 48(5): 052107, May 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v48/i5/p052107_s1.

Kunze:2010:LTM

[KGNS10] T. Kunze, S. Gemming, S. Numazawa, and M. Schreiber. Lowtemperature modeling for degenerate and frustrated Heisenberg systems with anisotropy. Computer Physics Communications, 181(4):806–812, April 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509004056.

Kramers:1925:SSD

[KH25] H. A. Kramers and W. Heisenberg. Über die Streuung von Strahlung durch Atome. (German) [On the scattering of radiation by atoms]. Zeitschrift für Physik, 31(1):681–708, February/April 1925. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.springer.

com/article/10.1007/BF02980624;   http://wokinfo.com/
media/pdf/heisenbergexample.pdf;   http://www.springerlink.
com/content/x2x7220805540747/.

Koppe:1947:SWS

[KH47] H. Koppe and W. Heisenberg. Die spezifische Wärme der Supraleiter nach der Theorie. (German) [The specific heat of the superconductor according to theory]. Annalen der Physik (1900), 436(7–8):405–414, 1947. ISSN 1521-3889.

Koeth:2019:TJU

[KH19] Timothy Koeth and Miriam Hiebert. Tracking the journey of a uranium cube: a mysterious object led two physicists to investigate the German quest and failure to build a working nuclear reactor during World War II. Physics Today, 72(5):36–43, May 1, 2019. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL https://physicstoday.scitation.org/doi/

10.1063/PT.3.4202. See [CR13] for more on missing uranium cubes, and [MWL^{+}11] for nuclear forensics of the reactor cubes.

Kienle:1943:BGR

[KHBF43] H. Kienle, W. Heisenberg, L. Bergmann, and F. Firbas. Besprechungen. (German) [Reviews]. Naturwissenschaften, 31(21– 22):250–254, ???? 1943. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.springerlink.

com/content/gg477j8075340344.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Katsura:1964:LHM

[KI64] Shigetoshi Katsura and Sakari Inawashiro. Linear Heisenberg model of ferro- and antiferromagnetism. Journal of Mathematical Physics, 5(8):1091–1106, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i8/p1091_s1.

Kresic-Juric:2006:BFP

[KJ06] Sasa Kresić-Jurić. On the Birkhoff factorization problem for the Heisenberg magnet and nonlinear Schrödinger equations. Journal of Mathematical Physics, 47(6):063501, June 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i6/
p063501_s1.

Kitanine:2009:TLF

[KKM^{+}09] N. Kitanine, K. K. Kozlowski, J. M. Maillet, N. A. Slavnov, and V. Terras. On the thermodynamic limit of form factors in the massless XXZ Heisenberg chain. Journal of Mathematical Physics, 50(9):095209, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v50/i9/p095209_s1.

Klein:1954:BRBa

[Kle54] Martin J. Klein. Book review: Nuclear Physics by W. Heisenberg. Philosophy of Science, 21(3):270, July 1954. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://

www.jstor.org/stable/185289.

Klein:1978:WAB

[Kle78] Abraham Klein. WKB approximation for bound states by Heisenberg matrix mechanics. Journal of Mathematical Physics, 19(1): 292–297, January 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i1/p292_s1.

Kleinknecht:2017:EHB

[Kle17] Konrad Kleinknecht. Einstein und Heisenberg: Begründer der modernen Physik. (German) [Einstein and Heisenberg: Founders of Modern Physics]. Verlag W. Kohlhammer, Stuttgart, Germany, 2017. ISBN 3-17-032385-7 (paperback), 3-17-032388-1 (e-book). 221 pp. LCCN QC15 .K54 2017.

Klotz:1993:GFH

[Klo93] Irving Klotz. Germans at Farm Hall knew little of A-bombs. Physics Today, 46(10):11, 13, 15, 135, October 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://www.physicstoday.org/resource/1/phtoad/v46/i10/
p11_s1.

Klotz:1996:BP

[Klo96] Irving M. Klotz. Bending perception: [Nazi Science: Myth, Truth, and the German Atomic Bomb. By Mark Walker. Plenum: 1995. Pp. 325. $28.95]. Nature, 379(6564):410–412, February 1, 1996. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Klein:1980:VPP

[KLV80] Abraham Klein, Ching Teh Li, and Moyez Vassanji. Variational principles for particles and fields in Heisenberg matrix mechanics. Journal of Mathematical Physics, 21(10):2521–2527, October 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i10/p2521_s1.

Koornwinder:1989:MPP

[Koo89] Tom H. Koornwinder. Meixner–Pollaczek polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 30(4):767– 769, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i4/p767_s1.

Kornelson:2005:LSL

[Kor05] Keri A. Kornelson. Local solvability of Laplacian difference operators arising from the discrete Heisenberg group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??): 598–615, ???? 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karlsch:2007:WHB

[KP07] Rainer Karlschh and Heiko Petermann, editors. Für und Wider “Hitlers Bombe”: Studien zur Atomforschung in Deutschland. (German) [For and against “Hitler’s bomb”: studies on nuclear research in Germany], volume 29 of Cottbuser Studien zur Geschichte von Technik, Arbeit und Umwelt. Waxmann, Münster, Germany, 2007. ISBN 3-8309-1893-3 (paperback), 3-927359-21-1. 349 pp. LCCN QC773.3.G3 F87 2007.

Kernbauer:2020:POM

[KPMN20] Alois Kernbauer, Manfred Popp, Frank Ulrich Müller, and Hermann Nicolai, editors. Paul Otto Müller: Schrödingers talentierter Schüler: seine Dissertation zur Allgemeinen Relativitätstheorie, seine Mitarbeit am Uranprojekt, volume 50 of Publikationen aus dem Archiv der Universität Graz. ADEVA, Akademische Drucku. Verlagsanstalt, Graz, Austria, 2020. ISBN 3-201-02051-6. 242 pp. LCCN ????

Kurtsiefer:1995:HMA

[KPS^{+}95] C. Kurtsiefer, T. Pfau, S. Spalter, C. R. Ekstrom, and J. Mlynek. A Heisenberg microscope for atoms. In Greenberger and Zeilinger [GZ95], pages 162–172. CODEN ANYAA9. ISBN 0-89766-921-5, 0-89766-922-3 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). LCCN Q11 .N5 vol. 755; QC173.96.

Kragh:1984:EMF

[Kra84] Helge Kragh. Equation with the many fathers: the Klein–Gordon equation in 1926. American Journal of Physics, 52(11):1024–1033, November 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Kragh:1986:BRD

[Kra86a] Helge Kragh. Book review: David Cassidy & Martha Baker. Werner Heisenberg. A Bibliography of his Writings (Berkeley Papers in History of Science, IX), Berkeley: Office for History of Science and Technology, University of California, 1984. pp. 153. W. Heisenberg. Gesammelte Werke/Collected Works, Series B. Eds, W. Blum, H.-P. Dürr and H. Rechenberg. Berlin, Heidelberg, New York, Tokyo: Springer-Verlag, 1984. Pp. 947. Elisabeth Heisenberg. Inner Exile. Recollections of a Life with Werner Heisenberg. Translated by S. Cappellari and C. Morris, Boston, Basel, Stuttgart: Birkhäuser, 1984. Pp. 170. British Journal for the History of Science, 19(2):232, July 1986. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4026614.

Kramish:1986:GGU

[Kra86b] Arnold Kramish. The Griffin: The Greatest Untold Espionage Story of World War II. Houghton-Mifflin, Boston, MA, USA, 1986. ISBN 0-395-36318-7. viii + 294 + 12 pp. LCCN D810.S8 R644 1986.

Kramish:1987:GPR

[Kra87] Arnold Kramish. Der Greif: Paul Rosbaud — der Mann, der Hitlers Atompläne scheitern liess. (German) [The Griffin: Paul Rosbaud — the man, let Hitler’s nuclear plans fail]. Kindler Verlag, München, Germany, 1987. ISBN 3-463-40060-X. 352

pp. LCCN D810.S8 R636. URL http://digitool.hbz-nrw.de:
1801/webclient/DeliveryManager?pid=3006542.

Kramish:1993:PHE

[Kra93] Arnold Kramish. Powers on Heisenberg: Embellishments on the Lesart. American Scientist, 81(??):479–480, ???? 1993. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Kragh:1995:AMW

[Kra95] Helge Kragh. Arthur March, Werner Heisenberg, and the search for a smallest length. Revue d’Histoire des Sciences, 48(4):401– 434, October 1995. CODEN RHSAAM. ISSN 0151-4105 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23633252.

Kragh:1996:BRBa

[Kra96] Helge Kragh. Book review: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Volume 3: 1940–1949 by Wolfgang Pauli; Karl von Mayenn. Isis, 87(1):196, March 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235799.

Kragh:2007:BRM

[Kra07] Helge Kragh. Book review: Matthias Dörries (ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Berkeley Papers in History of Science Vol. 20. Berkeley, CA: Office for History of Science and Technology, 2005. Pp. viii + 195. ISBN 0-9672617-2-4. $12.00 (paperback). British Journal for the History of Science, 40(1):115–116, March 2007. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://

www.jstor.org/stable/4500686.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN QC16.H35 W474 2005. URL http://www.loc.

gov/catdir/toc/fy0613/2006402966.html.

Klein:1992:SGA

[KS92] D. J. Klein and W. A. Seitz. Symmetric-group algebraic variational solutions for Heisenberg models at finite temperature. International Journal of Quantum Chemistry, 41(1):43–52, January 5, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klein:1983:VLS

[KSGB^{+}83] D. J. Klein, W. A. Seitz, M. A. Garcia-Bach, J. M. Picone, and D. C. Foyt. Variational localized-site cluster expansions. XI. Coherent state derivation of effective Heisenberg Hamiltonians. International Journal of Quantum Chemistry, 24(S17):555–571, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Katsura:1978:AHP

[KSSA78] Shigetoshi Katsura, Masaru Sampei, Masayuki Suzuki, and Yoshihiko Abe. Application of the HLISP program for the irreducible representation matrix of the permutation groups and the partition function of the dilute linear Heisenberg chain. Computer Physics Communications, 13(5–6):311–315, January/ February 1978. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0010465578900280.

Klein:1979:SQN

[KtL79] Abraham Klein and Ching teh Li. Semiclassical quantization of nonseparable systems. Journal of Mathematical Physics, 20(4): 572–578, April 1979. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v20/i4/p572_s1.

Kvasnica:1976:HWK

[Kva76] J. Kvasnica. Heisenberg, W K died. Ceskoslovensky Casopis pro Fyziku Sekce A, 24(4):421–423, ???? 1976. CODEN CKCFAH. ISSN 0009-0700 (print), 1804-8536 (electronic).

Khoo:1976:MTS

[KW76] I. C. Khoo and Y. K. Wang. Multiple time scale analysis of an anharmonic crystal. Journal of Mathematical Physics, 17(2):222– 227, February 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i2/p222_s1.

Kursunoglu:1987:RAG

[KW87] Behram Kurşunoğlu and Eugene Paul Wigner, editors. Reminiscences about a great physicist: Paul Adrien Maurice Dirac. Cambridge University Press, Cambridge, UK, 1987. ISBN 0-521-34013-6. xviii + 297 pp. LCCN QC16.D57 R46 1987. URL http:/

/adsabs.harvard.edu/abs/1987ragp.book.....D;   http://
www.loc.gov/catdir/description/cam023/86033409.html;
http://www.loc.gov/catdir/toc/cam028/86033409.html.

Kleint:1993:WHL

[KW93]  Christian Kleint and Gerald Wiemers, editors.  Werner

Heisenberg in Leipzig, 1927–1942, volume 58(2) of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse, 0365-6470. Akademie-Verlag, Berlin, Germany, 1993. ISBN 3-05-501585-1. 263 pp. LCCN QC16.H35 W48 1933. EUR$607 (!).

Kato:2002:DCT

[KW02] Go Kato and Miki Wadati. Direct calculation of thermodynamic quantities for the Heisenberg model. Journal of Mathematical Physics, 43(10):5060–5078, October 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kato:2003:DCF

[KW03] Go Kato and Miki Wadati. A direct calculation of the free energy from the Bethe ansatz equation for the Heisenberg model. Journal of Mathematical Physics, 44(9):4189–4200, September 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karlsch:2005:NLH

[KW05] Rainer Karlsch and Mark Walker. New light on Hitler’s bomb. Physics World, 18(6):14–18, June 2005. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://physicsworldarchive.iop.org/full/pwa-pdf/
18/6/phwv18i6a23.pdf.

Lacki:2002:OAAa

[Lac02a] J. Lacki. Observability, Anschaulichkeit and abstraction: a journey into Werner Heisenberg’s science and philosophy. Fortschritte der Physik = Progress of Physics, 50(5–7):440–458, ???? 2002. CODEN FPYKA6. ISSN 0015-8208 (print), 1521-3978 (electronic).

Lacki:2002:OAAb

[Lac02b] Jan Lacki. Observability, Anschaulichkeit and abstraction: A journey into Werner Heisenberg’s science and philosophy. In Papenfuss et al. [PLS02], pages 11–29. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Lakhtakia:1996:MMH

[Lak96] A. (Akhlesh) Lakhtakia, editor. Models and Modelers of Hydrogen: Thales, Thomson, Rutherford, Bohr, Sommerfeld, Goudsmit, Heisenberg, Schrödinger, Dirac, Sallhofer. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1996. ISBN 981-02-2302-1. xv + 424 pp. LCCN QC171.2 .M63

1996. URL http://www.worldscientific.com/worldscibooks/
10.1142/2788.

Lalli:2015:RHA

[Lal15] Roberto Lalli. Reviews: Heisenberg in the Atomic Age: Science and the Public Sphere — by Cathryn Carson. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 57(4):260–261, November 2015. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Lande:1959:HCW

[Lan59] Alfred Landé. Heisenberg’s contracting wave packets. American Journal of Physics, 27(6):415–417, September 1959. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/27/415/1.

Landea:1972:EQG

[Lan72] Alfred Landea. Einheit in der Quantenwelt (Gegen den Bohr– Heisenberg’schen Positivismus). (German) [Unity in the quantum world (through Bohr–Heisenberg positivism)]. Dialectica: International Review of Philosophy of Knowledge, 26(2):115–130, June 1972. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic).

Landsman:2002:BRG

[Lan02] N. P. Landsman. Book review: Getting even with Heisenberg: P. L. Rose, Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture , University of California Press, Berkeley, 1998, xx + 352 pp. $35, £21.95, ISBN 0-7923-3794-8, hbk. Studies in History and Philosophy of Modern Physics, 33(2):297– 325, June 2002. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S1355219802000151. See [Sch96b].

Landsman:2008:BRM

[Lan08] N. P. Landsman. Book review: Matthias Dörries (Ed.), Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting Between Niels Bohr and Werner Heisenberg. Office for History of Science and Technology, University of California, Berkeley, ISBN 0-9672617-2-4, 2005 (viii + 195 pp., $12.00 pbk). Studies in History and Philosophy of Modern Physics, 39 (2):462–464, May 2008. CODEN ???? ISSN 1355-2198 (print),

1879-2502 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S1355219807001025. See [Dör05].

Laplante:1990:HU

[Lap90] P. Laplante. Heisenberg uncertainty. ACM SIGSOFT Software Engineering Notes, 15(5):21–22, October 1990. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Laplante:1991:HUP

[Lap91] P. A. Laplante. The Heisenberg uncertainty principle and the halting problem. ACM SIGACT News, 22(3):63–65, Summer 1991. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Larmor:1930:SPU

[Lar30] Joseph Larmor. The scientific Principle of Uncertainty. Nature, 125(3149):345, March 8, 1930. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v125/n3149/pdf/125345a0.pdf.

Lassing:2013:FIS

[Läs13] Volker Lässing. Forschung im Schatten der Zollernburg: die Kaiser-Wilhelm-Institute und ihre Nobelpreisträger in Hechingen, Haigerloch und Tailfingen. (German) [Research in the shadow of the Zollernburg: the Kaiser Wilhelm Institutes and their Nobel Prize winners in Hechingen, Haigerloch and Tailfingen]. CM-Verlag, Albstadt, Germany, 2013. ISBN 3-939219-02-9 (hardcover). 224 pp. LCCN Q180.6.G32 K3554 2013.

Laurence:1958:IMW

[Lau58a] William L. Laurence. An indictment of the men who made the bomb: Brighter Than a Thousand Suns. New York Times, ?? (??):BR3, October 12, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/114521964/. Review of [Jun58].

Laurence:1958:R

[Lau58b] William L. Laurence. A reply. New York Times, ??(??): BR52, November 9, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/114361108/.

Laubichler:2007:BRM

[Lau07] Manfred D. Laubichler. Book review: Matthias Dörries: Michael Frayn’s Copenhagen in Debate: Historical Essays and Documents on the 1941 Meeting between Niels Bohr and Werner Heisenberg. Isis, 98(2):401–402, June 2007. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/521471.

Lawson:1972:OEG

[LB72] J. O. Lawson and S. J. Brient. On obtaining the exact Green’s function solution for the two-spin-1/2 Heisenberg ferromagnet. American Journal of Physics, 40(11):1643–1646, November 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/40/1643/1.

Lande:1969:AQG

[LBBH69] Alfred Landé, Max Born, Walter Biem, and W. Heisenberg. Auffassungen über die Quantentheorie. (German) [Conceptions about quantum theory]. Physik Journal, 25(3):105–113, March 1969. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Leibfried:2004:THL

[LBS^{+}04] D. Leibfried, M. D. Barrett, T. Schaetz, J. Britton, J. Chiaverini, W. M. Itano, J. D. Jost, C. Langer, and D. J. Wineland.

Toward Heisenberg-limited spectroscopy with multiparticle entangled states. Science, 304(5676):1476–1478, June 4, 2004. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/304/5676/
1476.full.pdf.

Liu:1971:ARS

[LC71] C. J. Liu and Yutze Chow. An algebraic realization of spinwave theory. Journal of Mathematical Physics, 12(10):2144–2150, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i10/p2144_s1.

Loly:1966:RAD

[LD66] P. D. Loly and S. Doniach. Removal of an apparent discrepancy between calculations of Dyson and of Oguchi for the Heisenberg ferromagnet. Physical Review, 144(??):319–??, April 8, 1966. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.144.319.

Levitt:1995:BRJ

[Lev95] Norman Levitt. Book review: James T. Cushing, Quantum Mechanics: Historical Contingency and the Copenhagen Hegemony. Physics Today, 48(11):84–85, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v48/i11/p84/s1.

Links:1999:BAS

[LFK99] Jon Links, Angela Foerster, and Michael Karowski. Bethe ansatz solution of a closed spin 1 XXZ Heisenberg chain with quantum algebra symmetry. Journal of Mathematical Physics, 40(2):726– 735, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lande:1924:TMH

[LH24] A. Landé and W. Heisenberg. Termstruktur der Multipletts höherer Stufe. (German) [Term structure of higher level multiplets]. Zeitschrift für Physik, 25(1):279–286, December 1924. CODEN ZEPYAA. ISSN 0939-7922 (print),

1431-5831 (electronic).  URL http://link.springer.com/
article/10.1007/BF01327529;   http://www.springerlink.
com/content/mj1642725246154k/.

Lehmann-Haupt:1993:BTD

[LH93] Christopher Lehmann-Haupt. Books of the times: Did a German scientist prevent catastrophe in World War II?: Heisenberg’s War: The Secret History of the German Bomb By Thomas Powers Illustrated. 608 pages. Alfred A. Knopf. New York Times, ?? (??):C13, March 8, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/109213855/.

Liu:2001:HAQ

[LH01] Q. H. Liu and B. Hu. The hydrogen atom’s quantum-to-classical correspondence in Heisenberg’s correspondence principle. Journal of Physics A (Mathematical and General), 34(28):5713–5719, ???? 2001. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Libby:1968:BRB

[Lib68] W. F. Libby. Book review: The German Atomic Bomb. The History of Nuclear Research in Nazi Germany. David Irving. Simon and Schuster, New York, 1968. 329 pp., illus. $6.95. Science, 160 (3824):175–??, April 12, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/160/3824/175.1.citation.

Licata:2016:QFH

[Lic16] Ignazio Licata. Quantum fractals: from Heisenberg’s uncertainty to Barnsley’s fractality, by Arkadiusz Jadczyk. Scope: monograph. level: postgraduate, researcher. Contemporary Physics, 57(2):272– 274, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lightman:2005:DGBb

[Lig05] Alan P. Lightman. The discoveries: great breakthroughs in twentieth-century science. Pantheon, New York, NY, USA, 2005. ISBN 0-676-97789-8, 0-375-42168-8 (hardcover). xviii + 553 + 16 pp. LCCN Q180.55.D57 .L53. URL http://www.loc.gov/

catdir/enhancements/fy0623/2005040854-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0623/2005040854-d.
html;   http://www.loc.gov/catdir/enhancements/fy0623/
2005040854-t.html.

Lindley:2007:UEH

[Lin07] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Doubleday, Garden City, NY, USA, 2007. ISBN 0-385-51506-5. vii + 257 pp. LCCN QC174.17.H4

L56 2007. URL http://www.loc.gov/catdir/toc/ecip0614/
2006017029.html.

Lindley:2008:UEH

[Lin08] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Anchor Books, New York, NY, USA, 2008. ISBN 1-4000-7996-9. viii + 257 pp. LCCN

QC174.17.H4 L56 2007. URL http://www.loc.gov/catdir/toc/
ecip0614/2006017029.html.

Liu:1972:PCA

[Liu72] C. J. Liu. Perturbation calculation of an algebraic realization of spin waves. Journal of Mathematical Physics, 13(2):248–252, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p248_s1.

Liu:2008:HCM

[LJ08] Dong Liu and Cuipo Jiang. Harish-Chandra modules over the twisted Heisenberg–Virasoro algebra. Journal of Mathematical Physics, 49(1):012901, January 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v49/i1/p012901_s1.

Lederer:1977:ZEH

[LL77] Erich Lederer and W. Lohmann. Zwei Erinnerungen an Heisenberg. (German) [Two remembrances of Heisenberg]. Physikalische Blätter, 33(2):89, February 1977. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19770330209/
abstract.

Lipinski:1982:DTF

[LL82] Józef Lipiński and Jerzy Leszczyński. Derivation of theoretical formulas for resonance integrals from Heisenberg equation of motion. International Journal of Quantum Chemistry, 22(2):253–263, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levy-Leblond:1986:CQP

[LL86] Jean-Marc Levy-Leblond. Correlation of quantum properties and the generalized Heisenberg inequality. American Journal of Physics, 54(2):135–136, February 1986. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/?AJP/54/135/1. See note [May87].

Lewis:1929:QLU

[LM29] G. N. Lewis and J. E. Mayer. The quantum laws and the uncertainty principle of Heisenberg. Proceedings of the National Academy of Sciences of the United States of America, 15(2):127– 139, February 15, 1929. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.org/

stable/85605.

Lahti:1987:HIC

[LM87] Pekka J. Lahti and Maciej J. Maczyński. Heisenberg inequality and the complex field in quantum mechanics. Journal of Mathematical Physics, 28(8):1764–1769, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i8/
p1764_s1.

Leibfried:2015:BWE

[LMOS15] Stephan Leibfried, Christoph Markschies, Ernst Osterkamp, and Günter Stock. Berlins wilde Energien: Porträts aus der Geschichte der Leibnizschen Wissenschaftsakademie. (German) [Berlin’s Wild Energies: Portraits from the History of the Leibniz Science Academy]. De Gruyter Akademie Forschung, Berlin, Germany, 2015. ISBN 3-11-037598-2 (hardcover). 517 pp. LCCN AS182.B38 B477 2015.

Lee:2006:MHG

[LN06] James R. Lee and Assaf Naor. L_{p} metrics on the Heisenberg group and the Goemans–Linial conjecture. In IEEE [IEE06], pages 99–108. ISBN 0-7695-2720-5, 0-7695-2362-5. ISSN 0272-5428. LCCN QA76 .S974 2006. URL http://ieeexplore.ieee.

org/servlet/opac?punumber=4031329. IEEE Computer Society

Order Number P2720.

Lochak:2007:COL

[Loc07] Georges Lochak. A complementary opposition: Louis de Broglie and Werner Heisenberg. In Evans and Thorndike [ET07], pages 73–80. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. LCCN QC174.12 .Q346 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006934045-b.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006934045-d.html;   http://www.loc.gov/catdir/toc/fy0708/
2006934045.html.

Loeb:1963:HUP

[Loe63] Arthur L. Loeb. The Heisenberg uncertainty principle. American Journal of Physics, 31(12):945, December 1963. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/31/945/1. Comments on two previous

derivations unknown to [Tya63].

Logan:1993:BMB

[Log93] Jonothan L. Logan. Bomb maker or bomb breaker? Boston Globe, ??(??):B43–B45, March 7, 1993. ISSN 0743-1791. Review of [Pow93b].

Logan:1996:CM

[Log96] Jonothan Logan. The critical mass: As Allied scientists established the feasibility of an atomic bomb, Germany’s leading theorist came to a mistaken conclusion. American Scientist, 84(3): 263–277, May/June 1996. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.jstor.org/

stable/29775672.

Logan:1999:BRP

[Log99] Jonothan Logan. Book review: Paul Lawrence Rose, Heisenberg and the Nazi Atomic Bomb Project: a Study in German Culture. Physics Today, 52(3):81–??, March 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v52/i3/p81/s1.

Logan:2000:RSN

[Log00] Jonothan Logan. Review: “A strange new quantum ethics” [Copenhagen. Michael Frayn. 132 pp. Methuen Publishing Ltd., 2000]. American Scientist, 88(4):356–359, July/August 2000. CO-DEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

URL https://www.jstor.org/stable/27858060.

Lorenz:2011:PP

[Lor15] Robert Lorenz. Protest der Physiker. transcript Verlag, Bielefeld, Germany, 20110615. ISBN 3-8376-1852-8, 3-8394-1852-6. LCCN DD259.4 .L674 2011eb.

Losee:1961:CMP

[Los61] John Price Losee, Jr. A Comparison of Methodological Principles Basic to the Quantum Mechanics of Bohr and Heisenberg, the Metaphysics of Emmet, and the Theology of Tillich. Ph.D dissertation, Drew University, Madison, NJ, USA, November 1961. 253 pp.

Low:1997:CRQ

[Low97] Stephen G. Low. Canonically relativistic quantum mechanics: representations of the unitary semidirect Heisenberg group, U(1, 3) ⊗_{s} H(1, 3). Journal of Mathematical Physics, 38(5):2197– 2209, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Logan:1991:LEH

[LRD^{+}91] Jonothan L. Logan, Helmut Rechenberg, Max Dresden, A. Van Der Ziel, and Mark Walker. Letters to the Editor: Heisenberg, Goudsmit and the German “A-bomb”. Physics Today, 44(5):13, 15, 90–92, 94–96, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/www.physicstoday.org/resource/1/phtoad/v44/i5/p13_s1;
 https://physicstoday.scitation.org/doi/abs/10.1063/1.

2810103. Letters in response to [Wal90a], including Walker’s rebuttal.

Logan:1993:HB

[LS93a] Jonothan L. Logan and Robert Serber. Heisenberg and the bomb. Nature, 362(6416):117, March 11, 1993. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v362/n6416/pdf/362117a0.pdf.

Logan:1993:SCH

[LS93b] Jonothan L. Logan and Robert Serber. Scientific correspondence: Heisenberg and the bomb. Nature, 362(6416):117, March 11, 1993. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v362/
n6416/full/362117a0.html.

Luck:1976:BCE

[LSK^{+}76] W. A. P. Luck, K. Strubecker, K. Kromphardt, A. Haug, H. Fahlenbrach, A. Unsold, B. L. van der Waerden, K. Strubecker, and W. Luck. Bücher: Carnap: Einführung in die Philosophie der Naturwissenschaft/Lecture Notes in Physics/Marion: Physics and the Physical Universe/Grau: Übungsaufgaben zur Quantentheorie/Berichte der Arbeitsgemeinschaft Magnetismus 1971/ Benz: Arnold Sommerfeld/Hermann: Heisenberg/Braun: Differential Equations and Their Applications/Eysenck: Die Ungleichheit der Menschen. Physikalische Blätter, 32(10):474–478, October 1976. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19760321012/abstract.

Li:2012:GTB

[LTG^{+}12] Hong-Zhe Li, Bo Tian, Rui Guo, Yu-Shan Xue, and Feng-Hua Qi. Gauge transformation between the first-order nonisospectral and isospectral Heisenberg hierarchies. Applied Mathematics and Computation, 218(15):7694–7699, April 1, 2012. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300312000513.

Luck:1982:HLS

[Luc82] Werner A. P. Luck. Heisenberg als Lehrer in schwieriger Zeit. (German) [Heisenberg as teacher in difficult times]. Physikalische Blätter, 38(5):130–131, May 1982. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19820380507/
abstract.

Lucas:2004:SCC

[Luc04] Amand A. Lucas. Scribbles that changed the course of human affairs. Classe des Sciences, Académie Royale de Belgique, Bruxelles, Belgique, 2004. ISBN 2-8031-0203-X. ISSN 0365-0936.

110 pp. LCCN ???? URL http://www2.academieroyale.be/
academie/documents/LucasScribbles13657.pdf.

Lucas:2005:BAC

[Luc05] Amand A. Lucas. Bombe atomique et croix gammée. (French) [The atomic bomb and Iron Cross], volume 22 (série 3) of Mémoires de la Classe des sciences. Collection in 8. Académie royale de Belgique, Bruxelles, Belgique, 2005. ISBN 2-8031-0221-8 (paperback). 119 pp. LCCN ???? URL http://catalogue.bnf.

fr/ark:/12148/cb422181177;   http://www2.academieroyale.
be/academie/documents/LucasBombeatomique13655.pdf.

Lucas:2007:RFH

[Luc07] Amand A. Lucas. Revisiting Farm Hall. Europhysics News, 38(4):25–29, July/August 2007. CODEN EUPNAS. ISSN 0531-7479 (print), 1432-1092 (electronic). URL http://www.

europhysicsnews.org/articles/epn/abs/2007/04/epn07402/

epn07402.html. See comment and reply [BL08].

Lucas:2011:BS

[Luc11] Amand Lucas. The Bomb and the Swastika. CreateSpace Independent Publishing, ????, 2011. 68 pp. URL http:

//www.amazon.com/The-Bomb-Swastikahistorys-scientists/

dp/1466426675. A play in four acts.

Lucas:2016:SHB

[Luc16] Amand A. Lucas. Les savants d’Hitler et la bombe atomique. (French) [Hitler’s scientists and the atomic bomb]. L’Académie en Poche. Académie Royale de Belgique, Bruxelles, Belgium, 2016. ISBN 2-8031-0464-4. 104 pp. LCCN ????

Ludovico:2001:EHR

[Lud01] Anna Ludovico. Effetto Heisenberg: la rivoluzione scientifica che ha cambiato la storia. (Italian) [The Heisenberg Effect: the scientific revolution that has changed history]. Temi del nostro tempo. Armando, Roma, Italy, 2001. ISBN 88-8358-182-2. 223 pp. LCCN QC174.17.H4 E34 2001.

Lund:1984:CHC

[Lun84] Fernando Lund. The continuous Heisenberg chain and constrained harmonic motion. Journal of Mathematical Physics, 25(2):256– 260, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i2/p256_s1.

Lurcat:1995:BNA

[Lur95] François Lurçat. La bombe nucléaire allemande et la revision de l’histoire. (French) [The German nuclear bomb and the rewriting of history]. Temps Modernes, 50(??):118–143, ???? 1995.

Lustig:2010:ERL

[Lus10] Harry Lustig. Essay review: The life and times of Werner Heisenberg. Physics in Perspective (PIP), 12(4):470–496, December 2010. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://www.springerlink.com/
content/h348l1l5x5955377/.

Ladenburg:1934:ANP

[LW34] Rudolf Ladenburg and Eugene Wigner. Award of the Nobel Prizes in Physics to Professors Heisenberg, Schrödinger and Dirac. The Scientific Monthly, 38(1):86–91, January 1934. CODEN SC-MOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL

http://www.jstor.org/stable/15534.

Liu:2010:WMT

[LWZ10] Dong Liu, Yuezhu Wu, and Linsheng Zhu. Whittaker modules for the twisted Heisenberg–Virasoro algebra. Journal of Mathematical Physics, 51(2):023524, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v51/i2/p023524_s1.

McMillan:1940:RE

[MA40] Edwin McMillan and Philip Hauge Abelson. Radioactive element 93. Physical Review, 57(12):1185–1186, June 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v57/i12/p1185_2. This

paper, on the discovery of neptunium, was the last paper on transuranic elements and fission in the Physical Review until the after the end of World War II in 1945; see [Wal04].

Maad:2005:SPH

[Maa05] Sara Maad. A semilinear problem for the Heisenberg Laplacian on unbounded domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(??):1279–1290, ???? 2005. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacLachlan:1958:BRW

[Mac58] James MacLachlan. Book review: Werner Heisenberg, Physics and Philosophy: The Revolution of Modern Science. Vol. 19 of World Perspectives. Physics Today, 11(9):36–??, September 1958. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v11/i9/p36/s2.

MacKinnon:1977:HMR

[Mac77] Edward MacKinnon. Heisenberg, models, and the rise of matrix mechanics. Historical Studies in the Physical Sciences, 8(??):137– 188, ???? 1977. CODEN HSPSAS. ISSN 0073-2672. URL http:

//www.jstor.org/stable/27757370.

MacDonald:1978:DH

[Mac78a] N. MacDonald. Dodging Heisenberg. Nature, 272(5655):667, April 20, 1978. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v272/n5655/pdf/272667a0.pdf.

MacKinnon:1978:BRB

[Mac78b] Edward MacKinnon. Book review: Werner Heisenberg: 1901– 1976 by Armin Hermann; Timothy Nevill. Isis, 69(1):149–150, March 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/230672.

MacPherson:1985:STF

[Mac85] Malcolm MacPherson. A small town with a footnote to history: Hidden below, the Nazi A-bomb lab Haigerloch. New York Times, ??(??):10, 47, November 3, 1985. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.

proquest.com/hnpnewyorktimes/docview/111210585.

MacPherson:1986:TBF

[Mac86] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Dutton, New York, NY, USA, 1986. ISBN 0-525-24409-3. xiii + 316 + 8 pp. LCCN QC773 .M24

1986. US$18.95. URL http://catalog.hathitrust.org/api/
volumes/oclc/12751401.html.

MacPherson:1987:TBF

[Mac87] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Berkley Books, New York, NY, USA, 1987. ISBN 0-425-10423-0. xi + 284 + 8 pp. LCCN QC773 .M24 1987.

Macrakis:1993:SSS

[Mac93] Kristie Macrakis. Surviving the swastika: scientific research in Nazi Germany. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1993. ISBN 0-19-507010-0. xii + 280 + 16 pp. LCCN Q180.G4 M3 1993. URL http://www.

h-net.org/review/hrev-a0a0m1-aa;   http://www.loc.gov/
catdir/enhancements/fy0604/93019919-d.html;   http://
www.loc.gov/catdir/enhancements/fy0724/93019919-b.html.

Maier:2007:GBW

[Mai07] Helmut Maier, editor. Gemeinschaftsforschung, Bevollmächtigte und der Wissenstransfer: die Rolle der Kaiser-Wilhelm-Gesellschaft im System kriegsrelevanter Forschung des Nationalsozialismus. (German) [Community research, plenipotentiaries and the transfer of knowledge: the role of the Kaiser Wilhelm Society in the system of Nazi-related research of war-relevant research], volume 17. Wallstein, Göttingen, Germany, 2007. ISBN 3-8353-0182-9 (paperback). LCCN Q180.G4 K38 2000 v.17; U395.G3 G45 2007.

Majorana:1933:KGN

[Maj33] Ettore Majorana. Über die Kerntheorie. (German) [On nuclear theory]. Zeitschrift für Physik, 82(3–4):137–145, March 1933. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01341484.

Manousakis:1991:SHA

[Man91] Efstratios Manousakis. The spin-1/2 Heisenberg antiferromagnet on a square lattice and its application to the cuprous oxides. Reviews of Modern Physics, 63(1):1–62, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
63.1;   http://rmp.aps.org/abstract/RMP/v63/i1/p1_1.

Manasreh:2011:ADH

[Man11] Omar Manasreh. Appendix A: Derivation of Heisenberg Uncertainty Principle. In Introduction to Nanomaterials and Devices, pages 420–423. Wiley, New York, NY, USA, 2011. ISBN 0-470-92707-0 (hardcover), 1-118-14837-1 (e-book), 1-118-14839-8 (ebook), 1-118-14840-1 (e-book), 1-118-14841-X (e-book).

Margenau:1931:UPF

[Mar31] Henry Margenau. The Uncertainty Principle and free will. Science, 74(1928):596, December 11, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/74/1928/596.1.full.pdf.  Com-

ments on [Com31].

Margolin:1973:BRB

[Mar73] Jean-Claude Margolin. Book review: Correspondance 1916-1955 by Albert Einstein, Max Born, Bertrand Russell, Werner Heisenberg, and Pierre Leccia. Les Études philosophiques, 1(??):75, January/March 1973. CODEN ???? ISSN 0014-2166 (print), 2101-0056 (electronic). URL http://

www.jstor.org/stable/20846392;   http://www.jstor.org/
stable/pdfplus/20846392.pdf.

March:1994:BRB

[Mar94] Robert H. March. Book review: The Farm Hall Transcripts, Introduced by Sir Charles Frank, ISBN 0-520-08499-3. Physics Today, 47(7):62–63, July 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v47/i7/p62_s1.

Marburger:2008:HDH

[Mar08] John H. Marburger III. A historical derivation of Heisenberg’s uncertainty relation is flawed. American Journal of Physics, 76(6): 585–587, June 2008. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?AJP/
76/585/1.

Massimi:2004:WDI

[Mas04] M. Massimi. What demonstrative induction can do against the threat of underdetermination: Bohr, Heisenberg, and Pauli on spectroscopic anomalies (1921–24). Synthese, 140(3):243–277, June 2004. CODEN SYNTAE. ISSN 0039-7857 (print), 1573-

0964 (electronic). URL http://link.springer.com/content/
pdf/10.1023/B%3ASYNT.0000031319.64615.49.pdf.

Masters:2014:WHP

[Mas14] Barry R. Masters. Werner Heisenberg’s path to matrix mechanics. Optics and Photonics News, 25(7):42–49, July/August 2014. CODEN OPPHEL. ISSN 1047-6938 (print), 1541-3721

(electronic). URL http://www.osa-opn.org/home/articles/
volume_25/july_august_2014/features/werner_heisenberg_
s_path_to_matrix_mechanics/.

Mathieu:1986:AOBe

[Mat86a] J.-P. Mathieu. Analyses d’ouvrages: Inner Exile. Recollections of a life with Werner Heisenberg par Elisabeth Heisenberg; S.

Cappellari; C. Morris; V. Weisskopf. Revue d’Histoire des Sciences, 39(2):188, avril–juin 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632728.

Mathieu:1986:AOBb

[Mat86b] J.-P. Mathieu. Analyses d’ouvrages: Werner Heisenberg. A bibliography of his writings, (^{≪}Berkeley Papers in History of Science^{≫}, IX) par David Cassidy; Martha Baker. Revue d’Histoire des Sciences, 39(1):87–88, January 1986. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632583.

Mathieu:1987:AOBd

[Mat87] J.-P. Mathieu. Analyses d’ouvrages: Gesammelte Werke, Collected Works, Series A/Part I: Original scientific Papers. Wissenschaftliche Originalarbeiten par Werner Heisenberg; W. Blum; H.-P. Dürr; H. Rechenberg. Revue d’Histoire des Sciences, 40 (3–4):395–396, juillet–décembre 1987. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.

jstor.org/stable/23632915.

Mayants:1987:NCQ

[May87] Lazar Mayants. Note on “Correlation of quantum properties and the generalized Heisenberg inequality” [Am. J. Phys. 54, 135 (1986)]. American Journal of Physics, 55(12):1063, December 1987. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://link.aip.org/link/?AJP/55/1063/1. See

[LL86].

Montwill:2011:QAD

[MB11a] Alex Montwill and Ann Breslin. The quantum adventure: does God play dice? World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). ix + 248 pp. LCCN QC174.12 .M66 2012.

Montwill:2011:WHE

[MB11b] Alex Montwill and Ann Breslin. Werner Heisenberg — an equation for uncertainty. In The quantum adventure: does God play dice? [MB11a], chapter 8, pages 109–124. ISBN 1-84816-647-8 (hardcover), 1-84816-648-6 (paperback), 1-84816-649-4 (e-book). LCCN QC174.12 .M66 2012.

Medeiros:1992:NSA

[MC92] Djalma Medeiros and G. G. Cabrera. Numerical study of antiferromagnetic Heisenberg chains for spin S = 1/2, using the Lanczos method. Journal of Magnetism and Magnetic Materials, 104–107 (part 2):799–800, February 2, 1992. CODEN JMMMDC. ISSN 0304-8853 (print), 1873-4766 (electronic). Proceedings of the International Conference on Magnetism, Part II.

Matsen:1997:FTMa

[MC97] F. A. Matsen and L. L. Campbell. The freeon theory of magnetism. I. the Heisenberg interaction. International Journal of Quantum Chemistry, 65(4):287–297, ???? 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=42766;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=42766&PLACEBO=IE.pdf.

McCloskey:1994:EUP

[McC94] Donald N. McCloskey. An economic Uncertainty Principle. Scientific American, 271(5):107, November 1994. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v271/n5/pdf/scientificamerican1194-107.pdf.

McCauley:2024:HRT

[McC24] Joseph L. McCauley. Heisenberg’s 1939 reactor theory, Serber’s 1943 Los Alamos Primer, and Heisenberg’s 1945 Farm Hall critical mass calculation. American Journal of Physics, 92(10):765–774, October 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). See comments and corrections [Ree24].

Mehra:1973:PCN

[Meh73] Jagdish Mehra, editor. The physicist’s conception of nature: Symposium on the Development of the Physicist’s Conception of Nature in the 20th century. Held at the International Centre for Theoretical Physics, Miramare, Trieste, Italy, 18–25 September 1972. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1973. ISBN 90-277-0345-0, 90-277-2536-5. LCCN QC173.96 .S95 1972. URL http://www.springer.

com/us/book/9789027703453.

Mehra:1975:SCP

[Meh75] Jagdish Mehra, editor. The Solvay conferences on physics: aspects of the development of physics since 1911. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1975. ISBN 90-277-0635-2. xxxii + 415 pp. LCCN QC1.S792 M43.

Mehra:1987:NBD

[Meh87] Jagdish Mehra. Niels Bohr’s discussions with Albert Einstein, Werner Heisenberg, and Erwin Schrödinger: the origins of the Principles of Uncertainty and Complementarity. Foundations of Physics, 17(5):461–506, May 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Menzel:1936:DJP

[Men36] Willi Menzel. Deutsche und jüdische Physik. (German) [German and Jewish physics]. Völkischer Beobachter, ??(??):??, January 29, 1936. See response [Hei36i].

Mendelsohn:1984:TTS

[Men84] Everett Mendelsohn, editor. Transformation and tradition in the sciences: essays in honor of I. Bernard Cohen. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26724-2. xiv + 577 pp. LCCN Q126.8 .T73 1984. URL http://www.loc.gov/

catdir/description/cam022/84007832.html;   http://www.
loc.gov/catdir/toc/cam028/84007832.html.

Merton:1938:SSO

[Mer38] Robert K. Merton. Science and the social order. Philosophy of Science, 5(3):321–337, July 1938. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL http://www.jstor.

org/stable/184838.

Mester:2004:GHI

[Mes04] F. Mester. Generalization of the Heisenberg indeterminacy principle and the implied search for Ultimate Reality and Meaning. Ultimate Reality and Meaning, 27(4):330–357, ???? 2004. CO-DEN ???? ISSN 0709-549X.

Meyer:1976:PPM

[Mey76] K. Meyer. Physiker und Philosoph — in memoriam Werner Heisenberg. (German) [Physicist and philosopher – in memoriam Werner Heisenberg]. Schweizerische Bauzeitung = Revue Polytechnique Suisse, 94(9):100–102, February 26, 1976. CODEN ????

ISSN 0036-7524. URL http://retro.seals.ch/digbib/view?
rid=sbz-002:1976:94::1164&id=hitlist&id2=&id3=.

Mark:1969:PMU

[MF69] Hans Mark and Sidney Fernbach, editors. Properties of matter under unusual conditions: in honor of Edward Teller’s 60th birthday. Interscience Publishers, New York, NY, USA, 1969. ISBN 0-470-56990-5. ix + 389 pp. LCCN Q171 .P98.

Martienssen:1995:BAF

[MGG^{+}95] W. Martienssen, S. Großmann, P. Grassberger, D. Meissner, W. Sandhas, H. Nicolai, D. Hoffmann, H. Kant, P. Richter, H. Atmanspacher, M. Eckert, H. Rechenberg, H. Maier-Leibnitz, and D. Haarer. Buchbesprechungen: Argyris/Faust: Die Erforschung des Chaos/Plaschko/Brod: Nichtlineare Dynamik, Bifurkation und Chaotische Systeme/Bunde/Havlin: Fractals in Science/Goetzberger/Voß/Knobloch: Sonnenenergie: Photovoltaik/Lindner: Grundkurs Theoretische Physik/Vilenkin/Shellard: Cosmic Strings and other Topological Defects/Weiss: Großforschung in Deutschland Geschichte des Hahn-Meitner-Instituts/Fölsing: Wilhelm Conrad Röntgen Aufbruch ins Innere der Materie/Brandmüller: Galilei und die Kirche Ein Fall und seine Lösung/Meÿenn: Wolfgang Pauli Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a. Band 3: 1940 49/Atmanspacher/Primas: Der Pauli-Jung-Dialog und seine Bedeutung für die moderne Wissenschaft/Cassidy: Werner Heisenberg Leben und Werk/Weinberg: The First Nuclear Era The Life and Times of a Technological Fixer/Wenske: Wörterbuch Chemie/Dictionary of Chemistry. Physikalische Blätter, 51(11):1101–1106, November 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19950511117/abstract.

Mielke:1980:ESN

[Mie80] Eckehard W. Mielke. On exact solutions of the nonlinear Heisenberg–Klein–Gordon equation in a space-time of constant spacelike curvature. Journal of Mathematical Physics, 21(3):543– 546, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v21/i3/p543_s1.

Mielke:1981:TES

[Mie81] Eckehard W. Mielke. Toward exact solutions of the nonlinear Heisenberg–Pauli–Weyl spinor equation. Journal of Mathematical Physics, 22(9):2034–2039, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v22/i9/p2034_s1.

Millet:1958:RPP

[Mil58] Louis Millet. Review: Les principes physiques de la théorie des quanta by W. Heisenberg; Louis de Broglie. Les Études philosophiques, 13(1):75, January/March 1958. ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.org/

stable/20842599.

Miller:1985:WHB

[Mil85] Arthur I. Miller. Werner Heisenberg and the beginning of nuclear physics. Physics Today, 38(11):60–??, November 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v38/i11/p60/s1.

Miller:1990:STY

[Mil90] Arthur I. Miller, editor. Sixty-two years of uncertainty: historical, philosophical, and physical inquiries into the foundations of quantum mechanics, volume 226 of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1990. ISBN 0-306-43608-6. LCCN QC173.96 .N385 1989.

Milburn:1997:SMQ

[Mil97] G. J. (Gerard J.) Milburn. Schrödinger’s machines: the quantum technology reshaping everyday life. W. H. Freeman and Company, New York, NY, USA, 1997. ISBN 0-7167-3106-1. xvi + 188 pp. LCCN QC174.13 .M55 1997. URL http://

www.loc.gov/catdir/bios/hol057/96051024.html;   http:/
/www.loc.gov/catdir/description/hol055/96051024.html.

Miller:2010:JPP

[Mil10] Arthur I. Miller. 137: Jung, Pauli, and the Pursuit of a Scientific Obsession. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33864-9. xxiii + 336 pp. LCCN QC16.P37 M55 2010.

Miller:2012:YWM

[Mil12] D. J. Miller. You are Wrong Mr Einstein! Newton, Einstein, Heisenberg and Feynman Discussing Quantum Mechanics, by Harald Fritzsch, Scope: general interest. Level: undergraduates and/ or general readers. Contemporary Physics, 53(5):437–439, 2012. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Millard:1971:ISL

[ML71] Kenneth Millard and Harvey S. Leff. Infinite-spin limit of the quantum Heisenberg model. Journal of Mathematical Physics, 12(6):1000–1005, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v12/i6/p1000_s1.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. Institute of Physics Publishers, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Mnatsakanova:1998:IRH

[MMSV98] M. Mnatsakanova, G. Morchio, F. Strocchi, and Yu Vernov. Irreducible representations of the Heisenberg algebra in Krein spaces. Journal of Mathematical Physics, 39(5):2969–2982, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Monod:2004:EHK

[Mon04] Emmanuel Monod. Einstein, Heisenberg, Kant: methodological distinction and conditions of possibilities. Information and Organization, 14(2):105–121, April 2004. CODEN ???? ISSN 1471-7727 (print), 1873-7919 (electronic). URL https://dblp.org/

db/journals/iando/iando14.html#Monod04.

Monk:2009:GIH

[Mon09] Ray Monk. Getting inside Heisenberg’s head. In Garry Hagberg, editor, A Companion to the Philosophy of Literature, pages 453– 464. Wiley-Blackwell, Malden, MA, USA, 2009. ISBN 1-4443-1559-5.

Mooney:1993:THUa

[Moo93a] Christopher F. Mooney. Theology and the Heisenberg Uncertainty Principle: I. The Heythrop Journal, 34(3):247–273, 1993. ISSN 0018-1196 (print), 1468-2265 (electronic).

Mooney:1993:THUb

[Moo93b] Christopher F. Mooney. Theology and the Heisenberg Uncertainty Principle: II. The Heythrop Journal, 34(4):373–386, 1993. ISSN 1468-2265.

Morrison:1947:ASG

[Mor47] Philip Morrison. Alsos: The story of German science. Bulletin of the Atomic Scientists, 3(12):354, 365, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [vL47] and reply [Mor48].

Morrison:1948:RDL

[Mor48] Philip Morrison. A reply to Dr. von Laue. Bulletin of the Atomic Scientists, 4(4):104, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab48, vL48b].

Morita:1971:TTS

[Mor71] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. Journal of Mathematical Physics, 12(10):2062–2069, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i10/
p2062_s1.

Morita:1972:GSD

[Mor72a] Tohru Morita. General structure of the distribution functions for the Heisenberg model and the Ising model. Journal of Mathematical Physics, 13(1):115–123, January 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i1/p115_s1.

Morita:1972:TTSa

[Mor72b] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. II. Journal of Mathematical Physics, 13(5):714–720, May 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i5/p714_s1.

Morita:1972:TTSb

[Mor72c] Tohru Morita. Two-time spin-pair correlation function of the Heisenberg magnet at infinite temperature. III. BCC lattice and arbitrary spin. Journal of Mathematical Physics, 13(9):1428–1435, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i9/p1428_s1.

Moritz:1997:GGH

[Mor97] H. Moritz. Gauss, Gödel, Heisenberg. In Natke and Ben-Haim [NBH97], pages 101–108. ISBN 3-05-501740-4. LCCN Q375 .U52

1997. URL http://cs.utep.edu/interval-comp/main.html;
http://www.vchgroup.de/books/tis/eng/3-05-501740-4.html.

Mosch:1942:BWH

[Mos42] E. Mosch. Buchbesprechung: Werner Heisenberg, Wandlungen in den Grundlagen der Naturwissenschaft. 6 Vorträge. 3. erw. Aufl. 95 S. Leipzig 1942, Verlag S. Hirzel. Preis kart. 3,50 M. Zeitschrift für Angewandte Mathematik und Mechanik, 22(5):300–301, 1942. CODEN ZAMMAX. ISSN 1521-4001.

Mott:1985:BRE

[Mot85] Nevill Mott. Book review: Elisabeth Heisenberg, Inner Exile: Recollections of a Life with Werner Heisenberg. Physics Today, 38(11):101–??, November 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v38/i11/p101/s1.

Mott:1977:WHD

[MP77] Sir Nevill Mott and Sir Rudolf Peierls. Werner Heisenberg. 5 December 1901–1 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23:212–251, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL

http://www.jstor.org/stable/769614.

Makhankov:1992:CCH

[MP92] V. G. Makhan′kov and O. K. Pashaev. Continual classical Heisenberg models defined on graded su(2, 1) and su(3) algebras. Journal of Mathematical Physics, 33(8):2923–2936, August 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v33/i8/
p2923_s1.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Masing:1924:BGR

[MPB^{+}24] G. Masing, Otto Poppenberg, M. Bergmann, F. Paneth, W. Roth, W. Heisenberg, et al. Besprechungen. (German) [Reviews]. Naturwissenschaften, 12(40):827–832, ???? 1924. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://

www.springerlink.com/content/t0556358756l77u1.

Mazzoccoli:2024:RHP

[MRV24] Alessandro Mazzoccoli, Jorge Andres Rivero, and Pierluigi Vellucci. Refining Heisenberg’s principle: a greedy approximation of step functions with triangular waveform dictionaries. Mathematics and Computers in Simulation, 225(??):165–176, November 2024. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037847542400185X.

Meyer:2015:TNE

[MS15] Carl H. Meyer and Günter Schwarz. The theory of nuclear explosives that Heisenberg did not present to the German military. Preprint 467, Max-Planck-Institut für Wissenschaftsgeschichte, Berlin, Germany, 2015. ii + 36 pp.

URL https://www.mpiwg-berlin.mpg.de/preprint/theory-
nuclear-explosives-heisenberg-did-not-present-german-
military.

Matsumoto:1980:EOQ

[MSUU80] H. Matsumoto, G. Semenoff, H. Umezawa, and M. Umezawa. Extended objects in quantum field theory. Journal of Mathematical Physics, 21(7):1761–1769, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v21/i7/p1761_s1.

Mumpower:2022:IST

[MTV22] Matthew R. Mumpower, Patrick Talou, and Ramona Vogt. Impact on science and technology. In Talou and Vogt [TV23], chapter 4, pages 401–464. ISBN 3-031-14544-5 (hardcover), 3-031-14545-3 (e-book), 3-031-14546-1, 3-031-14547-X. LCCN ????

Muller:1940:BVU

[Mül40] Paul Otto Müller. Eine Bedingung für die Verwendbarkeit von Uran als Sprengstoff. (German) [A condition for the usability of uranium as an explosive]. Report FA-002-482, Deutsches Museum Archiv, Berlin, Germany, May 1940.

Mayer:2011:NFM

[MWL^{+}11] K. Mayer, M. Wallenius, K. Lützenkirchen, J. Galy, Z. Varga, N. Erdmann, R. Buda, J.-V. Kratz, N. Trautmann, and K. Fifield. Nuclear forensics: A methodology applicable to nuclear security and to non-proliferation. Journal of Physics: Conference Series, 312(6):062003, September 2011. CODEN JPCSDZ. ISSN 1742-6588 (print), 1742-6596 (electronic).

Mayer:2015:UGN

[MWL^{+}15] Klaus Mayer, Maria Wallenius, Klaus Lützenkirchen, Joan Horta, Adrian Nicholl, Gert Rasmussen, Pieter van Belle, Zsolt Varga, Razvan Buda, Nicole Erdmann, Jens Volker Kratz, Norbert Trautmann, L. Keith Fifield, Stephen G. Tims, Michaela B. Fröhlich, and Peter Steier. Uranium from German nuclear power projects of the 1940s — a nuclear forensic investigation. Angewandte Chemie, International Edition, 54(45):13452–13456, September 2015. CO-DEN ACIEF5. ISSN 1433-7851 (print), 1521-3773 (electronic).

Nagel:2016:GDU

[Nag16] Günter Nagel. Das geheime deutsche Uranprojekt 1939– 1945 — Beute der Alliierten. (German) [The secret German uranium project 1939–1945 — Allied loot]. Heinrich-Jung-Verlagsgesellschaft mbH, Zella-Mehlis, Germany, 2016. ISBN 3-943552-10-1 (hardcover). 560 pp. LCCN QC773.3.G3 N35 2016.

Nardone:1993:HPQ

[Nar93] Pascal Nardone. Heisenberg picture in quantum mechanics and linear evolutionary systems. American Journal of Physics, 61(3): 232–237, March 1993. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?AJP/
61/232/1.

Nairz:2002:EVH

[NAZ02] Olaf Nairz, Markus Arndt, and Anton Zeilinger. Experimental verification of the Heisenberg uncertainty principle for fullerene molecules. Physical Review A (Atomic, Molecular, and Optical Physics), 65(3):032109, February 2002. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http:

//pra.aps.org/abstract/PRA/v65/i3/e032109.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57(6):546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/ 10.1103/PhysRev.57.546. This paper reports experiments that confirm the Bohr–Wheeler prediction [?] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Natke:1997:UMM

[NBH97] H. Gunther Natke and Yakov Ben-Haim, editors. Uncertainty: models and measures: proceedings of the international workshop held in Lambrecht, Germany, July 22–24, 1996, volume 99 of Mathematical research. Akademie-Verlag, Berlin, Germany, 1997. ISBN 3-05-501740-4. LCCN Q375 .U52 1997. URL

http://cs.utep.edu/interval-comp/main.html;   http://
www.vchgroup.de/books/tis/eng/3-05-501740-4.html.

Nelson:1972:ASK

[Nel72] C. A. Nelson. Algebraic solution for the Källén–Pauli state vectors in the V θ sector by a congruence transformation. Journal of Mathematical Physics, 13(7):1051–1056, July 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i7/p1051_s1.

Neufeld:1999:BRP

[Neu99] Michael J. Neufeld. Book review: Paul Lawrence Rose, Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture. Berkeley/Los Angeles/London: University of California Press, 1998, xx, 352 pages. $35.00 (cloth). Physics in Perspective (PIP), 1(2):227–228, June 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Newing:1935:UPZ

[New35] R. A. Newing. Uncertainty Principle and the zero-point energy of the harmonic oscillator. Nature, 136(3436):395, September 7, 1935. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v136/
n3436/pdf/136395b0.pdf.

Ni:2004:GMH

[Ni04a] Yilong Ni. Geodesics in a manifold with Heisenberg group as boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):566–589, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ni:2004:HKG

[Ni04b] Yilong Ni. The heat kernel and Green’s function on a manifold with Heisenberg group as boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(??):590–611, ???? 2004. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nicolacopoulos:1990:GSP

[Nic90] Pantelis Nicolacopoulos. Greek Studies in the Philosophy and History of Science, volume 121 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 94-010-7403-8, 94-009-2015-6 (e-book). ISSN 0068-0346. xviii + 439 pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-009-2015-6.

Nishina:1932:HDL

[Nis32] Yoshio Nishina. Heisenberg and Dirac’s lectures on the problems of quantum mechanics. Keimeika Gaiyo, 11(??):3–??, ???? 1932.

Napolitano:2011:IBQ

[NKD^{+}11] M. Napolitano, M. Koschorreck, B. Dubost, N. Behbood, R. J. Sewell, and M. W. Mitchell. Interaction-based quantum metrology showing scaling beyond the Heisenberg limit. Nature, 471(7339): 486–489, March 24, 2011. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v471/n7339/full/nature09778.html.

Noyes:1931:UP

[Noy31] W. A. Noyes. The Uncertainty Principle. Science, 74(1928):595– 596, December 11, 1931. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/74/1928/595.2.full.pdf. See [Com31, Dar31,

Mar31].

Nelson:1971:RBS

[NRS71] C. A. Nelson, A. K. Rajagopal, and C. S. Shastry. Relation between the singularities of the S matrix and the L^{2} class of solutions in potential theory. Journal of Mathematical Physics, 12(4): 737–741, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i4/p737_s1.

Nordmann:1995:OEW

[NW95] Alfred Nordmann and Hartmut Wickert. Operation Epsilon: oder wie die Deutschen doch noch ihre Atombombe erfanden. (German) [Operation Epsilon: or how the Germans still invented their bomb]. Niedersächsische Staatstheater, Hannover, Germany, 1995. 80 pp. LCCN ????

OBrien:2003:SLS

[OD03] Ted A. O’Brien and Ernest R. Davidson. Semiempirical local spin: Theory and implementation of the ZILSH method for predicting Heisenberg exchange constants of polynuclear transition metal complexes. International Journal of Quantum Chemistry, 92(3):294–325, April 5, 2003. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

OFlaherty:1992:WHN

[O’F92] James C. O’Flaherty. Werner Heisenberg on the Nazi revolution: Three hitherto unpublished letters. Journal of the History of Ideas, 53(3):487–494, July/September 1992. CODEN JHIDAT. ISSN 0022-5037 (print), 1086-3222 (electronic). URL https://

www.jstor.org/stable/2709890.

Olsen:1963:TGS

[Ols63] Arthur J. Olsen. Trackdown of the German scientist: Nazism and defeat scattered the leaders of a once-great scientific establishment. Herewith a review of where they went and what some of them did when they got there. New York Times, ??(??): 214, September 22, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/116463805/.

Omelyanovskii:1975:HDP

[Ome75] M. E. Omelyanovskii. Heisenberg and the development of physical concepts. Voprosy Filosofii, 1(??):88–96, ???? 1975. CODEN ???? ISSN 0042-8744.

Omelyanovskii:1979:SIR

[Ome79] M. E. Omelyanovskii. Some introductory remarks on article by Heisenberg, W., Sense and Meaning of Beauty in Exact Sciences. Voprosy Filosofii, 12(??):44–48, ???? 1979. CODEN ???? ISSN 0042-8744.

Oord:2009:CMF

[Oor09] Thomas Jay Oord, editor. Creation made free: open theology engaging science. Pickwick Publications, Eugene, OR, USA, 2009. ISBN 1-60608-488-7 (paperback). vii + 272 pp. LCCN BL240.3 .C743 2009.

Oren:1998:HMP

[Ore98] Dan A. Oren. Heisenberg meets photobiology? Science, 279 (5353):963, February 13, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/279/5353/963.8.full.

Olson:1998:BES

[OS98] Richard Olson and Roger Smith, editors. Biographical encyclopedia of scientists. Marshall Cavendish, New York, NY, USA, 1998. ISBN 0-7614-7064-6 (set), 0-7614-7065-4 (vol. 1), 0-7614-7066-2 (vol. 2), 0-7614-7067-0 (vol. 3), 0-7614-7068-9 (vol. 4), 0-7614-7098-0 (vol. 5). lv + 1460 + lxxx (five volumes) pp. LCCN Q141 .B532 1998.

Oexle:2003:HHA

[OS03] Otto Gerhard Oexle and Carola Sachse. Hahn, Heisenberg und die anderen: Anmerkungen zu ‘Kopenhagen’, ‘Farm Hall’ und ‘Göttingen’. (German) [Hahn, Heisenberg and the others: Comments on ‘Copenhagen ’,‘ Farmhall’ and ‘Göttingen ’], volume 9. Forschungsprogramm Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus, Berlin, Germany, 2003. 49 pp. LCCN ????

Odake:2007:EHO

[OS07] Satoru Odake and Ryu Sasaki. Exact Heisenberg operator solutions for multiparticle quantum mechanics. Journal of Mathematical Physics, 48(8):082106, August 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v48/i8/p082106_s1.

Osterath:2011:HCD

[Ost11] Brigitte Osterath. Heisenberg der Chemielehrer und Drogenkoch. (German) [Heisenberg: the chemistry teacher and drug chef]. Nachrichten aus der Chemie, 59(4):427–428, 2011. CODEN NACHFB. ISSN 1439-9598 (print), 1868-0054 (electronic).

Oechsner:1986:BKP

[OSWR86] H. Oechsner, A. Schlachetzki, H. Walther, and H. Rechenberg. Buchbesprechungen: Kirschner: Polarized Electrons at Surfaces/Paul, Teubner: Optoelektronische Halbleiterbauelemente/ Stitch, Bass: Laser Handbook, Vol. 4/Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u. a.: Teil II: 1930–1939. Physikalische Blätter, 42(7):252–253, July 1986. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19860420726/abstract.

Otremba:1979:WHA

[Otr79] Heinz Otremba. Werner Heisenberg: Atomphysiker u. Philosoph. (German) [Werner Heisenberg: atomic physicist and philosopher]. Echterhaus, Würzburg, West Germany, 1979. 59 pp.

Oppenheim:2010:UPD

[OW10] Jonathan Oppenheim and Stephanie Wehner. The Uncertainty Principle determines the nonlocality of quantum mechanics. Science, 330(6007):1072–1074, November 19, 2010. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/330/6007/1072.full.pdf.

Ozawa:2002:PMI

[Oza02] Masanao Ozawa. Position measuring interactions and the Heisenberg uncertainty principle. Physics Letters A, 299(1):1–7, June 24, 2002. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037596010200659X.

Ozawa:2003:PCH

[Oza03] Masanao Ozawa. Physical content of Heisenberg’s uncertainty relation: limitation and reformulation. Physics Letters A, 318(1– 2):21–29, November 3, 2003. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Packer:1988:SME

[Pac88] Judith A. Packer. Strong Morita equivalence for Heisenberg C^{∗}- algebras and the positive cones of their K_{0}-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 40(??):833–864, ???? 1988. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pagels:1982:CCQ

[Pag82] Heinz R. Pagels. The Cosmic Code: Quantum Physics as the Language of Nature. Simon and Schuster, New York, NY, USA, 1982. ISBN 0-671-24802-2. 370 + 8 pp. LCCN QC174.13 .P33 1982.

Pais:1991:NBT

[Pai91] Abraham Pais. Niels Bohr’s times: in physics, philosophy, and polity. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773

.P35 1991. US$35.00. URL http://www.loc.gov/catdir/
enhancements/fy0635/90027248-d.html.

Pandit:2002:HEC

[Pan02] Giridhari Lal Pandit. Heisenberg–Einstein context principle and the dynamic core-context of discovery in physics. In Papenfuss et al. [PLS02], pages 32–53. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15

2002. EUR 119.00. URL http://www.loc.gov/catdir/bios/
wiley046/2003268171.html;   http://www.loc.gov/catdir/
description/wiley0310/2003268171.html;   http://www.loc.
gov/catdir/toc/wiley031/2003268171.html.

Park:2006:BRG

[Par06] David Park. Book review: Gerald Holton, Victory and Vexation in Science: Einstein, Bohr, Heisenberg, and Others. Cambridge, Mass. and London: Harvard University Press, 2005. xiv + 229 pages. $35.00 (cloth). Physics in Perspective (PIP), 8(4):481–483, December 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Pauli:1931:BRW

[Pau31] Wolfgang Pauli. Book review: W. Heisenberg, Die physikalischen Prinzipien der Quantentheorie. Naturwissenschaften, 19(??):188– 189, ???? 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Pauli:1934:DPM

[Pau34] Wolfgang Pauli. Discussion of paper by M. Heisenberg on “La Structure du Noyau”. In Cockcroft et al. [CCJ^{+}34], pages 324– 325. LCCN ???? Publiés par la commission administrative de l’institut.

Pauli:1941:RFT

[Pau41] W. Pauli. Relativistic field theories of elementary particles. Reviews of Modern Physics, 13(3):203–232, July 1941. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
13.203;   http://rmp.aps.org/abstract/RMP/v13/i3/p203_
1.

Pauli:1993:PIG

[Pau93] Wolfgang Pauli. Postscript: “On the isospin group in the theory of the elementary particles”. In Blum et al. [BDR93], pages 337–351. ISBN 3-540-13848-X. The author reports: “As essential parts of the preprint [[HP93]] with the above title don’t any longer agree with my opinion, I am forced to give up the plan to publish a common paper with Heisenberg on the subject in question.”.

Pawlak:2002:BFB

[Paw02a] Alexander Pawlak. Bohrende Fragen: Bohr-Dokumente zum Treffen mit Heisenberg 1941 in Kopenhagen freigegeben. (German) [Nagging questions: Bohr documents on the meeting with Heisenberg in 1941 released in Copenhagen]. Physik Journal, 1(3):7–8, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1108683/issue.html.

Pawlak:2002:RHV

[Paw02b] Alexander Pawlak. Rückblick auf Heisenbergs verschränktes Leben. (German) [Review of Heisenberg’s entangled life]. Physik Journal, 1(1):8, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-

physik.de/details/physikjournalIssue/1089837/Issue_1_
2002.html.

Pawlak:2006:WHO

[Paw06] Alexander Pawlak. W. Heisenberg: Originaltonaufnahmen 1951–1967. (German) [W. Heisenberg: original recordings 1951– 1967]. Physik Journal, 5(3):63, ???? 2006. CODEN PJHOB2.

ISSN 1617-9439 (print), 1619-6597 (electronic). URL http:/

/www.pro-physik.de/details/physikjournalIssue/1089745/
Issue_3_2006.html.

Pawlak:2008:GSB

[Paw08] Alexander Pawlak. G. Schiemann: Werner Heisenberg. Physik Journal, 7(5):63, ???? 2008. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-

physik.de/details/physikjournalIssue/1089697/Issue_5_
2008.html.

Price:1977:UPF

[PC77] William Charles Price and Seymour S. Chissick, editors. The Uncertainty Principle and foundations of quantum mechanics: a fifty years’ survey. Wiley, New York, NY, USA, 1977. ISBN 0-471-99414-6. xvii + 572 pp. LCCN QC174.125 .U5.

Page:2014:TGI

[PC14] Julien Page and Gabriel Catren. Towards a Galoisian interpretation of Heisenberg indeterminacy principle. Foundations of Physics, 44(12):1289–1301, December 2014. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/s10701-014-9812-2.

Popp:2023:PGU

[PdK23] Manfred Popp and Piet de Klerk. The peculiarities of the German uranium project (1939–1945). Journal of Nuclear Engineering, 4 (3):634–653, September 2023. CODEN ???? ISSN 2673-4362.

Porsezian:1992:IAO

[PDL92] K. Porsezian, M. Daniel, and M. Lakshmanan. On the integrability aspects of the one-dimensional classical continuum isotropic biquadratic Heisenberg spin chain. Journal of Mathematical Physics, 33(5):1807–1816, May 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v33/i5/p1807_s1.

Pearce:1976:MCF

[Pea76] Paul A. Pearce. Monotonicity of correlation functions. Journal of Mathematical Physics, 17(5):744–746, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v17/i5/
p744_s1.

Peierls:1971:AGH

[Pei71] Rudolf Peierls. Atomic Germans [Heisenberg’s recollections]. New York Review of Books, ??(??):23–24, July 1, 1971. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1992:LHB

[Pei92a] Rudolf Peierls. Letters: Heisenberg’s bomb. New Scientist, 135(1842):??, October 10, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http:

//www.newscientist.com/article/mg13618426.100-letters-
heisenbergs-bomb.html. See [Cha92].

Peierls:1992:USH

[Pei92b] Rudolf Peierls. The uncertain scientist [a Heisenberg biography]. New York Review of Books, ??(??):43–45, April 23, 1992. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1993:SHG

[Pei93a] Rudolf Peierls. Als Student bei Heisenberg. (German) [As a Student of Heisenberg]. In Kleint and Wiemers [KW93], pages 104– 107. ISBN 3-05-501585-1. LCCN QC16.H35 W48 1933. EUR$607 (!).

Peierls:1993:BNW

[Pei93b] Rudolf Peierls. The bomb that never was: [review of Heisenberg’s War: The Secret History of the German Bomb Project, by Thomas Powers (Knopf, New York, 1993)]. New York Review of Books, ??(??):6–9, April 22, 1993. ISSN 0028-7504 (print), 1944-7744 (electronic).

Peierls:1997:AH

[Pei97] Sir Rudolf Ernst Peierls. Atomic Histories, volume 18 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1997. ISBN 1-56396-243-8 (hardcover). xvii + 378 pp. LCCN QC71 .P38 1997.

Perrin:1939:CRC

[Per39] Francis Perrin. Calcul relatif aux conditions éventuelles de transmutation en chaîne de l’uranium. (French) [Calculation of possible conditions on the uranium chain transmutation].

Comptes rendus de l’Académie des sciences, Paris, 208(18):1394–

1396, May 1, 1939. URL http://gallica.bnf.fr/ark:/12148/
bpt6k6238836c/f26.image.r=perrin%20calcul%20relative%
20aux%20conditions#.

Perlman:1964:ESH

[Per64] H. S. Perlman. Equivalence of the Schrödinger and Heisenberg pictures. Nature, 204(4960):771–772, November 21, 1964. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v204/
n4960/pdf/204771b0.pdf.

Perline:1988:SDP

[Per88] Ron Perline. Self-dual polynomials, Hermite matrices, and Heisenberg functionals. SIAM Journal on Matrix Analysis and Applications, 9(3):373–377, ???? 1988. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Pfeiffer:1969:AHW

[Pfe69] Heinrich Pfeiffer, editor. Alexander von Humboldt. Werk und Weltgeltung. (German) [Alexander von Humboldt. Work and international standing]. Piper, München, Germany, 1969. 505 pp. LCCN Q143.H9 A68.

Pfeiffer:1977:DUW

[Pfe77] Heinrich Pfeiffer, editor. Denken und Umdenken: zu Werk und Wirkung von Werner Heisenberg. (German) [Thinking and rethinking: the work and influence of Werner Heisenberg]. Piper, München, West Germany, 1977. ISBN 3-492-02275-8. 279 + 10 pp. LCCN QC16.H35 D46. DM28.00.

Pohl:2020:LWA

[PFH20] Norman Pohl, Michael Farrenkopf, and Friederike Hansell, editors. Lebenswerk Welterbe: Aspekte von Industriekultur und Industriearchäologie, von Wissenschafts- und Technikgeschichte: Festschrift für Helmuth Albrecht zum 65. Geburtstag. (German) [Lifework World Heritage: Aspects of industrial culture and industrial archeology, of the history of science and technology: Festschrift for Helmuth Albrecht on his 65th birthday]. GNT-Verlag, Berlin, Germany, 2020. ISBN 3-86225-534-4, 978-38622-5-1-2-0-9 hardbound. 478 pp. LCCN ????

Phillips:1992:HUB

[Phi92] Melba Phillips. Heisenberg and uncertainty: Book review: Uncertainty: The Life and Science of Werner Heisenberg, by David C. Cassidy. The Physics Teacher, 30(1):39–4, January 1992. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic). URL

http://tpt.aapt.org/resource/1/phteah/v30/i1/p39_s1.

Pauli:1979:WBB

[PHvMW01] Wolfgang Pauli, Armin Hermann, K. von Meyenn, and Victor Frederick Weisskopf, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, ua. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others], volume 2, 6, 11, 15, 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979–2001. ISBN 0-387-08962-4 (vol. 1). ???? pp. LCCN QC16.P37 A34.

Pickover:2008:AHL

[Pic08] Clifford A. Pickover. Archimedes to Hawking: laws of science and the great minds behind them. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-533611-9 (hardcover). ix + 514 pp. LCCN

Q175.32.R45 P53 2008. URL http://catdir.loc.gov/catdir/
toc/ecip087/2007051167.html;   http://www.gbv.de/dms/
goettingen/555633241.pdf;   http://www.oup.com/us/catalog/
general/subject/Physics/?view=usa&ci=9780195336115.

Porsezian:1991:DRS

[PL91] K. Porsezian and M. Lakshmanan. On the dynamics of the radially symmetric Heisenberg ferromagnetic spin system. Journal of Mathematical Physics, 32(10):2923–2928, October 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i10/
p2923_s1.

Plotnitsky:2010:EPB

[Plo10] Arkady Plotnitsky. Epistemology and Probability: Bohr, Heisenberg, Schrödinger, and the Nature of Quantum-Theoretical Thinking, volume 161 of Fundamental theories of physics. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 0-387-85333-2, 0-387-85334-0. xxxiii + 402 pp. LCCN QC173.98 .P56 2010. EUR 123.00, SFR 191. URL http://www.

gbv.de/dms/bowker/toc/9780387853338.pdf.

Papenfuss:2002:YWH

[PLS02] Dietrich Papenfuss, D. (Dieter) Lüst, and Wolfgang Schleich, editors. 100 Years Werner Heisenberg: works and impact. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2002. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). ix + 299 pp. LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Pollock:1995:HIP

[Pol95] M. D. Pollock. On the Heisenberg Indeterminacy Principle and the formation of structure in the universe. Modern Physics Letters A (MPLA), 10(7):539–547, March 7, 1995. CODEN MPLAEQ. ISSN 0217-7323 (print), 1793-6632 (electronic). URL http://www.

worldscientific.com/doi/abs/10.1142/S0217732395000570.

Polizzi:2000:BRW

[Pol00] Gaspare Polizzi. Book review: Werner Heisenberg, Lo sfondo filosofico della fisica moderna, a cura di, Giuseppe Gembillo e Enrico Antonio Giannetto, Palermo, Sellerio, 1999, 210 pp., L. 35.000 (ISBN 88-389-1450-8) (Nuovo Prisma, Serie Filosofia della scienza, 19). Nuncius, 15(1):452–455, ???? 2000. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00650.

Popp:2016:HAW

[Pop16a] Manfred Popp. Hitlers Atombombe — Warum es sie nicht gab. (German) [Hitler’s atomic bomb — why it did not exist]. Spektrum der Wissenschaft (German translation of Scientific American), ??(??):??, September 21, 2016. CODEN

SPEKDI. ISSN 0170-2971. URL https://www.spektrum.de/
news/hitlers-atombombe-warum-es-sie-nicht-gab/1423529.

Popp:2016:HAS

[Pop16b] Manfred Popp. Hitlers Atombombe. Störfall der Wissenschaftsgeschichte. (German) [Hitler’s atomic bomb. Accident in the history of science]. Spektrum der Wissenschaft (German translation of Scientific American), 12(16):12–21, ???? 2016. CO-DEN SPEKDI. ISSN 0170-2971. URL https://www.spektrum.

de/inhaltsverzeichnis/hitlers-atombombe-spektrum-der-
wissenschaft-12-2016/1373070.

Popp:2016:MDI

[Pop16c] Manfred Popp. Misinterpreted documents and ignored physical facts: The history of ‘Hitler’s Atomic Bomb’ needs to be corrected. Berichte zur Wissenschaftsgeschichte, 39(3):265–282, September 2016. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Popp:2018:WHD

[Pop18] Manfred Popp. Werner Heisenberg und das Deutsche Uranprojekt im Dritten Reich. (German) [Werner Heisenberg and the German uranium project in the Third Reich]. In K. Kleinknecht, editor, Quanten 6, pages 11–12. Hirzel, Stuttgart, Germany, 2018. ISBN 3-7776-2754-2. LCCN ????

Popp:2021:WHD

[Pop21] Manfred Popp. Why Hitler did not have atomic bombs. Journal of Nuclear Engineering, 2(1):9–27, March 2021. CODEN ???? ISSN 2673-4362.

Powers:1976:RII

[Pow76] Robert T. Powers. Resistance inequalities for the isotropic Heisenberg ferromagnet. Journal of Mathematical Physics, 17(10):1910– 1918, October 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v17/i10/p1910_s1.

Powers:1993:HKG

[Pow93a] Thomas Powers. Heisenbergs Krieg: die Geheimgeschichte der deutschen Atombombe. (German) [Heisenberg’s War: the secret history of the German atom bomb]. Hoffmann und Campe, Hamburg, Germany, 1993. ISBN 3-455-08479-6. 767 pp. LCCN QC16.H35 P6915 1993.

Powers:1993:HWS

[Pow93b] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Alfred A. Knopf, New York, NY, USA, 1993. ISBN 0-394-51411-4. xi + 607 pp. LCCN QC16.H35 P69 1993. US$27.50, CAN$34.50.

Powers:1994:HWS

[Pow94] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Back Bay Books, Boston, MA, USA, 1994. ISBN 0-316-71623-5. xi + 607 pp. LCCN QC16.H35 P69 1994. US$16.95.

Powers:2000:HWS

[Pow00] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 2000. ISBN 0-306-81011-5 (paperback). xi + 607 pp. LCCN QC16.H35 P69 2000.

Powers:2002:LEH

[Pow02] Thomas Powers. Letter to the Editor: Heisenberg and Bohr. New York Times, ??(??):A18, February 16, 2002. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/hnpnewyorktimes/docview/92195480/
13F386CFDB17FA2143.

Powers:2003:LCP

[Pow03] Thomas Powers. A letter from Copenhagen. New York Review of Books, 50(13):55–56, August 14, 2003. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://werner-

heisenberg.unh.edu;   http://www.nybooks.com/articles/
archives/2003/aug/14/a-letter-from-copenhagen/. Previ-

ously unpublished letter from German physicist Werner Heisenberg to his wife Elisabeth about his meeting with Niels Bohr in September 1941.

Pauli:1955:NBD

[PRW55] W. Pauli, L. Rosenfeld, and V. Weisskopf, editors. Niels Bohr and the development of physics: essays dedicated to Niels Bohr on the occasion of his 70th birthday. McGraw-Hill, New York, NY, USA, 1955. vii + 195 pp. LCCN QC71 .P3 1955.

Percival:1969:APM

[PS69] Ian C. Percival and H. Henry Stroke. Atomic physicists meet for Arnold Sommerfeld centennial. Physics Today, 22(2):99–107, February 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Pu:2002:AOB

[PSZM02] Han Pu, Chris Search, Weiping Zhang, and Pierre Meystre. Atom optics from de Broglie waves to Heisenberg ferromagnets. In Papenfuss et al. [PLS02], pages 235–240. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Puga:1980:TEH

[Pug80] M. W. Puga. Thermal excitations in Heisenberg-xy systems. Journal of Mathematical Physics, 21(8):2307–2312, August 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v21/i8/p2307_s1.

Pauli:1993:WBB

[PvM93] Wolfgang Pauli and Karl von Meyenn, editors. Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band III: 1940– 1949. (German) [Scientific correspondence with Bohr, Einstein, Heisenberg, and others. Volume 3: 1940–1949]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-78802-6. ???? pp. LCCN ????

Pyenson:1981:BRB

[Pye81] Lewis Pyenson. Book review: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Vol. I: 1919– 1929 by Wolfgang Pauli; A. Hermann; K. V. Meyenn; V. F. Weisskopf. Isis, 72(3):524–525, September 1981. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/230307.

Rabinowitch:1948:CE

[Rab48] Eugene Rabinowitch. Comment by the Editor. Bulletin of the Atomic Scientists, 4(4):104–105, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor48, vL48b].

Radkau:1983:AKD

[Rad83] Joachim Radkau. Aufstieg und Krise der deutschen Atomwirtschaft: 1945–1975; verdrängte Alternativen in der Kerntechnik und der Ursprung der nuklearen Kontroverse. (German) [The Rise and Crisis of the German Nuclear Industry 1945–1975: Repressed alternatives in nuclear technology and the origin of the nuclear controversy], volume 7756 of Rororo. Rowohlt-Taschenbuch-Verlag, Reinbek bei Hamburg, West Germany, 1983. ISBN 3-499-17756-0. 583 pp. LCCN ????

Radkau:1991:AKD

[Rad91] Joachim Radkau. Aufstieg und Krise der deutschen Atomwirtschaft 1945–1975: verdrängte Alternativen in der Kerntechnik und der Ursprung der nuklearen Kontroverse. (German) [The Rise and Crisis of the German Nuclear Industry 1945–1975: Repressed alternatives in nuclear technology and the origin of the nuclear controversy]. Repro Pfeffer, Niedernberg, Germany, 1991. 583 pp. LCCN ????

Rashid:1978:ISG

[Ras78] M. A. Rashid. The intelligent states. I. Group-theoretic study and the computation of matrix elements. Journal of Mathematical Physics, 19(6):1391–1396, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v19/i6/p1391_s1.

Razavy:2011:HQM

[Raz11] Mohsen Razavy. Heisenberg’s quantum mechanics. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2011. ISBN 981-4304-11-5 (paperback), 981-4304-10-7. xix + 657 pp. LCCN QC174.12 .R39 2011.

Rutherford:1930:RRS

[RCE30] Sir Ernest Rutherford, James Chadwick, and C. D. (Charles Drummond) Ellis. Radiations from radioactive substances. Cambridge University Press, Cambridge, UK, 1930. xi + 1 + 588 pp. LCCN QC721 .R94.

Rozema:2012:PNV

[RDM^{+}12a] Lee A. Rozema, Ardavan Darabi, Dylan H. Mahler, Alex Hayat, Yasaman Soudagar, and Aephraim M. Steinberg. Publisher’s note: Violation of Heisenberg’s Measurement-Disturbance Relationship by Weak Measurements [Phys. Rev. Lett. 109, 100404 (2012)]. Physical Review Letters, 109(18):189902, November 2012. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.109.189902. See [RDM^{+}12b].

Rozema:2012:VHM

[RDM^{+}12b] Lee A. Rozema, Ardavan Darabi, Dylan H. Mahler, Alex Hayat, Yasaman Soudagar, and Aephraim M. Steinberg. Violation of Heisenberg’s measurement-disturbance relationship by weak measurements. Physical Review Letters, 109(10):100404:1–100404:5, September 2012. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/

doi/10.1103/PhysRevLett.109.100404. See [Ano12] and correc-

tion [RDM^{+}12a].

Rotger:1995:AFU

[RDP^{+}95] A. Rötger, E. Dreisigacker, G. Zu Putlitz, E. Dreisigacker, G. Fasol, A. Rötger, G. Schäfer, A. Rötger, H. Rollnik, G. Röpke, and D. Shirkov. Aktuelles: Fachhochschulen, Umwelt- und Informationstechnik: Forschungsminister Rüttgers stellt Schwerpunkte seiner künftigen Arbeit vor/Wir fangen an, relevant zu werden: Feier im Magnus-Haus und Festkolloquium in der TU Berlin zum 150. Jahrestag der Gründung der Physikalischen Gesellschaft/Die moderne Universität Wie wird sie zu einem Centre of Excellence ?: Bericht von der Nature’s 4th International Conference in Japan,

12. 13. Dezember 1994 in Tokio/Auseinandergelebt: Aktionspapier des Zentralverbandes Elektrotechnik und Elektronikindustrie benennt Themen für Innovationsförderung/1994: Starke Umsatzzuwächse auf dem Halbleitermarkt/Röntgenspiegel aus Metall-Keramik-Multilagen/Neues von den Fullerenen Schlitten auf Nano-Kugellager/Bislang stärkster Hinweis auf Existenz eines massereichen Schwarzen Lochs/Neuer Plasmazustand entdeckt: Fusionsexperiment ASDEX upgrade gelingt sanfte Energieabfuhr/ Poröses Silizium Vorsicht geboten/Drei Jahre Heisenberg–Landau-Programm. Physikalische Blätter, 51(3):152–190, March 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19950510303/abstract.

Rechenberg:1981:UTE

[Rec81a] Helmut Rechenberg. Unified theories of elementary particles: Critical assessment and prospects: International symposium in memory of Werner Heisenberg. Physikalische Blätter, 37(11): 347–348, November 1981. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19810371108/abstract.

Rechenberg:1981:WHK

[Rec81b] Helmut Rechenberg. Werner Heisenberg und das Kaiser-Wilhelm- (Max-Planck-)Institut für Physik. Physikalische Blätter, 37(12): 357–364, December 1981. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19810371206/abstract.

Rechenberg:1984:WHL

[Rec84] H. Rechenberg. Werner Heisenberg as lecturer and author of scientific review articles and books. In Blum et al. [BDR84e], pages 17– 20. ISBN 3-642-61743-3 (print), 3-642-61742-5 (e-book). LCCN

???? URL http://link.springer.com/chapter/10.1007/978-
3-642-61742-3_2/.

Rechenberg:1985:AFI

[Rec85a] H. Rechenberg. Aktuelles: Fraunhofer-Institut für Lasertechnologie/Werner Heisenberg eine Diskussion/Neutronen für die Industrie/Akademie der Wissenschaften in West-Berlin geplant/ Strahlenbelastung in Wohnhäusern. Physikalische Blätter, 41 (2):50–56, February 1985. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19850410214/abstract.

Rechenberg:1985:BRB

[Rec85b] Helmut Rechenberg. Book review: Werner Heisenberg: A Bibliography of His Writings by David Cassidy; Martha Baker. Isis, 76(3):400–401, September 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/232876.

Rechenberg:1986:BRB

[Rec86] H. Rechenberg. Book review: Scientific Correspondence with Bohr, Einstein, Heisenberg and Others. Volume II: 1930–1939 by Wolfgang Pauli; Karl von Meyenn; Armin Hermann; Victor F. Weisskopf. Isis, 77(2):387–388, June 1986. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/232721.

Rechenberg:1991:ELMa

[Rec91a] H. Rechenberg. Ehrung für Lise Meitner und Werner Heisenberg: Aufnahme der Büsten von Lise Meitner und Werner Heisenberg in den Ehrensaal des Deutschen Museums. (German) [Tribute to Lise Meitner and Werner Heisenberg: Inauguration of the busts of Lise Meitner and Werner Heisenberg in the Hall of Honor of the German Museum]. Physikalische Blätter, 47(8):773, August 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19910470810/abstract.

Rechenberg:1991:ELMb

[Rec91b] H. Rechenberg. Ehrung für Lise Meitner und Werner Heisenberg: Aufnahme der Büsten von Lise Meitner und Werner Heisenberg in den Ehrensaal des Deutschen Museums. (German) [Tribute to Lise Meitner and Werner Heisenberg: Inauguration of the busts of Lise Meitner and Werner Heisenberg in the Hall of Honor of the German Museum]. Physik Journal, 47(8):773, August 1991. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

Rechenberg:1991:EWH

[Rec91c] H. Rechenberg. Erinnerungen an Werner Heisenberg. (German) [Remembrances of Werner Heisenberg]. Physikalische Blätter, 47(12):1078–1081, December 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19910471214/
abstract.

Rechenberg:1991:HGG

[Rec91d] H. Rechenberg. Heisenberg, Goudsmit and the German A-Bomb. Physics Today, 44(5):91–92, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v44/i5/p13_s1.

Rechenberg:1992:WHD

[Rec92] Helmut Rechenberg. Werner Heisenberg: Deutsche und jüdische Physik. (German) [Werner Heisenberg: German and Jewish physics]. Piper, Munich, Germany, 1992. ISBN 3-492-11676-0. 211 pp. LCCN ????

Rechenberg:1993:WHB

[Rec93] Helmut Rechenberg. Werner Heisenberg — bibliography. In Blum et al. [BDR93], pages 645–700. ISBN 3-540-13848-X. Dated 1992. See also [CB84].

Rechenberg:1996:WHF

[Rec96] Helmut Rechenberg. Werner Heisenberg und das Forschungsprogramm des Kaiser-Wilhelm-Instituts für Physik (1940–1948). (German) [Werner Heisenberg and the research program of the Kaiser Wilhelm Institute for Physics (1940–1948)]. In vom Brocke and Laitko [vBL96a], pages 245–262. ISBN 3-11-015483-8 (vol. 1). LCCN Q180.G42 K35 1996. URL http://

www.degruyter.com/viewbooktoc/product/146831;   http:/
/www.reference-global.com/doi/book/10.1515/9783110802443.

Rechenberg:2001:GQG

[Rec01] Helmut Rechenberg. Das große Quantenei: Zum 100. Geburtstag von Werner Heisenberg (1901–1976). (German) [The big quantum egg: on the 100th birthday of Werner Heisenberg]. Physikalische Blätter, 57(12):59–63, December 2001. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.20010571218/
abstract.

Rechenberg:2002:AKW

[Rec02a] Helmut Rechenberg. Vom Atomkern zum kosmischen Wirbel Dem Physiker, Philosophen und Friedensforscher Carl Friedrich von Weizsäcker zum 90. Geburtstag mit herzlichen Wünschen gewidmet. (German) [From the nucleus to the cosmic vortex: The physicist, philosopher and peace researcher, Carl Friedrich von Weizsäcker’s 90th birthday, dedicated with best wishes]. Physik Journal, 1(6):59–61, ???? 2002. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-

physik.de/details/articlePdf/1108561/issue.html.

Rechenberg:2002:WHS

[Rec02b] Helmut Rechenberg, editor. Werner Heisenberg (1901–1976): Schritte in die neue Physik; [Begleitbuch zur Ausstellung zum 100. Geburtstag von Werner Heisenberg]. (German) [Werner Heisenberg (1901–1976): Steps into new physics [Introductory book for the exhibition for the 100th birthday of Werner Heisenberg]], volume 2 of Veröffentlichungen des Universitätsarchivs Leipzig. Sax-Verl., Beucha, Germany, second edition, 2002. ISBN 3-934544-25-8. 152 pp. LCCN ???? EUR 10.00,

SFR 19.00.  URL http://www.gbv.de/du/services/agi/
DADD826AB2E19CF5C1256D2E004F95CC/420000071853.

Rechenberg:2003:BHL

[Rec03] Helmut Rechenberg. Werner Heisenberg — Leben und Werk, von D. C. Cassidy. Spektrum, Heidelberg 1995. 808 S., 19 Abb., gebunden, ISBN 3-86025-315-8. Physik Journal, 2(10): ??, September 18, 2003. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic). URL http://www.pro-

physik.de/details/rezension/1109153/Werner_Heisenberg_
-_Leben_und_Werk.html. See [Cas95].

Rechenberg:2009:WHS

[Rec09] Helmut Rechenberg, editor. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die “fröhliche Wissenschaft” (Jugend bis Nobelpreis). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-69221-5 (print), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN

QC16.H35 2009.  URL http://ebooks.ciando.com/book/
index.cfm/bok\_id/46829;   http://www.ciando.com/img/
books/width167/3540692223\_k.jpg;   http://www.ciando.
com/pictures/bib/3540692223bib\_t\_1.jpg.

Rechenberg:2010:WHS

[Rec10] Helmut Rechenberg. Werner Heisenberg — die Sprache der Atome: Leben und Wirken — eine wissenschaftliche Biographie: die ,,fröhliche Wissenschaft“(Jugend bis Nobelpreis). (German) [Werner Heisenberg — the language of atoms: Life and Work — a scientific biography: the “Happy Science” (youth to Nobel Prize)]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-69221-5 (hardcover), 3-540-69222-3 (e-book). xxi + 1001 pp. LCCN QC16.H35 R434 2010. URL http://ebooks.ciando.

com/book/index.cfm/bok_id/46829;   http://www.ciando.
com/img/books/width167/3540692223_k.jpg;   http://www.
ciando.com/pictures/bib/3540692223bib_t_1.jpg.

Reed:2019:NGU

[Ree19] B. Cameron Reed. A note on German uranium stores. Physics Today, 72(11):10, November 2019. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL https://

physicstoday.scitation.org/doi/pdf/10.1063/PT.3.4330.

Reed:2024:HB

[Ree24] B. Cameron Reed. On Heisenberg and the bomb. American Journal of Physics, 92(12):908–909, December 2024. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). This short paper corrects some errors in [McC24] and supplies additional historical material.

Rechenberg:1991:BBL

[RFS^{+}91] H. Rechenberg, A. Förster, W. Stamm, D. Stauffer, H. Rauch, W. Weppner, H. W. Schrötter, and W. Zinth. Buchbesprechung: Broglie, Lochak: Heisenberg’s Uncertainties and the Probabilistic Interpretation of Wave Mechanics/Spatschek: Theoretische Plasmaphysik Eine Einführung/Gaisser: Cosmic Rays and Particle Physics/Hüfner: Nur ein Störfall /Weisbuch: Complex Systems Dynamics/Audretsch u. Mainzer: Wieviele Leben hat Schrödingers Katze?/Herder Lexikon Physik/Kudo u. Fueki: Solid State Ionics/Kovetz: The Principles of Electromagnetic Theory/ Brückner, Feller u. Grummt: Applications of Time-Resolved Optical Spectroscopy. Physikalische Blätter, 47(9):857–859, September 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19910470915/abstract.

Riewe:1972:QDH

[RG72] F. Riewe and A. E. S. Green. Quantum dynamics of higherderivative fields. Journal of Mathematical Physics, 13(9):1368– 1374, September 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i9/p1368_s1.

Reiprich:2002:WHG

[RH02] Kurt Reiprich and Werner Heisenberg, editors. Werner Heisenberg: zum 100. Geburtstag. (German) [Werner Heisenberg: 100th Birthday]. Sonderheft // Rohrbacher Kreis. Rosa-Luxemburg-Stiftung Sachsen, Leipzig, Germany, 2002. ISBN 3-89819-116-8. 64 pp. LCCN ????

Rheault:1991:ADP

[Rhe91] Paul Rheault. Autour du principe de realitè physique quantique: Études sur Planck, Bohr et Heisenberg. M.A. dissertation, Université Laval, Sherbrooke, Quebec, Canada, 1991. 104 pp.

Rhodes:1986:MAB

[Rho86] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 1986. ISBN 0-671-44133-7 (paperback), 0-671-65719-4 (hardcover). 886 + 42 pp. LCCN QC773 .R46 1986.

Rigden:1987:EHF

[Rig87] John S. Rigden. Editorial: Heisenberg, February 1927, and physics. American Journal of Physics, 55(2):107, February 1987. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/55/107/1.

Rigden:2002:HEE

[Rig02] John S. Rigden. Hydrogen: the essential element. Harvard University Press, Cambridge, MA, USA, 2002. ISBN 0-674-00738-7. 280 pp. LCCN QD181.H1 R54 2002. URL http://catdir.loc.

gov/catdir/toc/fy022/2001051708.html.

Ringer:1961:GUC

[Rin61] Fritz Franz Klaus Ringer. The German universities and the crisis of learning, 1918–1925. Ph.D. dissertation, Harvard University, Cambridge, MA, USA, 1961. URL http://search.proquest.

com/docview/302053847?.

Ringer:1969:DGM

[Rin69] Fritz K. Ringer. The decline of the German mandarins: the German academic community, 1890–1933. Harvard University Press, Cambridge, MA, USA, 1969. 528 pp. LCCN LA727 .R47.

Ringer:1990:DGM

[Rin90] Fritz K. Ringer. The decline of the German mandarins: the German academic community, 1890—1933. University Press of New England, Hanover, NH, USA, 1990. ISBN 0-8195-6235-1. xiv + 528 pp. LCCN LA727 .R47 1990.

Rechenberg:2005:WHI

[RKW05] Helmut Rechenberg, Christian Kleint, and Gerald Wiemers, editors. Werner Heisenberg im Spiegel seiner Leipziger Schüler und Kollegen. (German) [Werner Heisenberg as seen by his Leipzig pupils and colleagues]. Leipziger Uni-Vlg, Leipzig, Germany, 2005. ISBN 3-86583-079-X (paperback). 235 pp. LCCN ???? EUR 24.00.

Reis:2009:NSD

[RMRP09] Helton J. Reis, Marat A. Mukhamedyarov, Albert A. Rizvanov, and Andras Palotas. In the name of science: Don’t tamper with the deceptive truth. Mini Reviews in Medicinal Chemistry, 9(14): 1676–1677, December 2009. CODEN MMCIAE. ISSN 1389-5575 (print), 1875-5607 (electronic). URL http://www.eurekaselect.

com/86040/article.

Robinson:1969:TEV

[Rob69] M. C. Robinson. A thought experiment violating Heisenberg’s uncertainty principle. Canadian Journal of Physics = Journal canadien de physique, 47(9):963–967, ???? 1969. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). See remarks [BCL^{+}69].

Robinson:1980:PNT

[Rob80] M. C. Robinson. Proposed new test of Heisenberg Principle. Journal of Physics A (Mathematical and General), 13(3):877–880, March 1980. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Robinson:1982:RCP

[Rob82] M. C. Robinson. Reply to comments on ‘On a proposed new test of Heisenberg’s principle’. Journal of Physics A (Mathematical and General), 15(10):3379–3380, October 1982. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rogers:2013:NDY

[Rog13] J. D. Rogers. The neutron’s discovery — 80 years on. Physics Procedia, 43:1–9, 2013. CODEN PPHRCK. ISSN 1875-3892. URL

http://adsabs.harvard.edu/abs/2013PhPro..43....1R.

Rohrlich:1985:BRB

[Roh85] Fritz Rohrlich. Book review: Inner Exile, Recollections of a Life with Werner Heisenberg, by Elisabeth Heisenberg. American Journal of Physics, 53(3):286, March 1985. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/?AJP/53/286/1.

Rosenfeld:1960:HPP

[Ros60] L. Rosenfeld. Heisenberg, physics and philosophy. Nature, 186 (4728):830–831, June 11, 1960. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v186/n4728/pdf/186830a0.pdf.

Rosenfeld:1979:BRR

[Ros79] Léon Rosenfeld. Berkeley Redivivus (review of W. Heisenberg’s Natural Law and the Structure of Matter). In Cohen and Stachel [CS79], chapter II.23, pages 686–687. ISBN 90-277-0651-4, 90-277-0652-2 (paperback), 94-009-9349-8 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 21 QC7. URL https://link.springer.

com/book/10.1007/978-94-009-9349-5.

Rose:1998:HNA

[Ros98] Paul Lawrence Rose. Heisenberg and the Nazi atomic bomb project: a study in German culture. University of California Press, Berkeley, CA, USA, 1998. ISBN 0-520-21077-8, 0-585-32190-6 (e-book). xx + 352 pp. LCCN QC16.H35 R67 1998. URL

http://preterhuman.net/texts/religion.occult.new_age/
occult.conspiracy.and.related/Rose,%20Paul%20Lawrence%
20-%20Heisenberg%20and%20the%20Nazi%20Atomic%20Bomb%20Project.
pdf;   http://www.scribd.com/doc/176628605/Rose-Paul-
Lawrence-Heisenberg-and-the-Nazi-Atomic-Bomb-Project.

Rose:2000:FCP

[Ros00] Paul Lawrence Rose. Frayn’s Copenhagen plays well, at history’s expense. The Chronicle of Higher Education, ??(??): ??, May 5, 2000. ISSN 0009-5982 (print), 1931-1362 (elec-

tronic). URL https://www.chronicle.com/article/Frayns-
Copenhagen-Plays/22341.

Roth:1993:HUP

[Rot93] Wolff-Michael Roth. Heisenberg’s Uncertainty Principle and interpretive research in science education. Journal of Research in Science Teaching, 30(7):669–680, 1993. CODEN JRSTAR. ISSN 0022-4308 (print), 1098-2736 (electronic).

Rovelli:2020:H

[Rov20] Carlo Rovelli. Helgoland, volume 756 of Piccola biblioteca Adelphi. Adelphi edizioni, Milano, Italy, 2020. ISBN 88-459-3505-1. 227 pp. LCCN QC174.123 .R68 2020.

Rovelli:2021:HMS

[Rov21] Carlo Rovelli. Helgoland: Making Sense of the Quantum Revolution. Riverhead Books, New York, NY, USA, 2021. ISBN 0-593-32888-4 (hardcover), 0-593-32890-6 (e-book). xviii + 235 pp. LCCN QC173.96 .R6813 2021. Translation to English by Erica Segre and Simon Carnell.

Rowlands:2009:LC

[Row09] Peter Rowlands. The Liverpool cyclotrons. IOP History of Physics Newsletter, 25:31–46, February 2009. ISSN 1756-

168X. URL http://www.iop.org/activity/groups/subject/
hp/newsletter/archive/file_64509.pdf.

Roychoudhuri:1978:HMM

[Roy78] Chandrasekhar Roychoudhuri. Heisenberg microscope — misleading illustration. Foundations of Physics, 8(11–12):845–849, December 1978. CODEN FNDPA4. ISSN 0015-9018 (print),

1572-9516 (electronic). URL http://www.springerlink.com/
content/t58j8k8m73666lu7/.

Rozental:1967:NBH

[Roz67] S. (Stefan) Rozental, editor. Niels Bohr: his life and work as seen by his friends and colleagues. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. 355 pp. LCCN QC16.B63 N53.

Rosenfeld:1964:NBH

[RRK^{+}64] Léon Rosenfeld, Erik Rüdinger, Oskar Klein, Werner Heisenberg Hendrik G. B. Casimir, Otto Robert Frisch, Stefan Rozental, Aage Bohr, Abraham Pais, Jørgen Kalckar, Christian Møller, Mogens Pihl, Viktor F. Weisskopf, Johannes Pedersen, Viggo Kampmann, Richard Courant, Paul A. M. Dirac, Hans Henrik Koch, William Scharff, Mogens Andersen, Hans Bohr, and Niels Bohr. Niels Bohr: Hans liv or virke fortalt af en kreds af venner og medarbejdere. (Danish) [Niels Bohr: His life and works told by a group of friends and co-workers]. J. H. Schultz Forlag, Copenhagen, DK, 1964. 341 pp. LCCN ????

Ron:1994:FSB

[RS94] Amos Ron and Zuowei Shen. Frames and stable bases for subspaces of L_{2}(R^{d}) the duality principle of Weyl–Heisenberg sets. In Brown et al. [BCEP94], pages 422–425. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Rigden:2011:BNG

[RS11a] John S. Rigden and Roger H. Stuewer. Book notes: [Galileo’s Two Chief World Systems and Two New Sciences; Harald Fritzsch, You Are Wrong, Mr Einstein!]. Physics in Perspective (PIP), 13(3): 373–375, September 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0067-4.

Rogel-Salazar:2011:HQM

[RS11b] J. Rogel-Salazar. Heisenberg’s Quantum Mechanics, by Mohsen Razavy, Scope: advanced textbook. Level: advanced undergraduate. Contemporary Physics, 52(6):620–621, 2011. CODEN CT-PHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Rudner:1958:BRB

[Rud58] Richard Rudner. Book review: Physics and Philosophy. The Revolution in Modern Science, Werner Heisenberg. Science, 128 (3325):649–652, September 19, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://labs.

adsabs.harvard.edu/ui/abs/1958Sci...128..649H;   http:
//www.sciencemag.org/content/128/3325/649.

Rurup:1979:WGB

[Rür79] Reinhard Rürup, editor. Wissenschaft und Gesellschaft: Beiträge zur Geschichte der Technischen Universität Berlin, 1879–1979. (German) [Science and Society: Contributions to the History of the Technical University of Berlin, 1879–1979]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-09672-8, 3-540-09672-8. xv + 610 (vol. 1), vi + 273 (vol. 2) pp. LCCN T173 .B51938 1979.

Rushbrooke:1964:TRD

[Rus64] G. S. Rushbrooke. On the theory of randomly dilute Ising and Heisenberg ferromagnetics. Journal of Mathematical Physics, 5 (8):1106–1116, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v5/i8/p1106_s1.

Rose:1992:DHM

[RW92] Paul Lawrence Rose and Mark Walker. Did Heisenberg misconceive A-Bomb? Physics Today, 45(2):126–??, February 1992. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v45/i2/p126/s1.

Renneberg:1994:STN

[RW94] Monika Renneberg and Mark Walker, editors. Science, Technology, and National Socialism. Cambridge University Press, Cambridge, UK, 1994. ISBN 0-521-40374-X (hardcover), 0-521-52860-7 (paperback). xix + 422 pp. LCCN Q127.G3 S36 1994. URL http:

//www.loc.gov/catdir/description/cam025/92041633.html;
 http://www.loc.gov/catdir/samples/cam031/92041633.html;
 http://www.loc.gov/catdir/toc/cam028/92041633.html.

Rechenberg:2002:WHG

[RW02] Helmut Rechenberg and Gerald Wiemers, editors. Werner Heisenberg: Gutachten- und Prüfungsprotokolle für Promotionen und Habilitationen (1929–1942). (German) [Werner Heisenberg: Appraisal and test certificates for graduations and habilitations], volume 29 of Berliner Beiträge zur Geschichte der Naturwissenschaften und der Technik. ERS Verlag, Berlin, Germany, second edition, 2002. ISBN 3-928577-43-3. 266 pp. LCCN QC16.H35 W475 2002.

Sallhofer:2001:PFD

[Sal01] Hans H. Sallhofer. Der Physikerstreit: ein fiktiver Diskurs mit Schrödinger und Heisenberg. (German) [The physicist dispute: a fictional discourse between Schrödinger and Heisenberg]. Universitas, München, Germany, 2001. ISBN 3-8004-1425-2. 312 pp. LCCN QC174.12 .S2555 2001.

Salvia:2019:ECP

[Sal19] Stefano Salvia. Embattled cooperation(s): Peaceful atoms, pacifist physicists, and partisans of peace in the Early Cold War (1947–1957). Physics in Perspective (PIP), 21(1):43–62, March 2019. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Santilli:1981:GHU

[San81] R. M. Santilli. Generalization of Heisenberg Uncertainty Principle for strong interactions. Hadronic Journal, 4(3):642–657, ???? 1981. CODEN HAJODX. ISSN 0162-5519.

Sanders:2002:HGB

[San02] T. M. Sanders. Heisenberg and the German bomb. Contemporary Physics, 43(5):401–403, 2002. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Saperstein:2010:BRD

[Sap10] Alvin M. Saperstein. Book review: David C. Cassidy, Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb. Physics Today, 63(1):49–??, January 2010. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v63/i1/p49/s1.

Schliesser:2009:RSC

[SAR^{+}09] A. Schliesser, O. Arcizet, R. Rivière, G. Anetsberger, and T. J. Kippenberg. Resolved-sideband cooling and position measurement of a micromechanical oscillator close to the Heisenberg uncertainty limit. Nature Physics, 5(7):509–514, June 7, 2009. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (elec-

tronic). URL http://www.nature.com/nphys/journal/v5/n7/
full/nphys1304.html.

Schrodinger:1926:VHB

[Sch26] Erwin Schrödinger. Über das Verhältnis der Heisenberg–Born– Jordanschen Quantenmechanik zu der meinem. (German) [On the relation of Heisenberg–Born–Jordan quantum mechanics to mine]. Annalen der Physik (1900), 384(8):734–756, ???? 1926. ISSN 1521-3889.

Schrodinger:1927:AWG

[Sch27] Erwin Schrödinger. Abhandlungen zur Wellenmechanik. (German) [Essays on wave mechanics]. Johann Ambrosius Barth Verlag, Leipzig, Germany, 1927. ix + 169 pp. LCCN QC174.2 .S3 1927.

Schrodinger:1930:HUG

[Sch30] Erwin Schrödinger. Zum Heisenbergschen Unschärfeprinzip. (German) [On Heisenberg’s Uncertainty Principle]. Ständiger Beobachter der Preussischen Akademie der Wissenschaften, Phys.-math. Klasse, Sitzungsberichte, ??(??):296–303, ???? 1930. CODEN SPWPAI. ISSN ????

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schlegel:1959:BRW

[Sch59] Richard Schlegel. Book review: Werner Heisenberg (author) and Arnold J. Pomerans (translator), The Physicist’s Conception of Nature. Physics Today, 12(6):54–??, June 1959. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v12/i6/p54/s1.

Schrodinger:1966:HBJ

[Sch66] Erwin Schrödinger. A Heisenberg–Born–Jordan-féle kvantummechanika viszonya az enyémhez. (Hungarian) [The relation of Heisenberg–Born–Jordan quantum mechanics to mine]. Magyar Fizikai Folyóirat, 14(??):359–374, ???? 1966. CODEN MGFFAC. ISSN 0025-0104.

Stockmann:1978:BFB

[SCH^{+}78] F. Stöckmann, M. Cardona, K. Hecht, Robert Gerwin, and H. Sachsse. Bücher Filme: Behrens: Physikdaten/Behrens u. Ebel: Physikdaten/Sexl: Physik/Hermann: Die Jahrhundertwissenschaft Werner Heisenberg und die Physik seiner Zeit/ Storck: Einführung in die Philosophie der Technik. Physikalische Blätter, 34(6):289–291, June 1978. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19780340612/
abstract.

Scharf:1982:MWH

[Sch82] Joachim-Hermann Scharf. In memoriam Werner Heisenberg aus Anlaß seines 80. Geburtstages. (German) [In memory of Werner Heisenberg on the occasion of his 80th birthday], volume 55 of Deutsche Akademie der Naturforscher Leopoldina (Halle, Saale): Nova acta Leopoldina / Neue Folge; 248. Barth, Leipzig, East Germany, 1982. 136 pp. LCCN ????

Schweber:1986:ETR

[Sch86] S. S. Schweber. The empiricist temper regnant: Theoretical physics in the United States 1920–1950. Historical Studies in the Physical and Biological Sciences, 17(1):55–98, ???? 1986. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/stable/27757575.

Schroeder:1991:FCP

[Sch91a] Manfred Schroeder. Fractals, Chaos, Power Laws. W. H. Freeman and Company, New York, NY, USA, 1991. ISBN 0-671-74217-5, 0-7167-2136-8, 0-7167-2357-3. xviii + 429 pp. LCCN QD921 .S3 1990.

Schweber:1991:BRB

[Sch91b] S. S. Schweber. Book review: Gesammelte Werke/Collected Works by Werner Heisenberg; W. Blum; H.-P. Dürr; H. Rechenberg. Isis, 82(1):159–160, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/233580.

Schwinger:1994:QUP

[Sch94] Julian Schwinger. Quantum Uncertainty Principle : No loopholes. Science, 264(5167):1830–1831, June 24, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/264/5167/1830.full.pdf.

Schirmacher:1996:GES

[Sch96a] Wolfgang Schirmacher, editor. German essays on science in the 20th century, volume 82 of German library. Continuum, New York, NY, USA, 1996. ISBN 0-8264-0746-3 (hardcover), 0-8264-0747-1 (paperback). xviii + 314 pp. LCCN QC9.G3 G47 1996.

Schroeck:1996:QMP

[Sch96b] F. (Franklin) Schroeck. Quantum mechanics on phase space, volume 74 of Fundamental theories of physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-3794-8. xvi + 668 pp.

LCCN QC174.12 .S37 1996. URL http://www.loc.gov/catdir/
enhancements/fy0822/95025631-d.html;   http://www.loc.
gov/catdir/enhancements/fy0822/95025631-t.html.

Schmudgen:1999:ORD

[Sch99] Konrad Schmüdgen. Operator representations of a q-deformed Heisenberg algebra. Journal of Mathematical Physics, 40(9):4596– 4605, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schrag:2000:HHP

[Sch00a] Peter Schrag. History’s Heisenberg Principle. The American Prospect, 11(18):40–??, November 9, 2000. CODEN APROEY. ISSN 1049-7285. URL http://connection.

ebscohost.com/c/articles/3342279/historys-heisenberg-
principle;   http://prospect.org/article/historys-heisenberg-
principle.

Schwarzschild:2000:BHS

[Sch00b] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i5/
p51/s1.

Schaaf:2001:HHB

[Sch01] Michael Schaaf. Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. (German) [Heisenberg, Hitler and the Bomb: Conversations with witnesses]. Verlag für Geschichte der Naturwissenschaften und der Technik, Berlin, Germany, 2001. ISBN 3-928186-60-4 (paperback). 148 pp. LCCN QC16.H35 S32 2001.

Schmidt:2002:EII

[Sch02a] Andreas U. Schmidt. Erratum: “Infinite infrared regularization and a state space for the Heisenberg algebra”, [J. Math. Phys. 43, 243 (2002)]. Journal of Mathematical Physics, 43(6):3412, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [Sch02b].

Schmidt:2002:IIR

[Sch02b] Andreas U. Schmidt. Infinite infrared regularization and a state space for the Heisenberg algebra. Journal of Mathematical Physics, 43(1):243–259, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [Sch02a].

Schweber:2002:EFQ

[Sch02c] Silvan S. Schweber. Enrico Fermi and quantum electrodynamics, 1929–32. Physics Today, 55(6):31–36, June 2002. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v55/i6/p31_s1.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schirrmacher:2005:DMA

[Sch05] Arne Schirrmacher, editor. Dreier Männer Arbeit in der frühen Bundesrepublik: Max Born, Werner Heisenberg und Pascual Jordan als politische Grenzgänger. (German) [Three men’s work in the early Federal Republic: Max Born, Werner Heisenberg and Pascual Jordan as political border crossers], volume 296 of Preprint / Max-Planck-Institut für Wissenschaftsgeschichte. Max-Planck-Inst. für Wissenschaftsgeschichte, Berlin, Germany, 2005. 50 + 6 pp. LCCN ????

Schirrmacher:2007:PPF

[Sch07] Arne Schirrmacher. Physik und Politik in der frühen Bundesrepublik Deutschland. Max Born, Werner Heisenberg und Pascual Jordan als politische Grenzgänger. (German) [Physics and politics in the early Federal Republic Germany. Max Born, Werner Heisenberg and Jordan as a political frontier]. Berichte zur Wissenschaftsgeschichte, 30(1):13–31, March 2007. CODEN BE-WID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Scharf:2011:JHL

[Sch11a] Rainer Scharf. Jenseits des Heisenberg-Limits. (German) [Beyond the Heisenberg limit]. Physik Journal, 10(??):??, March 24, 2011. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-

6597 (electronic). URL http://www.pro-physik.de/details/
news/1111037/Jenseits_des_Heisenberg-Limits.html. See

[NKD^{+}11].

Schwarz:2011:RND

[Sch11b] Stephan Schwarz. Regimmotstånd? Deutsche Physik-affären. (Swedish) [Regime opposition? German Physics affairs]. KOS-MOS 2011, ??(??):79–105, ???? 2011. URL https://docplayer.

se/108769252-Regimmotstand-deutsche-physikaffaren.html.

Schwarz:2011:STN

[Sch11c] Stephan Schwarz. Science, technology, and the Niels Bohr Institute in Occupied Denmark. Physics in Perspective (PIP), 13(4): 401–432, December 2011. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-011-0059-4.

Schafer:2013:NCW

[Sch13] Wolf Schäfer. Negative charge: What strained the relationship between two of Germany’s most respected scientific thinkers. The Berlin Journal, (25):8–11, Fall 2013. URL https://www.

stonybrook.edu/commcms/sbigs/pdfs/negitiveCharge.pdf.

Schwarz:2015:DKH

[Sch15] Stephan Schwarz. Dramatik kring ett Heisenberg-brev från krigstidens Berlin. (Swedish) [Drama about a Heisenberg letter from wartime Berlin]. KOSMOS 1914, 15(??):109–123, 2015.

Schwarz:2022:DAW

[Sch22] Stephan Schwarz. Drama around a wartime Heisenberg letter. Physics in Perspective (PIP), 24(1):72–92, March 2022. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s00016-021-00285-1.

Scull:1975:HSS

[Scu75] Sidney C. Scull. Heisenberg subgroups of semisimple Lie groups. Journal of Mathematical Physics, 16(8):1646–1648, August 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i8/p1646_s1.

Seidel:1990:BRB

[Sei90a] Robert Seidel. Book reviews: Books on the bomb. Atomic Bomb Scientists: Memoirs, 1939-1945 by Joseph J. Ermenc. The End of the World That Was: Six Lives in the Atomic Age by Peter Goldman. Manhattan: The Army and the Atomic Bomb by Vincent C. Jones. Day of the Bomb: Countdown to Hiroshima by Dan Kurzman. The General and the Bomb: A Biography of General Leslie R. Groves, Director of the Manhattan Project by William Lawren. Time Bomb: Fermi, Heisenberg, and the Race for the Atomic Bomb by Malcolm C. MacPherson. The Making of the Atomic Age by Alwyn McKay. The Road to Trinity: A Personal Account of How America’s Nuclear Policies Were Made by K. D. Nichols. The Making of the Atomic Bomb by Richard Rhodes. Stallion Gate by Martin Cruz Smith. The Atomic Scientists: A Biographical History by Henry A. Boorse. Lloyd Motz, and Jefferson Hane Weaver. Forging the Atomic Shield: Excerpts from the Office Diary of Gordon E. Dean by Gordon E. Dean and Roger M. Anders. The Nuclear Oracles: A Political History of the General Advisory Committee of the Atomic Energy Commission, 1947-1977 by Richard T. Sylves. Better a Shield Than a Sword by Edward Teller. Klaus Fuchs, Atom Spy by Robert Chadwell Williams. Justice Downwind: America’s Atomic Testing Program in the 1950s by Howard Ball. The Atomic Papers: A Citizen’s Guide to Selected Books and Articles on the Bomb, the Arms Race, Nuclear Power, the Peace Movement, and Related Issues by Grant Burns. Physics, Technology and the Nuclear Arms Race by D. W. Hafemeister and D. Schroeer. Under the Cloud: The Decades of Nuclear Testing by Richard L. Miller. Bombs in the Backyard: Atomic Testing and American Politics by A. Costandina Titus. Nuclear Fear: A History of Images by Spencer R. Weart. Isis, 81(3):519–537, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233429.

Seigfried:1990:AQP

[Sei90b] Hans Seigfried. Autonomy and quantum physics: Nietzsche, Heidegger, and Heisenberg. Philosophy of Science, 57(4):619–630, December 1990. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL http://www.jstor.org/stable/187764.

Seife:2005:DDP

[Sei05] Charles Seife. Do deeper principle’s underlie quantum uncertainty and nonlocality? Science, 309(5731):98, July 1, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/309/5731/
98.full.pdf.

Semmes:2003:IHG

[Sem03] Stephen Semmes. An introduction to Heisenberg groups in analysis and geometry. Notices of the American Mathematical Society, 50(6):640–646, June/July 2003. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Serwer:1977:UZP

[Ser77] Daniel Serwer. Unmechanischer Zwang: Pauli, Heisenberg, and the rejection of the mechanical atom, 1923–1925. Historical Studies in the Physical Sciences, 8(??):189–256, ???? 1977. CO-DEN HSPSAS. ISSN 0073-2672. URL http://www.jstor.org/

stable/27757371.

Seth:2010:CQA

[Set10] Suman Seth. Crafting the quantum: Arnold Sommerfeld and the practice of theory, 1890–1926. Transformations. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-01373-8 (hardcover). viii + 378 pp. LCCN QC16.S76 S48 2010.

Seth:2011:HPS

[Set11] Suman Seth. Heisenberg in the “Public sphere”. Science, 331(6020):1012, February 25, 2011. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/331/6020/1012.1.full.pdf.

Suzuki:1971:ZPF

[SF71] Masuo Suzuki and Michael E. Fisher. Zeros of the partition function for the Heisenberg, ferroelectric, and general Ising models. Journal of Mathematical Physics, 12(2):235–246, February 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v12/i2/p235_s1.

Sommerfeld:1922:IMI

[SH22a] A. Sommerfeld and W. Heisenberg. Die Intensität der Mehrfachlinien und ihrer Zeemankomponenten. (German) [The intensity of line multiplets and their Zeeman components]. Zeitschrift für Physik, 11(1):131–154, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01328408;   http://www.
springerlink.com/content/mv9248853766w370/.

Sommerfeld:1922:BRR

[SH22b] A. Sommerfeld and W. Heisenberg. Eine Bemerkung über relativistische Röntgendubletts und Linienschärfe. (German) [A note on relativistic X-ray doublets and line sharpness]. Zeitschrift für Physik, 10(1):393–398, December 1922. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic). URL http://link.

springer.com/article/10.1007/BF01332582;   http://www.
springerlink.com/content/t5l853jr88m10630/.

Szilard:1955:CBA

[SH57] Leo Szilard and Werner Heisenberg. [Correspondence between the authors]. Four letters in English and German in the Leo Szilard Papers archive., 1955–1957. URL http://library.ucsd.edu/

dc/object/bb3278005z.

Salam:1991:UFF

[SHD91] Abdus Salam, W. Heisenberg, and P. Dirac. La unificación de las fuerzas fundamentales. (Spanish) [Unification of fundamental forces]. Gedisa, Barcelona, Spain, 1991. ISBN 84-7432-413-0.

???? pp. LCCN ???? URL http://148.201.96.14/dc/ver.
aspx?ns=000033655;   http://en.scientificcommons.org/

8713570. Notes by J. Evans and G. Watts. Preface by John C. Taylor. Translated to Spanish by Alberto L. Bixio.

Shea:1983:OHR

[She83a] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1584-X. 252 pp. LCCN QC773 .O87 1983. URL http://

link.springer.com/book/10.1007/978-94-009-7133-2.

Shea:1983:OHRb

[She83b] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1584-X. 252 pp. LCCN QC773 .O87 1983. URL http://

link.springer.com/book/10.1007/978-94-009-7133-2.

Sakamoto:1973:CIE

[SHI73] Kazuyoshi Sakamoto, Taisen Hayashi, and Y. J. I’Haya. Core integrals of 2pπ-electron evaluated from Heisenberg’s equation of motion. International Journal of Quantum Chemistry, 7(S7):337– 345, January 21–27, 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Shimony:1978:HD

[Shi78] Abner Shimony. On Heisenberg’s discovery. Science, 199(4325): 168–169, January 13, 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/199/4325/168.2.full.pdf.

Shimony:1983:RPB

[Shi83] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Laudan [CL83], chapter 11, pages 209–221. ISBN 94-009-7055-2, 94-009-7057-9 (e-book). ISSN

0068-0346. LCCN ???? URL https://link.springer.com/
chapter/10.1007/978-94-009-7055-7_11.

Shimony:1986:RPB

[Shi86] Abner Shimony. Reflections on the philosophy of Bohr, Heisenberg, and Schrödinger. In Cohen and Wartofsky [CW85], chapter 20, pages 305–317. ISBN 90-277-1971-3, 94-009-5345-3 (ebook). ISSN 0068-0346. LCCN ???? URL https://link.

springer.com/chapter/10.1007/978-94-009-5345-1_20.

Shifman:2002:HS

[Shi02] Mikhail A. Shifman. From Heisenberg to supersymmetry. In Papenfuss et al. [PLS02], pages 123–132. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15

2002. EUR 119.00. URL http://www.loc.gov/catdir/bios/
wiley046/2003268171.html;   http://www.loc.gov/catdir/
description/wiley0310/2003268171.html;   http://www.loc.
gov/catdir/toc/wiley031/2003268171.html.

Silva:2013:WHT

[Sil13] Ignacio Silva. Werner Heisenberg and Thomas Aquinas on natural indeterminism. New Blackfriars, 94(1054):635–653, 2013. ISSN 0028-4289 (print), 1741-2005 (electronic).

Simon:1979:MMF

[Sim79] Walther C. Simon, editor. Mensch und Musik: Festschrift für Herbert von Karajan. (German) [Man and Music: Festschrift for Herbert von Karajan]. O. Müller, Salzburg, Austria, 1979. ISBN 3-7013-0578-1. 108 pp. LCCN ML3800 .M33 1979.

Sime:1989:LMD

[Sim89] Ruth Lewin Sime. Lise Meitner and the discovery of fission. Journal of Chemical Education, 66(5):373–375, May 1989. CODEN JCEDA8. ISSN 0021-9584 (print), 1938-1328 (electronic). URL

http://pubs.acs.org/doi/abs/10.1021/ed066p373.

Sime:2000:STE

[Sim00] Ruth Lewin Sime. The search for transuranium elements and the discovery of nuclear fission. Physics in Perspective (PIP), 2(1):48–62, March 2000. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.springerlink.

com/content/nr2t13tndn6t9t72/.

Sime:2007:UGS

[Sim07] Ruth Lewin Sime. “Die ‘Uranspaltung’ hat da die ganze Situation gerettet”: Otto Hahn und das Kaiser-Wilhelm-Institut für Chemie im Zweiten Weltkrieg. (German) [“Uranium fission has saved the whole situation”: Otto Hahn and the Kaiser Wilhelm Institute for Chemistry in the Second World War]. In Maier [Mai07], pages 268–304. ISBN 3-8353-0182-9 (paperback). LCCN Q180.G4 K38 2000 v.17; U395.G3 G45 2007.

Sime:2010:IHN

[Sim10] Ruth Lewin Sime. An inconvenient history: the nuclear-fission display in the Deutsches Museum. Physics in Perspective (PIP), 12(2):190–218, June 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.springerlink.

com/content/w716842562715257/.

Sime:2012:PFO

[Sim12] Ruth Lewin Sime. The politics of forgetting: Otto Hahn and the German Nuclear-Fission Project in World War II. Physics in Perspective (PIP), 14(1):59–94, March 2012. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-011-0065-6;
http://www.springerlink.com/content/k12202vg92147h68/.

Seitz:1973:CHS

[SK73] W. A. Seitz and D. J. Klein. Computations on Heisenberg Spin Models. International Journal of Quantum Chemistry, 7(4):647– 665, July 1973. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Skala:2009:HUR

[SK09] L. Skála and V. Kapsa. Heisenberg uncertainty relations can be replaced by stronger ones. International Journal of Quantum Chemistry, 109(8):1626–1630, ???? 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stanley:1970:DRT

[SL70] H. Eugene Stanley and M. Howard Lee. Diagrammatic representation of the two-spin correlation function for the generalized Heisenberg model. International Journal of Quantum Chemistry, 5(S4):407–418, January 19–24, 1970. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory and Quantum Biology.

Slater:1972:SEH

[Sla72] John C. Slater. Statistical exchange and the Heisenberg exchange integral. Journal de physique, 33(C3):7–11, ???? 1972. CODEN JOPQAG. ISSN 0302-0738.

Stark:1941:JDP

[SM41] Johannes Stark and Wilhelm Müller, editors. Jüdische und deutsche Physik: Vorträge zur Eröffnung des Kollegiums für theoretische Physik an der Universität München. (German) [Jewish and German Physics: Lectures at the opening of the College of Theoretical Physics at the University of Munich]. Heling, Leipzig, Germany, 1941. 55 pp. LCCN ????

Ma:2011:QSW

[sMDN^{+}11] Xiao song Ma, Borivoje Dakic, William Naylor, Anton Zeilinger, and Philip Walther. Quantum simulation of the wavefunction to probe frustrated Heisenberg spin systems. Nature Physics, 7(5): 399–405, February 20, 2011. CODEN NPAHAX. ISSN 1745-2473 (print), 1745-2481 (electronic). URL http://www.nature.com/

nphys/journal/v7/n5/full/nphys1919.html.

Stelzer:1977:FQE

[SME77] M. Stelzer, J. Mühlstein, and V. Ernst. A finite quantum electrodynamical base for many photon collective phenomena. Journal of Mathematical Physics, 18(10):1966–1977, October 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v18/i10/
p1966_s1.

Sanchez-Marin:1987:ASH

[SMMM87] J. Sanchez-Marin, J. P. Malrieu, and D. Maynau. Approximate solutions of Heisenberg Hamiltonians. International Journal of Quantum Chemistry, 31(6):903–925, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smorodinskii:1992:EPP

[Smo92a] Ya. A. Smorodinskii. Eopredelennost”: Pobedy i porazhenija Vernera Geĭzenberga. (Russian) [Uncertainty: The triumphs and defeats of Werner Heisenberg]. Uspekhi Fizicheskikh Nauk, 162 (11):201–205, November 1992. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/en/

articles/1992/11/f/. Review of [Cas92c].

Smorodinskii:1992:LWH

[Smo92b] Ya. A. Smorodinskii. The legacy of Werner Heisenberg. Soviet Physics. Uspekhi, 35(1):50–52, January 1992. CODEN SOPUAP. ISSN 0038-5670. Review of [BDR84a, BDR84b].

Smorodinskii:1992:NVG

[Smo92c] Ya. A. Smorodinskii. Nasledie Vernera Geĭzenberga. (Russian) [The legacy of Werner Heisenberg]. Uspekhi Fizicheskikh Nauk, 162(1):141–145, January 1992. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/en/

articles/1992/1/e/.

Smorodinskii:1992:UTD

[Smo92d] Ya. A. Smorodinskii. Uncertainty: The triumphs and defeats of Werner Heisenberg. Soviet Physics. Uspekhi, 35(11):999–1001, November 1992. CODEN SOPUAP. ISSN 0038-5670. URL http:

//ufn.ru/en/articles/1992/11/f/. Review of [Cas92c].

Smorodinski:1993:HDB

[Smo93] Yakov A. Smorodinski. Heisenberg und Dirac: Die Bedeutung des Schönen in der Naturwissenschaft. (German) [Heisenberg und Dirac: The meaning of beauty in science]. Physikalische Blätter, 49(5):436–438, May 1993. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19930490518/abstract.

Snoek:1949:WHT

[Sno49] J. L. Snoek. The Weiss–Heisenberg theory of ferro-magnetism and a new rule concerning magnetostriction and magnetoresistance. Nature, 163(4152):837–838, May 28, 1949. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v163/
n4152/pdf/163837b0.pdf.

Snygg:1982:HPD

[Sny82] John Snygg. The Heisenberg picture and the density operator. American Journal of Physics, 50(10):906–909, October 1982. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/50/906/1.

Sommerfeld:1936:RHW

[Som36] Arnold Sommerfeld. Rezensionen: Heisenberg, Wandlungen in den Grundlagen der Naturwissenschaften; Röntgens Briefe an Zehnder. (German) [Reviews: Heisenberg, Changes in the Fundamentals of Natural Science; Röntgen’s letters to Zehnder]. Physikalische Zeitschrift, 37(??):416–??, ???? 1936. CODEN PHZTAO. ISSN 0369-982X.

Sorokin:2002:HAS

[Sor02] Dmitri Sorokin. The Heisenberg algebra and spin. In Papenfuss et al. [PLS02], pages 295–299. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15

2002. EUR 119.00. URL http://www.loc.gov/catdir/bios/
wiley046/2003268171.html;   http://www.loc.gov/catdir/
description/wiley0310/2003268171.html;   http://www.loc.
gov/catdir/toc/wiley031/2003268171.html.

Spielberg:1998:WHR

[Spi98] Ana Elisa Spielberg. Werner Heisenberg: Reflexiones sobre pragmatismo y positivismo. (Spanish) [Werner Heisenberg: Reflections on pragmatism and positivism]. In The Paidea Project On-Line: Twentieth World Congress of Philosophy, in Boston, Massachusetts from August 10–15, 1998, page ?? ????, ????, 1998.

URL https://www.bu.edu/wcp/Papers/Scie/ScieSpie.htm.

Strubecker:1964:BSE

[SPS^{+}64] K. Strubecker, M. Päsler, F. Stöckmann, G. Grau, G. W. Schumann, A. Moghissi, F. Stockmann, W. Kleen, J. Demny, W. Braunbek, Dore Brüche, W. Finkeinburg, and W. Heiser. Bücher: Schäfke: Einführung in die Theorie der speziellen Funktionen der mathematischen Physik/Greub: Linear Algebra/ Bermant: A Course of Mathematical Analysis/Bopp: Werner Heisenberg und die Physik unserer Zeit/Sproull: Modern Physics/ Sharman: Vibrations and Waves/Heimendahl: Licht und Farbe/ Disposal of Radioactive Wastes into Fresh Water/Taylor: Semiconductor Particle Detectors/Gärtner: Einführung in die Physik des Transistors/Beiträge zur Metallphysik I/Westphal: Kleines Lehrbuch der Physik/Brüche: Physikertagung Stuttgart 1962/ Gray: American Institute of Physics Handbook/Brandt: Forschen und Gestalten/Kromphardt: Vom Zukunftsauftrag der neuen Universität/Kommentar: Deutsches Atomgesetz und Strahlenschutzrecht. Physikalische Blätter, 20(1):36–42, January 1964. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.19640200110/abstract.

Squires:1967:HFT

[Squ67] E. J. Squires. Heisenberg field theory. Nature, 214(5091):949– 950, May 27, 1967. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v214/n5091/pdf/214949b0.pdf.

Stopes-Roe:1949:UP

[SR49] H. V. Stopes-Roe. The Uncertainty Principle. Nature, 164(4162): 245–246, August 6, 1949. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v164/n4162/pdf/164245a0.pdf.

Serber:1992:APF

[SR92] R. (Robert) Serber and Richard Rhodes. The Los Alamos primer: the first lectures on how to build an atomic bomb. University of California Press, Berkeley, CA, USA, 1992. ISBN 0-520-07576-5. xxxiii + 98 + 8 pp. LCCN QC773.A1 S47 1992. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http://
www.loc.gov/catdir/bios/ucal051/91014068.html;   http:
//www.loc.gov/catdir/description/ucal041/91014068.html.

Sauer:2016:PHC

[SS16] Tilman Sauer and Raphael Scholl, editors. The Philosophy of Historical Case Studies, volume 319 of Boston Studies in the Philosophy and History of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346. viii + 296 pp.

LCCN Q126.8 .P45 2016. URL https://link.springer.com/
book/10.1007/978-3-319-30229-4.

Simon:1965:ESH

[ST65] J. M. Simon and J. G. Taylor. Existence of the Schrödinger and Heisenberg pictures. Nature, 205(4978):1305–1306, March 27, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v205/
n4978/pdf/2051305a0.pdf.

Stark:1936:PLD

[Sta36] Johannes Stark. Philipp Lenard als deutscher Naturforscher. (German) [Philipp Lenard as German natural scientist]. Nationalsozialistische Monatshefte, 71(??):106–111, January 1936.

Stark:1937:WJW

[Sta37] Johannes Stark. ‘Weiße Juden’ in der Wissenschaft. (German) [‘White Jews’ in science]. Das Schwarze Korps, ??(??):??, July

15, 1937. URL https://uni-tuebingen.de/fileadmin/Uni_
Tuebingen/Fakultaeten/MathePhysik/Institute/IAP/Forschung/
MOettel/Geburt_QM/stark_weisse_juden_1937.html.

Stachel:2009:GC

[Sta09a] John J. Stachel. Going Critical: Volume One: The Challenge of Practice, volume 201 of Boston Studies in the Philosophy of Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-1308-6 (hardcover). 407 (est.) pp. LCCN ????

Stapp:2009:MM

[Sta09b] Henry Stapp. Matrix mechanics. In Greenberger et al. [GHW09], pages 368–371. ISBN 3-540-70626-7, 3-540-70622-4. LCCN C174.12 .C66 2009.

Staley:2011:BRD

[Sta11a] Richard Staley. Book review: Discontinuous memory and the making of quantum mechanics: Quantum Theory at the Crossroads: Reconsidering the 1927 Solvay Conference by Guido Bacciagaluppi and Antony Valentini; Beyond Uncertainty: Heisenberg, Quantum Physics, and the Bomb by David C. Cassidy; Crafting the Quantum: Arnold Sommerfeld and the Practice of Theory, 1890–1926 by Suman Seth. Historical Studies in the Natural Sciences, 41(4): 447–456, November 2011. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/
10.1525/hsns.2011.41.4.447. See [BV09, Cas09, Set10].

Stanley:2011:BRK

[Sta11b] Matthew Stanley. Book reviews: Kristian Camilleri, Heisenberg and the Interpretation of Quantum Mechanics: The Physicist as Philosopher. Cambridge: Cambridge University Press, 2009. Pp. xii + 199. ISBN 978-0-521-88484-6. £45.00 (hardback). Cathryn Carson, Heisenberg in the Atomic Age: Science and the Public Sphere. Cambridge: Cambridge University Press, 2010. Pp. xxiv + 541. ISBN 978-0-521-82170-4. £55.00 (hardback). British Journal for the History of Science, 44(2):308–309, 2011. CODEN BJH-SAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Storey:1994:PDU

[STCW94] Pippa Storey, Sze Tan, Matthew Collett, and Daniel Walls. Path detection and the Uncertainty Principle. Nature, 367(6464):626– 628, February 17, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v367/n6464/pdf/367626a0.pdf.

Stimson:1947:DUA

[Sti47] Henry L. Stimson. The decision to use the atomic bomb. Bulletin of the Atomic Scientists, 3(2):37–41, 66–67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stroeker:1962:BRW

[Str62a] E. Stroeker. Book review: Wandlungen in den Grundlagen der Naturwissenschaft (Changes in the Fundamental Concepts of Natural Sciences). Ten Lectures by W. Heisenberg. Angewandte Chemie, International Edition in English, 1(4):224, 1962. CO-DEN ACIEAY. ISSN 0570-0833.

Stroker:1962:LPP

[Str62b] E. Ströker. Literatur: Physik und Philosophie, von W. Heisenberg. S. Hirzel Verlag, Stuttgart 1959. 1. Aufl., 201 S., DM 10. Angewandte Chemie, 74(9):333–334, 1962. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Stroker:1962:LWG

[Str62c] E. Ströker. Literatur: Wandlungen in den Grundlagen der Naturwissenschaft. Zehn Vorträge, von W. Heisenberg. S. Hirzel Verlag, Stuttgart 1959. 9. Aufl., 183 S., geb. DM 12. Angewandte Chemie, 74(10):358, 1962. CODEN ANCEAD. ISSN 1521-3757.

Strickland:2011:WSC

[Str11] Jeffrey Strickland. Weird scientists — the creators of quantum physics. Lulu.com, ????, 2011. ISBN 1-257-97624-9. LCCN ????

Stuewer:1985:BNF

[Stu85] Roger H. Stuewer. Bringing the news of fission to America. Physics Today, 38(10):48–56, October 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

www.physicstoday.org/resource/1/phtoad/v38/i10/p48_s1.

Sudo:2006:SDH

[Sud06] Takahiro Sudo. Symmetries on the discrete Heisenberg group C^{∗}- algebra. Journal of Mathematical Physics, 47(3):033512, March 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v47/i3/p033512_s1.

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp. LCCN QC174.1

.A85 1972.  URL http://hooke.lib.cam.ac.uk/cgi-bin/
bib_seek.cgi?cat=ul&bib=1733506;   http://www.loc.gov/
catdir/enhancements/fy1001/72075298-d.html;   http://
www.loc.gov/catdir/enhancements/fy1001/72075298-t.html.

Sachse:2005:PSW

[SW05] Carola Sachse and Mark Walker, editors. Politics and science in wartime: comparative international perspectives on the Kaiser Wilhelm Institute, volume 20 of Osiris. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-73329-7, 0-226-73328-9. ISSN 0369-7827 (print), 1933-8287 (electronic). iv + 310 pp. LCCN Q49.K14174 P65 2005; AS182 .K357 2005.

Sweet:1993:BRU

[Swe93] William Sweet. Book reviews: Uncertainties: Heisenberg’s War: The Secret History of the German Bomb, by Thomas Powers; Uncertainty: The Life and Science of Werner Heisenberg, by David C. Cassidy. Bulletin of the Atomic Scientists, 49(7):50– 52, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wei93].

Saenz:1973:TMB

[SZ73] A. W. Sáenz and W. W. Zachary. Two-magnon bound states in Heisenberg ferromagnets. Journal of Mathematical Physics, 14 (12):1837–1852, December 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v14/i12/p1837_s1. See erra-

tum [SZ74].

Saenz:1974:ETM

[SZ74] A. W. Saénz and W. W. Zachary. Erratum: “Two-magnon bound states in Heisenberg ferromagnets”. Journal of Mathematical Physics, 15(10):1819, October 1974. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v15/i10/

p1819_s1. See [SZ73].

Tanona:2004:UBR

[Tan04] Scott Tanona. Uncertainty in Bohr’s response to the Heisenberg microscope. Studies in History and Philosophy of Modern Physics, 35(3):483–507, September 2004. CODEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S1355219804000371.

Taubes:1994:HHE

[Tau94] Gary Taubes. Heisenberg’s heirs exploit loopholes in his law. Science, 263(5152):1376–1377, March 11, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/263/5152/1376.full.pdf.

Tsai:2002:DCB

[TBL02] Shan-Ho Tsai, Alex Bunker, and D. P. Landau. Dynamic critical behavior of the classical anisotropic BCC Heisenberg antiferromagnet. Computer Physics Communications, 147(1–2):97–100, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002242.

Teller:1930:WGH

[Tel30] Eduard Teller. Über das Wasserstoffmolekülion. (German) [On the hydrogen molecule ion]. Zeitschrift für Physik, 61(7–8):458–480, July 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/

10.1007/BF01330302. Submitted 6 February 1930 as Ph.D. dissertation at Universität Leipzig, under the direction of Professor Werner Heisenberg. Oral examination on 28 February 1930. Doctoral certificate awarded 19 May 1930.

Teller:1961:QME

[Tel61] E. Teller. Der quantenmechanische Meßprozess und die Entropie. (German) [The quantum-mechanical measurement process and entropy]. In ????, editor, Werner Heisenberg und die Physik unserer Zeit, pages 90–92. ????, ????, 1961. LCCN ????

Teller:1971:VWG

[Tel71] E. Teller. Die Verantwortlichkeit des Wissenschaftlers in der Gesellschaft. (German) [The responsibility of the scientist in society]. In ????, editor, Quanten und Felder, Physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg, pages 93–100. ????, ????, 1971. ISBN ???? LCCN ????

Teller:1976:OWH

[Tel76] Edward Teller. Obituary: Werner Heisenberg. Nature, 260(5552): 657–658, April 15, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v260/n5552/pdf/260657a0.pdf.

Teller:1993:EBH

[Tel93] Edward Teller. Erinnerungen und Bilder: Heisenberg in Leipzig. (German) [Memories and pictures: Heisenberg in Leipzig]. In Kleint and Wiemers [KW93], pages 108–109. ISBN 3-05-501585-1. LCCN QC16.H35 W48 1933. EUR$607 (!).

Teller:19xx:HBA

[Telxx] Edward Teller. Heisenberg, Bohr and the atomic bomb. Video in-

terview (6m3s)., 19xx. URL https://www.youtube.com/watch?
v=ClhLnZtgcsE.

Teller:2001:HAS

[Tel01] Edward Teller. Heisenberg hat die Atombombe sabotiert. (German) [Heisenberg sabotaged the atomic bomb]. In Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. (German) [Heisenberg, Hitler and the Bomb: Conversations with witnesses] [Sch01], page 114. ISBN 3-928186-60-4 (paperback). LCCN QC16.H35 S32 2001.

Temme:1991:SGP

[Tem91] F. P. Temme. The structure and Gel′fand patterns inherent in the Heisenberg supergenerator algebra of dual Liouville-space |kq{K̃.}(k_{1} −k_{n}): [λ̃], S_{n}⟩⟩ tensors and the Cayley algebra of scalar invariants over a (k_{1} − k_{n}) field. Journal of Mathematical Physics, 32(6):1638–1650, June 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v32/i6/p1638_s1.

Terry:1982:AFC

[Ter82] Paul Terry. The asymptotic form of the continuum wavefunctions and redundant poles in the Heisenberg condition. Journal of Mathematical Physics, 23(1):87–91, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v23/i1/p87_s1.

Teichmann:2002:PPM

[TEW02] Jürgen Teichmann, Michael Eckert, and Stefan Wolff. Physicists and physics in Munich. Physics in Perspective (PIP), 4 (3):333–359, August 2002. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-002-8372-6.

Thomas:1980:QHF

[Tho80] Lawrence E. Thomas. Quantum Heisenberg ferromagnets and stochastic exclusion processes. Journal of Mathematical Physics, 21(7):1921–1924, July 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v21/i7/p1921_s1.

Thomas:1997:FCW

[Tho97] Robert McG. Thomas, Jr. Frederic A. C. Wardenburg 3d, 92, war hero. New York Times, ??(??):44, August 17, 1997. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/109801785/.

Tiwary:2014:HTU

[Tiw14] Deeptiman Tiwary. Heisenberg’s theory of uncertainty based on Vedas, Rajnath Singh says. The Times of India, ??(??):??,

November 16, 2014. URL http://timesofindia.indiatimes.
com/india/Heisenbergs-theory-of-uncertainty-based-on-
Vedas-Rajnath-Singh-says/articleshow/45163343.cms.

Torresani:1991:WAR

[Tor91] Bruno Torrésani. Wavelets associated with representations of the affine Weyl–Heisenberg group. Journal of Mathematical Physics, 32(5):1273–1279, May 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v32/i5/p1273_s1.

Tokmachev:2009:EGH

[TR09] A. M. Tokmachev and M. A. Robb. Efficient generation of Heisenberg Hamiltonian matrices for VB calculations of potential energy surfaces. International Journal of Quantum Chemistry, 109(3): 401–413, ???? 2009. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trenn:1978:BRB

[Tre78] Thaddeus J. Trenn. Book review: Die Jahrhundertwissenschaft: Werner Heisenberg und die Physik seiner Zeit by Armin Hermann. Isis, 69(2):315–316, June 1978. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/230485.

Treder:2001:BRN

[Tre01] Hans-Jürgen Treder. Book review: Naturwissenschaft und Religion (Versuch einer Verhältnisbestimmung) Science Edition by Wilfried Schröder, Max Planck, Werner Heisenberg. Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte, 85(1):119–120, ???? 2001. CODEN SUARAH. ISSN 0039-4564. URL http:

//www.jstor.org/stable/20777800.

Trifonov:2002:GHU

[Tri02] D. A. Trifonov. Generalizations of Heisenberg uncertainty relation. European Physical Journal B: Condensed Matter and Complex Systems, 29(2):349–353, September 11, 2002. CODEN EPJBFY. ISSN 1434-6028 (print), 1434-6036 (electronic).

Trubatch:1972:EBS

[Tru72] S. L. Trubatch. Erroneous bound state conditions from an algebraic misrepresentation of spin wave theory. Journal of Mathematical Physics, 13(12):1869, December 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v13/i12/p1869_s1.

Truong:1975:WQA

[Tru75] T. T. Truong. Weyl quantization of anharmonic oscillators. Journal of Mathematical Physics, 16(5):1034–1043, May 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i5/
p1034_s1.

Turner:1946:AE

[Tur46a] Louis A. Turner. Atomic energy from U^{238}. Physical Review (2), 69(7–8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/ 10.1103/PhysRev.69.366. Publication voluntarily delayed six years because of war-time secrecy.

Turner:1946:AEU

[Tur46b] Louis A. Turner. Atomic energy from U^{238}. Physical Review (2), 69(7–8):366, April 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/ 10.1103/PhysRev.69.366. Publication voluntarily delayed six years because of war-time secrecy.

Turchetti:2006:BRG

[Tur06] Simone Turchetti. Book review: Gerald Holton, Victory and Vexation in Science. Einstein, Bohr, Heisenberg and Others. Cambridge, Mass.: Harvard University Press, 2005. xiv + 230 pp., ill., ISBN 0-674-01519-3. Nuncius, 21(1):191–192, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00401.

Tuttle:2005:HYG

[Tut05] Jon Tuttle. How you get that story: Heisenberg’s Uncertainty Principle and the literature of the Vietnam War. The Journal of Popular Culture, 38(6):1088–1098, November 2005. ISSN 0022-3840 (print), 1540-5931 (electronic).

URL http://onlinelibrary.wiley.com/journal/10.1111/%
28ISSN%291540-5931.

Talou:2023:NFT

[TV23] Patrick Talou and Ramona Vogt, editors. Nuclear Fission Theories, Experiments and Applications. Springer International Publishing, Cham, Switzerland, 2023. ISBN 3-031-14544-5 (hardcover), 3-031-14545-3 (e-book), 3-031-14546-1, 3-031-14547-X. xvi + 472 + 242 pp. LCCN ????

Telegdi:1991:BRW

[TW91] Valentine L. Telegdi and Victor F. Weisskopf. Book review: W. Blum, H.-P. Dürr, and H. Rechenberg, Heisenberg: Gesammelte Werke [Collected Works]. Physics Today, 44(7):55–58, July 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/phtoad/v44/i7/p55/
s1.

Trischler:2010:PPR

[TW10] Helmuth Trischler and Mark Walker, editors. Physics and Politics: Research and Research Support in Twentieth Century Germany in International Perspective, volume 5 of Beiträge zur Geschichte der Deutschen Forschungsgemeinschaft. Franz Steiner Verlag, Stuttgart, Germany, 2010. ISBN 3-515-09601-9 (paperback). ISSN 1861-1478. 285 pp. LCCN Q180.G4 P49 2010.

Tyagi:1963:NDH

[Tya63] N. K. Tyagi. New derivation of the Heisenberg uncertainty principle. American Journal of Physics, 31(8):624, August 1963. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/31/624/1. See [Loe63]

for pointers to two alternate derivations.

Ullmann-Margalit:1992:SEB

[UM92] Edna Ullmann-Margalit. The Scientific Enterprise: The Bar-Hillel Colloquium: Studies in History, Philosophy, and Sociology of Science, Volume 4, volume 146 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 94-010-5190-9, 94-011-2688-7 (e-book). ISSN 0068-0346. ix + 302 pp. LCCN

Q174 .S354 1992. URL https://link.springer.com/book/10.
1007/978-94-011-2688-5.

Unruh:1972:GFT

[Unr72] Henry Unruh, Jr. Green’s function treatment of short linear Heisenberg chains. American Journal of Physics, 40(6):873–877, June 1972. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://link.aip.org/link/?AJP/40/873/1.

Ulbig:1995:AGW

[URWB95] S. Ulbig, A. Rötger, H. Walther, and L. Behlau. Aktuelles: Ein Gedankenexperiment wird Wirklichkeit: Auf dem Weg zum Heisenberg-Mikroskop für Atome/Neutrinooszillationen: Hinweise, aber (noch) keine Beweise: GALLEX besteht Test mit künstlicher Neutrinoquelle/Atomare Auflösung in RTMinduzierter Photonenemission/Schnellere Transistoren durch Einsatz eines Rasterkraftmikroskops/Magnus-Haus feierlich eingeweiht/ Photonik Fundament der Datenautobahn: Ein Informationsseminar im Heinrich-Hertz-Institut, Berlin/Fraunhofer-Gesellschaft investiert in Dresdener Materialforschung/Wer darf die schwersten Elemente benennen?/Wissenschaftsrat empfiehlt BESSY-II für die Blaue Liste/SSC-Physiker ein Jahr nach dem großen Knall/The Future of Spectroscopy: Eine Konferenz zu Ehren von Gerhard Herzberg/EU aktuell/Koordinierung durch Zusammenarbeit/Projekt- und Antragsvarianten im Programm Informationstechnologie/Ausschreibung der spezifischen Programme/ Neue Kommission der EU. Physikalische Blätter, 51(1):6–20, January 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19950510103/abstract.

VanVleck:1941:NLT

[Van41] J. H. Van Vleck. Note on Liouville’s Theorem and the Heisenberg Uncertainty Principle. Philosophy of Science, 8(2):275–279, April 1941. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/184622.

Villani:2015:RLQ

[VB15] Cédric Villani and Edmond Baudoin. Les rêveurs lunaires: quatre génies qui ont changé l’histoire. (French) [The lunar dreamers: four geniuses who changed history]. Gallimard, Paris, France, 2015. ISBN 2-07-066593-3 (paperback). 179 + 12 pp. LCCN ????

vomBrocke:1996:KWM

[vBL96a] Bernhard vom Brocke and Hubert Laitko, editors. Die Kaiser-Wilhelm-/Max-Planck-Gesellschaft und ihre Institute: Studien zu ihrer Geschichte. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1996. ISBN 3-11-015483-8 (vol. 1). xiii + 672 pp. LCCN Q180.G42 K35 1996. URL http://

www.degruyter.com/viewbooktoc/product/146831;   http:/
/www.reference-global.com/doi/book/10.1515/9783110802443.

vomBrocke:1996:KWS

[vBL96b] Bernhard vom Brocke and Hubert Laitko, editors. Die Kaiser-Wilhelm-/Max-Planck-Gesellschaft und ihre Institute: Studien zu ihrer Geschichte. (German) [The Kaiser Wilhelm/Max Planck Society and their institutions: studies on their history]. Walter de Gruyter, New York, NY, USA and Berlin, Germany, 1996. ISBN 3-11-015483-8 (vol. 1). xiii + 672 pp. LCCN Q180.G42 K35

1996. URL http://www.degruyter.com/viewbooktoc/product/
146831;   http://www.reference-global.com/doi/book/10.
1515/9783110802443.

VanOosten:1995:HE

[VBN95a] A. B. Van Oosten, R. Broer, and W. C. Nieuwpoort. Heisenberg exchange in La_{2}CuO_{4}. International Journal of Quantum Chemistry, 56(S29):241–243, February 25, 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods.

VanOosten:1995:HEC

[VBN95b] A. B. Van Oosten, R. Broer, and W. C. Nieuwpoort. Heisenberg exchange in La_{2}CuO_{4}. International Journal of Quantum Chemistry, 56(S29):241–243, ???? 1995. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods.

vanOosten:1995:HEL

[vBN95c] A. B. van Oosten, R. Broer, and W. C. Nieuwpoort. Heisenberg exchange in La_{2}CuO_{4}. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 29:241–??, 1995. CODEN IJQSDI. ISSN 0161-3642.

vonBorzeszkowski:1988:HCL

[vBW88] Horst-Heino von Borzeszkowski and Renate Wahsner. Heisenberg’s cut and the limitations of its movability. Annalen der Physik (1900), 500(8):522–528, 1988. ISSN 1521-3889.

vanCalmthout:2016:SGJ

[vC16] Martijn van Calmthout. Sam Goudsmit: de jacht op de atoombom van Hitler. Meulenhoff, Amsterdam, The Netherlands, 2016. ISBN 90-290-8958-X (paperback), 94-023-0744-3 (e-book). 284 + 16 pp. LCCN DS135.N6 G656 2016.

vanCalmthout:2018:SGH

[vCH18] Martijn van Calmthout and Michiel Horn, editors. Sam Goudsmit and the hunt for Hitler’s atom bomb. Prometheus Books, Amherst, NY, USA, 2018. ISBN 1-63388-450-3 (hardcover), 1-63388-451-1 (e-book). 243 + 14 pp. LCCN DS135.N6 G65613 2018.

vanDongen:2015:CHU

[vD15] Jeroen van Dongen. Communicating the Heisenberg uncertainty relations: Niels Bohr, complementarity and the Einstein–Rupp experiments. In Aaserud and Kragh [AK15], pages 310–346. ISBN 87-7304-387-7. ISSN 1904-5514. LCCN QC172 .O63

2015. URL http://www.ijqf.org/wps/wp-content/uploads/
2015/06/Bohr-Book.pdf.

vanderWaerden:1967:SQM

[vdW67] B. L. (Bartel Leendert) van der Waerden, editor. Sources of Quantum Mechanics. Classics of science. North-Holland Publishing Co., Amsterdam, The Netherlands, 1967. xi + 430 pp. LCCN QC174.12 S655.

vanderWaerden:1968:SQM

[vdW68] B. L. (Bartel Leendert) van der Waerden, editor. Sources of quantum mechanics, volume 5 of Classics of science. Dover, New York, NY, USA, 1968. ISBN 0-486-61881-1. xi + 430 pp. LCCN QC174.1 .W3 1968.

vonLaue:1947:KDP

[vL47] Max von Laue. Die Kriegstätigkeit der deutschen Physiker. (German) [The war work of the German physicist]. Physikalische Blätter, 3(12):424–425, December 1947. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http:/

/onlinelibrary.wiley.com/doi/10.1002/phbl.19470031206/

abstract. Response to [Mor47]. See also reply [Mor48].

vonLaue:1948:KDP

[vL48a] Max von Laue. Die Kriegsaktivitätkeit der deutschen Physiker. (German) [The wartime activities of the German physicist]. Physikalische Blätter, 4(??):424–425, ???? 1948. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). Cited vaguely in [Wal90a, page 59].

vonLaue:1948:WAG

[vL48b] Max von Laue. The wartime activities of German scientists. Bulletin of the Atomic Scientists, 4(4):103, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Mor48], comment [Rab48], and further reply [vL48a].

Vyatchanin:1997:HMQ

[VL97] S. P. Vyatchanin and A. Y. Lavrenov. Heisenberg microscope and quantum variation measurement. Physics Letters A, 231(1– 2):38–46, ???? 1997. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

vonMeyenn:1993:WPW

[vM93] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band III: 1940– 1949. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume III: 1940–1949, volume 11 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-54911-0. ISSN 0172-6315. lxiv + 1070 pp. LCCN QC16.P37

A34. URL http://www.springer.com/physics/book/978-3-
540-54911-6.

vonMeyenn:1996:WPW

[vM96] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil I: 1950–1952. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part I: 1950–1952, volume 14 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-59442-6. ISSN 0172-6315. xxxvii + 968 pp. LCCN QC16.P37 A34. URL http://www.

springerlink.com/content/978-3-540-78803-4.

vonMeyenn:1999:WPW

[vM99] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil II: 1953–1954. Scientific Correspondence with Bohr, Einstein, Heisenberg, a.o. Volume IV, Part II: 1953–1954, volume 15 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-64312-5. ISSN 0172-6315. xxxv

+ 1100 pp. URL http://www.springer.com/physics/book/
978-3-540-64312-8.

vonMeyenn:2001:WPW

[vM01] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil III: 1955–1956 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part III: 1955–1956, volume 17 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-67591-4, 3-540-78805-0. ISSN 0172-6315. lxv + 994 pp. LCCN QC16.P37 A34.

vonMeyenn:2002:BSA

[vM02] Karl von Meyenn. Vom Bohr-Sommerfeldschen Atom zu Heisenbergs Hexeneinmaleins. Zum 100. Geburtstag von Werner Heisenberg. (German) [From Bohr–Sommerfeld’s atom to Heisenberg’s magic formula. The 100th birthday of Werner Heisenberg]. Naturwissenschaftliche Rundschau, 55(2):69–81, 2002. CODEN NARSAC. ISSN 0028-1050.

vonMeyenn:2005:HZP

[vM05a] Karl von Meyenn. Heisenbergs Zusammenarbeit mit Pauli während der Leipziger Jahre. (German) [Heisenberg’s collaboration with Pauli during the Leipzig years]. In Leipziger Festschrift zu Werner Heisenbergs 100. Geburtstag. (German) [Leipzig Celebrations on Werner Heisenberg’s 100th Birthday], pages 58–81. Berichte der Sächsischen Akademie der Wissenschaften, Leipzig, Germany, 2005. ISBN ???? LCCN ????

vonMeyenn:2005:WPW

[vM05b] Karl von Meyenn, editor. Wolfgang Pauli: Wissenschaftlicher Briefwechsel Mit Bohr, Einstein, Heisenberg u.a. Band IV, Teil IV: 1957–1958 Scientific Correspondence With Bohr, Einstein, Heisenberg, a.o. Volume IV, Part IV: A:1957, B:1958, volume 18 of Sources in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 3-540-40296-9. ISSN 0172-6315. xl + 1585 pp. LCCN QC16.P37 A34. URL http://www.

springerlink.com/content/978-3-540-26832-1.

vonMeyenn:1985:WPW

[vMHW85] Karl von Meyenn, Armin Hermann, and Victor F. (Victor Frederick) Weisskopf, editors. Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Band II, 1930– 1939. Scientific Correspondence with Bohr, Einstein, Heisenberg a.o. Volume II: 1930–1939, volume 6 of Sources in the History of Mathematics and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13609-6 (New York), 3-540-13609-6 (Berlin). xxxix + 783

pp. LCCN QC16.P37 A34. URL http://www.springerlink.
com/content/978-3-540-13609-5.

vonRauchhaupt:2000:BSK

[vRFS^{+}00] Ulf von Rauchhaupt, Harald Fritzsch, Klaus Stierstadt, Rainer Kayser, Klaus Stierstadt, Rainer Kayser, Norbert Dragon, Tilman Pfau, Christoph Bennemann, and Holger Perlt. Bücher und Software: Krafft: Vorstoß ins Unerkannte. Lexikon großer Naturwissenschaftler/Weizsäcker: Große Physiker. Von Aristoteles bis Werner Heisenberg/Zee: Fearful Symmetry. The Search for Beauty in Modern Physics/Schroeder: An Introduction to Thermal Physics/Fraser, Lillestol u. Sellevag: Auf der Suche nach dem Unendlichen/Rebhan: Theoretische Physik Band 1: Mechanik, Elektrodynamik, Relativitätstheorie, Kosmologie/Meschede: Optik, Licht und Laser/Mantegna u. Stanley: An Introduction to Econophysics. Correlations and Complexity in Finance/Maple 6. Physikalische Blätter, 56(9):75–79, September 2000. CO-DEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
phbl.20000560918/abstract.

vonSchirach:2014:NPH

[vS14] Richard von Schirach. Die Nacht der Physiker: Heisenberg, Hahn, Weizsäcker und die deutsche Bombe. (German) [The night of the physicists: Heisenberg, Hahn, Weizsäcker and the German bomb], volume 61642 of rororo sachbuch. Rowohlt Taschenbuch, Reinbek, Germany, 2014. ISBN 3-499-61642-4. 256 pp. LCCN ???? URL

http://d-nb.info/1044314834/04.

vanVliet:1978:LRT

[vV78] K. M. van Vliet. Linear response theory revisited. I. The manybody van Hove limit. Journal of Mathematical Physics, 19(6): 1345–1370, June 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v19/i6/p1345_s1.

vonWeizsacker:1935:TKG

[vW35] C. F. von Weizsäcker. Zur Theorie der Kernmassen. (German) [On the theory of nuclear mass]. Zeitschrift für Physik, 96(7–8):431– 458, July 1935. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-

5831 (electronic). URL http://link.springer.com/article/
10.1007/BF01337700.

vonWeizsacker:1940:MEU

[vW40] Carl Friedrich von Weizsäcker. Eine Möglichkeit der Energiegewinnung aus U 238. (German) [A possibility of generating energy from U 238]. Report AMPG I-34-120, Archiv der MPG, Berlin, Germany, 1940.

vonWeizsacker:1941:BVN

[vW41a] Carl Friedrich von Weizsäcker. Bericht über die Vortragsreise nach Kopenhagen vom 19.–24.3.41. (German) [Report about the lecture tour to Copenhagen from March 19-24, 1941]. Report 34, Archives of the Max Planck Society, Berlin, Germany (AMPG). I. Abteilung, Berlin, Germany, 1941. ?? pp.

vonWeizsacker:1941:KUM

[vW41b] Carl Friedrich von Weizsäcker. Kernenergieerzeugung aus dem uranisotop der masse 238 und anderen schweren elementen. (herstellung und verwendung des elements 94). (German) [Nuclear energy production from the uranium isotope of mass 238 and other heavy elements. (Production and use of element 94)]. Report 34, Archives of the Max Planck Society, Berlin, Germany (AMPG). I. Abteilung, Berlin, Germany, June 6, 1941. ?? pp.

vonWeizsacker:1948:STG

[vW48] C. F. von Weizsäcker. Das spektrum der turbulenz bei großen Reynoldschen zahlen. (German) [The spectrum of turbulence in large Reynolds numbers]. Zeitschrift für Physik, 124(7–12):614– 627, ???? 1948. CODEN ZEPYAA. ISSN 0044-3328. URL https:

//link.springer.com/article/10.1007/BF01668898.

vonWeizsacker:1957:SGA

[vW57] Carl Friedrich von Weizsäcker. Should Germany have atomic arms. Bulletin of the Atomic Scientists, 13(8):283–286, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Followup on [BBF^{+}57a].

vonWeizsacker:1988:BGC

[vW88] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. C. Hanser, Munich, West Germany, second edition, 1988. ISBN 3-446-14649-0. 476 pp. LCCN ????

vonWeizsacker:1991:BGC

[vW91] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. Deutscher Taschenbuch Verlag, Munich, West Germany, 1991. ISBN 3-423-11388-X. 476 pp. LCCN ????

vonWeizsacker:2002:GPA

[vW02] Carl Friedrich von Weizsäcker. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg], volume 33078 of dtv. Deutscher Taschenbuch-Verlag, München, Germany, unabridged edition, 2002. ISBN 3-423-33078-3. 376 pp. LCCN QC6. EUR-D 12.50.

vonWeizsacker:1941:KBE

[vW1 ] Carl Friedrich von Weizsäcker. Kurzer Bericht u ber die eventuelle praktische Auswirkung der Uranuntersuchungen auf Grund einer Ru cksprache mit Dr. Diebner. (German) []. Report 34, Archives of the Max Planck Society, Berlin, Germany (AMPG). I. Abteilung, Berlin, Germany, 1941 (??). ?? pp.

vWeizsacker:1993:CFW

[vWL93] C. F. Freiherr v. Weizsäcker and Konrad Lindner. Carl Friedrich von Weizsäcker über sein Studium in Leipzig. (German) [Carl Friedrich von Weizsäcker on his studies in Leipzig]. NTM Zeitschrift für Geschichte der Wissenschaften, Technik und Medizin, 1(1):3–18, December 1993. CODEN NTMSBJ. ISSN 0036-6978 (print), 1420-9144 (electronic). URL http://link.

springer.com/article/10.1007/BF02914089.

vonWeizsacker:2004:GPA

[vWR04] Carl Friedrich von Weizsäcker and Helmut Rechenberg. Grosse Physiker: von Aristoteles bis Werner Heisenberg. (German) [Great Physicists: from Aristole to Werner Heisenberg]. Marix, Wiesbaden, Germany, 2004. ISBN 3-937715-46-0. 375 pp. LCCN ????

vonWeizsacker:1977:WH

[vWvdW77] Carl Friedrich Freiherr von Weizsäcker and B. L. (Bartel Leendert) van der Waerden. Werner Heisenberg, volume 228 of Reihe Hanser. Hanser, München, West Germany, 1977. ISBN 3-446-12354-7. 103 pp. LCCN QC16.H35 W44.

Vyatchanin:1999:QVM

[Vya99] S. P. Vyatchanin. Quantum variation measurement and a Heisenberg microscope. Optics and Spectroscopy, 87(4):532–540, ???? 1999.

Vinet:2009:AHW

[VZ09] Luc Vinet and Alexei Zhedanov. Automorphisms of the Heisenberg–Weyl algebra and d-orthogonal polynomials. Journal of Mathematical Physics, 50(3):033511, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v50/i3/
p033511_s1.

Wilber:1987:CCE

[W^{+}87] Ken Wilber et al., editors. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, 1987. ISBN 84-7245-172-0. 298 pp. LCCN ????

Wagner:1981:LLH

[Wag81] Siegfried Wagner. Lavoisier, Lichtenberg und Heisenberg. Berichte zur Wissenschaftsgeschichte, 4(3–4):127–141, ???? 1981. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Wallace:1983:HQ

[Wal83] Bryan G. Wallace. Heisenberg and QCD. Physics Today, 36 (9):111–??, September 1983. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v36/i9/p111/s1.

Walker:1989:GNS

[Wal89a] Mark Walker. German National Socialism and the quest for nuclear power, 1939–1949. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36413-2. x + 290 pp.

LCCN TK1078 .W35 1989. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;   http://www.loc.gov/catdir/
description/cam032/88036458.html;   http://www.loc.gov/
catdir/toc/cam031/88036458.html.

Walker:1989:NSG

[Wal89b] Mark Walker. National Socialism and German physics. Journal of Contemporary History, 24(1):63–89, January 1989. ISSN 0022-0094 (print), 1461-7250 (electronic). URL http://www.jstor.

org/stable/260700.

Walker:1990:HGG

[Wal90a] Mark Walker. Heisenberg, Goudsmit and the German Atomic Bomb. Physics Today, 43(1):52–60, January 1990. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://www.physicstoday.org/resource/1/phtoad/v43/i1/

p52_s1. See responses and rebuttal [LRD^{+}91].

Walker:1990:LDA

[Wal90b] Mark Walker. Legenden um die deutsche Atombombe. (German) [Legends about the German atomic bomb]. Vierteljahrshefte für Zeitgeschichte, 38(1):45–74, January 1990. CODEN ???? ISSN 0042-5702 (print), 2196-7121 (electronic). URL https://www.

ifz-muenchen.de/heftarchiv/1990_1.pdf.

Walker:1991:PS

[Wal91] Mark Walker. The physics of solitude. Nature, 354(6352):365–366, December 5, 1991. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v354/n6352/pdf/354365a0.pdf. Review of [Cas92c].

Walker:1992:PPW

[Wal92] Mark Walker. Physics and propaganda: Werner Heisenberg’s Foreign Lectures under National Socialism. Historical Studies in the Physical and Biological Sciences, 22(2):339–389, ???? 1992. CO-DEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/stable/27757685.

Walker:1993:SDA

[Wal93] Mark Walker. Selbstreflexionen deutscher Atomphysiker: die Farm Hall-Protokolle und die Entstehung neuer Legenden um die ,,deutsche Atombombe“. (German) [Self-reflections of German nuclear physicists: the Farm Hall protocols and the creation of new legends “The German atomic bomb”]. Vierteljahrshefte für Zeitgeschichte, 41(4):519–542, October 1993. CODEN ???? ISSN 0042-5702 (print), 2196-7121 (electronic). URL https://www.

ifz-muenchen.de/heftarchiv/1993_4.pdf.

Walker:1995:NSM

[Wal95] Mark Walker. Nazi science: myth, truth, and the German atomic bomb. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-44941-2. viii + 325 pp. LCCN Q127.G3 W35 1995.

Walker:1999:HNA

[Wal99] J. Samuel Walker. Heisenberg and the Nazi Atomic Bomb Project: a Study in German Culture (review). Technology and Culture, 40(4):910–911, October 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33462.

Walker:2003:OHV

[Wal03] Mark Walker. Otto Hahn: Verantwortung und Verdrängung. (German) [Otto Hahn: responsibility and repression]. Forschungsprogramm ,,Geschichte der Kaiser-Wilhelm-Gesellschaft im Nationalsozialismus” (Research Program “History of the Kaiser Wilhelm Society in the National Socialist Era”), Berlin, Germany,

2003. 64 pp. LCCN ???? URL www.mpiwg-berlin.mpg.de/KWG/
Ergebnisse/Ergebnisse10.pdf.

Wald:2004:FLP

[Wal04] Chelsea Wald. Focus: Landmarks: The Physical Review’s explosive secret. Physical Review Focus, 14(??):17–??, ???? 2004. CODEN PRFHAQ. ISSN 1539-0748. URL http://physics.aps.

org/story/print/v14/st17.

Walker:2005:WKR

[Wal05] Mark Walker. Eine Waffenschmiede? Kernwaffen- und Reaktorforschung am Kaiser-Wilhelm-Institut für Physik. (German) [An armory? Nuclear and reactor research at the Kaiser Wilhelm Institute for Physics]. Ergebnisse, 26(??):1–52, 2005. URL http:

//www.mpiwg-berlin.mpg.de/KWG/Ergebnisse/Ergebnisse26.

pdf. Translation to German by Birgit Kolboske.

Walker:2006:OHR

[Wal06] Mark Walker. Otto Hahn: Responsibility and repression. Physics in Perspective (PIP), 8(2):116–163, May 2006. CODEN PHPEF2.

ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-006-0277-3.

Walker:2007:WKR

[Wal07] Mark Walker. Eine Waffenschmiede? Kernwaffen- und Reaktorforschung am Kaiser-Wilhelm-Institute für Physik. (German) [An armory? Nuclear and reactor research at Kaiser Wilhelm Institute for Physics]. In Maier [Mai07], pages 352–394. ISBN 3-8353-0182-9 (paperback). LCCN Q180.G4 K38 2000 v.17; U395.G3 G45 2007.

Walker:2008:BRW

[Wal08] Mark Walker. Book reviews: When physics, politics, and consciousness collide: Faust in Copenhagen: A Struggle for the Soul of Physics, by Gino Segrè. The Mental Aftermath: The Mentality of German Physicists, 1945–1949, by Klaus Hentschel. Quantum Computer Science: An Introduction, by N. David Mermin. Quantum Information: An Overview, by Gregg Jaeger. Physics Today, 61(3):53–54, 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v61/i3/p53/s1.

Walker:2017:PHG

[Wal17] Mark Walker. Physics, history, and the German atomic Bomb. Berichte zur Wissenschaftsgeschichte, 40(3):271–288, September 2017. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Walker:2019:CR

[Wal19] Mark Walker. Copenhagen revisited. In Björkman et al. [BLW19], pages 230–247. ISBN 0-8153-9474-8 (hardcover), 1-351-18508-X (mobi), 1-351-18509-8 (e-pub), 1-351-18510-1 (adobe), 1-351-18511-X (e-book). LCCN DD256.6 .H58 2019.

Walker:2020:BGA

[Wal20a] Mark Walker. A biography of the German atomic bomb. In Forstner and Walker [FW20], pages 163–177. ISBN 3-030-48508-0 (print), 3-030-48509-9 (e-book), 3-030-48510-2, 3-030-48511-0.

LCCN QC16.E5 .B564 2020. URL http://www.springerlink.
com/content/978-3-030-48509-2.

Walker:2020:WHV

[Wal20b] Mark Walker. Werner Heisenberg’s visit to Cracow, November 1943. In Pohl et al. [PFH20], pages 147–153. ISBN 3-86225-534-4, 978-38622-5-1-2-0-9 hardbound. LCCN ????

Walker:2022:DWH

[Wal22] Mark Walker. Did Werner Heisenberg understand how atomic bombs worked? Berichte zur Wissenschaftsgeschichte, 45(1–2): 219–244, June 2022. CODEN BEWID8. ISSN 0170-6233 (print), 1522-2365 (electronic).

Walker:2024:HHA

[Wal24a] Mark Walker. The historiography of “Hitler’s atomic bomb”. Physics in Perspective (PIP), 26(1):18–41, March 2024. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL https://link.springer.com/article/10.1007/s00016-
024-00309-6.

Walker:2024:HAB

[Wal24b] Mark Walker. Hitler’s atomic bomb history, legend, and the twin legacies of Auschwitz and Hiroshima. Cambridge University Press, Cambridge, UK, 2024. ISBN 1-009-47928-8. 380 pp. LCCN D810.S2 W35 2024.

Wang:1998:HCS

[Wan98] You-De Wang. Heisenberg chain systems from compact manifolds into S2. Journal of Mathematical Physics, 39(1):363–371, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:2001:PSS

[Wan01] Wei Wang. Positive solution of a subelliptic nonlinear equation on the Heisenberg group. Bulletin canadien de mathématiques = Canadian Mathematical Bulletin, 44(??):346–354, ???? 2001. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Warner:1990:HP

[War90] Constance A. Warner. Heisenberg’s principles. Nature, 345(6277): 656, June 21, 1990. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v345/n6277/pdf/345656d0.pdf.

Wattenberg:2001:BRP

[Wat01] Albert Wattenberg. Book review: Per F. Dahl, Heavy Water and the Wartime Race for Nuclear Energy. Bristol and Philadelphia: Institute of Physics, 1999, xvi + 399 pages. $60.00 (Hardbound). Physics in Perspective (PIP), 3(1):131–132, March 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wickramasekara:2003:RSR

[WB03] S. Wickramasekara and A. Bohm. Representation of semigroups in rigged Hilbert spaces: Subsemigroups of the Weyl–Heisenberg group. Journal of Mathematical Physics, 44(2):930–942, February 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Williams:1984:AAD

[WC84] Robert Chadwell Williams and Philip L. (Philip Louis) Cantelon, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present, 1939–1984. University of Pennsylvania Press, Philadelphia, 1984. ISBN 0-585-17297-8 (e-book), 0-8122-7920-4, 0-8122-1169-3 (paperback). xv + 333 pp. LCCN U264 .A44 1984.

Wilber:2002:CCE

[WdC02] Ken Wilber and Pedro de Casso. Cuestiones cuánticas escritos místicos de los físicos más famosos del mundo. (Spanish) [Quantum issues: mystical writings of the world’s most famous physicists]. Kairós, Barcelona, Spain, seventh edition, 2002. ISBN 84-7245-172-0. 296 pp. LCCN ????

Weinreich:1946:HPP

[Wei46] Max Weinreich. Hitler’s professors; the part of scholarship in Germany’s crimes against the Jewish people. [YIVO English translation series]. Yiddish Scientific Institute-YIVO, New York, NY, USA, 1946. 3 + 5–29 pp. LCCN DS135.G33 W37.

Weinstock:1985:BRB

[Wei85] Robert Weinstock. Book review: Inner Exile: Recollections of a Life with Werner Heisenberg by Elisabeth Heisenberg; S. Cappellari; C. Morris. Isis, 76(2):283, June 1985. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231816.

Weisskopf:1993:LGH

[Wei93] Victor Weisskopf. Letter: Giving Heisenberg his due. Bulletin of the Atomic Scientists, 49(9):53, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe93].

Weiss:1998:WPH

[Wei98] Peter Weiss. Wave or particle? Heisenberg, take a hike! Science News (Washington, DC), 154(10):149, 1998. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Weinreich:1999:HPP

[Wei99] Max Weinreich. Hitler’s professors: the part of scholarship in Germany’s crimes against the Jewish people. Yale University Press, New Haven, CT, USA, 1999. ISBN 0-300-05387-8 (paperback). xi + 291 pp. LCCN DS146.G4 W3913 1999.

Wells:1992:LHB

[Wel92] Alan Wells. Letters: Heisenberg’s bomb. New Scientist, 135(1842):50, October 10, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL http:

//www.newscientist.com/article/mg13618426.200-letters-
heisenbergs-bomb.html. See [Cha92].

Weyl:1931:GQ

[Wey31a] Hermann Weyl. Gruppentheorie und Quantenmechanik. S. Hirzel, Leipzig, Germany, second edition, 1931. xi + 366 pp. LCCN QA171 .W5 1931.

Weyl:1931:TGQ

[Wey31b] Hermann Weyl. The Theory of Groups and Quantum Mechanics. Methuen and Co. Ltd., London, UK, 1931. xxii + 1 + 422 + 1 pp. LCCN QA171 .W54. Translated from the second (revised) German edition to English by H. P. (Howard Percy) Robertson.

Weyl:1950:TGQ

[Wey50] Hermann Weyl. The Theory of Groups and Quantum Mechanics. Dover, New York, NY, USA, 1950. xxii + 422 pp. LCCN QA171 .W54. Translated from the second (revised) German edition to English by H. P. (Howard Percy) Robertson.

Werner:2019:UHT

[WF19] Reinhard F. Werner and Terry Farrelly. Uncertainty from Heisenberg to today. Foundations of Physics, 49(6):460–491, June 2019. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Wigner:1996:SWJ

[WG96] E. P. Wigner and H. H. Goldstine. The scientific work of John von Neumann. In Wigner [Wig96m], pages 123–126. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Whitfield:2007:WMR

[WH07] Donald Whitfield and James L. Hicks, editors. What’s the matter?: readings in physics. Great Books Foundation, Chicago, IL, USA, 2007. ISBN 1-880323-91-5. xx + 540 pp. LCCN QC7.5 .W53 2007.

URL http://store.greatbooks.org/adu-wtm.html. Foreword

by Alan Lightman.

Winterberg:1996:MHN

[WHF^{+}96] Friedwardt Winterberg, Günter Herrmann, Igor Fodor, Lincoln Wolfenstein, and Mark E. Singer. More on how Nazi Germany failed to develop the atomic bomb. Physics Today, 49(1):11–83, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Whitaker:1996:EBQ

[Whi96] Andrew Whitaker. Einstein, Bohr, and the quantum dilemma. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-48220-8 (hardcover), 0-521-48428-6 (paperback). xvii + 349 pp.

LCCN QC174.12 .W48 1996. URL ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;   http://www.cambridge.org/
uk/catalogue/catalogue.asp?isbn=0521484286;   http://
www.loc.gov/catdir/description/cam027/95018270.html;
http://www.loc.gov/catdir/toc/cam021/95018270.html.

Wightman:1967:BRW

[Wig67] Arthur S. Wightman. Book review: Werner Heisenberg, Introduction to the Unified Field Theory of Elementary Particles. Physics Today, 20(12):77–??, December 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/phtoad/v20/i12/p77/s1.

Wigner:1976:OWK

[Wig76] Eugene P. Wigner. Obituary: Werner K. Heisenberg. Physics Today, 29(4):86–87, April 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v29/i4/p86/s1.

Wigner:1996:ADM

[Wig96a] E. P. Wigner. Address delivered to the Memorial Meeting [for Paul Dirac]. In Wigner [Wig96m], pages 214–215. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:ABE

[Wig96b] E. P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Wigner [Wig96m], pages 133–138. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:BJN

[Wig96c] E. P. Wigner. Biography of John von Neumann. In Wigner [Wig96m], pages 209–210. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CRA

[Wig96d] E. P. Wigner. Concluding remarks (address at the Dirac Symposium). In Wigner [Wig96m], pages 195–196. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EM

[Wig96e] E. P. Wigner. Einstein — A memoir. In Wigner [Wig96m], page 197. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EF

[Wig96f] E. P. Wigner. Enrico Fermi (1901–1954). In Wigner [Wig96m], pages 115–119. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:EAE

[Wig96g] E. P. Wigner. Erinnerungen an Albert Einstein. (German) [Memories of Albert Einstein]. In Wigner [Wig96m], pages 198–200. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:JN

[Wig96h] E. P. Wigner. John von Neumann (1903–1957). In Wigner [Wig96m], pages 127–130. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:LS

[Wig96i] E. P. Wigner. Leo Szilard (1898–1964). In Wigner [Wig96m], pages 139–149. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:NLE

[Wig96j] E. P. Wigner. New light on Einstein letter — an interview with E. P. Wigner and A. M. Weinberg by D. Sundberg. In Wigner [Wig96m], pages 216–218. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:RPD

[Wig96k] E. P. Wigner. Remembering Paul Dirac. In Wigner [Wig96m], pages 219–230. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:TYK

[Wig96l] E. P. Wigner. Thirty years of knowing Albert Einstein. In Wigner [Wig96m], pages 201–208. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96m] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Willers:1936:BWHa

[Wil36a] Fr. A. Willers. Buchbesprechung: W. Heisenberg, Wandlungen in den Grundlagen der Naturwissenschaften. Zwei Vorträge. 45 S. Leipzig 1935, Verlag S. Hirzel. Preis kart. 2 M. Zeitschrift für Angewandte Mathematik und Mechanik, 16(1):62–63, 1936. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

Willers:1936:BWHb

[Wil36b] R. A. Willers. Buchbesprechung: W. Heisenberg, E. Schrödinger, P. A. M. Dirac, Die moderne Atomtheorie. Die bei der Entgegennahme des Nobelpreises 1933 in Stockholm gehaltenen Vorträge. 45 S. u. 6 Fig. Leipzig 1934, Verlag von S. Hirzel. Preis kart. 2,50 M. Zeitschrift für Angewandte Mathematik und Mechanik, 16(2): 126, 1936. CODEN ZAMMAX. ISSN 1521-4001.

Williams:1984:CSH

[Wil84] B. G. Williams. Compton scattering and Heisenberg’s Microscope revisited. American Journal of Physics, 52(5):425–430, May 1984. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Williams:1994:BRB

[Wil94] Robert C. Williams. Book review: Heisenberg’s War: The Secret History of the German Bomb. By Thomas Powers. (New York: Knopf, 1993. xii, 607 pp. $27.50, ISBN 0-394-51411-4). The Journal of American History, 80(4):1521–1522, 1994. ISSN 0021-8723 (print), 1945-2314 (electronic). URL http://jah.

oxfordjournals.org/content/80/4/1521.2.short.

Williams:1999:HUP

[Wil99] M. Williams. The Heisenberg uncertainty principle and anaesthetic performance. Anaesthesia, 54(5):513–514, ???? 1999. CO-DEN ANASAB. ISSN 0003-2409 (print), 1365-2044 (electronic).

Wiseman:1998:EHM

[Wis98] H. M. Wiseman. Extending Heisenberg’s measurement-disturbance relation to the twin-slit case. Foundations of Physics, 28(11):1619– 1631, ???? 1998. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Witte:1983:BJH

[Wit83] H. Witte. Buchbesprechung: J. H. Scharf (Ed.): In memoriam Werner Heisenberg, Nova Acta Leopoldina, Abhandlungen der Deutschen Akademie der Naturforscher Leopoldina. Neue Folge Nr. 248, Band 55. Halle (Saale) 1982. Kommissionsverlag Johann Ambrosius Barth, Leipzig. 136 Seiten, Preis: DM 43. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 1963–1990, 87(8): 704, 1983. CODEN BBPCAX. ISSN 0005-9021.

Witze:2012:MEA

[Wit12] Alexandra Witze. Matter & energy: Adding precision to uncertainty: Heisenberg’s famous physics principle gets refinements. Science News (Washington, DC), 182(8):15, 2012. CODEN SC-NEBK. ISSN 1943-0930.

Weidlich:1992:BGG

[WMG^{+}92] W. Weidlich, K. V. Meyenn, J. Geurts, W. Priester, E. Bauer, and H. W. Spiess. Buchbesprechungen: Gierer: Die gedachte Natur/ Cassidy, Freeman: Uncertainty. The Life and Science of Werner Heisenberg/Werner, Weber, Rühle: Halbleiter in Forschung und Technik/Maran: The Astronomy and Astrophysics Encyclopedia/Watts: Introduction to Surface Analysis by Electron Spectroscopy/Callaghan: Principles of Nuclear Magnetic Resonance Microscopy. Physikalische Blätter, 48(7–8):624–626, July/August 1992. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19920480743/abstract.

Wolff:2014:HOP

[Wol14] Johanna Wolff. Heisenberg’s observability principle. Studies in History and Philosophy of Modern Physics, 45(??):19–26, February 2014. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S135521981300107X.

Woo:1977:PCT

[Woo77] G. Woo. Pseudoparticle configurations in two-dimensional ferromagnets. Journal of Mathematical Physics, 18(6):1264–1266, June 1977. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v18/i6/p1264_s1.

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Walker:1992:FHT

[WR92] Mark Walker and Helmut Rechenberg. Farm-Hall-Tonbänder: Über die Uranbombe: Werner Heisenbergs abgehörter Vortrag vom 14. August 1945 in Farm Hall. Physikalische Blätter, 48(12): 994–1001, December 1992. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic).

Wreszinski:1975:PSG

[Wre75] Walter F. Wreszinski. On properties of the singularity of the ground state in certain classical Heisenberg models. Journal of Mathematical Physics, 16(9):1857–1859, September 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i9/
p1857_s1.

Wuthrich:2016:HUC

[Wüt16] Adrian Wüthrich. Heisenberg’s Umdeutung: A case for a (quantum-)Dialogue between history and philosophy of science. In Sauer and Scholl [SS16], chapter 14, pages 285–296. ISBN 3-319-30227-2, 3-319-30229-9 (e-book). ISSN 0068-0346. LCCN

Q126.8 .P45 2016. URL https://link.springer.com/chapter/
10.1007/978-3-319-30229-4_14.

Wheeler:1983:QTM

[WZ83] John Archibald Wheeler and Wojciech Hubert Zurek, editors. Quantum Theory and Measurement. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1983. ISBN 0-691-08315-0, 0-691-08316-9 (paperback). xxviii + 811 pp. LCCN QC174.125 .Q38 1983. US$60.00, US$19.50.

Yafaev:2000:HM

[Yaf00] Dmitri R. Yafaev. The Heisenberg model. Lecture Notes in Mathematics, 1735:137–144, 2000. CODEN LNMAA2. ISBN 3-540-67587-6 (print), 3-540-45170-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0105547/.

Yam:2002:NSN

[Yam02a] Philip Yam. News scan: In no uncertain terms [letter of Niels Bohr to Werner Heisenberg]. Scientific American, 286(4):30, April 2002. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.sciam.com/2002/0402issue/
0402numbers.html.

Yamazaki:2002:YWHa

[Yam02b] Kazuo Yamazaki. 100 years Werner Heisenberg works and impact. Fortschritte der Physik = Progress of Physics, 50(5–7):459–460, 2002. CODEN FPYKA6. ISSN 1521-3978.

Yamazaki:2002:YWHb

[Yam02c] Kazuo Yamazaki. 100 years Werner Heisenberg works and impact. In Papenfuss et al. [PLS02], pages 30–31. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Yang:2013:SPI

[Yan13a] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing, Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.com/
worldscibooks/10.1142/8640.

Yang:2013:WHC

[Yan13b] Chen Ning Yang. Werner Heisenberg (1901–1976). commentary. In Selected papers II, with commentaries [Yan13a], pages 248–253. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). LCCN QC21.3. URL http://

www.worldscientific.com/doi/abs/10.1142/9789814449021_
0032.

Yeh:1970:SSO

[Yeh70] Robert H. T. Yeh. Some simple observations on Griffiths’ theorems for the classical Heisenberg model. Journal of Mathematical Physics, 11(4):1317–1319, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v11/i4/p1317_s1.

Zak:1996:DWH

[Zak96] J. Zak. Discrete Weyl–Heisenberg transforms. Journal of Mathematical Physics, 37(8):3815–3823, August 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zandonella:2002:NUH

[Zan02a] Catherine Zandonella. No uncertainty over Heisenberg’s role in Nazi bomb. New Scientist, 173(2330):15–??, February 16, 2002. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500

(electronic). URL http://www.newscientist.com/article/
mg17323302.000-no-uncertainty-over-heisenbergs-role-in-
nazi-bomb.html.

Zandonella:2002:UHB

[Zan02b] Catherine Zandonella. Uncertainty over Heisenberg’s bomb making ends. New Scientist, ??(??):??, February 7, 2002. CO-DEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (elec-

tronic). URL http://www.newscientist.com/article/dn1897-
uncertainty-over-heisenbergs-bomb-making-ends.html.

Zeldovich:1987:HIS

[Zel87] Ya. B. Zel’dovich. Heisenberg and isotopic spin. Physics Today, 40(4):124–??, April 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v40/i4/p124/s2.

Zhou:1996:HLG

[ZG96] Huan-Qiang Zhou and Xi-Wen Guan. Hidden local gauge invariance in the one-dimensional Heisenberg XXZ model with the general boundary terms. Journal of Mathematical Physics, 37(1):227– 232, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zupancic:1965:CRM

[Žup65] A. O. Župančič. Creation rate of matter and the Heisenberg Uncertainty Principle. Nature, 207(4994):279, July 17, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v207/
n4994/pdf/207279a0.pdf.